A Complete Bibliography of Publications in Marine Policy (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

14 August 2025 Version 1.66

Title word cross-reference

30 × 30 [LMC^{+}23b, ÓHCÓ24]. ^{th} [HAD24]. _{2} [EV22, GZW^{+}20, GICM25, MVSP^{+}21, OOGK^{+}24, SSK23]. Q [POM24]. × [EHÓ25].

İskenderun [CA21].

0 [LK24b, Tan23a, WQ25].

10.3389/978 [LK24b]. 107 [GZW^{+}20]. 11 [GA20]. 115 [EHM^{+}22, Mot22]. 124 [SDG^{+}22]. 126 [TGM22]. 13th [WFP25]. 14 [Gul20, Tho25]. 14.7.1 [Liu23]. 143 [SCSR^{+}23]. 14801 [Ber25]. 151 [RP23a]. 155 [WWG^{+}25]. 160 [Smi23b]. 168 [CBH^{+}24b, SCM^{+}25]. 169 [SC25]. 170 [DHU25]. 171 [ZGR^{+}25a]. 180 [BJ25, FSRC25a, MHS^{+}25a]. 19 [AEL^{+}23, ASGZ22, ASG^{+}24, ABME22, BRM^{+}21, BSI^{+}21, BASM22, CDNS23, CCKK22, CRCP^{+}23, De 21, DME23, Ero22, FTM^{+}22, FGPPGG22, GPTASM23, GABSD21, GNMGWM21, GMCMJA^{+}21, GS21, HDCH21, HWC^{+}25, HHB^{+}21, HBP^{+}21, Isr22, KFP20, KdBK25, KAC22, Kob22, LWSS21, LSB^{+}21, LC20b, LEERRM^{+}21, LKE^{+}22, MCBN22, MÇ23, MAK22, MGK^{+}25, MMK^{+}22, NK21, NGRM^{+}22, NVP^{+}23, NAS^{+}22, PGK22, PAK22, PBD^{+}23, RSK^{+}25, RDD^{+}21, SVV24, SSB^{+}22, SCH^{+}24b, SdCMN^{+}22, SCLM^{+}21, SSP^{+}21a, SNK^{+}20, SRH^{+}22, SSP^{+}21b, SAB^{+}24, Tan22, TAK^{+}22, VMM21, VDP^{+}22, VMJ25, VMP^{+}24, WHL22, WPWM^{+}22, XSCL21, YCBW21, ZLG^{+}24, ZTZ^{+}21, ZTW23, ZTM^{+}23, ZT23a, ZCS^{+}23]. 1949 [Kay25]. 1960s [GMVM23, Lee24].

2-8325 [LK24b]. 20 [HAD24, TMAS24]. 20-year [VBd^{+}21]. 2000 [PdBA^{+}25]. 2000s [Lee24]. 2001 [LC23]. 2005 [HMM^{+}24, MK24]. 2006 [CK23, KBC25]. 2018 [Kay25, ZPX^{+}22]. 2020 [Ano20a, Ano20b, Ano20o, Ano20p, Ano20r, Ano20q, Ano20s, Ano20t, Ano20u, Ano20v, Ano20w, Mar21a]. 2021 [Ano21a, Ano21b, Ano21c, Ano21p, Ano21q, Ano21s, Ano21r, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, SWGZ22]. 2022 [ALLS23, Ano22a, Ano22b, Ano22c, Ano22p, Ano22q, Ano22s, Ano22r, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x]. 2023 [Ano23a, Ano23b, Ano23p, Ano23s, Ano23r, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, BÖ24, CBL24, YCL25]. 2024 [Ano24a, Ano24b, Ano24c, Ano24q, Ano24s, Ano24r, Ano24t, Ano24u, Ano24v, Ano24w, Ano24x]. 2025 [Ano25b, Ano25a, Ano25o, Ano25q, Ano25p, Ano25r, Ano25s, Ano25t, Ano25u, Ano25v]. 21st [PZPB21, SF22, CTH^{+}23]. 2nd [Tan23b].

3 [Cha20b, Lv20]. 305 [Smi22]. 34 [SFGEPC22]. 3537 [LK24b]. 3537-0 [LK24b].

4 [Tan23b]. 4.0 [KBP24]. 450 [Smi20]. 48.3 [CHC^{+}21]. 49.99 [Rub24]. 492pp [SSSL25]. 4China [WFP25].

5.1 [ALLS23].

6 [LPO25, Niu22]. 65-day [BSI^{+}21].

7 [Ngu23]. 77 [dlC24].

82 [Ege23].

9 [LL24a, Mar21a]. 97 [Mal21]. 978 [Cha20b, LPO25, LL24a, Lv20, Ngu23, Niu22, Tan23a, Tan23b, WQ25]. 978-0-19-776715-3 [Ste25]. 978-1-032-20442-0 [WQ25]. 978-1-032-27099-9 [LL24a]. 978-1-108-84226-6 [Niu22]. 978-1-138-38814-7 [Ngu23]. 978-1-316-51742-0 [Tan23a]. 978-1-78643-673-3 [Cha20b, Lv20]. 978-1-78990-907-4 [SSSL25, Tan23b].

978-3-031-10126-7 [Rub24]. 978-3-031-16276-3 [NJW25]. 978-3-031-20739-6 [Sta23a]. 978-3-031-20740-2 [Sta23a]. 978-3-031-20742-6 [Sta23a]. 978-3-031-75177-6 [LPO25]. 978-90-823810-4-7 [HH21].

A.-K [Sta23a]. Aarhus [RPY^{+}24]. abalone [HCSBM^{+}23]. Abandoned [GRS^{+}23, RCM24, GHW^{+}22, GAD^{+}23, HGG^{+}22a, Lov23, NMN^{+}25, Sam22]. abatement [NNR^{+}21, SBW^{+}21]. abating [Osm23]. Aberdeen [BHCL23]. ability [FWL^{+}22, GCV^{+}21, Mag23]. Abingdon [WQ25]. ABMTs [TCZ21, TG25a]. aboard [MLE^{+}24]. Aboriginal [DF24]. abundance [AFRFGP25, DT23, DVRK25, GSAF21, MJ20, SGPF21]. abuses [APGWTM^{+}24, WS23]. academic [GLB^{+}24, VAS^{+}25]. ACAP [BKW^{+}24]. Accelerating [BDM21, Cha22]. acceptability [EV22, HHEJB20, MHF^{+}25, RF25, SMC^{+}21]. acceptable [HWT23]. Acceptance [ZMT^{+}25, KNY20, LYQM21]. Access [BBS^{+}21, BCG^{+}24, GBLBV20, MKS21, AKF^{+}25, Art20, DTBP21, FHBS23, KBH^{+}24, KL21b, LP20, MCS25b, NM23, ORdS25, OA20, PFG^{+}21, PGK22, SPFA^{+}20, SCF^{+}22, SAC25, AGK22]. accessible [LHV^{+}20, PCHD^{+}25, SMC^{+}25]. Accidents [Cha20b, FFE20, FFE^{+}24, KLK23, RE23, YHU22]. accomplish [ALLS23]. Accord [HvdBH23]. according [dSJNS20, KS22]. accountability [BDC^{+}21]. Accounting [DA24, CVH^{+}20, DJC^{+}23, GPA^{+}22, GGP^{+}22, HCDN24, KSO^{+}25, MMNC22, PMN^{+}22, TGP^{+}24]. accounts [CQAS^{+}25, CNG^{+}23, GMG^{+}23, NMCL24]. accumulated [vdBSvdV^{+}23]. accuracy [DQJ23, dSOW^{+}21]. accurate [ZLD^{+}21]. achievable [HRC^{+}24]. achieve [CMOB^{+}20, LZL25, PHD^{+}24, TFH^{+}23]. achievement [PWL^{+}23a]. Achievements [GJS22, CCX23, Gul20, Tan24, Fis22]. Achieving [AAA^{+}24, BPA^{+}22, MMSB^{+}20, PDB^{+}23, Puh23, EHÓ25, GKP^{+}22, HvPPV20b, Kob22, OKNB22, OGS24, Wan24a, DUYEA21]. Achilles [BD23]. acidification [AHN^{+}23, GBC^{+}20, WWS^{+}25]. acknowledge [VBd^{+}21]. acoupa [dSJNS20, dSJNS20]. acoustic [KDM^{+}20, MGBS24, WCF^{+}21]. across [BDC23, BTN^{+}24, DDAB^{+}23, DPG^{+}22, FRVB22, FIB25, FLLML25, GSAF21, HHEJB20, KMTvT25, KBD^{+}20, LJ20, Lot22, MWM^{+}23a, OPL20, RKKM24, SHPW21, SHP^{+}21, SPNP24, SOLN20, TOHK20, XWH^{+}22, YT20]. Act [BGK^{+}22, HMM^{+}24, WMKG21, BMR23, Kay25, Squ21]. Actiniaria [YNZ22]. Action [FS22, MPSF20, NJW25, AHB^{+}23, BGK^{+}22, BCB20, CWLD25, CCG^{+}25, KK22, KH24, LG24, Lia25, LA23, MWM^{+}23a, Moh23, OGGC20, PN20, PKB^{+}23, WNJ25, WW21a, WWD23, WWS^{+}25, Zhu22, vdBSvdV^{+}23]. action-oriented [WNJ25]. actionable [SB20]. Actioning [HNA25]. actions [BDM21, COYS24, Joh25, Li23a, MB22, PCG^{+}23, TBAR^{+}25, WYH21, dlPC24]. Activating [Men23]. active [CCK^{+}21, DXS^{+}21, GWAPRPRU22, JG20, SKL^{+}20]. activities [ABI24, AP24, Cao25, CCBB^{+}25, CWR20, DNR21, Fei20, GCL21, Gug22, HHBC20, Kim23, LCL20, LWX^{+}20, NWS^{+}24, OAAMS24, Puh23, QWBY21, Rog24, SOT^{+}22, SvLGK24, SV22, SSMG21, SHAC23, TLT^{+}22, YG20, YSEK22, Zha25]. activity [CLS^{+}23, EEIER25, KDM^{+}20, NAN^{+}24, Okş23, SGB21, Shi24a, SVP^{+}23, War20, Yu22]. activity-based [Okş23]. actor [QRK^{+}23]. actor-driven [QRK^{+}23]. actors [BPA^{+}22, BE20b, CDNS23, CC23, SFB24, YWYL22, KWMM20a, MAMCHP^{+}20]. adaptable [JvH24]. Adaptation [CDNS23, BOOO23, BG20b, Bir20, BB20, BGQL24, BvdH23, CBAPM20, DMK24, DP23, FAMW^{+}25, FGPVPP^{+}24, GPI^{+}24, GOH^{+}20, JTR^{+}20, MKM23, O’D21, OCD^{+}20, OGGC20, PCCOS24, RKR20, RHL20, SMLRB^{+}21, SSF23, TRN^{+}24, VHSV20, WWS^{+}25]. adaptations [BRM^{+}21, IRK^{+}21]. adaptative [dMVJdAdSM25]. adapted [BCR25]. Adapting [GDB^{+}24, LMSLMQ21, CMWT24, SOO20, SGA^{+}21]. Adaption [RHO24]. Adaptive [Dun21a, Kim23, VLG^{+}24, WGBS25, dV21, BBJ^{+}22, CD25, CCG^{+}25, Cra20, FSF^{+}21, GLB^{+}24, HCWC21, RRS25, TFH^{+}23, WMS^{+}25, WGP^{+}21]. adaptiveness [SSL25]. added [WCC^{+}20]. Adding [KY20, LTW23]. address [APGWTM^{+}24, Ber24, LTJ^{+}23, McC24, PFWM23, ST24, SHZ^{+}21, Tel23b, VHCJdYC^{+}24]. Addressing [GTMP20, JSW22b, LJ23, OPL20, SGB21, CWKW20, FWGM24, SRY24, VMV^{+}24a, WW25, Wen20, GSM20, PMM^{+}21]. Adjacency [MS20]. adjacent [BDMW22, QTPL21, Tod22b]. Adjustment [SWXM21, BBK21a, CQ20]. adjustments [XL25]. administrative [AKFY22, Oan22, Ros22, Sak21]. ado [OSV21]. adopt [LPMS24, PDD^{+}25]. adopted [BRP^{+}23, MB22]. adopting [CGGA20, KSS23]. Adoption [dlPdlLLC^{+}22, FLLML25, GMC24, Lin25, MB22, NCD23, TM23b, WMRB22, XLW24]. Adriatic [VJ24, BLM^{+}25, SPN^{+}21]. advance [KFP20, KSB^{+}23a, LGW^{+}21, SC24, SC25]. Advanced [WKK^{+}25]. Advances [SS23, TLT^{+}22, GWÁV^{+}25, KSB^{+}23a, dS21]. Advancing [ANC^{+}25, Joh25, KBH^{+}24, SMB^{+}25, OWC^{+}24]. Advantages [Jen23]. adverse [BMR20, CDC24]. advice [AMR21b, CPR21, CBB^{+}24, PCCR^{+}25, BKW^{+}24]. advisory [GLT23, LNRM23, QSC24]. advocacy [Shi23b]. Aegean [SDL^{+}22]. aerial [SSOP^{+}21]. affairs [LTSL22, Tod25, Wil21b]. affect [ÁCT20, HvPPV20a, JD20, NAN^{+}24, RJ21, WYD^{+}22]. affected [LWSS21, MNGM24, OGS25, RCT20, SVV24]. affecting [Cao25, LSY20, ZZW^{+}21, ZZD^{+}23]. affiliations [CT24]. Africa [AAA^{+}24, AWBC20, AN25, AGA^{+}21, BDVR25, BRM^{+}21, BBHN20, BTN^{+}24, CES20, hBDF24, DAN^{+}23, DSV^{+}25, Ege23, EJC24, FTGD^{+}20, FWSS^{+}25, FMLMM23, GSG^{+}24, MZVC25, MF22, MBG^{+}24a, NRF21, NAS^{+}22, OYECdOP24, SRvNN21, SON21, SS21b, SSP^{+}21a, SN23, UOO20, WLCG23]. African [BCBS22, BGMCG24, OYKBA22, Sen25, TNE^{+}24]. after [ABC^{+}24, BMGDS^{+}20, Lia25, MHT^{+}24, NAP^{+}21, PBG^{+}22, SBG^{+}25, SMC^{+}21, SFB24, TMDN23, ZTM^{+}23]. aftermath [BVvKL21]. Afuera [dPJ21]. again [GRvdM24]. against [ABME22, BSI^{+}21, Gar22, GY23, Gul23, HJPL23, KAN^{+}23, Ma20, MR24, Oan22, PGSS22, Poo23]. age [HVN21, Ros25, SBGR24]. agencies [Han21, Min24, Pan25a]. agency [BEM^{+}20]. agenda [GEB^{+}21, LMC^{+}23b, TSH^{+}22a, ZLY23]. agent [MF23]. agents [GC22]. agglomeration [ZW21]. Aggregating [CMT^{+}23, MSZ^{+}23, DPB^{+}23, EMH^{+}23, WLRH^{+}20, ZMM^{+}23]. Agoa [MLJ^{+}22]. Agreement [ALLS23, SLP^{+}22, WS24, Bal22, CJK24, CMSN^{+}22, Cla20, CM25, DAB^{+}25, Fri24, GNJR24, HGL^{+}20, Kim20, LX24, Liu24, Mar21a, MH23, Sas25a, Shi20a, dlC24, BJH21, EB23, Has23, Kim21, Kim24, LZ24, Moh23, Mol24, Tan24, TG25a, VSV^{+}23, VYB20, Wan23a, WW25, ZL24]. agreements [Aqu23, PK20b]. agricultural [ZCG^{+}23]. Agriculture [Gug22]. agrifood [MdCLL^{+}23]. ahead [CCNGS25, CQW24, GJS22, NCE21]. AHOY [KDS22]. AHP [AHBAKH24, CA21, DG20, LZL23]. AHP-entropy-based [DG20]. AI [KMSN25]. AI-driven [KMSN25]. Aichi [GA20]. aids [Kar22c]. aim [CLD^{+}24]. aiming [NW20]. Air [VGK23, BG20a, CAC23, CZW24, HKMS20, KSS23, Kon20, SOS20, VMV^{+}24a]. AIS [ATF^{+}21, CLS^{+}23, HWM^{+}25, MLJ^{+}22, SSOP^{+}21]. Akili [CMWT24]. al [CBH^{+}25, LLM^{+}23, MJMAA22, MJ23, SCD21, SCM^{+}25, TGM22]. Alaska [Bir20, Bra22, DCC^{+}20, GBSC22, SMF^{+}23, SCH^{+}24b, Szy20, SSF23]. albacore [Abo23]. albatross [DPG^{+}22]. Albert [GAA^{+}20]. Alboran [EEIER25]. ALDFG [GRS^{+}23]. Aleutians [Ree20]. algae [GBLBV20]. algal [EHLT21, SLMG^{+}21]. Algeria [GDY25]. algorithm [KP24]. alien [DUYEA21, OGS25, SSDQ20]. alienation [TNY23]. align [CPR21]. Aligning [LIK^{+}24, SPHO23]. alignment [MMG^{+}25]. Aline [SSSL25, Tan23b]. Aliwal [FMLMM23]. alleviate [YSWW21]. alleviation [OWS20]. alliances [GR25]. Allocating [DMT25]. Allocation [FDF^{+}21, HGSD23, HM22b, MKS21, Per24, Ren21b, TKP23]. allocations [AGN^{+}20]. allow [PCCR^{+}25]. Allowable [DSJ^{+}23, ZLD^{+}21, EJC24, Sta23b]. allowed [WL24a]. along [AVC24, Alf20, CDRS22, EEIER25, FFE^{+}24, GSA23, HLLK22, KOW^{+}25, KSV^{+}23, MDLA21, OWCFDC21, RPK23, RSLR24, ZWG^{+}21, ZNA^{+}23]. Alopias [MH20b]. already [Men23]. also [LZZ22]. Alternating [SAB^{+}24]. Alternative [GM21, HCDN24, WL20, CDS^{+}25, Gor21, Han21, HRS^{+}24, KGW24, Lat25, MF23, MCTN23, Pha20, WZX24, WMS21, dSOW^{+}21]. alternatives [PCD^{+}23, dS22a, dJRFVOA24]. Amanda [Rub24]. Amapá [JAdS^{+}20]. amateur [DD22]. Amazon [AML^{+}21, GdCC^{+}25, JAdS^{+}20, dJdMCL^{+}24]. Ambarlı [Okş23]. ambiguity [BvdH23, KvTNL20, Li23b]. ambition [DAB^{+}25, Obu20]. ambitions [Pan25b]. ambitious [RSS^{+}20]. Ambon [LLLF22]. Ameghinomya [BRR^{+}22]. amended [Lia25]. Amendment [ML22, LH24]. amendments [VDP^{+}22, Kay25]. amenity [DR25]. America [AFR25, AE23, GGMCMM^{+}24, KABT21, MBC^{+}24, San25]. American [MGPR^{+}25, BKG^{+}25, CMS^{+}25, MM20, PBH^{+}23, SK24, dOPM^{+}20]. americanus [MM20]. Americas [OMC^{+}22]. amid [HCW^{+}23, MYGM25]. amidst [SNK^{+}20, SAB^{+}24]. ammonia [FLLML25]. among [BPA^{+}22, BA25b, CLW^{+}21, DDD22, GPI^{+}24, GWCD22, IOHC22, Lev20, MRKM23, MJ23, MMK^{+}22, ONN^{+}20, PCCR^{+}25, SBG^{+}25, SKS21, TRN^{+}24, TNY23, TSWC23, VM21, WSD21, ZMT^{+}25]. amount [HGG^{+}22a]. Ampat [SWI^{+}22, WMRB22]. Amphiprion [YNZ22]. Amsterdam [De 20, DPG^{+}22]. Amundsen [BAAH24a]. Anadromous [SCBdA20]. analyse [FFD^{+}23]. analyses [GÁPSAP^{+}23]. Analysing [VGA25]. Analysis [BRKS25, CSMMC^{+}25, CJ21, FFE^{+}24, HZ23, HCSBM^{+}23, JL21c, Kao23, KLS^{+}25, MCRPC20, PKC23, PCD^{+}23, RJAT23, Şan20, SFGEPC22, SV22, YXDL24, dRA25, AHBAKH24, AN25, ASKI20, AOO^{+}23b, AH23, AVC24, AGA^{+}21, AONF21, ANR^{+}22, BDVR25, BG20a, BDLO23, BISSBdlMM22, BDCJ22, BGK^{+}22, BMC20, BBBS22, BGQL24, BYL25, CSLGFV20, CMC^{+}23, CQ20, CDHY22, CLL^{+}22b, CL24, CCCC21, COSS21, CLB^{+}23, CEMRMB^{+}24, DGG22, DC20, DAN^{+}23, DLX24, ETB25, EAGV20, FY24, FFE20, FGPPGG22, FGC^{+}24, FS21, For22, GR25, GEGRdL24, GMM^{+}20, GGZRAHVS25, GPM23, GBC21a, GBC21b, GY23, GMRSP23, GOH^{+}20, HW22, HSGNMGC21, HNMSG21, HSHS24, HHL^{+}24, Isl25, JY20, JLY23, JWX^{+}20, JHL21, Jon21, JKKT20, KV20, KCB24, KJE21a, KJE21b, KLK23, KdBK25, KOW^{+}25, KF23, KABT21, LHW21, LWH^{+}24, LLW22, LXW^{+}21, LLL^{+}24, LTJ^{+}21, LJRHH21]. analysis [Ma20, MR24, Mar21a, MBORGN24, MF20, MGF20, MCKZ21, Mur21b, NCF25, ORSB22, PLLJ24, PHP25, PCSS^{+}21, QTD21, RZ21a, RS24, RHL20, SRvNN21, SB25, SC21, STIS20, Sha20, SGK21, Shi20a, SKL^{+}20, SJ21, SWH^{+}23, Smy24, Sol24, SLGAG24, SPT^{+}20, SCY24, SADP25, SLS^{+}20, STF21, Tar25, TJ21a, TMG24, TMM^{+}23, TMS^{+}22, TCT24, TPG^{+}23, VMC^{+}21, VHCJdYC^{+}24, VVR^{+}22, WFR23, WZZC21, WSD21, WW22, WZL25, WMS^{+}25, WL24a, XSCL21, YHU22, YT20, Yu22, YWYL22, YYD23, YZL24, ZLG^{+}24, ZCCQ22a, ZT21, ZZD^{+}23, ZWC^{+}23, ZLH21, dKMK21, dPJ21, DFC^{+}20, WYD^{+}22]. analytic [HF20]. analytical [Min24, TJ21b]. analyzed [FLC^{+}23]. Analyzing [EMS^{+}23, RBA23, VRBBM^{+}24, YU20, ABRA20, Chr20]. Anambas [AGSI22]. anchor [AKL^{+}22]. anchored [DPB^{+}23]. anchoring [GBFMT23]. Anchors [MUR^{+}21a]. Anchoveta [CDG^{+}21, DGH25]. anchovy [CC21, DGG22, SWI^{+}22, UISM^{+}23]. Andalusia [CRCP^{+}23]. Andaman [TJS22]. Andreas [Smi23b]. Andres [PCCOS24]. anemone [YNZ22]. Angeles [AKF^{+}25, WEF^{+}21]. Angler [BST23, CD24, GCM^{+}23, CLL20, GS21]. anglers [SSA^{+}23]. angling [ES22, GS21, QCM^{+}20, WDC^{+}20]. Anglophone [BDVR25]. Anguilla [FRDK24]. animal [HACCM22, SRLLG23]. Ankobohobo [PLLJ24]. Anne [CJ21]. Annex [VSV^{+}23]. annual [DDAB^{+}23, ZTA23]. anomalies [BK23]. answer [HMG^{+}21a]. answers [WLCC23]. Antarctic [BAAH24b, GB21, AM22, ADB^{+}22, BL23a, CMC22, CBB^{+}24, DKW20, DOBE^{+}25, lTS25, Gar20, Hon21, Liu25, TLYC20, TH22, Tra23b, TFJR25, VDOE^{+}24, WS24, ZZMG^{+}24]. Antarctica [TBH^{+}21, BAAH24a, BBK^{+}21b, lTS25, KGW20, SB20, Tod25]. Anthropocene [AHB^{+}23, BAF^{+}23, Ber24, Dun21b]. Anthropogenic [YSEK22, ASG^{+}24, FMGV20, MNGM24, MG22, YG20]. Anthropopause [YNZ22]. Anticipated [SR22, SLS25, TKP23]. Antigua [SWF^{+}24]. antiqua [BRR^{+}22]. Anzac [Dun24]. AOPS [Laj21]. Aotearoa [HNB^{+}25, LLT^{+}20, MAR^{+}20b, PCN^{+}24, RRM^{+}24]. apparel [VVR^{+}22]. appendix [Sel20, HS23]. applicants [BI21]. Application [FFE^{+}24, GCG^{+}21, POM24, VRBC^{+}21, ADK^{+}25, BMB^{+}24, CMBB^{+}25, FWSS^{+}25, FAC^{+}23, GICM25, HJLN24, JH24, KGL22, MXZ23, MFRQ23, Oan22, SX25, Shi24b, TOAPJ21, Zhu23, AE21, CBAPM20, SKL^{+}20, SRSGMGG21b, XWL^{+}20]. applications [AdSMS21]. applied [ABVPM20, dlFGFFR^{+}20, SRSGMGG21a]. Applying [BLM^{+}25, JL21d, QLU^{+}23, VBH^{+}23, WWC23, YJG24, QUK21, YCF24, DAA21, WHPF20]. appraisal [CK23, COM20, HF20, HAsR24, LX24, TK23]. approach [AHBAKH24, AMK^{+}21, AVF^{+}23a, Alf20, ASM21a, ACG21, BRR^{+}22, CBO^{+}21, Can25, CMR21, CPBO20, CD25, CTH^{+}21, CAM^{+}23, CAC23, Cha24a, CMWZ24, CGGA20, CTH^{+}23, CSH23, CHB23, EMN^{+}24, FGK^{+}20, FMCT^{+}21, FFD^{+}23, GSGOP^{+}21, dlFGFFR^{+}20, dHJTGO21, GVB^{+}25, GMG^{+}23, GJY23, GR20b, GSB20, GML^{+}24, HvT22, Han20, HAsR24, HKZ25, Hor25, HGL^{+}20, IO23, JXWL25, KA24, LMC^{+}23a, LZL23, LSC25, LKE^{+}22, LFVCF24, MGS21, MHT^{+}24, MSBA22, MPN^{+}24, NGP^{+}21, NSPRSS20, OKR^{+}20, Osm23, OAM20, PVH^{+}21, PBT^{+}20, PMAVM20, PS21, RBB^{+}24a, RKS^{+}23, SWI^{+}22, Sep25, SEG^{+}22, SOLN21, Sta23b, THP22a, UE25, VGS^{+}20, VRBC^{+}21, VBC^{+}20, WH20, YHX24, ZMR20, ZCG^{+}23, ZGY21, JL21d, LPMS24]. approaches [APGWTM^{+}24, BDG^{+}25, BB20, BEH^{+}21, DSL^{+}21, GCS21, HCDN24, Joh25, KAN^{+}23, Kul23, LRD^{+}22, MC25, MM23, NCD23, OZF^{+}25, QLU^{+}23, QRK^{+}23, SG20c, THF^{+}25, TFMT^{+}23, VBH^{+}23, WWD24, WCF^{+}21, YJG24, ZWJvL20, dSOW^{+}21]. appropriate [GBDC24, Kai21]. approval [Kak20]. April [Ano21a, Ano22a, Ano23a, Ano24a, Ano25a]. Aquaculture [LG24, MF20, OOGA22, TVN23, YXX^{+}24, Åm21, AHBAKH24, AHENNA^{+}24, AOO^{+}23a, AO20, AELR^{+}20, ASG^{+}24, BVvKL21, BE20a, BE20b, BBFA^{+}24, BDLO23, BBGHR23, BÅ25a, BABC20, BBO21, BLS21, BTE21, BHK^{+}25, CFM^{+}21, CMR21, CGGA20, CJH24, CVA22a, CVA22b, CFA^{+}24, CWMM^{+}21, CSRG25, Ero22, EMS^{+}23, FSZ25, FBC^{+}23, GVB^{+}25, HJLN24, HS20, Han21, HHB^{+}21, HCSBM^{+}23, Her22, HSHS24, IKB21, JFB^{+}24, JY20, JKV^{+}22, KCCC^{+}21, Kob22, KBD^{+}20, LGF24, LFPBD^{+}20, LAF^{+}24, LLL23, LFVCF24, MGS21, MCPS20, MVSP^{+}21, MMF^{+}25, MHF^{+}25, MM22, MMNC22, MBC^{+}24, MFF^{+}23, NSZ^{+}21, NVP^{+}23, OOSG21, POM24, PŚ22, RMF^{+}21, RFG24, RF25, RBNE20, SYA25, Sah22, SZ20, Squ20, TCS^{+}20, TFF22, TTL^{+}25, VMD^{+}24, VSPMS^{+}22, WOR^{+}25, WFG22a, WMKG21, YHU22, YK20, YTG^{+}23, KBP24, TCF^{+}25]. aquarium [AP23, SC21]. aquatic [BRM^{+}21, BLM^{+}25, MCBN22, QL24, Sol24, SLO^{+}24, SN23, SSP^{+}21b, VGS^{+}20]. Arab [Gul20, Squ20, Cha24a]. Arabian [DMS^{+}24, SONB^{+}22]. Arabian/Persian [SONB^{+}22]. Archipelagic [VKS23]. Archipelago [HF20, MIH^{+}24, ASM21a, ABH25b, CWB^{+}20, DSD^{+}24, HS24, TOAPJ21]. archipelagos [GMGLV21]. architecture [Kim24]. Arctic [Mol24, VYB20, WFP25, AHN25, Bal22, Boy21, BVT^{+}21, BTV^{+}22, CDRS22, De 22, Dyc24, FFE20, Geh23, Gun21, GGL22, HLLK22, JHL21, KKA^{+}21, Kra22, KY20, Laj21, Mol24, Osm23, Øst24, PH24, Pri22, QLZ^{+}24, RCT20, Tod22a, VYB20, WA24, WF25, WFP25, WCT^{+}20, WSDB20, YK21, ZXGL21]. Arctocephalus [PBH^{+}23]. ARDL [UE25]. Area [Cho22, FMLMM23, GBC21b, GDY25, HSSM24, LTJ^{+}21, ML22, RCM^{+}22, RJAT23, TJ21a, TCT21, VMC^{+}21, WJLSH21, ACA^{+}21, AHAB^{+}22, ASM^{+}21b, BBB^{+}24, BCR25, BCJ21, CAGBM22, CJ22, DLJ^{+}20, DKW20, DMHS25, DB20, DSC^{+}23, EAA^{+}22, Fei20, FWSS^{+}25, GBC21a, GMC24, HLW^{+}23, HPPH22, HJ25, HLM^{+}20, HCT^{+}21, Jae20, JPM^{+}23, JL21c, Kim21, KTC^{+}22, LLA^{+}25, LZRC23, MM21a, MCRPC20, MG25, MNS^{+}24, OSBA24, PLLJ24, PBH^{+}23, RP23a, Rus20, SAO^{+}21, SMLRB^{+}21, SPHO23, SJ21, SDG^{+}21, SDG^{+}22, SCC25, SB20, TLYC20, Tes20, TBCB25, Tra23a, VWFW22, VLG^{+}24, WF25, WD20, WMRB22, YT20, YCL22b, YC21b, dS22a, BAAH24b, CGMH24, FMGV20, Fri24, GABA^{+}22, PEM^{+}25, RHG^{+}25, RP23b, SOT^{+}22, VBd^{+}21, VCC^{+}21, War20, dS21]. area-based [DMHS25, EAA^{+}22, KTC^{+}22, WF25]. Areas [BTN^{+}24, Has23, HSGNMGC21, KJE21b, MFA^{+}24, MFG^{+}22, SFGEPC22, ABC^{+}25, ACA^{+}21, ASC25, BCW^{+}25, BBB25, BDM20, BPE^{+}25, BD23, BAAH24a, BWV25, CBO^{+}21, CHH^{+}25, CMGF20, CCBB^{+}25, CLRW22, Cha21, CMWZ24, CLU^{+}24, DDAB^{+}23, Col20, CWR20, CMOJ^{+}24, De 20, Dua22, DSV^{+}25, EV22, FDGM24, FTGD^{+}20, FMS^{+}22, GFD21, GNJR24, GM25a, Gar20, GCMT20, GBLBV20, GCL21, GBFMT23, GBC21b, GWAPT21, GMVM23, HJ21, HLW^{+}23, HM22a, HCDN24, JWX^{+}20, JL21b, KFM^{+}20, KDM^{+}20, KVL^{+}23, KOCL23, cL22, Liu24, MSBA22, MCR22, MPSF20, MMO^{+}24, NNM^{+}21, ONA21, OKH^{+}24, OZF^{+}25, OKR^{+}20, OCCSC23, PBG23, Pha20, dlÁPAFGO20, PL23, PdBA^{+}25, Pri22, PKMT23, RKC^{+}24, RGK^{+}20, RPY^{+}24, RKS^{+}23, SON21, Sol23a, SO20, SZ24, THT^{+}23, TBH^{+}21, Tod20, TEA^{+}22, TJ21b, TCT24, VHD20, VAG^{+}24, VBC^{+}20, WZZC21, WSD21, WHAD^{+}25]. areas [WGJ^{+}21, YG20, YCL22a, YSWW21, YD22, ZRR^{+}25, ZJ22, ZPX^{+}22, Zha22, AIG^{+}20, ÁFFSC20, AVF23b, BRKS25, BvBN^{+}20, BEH^{+}21, CCS23, Gel25, IMB^{+}24, LGLL24, MGPR^{+}25, NJBM^{+}23, ORSB22, QTBL23, RA20, RCM^{+}22, Teo23, TCD^{+}20, WB24]. Argentina [SV22, TDC21]. Argentine [SV22, SVG24, ZCC^{+}24]. arguments [BE20b, Sel20]. arid [AS23b]. armed [CLS20]. arrangement [Li23b, VJ24]. arrangements [AGSI22, Cla20, FFD^{+}23, HIP^{+}21, Kra22, MMSB^{+}20]. Arte [ACV^{+}21]. Article [WL20, Fla21, Guo20, Shi24a, ALLS23, Shi23a, WL20]. Artificial [OVSVPP^{+}22, TDP20, KBB^{+}25]. Artisanal [DRCdMCJ21, GRPRVG23, MSHK25, SBFHD^{+}23, APGWTM^{+}24, AACdlT23, ACV^{+}21, BCBS22, BBV^{+}25, BMK^{+}25, CSMMC^{+}25, CDNS23, hBDF22, DSD^{+}24, FA24, FLB^{+}21, GSAF21, GWAPRPRU22, JGWM^{+}23, JTR^{+}20, KWS20, KLSG21, KOW^{+}25, LASG^{+}25, MHS^{+}25a, MHS^{+}25b, CSdMR^{+}24, MS21, MH20b, OCD^{+}23, PRM^{+}20, PCCOS24, PV25, RPK23, RMBR22, SPG20, dSF24]. artistic [WPvO24]. ASC [GVL22]. Ascension [ASM^{+}21b]. ASEAN [SF22]. Asia [War23, BBC24, BRM^{+}21, Dav25, DAN^{+}23, FOKM23, JF24, MGDJ22, PGP22, PPvD^{+}24, PK20b, QK24, WM22a, YS21, AFHM23]. Asian [HH21, PKB^{+}23, GLAVL21, JL24, SYA25, Ton22b]. ask [BAJR24]. asked [RKC^{+}24]. asks [Has23]. aspect [Cha22]. aspects [BN20, CMC22, MLVT21, MSSS22, dJdMCL^{+}24, dMVJdAdSM25]. aspirations [GAD^{+}24]. assemblage [OKFN24]. assemblages [WB24]. assess [AFRFGP25, CSVC23, DDD22, EAGV20, GSAF21, GKP^{+}22, MLJ^{+}22, MLVT21, PCD^{+}23, QLU^{+}23, TLT^{+}22]. assessed [ASV^{+}21]. Assessing [AFR25, ADM^{+}25, AR23a, BJGG^{+}24, BUR^{+}22, BHKK23, CMR21, CACJR^{+}21, CA20, CWB^{+}20, CAM^{+}23, Cho22, DME23, EAAG21, FGPPGG23, FMGV20, FWSS^{+}25, For22, GAD^{+}20, HGH^{+}20, KDHL24, LK24a, LASG^{+}25, MBL24, NH20, NGT^{+}23, ONA21, OKH^{+}24, PTS^{+}25, PEM^{+}25, Pra20, QTBL23, RPK23, SSP^{+}21b, WRL23, WLS^{+}25, WML25, WABR^{+}22, ZABI25, dOPM^{+}20, FAGS21, GBDC24, MCS^{+}25a, MGBS24, WNMS^{+}24, AAKI21, ABM21, DKKS21, LTRA20, OPL20, Szy20, WGP^{+}21]. Assessment [AGM^{+}22, GAA^{+}20, HM22a, HF20, Koj23, LDL^{+}23, LFPBD^{+}20, MFC^{+}24, AANM^{+}22, ASC25, AP23, BB21, Bov22, BSK^{+}20, CAC23, CTCD21, Cha22, CJL^{+}23, CL20, CMSH25, CCOP25, DOBE^{+}25, DBS22a, DGIS^{+}22, Fuj21, GBM^{+}25, GSG^{+}24, VKL^{+}25, GJY23, GMCMJA^{+}21, HJLN24, HYN^{+}23, HBR21, HBR^{+}23, HAsR24, HMY^{+}23, HSHS24, JL24, KKAO^{+}25, KHH^{+}23, KMJ24, KM25, LLCH21, Lov23, MPD^{+}23, MYLV24, MVR21b, MPL^{+}23, MN23, Moh23, MIH^{+}24, Oan22, OKR^{+}20, OMC^{+}22, PDD^{+}25, PLWO24, PSMR20, PWL^{+}23b, QUK21, RXN23, SAO^{+}21, SC21, SNK^{+}20, WFK^{+}24, WHL22, WCL25, YTBM^{+}22, ZTZ^{+}21, ZCA^{+}23, PS21]. Assessments [SSAAB^{+}24, BBB23, CTH^{+}21, DQJ23, FOH^{+}20, KAA^{+}24, SSOP^{+}21, SDG^{+}21, SDG^{+}22, TFMT^{+}23, CDC20b]. Asset [HBCA^{+}24]. assets [BS22, FSKJ23]. assignments [DSC^{+}23]. assistance [Sar23]. associated [DWC^{+}20, MHS^{+}25a, MHS^{+}25b, PMAVM20, QWBY21, SKS21, SKW^{+}22, ZCA^{+}23]. Association [CJLJ24, LA23, CLL^{+}22b, SL23]. Associations [HVN21, Kay25, Sta25]. assumptions [LLM^{+}23]. Assuring [BBO21]. Asturias [dlFGFFR^{+}20, SPG20]. Asymmetric [JTdH25, ZLGX20]. at-sea [EHM^{+}20, EHM^{+}22]. Atacama [CBO^{+}21]. Atlantic [BdSL25, BUM^{+}24, CMI^{+}22, DFP^{+}25, FMGV20, MFT^{+}22, Mar20a, MF20, MAG^{+}22, RHST21, dOSTB^{+}20, WCF^{+}21, AIHR^{+}25, ÁFFSC20, Bjø23, CPV^{+}21, CCGM21, CNB^{+}23, CSGC22, DSL20, EB24, FRVB22, FSZ25, FTSG21, GSG20, Gra21, GCM^{+}23, GCC^{+}21b, GICM25, HBHJ20, KLSG21, LCL20, MSGAR20, MMGAA^{+}21, MM20, NOL^{+}23, PCG^{+}23, PBG23, SGR21, SGL^{+}23, Sha22, SNS23, SAS^{+}25b, SBT^{+}23, SGZH^{+}21, SGH22, SBW^{+}21, TCF^{+}25, TFF22, TCD^{+}20, VBH^{+}23, VLG^{+}24, YK20]. atlanticus [GCM^{+}23]. Atoll [YCL22b]. Attack [PB20, HPG20, Los23, Los24]. attempts [OYKBA22]. attitude [GBB^{+}21, KCCC^{+}21, KHMM20, LM24, RPK23]. Attitudes [PC20, AAKI21, CGAF25, CVA22a, DDD22, HWM22, Hos24, dOJCSLdSB21, ONN^{+}20, RKC^{+}24, RSH^{+}25, SAB^{+}20, WJB^{+}20, YTÇ22]. attitudes/behaviours [DDD22]. attitudinal [NGP^{+}21]. attractiveness [BGQ25]. attribute [TDC21]. attributes [CSVC23, CCOP25, OMC^{+}22, QWBY21]. auction [AMIM21, KV20]. Audience [WNJ25]. audiences [WWD23]. audio [WPvO24]. auditory [VDOE^{+}24]. augment [SCCW21]. augmented [UE25]. August [Ano20a, Ano21b, Ano22b, Ano23b, Ano24b, Ano25b]. Australasia [BBBS22]. Australia [BLL22, BR22, BBHN20, BBGG22, CCK^{+}21, DPP^{+}20, Dun24, ESD24, FFD^{+}23, HPG20, HWM22, HMG^{+}21a, HGH^{+}20, IMB^{+}24, Jon21, MYLV24, MCW^{+}22a, RKC^{+}24, RHC20, RDD^{+}21, SGLF^{+}23, SHAC23, THF^{+}25, DR22, WWG^{+}23, WWG^{+}25, YT20, YW24]. Australia-wide [HMG^{+}21a]. Australian [KB20, LHW21, MCW^{+}22b, MKS21, OAM20, OAM21]. Australianised [Dun24]. australis [PBH^{+}23]. Authorities [SCC25, GLSVMB25]. Authority [Sin21, ZRR^{+}25, BMM^{+}20, JPM^{+}23, WMM24, AR22, BHKJT23, BRM20, CM25, GSM20]. automated [KP24, SKL^{+}20]. automatic [AHT25, TMBM^{+}25]. Autonomous [DBZ24, ATM^{+}24, KZR22, Kle21, MNHS25, Ste23, dKMK21]. autonomy [Sar23]. autoregressive [SGK21]. availability [MAA^{+}24, MMNC22]. Available [RJAT23, MH24]. Aveiro [BASM22]. average [WL24a]. Aviv [PC20]. avoid [McC20]. avoidance [KP24]. Avoided [BRKS25]. avoiding [Sas25a]. Awareness [CZX24, BFdSdFL^{+}25, DG24b, DDD22, GCV^{+}21, HT23, HWMO20, Lev20, OCD^{+}20, Pra24, WNBHG24]. Away [JFB^{+}24, dO23]. axes [DAB^{+}25]. Azerbaijan [AB20]. Azores [ASM21a, CVCV^{+}22, CRIMLS24, CFS^{+}24, GMGLV22, MCRPC20].

b [Shi24a]. bêche [PSG25]. bêche-de-mer [PSG25]. Back [Bjø23, CBCA25]. background [LCL24, Sas25a]. Badouzi [CCCC21]. Bags [NW20]. Bahamian [Doe21]. Bahia [MN21, SBFHD^{+}23, GBAECCRM25]. bait [LMC^{+}23a]. baited [HMG^{+}21a]. Baja [GBAECCRM25, CSMMC^{+}25]. balance [AED20, LZL25, Pre21]. balanced [GVB^{+}25]. Balancing [EV22, HKZ25, HZC^{+}24]. Balearic [GBFMT23]. ballast [KLS22, Lia25, LCL24]. Baltic [PKH22, EVPMKP22, HJ25, Kiz24, LEVET21, MCKZ21, PS21, RCM24, RKS^{+}23, SLK^{+}23]. ban [BSI^{+}21, CHEC21, CD23, CLB^{+}20, DSB^{+}25, GTA^{+}20, KGP^{+}20, Lev20, Pen22, PBT^{+}20, PBG^{+}22, vLM22]. Banda [EABA^{+}22]. bane [MCBN22]. bang [KCH^{+}25]. Bangladesh [ABC^{+}25, ABH25a, ANC^{+}25, BUR^{+}22, BSI^{+}21, DGI^{+}24, DAAP21, HHB^{+}21, HBP^{+}21, HSN^{+}21, INW20, IKB21, IRK^{+}21, Isl25, LMR25, PPK23, RKR20, SRR^{+}21, SSP^{+}21b]. Bangladeshi [ASKI20]. Bank [APH24, GÁPSAP^{+}23, GÁSVPS^{+}25, WCF^{+}21, CLS^{+}23, Ken20]. banks [vLS24]. bans [CLB^{+}20, WLW^{+}22]. Barbuda [JMB^{+}20, SWF^{+}24]. Barclay [AFHM23]. barcoding [BdLAdC^{+}20]. Barents [BFMS21, DD22]. bargain [BHK^{+}25]. barnacle [GRA^{+}22]. Barrier [DR25, MHC^{+}25, DR22, Wat22, YHU22]. Barriers [KMS20, Lot22, MKM23, AHM22, BDHI21, CWMM^{+}21, ETB25, EFCU25, FSF^{+}21, FHBS23, KSS23, LHTR23, NCD23, OBNK24, SDI^{+}20, SBP24, SMB^{+}25, XZ23]. base [GSM20]. Based [HJPL23, Öİ24, WB24, AHENNA^{+}24, AKFY22, ABH25b, ASM21a, ACG21, ALF^{+}22, AGA^{+}21, BRSC^{+}24, BBV^{+}25, BMMF22, BDG^{+}25, BV21, CPV^{+}21, CLS^{+}23, CC21, COJE24, CMR21, CMT^{+}23, CdLN^{+}21, CSLGFV20, CS21, CLL^{+}22b, CJL^{+}23, CL24, CLU^{+}24, CBB^{+}24, CGAF25, CCCC21, CES20, Col20, CSVC23, CSL21, DMHS25, DMK24, DRY22, DBH^{+}21, DG20, EAGV20, EAA^{+}22, FSRC25a, FSRC25b, FMTA^{+}25, FHBS23, GFD21, GPG^{+}21, GLAVL21, GDB^{+}24, GSB20, Gug24, HFD^{+}20, HLW^{+}20, HGG^{+}22b, HC25, HVN21, HGG22c, HC21, HSHS24, HLO^{+}21, IRK^{+}21, JKV^{+}22, KS23, KOBM24, Kar22a, KMM24, KP24, KTC^{+}22, LLT^{+}20, LLW22, LYQM21, LJ22, LH25, LGP25, LXW^{+}21, Liu23, LLS^{+}20, LFVCF24, MR24, MUR^{+}21a, MLB^{+}25, MSB^{+}25, Moh23, NSPRSS20, Okş23, OCCSC23, OVSVPP^{+}22, PBT^{+}20, PGSS22, PGA^{+}20, PS21, QYD20, QTD21, QLU^{+}23, QRK^{+}23, RZ21a]. based [RHG^{+}25, RF25, RR20, RXN23, RXX24, RSFvH^{+}22, RMU^{+}22, RDM^{+}20, RKS^{+}23, RDD^{+}21, SVF20, SVP^{+}23, STKC22, TM23a, THP22a, TCT21, Ton24, VRBC^{+}21, VBH^{+}23, VJÓK25, WLQD21, WHPF20, WLX21, WYD^{+}22, WHL22, WF25, XHZ^{+}21, YWY20, YYD23, ZZH21, ZCCQ22a, ZY21, ZCS^{+}23, Zhu22, ZSWJ23, dOPM^{+}20, vTHS21, QS21, FMGV20, LPMS24, OWC^{+}24]. baseline [DLJ^{+}20, GRS^{+}23, CSdMR^{+}24, MBC^{+}24, TPG^{+}23, Wat22]. baselines [EMH^{+}23, VKS23]. basepoints [VKS23]. bases [Kuł22, YYM25, SV22]. Basin [JZW21, SBP^{+}25, TCF^{+}25, HGQJD24, Kas21, SV22]. basis [CC23, CDY24, TML^{+}20, YLSC23]. Basque [RHO24]. bass [CAM^{+}23, CMS^{+}25]. bathers [AMLC^{+}23]. batoids [PMFR^{+}22]. Batten [SMF^{+}23]. Bay [AAKI21, Bra22, DSQ^{+}22, HM21, Jon21, KJE21b, MF20, MMNC22, MGNP25, SGPF21, SHAC23, YTG^{+}23, ADR^{+}21, DA24, DMT25, Gir22, IBC^{+}20, PGA^{+}20, PMT^{+}20, SL20, SJ21, THGR^{+}20, UISM^{+}23, WLX21, dZ20]. Bayes [KA24]. Bayesian [PBT^{+}20]. BBNJ [Bal22, CJK24, Cla20, CM25, De 22, Dua22, Fri24, GNJR24, HHD20, HGL^{+}20, Kim24, Kra22, LV23, LX24, LZ24, Liu24, Lot23, MH23, MS20, MFH^{+}20, Shi20a, Tan24, TG25a, Wan23a, WS24, WW25, ZL24, dlC24]. Be [HNA25, BdLAdC^{+}20, BCR25, BRKS25, CSM^{+}25, CMWT24, Lev20, Pha20, SWF^{+}24, Sin21, Sol23b, Ste23, WL24a]. Beach [BISSBdlMM22, CDP22, SAB^{+}20, Cat20, CLC21, CMSH25, KL21b, Lev20, PBG^{+}22, TNE^{+}24, YSEK22]. beach-activities [YSEK22]. Beach-user [SAB^{+}20]. beaches [Che21a, CMSH25, SMC^{+}25, WFR23]. beachgoers [VRBBM^{+}24]. beacons [BMRM23]. bear [Geh23]. beauty [CDP22, JXWL25, WFR23]. because [SCF^{+}22]. becomes [AH21]. Becoming [AO20, Shi20b]. beds [KWY^{+}22, MR24]. been [Gal22, HWMO20, ZD20b]. before [ABC^{+}24]. begins [TM23b]. behavior [DNR21, Fen23, FR24, HZ23, dSJNS20, LLJZ22, MLM23, SKS21, VRBBM^{+}24, WKO^{+}22, WFG^{+}22b, XXLW24, XLW24, DLYM22]. Behavioral [BC21, Muk21, GPM20, NGP^{+}21]. behaviors [Hos24, Ngu24]. behaviour [CBR21, LYQM21, SFGEPC22, WKN^{+}25, YBW20]. Behavioural [DOBE^{+}25]. behaviours [DDD22, PC20]. behind [Lee24, YL21]. Beibu [TLT^{+}22, WLW^{+}22]. being [AVF^{+}23a, BBB25, BCG^{+}24, GCF^{+}23, GGZRAHVS25, MNGM24, MTC^{+}23, RGH^{+}22, SCF^{+}22]. beliefs [Fro21]. Belize [AGK22, Mur21b]. Belizean [GABA^{+}23]. Bellingshausen [BAAH24a]. Below [NSN22, Gul20, MH20a]. Belt [HH21, TBPC20]. beluga [TMBM^{+}25]. Benchmarking [WD23]. Beneficial [PGK22, GOH^{+}20, SOK^{+}22]. benefit [AH23, BTE21, EECE^{+}24, LWX^{+}20, MR24, MB21, SC21, VDN^{+}24]. Benefits [BBB^{+}20, KTC^{+}22, PSTW21, BBS^{+}21, BvBN^{+}20, BS22, DR25, DD22, JKKT20, KJS^{+}22, LIK^{+}24, MRKM23, MCS25b, PFG^{+}21, QUK21, SBDC24, Squ20, TSA^{+}22, TFH^{+}23, TCS^{+}20, WWG^{+}23, WWG^{+}25, WLCC23, WBPG20, dlC24]. Bengal [CDP22]. Benin [GSG^{+}24]. benthic [DGIS^{+}22, DS22b, FARBV^{+}24, GSAF21, NJBM^{+}23]. benthopelagic [CDC20a]. Bering [BRKS25, Tod22b]. Bermuda [WLRH^{+}20]. Bern [SAS24]. berteroana [ÁTGCE^{+}25]. Best [BKW^{+}24, BPJ^{+}24, KCH^{+}25, Squ21, ABM21, CN24, GG20, GMG^{+}24, KSO^{+}25, MHT^{+}24, MF22, MWM^{+}23a, MH24, OGM24]. best-worst [MHT^{+}24]. Better [WNBHG24, CGB^{+}25, CAHR^{+}24, FMLMM23, GMRSP23, OBNK24, SOO20, TRL^{+}25]. Between [hBDF22, ABM21, BCN23, BDDE20, BMK^{+}25, BMB^{+}21, CJLJ24, CSL21, DC23, DAAP21, EHÓ25, GPA^{+}22, GR25, GLAVL21, dHJTGO21, GdCC^{+}25, GCL21, GVH^{+}23, GGL22, HTP^{+}23, Hor23, HWY22, IGO21, JBGC25, LJ23, LZL25, LHPJ23, LZRC23, MZ24, Mar21a, MMG^{+}25, PDB^{+}23, PSPR20, PL20b, PGM^{+}24, PdBA^{+}25, PKMT23, RTMCVMDE22, RHST21, SL23, SLB21, SSA^{+}23, Shi23a, Shi24a, SS23, ST24, SSMG21, Tar25, TRL^{+}25, TG25a, TH22, TRN^{+}24, Tra23a, TFJR25, TJ21b, VGK23, WS24, WB24, Zha22, ZX24, ZY23, ZCC^{+}24, vdGD21]. Beyond [BWV25, CMD^{+}20, FDF^{+}21, Has23, Kıl24a, MTN23, SDG^{+}21, SDG^{+}22, Teo23, TM23b, TCD^{+}20, VM21, WBPG20, dZ20, AS23a, BLL22, BBO21, CMOB^{+}20, Cla20, CWR20, CM25, DMNT20, EAA^{+}22, FARBV^{+}24, FMS^{+}22, GNJR24, Gar20, GAN^{+}25, HJ21, HM22a, HGL^{+}20, MDJ^{+}22, MH24, MCR22, NRF21, OKR^{+}20, PSTW21, Pri22, QK24, SBM^{+}25a, TTA^{+}20, TD21, TMDN23, VHD20, VBC^{+}20, WF25, WPWM^{+}22, WGJ^{+}21, YG20, ZJ22, vNdB25, Sam22]. Bhitarkanika [KHMM20]. bias [CBH^{+}24b, CBH^{+}24a, CBH^{+}25, PBG23, RZ21b, SCM^{+}25, TDP20, WFR23]. biased [KSB^{+}23b, Ren21a]. Bibliography [Smi22, Smi23b]. bibliometric [AN25, BBP25, BMR23, CTCD21, DLX24, HHL^{+}24, RXN23, TMM^{+}23, WHL22, WNDZ23]. bibliometric-based [RXN23, WHL22]. bicinctus [YNZ22]. bid [SKE22]. Bigi [DB20]. Bill [Dbo22, NW20]. billfish [KWS20, KOW^{+}25]. BIMCO [Bov22]. binding [CWR20, HvT22, LR20, Pri22]. Bio [DAAP21]. Bio-economic [DAAP21]. Biobío [FMTA^{+}25]. biocultural [CBCA25]. Biodegradability [DFF23]. Biodegradable [ZMM^{+}23, EMH^{+}23, MSZ^{+}23, MZT^{+}23, SGL20]. Biodiversity [Has23, BG25, BLL22, CM25, DMNT20, EHÓ25, FWGM24, FFGJ22, GNJR24, GABA^{+}22, HJLN24, HZC^{+}24, HGL^{+}20, MDJ^{+}22, MH24, PKB^{+}21, PT23, Pri22, RSS^{+}20, SSAAB^{+}24, TD21, TMDN23, TV21, TFJR25, DR22, WBB^{+}25, YJG24, MTN23, OGS25, TM23b]. bioeconomic [BCBS22, Mot20, Mot22, PFWM23]. Biological [MG22, SLP^{+}22, FMS^{+}22, HM22a, LHV^{+}20, SMF^{+}23, TDC21]. Biologically [MFA^{+}24]. Biology [dSJNS20, GSA23]. bioluminescent [SWW^{+}21]. biomass [MDG^{+}23, PMAVM20]. bioprospecting [WS24, ZL24]. Biosecurity [MBC^{+}24, CH25, KCCC^{+}21, SSDQ20, VWJMAJ20]. Biosphere [MMSB^{+}20]. biota [CDC20a, JHHG23]. bipartisan [ADWP24]. Bird [SEG^{+}22]. birds [BEM^{+}20, Ber23, ZCC^{+}24]. Biscay [ADR^{+}21, DMT25, IBC^{+}20, PGA^{+}20, PMT^{+}20, UISM^{+}23]. bite [SAB^{+}20]. bivalve [BHB^{+}24, NNM^{+}21, PL23, dJdMCL^{+}24]. Bjørn [Fla21]. Black [SSZ^{+}22, AHN25, GEB^{+}21, MJMAA21, QLZ^{+}24, SCD21, TGM22, BMM^{+}24, CMS^{+}25, DGG22, DUYEA21, ES22, MCS^{+}25a, MÇ23, MHF^{+}25, Şan20, SRS^{+}25]. blackspot [LT24]. blanket [CLB^{+}20]. bleaching [DDAB^{+}23]. blind [CN23, NBB^{+}23]. Blockchain [NCD23, How20, Lin25, LFVCF24, MBL24, ZSLY20]. blockchain-based [LFVCF24]. bloom [XLLS21]. blooms [EHLT21, SLMG^{+}21]. Blue [AC23, BBWC21, BE20c, CS21, CMBB^{+}25, Dyr25, HSGS^{+}23, LSOP24, NJW25, PCHD^{+}25, SLW^{+}23, SH24b, Tho22, TFMT^{+}23, AFR25, ARP25, Als20, AHB^{+}23, ANC^{+}25, ANR^{+}22, ABC^{+}24, BAVV24, BLM^{+}25, BGQ25, Bre25, CBC24, CSM^{+}25, DKKS21, DGI^{+}24, DČFW23, DSL20, DPSF25, FFD^{+}23, FAB^{+}23, GMFD25, GdACT20, HJ21, Has20, HBCA^{+}24, HSN^{+}21, HH25, KCB24, Kıl24a, Kıl24b, KNY22, KWY^{+}22, LMR25, Lou22, MMB^{+}23, MMF^{+}25, MFC^{+}24, MLB^{+}25, MCKZ21, MCTZ24, NCF25, NO25, OWC^{+}24, OCZ^{+}22, PSD23, Puh23, QZZX20, Qi22, Qi24, QUK21, QK24, RRM^{+}24, San25, SDI^{+}20, SHPW21, SSA^{+}23, SWHT22, SCB25, SR22, SOK^{+}22, TFH^{+}23, TT22, TAP^{+}23, WXLD21, WLQD21, Yad24, YCF24, AC23, CFM^{+}21, DZZC24, GFL^{+}21, HvT22, HWC^{+}25, IMA22, LGP25, LMMT23, MBG^{+}24a, MFC^{+}24, MY24, MSHK25, SCJ22, SWW^{+}21, WMKG21, ZFGL20]. blue-tourism [GMFD25]. bluefin [Bjø23, GSG20, HWT23, SAS^{+}25b]. bluefish [ES22]. blueprint [HH25, SH24b]. bluewashing [Kıl24a]. board [ASKI20, MRD^{+}20, OVA22, SSDQ20, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano21o, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano22o, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24o, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano25c, Ano25d, Ano25e, Ano25f, Ano25g, Ano25h, Ano25i, Ano25j, Ano25k, Ano25l, Ano25m]. boat [AFRFGP25, CRJ22, Kle21, MAR^{+}20b, RDD^{+}21, SCGC20]. boat-based [RDD^{+}21]. boater [DWC^{+}20]. boating [BAJR24, DWC^{+}20, PHF20, VMC^{+}21]. boats [ÁCT20, WDC^{+}20, BEM^{+}20]. bodies [BLL22, BJ25, Bet25, SHFP21]. Body [GNJR24]. Bohai [LYHL23, LSH21]. bolts [SG20a]. bond [Kıl24b]. bonds [Kıl24a, NCF25, Tho22]. Bonferroni [CMR21]. Bonn [VSV^{+}23]. Book [APH24, AFHM23, Cha20b, De 20, HH21, LPO25, LL24a, Lv20, Ngu23, Niu22, NJW25, Rub24, Smi20, Smi22, Smi23a, Smi23b, SSSL25, Sos21a, Sta23a, Ste25, Tan23a, Tan23b, WQ25]. Boom [BDD^{+}21]. Boon [MCBN22]. boost [Pan25b]. border [AHR24, De 21, FKNG25, GSGOP^{+}21, IO23, LCL24, MMK^{+}22, Tse21]. border-crossing [Tse21]. borders [KMTvT25]. Bosco [Smi22]. Bosomtwi [GAA^{+}20]. Bosomtwi-Sam [GAA^{+}20]. both [CWLD25, VHD20, WWD23]. Bothnian [EVPMKP22]. bottlenose [AVC24]. Bottom [FARBV^{+}24, ADM^{+}25, BMGDS^{+}20, CGB^{+}25, DAA21, DDF^{+}20, DGIS^{+}22, GRSRSN^{+}24, LCAX21, MVR21b, NSK20, RQB^{+}20, SDG^{+}23, SLB21, SC22, ZV20]. bottom-up [SLB21]. boundaries [BDC23, SDG^{+}21, SDG^{+}22]. Boundary [Smi23b, hC24, Fra23, GG22, LLT^{+}20, MC25, Øst20, PKH22, PK20b, RB22, SHPW21]. boundary-crossing [LLT^{+}20]. box [GEB^{+}21, Tod20]. brasiliensis [dJdMCL^{+}24]. brave [KFP20]. Brazil [LFLF22, SBP^{+}25, dOSTB^{+}20, AFMdSdTdCC22, AML^{+}21, BNF21, BdLAdC^{+}20, BMK^{+}25, CMI^{+}22, dFVCGdS21, DMd^{+}20, FSF^{+}21, GCMT20, Gon25, HSM^{+}23, JAdS^{+}20, dSJNS20, KHdSM^{+}22, MG25, MIMS^{+}20, MN21, MMS24, ORSB22, PMC24, PBH^{+}23, dSRKTC24, RCM^{+}22, SdACdSLdAS24, SA21, SBM^{+}25a, SBFHD^{+}23, SSB^{+}22, dSTHL^{+}24, dS21, dSF24, dSdLSGB22]. Brazilian [CSGCD25, CCOP25, GCF^{+}23, GdACT20, MM21a, Mar20a, NH20, NAP^{+}21, NBB^{+}23, PDS24, SSZ^{+}24, SS23, VBd^{+}21, dS21, dS22a]. break [MMS24, WCF^{+}21]. Breaking [AC23]. breeding [OKH^{+}24, WCT^{+}20]. Brexit [BFMS21, BM21, DGS^{+}24, EB24, FDF^{+}21, Fox22, KFP20, ZX24]. bridge [FS21, Osm23]. Bridging [GLB^{+}24, TMAS24, VLL^{+}24, WMRB22]. Brief [CMWZ24]. bright [CSGCD25]. Bringing [BOBMG25]. brink [Bjø23]. Bristol [Bra22, DA24]. Britain [TG25b]. British [DBS^{+}22b, SSB^{+}23, WJLSH21]. broad [HMG^{+}21a]. broad-scale [HMG^{+}21a]. Broadening [SS21a]. broader [CMS^{+}21]. Brochmann [Fla21]. Brooks [Sos21a]. Brown [VMV24b, GABSD21]. Brushett [Rub24]. BRUVs [HMG^{+}21a]. buck [KCH^{+}25]. buds [PBG^{+}22]. Bueger [Ste25]. build [Jon21, MZ24, PALB20, SCLM^{+}21]. Building [DB20, How20, MPN^{+}24, PBC^{+}24, RJAT23, VAG^{+}24, XT23, AKFY22, AAA^{+}21, ACG21, FMS^{+}22, PTCC24, TYS24, TZC23, Ton24, VHD20, YCL25, WNBRSG22]. Bulgaria [SSZ^{+}22]. Bulgarian [SRS^{+}25]. bullying [MLE^{+}24, ÖB22]. bumpy [CMC24]. bundles [WTO24]. bunkering [HPS23, KGW24, MBL24]. Busch [Niu22]. Business [LEVET21, BDR20, LH25, MLVT21]. business-oriented [MLVT21]. bust [BDD^{+}21]. buy [CDB20]. buy-in [CDB20]. by-catch [MSGAR20]. by-products [LAF^{+}24]. Bycatch [AIHR^{+}25, GCB^{+}23, RQB^{+}20, WLW^{+}22, ADM^{+}25, BKW^{+}24, BdSL25, BEM^{+}20, BDG^{+}25, CPBO20, CLB^{+}23, DER^{+}21, DBB^{+}22, HBHJ20, JSW22b, MHS^{+}25a, MHS^{+}25b, MM23, MWB^{+}21, SBW^{+}21]. Bycatch-neutral [GCB^{+}23]. bycatches [BUM^{+}24]. Byron [ESD24].

CA [AS23b, KL21b, WWS^{+}25, NHD23]. CA-Markov [AS23b]. Cable [Ros25, Can25, Pan25b]. cabled [AIG^{+}20]. cables [Dav25, Gov25, TG25b, Wat25]. Cabo [HSGNMGC21, HSGNMGC21]. cabotage [CC22]. Cacão [BdLAdC^{+}20]. Cadiz [GMPT^{+}20, GTA^{+}20]. cage [HACCM22, SMHPMO^{+}21]. Cages [DMS^{+}24]. calculation [CA21, VGK23]. Caledonia [LDAC20, DPP^{+}20]. Caledonian [GA20]. calendars [HDCH21]. California [CSMMC^{+}25, GBAECCRM25, ABDRV^{+}23, AKF^{+}25, ALLR^{+}24, BCMCM21, OKS^{+}22, PLOG24, QCM^{+}20, RC22, VWFW22, WEF^{+}21]. call [AVF^{+}23a, CCB^{+}21, FFYE21, PdBA^{+}25, VJL^{+}21]. Callinectes [ABC^{+}24, BLM^{+}25]. Calling [AGSI22]. Cambodia [LHTR23]. Cambridge [Niu22, Smi20, Tan23a]. cameras [AIG^{+}20]. Cameron [Ngu23]. Cameroon [NAS^{+}22, WN25]. campaigns [WWD23]. Campeche [GÁPSAP^{+}23, GÁSVPS^{+}25, dRA25]. Campos [SBP^{+}25]. Can [AACdlT23, AAW^{+}23, Ber24, Fen23, Gir22, LZL25, NCF25, NWZ^{+}21, OAdJC^{+}25, PHD^{+}24, PMN^{+}22, RM21b, SAS^{+}25b, Sol23b, Ste23, TLYC20, WS21, WKH22, WL24a, YHX24, YSWW21, BEM^{+}20, BdLAdC^{+}20, BCR25, CBR21, CSM^{+}25, CMS^{+}21, MWB^{+}21, RLS20, RMBR22, SWF^{+}24, SWHT22, SSAAB^{+}24, TSA^{+}22, TFH^{+}23, WNJ25, WHAD^{+}25, YFD^{+}21, dlPC24, CDC20b, HHEJB20, Pha20]. Canada [BBB^{+}24, BBS^{+}21, BMRM23, BMB^{+}21, CA20, CMNH21, CGP23, DF24, FIB25, JKCY25, JHHG23, JSW22b, KDH^{+}21, Laj21, LTRA20, NGRM^{+}22, OCT^{+}22, Sha22, SSB^{+}23, VAG^{+}24, WA24, WJLSH21, WH20, ZSWJ23]. Canadian [AMR21a, AMR21b, CMS20, CLD^{+}24, DKKS21, Gor21, HBHJ20, WMKG21]. Canary [GMCMJA^{+}21, MBORGN24]. canned [Kaw22]. canning [FGPVPP^{+}24, dHJTGO21, RAFLRP21]. canyon [SCA^{+}22]. capabilities [WTO24]. Capability [GDLC20, TMAS24]. capacities [BBS^{+}21]. Capacity [LLLF22, LLCH21, OCD^{+}23, VHD20, ACG21, BXV20, CHN^{+}24, FR24, FAC^{+}23, FMS^{+}22, GLB^{+}24, GBFMT23, HCWC21, JDF^{+}23, Kim23, OEV23, PHF20, PTCC24, RRS25, SMHPMO^{+}21, WGP^{+}21, ZRR^{+}25]. Cape [LA23, SAO^{+}21, Aqu23, ESD24]. Capital [JL21d, GMFD25, RAFLRP21, RC22, SBP^{+}25, SZ20, Sol23b, SOLN20, TMAS24, TV21, WTO24, XXLW24, CLZ23]. captain [Ste23]. captains [CSGC22]. capture [CCKK22, GSA23, LYLZ25, SHS^{+}25, TMS^{+}22]. captured [TSS^{+}24]. Capturing [IO23, MHEW20, HKKY22]. Carbon [Okş23, WW22, AHN25, AFR25, ANR^{+}22, CDHY22, CDS^{+}25, DGG22, DXS^{+}21, DG24b, Dom23, DPSF25, FAB^{+}23, GZW^{+}20, HZ23, HSGS^{+}23, KNY22, KWY^{+}22, NHD23, QLZ^{+}24, QTBL23, QUK21, QK24, RLRS21, SLL^{+}24, SBM^{+}25a, SGA^{+}21, SLW^{+}23, SOK^{+}22, TBV^{+}23, TMTC20, TAP^{+}23, WXLD21, WLQD21, WKK^{+}25, WL22b, WLXY22, ZFGL20]. carbonate [PCG^{+}23]. carding [KL21a, KL24]. care [VKN20]. careful [HNA25]. cargoes [PPvD^{+}24]. Caribbean [BJH21, BJ25, Bet25, GdCC^{+}25, GGMCMM^{+}24, HGQJD24, Has23, Lov23, Mah22, Moh23, SWF^{+}24, Tho25, WLRH^{+}20]. CARICOM [Has20]. Carolin [WQ25]. Carolina [ALML23, CHN^{+}24]. carrying [CHN^{+}24, FR24, FAC^{+}23, GBFMT23, PHF20, SMHPMO^{+}21]. cascading [KSÖB25]. Case [CJ23, HS23, LA23, LAF^{+}24, MLJ^{+}22, PKC23, RJAT23, SWI^{+}22, SN23, SSAAB^{+}24, APV^{+}20, ASV^{+}21, AOO^{+}23b, AKF^{+}25, AACdlT23, ACV^{+}21, Aqu23, AS23a, AGSI22, AR23a, ASGZ22, BSBD21, BE20a, BDLO23, BMGDS^{+}20, BJGG^{+}24, BDDE20, BLM^{+}25, BUM^{+}24, BMB^{+}21, CVCV^{+}22, CC21, Cao25, CDC^{+}25, CWB^{+}20, CRIMLS24, CMC^{+}23, CJK24, CDHY22, CDC24, CTIH21, CLC21, CCCC21, CCKK22, CES20, CLB^{+}20, CVA22b, CFS^{+}24, dFVCGdS21, Cro21, COJZ^{+}21, DRCdMCJ21, DSQ^{+}22, DUYEA21, DFF23, DG20, DMT25, Dun21b, EAGV20, EVPMKP22, FTGD^{+}20, FY24, FGPPPV21, FCG25, Fro21, FMWW21, FMS^{+}22, GPA^{+}22, GMFD25, GBAECCRM25, Geh23, GPM20, GPM23, GCC^{+}21b, GBDC24, GAA^{+}20, Han21, HWT24, HLW^{+}23, HRR^{+}20, HHB^{+}21, Hor23, HCT^{+}21, IBC^{+}20, JY20, JTR^{+}20, JAH^{+}24, JLY23, JTdH25, JAdS^{+}20, JG23, JMB^{+}20, JL21c, Kim23, KdBK25, Koe22]. case [Kon20, KABT21, LDC^{+}24, LrS22, cL22, LL25, LWG23, LTJ^{+}23, LJ22, LH25, LJL22, LLSY22, LLA^{+}25, LJ21, LSY20, LSOP24, MC25, MAA^{+}24, MVSP^{+}21, MSBA22, MFAH25, MLB^{+}25, MMSB^{+}20, MGF20, MCTZ24, MBL24, NHD23, NNRS24, NEH^{+}23, OCD^{+}20, Okş23, OVSVPP^{+}22, Øst24, OAM21, PCA21, PRM^{+}20, PKR^{+}22, PL24a, PESO24, PBG23, PGM^{+}24, PMR23, PC20, Qi24, RBA23, RMF^{+}21, RBB^{+}24a, RYL25, RCM24, RDM^{+}20, RNGRC22, Rus20, SPG20, SGR21, SVV24, SSZ^{+}24, SMYH22, SCGC20, SCCW21, SOS20, SVG24, SGZH^{+}21, SCLM^{+}21, SSW^{+}25, SZBA20, SSZ^{+}22, SRSGMGG21b, SSL25, SLK^{+}23, TM23a, Tei23, TBH^{+}21, THGR^{+}20, TSH22b, TLT^{+}22, Tru23, TDP20, TYS22, UISM^{+}23, VGK23, WD23, WJB^{+}20, WW21a, WSD21, WRL23, WML25, WCL25, WFG^{+}22b, WKBS20, WDC^{+}20, WLW^{+}22, WWC23, YFD^{+}21, YJG24, YS21, YCL22b, YD22, YXX^{+}24, YC21b, ZJ20]. case [ZCCQ22a, ZCCQ22b, dJdMCL^{+}24, vLS24, vNdB25, AANM^{+}22, CMOJ^{+}24, GLADGGVL23, IKB21, LDAC20, ZWJvL20]. case-study [CRIMLS24, Cro21]. cases [dlÁPAFGO20, STG^{+}20]. cash [ABRA20, BK23, NRYG20]. Caspian [AB20, Pie21]. Cassandra [Sos21a]. Cassurubá [MN21]. Castellón [DL21]. castle [FTG^{+}23]. casualty [FFE^{+}24]. Catalan [CGB^{+}25]. Catalonia [GPM23]. catalyze [FLB^{+}21]. catalyzing [MGDJ22]. Catch [And24, DSJ^{+}23, HBJ21, PHD^{+}24, ZLD^{+}21, AHT25, BXV20, EJC24, GÁPSAP^{+}23, GMM^{+}20, HNA25, MLMD21, MMM^{+}22, Mat20, MAK22, MAG^{+}22, MSGAR20, STIS20, SCSR^{+}22, SCSR^{+}23, SVV24, TSLK23, TYS22, YCL22a, dlPC24]. Catch-rate [PHD^{+}24]. catches [BSLRR24, CPR21, CHB^{+}20, FZ25, IBC^{+}20, KMG^{+}20, MH20a, SFGEPC22, SIW^{+}21, Sta23b, VAMR^{+}20]. Catching [Mar21b]. categorising [SRS23]. category [Kai21]. Caught [ASPT24, JKCY25, SH22, YS21, BdSL25, MM22]. causality [dHJTGO21, Tar25]. causative [LM24]. cause [De 21, RE21]. caused [KLK23, MRKM23, SL20, WABR^{+}22, dOPM^{+}20]. Causes [GAD^{+}23, DT23, NBB^{+}25]. causing [MÇ23]. Cavallo [HNMSG21]. Cayman [BDMW22, Dun21a]. CCAMLR [CHC^{+}21, SB20, TBH^{+}21, Wen20]. CCG [Zha25]. cell [CGAF25, LJ22]. cell-based [CGAF25, LJ22]. cells [LDC^{+}24]. Cellular [Tel23a]. Celtic [FGPPGG22, CRP^{+}20, FGPPGG23]. census [APF22]. centered [SDG^{+}21, SDG^{+}22]. Centering [BJA^{+}24]. Central [Abo23, AKF^{+}25, HAD24, MAG^{+}22, TCS^{+}20, CS21, FMCT^{+}21, MHS^{+}25a, MHS^{+}25b, PCHD^{+}25, RNGRC22, AS23a, CMC^{+}23, COJZ^{+}21, DMd^{+}20, Mol24, MGPR^{+}25, VYB20]. central-southern [FMCT^{+}21]. Centralization [Sak21]. centre [LMMT23]. Centring [YCF24]. centuries [Ken20, TUBA22]. Century [CTH^{+}23, PZPB21, SF22]. CEO [VJO23, VJÓK25]. Cerastoderma [BASM22]. Certifiability [NRF21]. certification [AO20, BI21, BLM^{+}25, BGMCG24, GBDC24, HPG20, LJ21, MLA^{+}24, NK21, NRF21, PCK22, RFG24, Sah22, WS21, WTO24]. certifications [JAK^{+}23]. certified [AO20, LLM^{+}23, SPG20, TOAPJ21, TOL^{+}22, WEF^{+}21]. cessation [LBL21]. Cetacean [OWCFDC21, DER^{+}21, DHU24, DHU25, PEM^{+}25, RDM^{+}20, SPJB22]. cetacean-based [RDM^{+}20]. Cetaceans [WtDD^{+}21, BBBS22, DBB^{+}22, KDH^{+}21, MLJ^{+}22, MIMS^{+}20]. Cg [Yu22]. Cg-18 [Yu22]. Chagos [HS24]. chain [ALM^{+}21b, Bot20, CSMMC^{+}25, CJN25, CLB^{+}23, CSCGC20, dFVCGdS21, CEMRMB^{+}24, FGC^{+}24, FLLML25, GSG^{+}24, GNMGWM21, How20, HP20, JTdH25, JAdS^{+}20, JKKT20, KWMM20a, KWMM20b, KOW^{+}25, Knu25, LCL20, LWH^{+}24, Lin25, LLL^{+}24, LSOP24, MGS21, MMM^{+}22, MMW^{+}23, NAS^{+}22, PCSS^{+}21, Seu22, SdCMN^{+}22, THP^{+}22b, TGHF^{+}22, TPG^{+}23, VMJ25, VKN20, WCC^{+}20, WEF^{+}21, YTG^{+}23, ZW21, ZCSK23, dlPC24, BMR20, dS22a]. chains [AANM^{+}22, BRM^{+}21, CTDY25, DLJ^{+}20, FFYE21, GKO^{+}25, HRC^{+}24, MFAH25, RMU^{+}22, TOHK20, VAS^{+}25, VVR^{+}22, WKBS20, WMS21]. Challenge [WQ25, AML^{+}21, Kaw22, LC20a, SH23, WEF^{+}21, YK21, YYL20, AMK^{+}21]. challenged [BCR25, TJ21a]. Challenges [SONB^{+}22, CMC24, CJH24, DF24, DBZ24, GG20, GMGLV21, JF24, KWF^{+}21, KMSN25, KHdSM^{+}22, KOCL23, LGP25, Liu23, PDS24, RWVH21, RCM^{+}22, Sha22, SCGC20, SF22, Tel23a, WD20, WdSRY23, dRA25, ASPT24, AM22, BH23a, Ber24, BI21, CC22, CMMVdPV25, CWKW20, CGGA20, CSRG25, Cra20, Dav25, DQJ23, hBDF24, DG23, DLL25, EMN^{+}24, Fis22, FMLMM23, FLLML25, GAGLS25, GSR21, GN22, GE23, GGMCMM^{+}24, GJS22, HIP^{+}21, HFD^{+}20, HZC^{+}24, HLM^{+}20, INW20, Kob22, Koj23, Kul23, KSV^{+}23, LL25, LMSLMQ21, LWH^{+}24, LJ23, LCL24, LF25, LFVCF24, LL24b, Mag23, MG25, MTS^{+}24, MIH^{+}24, NSZ^{+}21, NJBM^{+}23, OPL20, OWC^{+}24, OYECdOP24, PVH^{+}21, QK24, RBTW24, RFPS20, RSFvH^{+}22, SAA^{+}21, Sas25a, SSZ^{+}24, SEG^{+}22, TNE^{+}24, Tho25, Tod22a, WvdMG^{+}21, Wan24b, WWG^{+}23, WWG^{+}25, WRV20, Wen20, Wil21b, WLXY22]. challenges [Wu22, YL21, YYM25, ZZH21, ZSZ20, ZSLY20, CQW24, CLS20, CWR20, JL21c, SIW^{+}21, WMKG21]. Challenging [AHENNA^{+}24, Cro21, KLSG21, MFT^{+}22]. Cham [AFHM23, KJE21b, LPO25, Rub24, Sta23a]. Chamelea [BSBD21]. Chang [LK24b]. Change [Kar23, LDC^{+}20, Niu22, Smi20, vHSSK20, AMLC^{+}23, BG20b, BSM23, BGQL24, Bre25, BvdH23, CHH^{+}25, CWB^{+}20, CWKW20, CJLJ24, CWKE20, De 21, DP23, EAGV20, FZ25, FAMW^{+}25, FWGM24, FRLMW25, GCV^{+}21, GLB^{+}24, GB21, GRvdM24, HCWC21, HBR^{+}23, Jon21, KSÖB25, KOBM24, Ken20, KDHL24, KHH^{+}23, KWK^{+}25, Kon20, KDVR20, LMSLMQ21, MR24, Mal21, MGDJ22, Mat20, McC24, MHN^{+}20, MYGM25, O’D21, OGGC20, PBS^{+}23, PK23, QSC24, QF20, RB20, RHO24, RDLA^{+}20, RHL20, SOO20, SMLRB^{+}21, SBD^{+}23, SBM^{+}25b, SGL^{+}23, SRH23, SOK^{+}22, SN23, SHFP21, SYY24, TRN^{+}24, TKP23, TMTC20, VKS23, VLL^{+}24, WNJ25, WWD24, WW25, WWD23, WtDD^{+}21, WGP^{+}21, YG20, dSdLSGB22, vLS24]. Changes [GS21, CSdMR^{+}24, AFMdSdTdCC22, BdSL25, BASM22, CBR21, CLL20, CAM^{+}23, FRVB22, GTMP20, HCW^{+}23, KPN20, KMJ24, MdCLL^{+}23, MK24, NHA23, OCZ^{+}22, PSG25, SSK23, TLT^{+}22, TGHF^{+}22, VGC^{+}22, XWH^{+}22, ZZH21, dSF24, dSVT^{+}25]. Changing [CMOB^{+}20, DKKS21, Fen23, MSO23, ZZMG^{+}24, GMG^{+}23, HSH^{+}21, HACCM22, TJS^{+}25, Tra23a, Wen20, WGBS25, YMG23]. Channel [DFF23]. channels [ZZW^{+}21]. Chanos [SGPF21]. Chaos [ABME22]. Characterising [FMTA^{+}25, KDVR20]. Characteristics [GSB24, YYD23, EdLP^{+}21, FLC^{+}23, JKV^{+}22, MHS^{+}25a, MHS^{+}25b, WL22b, ZSZ20, ZWC^{+}23, CMS20, ZLZ21]. Characterization [SVP^{+}23, VMC^{+}21]. charter [Bov22, CRJ22, GGZRAHVS25, WDC^{+}20]. Charting [MBDG23, TNE^{+}24, Wan24a]. charts [Alm21a]. check [dOPM^{+}20]. Checking [BASM22]. Cheilinus [HS23]. Cheltenham [Smi23b]. chemical [ALML23, Lev20, MLM23]. Chiang [LK24b]. chicken [HM22b]. child [TOL^{+}22]. Children [KL21b]. Chile [DSQ^{+}22, PMR23, AELR^{+}20, AS23a, ÁTGCE^{+}25, BVvKL21, BBFA^{+}24, BMH^{+}22, BRR^{+}22, CBO^{+}21, CBC24, FMCT^{+}21, FMTA^{+}25, GEGRdL24, GRPRVG23, GCV^{+}20, MHS^{+}25a, MHS^{+}25b, MS21, RMBR22, Sep25, TBAR^{+}25, VMV24b]. Chilean [AS23a, BBGHR23, CC21, GSAF21, LKE^{+}22, PV25, RQB^{+}20, SCLM^{+}21, WFG^{+}22b]. Chiloé [DSQ^{+}22, BMH^{+}22, PMR23]. China [AP24, CHB^{+}20, CDY24, COYS24, CZX24, Fan24, GC22, Guo20, HJPL23, HH21, HGLR22, Hsi20, Ngu24, PL20a, PWL^{+}23b, Sas25a, SW22, SHZC21, Zha21, ZFGL20, ZLGX20, Zou23, AWX21, BHdMCM^{+}21, CC22, CC23, CLL22a, CLZ23, CQ20, CDHY22, CMWZ24, COYS24, CDC24, CSHO^{+}20, DLX21, DSJ^{+}23, DG20, Dua22, FXC^{+}22, FZYF23, FZL25, FWL^{+}22, FL24, GY23, GW21, Guo20, GGL22, HJPL23, HPH22, Hon21, HLM^{+}20, HSHS24, JL21a, JLY23, JWX^{+}20, JF20, JL21b, JZW21, JF22, JH24, JK22, KDHL24, KGP^{+}22, LDC^{+}20, LXSX21, LW22, LLHG23, Li23b, LMC^{+}23b, LC23, LYQM21, LXH^{+}22, LYL^{+}22, LDL^{+}23, Lin24, LWX^{+}20, Liu20, Liu21, LXW^{+}21, LJL22, LHPJ23, LH24, LCL24, LYLZ25, MCL22, MZ24, MZZ24, MX21, Min24, MY24, NWZ^{+}21, PH24, PZZ^{+}21, Qi24, QYD20, QTD21, QWBY21, RZ21b, RJ21, Ren21a, Ren21b, RXN23, RXX24, RYL25, Sha20, SC22, SLY21, Shi23b, SGWL21, SF22, SFB24, SC24, SC25]. China [SGRW22, SWXM21, SWW^{+}21, STKC22, SCY24, SYY24, SH25, TLYC20, Tan22, TZZ^{+}21, TKC23, TBPC20, WDG^{+}20, WSD21, WCZS21, WLX21, WW22, Wan24a, WZX24, Wan24c, WKK^{+}25, WdSRY23, WCL25, WFP25, WLW^{+}22, WL22b, WL24a, WL24b, XL25, XZ23, XLLS21, XHY22, XT23, XLW24, YWY20, YCL25, YSWW21, YW20, YYL20, YH20, YD22, YWYL22, YYD23, YLSC23, YZL24, YYM25, YXX^{+}24, ZV20, ZL20, ZTZ^{+}21, ZMZ23, ZMT^{+}25, ZJ20, ZPX^{+}22, Zha22, ZCCQ22a, ZCCQ22b, ZZW^{+}21, ZT21, ZLZ21, ZZD^{+}23, ZWC^{+}23, ZCSK23, ZCS^{+}23, Zho23, ZLH21, ZLD^{+}21, ZX24, Zou23, dlC24, WRL23]. Chinese [Alf20, CWKW20, Cha22, CDY24, CL24, CZX24, CZW24, CZ24, DLYM22, DLL25, LC20a, LK24a, LZL25, Lia25, LCAX21, LL24b, Mar21b, MLAS23, MLM23, Pan25a, SGK21, SHZ^{+}21, SFN^{+}21, YTG^{+}23, ZWJvL20, ZZH21, ZTW23, ZSZ20, ZWXL21]. chitin [VAS^{+}25]. chitin/chitosan [VAS^{+}25]. chitosan [VAS^{+}25]. choice [AFRFGP25, CDD^{+}23, CLL20, DMK24, Faj22, FRDK24, KNY22, PDD^{+}25, SNP21, SL20, TOAPJ21, THT^{+}23, Tra23a, WLX21, WYD^{+}22]. chokes [CBR21]. cholera [HO25]. chondrichthyans [MT20]. Chondrichthyes [KWK^{+}25]. Chorrera [GBC21b]. Christian [Ste25]. chronic [PGK22]. chum [KS23]. ciguatera [Nal24]. Circular [CSRG25, SBGR24]. circulation [Cha22]. Circumpolar [BB20]. CITES’ [Mag23, HS23, Kai21, Sel20, SMW^{+}21]. cities [Bre25, GPM20, LZL25, LHPJ23, RZ21b, RJ21, RXX24]. Citizen [SCA^{+}20, CWLD25, GS21, GVH^{+}23, HLCM21, HMG^{+}21b, TSLK23, WFG^{+}22b]. City [PCSS^{+}21, TM23a, DRCdMCJ21, SdCMN^{+}22, VRBBM^{+}24, TCT21]. civilization [LYL^{+}22, RXX24, TZC23]. claim [SCF^{+}22, SMCW22]. claims [GG22, LZX24]. clam [BSBD21, BRR^{+}22, CCG^{+}25, LCMBGP22, SVP^{+}23]. clarifications [dS22a]. Clarifying [TAP^{+}23]. clarity [Ske24]. Clark [De 20]. Class [Dun24]. classification [HLW^{+}23, HLO^{+}21, PL23]. clauses [Bov22, LK24a]. Clean [ZABI25, BST23, CCX23, TSLK23, THS^{+}24, XLW24]. clean-up [TSLK23]. cleaner [TMAS24]. cleaning [CMI^{+}22]. clearance [AHR24]. Clew [MMNC22]. client [RMU^{+}22]. Climate [CCNGS25, FAMW^{+}25, GB21, Kar23, MHN^{+}20, Niu22, OOGK^{+}24, Pre21, SN23, VKS23, AH23, BG20b, Bir20, BB20, BA25b, BvdH23, CHH^{+}25, CWB^{+}20, CAVJ^{+}20, CWKW20, CJLJ24, CWKE20, DKKS21, FZ25, FWGM24, FRLMW25, GPI^{+}24, GLB^{+}24, GLT23, HCWC21, HSH^{+}21, HBR^{+}23, HGLR22, HSM^{+}23, HSGS^{+}23, Jon21, KS23, KOBM24, KHH^{+}23, Kon20, KDVR20, LMSLMQ21, LGW^{+}21, MR24, Mal21, MKM23, MWM^{+}23a, MN23, McC24, Moh23, MYGM25, O’D21, OCD^{+}20, OGGC20, PK23, PCCOS24, QSC24, RRS25, RB20, RHO24, RDLA^{+}20, RHL20, SOO20, SMLRB^{+}21, SBM^{+}25b, SGL^{+}23, SOK^{+}22, SHFP21, SSF23, Tel23b, TFH^{+}23, TLT^{+}22, TJS^{+}25, TRN^{+}24, TKP23, VLL^{+}24, WWD24, WW25, WWD23, Wen20, WHAD^{+}25, WtDD^{+}21, YG20, dSdLSGB22, Smi20]. climate-adaptive [TFH^{+}23]. climate-ready [LGW^{+}21, MWM^{+}23a, RRS25]. climate-resilient [VLL^{+}24]. climate-responsiveness [DKKS21]. climate-shifted [Tel23b]. Climate-smart [CCNGS25, OOGK^{+}24]. climates [DKKS21]. climatic [IRK^{+}21, SDL^{+}22]. clock [AR22]. close [HSSM24]. closed [GM25b, MLMD21, MMM^{+}22, MMC^{+}22, MAA^{+}25, OA20]. closer [LC23]. Closing [BÅ25a, GR20b, MCPS20]. closure [MNS^{+}24, Pra20]. closures [Ber23, De 21]. Clownfish [YNZ22]. club [GR20b]. clubs [Sak21]. clusters [SPT^{+}20, ZYTT21a, ZYTT21b]. CO [GZW^{+}20, BLY^{+}24, BAJR24, HYN^{+}23, LPMS24, MSBA22, SB20, Ton24, CPV^{+}21, CdLN^{+}21, CD25, CCP25, CSL21, DL21, DBS^{+}22b, Dun21b, EMS^{+}23, FMTA^{+}25, GR25, Gar20, GMMMRC^{+}25, GRPRVG23, GWAPRPRU22, GEB^{+}21, HWT24, HW22, INW20, KCH^{+}25, Kim23, KJS^{+}22, LJ22, LA23, MRD^{+}20, MAR^{+}20b, MGNP25, PL20b, PWL^{+}23b, RHG^{+}25, SCJ22, SSZ^{+}24, SPHO23, VGA25, WJB^{+}20, WBPG20, ZRR^{+}25, GICM25, MVSP^{+}21, OOGK^{+}24, SSK23]. co-benefits [WBPG20]. Co-building [Ton24]. Co-creation [LPMS24]. co-designing [RHG^{+}25]. co-development [CPV^{+}21, EMS^{+}23]. co-evolution [LJ22]. co-exist [Gar20]. co-existence [WJB^{+}20]. co-existing [GRPRVG23]. co-governance [MAR^{+}20b]. co-led [DBS^{+}22b]. Co-location [BLY^{+}24]. co-managed [CdLN^{+}21, Dun21b, ZRR^{+}25]. Co-management [HYN^{+}23, MSBA22, CD25, CCP25, CSL21, DL21, FMTA^{+}25, GMMMRC^{+}25, GWAPRPRU22, HWT24, HW22, INW20, Kim23, KJS^{+}22, LA23, MRD^{+}20, MGNP25, PL20b, SSZ^{+}24, SPHO23, VGA25]. co-operation [PWL^{+}23b]. Co-produced [BAJR24, GEB^{+}21]. Co-production [SB20, KCH^{+}25, SCJ22]. co-sponsorship [GR25]. coal [SBGR24]. coalitions [VJ24]. Coast [BBS^{+}21, KABT21, LH24, SBFHD^{+}23, TCS^{+}20, WB24, ZABI25, AB20, AVC24, ADR^{+}21, BOOO23, CHEC21, CACJR^{+}21, CHN^{+}24, DXS^{+}21, GSA23, GBFMT23, HDAPG^{+}24, JAdS^{+}20, dSJNS20, KWS20, KOW^{+}25, KSV^{+}23, LCL20, MR24, MDLA21, Mba22, OWS20, Pan25a, PMC24, RPK23, RSLR24, SRR^{+}21, WDC^{+}20, dJdMCL^{+}24, DBS^{+}22b, DP23, EJC24, GC22, Lia25, Lin24, MBB^{+}22, RTMCVMDE22]. Coastal [Bir20, BB20, CAGBM22, COM20, EVPMKP22, HM21, KWG23, OC24, O’D21, PFM20, RS24, SML^{+}25, SH22, TPAC24, YMG23, dSF24, AN25, ALML23, AH23, Alf20, AC21, BDVR25, BOOO23, BG20b, BSM23, Bea22, BMR23, BASM22, BBGG22, BYL25, CBO^{+}21, CBCA25, CSV^{+}24, CDP22, CVH^{+}20, CQ20, CLU^{+}24, CMWT24, CBD^{+}23, DC20, DMK24, DJC^{+}23, DHC22, DSQ^{+}22, DB20, DSV^{+}25, EdLP^{+}21, EMS^{+}23, FMGV20, FRVB22, FSF^{+}21, FAGS21, GPI^{+}24, GCF^{+}23, GSA23, GPM20, GOH^{+}20, Gro20, HWM^{+}25, HPC^{+}23, HC22, Hon20, HHEJB20, HBP^{+}21, HP23, HCDN24, IRK^{+}21, IGS^{+}23, Isl25, JL21b, JXWL25, KS22, KOBM24, KMM24, KBB^{+}25, KKAO^{+}25, KKO^{+}22, KL21b, LDG^{+}25, LSB^{+}21, LDAC20, cL22, LBWS23, LW22, LZL25, LLCH21, LTJ^{+}23, LGP25, LLSY22, Liu22, LLJZ22, LHPJ23, LLA^{+}25, LASG^{+}25, LRC24, MFG^{+}22, MKAF21, MAA^{+}25, MKSG20]. coastal [MYGM25, MS20, NNRS24, NAP^{+}21, NWZ^{+}21, NEH^{+}23, NBB^{+}23, OOGS20a, OOGS20b, OZF^{+}25, OGH24, ORdS25, OA20, OEA25, PBS^{+}23, dlÁPAFGO20, PMR23, RCT20, RMBS25, Ree20, RZ21b, RJ21, RXX24, RCM24, RGH^{+}22, RHL20, SA21, SVHV20, Sha20, SGK21, SBJF^{+}21, SCRCI24, SCCW21, SBT^{+}23, SGZH^{+}21, SGH22, SF22, SPA^{+}21, SA25, SNK^{+}20, SAB^{+}24, SLK^{+}23, TNE^{+}24, THT^{+}23, TLT^{+}22, TVN23, UOO20, VTRS20, DR22, VHSV20, VHCJdYC^{+}24, WKO^{+}22, WJB^{+}20, WSD21, WLX21, WYD^{+}22, WKK^{+}25, XWH^{+}22, YCL22a, YSWW21, Zha22, ZCG^{+}23, ZCC^{+}24, dSOW^{+}21, dSdLSGB22, dRA25, BMR23, NAP^{+}21, TCT21]. coastguard [CZ24]. coastline [OWCFDC21]. coasts [PBG^{+}22, RF25]. Cochrane [CBH^{+}25, SCM^{+}25]. cockle [BASM22]. Coco [GASLCP20]. cod [BDR20, BFMS21, EAGV20, SGL^{+}23, SGZH^{+}21, SGH22]. Coda [HNMSG21]. Code [FFE20, KA24, KOBS21, Sin21]. Code-sourced [KA24]. coexistence [SCB25]. cogens [Kuł22]. cognitions [WNJ25]. cognitive [NSM25]. coherence [ABH25b, BTE21, PDB^{+}23, Tar25, VHCJdYC^{+}24]. coin [dOPM^{+}20]. Cold [MSGAR20, Rog24]. Cold-water [MSGAR20]. Collaboration [HWY22, WP23, CdLN^{+}21, CS21, KMT^{+}21, MIMS^{+}20, MMG^{+}25, MHEW20, PTS^{+}25, TV21]. collaborative [CMS20, Can25, CSV^{+}24, DSL^{+}21, DBS^{+}22b, HC22, LJ23, MMSB^{+}20, VAG^{+}24, DB20]. collagen [VAS^{+}25]. collapse [Ber22, LT24]. collapses [VGC^{+}22]. collecting [GRS^{+}23]. collection [BMC20, CLL22a]. collective [BCB20, PN20, SVF20]. collegial [GR20a]. collision [KP24]. collisions [HSH^{+}21, SNP21]. Colombia [GMVM23, LZRC23, PCCOS24, PMFR^{+}22, SCRCI24]. Colombian [MVR21a, MVR21b]. colonial [Bra22, LTRA20]. colonial-nesting [LTRA20]. colonialism [PCN^{+}24]. colonization [Qaf24]. colour [SKL^{+}20]. COLREGs [KP24]. Columbia [SSB^{+}23, WJLSH21, DBS^{+}22b]. comanagement [Col20]. combat [BJH21, BJ25, Bet25, HFY^{+}24, Men23, Tan25]. combating [AHT25, And24, Wan23b, YW24]. Combatting [VVR^{+}22, HM22c]. combined [EMN^{+}24, HAsR24, KDS22, OCCSC23]. Combining [GMRSP23, MFG^{+}23, SVKP21, CBL24, PBT^{+}20]. combustion [GZW^{+}20]. comes [FR23]. commencing [TGP^{+}24]. Commentary [LLM^{+}23, Shi24a, SGRW22]. comments [ML22]. Commercial [Shi22, ADI^{+}25, AP23, BC20b, BDD^{+}21, CJH24, CSGC22, DSB21, ESD24, dlFGFFR^{+}20, GSG20, GWCD22, HKMS20, KCT^{+}24, MRKM23, MCW^{+}22b, MGF20, MIH^{+}24, MNHS25, MHEW20, OCD^{+}20, SS22, SHP^{+}21, SBW^{+}21, SKW^{+}22, SPMN23, SO20, VÖB20a, VÖB20b, VSD^{+}23, VWJMAJ20, WGBS25, YTBM^{+}22]. commercial-recreational [GSG20]. commercially [RZ21a]. Commission [Kar23, SOT^{+}22, Abo23, BAAH24b, GB21, HAD24, Kuł22, SBS21]. commitment [OGS25]. Commitments [HBL^{+}21, Cha24b, LBL21]. Committee [CR23]. committees [PCCR^{+}25]. commodities [KgJqY^{+}24]. ’Common [CAHR^{+}24, BBB23, CdJK24, DMG^{+}23, PV23, AELR^{+}20, BDDE20, Che21b, DMT25, FRCR24, LX24, OKR^{+}20, RAMEG24, SPG20, STS^{+}23, dS21, HBCA^{+}24, SPFA^{+}20]. commons [CBCA25, HBCA^{+}24, PR24, SII22, VHS^{+}20, Zha21]. communality [NGP^{+}21]. Communicating [AMK^{+}21, Gel25, PBG23, TDC21]. communication [BDLO23, BMMF22, CFA^{+}24, MBB^{+}22, MT20, ÓHCÓ24, Pau21, TFF22, WNJ25]. communitie [SML^{+}25, YMG23]. Communities [GPI^{+}24, BCR25, BB20, CMWT24, DC20, DP23, Fro21, GCF^{+}23, GPTASM23, GM25b, HHEJB20, HBP^{+}21, HC21, HM21, IRK^{+}21, IGS^{+}23, IAB^{+}20, JTR^{+}20, JB25, KOBM24, KHMM20, LSB^{+}21, Liu22, MKM23, MRD^{+}20, MTC^{+}23, NRYG20, NEH^{+}23, Pha20, RHG^{+}25, RM21b, RGH^{+}22, SOP^{+}22, SSB^{+}22, SSP^{+}21a, SNK^{+}20, THT^{+}23, UOO20, VHS^{+}20, WKO^{+}22, WJLSH21, WNMS^{+}24, WGP^{+}21, dlPdlLLC^{+}22, CFM^{+}21]. Community [AAKI21, GLAVL21, HWM22, JY20, SAO^{+}21, SSL25, AKF^{+}25, BJA^{+}24, BMM^{+}20, CFM^{+}21, Col20, CBD^{+}23, DBH^{+}21, DCC^{+}20, EMN^{+}24, FGF^{+}24, FSRC25a, FSRC25b, GFD21, GE23, HFD^{+}20, HCWC21, KMM24, KBH^{+}24, LJRHH21, LLS^{+}20, MCBN22, MFM21, MCS25b, MM21b, NHD23, RHG^{+}25, RF25, Ree20, RC22, SSZ^{+}24, SGA^{+}21, SCH^{+}24b, SVKP21, SSF23, VLG^{+}24, Wan23a, XT23, ZCZ20, dMVJdAdSM25, vLS24, Has23, MZ24]. Community-based [GLAVL21, Col20, DBH^{+}21, FSRC25a, FSRC25b, GFD21, HFD^{+}20, KMM24, LLS^{+}20, RHG^{+}25]. community-driven [AKF^{+}25, BMM^{+}20]. community-led [VLG^{+}24]. Community-level [JY20]. companies [BHEE22, BK23, DSWB20, KLS22, LCL20, RAFLRP21, WW21a, ZWJvL20]. Company [HJLN24, FGPPPV21, HT23, LrS22, Ree20]. Comparative [AP23, GGZRAHVS25, RHL20, BDVR25, COYS24, CSHO^{+}20, KGP^{+}22, LHW21, MBORGN24, MB21, NSZ^{+}21, NNRS24, OC20, QTD21, VVR^{+}22, WL24a, ZTM^{+}23, dKMK21]. compared [GCC^{+}21b, SC21]. compares [HWU^{+}23]. Comparing [Cha24b, dlFGFFR^{+}20, HW22, SW22, WJLSH21, ADR^{+}21, BBC24, PSPR20]. Comparison [KB20, KP24, ZWJvL20, GLAVL21, GMM^{+}20, PWL^{+}23a, TOHK20, Zha22]. comparisons [GVH^{+}23, LCAX21]. compatibility [Per24]. compatible [GLS^{+}21, JZW21]. compels [ADWP24]. Compendium [NJW25]. Compensation [YTÇ22, HRR^{+}20, HJ25, JF22, LYHL23, RXN23, YSL^{+}25]. competences [PBC^{+}24]. competencies [Yad24]. Competing [DAB^{+}25, BC21, Tea20]. competition [TG25b, WP23]. competitive [ZT21]. competitiveness [BHEE22, DRCdMCJ21, KKDK20, ZCCQ22b]. complementary [OOGK^{+}24, PGM^{+}24]. complete [Sin21]. Complex [MN21, DJC^{+}23, KLK23, KY20, LV23, MEG^{+}21, Tea20, TJ21b, VYB20, WCF^{+}21, XHZ^{+}21, YK21]. complexities [O’D21]. complexity [Gel25, SRS23]. Compliance [BD23, BCABG24, BC20b, FMCD20, AL20, BFdSdFL^{+}25, Dam24, DNR21, EHM^{+}20, EHM^{+}22, GDLC20, GRPE^{+}21, GHK20, Lin23, Liu25, LFVCF24, MJ23, MJ20, MAK22, PMK^{+}21, PMAVM20, RSH^{+}25, STS^{+}23, SOLN21, BC21]. compliant [WH20]. component [EAA^{+}21]. Components [SLO^{+}24, FZYF23, KBD^{+}20]. composition [BMC20, GÁPSAP^{+}23, PBG^{+}22, TSLK23]. comprehensive [AHBAKH24, AC23, GBM^{+}25, GSG^{+}24, JCMD20, KMSN25, LWX^{+}20, Ren21b, SADP25]. compromised [Mur21b]. Compulsory [Set23]. computable [WHPF20]. computer [VAMR^{+}20]. Con [KJE21a]. concealed [APGWTM^{+}24]. concentration [BADO20, FGPPPV21, SS22]. concept [BAVV24, RKS^{+}23, Wan23a, COJZ^{+}21]. concepts [JL21d]. conceptual [HWC^{+}25, LLL23, Qi22, SWHT22, ZYTT21b]. Conceptualising [BE20c]. Conceptualizing [Day24, RKKM24, SB25]. concern [LX24]. concerning [BFdSdFL^{+}25, BISSBdlMM22, Wan24b]. Concerns [CJ22, DGG22, AKFY22, CMS^{+}25, DBB^{+}22, MFG^{+}23, MTC^{+}23, PZZ^{+}21, SLS25]. concessions [VSD^{+}23]. conch [MVR21b]. conclusions [MGBS24]. Condition [MAG^{+}22, CQAS^{+}25, MKB21, VHSV20, vHSSK20]. conditioned [YMG23]. conditions [BLY^{+}24, BTE21, BvdH23, CJN25, CDG^{+}21, Col20, GAMR20, HSH^{+}21, HTH25, SOS^{+}21b, VM21, ZBF^{+}22]. conducted [KOBS21, Zha25]. confer [FTM^{+}22]. Conference [DRY22, DČFW23, DC23, DZZC24, LLHG23, RBB^{+}24b, TZC23, Ton22a, Ton22b, TC23, Ton24, War23, ABVPM20, AJDA24, DMNT20, MDJ^{+}22, MdCLL^{+}23, SCA^{+}22, TMDN23, WFP25, Cha24b, CLS20, CS20, HGG22c, LSH21, QYX21, RFPS20, SHZC21, YC21a, ZCZ20]. confidentiality [SCZ^{+}24]. configurations [SCSR^{+}22, SCSR^{+}23]. Conflict [EV22, Smy24, SCB25, CLS20, CVA22b, CFA^{+}24, Fen23, GdCC^{+}25, KHMM20, KL21b, MHN^{+}20, PMC24, STG^{+}20, TNY23, VGAH^{+}21, VLS^{+}23]. Conflicting [SPHO23, PFWM23, Gal21]. Conflicts [TFJR25, CMGF20, KWS20, KOCL23, Li23b, MN21, NWS^{+}24, ONA21, PMC24, Sas25a, dSOW^{+}21, Bre25, GCL21]. Conformity [WKO^{+}22]. confronting [GE23]. Congo [Kas21]. Congress [ADWP24]. connect [TG25b]. connected [VMV24b]. connectedness [BA25b]. Connecting [EFCU25, MWM^{+}23b, ADB^{+}22, YS21]. connections [KKA^{+}21]. connectivity [GLPM24, Jen21, KTC^{+}22, Lot22]. consciousness [MCL22].

Consensus [TYS24, EV22, GV22, MDJ^{+}22, TMDN23, VGA25]. Consequence [BVT^{+}21]. Consequences [BNF21, BFMS21, EEIER25, HvT22, JPM^{+}23, Kiz24, Kuł22, XL25, dS21]. consequentiality [XLLS21]. Conservation [BAAH24b, FMWW21, GB21, RDT^{+}23, SCC25, SLP^{+}22, YC21b, AAA^{+}24, AAKI21, AML^{+}21, AVF23b, AAFK^{+}21, AGSI22, AdSMS21, BG25, BCW^{+}25, BC20a, BMB^{+}24, BdSL25, BUR^{+}22, BGK^{+}22, BBHN20, BMB^{+}21, CMS20, CMMVdPV25, CDC^{+}25, CBCA25, CMC24, CHB^{+}20, CMWN^{+}23, COSS21, CBH^{+}24b, CBH^{+}24a, CBH^{+}25, CSGCD25, CTC^{+}23, DMHS25, DH22, lTS25, DBH^{+}22, DPG^{+}22, EHÓ25, EAA^{+}21, EAA^{+}22, EABA^{+}22, ESD24, FWGM24, FWSS^{+}25, FSZ25, FRDK24, FMS^{+}22, GNJR24, GdCC^{+}25, GLPM24, GWÁV^{+}25, Geh23, GBB^{+}21, GAN^{+}25, GMGLV21, GMGLV22, GMMMRC^{+}25, GEB^{+}21, Har22, HCG^{+}25, HMG^{+}21b, HDAPG^{+}24, HNB^{+}25, HGG22c, How20, HCT^{+}21, HH25, Kai21, Kar22c, KSB^{+}23b, KHMM20, KKAO^{+}25, KGL22, KVL^{+}23, LDL21, LMC^{+}23b, Liu20, LM24, MFT^{+}22, MBDG23, MJ20, Mba22, McC20, MWM^{+}23b, MMM^{+}25, MMG^{+}25, MGPR^{+}25, NGP^{+}21]. conservation [OT23, OWCFDC21, OVSVPP^{+}22, PCG^{+}23, PBG23, PT23, PTS^{+}25, PC20, Pri22, RPK23, RSS^{+}20, Sch24a, Sen25, SEG^{+}22, SCM^{+}25, SMW^{+}21, SWHT22, SGA^{+}21, SBDC24, SLW^{+}23, SVKP21, SLH22, SSAAB^{+}24, SZ24, Teo23, Tho22, TEA^{+}22, TMTC20, TMS^{+}22, TMBM^{+}25, VJL^{+}21, VPB^{+}24, VAG^{+}24, VDN^{+}24, WKO^{+}22, WLS^{+}25, WLCG23, YJG24, ZZMG^{+}24, ZCC^{+}24, Mur21b, TBPC20]. Conservationist [Mar20a]. conserve [Ber23]. Consideration [FAC^{+}23, ASG^{+}24, MMNC22, Min24, OCT^{+}22]. Considerations [FRVB22, KCH^{+}25, KFS^{+}22, Sar23, CMD^{+}20, HKKY22, JACD20, LSD^{+}22, OI23, THGR^{+}20, VMC^{+}21, WCC^{+}20]. considered [Pha20, VDOE^{+}24]. Considering [VHS^{+}20, CZW24, CSGCD25, ONA21, OKH^{+}24, SKE22]. consistency [ABM21, CA20, MPD^{+}23]. consistent [CMR21]. consolidation [SS22]. constabulary [Day24]. Constantinople [Qaf24]. constraining [SGL^{+}21]. constraints [BTV^{+}22, VDN^{+}24, Wan24b, WH20, Niu22]. construct [XXMZ23]. Constructing [MCL22, CKY20]. Construction [LL24a, Bau20, CLZ23, Los23, SHZC21, YLSC23]. Constructive [GR20a, CMSN^{+}22]. consultants [OOGS20a]. consultation [PVH^{+}21]. Consultations [BPJ^{+}24]. Consumer [WMS21, CMR21, CGAF25, CLCV21, DLYM22, HT23, MLM23, WM21]. consumerism [HWMO20]. consumers [AAA^{+}24, HWMO20, Nal24]. consumption [BABC20, BBHN20, CDC^{+}25, HKMS20, HDAPG^{+}24, HM21, LLW22, MFF^{+}23, NGA23, PESO24, Sol24, YFD^{+}21, ZY21]. consumption-based [LLW22]. contagious [Bov22]. container [BHEE22, COJZ^{+}21, JHL21, KKDK20, KMM21, KABT21, NCD23, Okş23, SSK23, WW21a, ZLWY22]. containing [Lev20]. containment [SOK^{+}22]. contaminated [LWZ23, MXZ23, WL22a]. Contamination [APH24, WM21]. contemporary [HLW^{+}20, Isl25]. contested [PFM20]. context [CK23, FGPPPV21, GTTM21, HZ23, IAS25, JF20, KCB24, LDAC20, RRR^{+}23, SRS^{+}25, Wat22]. contexts [MAK22]. contextual [BOBMG25]. continental [CHEC21, GdCC^{+}25, KS22, MFRQ23, MB22, dS21, dS22a]. continents [CMGF20]. contingent [CKY20, LrS22, ZMR20]. continue [dMLB^{+}20]. continues [RE23]. continuous [VRBC^{+}21]. continuum [Gue24]. contract [LC20a]. contracting [GR25]. Contractor [Fei20]. contracts [LK24a]. contribute [BBHN20, PMN^{+}22]. contributes [WFK^{+}24]. Contributing [RBB^{+}24b, LGF24, RE23]. Contribution [LRN^{+}22, NRYG20, RGH^{+}22, AOS^{+}24, CJN25, FGC^{+}24, GCF^{+}23, dlFGFFR^{+}20, GDY25, Gul23, KL22, WDC^{+}20, ZCCQ22a, ZT21, dSdLSGB22]. Contributions [SVG24, BBC24, Ege23, Moh23, PJN24, QUK21, VCC^{+}21, ZT23b]. contributory [CT24]. Control [CGMH24, DNR21, Squ21, CC21, CWR20, GMC24, GML^{+}24, HSdT21, JWX^{+}20, KLS22, LMT21, MX21, Mot20, Mot22, SGB21, Şan20, Tes20, WZZC21, YZG23, BCABG24, Kar22a, XWL^{+}20]. controlled [HACCM22, KZR22, Stø21]. controlling [Zhu23]. controls [BSSL^{+}20, KA24]. Controversial [CNB^{+}23, OOGS20a]. Controversies [LH24]. Controversy [KL21b, HM21]. conundrum [CdJK24]. convenience [Fel22, PSMR20]. Convention [CK23, GB21, KBC25, SOT^{+}22, WL20, CQW24, KS22, KCB24, KGW20, LC23, WL24b, RCH22, CCX23, Gul23, MG22, RPY^{+}24, SAS24]. conventional [BMB^{+}21]. Convergence [Sol24, WLS^{+}25]. Convergent [JL21d]. cooperate [LWZ23, WL22a]. Cooperation [EB23, SWGZ22, VGAH^{+}21, Dav25, FOKM23, GSGOP^{+}21, GBCNCL^{+}21, GY23, Hsi20, JF24, PH24, Poo23, RCT20, SRY24, TCZ21, Tel23b, Ton22b, Ton24, WRL23, WFP25, XHZ^{+}21, YC21a, ZFGL20, ZLGX20, MY24]. Cooperative [Kay25, BLM24, MHEW20, TSS^{+}24, WJLSH21, ZHL^{+}23, BDD^{+}21, PSTW21]. cooperatives [Sol23b]. coordinating [SZBA20, YK21]. Coordination [Bre25, HCT^{+}21, KMTvT25, LSH21, LHPJ23, MMSB^{+}20, Smy24, TCZ21]. cope [GTMP20, RKR20]. Coping [GRA^{+}22, SAB^{+}24, ZTZ^{+}21]. copper [VMV24b]. Coral [CLRW22, RBEA^{+}22, SVF20, WWD24, ADWP24, AHN^{+}23, BDMW22, BS22, CWB^{+}20, DDAB^{+}23, Dun21a, Dun21b, GA20, GN22, HVS22, HF20, LRD^{+}22, Lev20, McC20, McC21, MDG^{+}23, NTH^{+}21, SKE22, SPCCvT21, YFD^{+}21]. corals [AKFY22, MSGAR20]. cordatus [dFVCGdS21]. core [EdLP^{+}21, HPS23, Pan25a]. Cork [OGGC20]. coronavirus [Gor21, IKB21, SFB24]. corporate [KLS^{+}25, LZX24, RFG24, UOO20, WS23]. corpus [CL24]. corpus-based [CL24]. Corrected [LLM^{+}23]. corrective [KH24]. correlated [JY20]. Corridors [Wan23c]. Corrigendum [BJ25, CBH^{+}24b, DHU25, EHM^{+}22, FSRC25a, GZW^{+}20, Mal21, MHS^{+}25a, Mot22, RP23a, SCSR^{+}23, SDG^{+}22, SC25, WWG^{+}25, ZGR^{+}25a]. corroborates [OOM^{+}25]. corruption [Gar24]. Coryphaena [MAG^{+}22]. Cost [MPL^{+}23, AH23, CdLN^{+}21, GBM^{+}25, JV20, KCH^{+}25, KABT21, MR24, OAM21, SYA25, SC21, SBW^{+}21, WBH^{+}23, ZMR20, ZNA^{+}23]. cost-benefit [AH23, MR24]. cost-effective [CdLN^{+}21]. cost-efficiency [KCH^{+}25]. Costa [WB24, PSPR20, RJAT23, Teo23, UPBB25]. costly [BXV20, COJS^{+}22]. costs [APGWTM^{+}24, BBB^{+}20, BvBN^{+}20, CCGM21, CvHF^{+}24, CBAPM20, DD22, NSPRSS20, RLRS21, VGK23]. cotton [PBG^{+}22]. could [GLS^{+}21, SLMG^{+}21]. Council [BDC23, BI21, JAK^{+}23, LNRM23, SK20a, BPJ^{+}24, CN24, Dyc24, GDB^{+}24, LJ21, MPH^{+}23, NRF21, RCT20, RROP23, Sta23b, TOL^{+}22]. count [KPN20]. counter [Sel20]. countermeasures [LYHL23, SOK^{+}22, YWY20]. Counterproductive [GM25b]. Countries [OAAMS24, SLGAG24, AB20, AGA^{+}21, BHdMCM^{+}21, CJK24, GLAVL21, IGO21, IKB21, Kak20, LCC20, MAA^{+}24, NRF21, PMK^{+}21, SYA25, Sol24, SF22, TTA^{+}20, TCF^{+}25, dKMK21, MZ24]. Country [IMB^{+}24, Gar22, Her22, HLO^{+}21, MCKZ21, Shi20b, TNE^{+}24, TTA^{+}20]. country-level [MCKZ21]. County [LA23, CDC24]. county-level [CDC24]. coupling [LHPJ23, WSD21]. Cournot [Han20]. course [BH23a, BB21, Wan24a]. court [NT20, Pen22]. courts [Pat22]. coverage [SMW^{+}21]. COVID [AEL^{+}23, ASGZ22, ASG^{+}24, BRM^{+}21, BSI^{+}21, BASM22, CDNS23, CCKK22, CRCP^{+}23, De 21, DME23, Ero22, FTM^{+}22, FGPPGG22, GPTASM23, GABSD21, GNMGWM21, HDCH21, HWC^{+}25, HHB^{+}21, HBP^{+}21, Isr22, KFP20, KdBK25, KAC22, Kob22, LWSS21, LSB^{+}21, LBL21, LC20b, LEERRM^{+}21, LKE^{+}22, MCBN22, MAK22, MGK^{+}25, NK21, NGRM^{+}22, NAS^{+}22, PGK22, PAK22, PBD^{+}23, RSK^{+}25, RDD^{+}21, SVV24, SSB^{+}22, SCH^{+}24b, SdCMN^{+}22, SCLM^{+}21, SSP^{+}21a, SNK^{+}20, SSP^{+}21b, SAB^{+}24, Tan22, TAK^{+}22, VMM21, VMJ25, VMP^{+}24, WHL22, WPWM^{+}22, XSCL21, YCBW21, ZLG^{+}24, ZTZ^{+}21, ZTW23, ZTM^{+}23, ZT23a, ZCS^{+}23, ABME22, GMCMJA^{+}21, GS21, MÇ23, MMK^{+}22, NVP^{+}23, SRH^{+}22, VDP^{+}22]. COVID-19 [AEL^{+}23, ASGZ22, ASG^{+}24, BRM^{+}21, BSI^{+}21, BASM22, CDNS23, CCKK22, CRCP^{+}23, De 21, DME23, Ero22, FTM^{+}22, FGPPGG22, GPTASM23, GNMGWM21, HDCH21, HHB^{+}21, HBP^{+}21, Isr22, KFP20, KdBK25, KAC22, Kob22, LSB^{+}21, LC20b, LEERRM^{+}21, LKE^{+}22, MCBN22, MAK22, MGK^{+}25, NK21, NGRM^{+}22, NAS^{+}22, PGK22, PAK22, PBD^{+}23, RSK^{+}25, RDD^{+}21, SVV24, SSB^{+}22, SCH^{+}24b, SdCMN^{+}22, SSP^{+}21a, SNK^{+}20, SSP^{+}21b, SAB^{+}24, Tan22, TAK^{+}22, VMM21, VMJ25, VMP^{+}24, WHL22, WPWM^{+}22, XSCL21, YCBW21, ZLG^{+}24, ZTZ^{+}21, ZTW23, ZTM^{+}23, ZT23a, ZCS^{+}23, ABME22, GMCMJA^{+}21, GS21, MÇ23, MMK^{+}22, NVP^{+}23, SRH^{+}22, VDP^{+}22]. COVID-19-affected [LWSS21]. COVID’19 [Kiz24]. CPUE [FSRC25a, FSRC25b]. Crab [LSOP24, ADR^{+}23, ABC^{+}24, BLM^{+}25, dFVCGdS21, DD22, Gor21, KKA^{+}21, KV20, LMC^{+}23a, LJRHH21, PLLJ24, RKR20, SH24a, VKN20, YSEK22]. crabs [MMB^{+}23, OCZ^{+}22]. Crafting [GdACT20, Wil20]. crass [NW20]. Creating [RB20, CMWN^{+}23]. creation [hBDF24, GV22, LPMS24, SOK^{+}22]. creative [Gel25]. credit [AACdlT23]. crediting [KWY^{+}22]. crew [CSGC22, De 21, HCW^{+}23, YS23]. crews [Pau21]. crime [Álv23, BE20c, Gar24, PGP22, PHP25, SKL23, TCT24, Urb22]. Crimes [HFY^{+}24, SPNP24]. Criminal [Tei23, Álv23, Pat22, Ros22]. criminalise [Faj22]. crises [BMH^{+}22, HSM^{+}23, LCL20]. crisis [AVF23b, BVvKL21, De 21, FGL^{+}21, FG25, Hon20, McC20, Tan22, WKBS20, XL25]. criteria [ASV^{+}21, ABM21, CPBO20, EAGV20, GCG^{+}21, HW22, OCCSC23, SLK^{+}23, TMS^{+}22]. Critical [BL23b, CLP21, DL22, LTJ^{+}21, MFG^{+}22, NAP^{+}21, Åm21, CT24, CLD^{+}24, EAA^{+}21, GLSVMB25, JXWL25, Kıl24a, LX24, PMN^{+}22, SRvNN21, SMDM20, SM23, YT20, ZSLY20, ZLY23]. critically [GPH^{+}21]. critique [CMSN^{+}22]. critiques [YCF24]. Croakers [CLU^{+}24]. Croatian [PBD^{+}23]. crocodile [KHMM20]. Croix [GMRSP23]. Cros [HNMSG21]. Cross [Tel23b, CSM^{+}25, CSGCD25, FKNG25, GSGOP^{+}21, IO23, LCL24, MMK^{+}22, PWL^{+}23a, SL23, ZZMG^{+}24]. cross-border [FKNG25, GSGOP^{+}21, LCL24]. cross-national [SL23]. cross-regional [PWL^{+}23a]. cross-scale [CSM^{+}25, ZZMG^{+}24]. cross-shelf [CSGCD25]. crossing [LLT^{+}20, Tse21]. crowding [AFRFGP25]. Crowdsourced [SR22]. crucial [BFdSdFL^{+}25]. Cruise [ZCSK23, Kiz24, LMT21, LZZ22, LLHG23, LC20b, Set23, SCY24, WCC^{+}20, YCBW21, ZCS^{+}23]. crustacean [ADM^{+}25]. CSR [UOO20]. Cu [KJE21b]. Cuba [AVPAFM^{+}22]. cucumber [GÁPSAP^{+}23, HFD^{+}20, Pra20, SC24, SC25]. cucumbers [PSG25]. cultivation [BvdBvdFK20, Chr20, GBLBV20, YZH22]. Cultural [BTN^{+}24, GCL21, LC23, MWWS21, WM22a, War23, BEH^{+}21, CH25, DBH^{+}22, EHÓ25, FFYE21, GWÁV^{+}25, GW21, HHL^{+}24, KOCL23, LHW21, LKB^{+}23, Lin24, ML22, PT23, PPvD^{+}24, QK24, SSA^{+}23, SPHO23, SRS23, VDN^{+}24, ZCG^{+}23, HVS22]. culturally [DSV^{+}25]. culture [Lin23, LSC25, MCL22, MWM^{+}23b, SL23]. cumulative [CSGCD25, KSO^{+}25, OCT^{+}22, SBT^{+}23, HCWC21]. cure [LC20a]. curfews [MAK22]. Curieuse [COSS21]. curled [BBB23]. Current [BLY^{+}24, CLS20, CLZ23, DLX24, JTR^{+}20, Kul23, LYHL23, LSD^{+}22, OKS^{+}22, QLZ^{+}24, SKW^{+}22, TYS24, Tod25, VSV^{+}23, VCC^{+}21, WNDZ23, WM22b, ZGR^{+}25a, ZGR^{+}25b, ZMM^{+}23, BB21, BHKJT23, BRM20, CJN25, CCB^{+}21, DL21, DFC^{+}20, Kai21, MX21, NGT^{+}23, OZF^{+}25, PCG^{+}23, PdBA^{+}25, TAK^{+}22, WSDB20, GBC^{+}20, Kas21]. currents [WKK^{+}25]. curriculum [LL25]. curve [TMAS24]. customary [Kar22c, Mba22, SRH^{+}22, ZY23]. cuttlefish [GPM23]. CVM [LrS22]. Cyber [Lv20, SM23, LF25]. Cyber-seaworthiness [SM23]. cybernetics [SOS20]. cybersecurity [Kar22b]. cycle [LGP25, RS24, YS21]. cyclone [MDLA21]. cyclones [SNK^{+}20]. cylinder [GMM^{+}20]. Cynoscion [dSJNS20].

dagaa [IOHC22]. Dai’ichi [Mab24]. daily [MLM23]. Dakar [CDNS23]. damage [Fei20, JF22, SL20, Tel23b, ZMR20]. dams [DDGdS^{+}21]. danger [SBG^{+}24]. dangerous [RE21]. Danish [HVN21, NHA23, NSK20, SII22].

Danish-flagged [HVN21]. Danube [Tea20]. Dao [KJE21a]. Dar [SML^{+}25]. darkness [dSRKTC24]. Darwin [TOAPJ21]. dashed [Gul23]. Data [BCABG24, MCS^{+}25a, SLSV25, VBC^{+}20, Alm21a, ASM^{+}21b, ATF^{+}21, AGA^{+}21, BSLRR24, BMC20, CRP^{+}20, CLS^{+}23, CGB^{+}25, CPBO20, CLL22a, CDB20, CNG^{+}23, DBS22a, DGMF^{+}20, FTW^{+}23, Fuj21, GRS^{+}23, GR20a, GCS21, GAD^{+}20, HWM^{+}25, HMG^{+}21a, KA24, KMG^{+}20, KPN20, KBP24, LWSS21, LHV^{+}20, LLLF22, MZVC25, MLJ^{+}22, MFG^{+}23, MMNC22, MNS^{+}24, MAG^{+}22, MSGAR20, OC24, PBS^{+}23, PWL^{+}23b, RZ21a, Şan20, SRR^{+}21, SCD21, STIS20, SSOP^{+}21, SWH^{+}23, SSMG21, SCZ^{+}24, TLT^{+}22, TBAR^{+}25, TK23, TMBM^{+}25, WBB^{+}25, WCF^{+}21, XSCL21, XWH^{+}22, YTG^{+}23, ZCG^{+}23, vSHHT20, PCD^{+}23]. Data-driven [VBC^{+}20, GR20a, KA24, KBP24]. data-limited [CDB20, GCS21, PWL^{+}23b]. data-poor [Alm21a, BSLRR24, PBS^{+}23, YTG^{+}23]. data-scarse [CPBO20]. date [GJS22]. Davao [MLMD21, MMC^{+}22]. David [LK24b, Smi22, Smi23a]. day [BSI^{+}21, Fra23]. de-privatization [CAHR^{+}24]. DEA [PCD^{+}23, FMGV20]. DEA-based [FMGV20]. deadline [Sin21]. Dealing [MSGAR20]. death [CJ23]. debate [BE20b, MT20, SS21a]. debates [BÅ25a, Wan23b]. debris [SONB^{+}22, DT23, HLCM21, HWY22, SGLF^{+}23, SSL25, Wu22, XXMZ23]. Decadal [PSG25, PLWO24]. Decade [MZ24, Pol23, PCHD^{+}25, DG24b, Ngu24, SBG^{+}25, VMC^{+}21]. decades [BBFA^{+}24, Ehl21b, GV22, KKO^{+}22, LJ20, MZVC25, ZV20]. Decarbonisation [Ada22, vLM22, BH23a]. decarbonising [RSN25]. decarbonization [CTR24, Dom23, WZX24]. deceived [LBL21]. December [Ano20b, Ano21c, Ano22c, Ano23c, Ano24c]. Decent [LFLF22, LSD^{+}22, TJS^{+}25, YTÇ22]. decentralisation [JL21c]. decentralization [Sak21]. decentralized [MMSB^{+}20]. decides [Pre21]. Deciphering [SLY25, Yad24]. Decision [GSG20, AMR21b, BL23a, BBBS22, CDB20, DJC^{+}23, DBBL21, FRM22, GCP^{+}20, GSM20, HDH21, HDCH21, HW22, JKV^{+}22, JKCY25, KVL^{+}23, MXZ23, Mab24, MWM^{+}23b, MKB21, MMGAA^{+}21, PHF20, RB20, SCZ^{+}24, TDC21, Wat22, BBV^{+}25]. decision-makers [RB20]. Decision-making [GSG20, AMR21b, BL23a, BBBS22, DJC^{+}23, FRM22, GCP^{+}20, GSM20, JKCY25, Mab24, MWM^{+}23b, MKB21, MMGAA^{+}21, TDC21, BBV^{+}25]. decision-support [CDB20]. decisions [BC20b, EJC24, LMC^{+}23a, MKSG20, Shi24a, Sta23b, SN23, ZCG^{+}23]. deck [KZR22]. decline [Alm21a, ALLR^{+}24, EEIER25, FA24, Fla21, IAB^{+}20, LX24, MZVC25, OOSG21]. declines [SKS21]. declining [BB21, SSH^{+}23]. Decoding [BCABG24]. decommissioning [JvH24, TN21]. Decomposing [Seu22]. decomposition [LLW22, ZZD^{+}23]. decoupling [WW22]. decreasing [ZD20a]. deemed [TW22]. Deep [CHEC21, HZC^{+}24, Kak20, Wil21a, AGM^{+}22, AFHHD22, ADM^{+}25, AR22, BDCJ22, BHKJT23, BRM20, CDD^{+}23, CGHV20, CDC20a, CDC20b, CL20, Cra20, DFC^{+}20, DR25, Fei20, GG20, HSSM24, HSK^{+}23, HSHS24, Jae20, JACD20, KAA^{+}24, MM21b, MSGAR20, OKR^{+}20, OVA22, PL24b, PLHK^{+}24, RFPS20, Squ21, SCA^{+}22, TML^{+}20, VJL^{+}21, WSZ^{+}25, WLCC23, Wil20, Wil22, vdGD21, KGW20]. deep-ocean [AFHHD22]. Deep-sea [CHEC21, HZC^{+}24, Kak20, ADM^{+}25, AR22, BHKJT23, CDD^{+}23, CDC20b, HSSM24, HSHS24, JACD20, KAA^{+}24, MSGAR20, OKR^{+}20, PL24b, PLHK^{+}24, VJL^{+}21, WSZ^{+}25, WLCC23, vdGD21, KGW20]. deep-seabed [AGM^{+}22, HSK^{+}23]. deep-water [BDCJ22]. deepest [VJL^{+}21]. defects [CLL^{+}22b]. defences [COM20]. defend [HPG20]. deficiencies [BWV25]. deficiency [MÇ23]. deficit [OCD^{+}23]. deficits [CGP23]. defined [MKSG20, VvSvH^{+}22]. Defining [BPJ^{+}24, JG20, CVA22b, QCM^{+}20]. definition [GA20, Urb22]. Definitions [SRH23, SB25]. degazettement [ACA^{+}21]. degradation [LYLZ25]. degradations [BCS^{+}24]. degrade [PK23]. Déjà [SBT^{+}23]. del [GASLCP20, MMK^{+}23, MNGM24, TPAC24]. Delayed [AMR21b]. deliberate [VJL^{+}21]. Deliberating [HVS22]. Deliberative [BBGG22, SL20]. delimitation [Mar21a]. delimited [CZ24]. Delimiting [PMR23]. Delineating [AR23b]. Delivering [SCC25, Tel23a]. delivery [RFG24]. Delphi [HRS^{+}24, LZL23, MBDG23]. Delphi-AHP [LZL23]. Delta [BSBD21, dFVCGdS21, NHD23, Tea20, QYX21, UOO20, WHPF20]. demand [ABC^{+}24, BHdMCM^{+}21, CLW^{+}21, DHC22, HCSBM^{+}23, Kaw22, Tan22, TOAPJ21, TCC^{+}22]. DEMATEL [DL22]. demersal [BRSC^{+}24, GMM^{+}20, HKZ25, LSGG^{+}20, NHA23, OCCSC23, SDL^{+}22]. democracy [AELR^{+}20, AE23, BBGG22, PESO24]. democratic [VGA25]. Democratising [MF25]. Demographic [PSA^{+}25]. demography [GS21]. Demolition [TMG24]. demonstrates [CFA^{+}24]. Demonstration [LZL25, AC21, CKY22, RXX24]. Denmark [LDC^{+}24, NAN^{+}24]. density [PHF20]. dependence [TUBA22]. dependency [LP20]. dependent [Ber22, FRLMW25, IRK^{+}21, MAK22, WL22b]. depoliticization [CFA^{+}24]. deposition [RB22]. deposits [JG20, VCC^{+}20]. depredation [CRJ22]. deregulation [GWCD22]. Derelict [ABDRV^{+}23, DT23, ADR^{+}23, DSB21, Lov23]. derived [Öİ24, SOLN21, VAS^{+}25]. desalinated [LYQM21]. description [MMS24]. design [BBB^{+}24, Dun24, Fen23, FSRC25a, FSRC25b, GLPM24, MSZ^{+}23]. designation [CMWZ24, Cho22, KOCL23, Tes20, WD20]. Designing [DN20, OOGA22, RBB^{+}24a, ABM21, JZW21, RHG^{+}25]. designs [KvTNL20, SNP21]. desirability [PSMR20]. despite [HSM^{+}23, SSH^{+}23]. destinations [LLJZ22]. destructive [ABH25a, SBFHD^{+}23]. detail [MDJ^{+}22]. detect [GAD^{+}20]. detention [CLL^{+}22b, KA24]. detentions [MÇ23]. Determinants [SYA25, SZ20, WW21a, CACJR^{+}21, HPS23, Kao23, KPN20, KJS^{+}22, LCC20, Lin25, MNHS25, ZZD^{+}23]. Determination [Kar22a, BRR^{+}22, MMAP25]. determined [GÁPSAP^{+}23, MGS21, Moh23]. Determining [ADR^{+}23, HNA25, Öİ24]. developed [GBDC24, HHH^{+}23, TTA^{+}20]. Developing [BAVV24, BTN^{+}24, CC23, CMT^{+}23, CTIH21, IGS^{+}23, Moh23, SGL^{+}21, Sta25, Tho25, CdLN^{+}21, CJK24, Gun21, GML^{+}24, IKB21, KMS20, MAA^{+}24, MCW^{+}22b, NSN22, NW20, NRF21, ONA21, RLRS21, SAA^{+}21, TTA^{+}20, VSV^{+}23, WH20, BGLTC21]. Development [ASGS23, BCN23, DQJ23, LZL25, MZ24, RXN23, Run24, WM22b, vSHHT20, Alf20, AVPAFM^{+}22, Aqu23, AJDA24, BLS21, BTE21, BYL25, CPV^{+}21, CA20, CBC24, CLZ23, CLW^{+}21, CJL^{+}23, COSS21, CHN^{+}24, DKW20, DČFW23, DZZC24, Ege23, EMN^{+}24, EVPMKP22, EMS^{+}23, FXC^{+}22, FZ25, GLAVL21, GMG^{+}23, GLSVMB25, GW21, HS20, HFD^{+}20, Han21, HPH22, HSK^{+}23, Hor25, HSN^{+}21, HC21, HWY22, IMA22, JC23, JL21a, JLY23, JD20, JV20, KWF^{+}21, Kim20, KWMM20a, KGP^{+}22, KSO^{+}25, LHTR23, LHV^{+}20, Lee24, LZZ22, LLHG23, Li23a, LLLF22, LJ22, LJL22, LLL^{+}24, LLA^{+}25, ML21, MF22, MBG^{+}24a, MG22, MHF^{+}25, MPH^{+}23, MMNC22, MBC^{+}24, MPSF20, MCTZ24, MYGM25, Mur21b, NSN22, NSPRSS20, OZF^{+}25, PESO24, PMN^{+}22, PDJ22, QZZ20, QSC24, QYX21, RA20, RJ21, Ren21a, RXX24, RYL25, RBB^{+}24b, Sak21, SdACdSLdAS24, SVHV20, SWHT22, SGWL21, SHZ^{+}21]. development [SSZ^{+}22, THP22a, Ton22a, Ton22b, UOO20, VÖB20a, VÖB20b, WLY21, WYD^{+}22, WNDZ23, WA24, WMM24, WdSRY23, WFP25, YK20, YSWW21, YZH22, YYM25, YXX^{+}24, YC21b, ZWG^{+}21, Zha22, ZCCQ22b, ZZW^{+}21, ZLZ21, ZWC^{+}23, Zho23, ZT23b, dSF24, LMSLMQ21]. Developments [Qaf24, AE23, AR22, Fra23, GAD^{+}24, Guo20, MX21, YCL25]. deviation [RE21]. device [RQB^{+}20, WLRH^{+}20]. devices [DPB^{+}23, EMH^{+}23, MKM^{+}20, MM23, ZMM^{+}23, CMT^{+}23, MSZ^{+}23]. devil [HS24]. diagnose [STKC22]. diagnosis [CES20]. Diagnostic [CEMRMB^{+}24]. dichotomy [KLSG21]. did [Kim20, KWK^{+}25]. die-off [SOP^{+}22]. Diego [PB20]. diesel [dSTHL^{+}24]. difference [RM21b, WZZC21]. difference-in-differences [WZZC21]. Differences [BMB^{+}21, MÇ23, SCH^{+}24b, BDDE20, JLY23, PSPR20, RYL25, WZZC21]. Different [HC21, RF25, Gue23, HRS^{+}24, Her22, MMW^{+}23, PMK^{+}21, RHST21, VPRR^{+}21]. Differential [CRCP^{+}23]. differentiated [Che21b]. difficulties [LL24b]. diffusion [ZLPT23]. Digital [CTDY25, ETSKA25, LR20, HPH22, Jen21, LEERRM^{+}21, Wan24c]. Dikkowita [JAH^{+}24]. dilemma [RSAP22, SVF20, Zha21]. Dilemmas [GY23, SH25]. dimension [KBD^{+}20, MCW^{+}22a, PSA^{+}25, RSAP22, ZCC^{+}24, Cha20b]. dimensional [LRD^{+}22, Sen25, YZL24]. dimensions [BDM20, Har22, MK24]. Diomede [Tod22b]. dioxide [GZW^{+}20]. diplomacy [Geh23, Hon21, MC25, PCHD^{+}25, TLYC20, TBCB25]. Dipturus [GPH^{+}21]. Dire [Tod22a]. direct [AOS^{+}24, AED20, BBC24, NTH^{+}21, SRSGMGG21a]. Direction [MDJ^{+}22, YZL24]. directional [OAM20]. directionality [BBGHR23]. directions [AN25, Bre25, CTDY25, FZL25, Fuj21, NJBM^{+}23, PSD23, RXN23, WNDZ23]. Directive [GKP^{+}22, PDB^{+}23, ZGR^{+}25a, BN20, GBM^{+}25, Puh23, ANS^{+}20, KWK^{+}25, MPD^{+}23, MFRQ23, Mor20, PKB^{+}21, VRBC^{+}21, VBH^{+}23, ZGR^{+}25b]. Direnzo [Lv20]. disadvantages [Jen23]. disaggregated [Sol24, SLGAG24]. disappearance [CJ23]. Disaster [IGS^{+}23, SWF^{+}24, DMd^{+}20, KWG23, LDAC20, OGS24, SOP^{+}22, dOSTB^{+}20, YSWW21]. disasters [FAGS21, RMBR22]. Discard [GMPT^{+}20, PBT^{+}20, ABVPM20, GTA^{+}20]. discarded [GRS^{+}23, GHW^{+}22, GAD^{+}23, HGG^{+}22a, Lov23, NMN^{+}25, NSK20, RCM24]. discarding [VJL^{+}21]. discards [BMGDS^{+}20, CBR21, GMM^{+}20, SG20c]. Discharge [PL24a, FL24, LWZ23, Lia25, QL24, WZL25, XTY22]. disclosure [ZLWY22]. discolorations [SCA^{+}20]. discontinuity [FSRC25a, FSRC25b]. discounts [PYVF24]. Discourse [LSC25, BBP25, SG21, SH22]. Discourse-historical [LSC25]. discourses [CBC24, Los24, SPCCvT21, Tea20, VMD^{+}24]. discovered [WWG^{+}23, WWG^{+}25]. Discrepancies [CL24]. discrete [CDD^{+}23, DMK24, TOAPJ21]. discretionary [OAM21]. Discursive [VJ24]. discussion [CMWZ24, Dua22, JL21c]. discussions [HGSD23]. disease [LMT21, MAK22]. diseases [Bov22]. Disentangling [HTP^{+}23, Mil22, SBS21]. disincentives [BC21]. disparities [GWCD22]. Disparity [BCABG24]. dispatching [Min24]. dispersed [PPvD^{+}24]. Displacement [SKS21, Tru23]. dispossession [Bra22]. disputant [Ngu24]. dispute [HJPL23, Ngu24, Pan25a, PL24a]. disputed [Cha21, SCGC20, Zha25]. Disputes [Smi23b, CL24, GG22, HGLR22, MC25, NT20, Øst20, PK20b, Shi20a, ZX24, Zou23]. Disrupting [JD20]. disruptions [IGS^{+}23, VMJ25]. disservices [RFG24]. distance [FMCD20, OAM20, vRB24]. distant [AP24, HCS21, Shi23b, TFP25, VMV24b, YS23, YL21, YYM25, Zho23]. distant-water [YYM25]. distinct [Yu22]. Distinction [Shi24a]. distraction [WWD23]. distress [AONF21]. Distribution [AKFY22, AMIM21, AGN^{+}20, AS23b, Che21a, FIB25, GASLCP20, HRC^{+}24, MCS25b, QF20, SSV^{+}23, TZM^{+}22, TV21]. distributional [Muk21, VLL^{+}24]. distributions [CD24, PdBA^{+}25]. distributive [HNA25]. District [MKAF21, TCT21, YC21b, CDP22]. districts [KMM24]. Disturbance [DPSF25]. disturbances [Dam24]. disturbing [WWCS21]. dive [DR25]. diver [ALLR^{+}24, FR24, LM24]. divergent [JL21d]. divers [AHN^{+}23, HMG^{+}21b]. Diverse [BLS21, DSL^{+}21, DCC^{+}20, LIK^{+}24, SHP^{+}21]. diversification [NCF25, SBGR24, SBD^{+}23]. diversify [CWMM^{+}21]. Diversity [GGMCMM^{+}24, FMS^{+}22, HM22a, HMG^{+}21b, LGF24, RLS20, WvdMG^{+}21, MG22]. divide [LEERRM^{+}21]. Diving [AAFK^{+}21, CACJR^{+}21, GMGLV21, GMGLV22, HACCM22, MMO^{+}24, SMHPMO^{+}21]. DNA [BdLAdC^{+}20, CS21, GCC^{+}21a, YJG24]. Do [CJN25, HC25, JL21b, THT^{+}23, ALML23, BEM^{+}20, MS21, Øst20, RJ21, WLQD21, WYD^{+}22, BHEE22, CPR21]. dockside [KMG^{+}20]. documented [BFSP25]. Documenting [DLJ^{+}20, LA23]. documents [AKFY22, CQW24].

Does [ÁCT20, ADB^{+}22, CLL20, HPH22, HvPPV20a, LWG23, Lin25, LJL22, MAA^{+}24, NAN^{+}24, UOO20, GSA23, LG24, PBG^{+}22, QZZ20, SMW^{+}21, ZLZ21, RHC20, TJZ^{+}21]. doesn’t [SCF^{+}22]. doi [LK24b]. Doing [MFG^{+}22, TPAC24, GB21]. dolphin [AVC24, STS^{+}23]. dolphinfish [APGWTM^{+}24, MAG^{+}22]. Domain [CZX24]. Dome [RJAT23, Teo23]. domestic [Cha22, LMC^{+}23b, NNR^{+}21]. dominance [PBG^{+}22, VJ24]. Dominica [SWF^{+}24]. Dominican [PSPR20]. Donax [SVP^{+}23]. Donggang [LSY20]. Dongsha [YCL22b]. don’t [BEM^{+}20]. Dosidicus [CEMRMB^{+}24]. double [Gug24]. double-edged [Gug24]. doubled [SLL^{+}24]. doubloons [LHW21]. down [AFMdSdTdCC22, RC22, SLB21, SMF^{+}23]. Downeast [FSZ25]. downgrading [ACA^{+}21]. downs [CSQ^{+}24]. downscaling [MMNC22]. downsizing [ACA^{+}21]. downstream [dlPC24]. DPSIR [ADK^{+}25, CBO^{+}21, QUK21, QLU^{+}23]. draft [CMSN^{+}22, SK20b, ML22]. dragon [AWX21]. Dragons [LHW21]. drawing [QZZX20]. dredged [SOK^{+}22]. dressing [Lot23]. dried [PSG25]. drifting [CMT^{+}23, EMH^{+}23, MZT^{+}23, ZMM^{+}23]. drilling [JF22]. drive [BBT^{+}25]. driven [AKF^{+}25, BMM^{+}20, CD25, For22, GR20a, KA24, KMSN25, KBP24, LTJ^{+}23, QRK^{+}23, VBC^{+}20]. driver [JDF^{+}23, CSdMR^{+}24]. Drivers [Ken20, VLS^{+}23, WLXY22, CCCC21, GHW^{+}22, JAdS^{+}20, KL21b, LXH^{+}22, MGS21, MLA^{+}24, PDD^{+}25, PLHK^{+}24, SDI^{+}20, SCBB^{+}22, SRH23, SLY25, THS^{+}24]. drives [CDC^{+}25, SRLLG23]. Driving [DLX21, FZYF23, TJ21a, WFG22a, ZWXL21]. drone [CMSH25, SAB^{+}20]. drones [Kle21]. drop [GCC^{+}21a]. Drumhiller [Lv20]. dual [Cha22, FMTA^{+}25, WKK^{+}25]. due [Gor21, GWAPRPRU22, MS20, SKS21, SLMG^{+}21]. dugong [WLS^{+}25]. dumping [CMWZ24]. Dungeness [ADR^{+}23]. duplication [CCB^{+}21]. Duration [LCC20]. during [ASGZ22, BvdBvdFK20, Ber22, Cha21, hC24, EHLT21, FTM^{+}22, GS21, Isr22, JG23, MÇ23, NGRM^{+}22, SPJB22, SOS^{+}21b, TAK^{+}22, VDP^{+}22, WPWM^{+}22, XWH^{+}22, ZTW23, ZTM^{+}23]. Dutch [CWKE20, HvdBH23, MR24, SZBA20, SVKP21, vHSSK20]. duty [Dev22, LWZ23, NOB22, PMFR^{+}22, WL22a]. dykelands [ZCG^{+}23]. Dyking [dZ20]. Dynamic [HSH^{+}21, MWB^{+}21, Tar25, BA25b, CMD^{+}20, FDGM24, FWL^{+}22, Ren21b, YK21, ZZMG^{+}24]. dynamics [ASPT24, AH21, BDM21, CDHY22, GFD21, JAdS^{+}20, Kas21, LT24, MMK^{+}22, NO25, PGSFRMHH24, SN23, SYY24, SLY25, TTL^{+}25, YMG23]. dysfunctional [LG24].

e-logbooks [CLS^{+}23]. Early [SK24, SLO^{+}24, SNK^{+}20, WCZS21, YC21b, INW20]. earnings [SBP^{+}25]. Earth [VJL^{+}21]. Earthscan [Smi23a]. East [Gra21, GC22, PGP22, PK20b, ADR^{+}21, ÁFFSC20, BTN^{+}24, COYS24, FFE^{+}24, HGLR22, SRR^{+}21, Sas25a, TLYC20, VBH^{+}23]. Eastern [CDG^{+}21, YMG23, CLRW22, CDP22, PDJ22, WWG^{+}23, WWG^{+}25, DMS^{+}24, KLS^{+}25, LTW23, LTRA20, Lov23, Mah22]. EBM [GCP^{+}20]. eBook [Sta23a]. Ebro [BSBD21]. EBSAs [MFA^{+}24]. ECAN [MFG^{+}22]. eco [Abo23, AH21, Jon21, JAK^{+}23, RFG24, SH25, WWC23, DČFW23]. eco-certification [RFG24]. eco-environment [SH25]. eco-label [AH21]. eco-labelling [JAK^{+}23, WWC23]. eco-scalar [Abo23]. ecolabels [CGGA20, GCV^{+}21]. Ecological [AIG^{+}20, ABH25a, ABM21, DG20, EMN^{+}24, FAC^{+}23, FBC^{+}23, GCS21, LDL^{+}23, MMM^{+}25, TZC23, AE21, Alm21a, AELR^{+}20, BC20a, BMK^{+}22, BCS^{+}24, CH25, CAVJ^{+}20, CVA22b, DGH25, DFC^{+}20, DBS22a, DR25, DRY22, DL22, Dun21b, EJC24, FAGS21, GAN^{+}25, GABA^{+}23, GTMP20, HRR^{+}20, HMG^{+}21a, HJ25, JTR^{+}20, JF22, JMB^{+}20, KKA^{+}21, Kar22c, Kıl24b, LPMS24, LGLL24, LMC^{+}23a, LYHL23, LZL25, LYL^{+}22, MMB^{+}23, CSdMR^{+}24, MYGM25, NGT^{+}23, OCCSC23, PBS^{+}23, PWL^{+}23a, PLWO24, QS21, RXN23, RXX24, RDLA^{+}20, SOO20, SRR^{+}21, SBD^{+}23, SCCW21, SVP^{+}23, STF21, VGC^{+}22, VAG^{+}24, WLX21, WYD^{+}22, XXLW24, YSL^{+}25, dRA25, NAP^{+}21]. ecologicalization [XHY22]. ecologically [KVL^{+}23, MFA^{+}24]. ecology [CCP25, HKZ25, dSJNS20, WCZS21, WCF^{+}21]. ecology-economy [WCZS21]. Economic [BDR20, Cha20a, DMK24, DMS^{+}24, FY24, GMGLV22, JGWM^{+}23, KL24, Kiz24, LrS22, LZL25, PV23, QTPL21, SGWL21, SOLN20, SPA^{+}21, TM23a, TGP^{+}24, AP24, AB20, AE21, ADR^{+}21, AVPAFM^{+}22, AVF23b, BCW^{+}25, BBFA^{+}24, BBC24, BDCJ22, CH25, CMC22, CGMH24, CDHY22, CLP21, DGG22, DRCdMCJ21, DAAP21, DME23, Ero22, EB24, EAA^{+}21, FZ25, FGPPGG22, FGPPGG23, FGC^{+}24, FFYE21, FMLMM23, Fro21, GLADGGVL23, GASLCP20, GWCD22, HFD^{+}20, HBP^{+}21, JL21a, JK22, JKKT20, KKA^{+}21, LP20, LXSX21, LW22, LXW^{+}21, LMR25, LASG^{+}25, LFPBD^{+}20, LRC24, Ma20, MYLV24, MMF^{+}25, OKFN24, PKR^{+}22, PALB20, PDJ22, Puh23, Qi22, Qi24, RGK^{+}20, SBD^{+}23, SCBB^{+}22, Sha20, SBW^{+}21, Squ20, SYY24, SRSGMGG21a, SRSGMGG21b, TMS^{+}22, TOHK20, VMV24b, VSD^{+}23, VPRR^{+}21, VDN^{+}24, WDG^{+}20, WCC^{+}20, WW22, WKK^{+}25, WB24, WCL25, WDC^{+}20, YSWW21, ZFGL20]. economic [ZT21, ZNA^{+}23, ZCA^{+}23, dOPM^{+}20, AAFK^{+}21, GWAPRPRU22, NAP^{+}21, SMC^{+}25]. economics [BHCL23, DMS^{+}24, Gro20, TW22]. economies [AED20, DCC^{+}20, dlFGFFR^{+}20, LRN^{+}22, RLRS21]. Economy [CMBB^{+}25, TFMT^{+}23, AAA^{+}21, AFR25, Als20, ANC^{+}25, BG20a, BBWC21, BAVV24, BGQ25, CS21, CDC24, CSRG25, DKKS21, DGI^{+}24, FRCR24, FFD^{+}23, GMFD25, GdACT20, Has20, HKZ25, HSN^{+}21, HCDN24, KCB24, LXSX21, LWG23, LHPJ23, LMR25, Lou22, Mal21, MFC^{+}24, MLB^{+}25, MCTZ24, NO25, PSD23, QYX21, RJ21, RXX24, RRM^{+}24, SdACdSLdAS24, San25, SBGR24, SHPW21, SGK21, SX25, SCB25, SR22, SPH^{+}20, Tho22, TT22, WLY21, WCZS21, WNDZ23, YCF24, ZZD^{+}23, ZLH21, AC23, CFM^{+}21, DZZC24, HWC^{+}25, LMMT23, MBG^{+}24a, MSHK25, SGWL21, SH24b, WMKG21, NJW25]. Ecosystem [CVH^{+}20, FRVB22, HGG^{+}22b, HKZ25, KS23, KN21, RRR^{+}23, WHPF20, ABH25b, BCW^{+}25, BYL25, CDD^{+}23, CMI^{+}22, CD25, CTH^{+}21, CLU^{+}24, CWKE20, CSGCD25, CSL21, CNG^{+}23, DKW20, De 22, DJC^{+}23, DA23, DLX24, EAGV20, GPG^{+}21, GCC^{+}21a, HRR^{+}20, HF20, HC25, HPPH22, HPC^{+}23, HZC^{+}24, LLT^{+}20, LW22, LH25, Liu23, LRC24, MUR^{+}21a, MKAF21, MFC^{+}24, MWWS21, MKB21, MZT^{+}23, NO25, OVSVPP^{+}22, POM24, PGA^{+}20, PS21, QK24, QRK^{+}23, RFG24, RR20, RSFvH^{+}22, RKS^{+}23, RGH^{+}22, SAO^{+}21, SWI^{+}22, SL20, SAC25, SMC^{+}25, THGR^{+}20, Tra23a, WD23, WWD24, ZCG^{+}23, dSOW^{+}21, dO23, vTHS21, JL21d, LPMS24, OI23, OKS^{+}22, SLP^{+}22]. Ecosystem-based [HGG^{+}22b, KS23, WHPF20, ABH25b, CLU^{+}24, CSL21, EAGV20, GPG^{+}21, HC25, LLT^{+}20, LH25, Liu23, MUR^{+}21a, OVSVPP^{+}22, PGA^{+}20, PS21, QRK^{+}23, RR20, RKS^{+}23, LPMS24]. ecosystems [BMR20, BMMF22, COM20, DMG^{+}23, GCG^{+}21, HGG^{+}22b, KAA^{+}24, Lev20, Lot22, OKR^{+}20, QUK21, QLU^{+}23, QK24, SOK^{+}22, TMAS24, UPBB25, dO23]. ecotourism [CLW^{+}21]. ecotoxicological [UPBB25]. Ecuador [COJE24, DRCdMCJ21, LASG^{+}25, MTC^{+}23, vNdB25]. Ecuadorian [MTC^{+}23]. Ed [LPO25, Rub24, Smi20]. edge [Bre25]. edged [Gug24]. edited [SSSL25]. Edition [SSSL25, Tan23b]. Editorial [Als20, BG20b, AEJ24, ZT23b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano21o, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano22o, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano24o, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n, Ano25c, Ano25d, Ano25e, Ano25f, Ano25g, Ano25h, Ano25i, Ano25j, Ano25k, Ano25l, Ano25m]. Edmunds [Ste25]. eDNA [CCS23]. Eds [Cha20b, De 20, Lv20, NJW25, Sos21a, Tan23a, Tan23b]. educating [RB20]. education [FLB^{+}21, cL22, OBNK24, WSD21, WN25]. edule [BASM22]. Edward [Cha20b, Smi23b, SSSL25, Sos21a, Tan23b]. eel [KYMW21, MHT^{+}24, SK24, SFN^{+}21, WFK^{+}24]. eels [PHP25]. EEZ [ABI24, GAD^{+}24, Hor25, Rog24, TYS22]. EEZs [SPNP24]. Effect [CQ20, ASG^{+}24, CRCP^{+}23, DA23, GAMR20, HSHS24, PKR^{+}22, PAK22, QTPL21, RAB22, SKL^{+}20, XLLS21, XSCL21, YTÇ22, ZYTT21a]. Effective [CGAF25, CZ24, YCL22b, ABM21, AGM^{+}22, BvdH23, CJN25, CdLN^{+}21, DSB^{+}25, Dau23, EAA^{+}22, HHH^{+}23, JAH^{+}24, JH24, KMS20, MBDG23, OCD^{+}23, PKMT23, SCC25, TJ21a, VWFW22, WW21a, WMKG21, XXMZ23, ZD20b, ZBF^{+}22]. Effectively [DMHS25]. Effectiveness [OKNB22, SGLF^{+}23, ADK^{+}25, AHAB^{+}22, ADR^{+}23, BRKS25, CDC20b, CES20, Day24, Fel22, GM25a, GKP^{+}22, GDY25, KF23, RP23a, RP23b, SSDQ20, SFN^{+}21, SCY24, TAK^{+}22, VBd^{+}21, XWL^{+}20, YU20, ZPX^{+}22, ZCC^{+}24, ZABI25]. Effects [CCX23, GRSRSN^{+}24, GICM25, LT24, MJ20, OWS20, QS21, SHAC23, TTL^{+}25, WM21, BST23, CDNS23, CDC20a, CDB20, hBDF22, DME23, FGPPGG23, Fla21, GGL22, GLT23, HRR^{+}20, JK22, KBD^{+}20, LKE^{+}22, LM24, MVSP^{+}21, MGK^{+}25, NMN^{+}25, NUD^{+}23, OCT^{+}22, PGK22, Sha20, SSB^{+}22, SO20, TDC21, ZLG^{+}24]. efficacy [ALML23, HSHS24, PK23]. efficiencies [Cro21]. Efficiency [COJS^{+}22, GABSD21, ÁCT20, BDD^{+}21, DG23, GAMR20, JL21a, KCH^{+}25, LWG23, MWB^{+}21, QS21, Ren21b, SYA25, SLY21, VPRR^{+}21, XHY22, ZW21, ZLH21, OAM20]. efficient [DG24a, DG24b, PCCR^{+}25]. effort [AKF^{+}25, BCJ21, EECE^{+}24, HBJ21, MAG^{+}22, MM20, Obu20]. efforts [ABDRV^{+}23, CJLJ24, FMS^{+}22, LC23, MGPR^{+}25, NMN^{+}25, WW22]. egg [HM22b]. EGT [Alf20]. Egypt [AHBAKH24, AHENNA^{+}24, SK20b, TCC^{+}22]. Egyptian [ETSKA25]. EIA [LZ24]. Eilat [YNZ22]. eke [MAR^{+}20b]. elasmobranch [HHBC20, KKAO^{+}25, KGL22, BdLAdC^{+}20]. elasmobranchs [ADM^{+}25, GAD^{+}20, WLCG23]. electric [Pen22, RPK23]. Electronic [BDC^{+}21, GDLC20, MFG^{+}23]. element [VYB20, BAP21]. elementary [cL22]. eleven [CQ20]. Elgar [Cha20b, Smi23b, SSSL25, Sos21a, Tan23b]. Eliciting [CLB^{+}23, VDOE^{+}24]. eliminating [FR23]. Elise [Niu22]. Elmina [GE23]. Elsevier [De 20]. elusive [DGH25]. embayment [MCW^{+}22b]. Embracing [BvdH23, NMCL24, WSZ^{+}25]. Emergence [Sah22, PV25, SVF20]. emergency [BMRM23, LC20b]. Emergent [CMWT24, NWS^{+}24]. Emerging [CLRW22, MWM^{+}23a, ORdS25, RSS^{+}20, Dyr25, GPA^{+}22, HSGS^{+}23, LV23, Pan25b, TFMT^{+}23, YZH22, YCF24, Tan23a]. Emirates [Gul20, Squ20]. Emission [CGMH24, WLXY22, DXS^{+}21, GMC24, HRS^{+}24, JWX^{+}20, Okş23, Tes20, VGK23, WZZC21, WW22, Wan24a]. Emissions [LLW22, AHN25, Bau24, BAD21, CMC^{+}23, Che21b, CDHY22, GZW^{+}20, GICM25, HKMS20, HSSM24, Kon20, MVSP^{+}21, NNR^{+}21, OOGK^{+}24, QLZ^{+}24, SSK23, SADP25, VRBBM^{+}24, WKK^{+}25, WL22b, ZLPT23, Zhu22, Zhu23]. EMODnet [SSMG21]. emotion [BDLO23]. emotional [TNY23]. Emperor [BMR20]. Emphasis [Los23, AVPAFM^{+}22]. empire [TG25b]. Empirical [CC21, KF23, OGH24, AED20, CQ20, DMS^{+}24, Han20, cL22, LMR25, LJRHH21, NCD23, PSMR20, PPK23, RZ21a, XSCL21, ZYTT21b, ZCSK23, ZCS^{+}23, Pra20, RZ21b]. employed [KPN20]. employment [CDS^{+}25, LC20a, TTL^{+}25, WFG22c, WL24a]. Empowering [LLS^{+}20, SSB^{+}22, AHB^{+}23]. empowerment [FS22, dOJCSLdSB21]. Empty [HBL^{+}21]. enable [PGSFRMHH24]. enabled [Lin25, MWWS21]. enabler [WOG20]. enablers [MKM23]. Enabling [Col20, LMMT23, ZBF^{+}22, BLY^{+}24, MFA^{+}24, SGL^{+}21, CMOB^{+}20]. enact [Gal21]. Encounter [SHAC23]. encountered [JPM^{+}23]. Encouraging [HM22a]. end [BBBS22, WEF^{+}21]. end-of-life [BBBS22]. endangered [BFdSdFL^{+}25, BUM^{+}24, CMI^{+}22, Dam24, DPG^{+}22, GPH^{+}21, OVSVPP^{+}22, GMG^{+}24]. endemic [YSEK22]. ending [BBHN20, FTSG21]. endowment [RZ21b]. ends [MZ22]. endurance [KL21b]. energy [BST23, CA20, CKY22, DG23, DG24a, DG24b, GSR21, HHEJB20, JB25, KGP^{+}22, KMM21, KSO^{+}25, MBORGN24, MVSP^{+}21, MCRR23, MCTN23, OC20, ONA21, OKH^{+}24, SMB^{+}25, SZBA20, Wan24b, WNMS^{+}24, XLW24, vRB24, COJS^{+}22]. energy-efficient [DG24b]. energy-intensive [KMM21]. Enforcement [KAN^{+}23, Sha22, AGK22, ASM^{+}21b, Cha21, CZ24, DSB^{+}25, DG24b, FSKJ23, Hsi20, JH24, Lia25, Lin24, MWB^{+}21, MY24, Pan25a, Shi24a, Shi24b, SCZ^{+}24, SH25, Tan25, Tse21, VSV^{+}23, Wan24c, Wan24b, YW24, Zha25, ZX24]. engage [GSA23, MM23]. engaged [WWD23]. Engagement [MBB^{+}22, ZK21, BOOO23, BDC23, CRIMLS24, DG24a, EMN^{+}24, Has20, HMG^{+}21b, LHTR23, MHF^{+}25, MSHK25, OSBA24, RCM^{+}22, TMM^{+}23, WWD23]. engagements [WPvO24]. Engaging [GBSC22, RHG^{+}25, VPB^{+}24]. England [LL24a, SMB^{+}25, BD20, BHB^{+}24, CSGC22, DGS^{+}24, HHEJB20, SSDQ20, Smy24, WDC^{+}20]. English [ASKI20, CWMM^{+}21, NJBM^{+}23]. ENGOs [WYH21]. enhance [DMG^{+}23, GSM20, MAA^{+}24, TSA^{+}22]. Enhanced [EAA^{+}21, DSWB20, HKZ25, PSTW21, PdBA^{+}25]. enhancement [WCL25]. Enhancing [BG25, BMMF22, Can25, GLB^{+}24, HFY^{+}24, IAS25, MHF^{+}25, MMGAA^{+}21, PKH22, VSPMS^{+}22, BDR20, GGP^{+}22, MF22, SGH22]. Enjoy [dSRKTC24]. enough [GB21, MH24, SvLGK24, dlPdlLLC^{+}22]. Enrollment [JAK^{+}23]. ensemble [GCS21]. ensure [PK20a]. Ensuring [Liu25]. entangle [MTC^{+}23]. entanglements [AVF^{+}23a, MM20]. entangling [EMH^{+}23]. enterprise [LZZ22, Wil21b]. entitled [WL24a]. entrant [WFG22c]. entrepreneurialism [Zho23]. entrepreneurs [Knu25]. entropy [DG20]. entry [BDHI21]. Envelopment [PCD^{+}23]. Environment [APH24, LK24b, ADK^{+}25, AS23b, CA21, CCK^{+}21, CSHO^{+}20, DBBL21, DSV^{+}25, GPM20, Kuł22, KHLS23, LP21, MCBN22, MSSS22, MDG^{+}23, SH25, TMAS24, Tra23a, Dbo22, THF^{+}25]. Environmental [APH24, BAF^{+}23, CR23, CGHV20, CDC20b, DXS^{+}21, GG22, GKP^{+}22, GVR20, HM22a, JH24, JACD20, KMM21, LPO25, Lia25, MN21, PYVF24, RCM^{+}22, SGK21, SSSL25, Tan23b, YCL25, dMVJdAdSM25, AAKI21, AGM^{+}22, AR22, AKL^{+}22, AH21, BSY^{+}20, BDD^{+}21, BRM20, CMS20, CMMVdPV25, CS20, CQ20, DMd^{+}20, Dbo22, DUYEA21, Doe21, FY24, Fei20, FL24, GLADGGVL23, GG20, GR20a, GCC^{+}21a, GCL21, GMC24, Gul23, Guo20, HLW^{+}23, HSM^{+}23, HWY22, Jae20, JKV^{+}22, JCMD20, JL21b, KDHL24, Koj23, KDVR20, KBP24, LSH21, LC22, Liu25, LFVCF24, MLVT21, MPL^{+}23, MAK22, NNRS24, NWZ^{+}21, OKR^{+}20, OOGA22, PKH22, PC20, Puh23, RJ21, SC21, SCRCI24, SC24, SC25, Squ20, SH22, SRLLG23, TLT^{+}22, TOHK20, TML^{+}20, VGS^{+}20, VCC^{+}20, VBd^{+}21, WWCS21, WDG^{+}20, WCC^{+}20, War20, WB24, WFG^{+}22b, WN25, YJG24]. environmental [YS21, YZG23, ZW21, MFG^{+}22, PS21, dOSTB^{+}20, TGP^{+}24, TML^{+}20]. environmentally [FBC^{+}23, LLJZ22]. environments [AFHHD22, IMB^{+}24, SBJF^{+}21, SBT^{+}23, YMG23]. Envisioning [PCA^{+}23, VGC^{+}22]. EO [Ber25]. epidemic [LLHG23, LC20b, XSCL21]. Epinephelus [HDAPG^{+}24, RSLR24]. Epistemic [BC20a, VGA25]. equality [BGLTC21, Joh25, OKNB22, OBNK24, PGSFRMHH24]. equatorial [NOL^{+}23]. equilibrium [WLQD21, WHPF20]. equipment [HPH22, Okş23]. Equitable [TGF^{+}25, WLCC23, AVF^{+}23a, AKF^{+}25, BTE21, Dom23, GABA^{+}23, Gug24, JDF^{+}23, dlC24]. equitably [SCC25]. equity [ASC25, CMBB^{+}25, DGI^{+}24, FHBS23, GEGRdL24, HNA25, How20, KdBK25, LSOP24, MM22, OGS24, VFOA25]. era [KAC22, LLHG23, SS22, WSZ^{+}25, ZX24]. erosion [CLC21, FSF^{+}21, KBB^{+}25, NNRS24, SML^{+}25, TPAC24]. error [CBR21]. Errors [CBH^{+}24b, CBH^{+}24a, CBH^{+}25, SCM^{+}25]. ESC [FFYE21]. escalating [Kar22b]. escalation [KL21b]. escape [KM25, SKL^{+}20]. escaped [CAHR^{+}24]. Escapement [MKM^{+}20, ADR^{+}23, SKL^{+}20]. escapes [CAHR^{+}24, YHU22]. especially [ZJ22]. essential [LHV^{+}20]. established [QTPL21]. Establishing [AC23, KMTvT25, Liu24, LLA^{+}25, Wat22, WN25, BN20, DTBP21, HFY^{+}24, WWG^{+}23, WWG^{+}25]. establishment [CLL22a, KMM24, SAO^{+}21, TG25a, WF25]. estimate [GDB^{+}24, Kaw22, MNS^{+}24, SRSGMGG21a]. estimates [RWVH21]. Estimating [AOS^{+}24, BSLRR24, BBC24, CDD^{+}23, SRSGMGG21b, TSA^{+}22, WCT^{+}20, XSCL21, ZMR20, JV20]. Estimation [SKE22]. Estonia [LKB^{+}23]. Estonian [SK20a]. estuarine [AS23b, ZCC^{+}24]. estuary [DPP^{+}20, SCBdA20, SCA^{+}22, TMBM^{+}25]. ethical [GCL21]. ethnobiological [NEH^{+}23]. ethnography [GMRSP23]. ethos [Dun24]. EU [ABH25b, BDTD25, BGQ25, CHEC21, CCGM21, CdJK24, DL21, DN20, DMG^{+}23, EB23, Faj22, GMPT^{+}20, GCC^{+}21b, GICM25, IGO21, KL21a, KA21, LGP25, LFVCF24, MPD^{+}23, MGK^{+}25, MMGAA^{+}21, OGS25, OSV21, PKB^{+}21, PMK^{+}21, Pen22, dlÁPAFGO20, Puh23, SOLN20, Sta23b, SFN^{+}21, Tar25, THP^{+}22b, ZGR^{+}25a, ZGR^{+}25b, ZX24]. EU-level [CHEC21]. EU-Marine [MPD^{+}23]. EU28 [CMR21]. Eubalaena [MM20]. eucheumatoid [ADI^{+}25]. Eulachon [BMM^{+}20]. EUR [Rub24]. Europe [Smi23a, ALF^{+}22, BG20b, BB21, BRP^{+}23, dlFGFFR^{+}20, GGL22, LJ20, LJ23, OMC^{+}22, PY20, RHL20, SVV24]. European [RKKM24, Ada22, ARP25, Álv23, ACG21, Aqu23, ALF^{+}22, BDCJ22, BUM^{+}24, CVCV^{+}22, CAM^{+}23, CGGA20, DER^{+}21, FGPPGG23, FMGV20, FGC21, GBC^{+}20, GLAVL21, GSGOP^{+}21, GRA^{+}22, KL24, KvTNL20, KGP^{+}20, LFPBD^{+}20, MFRQ23, MGBS24, NVP^{+}23, OAdJC^{+}25, PMK^{+}21, PL23, PHP25, PV23, PŚ22, RPY^{+}24, SPFA^{+}20, SGR21, Sol24, SSMG21, SOLN21, TMG24, TDP20, VDP^{+}22, VSDR^{+}22, WBB^{+}25, ZWJvL20]. evacuations [BVT^{+}21, HVN21]. evaluate [BDG^{+}25, DQJ23, HLCM21, SA21, SNP21, SG20b]. Evaluating [AP24, DDAB^{+}23, CSL21, Dam24, GKP^{+}22, HBHJ20, KCCC^{+}21, MVR21a, MHS^{+}25a, MHS^{+}25b, MLM23, PESO24, RA20, RP23a, RP23b, RHC20, SII22, SHP^{+}21, Squ20, BTV^{+}22, PKB^{+}21, SLH22, SRS23, HM21, SSDQ20]. Evaluation [AVF23b, COJE24, FWL^{+}22, IBC^{+}20, JLY23, LWX^{+}20, LXW^{+}21, MBORGN24, ZCCQ22b, ASGS23, BDVR25, DSL20, DG20, EHM^{+}20, EHM^{+}22, HTH25, KMS20, KOBS21, LYL^{+}22, LYLZ25, LFLF22, MZ22, MM23, Mor20, NTH^{+}21, OWS20, Ren21b, SGWL21, TM23a, Tan24, TSS^{+}24, ZPX^{+}22, ZLD^{+}21, vdBSvdV^{+}23, YU20]. evasion [WOG20]. eve [AR22]. event [HGQJD24]. events [Cao25, SMKS22, SPA^{+}21]. Every [SCBB^{+}22]. everyday [dZ20]. Everything [FR23, SCF^{+}22]. Evidence [AHENNA^{+}24, AFR25, BK23, BWV25, CVA22a, DSD^{+}24, FSRC25a, FSRC25b, GÁSVPS^{+}25, LTW23, LZL25, NBB^{+}23, PGP22, PSA^{+}25, SBP^{+}25, SPG20, Tan22, TSS^{+}24, XLW24, ZT23a, CBB^{+}24, DMK24, DFP^{+}25, DPG^{+}22, DNR21, GAD^{+}23, GM25b, KSÖB25, MLB^{+}25, NRYG20, OGH24, QLU^{+}23, RJ21, TSH^{+}22a, YCF24, ZZW^{+}21, ZYTT21b, AAW^{+}23, CACJR^{+}21, CDG^{+}21, dHJTGO21, JL21b, LCC20, NAN^{+}24, THT^{+}23, UOO20]. Evidence-based [AHENNA^{+}24, CBB^{+}24, MLB^{+}25, QLU^{+}23]. evidence-informed [GAD^{+}23]. evidencing [CMS^{+}21]. Evolution [DSJ^{+}23, DLL25, LXH^{+}22, LHPJ23, QYD20, SS23, YYL20, YYM25, AWX21, CVA22b, FGPVPP^{+}24, LXSX21, LJ22, PN20, QZZX20, RS24, Sah22, SCY24, YYD23, ZWG^{+}21, ZWC^{+}23, HLM^{+}20, YH20, ZSZ20]. Evolutionary [WXLD21, Alf20, Bir20, DLX21, JWX^{+}20, SVHV20, WLQD21, WL22b, XHZ^{+}21]. Evolutions [DPP^{+}20, FGC21]. Evolving [CBC24, KDHL24, KY20, NEF23, VYB20]. ex [AMIM21]. ex-vessel [AMIM21]. examination [Kıl24a, SAB^{+}24, TAK^{+}22, ZD20a]. examine [CLB^{+}20, Ren21a]. Examining [AL20, CCCC21, Lin25, NHD23, SCSR^{+}23, SPJB22, WWD24, SRH^{+}22, AELR^{+}20, SCSR^{+}22]. example [AB20, BGMCG24, COJE24, Gir22, KMS20, LC20b, MMNC22, OSV21, Sam22, Wan23a, XL25]. Examples [BI21, GA20]. exceptionalism [PH24]. exchange [TRL^{+}25]. exclusion [vLS24]. exclusionary [CdJK24]. Exclusive [SOLN20, BCW^{+}25, CLP21, OKFN24]. executed [TRL^{+}25]. exempted [BBJ^{+}22]. exemption [LK24a]. Exemptions [Dom23]. exhaustion [TNY23]. exist [Gar20]. existence [WJB^{+}20]. Existing [COJS^{+}22, CSHO^{+}20, FGK^{+}20, FMS^{+}22, GRPRVG23, RB20]. expanding [BvBN^{+}20, CVA22b]. expansion [AHAB^{+}22, CFM^{+}21, Cha20a, CMS^{+}25, DPSF25, Dun21a, LHTR23, POM24, Tra23b]. expectation [RHC20]. expected [SSAAB^{+}24]. expedited [FXC^{+}22]. expenditure [BBV^{+}25, NO25]. experience [ACG21, BAVV24, HWY22, KKO^{+}22, LCMBGP22, MKS21, RF25, SBT21, THF^{+}25, ZLZ21, DFTK20, OGGC20, Tse21, YYL20]. experienced [AO20]. Experiences [YS23, CdJK24, CWLD25, INW20, LXH^{+}22, OBNK24, PGSFRMHH24, RB20, SCB25, ZTW23, LEVET21]. experiment [AFRFGP25, CDD^{+}23, CLL20, CKY20, DMK24, FRDK24, GdACT20, KNY22, MZT^{+}23, RXX24, SNP21, SL20, TOAPJ21, WLX21, WYD^{+}22]. Experimental [GM25b, NRYG20]. Experimenting [CQAS^{+}25]. Expert [CPV^{+}21, KHLS23, BVT^{+}21, FMTA^{+}25, OC20, dRA25]. expertise [ZD20b]. Explaining [IAB^{+}20, VM21]. explains [SOS20]. explanation [SOS25]. explanatory [BSM23]. explicit [JL24]. Exploitation [LPO25, ML22, ASPT24, GdCC^{+}25, Gue24, HS24, HSM^{+}23, LWX^{+}20, NOL^{+}23, PGSS22, War20, YS21]. exploited [MN23, OOM^{+}25]. Exploiting [SAS^{+}25b, BDDE20]. Exploration [BDC23, LPO25, LNRM23, BL22, MZZ24, NEF23, VJL^{+}21, WCL25]. Exploratory [SPT^{+}20, BvdBvdFK20, FS21, NGP^{+}21, TMT20]. Exploring [ÁTGCE^{+}25, CGP23, DGMF^{+}20, FZL25, GFD21, HWT23, HMG^{+}21b, KSÖB25, Kas21, LTSL22, LJ23, ML21, MMG^{+}25, PSD23, PL24b, Pra24, QRK^{+}23, RRR^{+}23, SBP24, SLSV25, SPNP24, SRS^{+}25, VWJMAJ20, WSD21, XZ23, YW20, dSdLSGB22, QCM^{+}20, RKC^{+}24, GSA23, KK22, KL21b]. export [FZYF23, VKN20]. exporting [BA25b]. exports [MFAH25]. expose [SFN^{+}21]. exposure [BBV^{+}25, CBR21, KDH^{+}21]. expressed [Wat25]. extended [DJC^{+}23, DLYM22, HLCM21, MLM23]. extension [MM23, TG25b, ZJ22]. extent [ABH25a, CQAS^{+}25, CMOJ^{+}24, LW22]. external [KA21, PGK22, SCLM^{+}21]. externalities [GAGLS25, GLMBMP25]. extinct [JG20, VCC^{+}20]. extinctions [OOM^{+}25]. extractive [ANC^{+}25, VVR^{+}22]. extractivism [Gal21]. Extraterritorial [JH24]. extreme [SPA^{+}21]. eye [BHEE22, NBB^{+}23].

Fabinyi [AFHM23]. face [AH23, CDC24, hBDF24, LL24b]. faced [DSL^{+}21, DSC^{+}23]. facilitate [AANM^{+}22, CJN25, Han21, OGH24, STIS20, SBJF^{+}21]. facilitated [LC22]. Facilitating [VKL^{+}25, KMT^{+}21]. facilitation [AHR24]. facilities [BÅ25a, Osm23, PB20]. Facing [AHB^{+}23, DC20, DFP^{+}25, DP23, MKM^{+}20]. Fact [HGSD23]. factor [DL22, RZ21b]. Factors [HMFH20, KWMM20b, LSY20, MLMD21, PPK23, AAKI21, Bau24, BC20b, CT24, CDC24, Dam24, lTS25, DLX21, FIB25, GLADGGVL23, HP23, IGO21, JZW21, KKDK20, LMR25, NEF23, PGSFRMHH24, Pra24, RE21, SGL^{+}21, TMS^{+}22, WFG22a, XHY22, XXLW24, ZWXL21, ZSLY20]. Factory [RBNE20]. Factory-made [RBNE20]. FAD [MSZ^{+}23, MZT^{+}23]. FADs [MZT^{+}23]. Fai [LK24b]. fail [CBR21, HM22c, MS21]. failed [Åm21, OSV21]. failure [GVL22, MF23, Pol23, SOS20]. failures [SSP^{+}21a]. fair [Dau23, dlC24]. faire [KK22, YL21]. fairwashing [SSW^{+}25]. fairways [PYVF24]. Falklands [AM22]. Falklands/Malvinas [AM22]. fall [GABSD21, SCF^{+}22]. fallout [HS22]. Family [MLVT21, PNS^{+}23, TNY23]. Family-run [MLVT21]. fan [SPCCvT21]. FAO [SFGEPC22]. farm [BST23, BBO21, CSLGFV20, FAC^{+}24, HRR^{+}20, JvH24, RBNE20, SBT21, TJZ^{+}21]. farm-induced [FAC^{+}24]. farmed [CAHR^{+}24, EHLT21, KM25, SLMG^{+}21, STG^{+}20]. farmer [AAW^{+}23, MSHK25]. farmers [ABC^{+}25, RKR20, TRN^{+}24]. farming [ADI^{+}25, ARP25, BE20a, CBC24, CSV^{+}24, FGPPPV21, Gir22, GVR20, HO25, HSN^{+}21, KHdSM^{+}22, KWY^{+}22, LG24, LJ22, MTS^{+}24, Mrd23, SCLM^{+}21, SLS^{+}20, YK20]. farms [BvdBvdFK20, BLY^{+}24, CCW21, DWC^{+}20, MMF^{+}25, PKMT23, SLS25, SVKP21, TYS24, WML25]. Faroe [Chr20, DAA21, EB24, Mrd23]. fast [LDG^{+}25]. fast-growing [LDG^{+}25]. fatalities [NSPRSS20]. fate [AIHR^{+}25]. Fatigue [RE23, ZTW23, ZTM^{+}23, CRIMLS24, PGK22, ZWJvL20]. fauna [SS23]. favour [dlPC24]. favourable [Che21b]. fear [Pol23]. fears [WM21].

Feasibility [HJ25, WDG^{+}20, ZL24, BDG^{+}25, MNHS25, ZFGL20]. feature [AH23]. features [PSTW21, YWY20]. February [Ano20o, Ano21p, Ano22p, Ano23p, Ano24p, Ano25n]. federal [ALML23, NH20, WEF^{+}21]. Federating [KBP24]. feedback [FTW^{+}23]. Feeding [AWX21]. ferruginea [OVSVPP^{+}22]. fertilizers [ADI^{+}25]. fiction [HGSD23]. field [PV25, Wil22]. fifth [TMDN23]. fifty [Smi23a, SRL25]. fight [Gar22, GY23, Oan22]. Fiji [GSB24, RGK^{+}20, TPG^{+}23]. Fijian [YS23]. fill [SSOP^{+}21]. Filtering [Wil22]. fin [TOAPJ21]. final [HHD20, MTN23]. finance [HH25, SLY21, TMAS24, WSD21]. finances [KJS^{+}22]. Financial [Ero22, PALB20, WLY21, AAW^{+}23, BDHI21, EP20, KB20, KWMM20b, KTI^{+}24, PESO24, PPK23, Tan22, TMAS24, ZZW^{+}21, ZLZ21]. financialization [SS22]. Financing [CTR24, BAVV24, SWHT22, TT22]. Finding [Gue24, Gug24, hC24]. Findings [KNY22, PBD^{+}23, TMT20]. Fine [EBB^{+}25]. Fine-scale [EBB^{+}25]. finfish [CGGA20, HHB^{+}21, KHdSM^{+}22, LHTR23]. Finland [PKH22]. finning [GBDC24]. Finnish [HvT22, HJ25]. Fins [BMF20, BDTD25]. FIPs [SSW^{+}25]. Firm [SSK23, BXV20, ZYTT21a]. firms [KLSG21, QZZ20]. First [DPG^{+}22, LPMG^{+}24, APF22, BBK21a, BBB^{+}24, BGK^{+}22, BMR23, GPH^{+}21, GBC21a, LKE^{+}22, MZT^{+}23, PBD^{+}23, SBT21, ZLD^{+}21]. fiscal [Chr20, JL21b]. Fish [LL25, MAK22, MMW^{+}23, OGH24, SBD^{+}23, SGZH^{+}21, SOS^{+}21b, ABC^{+}25, AED20, ABRA20, AP23, AFMdSdTdCC22, AAW^{+}23, BEM^{+}20, Bea22, BHdMCM^{+}21, BLM24, BA25b, BUM^{+}24, BTE21, COJE24, CMNH21, CDG^{+}21, CTDY25, DPB^{+}23, hBDDF22, DF24, EECE^{+}24, EMH^{+}23, FGPVPP^{+}24, Fox22, GSG^{+}24, GCV^{+}21, GVR20, GAA^{+}20, HNA25, Han20, HMG^{+}21a, HS23, HGH^{+}20, Hor25, IO23, JY20, KCT^{+}24, KL24, KM25, LG24, cL22, MFG^{+}23, MSO23, Mar21b, MN23, MDG^{+}23, MMK^{+}22, Nal24, NTH^{+}21, NAS^{+}22, OOSG21, OVA22, PT23, PDJ22, RAFLRP21, SCBB^{+}22, SKL^{+}20, SOS20, Sta25, SHFP21, dSTHL^{+}24, TCC^{+}22, TJZ^{+}21, UISM^{+}23, VM21, WB24, WLRH^{+}20, YFD^{+}21, YHU22, ZTZ^{+}21, ZMM^{+}23, CMT^{+}23, MSZ^{+}23, PMT^{+}20]. fish-exporting [BA25b]. Fisher [AGK22, AFMdSdTdCC22, BFdSdFL^{+}25, BUM^{+}24, HRMPCRRA22, LSOP24, MNGM24, PBS^{+}23, SMLRB^{+}21, AHT25, AACdlT23, AS23a, BEM^{+}20, CRIMLS24, CdJK24, DSD^{+}24, FKNG25, FSRC25a, FSRC25b, FMTA^{+}25, GSAF21, GRPRVG23, GMRSP23, KTI^{+}24, Li23b, NMN^{+}25, OCZ^{+}22, OSZ^{+}25, PGSS22, RPK23, RCM^{+}22, RSH^{+}25, SPG20, SBFHD^{+}23, Sol23b, SCBdA20, TSH22b, VBd^{+}21, VDP^{+}22, WKN^{+}25, YCL22a]. fisherfolk [CJN25, MAA^{+}25, RBA23]. Fisheries [ASGS23, ALLS23, AGA^{+}21, Bra22, CdJK24, CTH^{+}23, DL21, DMG^{+}23, EB24, EHM^{+}20, EHM^{+}22, FSKJ23, FLB^{+}21, FMCT^{+}21, Gal21, Gar24, GMG^{+}24, GML^{+}24, HAD24, HFY^{+}24, INW20, JC23, Kim21, KMJ24, Li23b, LAF^{+}24, Mol24, NSZ^{+}21, NCE21, Pan25a, Per24, PV23, Run24, SOT^{+}22, SCC25, SKW^{+}22, VDP^{+}22, VCC^{+}21, VLL^{+}24, VYB20, WKH22, ZL20, APV^{+}20, AMIM21, Abo23, AANM^{+}22, ASV^{+}21, ABH25a, AOO^{+}23b, Alm21a, Álv23, ACV^{+}21, AEL^{+}23, ABRA20, AFMdSdTdCC22, AVPAFM^{+}22, Aqu23, ADM^{+}25, AMR21b, ALF^{+}22, AR23a, AR23b, AEJ24, Art20, ASGZ22, AH21, AL20, BCW^{+}25, BSBD21, BRSC^{+}24, BPA^{+}22, BKW^{+}24, BSLRR24, Bal22, BBP25, BMB^{+}24, BdSL25, BDM20, BBC24, BEM^{+}20, BB21, BJGG^{+}24, Bea22, BSSL^{+}20, BHdMCM^{+}21, BHKK23, BBS^{+}21, BCG^{+}24, Ber24, BFSP25, BJ25, Bet25, BDD^{+}21, BBJ^{+}22, BLY^{+}24, BDL^{+}24]. fisheries [BASM22, BGMCG24, BK23, BSK^{+}20, BDC^{+}21, BAD21, CPR21, CdLN^{+}21, CGB^{+}25, CPBO20, CCGM21, CRIMLS24, CD25, CTH^{+}21, CAVJ^{+}20, CAM^{+}23, CHB^{+}20, CDG^{+}21, CDHY22, COYS24, CLU^{+}24, CMOB^{+}20, CMSN^{+}22, CES20, CBH^{+}24b, CBH^{+}24a, CBH^{+}25, CHC^{+}21, CCG^{+}25, CRCP^{+}23, CDB20, CSL21, DLJ^{+}20, DKKS21, DGH25, DQJ23, DH22, DDF^{+}20, DFP^{+}25, DVRK25, DAN^{+}23, DSL20, DGS^{+}24, DLL25, DCC^{+}20, DA24, DDGdS^{+}21, DFF23, DHR23, DRS^{+}25, EAAG21, ETSKA25, EAGV20, EB23, EECE^{+}24, EMS^{+}23, FZ25, FARBV^{+}24, FGF^{+}24, FRCR24, FS21, For22, FS22, FWL^{+}22, Fuj21, FHBS23, GSA23, GLAVL21, GLADGGVL23, GRSRSN^{+}24, Gar22, GSAF21, GLS^{+}21, GRA^{+}22, GDLC20, GCB^{+}23, GLB^{+}24, GAN^{+}25, GM21, GCL21, Gon25, GMMMRC^{+}25, GRvdM24, GBSC22, GMRSP23, GZW^{+}20, GMCMJA^{+}21, Gug22, GJS22, GICM25, GRPE^{+}21, GVH^{+}23]. fisheries [GV22, Guo20, GFL^{+}21, HMFH20, HDH21, HDCH21, HGSD23, HNA25, HWU^{+}23, HLW^{+}20, HYN^{+}23, Han21, HWT23, HWT24, Har22, HS22, HvdBH23, HF20, HC25, HW22, HZ23, HBR21, HBR^{+}23, HGLR22, HCSBM^{+}23, HTP^{+}23, HSdT21, HKZ25, HHH^{+}23, HT23, HvPPV20b, How20, Hsi20, HCS21, HBHJ20, HBL^{+}21, IKB21, IRK^{+}21, IAB^{+}20, JY20, JTR^{+}20, JSW^{+}22a, Jen23, JAdS^{+}20, JKKT20, KWS20, KWF^{+}21, KA21, KYMW21, KAA^{+}24, KMS20, Kao23, KOBM24, KFB^{+}20, KBH^{+}24, KB20, Kay25, KFP20, KSB^{+}23a, KSB^{+}23b, Ken20, KKAO^{+}25, Kıl24b, KPN20, Kim23, KHH^{+}23, KWMM20a, KWMM20b, KLSG21, Kob22, Koe22, KH24, KAN^{+}23, KSV^{+}23, KJS^{+}22, Lat25, LBL21, LRN^{+}22, LIK^{+}24, LBWS23, LLLF22, LCAX21, Liu23, LYLZ25, LA23, LSGG^{+}20, LASG^{+}25, LGW^{+}21, Lov23, LLS^{+}20, LSD^{+}22, MMM^{+}22, MMK^{+}23, MKM23, ML21, MLVT21, MF22, MMB^{+}23, MVR21a]. fisheries [MZ22, MCPS20, MWM^{+}23a, MHS^{+}25a, MHS^{+}25b, McC20, McC21, MAK22, MM22, MH20a, MKS21, MHN^{+}20, MMSB^{+}20, MAG^{+}22, MS21, MPSF20, Mrd23, MIH^{+}24, MSGAR20, NH20, NNM^{+}21, NHA23, NSK20, NAS^{+}22, OCZ^{+}22, OR21, OYKBA22, OGM24, OOGS20a, OOGS20b, OOGK^{+}24, OI23, OCD^{+}23, OCCSC23, OSZ^{+}25, OAM21, OMC^{+}22, OA20, OEA25, PCA21, PR24, PRM^{+}20, PFG^{+}21, PSPR20, PHD^{+}24, PDD^{+}25, PL20a, PVH^{+}21, PGSFRMHH24, Pen22, PCCR^{+}25, PWL^{+}23a, Pet22, PCK22, PJN24, PGGA22, PALB20, PCCOS24, PLOG24, PKS20, PFWM23, PKMT23, QTPL21, RSK^{+}25, RLS20, RRS25, RYS^{+}21, RCM24, RMU^{+}22, RROP23, RTMCVMDE22, RKS^{+}24, RHO24, RDLA^{+}20, RSH^{+}25, SAO^{+}21, SPFA^{+}20, SBP^{+}25, SOO20, SK20a, SK20b, SIW^{+}21, SWI^{+}22, Sas25a, SLB21, SCJ22, SGR21, SBD^{+}23, SBM^{+}25b, SVV24, SSZ^{+}24, Sen25, SMYH22]. fisheries [Seu22, SWF^{+}24, SCM^{+}25, SSH^{+}23, Shi23b, SKE22, SS22, SS21a, SSV^{+}23, Ske24, SHP^{+}21, Smi23a, Smy24, ST24, SLGAG24, SLSV25, SS21b, SPMN23, SGL^{+}21, SG21, SH22, SH23, SA25, SRH23, STKC22, SRH^{+}22, SHS^{+}25, SHFP21, SSP^{+}21b, SG20c, SKL23, STF21, Szy20, TNE^{+}24, Tan25, TTA^{+}20, Tel23b, THP22a, THGR^{+}20, TRL^{+}25, TSH22b, TLT^{+}22, TGF^{+}25, TOL^{+}22, TKP23, TEA^{+}22, Tra23a, TMTC20, TK22, TDP20, TTL^{+}25, UOO20, Urb22, VMD^{+}24, VSPMS^{+}22, VJÓK25, VGA25, VGC^{+}22, VSDR^{+}22, VGAH^{+}21, VLS^{+}23, WFK^{+}24, WS21, WHPF20, WKK^{+}25, WLRH^{+}20, WCN^{+}25, WH20, WWC23, WM22b, YCL25, YZL24, ZLG^{+}24, ZV20, Zha21, ZJ20, ZCCQ22b, ZX24, ZCC^{+}24, dJRFVOA24, dSVT^{+}25, vHSSK20, vdGD21, AMR21a, CN24, DDGdS^{+}21, DFTK20, EAAG21, FS22, GWAPT21, Kay25, LJL22, MPSF20, RM21b, WBPG20, Zho23, Fla21]. fisheries-based [KOBM24]. fisheries-management [BEM^{+}20]. Fisheries-Related [HFY^{+}24]. fisherman [LSY20]. Fishermen [Cao25, PLOG24, BC20b, BMK^{+}22, BGQL24, GWCD22, dSJNS20, LSY20, Mat20, NSM25, SKW^{+}22, TYS24, WGBS25, YSL^{+}25, ZWXL21, dOPM^{+}20]. Fishers [CAVJ^{+}20, CWMM^{+}21, Gue23, ONN^{+}20, AP24, APGWTM^{+}24, BBV^{+}25, BSI^{+}21, BAJR24, CPR21, CJ23, CRJ22, DMd^{+}20, DSB21, DME23, GGZRAHVS25, Har21, IOHC22, IAS25, JS23, KPN20, LBWS23, LEERRM^{+}21, MRKM23, Mar21b, MM23, MS21, Muk21, MSHK25, Nal24, OCD^{+}23, PRM^{+}20, Pra20, PDJ22, RTMCVMDE22, SSAP24, VPB^{+}24, WTO24, YS23, ZMT^{+}25]. fisherwomen [BBP25, FAMW^{+}25, SSB^{+}22]. Fishery [CMC^{+}23, JAH^{+}24, JK22, Mot20, WS23, AWX21, ABDRV^{+}23, AE21, APGWTM^{+}24, ASM21a, AP23, AMR21a, AONF21, ÁTGCE^{+}25, BCBS22, BLH^{+}20, BDHI21, Ber22, Ber23, BJ25, Bet25, BCJ21, BFMS21, BM21, Bjø23, BLM^{+}25, Bot20, BKG^{+}25, BRR^{+}22, CSMMC^{+}25, CC21, COJE24, CMR21, Cao25, CMC22, CdLN^{+}21, CCB^{+}21, Cha20a, CdJK24, CCKK22, CSHO^{+}20, CNB^{+}23, CHC^{+}21, CSCGC20, DGG22, DSL^{+}21, DRCdMCJ21, Doe21, DFF23, DMT25, DAAP21, DD22, EJC24, EVPMKP22, EMH^{+}23, FA24, FGPPGG22, FGPPGG23, FLB^{+}21, FRLMW25, FOH^{+}20, GMPT^{+}20, GRSRSN^{+}24, GdCC^{+}25, GCC^{+}21a, GCS21, GHW^{+}22, GSB24, GSG20, GABSD21, GM25b, GWAPRPRU22, GCM^{+}23, GNMGWM21, HGG^{+}22a, HYN^{+}23, HZ23, HvPPV20a, HC21, JGWM^{+}23, JL21a, JF20, KdBK25, KOW^{+}25, KMJ24, KN21, LCC20, LLM^{+}23, LMC^{+}23a, LBWS23, LK24a, LLL^{+}24]. fishery [LGW^{+}21, LJRHH21, LJ21, LFLF22, LCMBGP22, MRKM23, MZVC25, MVR21a, MVR21b, MJ20, MNS^{+}24, MLA^{+}24, MAG^{+}22, MGK^{+}25, MCS25b, MH20b, MEG^{+}21, MM20, NAN^{+}24, OKFN24, ODG25, OAM20, PHF20, PL24a, PLLJ24, PESO24, PL20b, PZZ^{+}21, Pra20, PWL^{+}23b, PTCC24, QWBY21, RAMEG24, RSLR24, dSRKTC24, RAB22, RQB^{+}20, RMBR22, Ros22, RHO24, RHST21, SII22, Sak21, SB25, SFGEPC22, SC21, SCBB^{+}22, SDL^{+}22, SC22, SCGC20, SMKS22, SAS^{+}25b, SCCW21, SCF^{+}22, SdCMN^{+}22, ST24, SGZH^{+}21, SGH22, SBW^{+}21, SSW^{+}25, SG20a, SH24a, Sta25, SWXM21, SLK^{+}23, TM23a, TSS^{+}24, TMT20, Tra23b, TFJR25, TYS22, TPG^{+}23, UISM^{+}23, VMV24b, VJ24, VJO23, WJB^{+}20, WW21b, WW22, WWG^{+}23, Wen20, WNMS^{+}24, WN25, WNBRSG22, WLW^{+}22, YTBM^{+}22, YML^{+}21, YYD23, YXDL24, ZXGL21, ZCCQ22a, ZSZ20, ZZW^{+}21, ZLZ21, ZLD^{+}21, ZX24, dSF24, Bra22]. Fishery [BPJ^{+}24, CDG^{+}21, DAA21, GGMCMM^{+}24, SSW^{+}25, Mot22, WWG^{+}25]. fishery-based [HC21]. fishery-independent [CdLN^{+}21]. fishery-specific [GHW^{+}22]. fishes [AGN^{+}20, LLLF22, OOM^{+}25, SO20]. Fishing [AHT25, AM22, AFHM23, BST23, CCBB^{+}25, GAN^{+}25, Lin23, MB22, NGRM^{+}22, OAAMS24, Rog24, SFGEPC22, SBM^{+}25b, WCN^{+}25, AGN^{+}20, ABH25a, AGS20, ÁCT20, And24, BOBMG25, BV21, BJH21, BJ25, Bet25, BSI^{+}21, BCJ21, BBJ^{+}22, BMK^{+}25, CRP^{+}20, CBR21, CDNS23, CLS^{+}23, Cao25, CJH24, CMWT24, CSGC22, DGG22, DSB^{+}25, DT23, hBDF22, DXS^{+}21, DGIS^{+}22, DLX21, DA23, DP23, DAHS^{+}23, DMT25, DS22b, EJC24, EECE^{+}24, Faj22, Fen23, FARBV^{+}24, FRM22, Fla21, FIB25, Fro21, GRS^{+}23, GCF^{+}23, GÁPSAP^{+}23, GÁSVPS^{+}25, dlFGFFR^{+}20, Gar24, GSG^{+}24, GHW^{+}22, GAD^{+}23, GAD^{+}20, GE23, GASLCP20, GPTASM23, GY23, GR20b, GABA^{+}22, GSB20, GAA^{+}20, HCWC21, HDAPG^{+}24, HBP^{+}21, HM22c, JBGC25, JS23, JL24, KL21a, KOBM24, Kar22c, Kas21, KgJqY^{+}24, Kim21, KLSG21, KLS^{+}25, KDVR20, KAN^{+}23, LMC^{+}23a, Lov23]. fishing [Ma20, MLMD21, MMM^{+}22, MMC^{+}22, MCBN22, MKM23, MFAH25, MJMAA21, CSdMR^{+}24, MAA^{+}25, MM23, MGF20, Mil22, MLAS23, MdCLL^{+}23, MMGAA^{+}21, MHEW20, NMN^{+}25, NWS^{+}24, NBB^{+}23, NSPRSS20, Oan22, OGS24, OVA22, OA20, PESO24, Pat22, Pen22, PGM^{+}24, PHP25, PMC24, PPK23, PMT^{+}20, PV25, QCM^{+}20, RAMEG24, Ree20, RWVH21, RC22, RCM24, RM21b, Ros22, RDD^{+}21, SPFA^{+}20, SL23, SIW^{+}21, SDI^{+}20, SCD21, SLL^{+}24, SCSR^{+}22, SCSR^{+}23, SSZ^{+}24, Sha22, SNS23, SCGC20, SLS25, SSB^{+}22, SKL^{+}20, SKS21, SCH^{+}24b, SBT21, SGH22, SLSV25, SSP^{+}21a, SH23, Sta23b, SRY24, SCBdA20, SSF23, TOAPJ21, Tei23, dSTHL^{+}24, TSH^{+}22a, TMAS24, TSLK23, TJS^{+}25, TFP25, Tse21, TK22, TGM22, UE25, VMM21, VSD^{+}23, VAMR^{+}20, VPRR^{+}21, VVR^{+}22, WKO^{+}22, WNMS^{+}24, WBH^{+}23, WN25, WGP^{+}21, WLW^{+}22, YS23, YL21, YW24, YYM25, ZMT^{+}25, dJdMCL^{+}24]. fishing [dMVJdAdSM25, dMLB^{+}20, dlPdlLLC^{+}22, GMMMRC^{+}25, Smi23a]. Fishpath [CDB20]. fit [GNJR24, MBL24, WS23]. fit-for-purpose [GNJR24]. Fits [Mag23]. Fittest [OYKBA22]. five [CMGF20, cL22, CMGF20]. fix [Abo23, FDF^{+}21]. Fixed [FDF^{+}21, OAM21]. flag [BISSBdlMM22, CT24, DBZ24, Fel22, Kar22a, PSMR20, dlÁPAFGO20]. flag-state [CT24]. flagged [BLH^{+}20, HVN21]. flagging [ZD20a]. flagging-out [ZD20a]. Flags [PSMR20, Fel22]. flapper [GPH^{+}21]. flavescens [PBH^{+}23]. fleet [AIHR^{+}25, AP24, ATF^{+}21, ASGZ22, BDCJ22, CWMM^{+}21, Fel22, dHJTGO21, HCWC21, SPG20, SLL^{+}24, SLSV25, TMG24, VSD^{+}23, ZD20a]. fleets [HBJ21, MPL^{+}23, MLAS23, NSPRSS20, TFP25]. flexibilities [May22]. Flexibility [WTO24]. flexible [DBS22a, PVH^{+}21, VLG^{+}24]. floating [AVC24, ADR^{+}21, CSLGFV20, OR21, WML25]. flood [CBAPM20]. Florianópolis [FSF^{+}21]. Florida [AP23, BDMW22, DMK24, FMWW21].

Flows [GMG^{+}23, JKKT20, vLS24]. flux [VLL^{+}24]. flying [GWAPRPRU22]. FMEA [CAC23]. focus [BSSL^{+}20, LBWS23, Lia25, ZZMG^{+}24, BC20b]. Focusing [Sas25a, YK20]. folks [RF25]. Follow [Hor23, BGMCG24]. following [FAGS21, LCL20]. follows [Tan24]. FONDEPES [AACdlT23]. FONOP [Tod22a]. Food [BvdBvdFK20, DMG^{+}23, GPI^{+}24, YH20, BMF20, BRM^{+}21, CSM^{+}25, CTIH21, CGHV20, hBDF24, FTM^{+}22, FDVP24, GSB24, JG23, KCT^{+}24, LLS^{+}20, MF22, NRYG20, OOGK^{+}24, PBT^{+}20, RBNE20, SNK^{+}20, SN23, SSP^{+}21b, TFH^{+}23, TRN^{+}24, TGHF^{+}22, TCC^{+}22, TMTC20, WMS21, ZWC^{+}23, Gug22]. footage [CMSH25]. footprint [DGG22, LDG^{+}25, SLL^{+}24, UE25]. footprints [Kas21]. forage [THGR^{+}20]. force [Cha21, Zha25, Ngu23]. forced [SRY24, TOL^{+}22, YS21]. forces [FZL25, TJ21a]. foreign [ABI24, AED20, Kao23, Mar20a, OOGS20a, PSMR20]. foresight [PMT^{+}20]. Foresighting [KFS^{+}22]. forest [BYL25, HGG^{+}22b, Isl25, QTBL23]. forests [CKY20]. Forging [CCNGS25]. Form [Tan24]. formal [GBSC22, PL20b, SPHO23]. formalities [AHR24]. formation [DGH25]. forming [Zou23]. formula [VGA25]. Forsake [dSRKTC24]. Forsskål [SGPF21]. fortunes [BMF20]. forum [NT20]. forward [CQW24, CHB^{+}20, Chr20, DSJ^{+}23, JPM^{+}23, Tod25, VBH^{+}23, VGC^{+}22, Wan24c, Wil21b]. fossil [HSM^{+}23, vLM22]. foster [GM21, GMG^{+}23]. fostering [GABA^{+}23]. fosters [HSM^{+}23]. found [Lev20]. Foundation [ZD20b]. foundations [GEB^{+}21, Sta23a]. Four [BBFA^{+}24, KMM24, Ken20, SDL^{+}22]. Fourier [UE25]. fourth [MDJ^{+}22, War23]. FPIs [AE21]. FQA [FDF^{+}21]. Fragmented [Jen21]. frames [TDC21]. Framework [ANS^{+}20, CJL^{+}23, VKL^{+}25, GKP^{+}22, GABA^{+}23, KWK^{+}25, MPD^{+}23, MFRQ23, PKB^{+}21, PDB^{+}23, SLP^{+}22, VRBC^{+}21, VBH^{+}23, ASV^{+}21, ADK^{+}25, BN20, BGLTC21, BDG^{+}25, BDL^{+}24, BTN^{+}24, CH25, CMT^{+}23, CD25, CDRS22, CSVC23, CCOP25, COJZ^{+}21, DQJ23, DBS22a, Dbo22, lTS25, DJC^{+}23, DSV^{+}25, FFGJ22, FWSS^{+}25, GBM^{+}25, GE23, GDB^{+}24, HJ21, HWC^{+}25, HGL^{+}20, KF23, LZL23, LSGG^{+}20, LJRHH21, LLS^{+}20, LLL23, MAR^{+}20b, Mor20, O’D21, OGGC20, PL23, Puh23, Qi22, QUK21, QLU^{+}23, SWHT22, SS22, SG20b, SLH22, SCZ^{+}24, SRSGMGG21a, STF21, TJ21b, VÖB20a, VÖB20b, WZX24, WZL25, WPWM^{+}22, ZCCQ22b, ZLY23, ZCA^{+}23, vTHS21]. Frameworks [KSO^{+}25, ATM^{+}24, BFdSdFL^{+}25, BUR^{+}22, GLS^{+}21, KH24, KM25, NCE21, OC20, OT23, PY20, SRvNN21]. Framing [TDC21, BVvKL21]. France [CQAS^{+}25, CBAPM20, DPP^{+}20, DMT25, LCL20, PdBA^{+}25, TLYC20]. Francis [Ngu23, WQ25]. Francisco [THGR^{+}20]. fraud [JG23, LFVCF24]. fraudulent [SFN^{+}21]. Free [FXC^{+}22]. Freedom [Ngu23]. freight [Tar25]. French [SO20, BG20a, DMG^{+}23, FLB^{+}21, KLSG21, Kow24, TMT20]. frequency [NCF25]. frequently [CMI^{+}22]. fresh [JTdH25, RBNE20]. Fried [Lv20]. frigates [Dun24]. frontier [KGW20]. Frontiers [LK24b]. frustration [Pol23]. FTCs [VSD^{+}23]. Fuel [BAD21, CZW24, DGG22, FLLML25, GZW^{+}20, GICM25, HKMS20, HSM^{+}23, LDC^{+}24, MF23, RAB22, SC22, VSD^{+}23, WZX24, ZY21, ZGY21]. fuels [CDS^{+}25, HRS^{+}24, KGW24, vLM22]. Fukushima [FL24, LWZ23, LCL24, MXZ23, Mab24, PL24a, WL22a, WZL25, XTY22]. Fulfiling [LWZ23]. Fulfilling [BHK^{+}25]. fulfilment [KL22]. Full [PFG^{+}21]. fully [BFSP25, VLG^{+}24]. function [OAM20, Tan24]. functional [Lam21, TZZ^{+}21]. functioning [Kiz24]. Fund [VDP^{+}22]. funded [MPSF20]. funding [CMC22, JV20]. further [KOBS21]. Future [CHH^{+}25, CMMVdPV25, KSB^{+}23a, MZ24, AKFY22, AN25, AS23b, CCNGS25, CDC20b, CTDY25, DC20, DVRK25, Dyc24, EJC24, FGL^{+}21, Gun21, HRR^{+}20, HNB^{+}25, JTR^{+}20, Kas21, Kaw22, KFS^{+}22, KSB^{+}23b, KgJqY^{+}24, Liu21, LF25, LSD^{+}22, MMB^{+}23, MCRR23, MG22, MGNP25, Nal24, NJBM^{+}23, NT20, OSZ^{+}25, PSD23, PCCR^{+}25, RB20, SC21, SBGR24, SGL^{+}21, TSH^{+}22a, Ton22a, Tra23a, TWK25, WNDZ23, Wan23a, WCN^{+}25, WM22b, XT23, YH20, ZCZ20, ZLY23, Zhu22, ZMM^{+}23, dKMK21, vdBSvdV^{+}23, Fuj21]. Future-proofing [CHH^{+}25]. futures [AJDA24, PCA^{+}23, RHST21, SDL^{+}22, dlPC24]. fuzzy [CMR21, FSRC25a, FSRC25b].

G [Smi20, TMAS24]. G-20 [TMAS24]. Gakkel [KY20]. Galang [WMS^{+}25]. Galapagos [CD23, CMC24, CD25, TOAPJ21, ZMR20, vNdB25, CAVJ^{+}20]. Galicia [HPC^{+}23, dlÁPAFGO20, SBM^{+}25b]. Galician [BSLRR24, FGPVPP^{+}24]. gallina [BSBD21]. Gambia [NRF21, SN23, AANM^{+}22]. game [BM21, DLX21, GSB20, HZ23, JWX^{+}20, JCMD20, XHZ^{+}21, ZLGX20, ZHL^{+}23, ZY21, ZGY21]. gamma [Kaw22]. gap [AMR21a, Dav25, HSSM24, JV20, LTSL22, Li23b]. Gaps [ÓHCÓ24, AGM^{+}22, CLP21, DVRK25, DA23, FAC^{+}24, FBC^{+}23, JAH^{+}24, Lov23, MMAP25, MAG^{+}22, NAP^{+}21, PMM^{+}21, QK24, SSOP^{+}21, SRH23, WLS^{+}25, XZ23, YU20, ZSWJ23]. Gargoor [HGG^{+}22a]. Gas [Cha20b, BYL25, Che21b, HKMS20, HSSM24, MYGM25, Zhu22, Zhu23]. Gastropoda [OVSVPP^{+}22]. Gata [HSGNMGC21]. Gathering [SCCW21]. GATT [WL20]. gayi [RQB^{+}20]. Gaza [Qaf24]. GCC [Als20, AED20]. Gdańsk [COJZ^{+}21]. gear [ABDRV^{+}23, ABH25a, DT23, DSB21, DA23, DFF23, Fen23, FIB25, GRS^{+}23, GHW^{+}22, GAD^{+}23, HGG^{+}22a, LTJ^{+}23, Lov23, McC21, MPH^{+}23, NMN^{+}25, RWVH21, RCM24, SCSR^{+}22, SCSR^{+}23, TSLK23]. gears [SDG^{+}23, SKL^{+}20]. Gender [AOO^{+}23b, BGLTC21, GEGRdL24, GFL^{+}21, LRN^{+}22, MSB^{+}25, ASC25, CN23, DRS^{+}25, Joh25, LTSL22, MLE^{+}24, ML21, MM22, MMK^{+}22, OKNB22, OBNK24, PGSFRMHH24, PCSS^{+}21]. Gender-based [MSB^{+}25]. gender-blind [CN23]. Genderizing [Szy20]. general [DC23, Koj23, Shi23a, WHPF20, WWD23, WL24a, PCSS^{+}21]. generalised [vTHS21]. generation [MGK^{+}25]. generic [BCR25]. genetic [HGL^{+}20, LR20]. genus [SFGEPC22]. geo [PMT^{+}20]. geo-foresight [PMT^{+}20]. geoduck [LCMBGP22]. geoeconomics [Gov25]. Geographic [dSOW^{+}21, FLC^{+}23, FAC^{+}23, JMB^{+}20, KBD^{+}20, VKS23]. Geographical [HKKY22, Gir22]. geometric [CMR21]. European [LBWS23]. geopolitical [BSY^{+}20, NWS^{+}24, Pan25b]. Geopolitics [Smi23b, Mar20a, MLAS23, Øst24, WRL23]. Georges [WCF^{+}21]. Georgia [AM22, CHC^{+}21, Tra23a]. Geospatial [FDGM24, MDLA21, SBP24]. German [GAD^{+}24, GABSD21, ZD20a, ZD20b]. Germany [LG24, SBM^{+}25a, ZD20b]. GES [HvT22]. Getting [CBCA25, Obu20, SSDQ20]. Ghana [AE21, AVF^{+}23a, ABRA20, AAW^{+}23, AONF21, GE23, GAA^{+}20, KDVR20, MAA^{+}25, NSZ^{+}21, OR21, OCD^{+}23, OA20, OEA25, SAO^{+}21, SCBB^{+}22, Sta25]. Ghanaian [TJZ^{+}21]. GHG [CMC^{+}23, CBL24, HRS^{+}24, SADP25]. Ghost [DA23, And24, LTJ^{+}23, MPH^{+}23, TSLK23, YSEK22]. Giang [Han21]. giant [CJCS^{+}22, LZZ22]. Gibraltar [STS^{+}23]. gigas [CEMRMB^{+}24]. Gilbert [MGNP25]. Gili [PN20]. gillnet [Ber23, TK22]. gillnets [SGL20]. Gilt [LHW21]. GIS [CA21, CSLGFV20, JKV^{+}22]. Giving [Har21]. glacialis [MM20]. glass [KYMW21, WFK^{+}24]. glauca [DSL20, SSA^{+}23]. gleaning [PJN24]. Global [BDL^{+}24, BPJ^{+}24, FWGM24, HBR^{+}23, HGLR22, KKA^{+}21, NC25, Pyć23, SKDK23, TBCB25, Ada22, Bau24, BGK^{+}22, BJ25, Bet25, BBT^{+}25, BABC20, BvBN^{+}20, CHH^{+}25, Cla20, CTDY25, CNG^{+}23, DAB^{+}25, Dun21b, EAAG21, FLC^{+}23, FFGJ22, FGL^{+}21, GPA^{+}22, Gar24, GRvdM24, HCW^{+}23, Han21, Has20, HHBC20, HBR21, HMY^{+}23, KKDK20, Kaw22, KL24, Knu25, KM25, LDAC20, LMC^{+}23b, MZZ24, MSP23, MLB^{+}25, OAdJC^{+}25, PR24, PBS^{+}23, Pet22, PCK22, QF20, QUK21, RU20, RP23a, RP23b, RWVH21, Seu22, SGL^{+}23, SSH^{+}23, SS21a, SSV^{+}23, SC24, SC25, SPNP24, SZ24, SLY25, Ton24, TFP25, VGS^{+}20, VHS^{+}20, WvdMG^{+}21, WW21a, WBH^{+}23, YHX24, CLZ23, HDHM^{+}20, NC25, RBTW24, SBDC24, TCZ21, GZW^{+}20]. Globalism [CSM^{+}25]. globalization [TMAS24]. globally [Gal22, QLU^{+}23]. glocal [SVF20]. goal [JV20, PFWM23, STS^{+}23]. Goals [BCN23, FWSS^{+}25, HSN^{+}21, NSN22, WCM^{+}24, YW20, TML^{+}20, vSHHT20]. goers [Lev20]. going [SCF^{+}22]. gold [FTG^{+}23, SAS^{+}25b]. golden [Dbo22]. Goliath [CNB^{+}23]. Gomez [CMWN^{+}23, DBH^{+}22, GWÁV^{+}25, WvdMG^{+}21]. Gone [FAC^{+}24]. Good [PK20b, AIHR^{+}25, BDTD25, CKY20, DPB^{+}23, DUYEA21, FRCR24, GG20, KMT^{+}21, MH24, NNRS24, Puh23, dSRKTC24, dRA25, GKP^{+}22, MMGAA^{+}21]. goodbye [BBBS22]. goods [YHX24]. Gorringe [CLS^{+}23]. got [ALML23]. govern [Smi22, ZRR^{+}25]. Governance [CJ21, GCC^{+}21b, HWU^{+}23, KJE21b, LK24b, LJRHH21, MMAP25, Rus20, Smi20, SWGZ22, Sta23a, SGL20, SG20b, TNE^{+}24, VHSV20, YK21, YC21b, dPJ21, ASV^{+}21, AVF^{+}23a, Alf20, ANR^{+}22, BL23a, BPA^{+}22, BPE^{+}25, BJGG^{+}24, BCB20, BJ25, Bet25, Bir20, BOW^{+}22, BHKJT23, BÅ25a, Boy21, CC23, CBCA25, CVH^{+}20, CBB^{+}24, CGHV20, CJ22, COSS21, CCG^{+}25, CGP23, CM25, CMOJ^{+}24, DPP^{+}20, DGI^{+}24, De 22, DFP^{+}25, hBDF24, DGS^{+}24, DB20, DCC^{+}20, DHR23, EdLP^{+}21, EFFP23, EFCU25, FZL25, FRM22, FL24, GGP^{+}22, Gal21, GCL21, GBC21a, GBC21b, GRvdM24, Gro20, GJS22, GFL^{+}21, HIP^{+}21, HJ21, HM22a, HvdBH23, HC25, HSdT21, HSGNMGC21, HNMSG21, HGL^{+}20, JKCY25, JCMD20, JL21c, Jon21, dOJCSLdSB21, JvH24, JXWL25, KA21, Kat20, Kim24, KY20, LLT^{+}20, LJ23, LSH21, LTJ^{+}21, LJ21, LSY20, MM21a, MZZ24]. governance [MHS^{+}25a, MHS^{+}25b, MAR^{+}20b, Mba22, MMSB^{+}20, Mur21b, OPL20, OYKBA22, OI23, OCD^{+}23, OMC^{+}22, PLLJ24, PFG^{+}21, PN20, PKH22, PMN^{+}22, Pet22, dlÁPAFGO20, Pri22, PPK23, PMT^{+}20, Pyć23, QLZ^{+}24, RBTW24, RU20, RKS^{+}24, SMDM20, SvLGK24, SA21, SVHV20, SLB21, SPHO23, SMYH22, SNS23, Shi24b, SJ21, Ske24, SC24, SC25, SKW^{+}22, SOLN20, SSP^{+}21a, SMCW22, SCA^{+}22, SHS^{+}25, SHZC21, TUBA22, Tea20, TSA^{+}22, TFH^{+}23, TJ21a, TMT20, Tod22b, TFJR25, VTRS20, VJ24, VHD20, VHS^{+}20, VYB20, Wan23c, WA24, WB24, WFP25, WKBS20, Wu22, XLLS21, XXMZ23, YHX24, YCL22b, YC21a, ZL20, ZLPT23, dZ20, dRA25, Dun21a, KJE21a]. governed [SCC25]. Governing [PR24, STG^{+}20, Tod25, Mab24, PTS^{+}25, SRvNN21, STF21, Sos21a]. Government [Fan24, YZG23, CdLN^{+}21, CS21, CVA22a, Fla21, HWT24, HZ23, HWY22, KWMM20a, LLCH21, NO25, THP22a, VGA25, WYH21, WXLD21, WLQD21]. government-led [HWT24]. governmental [CMS20, HWY22, KL22, Pet22, dOSTB^{+}20, WMRB22, ZRR^{+}25]. governments [JL21b]. grabbing [GAGLS25, Mal21]. gradient [CSGCD25]. Grand [Ken20, LA23]. Grande [Ken20, dS21]. Granger [dHJTGO21]. grata [RROP23]. gray [COYS24, CZ24]. graying [HCWC21]. Great [DR25, MHC^{+}25, MCL22, DR22, Wat22]. Greater [WKH22, SAO^{+}21]. Greece [GPM20, Mar21a, Sak21, SMC^{+}25, TTL^{+}25]. Greek [GAD^{+}20, TDP20]. Green [FLLML25, HWC^{+}25, LWG23, LC22, RZ21b, Ren21b, RXX24, RKS^{+}23, SYA25, SA25, TMAS24, Tho22, TC23]. greener [Wan24a]. greenhouse [Che21b, HKMS20, HSSM24, Zhu22, Zhu23]. Greenland [DHU24, DHU25, LJ21, SOS20, SSAP24]. grey [SLK^{+}23]. grid [RQB^{+}20]. grief [MMM^{+}25, Pol23]. Grieg [MF20]. Griffiths [HH21]. griseus [NGT^{+}23]. ground [Gue24, TMAS24, UE25]. groundfish [ASGZ22, LFLF22, MSGAR20, SCCW21]. Grounds [NSPRSS20, JL24, PESO24]. groundwork [LGLL24]. group [GPH^{+}21, HW22, LSY20, MPN^{+}24, dlC24]. Grouper [CNB^{+}23, ASM21a, HDAPG^{+}24, MFAH25, RYS^{+}21, RSLR24, Squ20]. groupers [dMLB^{+}20]. groups [BC20b, DSL^{+}21, EBB^{+}25, FS22, MPSF20]. grow [Ber25]. Growing [AML^{+}21, FDVP24, HWMO20, LDG^{+}25, SLK^{+}23]. growth [AFR25, AB20, BBFA^{+}24, BBWC21, BHB^{+}24, DGH25, Dyr25, EAA^{+}21, FZYF23, IAB^{+}20, JK22, LXSX21, LXW^{+}21, LMR25, MCKZ21, SDI^{+}20, Sha20, SGK21, SKE22, WW21b, WKK^{+}25, WB24, ZT21, ZZD^{+}23, HvT22]. Guaceto [Rus20]. Guadalupe [MAMCHP^{+}20, SMHPMO^{+}21]. Guadeloupean [MLJ^{+}22]. Guamuchal [GBC21b]. guaranteeing [WS23]. Guard [LH24, Pan25a, Lia25, Lin24]. Guardamar [TPAC24]. guards [GC22]. Guatemala [GBC21a, GBC21b, OWCFDC21]. Guérinière [CBAPM20]. Guianas [LFLF22]. Guidance [DSC^{+}23, GLPM24, vdBSvdV^{+}23]. guide [BBBS22, GCM^{+}23, Wan23c, IMA22, FWSS^{+}25, SSAAB^{+}24]. Guidelines [CMBB^{+}25, AdSMS21, MHF^{+}25, NH20, NCE21]. guides [SKS21]. Guilds [HRMPCRRA22]. Guinea [HFD^{+}20, Kas21, OYECdOP24, PGM^{+}24, SMKS22]. guitarfish [SCBB^{+}22]. Guitarfishes [SSH^{+}23, CJCS^{+}22, SCBB^{+}22]. Gujarat [KSV^{+}23]. Gulf [ABDRV^{+}23, Cha24a, GÁSVPS^{+}25, HGG^{+}22a, MLMD21, MMC^{+}22, QTBL23, WLW^{+}22, ALLR^{+}24, SONB^{+}22, BCMCM21, CLL20, CAGBM22, CCKK22, DSB^{+}25, EECE^{+}24, GMPT^{+}20, GTA^{+}20, GBSC22, GWCD22, GHK20, Hon20, Kas21, OYECdOP24, PKH22, PGM^{+}24, SLP^{+}22, SSF23, TBCB25, TLT^{+}22, VCC^{+}21, WML25]. gull [TDC21]. Gunther [Cha20b]. GVC [Seu22].

Habitat [MFRQ23, PBH^{+}23, BDL^{+}24, CLD^{+}24, EBB^{+}25, GA20, HLW^{+}23, Lot22, MKAF21, NBB^{+}25, TMBM^{+}25, WLS^{+}25, WD20]. habitats [DGIS^{+}22, FARBV^{+}24, MPD^{+}23, NJBM^{+}23, SSMG21, TBPC20]. Hadjimichael [Sta23a]. Haeundae [CLC21]. Hainan [ZJ20]. hake [BDCJ22, GNMGWM21, May22, MKM^{+}20, RQB^{+}20]. Hakodate [TM23a]. halibut [LJ21, SSAP24]. Halophytes [CLCV21]. hammerhead [ALLR^{+}24, Mag23, NC25]. hamper [OCD^{+}23]. Handbook [SSSL25, Tan23b]. Handbooks [SSSL25]. Handl [Cha20b]. handline [WTO24]. Handling [FDVP24, Okş23]. hands [KLS^{+}25]. Hani [YXX^{+}24]. hao [LL24a]. Haor [ABC^{+}25]. happen [NGRM^{+}22]. Harassment [MLE^{+}24, ÖB22, PK20a]. Harbour [JAH^{+}24, OGGC20, GAA^{+}20]. harbours [VGS^{+}20]. hard [DMK24]. Hardback [Rub24, NJW25, SSSL25]. Hardcover [Sta23a]. harmful [CMSN^{+}22, EHLT21, FR23, SS21a, SSV^{+}23, SLMG^{+}21, SLO^{+}24, VSDR^{+}22]. Harnessing [GAN^{+}25, KBB^{+}25, PRM^{+}20]. Harris [Smi20]. Harvest [HM22b, ABC^{+}25, CC21, CCK^{+}21, CLL20, GML^{+}24, JXWL25, Mot20, Mot22, Muk21, SBD^{+}23, Sta25]. harvester [MFG^{+}23, SVP^{+}23]. harvesting [BADO20, EHLT21, SVP^{+}23]. hashtag [WFR23]. hatches [SMF^{+}23]. Haugan [NJW25]. Hauraki [QTBL23]. havens [HMY^{+}23]. Hawai’i [MMAP25, Nal24, PLWO24, SC21, TG25b, ALML23, Cha20a, Lev20]. Hawaiian [BMR20]. Hazard [TSLK23, KWG23, MDLA21]. hazards [IRK^{+}21, TFMT^{+}23]. Head [Cat20, SEG^{+}22]. Health [MK24, Sha22, BDM21, Fen23, GWCD22, LCL20, PGK22, PK20a, Pyć23, SNS23, SHZ^{+}21, SKW^{+}22, TAK^{+}22, TSH22b, VDP^{+}22, WW21a, dO23, YCBW21]. healthier [dO23]. heart [HJ21]. heating [MBG^{+}24b]. heavily [GCV^{+}20]. heavily-used [GCV^{+}20]. hedging [HJPL23]. Hedonic [NTH^{+}21, Cat20]. heel [BD23]. hegemony [SHPW21]. help [Ber24, CBR21, CTC^{+}23, Fen23, MZ24, SLMG^{+}21, dlPC24]. helped [SSB^{+}22]. Helping [Bea22]. herald [GN22]. Heritage [BTN^{+}24, LC23, BEH^{+}21, CSH23, DBH^{+}22, EHÓ25, GWÁV^{+}25, GCL21, GW21, HMM^{+}24, HHL^{+}24, KOCL23, LHW21, LKB^{+}23, Lin24, ML22, PT23, PPvD^{+}24, SRS23, dS21, dJRFVOA24, WM22a, War23]. hermeneutical [SCJ22]. herring [CCP25, Hor23, THGR^{+}20]. heterogeneity [IGO21, JK22, SGK21, SSK23, WNM22]. heterogeneous [NNRS24]. Hexanchus [NGT^{+}23]. hidden [CvHF^{+}24, DBH^{+}22, KLS^{+}25]. hierarchical [DL22, MF23, PBT^{+}20, YYD23]. hierarchy [GCB^{+}23, HF20, JL21c]. High [CMSH25, HCG^{+}25, TCZ21, TBAR^{+}25, Alf20, AFHHD22, BMR20, CMD^{+}20, DAB^{+}25, DBH^{+}22, GR20b, Gul23, Lam21, LX24, MCPS20, MLAS23, MSGAR20, PCA^{+}23, RYL25, Ros22, SSH^{+}23, TSWC23, WvdMG^{+}21, WLY21, vdGD21]. high-ambition [DAB^{+}25]. high-modernist [Alf20]. high-quality [RYL25, WLY21]. High-resolution [CMSH25]. high-seas [AFHHD22, vdGD21]. higher [MM20, WSD21]. Highlighting [Nal24, AN25, VBC^{+}20]. highlights [PSTW21]. highly [BBK^{+}21b, VKN20]. hilsa [INW20, DAAP21]. Hilton [Cat20]. hinder [PGSFRMHH24]. hindering [Bea22]. hindrance [CTC^{+}23]. hippurus [MAG^{+}22]. hired [CSGC22]. historic [BMK^{+}22, SMKS22]. Historical [BdSL25, GMMMRC^{+}25, TBAR^{+}25, BMMF22, GABA^{+}22, GMVM23, Kow24, LSC25, SBGR24, SPHO23, WA24, dPJ21]. historically [MVR21b]. histories [WCN^{+}25]. history [ADB^{+}22, BLH^{+}20, Laj21, LLLF22, MN23, WWC23, vHSSK20]. holdings [BK23, SS22]. holistic [CH25, FOH^{+}20, FMCT^{+}21, JvH24, SEG^{+}22, THF^{+}25, ZZMG^{+}24]. Homarus [MM20]. Homer [Bir20]. Hondo [MTC^{+}23]. Honduras [CAGBM22]. Hong [KCT^{+}24, PSG25, Zhu22]. Hooks [MUR^{+}21a]. Hope [PCHD^{+}25, Pol23, TMDN23]. hopes [Gul23, MGNP25]. horizon [BWV25]. horizons [HRS^{+}24]. Hormigas [HSGNMGC21]. Hornidge [Sta23a]. horse [MKM^{+}20, SGR21]. Horticulture [CTIH21]. host [YNZ22]. hostility [PGP22]. Hot [SHFP21]. Hotspot [CRP^{+}20, CSGCD25, FSF^{+}21, GWÁV^{+}25, MN23, WvdMG^{+}21]. hotspots [AVC24, DFP^{+}25, RXN23]. Hou [LL24a]. hours [BBK21a]. Household [KJS^{+}22, KPN20, NRYG20, TRN^{+}24, TVN23]. households [KDVR20, SPA^{+}21, TSS^{+}24]. housing [CBAPM20]. Huaura [JTR^{+}20]. hub [HLLK22]. Human [BBB25, BDM20, KHMM20, MMG^{+}25, MGF20, SRY24, AVF^{+}23a, Bau24, BDM21, CK23, CLB^{+}20, CWR20, EMN^{+}24, FGF^{+}24, GCF^{+}23, GBAECCRM25, GE23, HJPL23, HACCM22, MIMS^{+}20, NOB22, OAAMS24, PCA^{+}23, RSAP22, SMC^{+}21, SSMG21, SPJB22, TMAS24, TLT^{+}22, TDC21, WPWM^{+}22, Who23, WKBS20, WBH^{+}23, BAP21, PDJ22]. human-animal [HACCM22]. Human-crocodile [KHMM20]. human-nature [PCA^{+}23]. Human-ocean [MMG^{+}25, HJPL23, WPWM^{+}22]. human-rights [OAAMS24]. humanitarian [De 21]. humanity [TSA^{+}22]. Humanizing [CHB23]. humankind [LX24]. Humpback [GCV^{+}20, FMCD20, GHK20, MCW^{+}22b]. humphead [HS23]. Humphreys [De 20]. hundred [Bau20, Liu21]. hundred-year [Bau20]. hunting [EVB21]. hunts [DHU24, DHU25]. hybrid [Cla20]. Hydrate [PY20]. hydro [MC25]. hydro-diplomacy [MC25]. Hydrocarbons [LPO25]. Hydrogen [LDC^{+}24]. hydrothermal [GCG^{+}21]. hydrothermally [VCC^{+}20]. hypoxia [Muk21].

ICCAT [GR25, TWK25]. ice [Dyc24, TBV^{+}23, WCT^{+}20]. ice-breeding [WCT^{+}20]. Iceberg [WD22]. icebreaking [BOW^{+}22]. Iceland [EHÓ25, FCG25, GV22, SDI^{+}20, VJO23]. Icelandic [BDDE20, BADO20, BAD21, EAGV20, Mat20, ÓHCÓ24, Run24]. ICES [LNRM23]. ichthyofauna [AVF23b]. ICJ [Shi23a]. ICZM [BMM^{+}24, OGGC20]. Identification [DA23, NBB^{+}25, TMBM^{+}25, DL22, FDGM24, Kuł22]. identified [ÓHCÓ24]. identify [ES22, HLW^{+}23, MFG^{+}23, dSOW^{+}21]. Identifying [FHBS23, GWAPT21, GKO^{+}25, KVL^{+}23, OCCSC23, PCCR^{+}25, PWL^{+}23a, QCM^{+}20, TGP^{+}24, WFG22a, GCG^{+}21, JHHG23, MG25]. if [PLHK^{+}24]. IFCAs [SCC25]. ignore [JGWM^{+}23]. II [HS23, Sel20]. II-listed [HS23]. II-Listing [Sel20]. III [Fra23]. ilisha [DAAP21]. Illegal [GASLCP20, KgJqY^{+}24, SH23, TSH^{+}22a, ABDRV^{+}23, AGS20, BSLRR24, BJH21, BJ25, Bet25, DGG22, EJC24, GÁSVPS^{+}25, GAD^{+}20, HS24, JGWM^{+}23, KL21a, KAN^{+}23, Ma20, MHT^{+}24, MJMAA21, NBB^{+}23, Pat22, PGM^{+}24, PHP25, SL23, SCD21, SBFHD^{+}23, Tan25, TGM22, VVR^{+}22]. Illuminating [TOL^{+}22]. illusion [SSW^{+}25]. illusions [RCH22]. illustrate [KBD^{+}20]. Imaginaries [ARP25]. Imanga [KDH^{+}21]. imbalance [SL23]. immigration [AHR24]. IMO [CBL24, BH23a, Bau24, BAP21, BÖ24, CR23, DG24b, FFE^{+}24, Kar22b]. Impact [IKB21, Lia25, MMM^{+}22, MCBN22, PBD^{+}23, RAFLRP21, SC22, SLY21, TSS^{+}24, WHL22, YSL^{+}25, AFR25, AP24, AED20, Álv23, AS23a, AH21, BG20a, BA25b, CHEC21, CGMH24, CRIMLS24, CAVJ^{+}20, CWKW20, COYS24, CBH^{+}24b, CBH^{+}24a, CBH^{+}25, CHC^{+}21, CM23, DGG22, DOBE^{+}25, DBS22a, FZ25, FR23, FS22, GMFD25, GMGLV22, GMCMJA^{+}21, GWCD22, Guo20, HJLN24, JL21a, JACD20, KAA^{+}24, Kaw22, Kim24, KSO^{+}25, LDL21, LZX24, LJL22, LMR25, MLVT21, MSSS22, MPL^{+}23, MX21, MYGM25, Muk21, NO25, OYKBA22, PGK22, QSC24, RSK^{+}25, RXX24, RDD^{+}21, SDG^{+}23, SC21, SCM^{+}25, SSP^{+}21a, SRSGMGG21a, Tan22, TSH22b, UOO20, WCC^{+}20, WRL23, Wan23c, XLW24, YFD^{+}21, Zha22, ZT23a, dOPM^{+}20, dlPdlLLC^{+}22, CDC20b]. impacting [NEF23, SMKS22]. Impacts [AIHR^{+}25, AEL^{+}23, CWKE20, DAHS^{+}23, FSRC25a, FSRC25b, GNMGWM21, HHB^{+}21, LKE^{+}22, MLE^{+}24, MCS25b, NAS^{+}22, RKR20, SdCMN^{+}22, VMP^{+}24, ZTZ^{+}21, ABH25a, AELR^{+}20, AR23a, AR23b, BMR20, BRM^{+}21, BSY^{+}20, BRP^{+}23, BASM22, CAM^{+}23, Cha20a, CSGCD25, CLD^{+}24, DPSF25, DDGdS^{+}21, Dun21a, DRS^{+}25, Ero22, EP20, EAGV20, VKL^{+}25, GAD^{+}23, GDB^{+}24, HGQJD24, HCWC21, HPC^{+}23, HZC^{+}24, HBP^{+}21, KAA^{+}24, KL24, Kob22, KHLS23, LSB^{+}21, MYLV24, MMB^{+}23, MCPS20, MN23, Mil22, MMK^{+}22, MZT^{+}23, Nal24, NVP^{+}23, OR21, PCG^{+}23, PV23, RGK^{+}20, RLRS21, SCSR^{+}22, SCSR^{+}23, SOP^{+}22, SBM^{+}25b, SBFHD^{+}23, SLS25, SBT21, SGFM^{+}21, SNK^{+}20, SRH^{+}22, SSP^{+}21b, SRSGMGG21b, dSTHL^{+}24, TSH^{+}22a, TLT^{+}22, TKP23, TGHF^{+}22, UPBB25, VMV24b, VAS^{+}25, WW25, WML25, WABR^{+}22, WNMS^{+}24, WLXY22, ZD20a, ZCS^{+}23, ZNA^{+}23, dSF24, dO23, LEERRM^{+}21]. impairs [RPY^{+}24]. impede [KS22]. impedes [TUBA22]. Impediments [LZ24, WH20, QLZ^{+}24]. imperiled [JKCY25]. impetus [KK22]. implement [CBL24, GTA^{+}20, HC22, MFRQ23, PŚ22, Tra23b]. Implementation [ANS^{+}20, COJS^{+}22, KWY^{+}22, WL24b, AIHR^{+}25, AHT25, AMR21a, ALF^{+}22, BKW^{+}24, BBB^{+}24, CZ24, DG23, DLX21, EdLP^{+}21, Fri24, GBCNCL^{+}21, HWU^{+}23, HZ23, INW20, Isl25, KCCC^{+}21, LZRC23, MLMD21, MMC^{+}22, MB22, NAP^{+}21, OPL20, OWS20, OSV21, PS21, QTD21, RR20, SBP24, SS21b, SGL20, SA25, SH24b, VBH^{+}23, VJO23, VvSvH^{+}22, YHU22, ZSLY20]. Implementing [ABH25b, BJH21, Cra20, DFTK20, LL25, LXH^{+}22, Moh23, TZZ^{+}21, THGR^{+}20, Tho25, ZGR^{+}25a, ZGR^{+}25b, DLJ^{+}20, HJ25, NCE21, RBB^{+}24a, SCRCI24, SAS24]. Implication [NK21, IMA22, MY24]. Implications [AHN25, ABC^{+}24, BYL25, CNB^{+}23, DER^{+}21, FMS^{+}22, KKAO^{+}25, Kra22, LBWS23, MNS^{+}24, MSB^{+}25, Puh23, TG25a, TKP23, WL20, Wat22, AKL^{+}22, BHdMCM^{+}21, BTV^{+}22, CDP22, CZX24, CNG^{+}23, CLCV21, DGI^{+}24, DFP^{+}25, DBH^{+}22, DPG^{+}22, DMS^{+}24, Dua22, EJC24, EB23, FR24, FRVB22, GJY23, Gun21, GHK20, HPS23, HWT24, Has20, HLLK22, Hor25, Hos24, JGWM^{+}23, KL21a, KSS23, LK24a, Liu23, MMW^{+}23, NUD^{+}23, PKC23, PL24a, Pra24, PMFR^{+}22, RLRS21, RB22, SDI^{+}20, STS^{+}23, SKW^{+}22, SGL20, SA25, TKC23, TCC^{+}22, TBV^{+}23, TWK25, THS^{+}24, VSD^{+}23, VMJ25, YBW20, YCBW21, ZZMG^{+}24, ZPX^{+}22, ZSLY20, ZY21, AGN^{+}20, Boy21, GMPT^{+}20, HWMO20, OWCFDC21, RGK^{+}20, TMTC20]. imply [SFN^{+}21]. Import [BGK^{+}22]. importance [AAFK^{+}21, CWLD25, CLB^{+}23, CBD^{+}23, CBL24, DGS^{+}24, MWM^{+}23b, NEH^{+}23, Sep25, SSL25, TN21, WvdMG^{+}21, WMS^{+}25]. importance-performance [SSL25, WMS^{+}25]. Important [PEM^{+}25, COJS^{+}22, MCTZ24, QLU^{+}23, MGPR^{+}25]. imports [AWX21, FGC^{+}24, WDG^{+}20]. imposition [Ros22]. improve [AACdlT23, AAW^{+}23, BGK^{+}22, CCS23, CDC20b, DRCdMCJ21, FLB^{+}21, GAN^{+}25, MPH^{+}23, MT20, MGNP25, OYKBA22, dSRKTC24, SSAAB^{+}24, WKH22, WBB^{+}25, ZCC^{+}24]. Improved [HRC^{+}24, AGK22, BBT^{+}25, BGMCG24, BDC^{+}21, DJC^{+}23, HBP^{+}21, YHU22]. Improvement [GGMCMM^{+}24, SSW^{+}25, AHAB^{+}22, FGF^{+}24, GRSRSN^{+}24, GWAPT21, KOBS21, WS23, ZGY21]. improvements [BCS^{+}24, SSW^{+}25, AO20]. improves [CMSH25, SDG^{+}21, SDG^{+}22]. Improving [AKF^{+}25, BHCL23, CJ22, GPM23, HCT^{+}21, KAA^{+}24, SCBdA20, AFHHD22, BDD^{+}21, GBB^{+}21, LZX24, MPD^{+}23, PTCC24, VHD20]. in-stream [CA20, JB25]. Inaction [JKCY25]. inactive [JG20, VCC^{+}20]. inadvertently [SLL^{+}24]. incentive [BDG^{+}25, JL21b, JZW21, XLW24]. incentive-based [BDG^{+}25]. incentive-compatible [JZW21]. incentives [APGWTM^{+}24, BC21, CGGA20, HWT23, OSV21, PGGA22]. inception [SBG^{+}25]. incidence [HGQJD24]. incident [CT24, SCGC20]. incidental [BUM^{+}24]. incidents [JG23, PKC23, Shi22]. Including [DSQ^{+}22, FTW^{+}23, TCZ21]. inclusion [AFHHD22, ML21, MLB^{+}25, PALB20, RBA23, RMF^{+}21, TEA^{+}22, dS21]. inclusive [AC23, CH25, LHTR23]. inclusivity [JAK^{+}23]. incoherence [EB24]. income [AS23a, FSRC25a, FSRC25b, IOHC22, KTI^{+}24, MDG^{+}23, Seu22, TVN23, UE25, YSL^{+}25]. incomplete [GOH^{+}20]. inconsistency [hBDF22]. Incorporating [Kar22c, KdBK25, ZCC^{+}24]. increase [CSH23, MM22, MWB^{+}21, SKE22]. Increased [Boy21, LJL22, SHP^{+}21]. increases [MJ20, MHN^{+}20]. Increasing [CVA22a, TJZ^{+}21, AH23, Sar23]. increasingly [TJ21a]. Indebting [Kıl24b]. independent [CdLN^{+}21]. Index [EV22, RP23a, Smi22, Smi23b, FMGV20, Sta23b, WD23, COJS^{+}22, LMSLMQ21, PDJ22, RP23b, SGWL21]. India [BPA^{+}22, CDP22, DT23, DMS^{+}24, DDD22, DAAP21, GSA23, GKO^{+}25, KHMM20, KSV^{+}23, MLB^{+}25, NRF21, Pan25b, PRM^{+}20, RBA23, RSK^{+}25, Sol23b, SPA^{+}21, TJS22]. Indian [KWF^{+}21, AIHR^{+}25, Alf20, BLH^{+}20, DXS^{+}21, GSA23, Kar23, MDG^{+}23, McC24, RPK23, SBS21, TGF^{+}25, Yad24]. Indicating [BE20a]. Indications [Gir22]. Indicator [Liu23, COJE24, CD24, TBV^{+}23]. Indicators [BRR^{+}22, BKG^{+}25, CMBB^{+}25, MVR21a, MKB21, RSLR24, WNMS^{+}24, DAA21, DDGdS^{+}21, EAAG21]. Indigenising [RRM^{+}24]. Indigenous [BMM^{+}20, CBCA25, IMB^{+}24, Lat25, LMMT23, MWM^{+}23b, MGNP25, AHB^{+}23, BCR25, BCS^{+}24, JXWL25, MCS^{+}25a, Mol24, SPN^{+}21, WCM^{+}24, BMB^{+}21, DBH^{+}21, DA24, GPI^{+}24, MRD^{+}20, VHS^{+}20, WF25, WJLSH21]. Individual [DP23, HvPPV20b, DBB^{+}22, DMT25, HWT23, HWT24, SSAP24, Zha22]. individuals [SOP^{+}22]. Indo [Can25, CZX24, Gov25, Pan25b, WLS^{+}25, Yad24]. Indo-Pacific [Can25, CZX24, Pan25b]. Indonesia [ASPT24, ABI24, AAKI21, AHAB^{+}22, APF22, AGSI22, AHR24, ANR^{+}22, CLRW22, DDAB^{+}23, DHR23, EAA^{+}22, EABA^{+}22, FKNG25, HLW^{+}20, HYN^{+}23, HF20, HAsR24, HKZ25, JSW^{+}22a, JPM^{+}23, KMG^{+}20, KgJqY^{+}24, LLLF22, MKAF21, MPN^{+}24, NRYG20, PN20, PMM^{+}21, PSA^{+}25, QLU^{+}23, RBEA^{+}22, RYS^{+}21, SMDM20, SIW^{+}21, SWI^{+}22, SLW^{+}23, SWH^{+}23, ST24, SGL^{+}21, SAB^{+}24, SSL25, TUBA22, TSS^{+}24, THT^{+}23, TRN^{+}24, TEA^{+}22, UE25, VKS23, WFK^{+}24, WMRB22, WMS^{+}25, YC21b, Zou23, vLS24]. Indonesian [ARP25, AFS^{+}23, DSWB20, Dun21b, LWSS21, LTW23, Lin20, LSOP24, PFWM23, RMU^{+}22, SHS^{+}25, VKS23, WTO24]. induced [FAC^{+}24]. industrial [AE21, CQ20, CCKK22, KLSG21, KLS^{+}25, LJL22, NSPRSS20, RHO24, Tar25, Tra23a, TFJR25, ZW21]. industrialization [AML^{+}21]. industries [LXSX21, TM23a, TFF22, VMV24b, YZH22].

Industry [KCB24, MTS^{+}24, TYS22, AOO^{+}23a, ABC^{+}24, Ber25, BDR20, BMH^{+}22, CBC24, Cha22, CQ20, CMS^{+}25, CGGA20, FGPVPP^{+}24, Fla21, FR23, Fro21, FBC^{+}23, GEGRdL24, Gar24, GG20, HPS23, HPG20, HJLN24, HHB^{+}21, HPH22, HKKY22, JFB^{+}24, JLY23, JD20, KSÖB25, KCCC^{+}21, KSS23, Kiz24, KF23, Koe22, LWSS21, LTW23, Lee24, LLHG23, LJ22, LJL22, LC22, LFPBD^{+}20, LKE^{+}22, MGF20, MMO^{+}24, NSN22, PAK22, PPK23, RAMEG24, RU20, Ren21a, Ren21b, RDM^{+}20, RAB22, SDI^{+}20, SLY21, SZ20, SLS^{+}20, SLS^{+}21, Ton22a, VJÓK25, WOR^{+}25, WGBS25, WN25, WL22b, XHY22, ZCCQ22b, ZZW^{+}21, ZWC^{+}23, ZSLY20, ZCS^{+}23, vLPBK21, SDI^{+}20, LJL22]. Industry-initiated [TYS22]. inequalities [IOHC22, SAC25]. inequality [Mag23, MLE^{+}24, NC25, TVN23, UE25]. inequities [MTC^{+}23]. inertia [SGR21]. infectious [Bov22]. infer [CAVJ^{+}20]. infinite [PCA^{+}23]. inflation [Tar25]. Influence [KTI^{+}24, WZZC21, ZW21, AB20, BAP21, BMR23, CFA^{+}24, Fri24, HT23, OAdJC^{+}25, PESO24, QL24, QF20, WDG^{+}20, WWD23, YXDL24]. influenced [YNZ22]. Influences [Lee24, YSEK22]. influencing [AAKI21, CDC24, Dam24, HMFH20, JZW21, KWMM20b, MLMD21, Pra24, PPK23, SYA25, XHY22, XXLW24]. inform [BBB23, BEM^{+}20, BDL^{+}24, MMNC22, NSM25, Nal24, PBS^{+}23, SMLRB^{+}21, SSOP^{+}21, SMW^{+}21, SSAAB^{+}24, TCT24, VCC^{+}20, WD20, WCN^{+}25, WNBRSG22, YJG24, ZCG^{+}23]. Informal [MJ23, Pau21, GBSC22, IO23, PL20b, vLS24]. informalisation [YS23]. informality [GWAPRPRU22]. Information [dSOW^{+}21, ATF^{+}21, CMR21, DQJ23, DSC^{+}23, FLC^{+}23, GSB20, LR20, NM23, RPY^{+}24, SDG^{+}21, SDG^{+}22, WM21]. informatization [JL21a]. Informed [AAA^{+}21, GAD^{+}23]. informing [WtDD^{+}21]. informs [BL23a, MDG^{+}23]. infrastructure [AH23, BL23b, BS22, CDRS22, HLLK22, RKS^{+}23, WN25, ZCG^{+}23]. infrastructures [Alf20]. infringement [Gar22]. ingredients [Lev20]. Ingvild [Niu22]. Initial [BDL^{+}24, RDD^{+}21, MMNC22, PWL^{+}23b, Tan24, ZTZ^{+}21]. initiated [TYS22]. Initiative [BDTD25, LMC^{+}23b, TSLK23, Wan23c, Ber23, DFTK20, TBPC20]. initiatives [MGDJ22, RSN25, VVR^{+}22]. injuries [Dev22]. injury [VDOE^{+}24]. injustice [WWCS21]. inland [OOSG21, PPK23, Zha22]. Innovating [GRvdM24]. Innovation [AOO^{+}23a, PVH^{+}21, Per24, BBGHR23, FWL^{+}22, GLS^{+}21, GVR20, Kow24, NWZ^{+}21, PSD23, Pen22, RJ21, Sha20, SLY21, SKDK23, Ton24, VMD^{+}24, VGA25, WSD21, ZW21]. innovations [HCW^{+}23, LMMT23, SDG^{+}23, SR22]. Innovative [ABVPM20, TT22, Wan24c, GMMMRC^{+}25]. inorganic [ADI^{+}25, TAP^{+}23]. Input [BSSL^{+}20, dlFGFFR^{+}20, BG20a, MGK^{+}25, SG21, SRSGMGG21b, TM23a, WB24, ZCCQ22a, ZT21, ZZD^{+}23]. Input-Output [dlFGFFR^{+}20, BG20a, SRSGMGG21b, ZT21, ZZD^{+}23]. inquiries [CVA22a]. insecurity [CMWT24, NRYG20]. inshore [BCJ21, CWMM^{+}21, SKS21, SSAP24, SCC25]. insight [OC20]. Insights [ADI^{+}25, AAA^{+}24, BOOO23, BSLRR24, Bir20, DT23, DBBL21, EHÓ25, FZL25, FGF^{+}24, FRLMW25, FRDK24, MKM23, MZ22, MPH^{+}23, OCD^{+}23, PMC24, UE25, WNBHG24, ABC^{+}24, CGHV20, GPG^{+}21, GCM^{+}23, OWC^{+}24, OGS24, RDD^{+}21, SYA25, SLY25, SKDK23, Tho22, VWJMAJ20, WKK^{+}25, BCBS22, BBS^{+}21, CBO^{+}21, CSV^{+}24, CBL24, GRA^{+}22, HMG^{+}21b, IRK^{+}21, JZW21, KK22, KDVR20, NCF25, SGR21, SL20, SSB^{+}22, SSMG21]. Inspection [XWL^{+}20, HFY^{+}24, Öİ24, Şan20]. inspections [Kar22a, Liu25]. inspiration [Ada22, LMT21]. instability [BLM24]. instigating [FGK^{+}20]. Institute [HH21]. Institution [BV21]. Institution-based [BV21]. Institutional [AHM22, BDHI21, Cla20, GLMBMP25, Kra22, SGR21, SG20a, SA25, XL25, YZH22, APGWTM^{+}24, AGSI22, BPA^{+}22, BSM23, BUR^{+}22, CDG^{+}21, Cra20, DGH25, DP23, GLADGGVL23, Joh25, RRR^{+}23, SL23, SVF20, SOS20, TGP^{+}24, WWD24, WH20]. Institutionalising [GNJR24]. Institutionalization [ONN^{+}20]. institutions [Ber22, GBSC22, HZ23, Kat20, Mab24, PL20b]. instrument [CWR20, ESD24, Gar20, LR20, Lou22, MFH^{+}20, Pri22]. Instruments [PŚ22, Aqu23, BSSL^{+}20, CBL24, Kar22b, PMFR^{+}22, RLS20, VHCJdYC^{+}24]. insurance [JF20, LK24a, Set23, YZL24, ZSZ20]. insurers [LK24a]. Intangible [HMM^{+}24]. integral [BRR^{+}22]. Integrated [HIP^{+}21, KSB^{+}23a, SCRCI24, AHM22, BTN^{+}24, BBGG22, EdLP^{+}21, VKL^{+}25, KABT21, LLCH21, LZL23, PMR23, SGFM^{+}21, TCS^{+}20, THF^{+}25, TFMT^{+}23, ZCA^{+}23, dSdLSGB22, CAGBM22]. Integrating [APV^{+}20, BEH^{+}21, CHH^{+}25, FGF^{+}24, LKB^{+}23, MM23, PHF20, PKS20, San25, SSF23, VJÓK25, CJH24, HvT22, MF25, SRR^{+}21, MLVT21, PFG^{+}21]. Integration [AS23b, JKV^{+}22, Koj23, Kon20, BC20a, GGP^{+}22, IOHC22, QYX21, RLS20, SRS^{+}25, VHSV20, WOR^{+}25, XXLW24]. integrity [FTW^{+}23, FOKM23, PSTW21]. intellectual [VAS^{+}25]. Intelligent [ZLWY22, THP22a, TC23]. intensity [SS21a, TBV^{+}23, BAD21]. intensive [KMM21]. intention [DLYM22, Lin25]. intentions [GPM20, MLM23]. Inter [Har22, HSdT21, PTS^{+}25, RKS^{+}24, SRY24]. inter-organizational [HSdT21, PTS^{+}25, RKS^{+}24]. Inter-state [Har22, SRY24]. interacting [POM24]. Interaction [ZCC^{+}24, Ber24, Fri24, Kar23, TDC21, YWYL22]. Interactions [RTMCVMDE22, TH22, BPA^{+}22, BL22, BMK^{+}25, ESD24, GBAECCRM25, HTP^{+}23, HACCM22, OPL20, RHST21, SVHV20, SMC^{+}21, SRS^{+}25, TG25a, WPWM^{+}22, BMM^{+}24, LPMS24]. interactive [CRP^{+}20, LJ23]. interacts [BOW^{+}22]. Interconnection [AB20]. interdisciplinary [SRH^{+}22]. interest [BAP21, Mar21a]. interests [Has23, SC24, SC25]. interface [CMS^{+}21, GSA23, PKH22]. interfere [CMOJ^{+}24]. intergovernmental [DMNT20, MDJ^{+}22, TMDN23]. interindustry [TTL^{+}25]. intermedius [GPH^{+}21]. International [CCX23, Cha20b, HCW^{+}23, Har22, HH21, KMT^{+}21, LNRM23, QLZ^{+}24, Shi23b, SSSL25, Tan23b, TCZ21, WL20, Wan23b, WQ25, War20, ZY23, Ber24, BBK^{+}21b, CZW20, Cha22, CMWN^{+}23, Che21b, CBB^{+}24, CWR20, Dom23, DC23, Gar22, GY23, Gra21, Gue23, HCSBM^{+}23, Hor25, Joh25, Kle21, KL22, KH24, Koj23, MAG^{+}22, NC25, Pri22, QSC24, Sam22, SSH^{+}23, Shi23a, Shi24a, SX25, SSK23, SK24, SLH22, TKC23, Ton22a, TC23, Urb22, VMV^{+}24a, WRL23, Wan23a, Wan24b, WZL25, WdSRY23, WL24b, XL25, ZT23b, vdGD21, AR22, BDC23, BHKJT23, BRKS25, BRM20, CM25, GSM20, KA24, Kuł22, Mar21a, MK24, PZPB21, RJAT23, Sin21]. Internationalization [LZ24]. internationally [BLM24]. Internet [AAW^{+}23, PGK22, TJZ^{+}21]. interoperable [LHV^{+}20, WBB^{+}25]. Interplay [Per24, Hor23, PL20b, SLB21]. Interplays [WS24]. Interpretation [PL23, WL20]. interpretations [BDM20]. intersection [BJ25, Bet25, GPA^{+}22, vdGD21]. Interstate [PGP22]. Intertidal [PJN24, ÁTGCE^{+}25]. intertwined [Tru23]. intervention [Fen23]. Interventionist [LTW23]. interventions [ALML23, BOOO23, FMWW21, GHW^{+}22, MHC^{+}25, MLB^{+}25, MS21, PBG23, SLB21, TCT24, VWJMAJ20]. interviews [GLB^{+}24]. intra [Har22]. intra-state [Har22]. introduced [SPN^{+}21]. Introducing [Liu22, TCT21, WCC^{+}20, Åm21, Jen23, NHA23, RAB22]. Introduction [Ehl21a, GAD^{+}23, PZPB21, SNS23, VFOA25, GRvdM24, EAAG21]. introductions [MCS^{+}25a]. intrusion [RKR20]. Inuit [BOW^{+}22, KDH^{+}21, SCF^{+}22]. invasions [SMF^{+}23]. Invasive [FCG25, MMB^{+}23, AC21, ABC^{+}24, BLM^{+}25, CAHR^{+}24, GTTM21, OGS25, SSDQ20, SMC^{+}21]. invertebrate [CCK^{+}21]. invertebrates [OVSVPP^{+}22]. Investigating [DSD^{+}24, GTTM21, TBV^{+}23, TTL^{+}25, KA24, SS22]. investigation [FFE^{+}24, Han20, VDN^{+}24]. Investing [NO25]. investment [AED20, FGPPPV21]. investments [BXV20, SDI^{+}20]. invisible [IO23]. involved [MMK^{+}22]. involvement [LEVET21, STIS20, vRB24]. Involving [DSL20, HS20, DBBL21]. iObserver [VAMR^{+}20]. IOCD [Bov22]. IPCC [CTH^{+}21]. Iphigenia [dJdMCL^{+}24]. IPMDA [CZX24]. Iran [AB20, PAK22]. Iranian [HGG^{+}22a]. Ireland [HFD^{+}20, AH23, AHM22, CSV^{+}24, CCW21, CM23, MMNC22, OCD^{+}20, OGGC20, dlÁPAFGO20, vRB24]. Irish [CPR21, OC20]. Irregular [MSSS22]. ISA [MB21]. ISAV [BVvKL21]. ISBN [Cha20b, HH21, LPO25, LL24a, Lv20, Ngu23, Niu22, NJW25, Rub24, SSSL25, Sta23a, Ste25, Tan23a, Tan23b, WQ25]. ISIS [PMT^{+}20]. ISIS-Fish [PMT^{+}20]. Isla [MTC^{+}23, GASLCP20, dPJ21]. Island [BAVV24, BMH^{+}22, Cat20, CTIH21, DRY22, KWG23, MAMCHP^{+}20, Moh23, SMHPMO^{+}21, SOP^{+}22, Tho25, WQ25, WMS^{+}25, AHM22, AJDA24, AL20, CDC24, COSS21, CFS^{+}24, DLX24, KMS20, NRYG20, NSN22, NNR^{+}21, PMR23, RMU^{+}22, SMKS22, TFH^{+}23, ASM^{+}21b, IGS^{+}23, Kak20, KK22, Los23, Los24, MZ24, OAAMS24, SCB25, WCN^{+}25, dJdMCL^{+}24]. island-based [RMU^{+}22]. Islands [LTJ^{+}23, PN20, SWW^{+}21, TJS22, ADI^{+}25, APF22, CDC24, DČFW23, OC24, THP22a, THT^{+}23, AGSI22, AM22, BDMW22, CTIH21, Chr20, DAA21, Dun21a, EB24, GBFMT23, GMGLV22, GMRSP23, GMCMJA^{+}21, MBORGN24, Mrd23, RGH^{+}22, SW22, Tod22b, Tra23a, YTBM^{+}22, ZMR20]. Islas [HSGNMGC21]. Israel [PC20, YNZ22, ZABI25]. Issue [EB23, VFOA25, BR22, CDC^{+}25, Ehl21a, GAD^{+}23, PZPB21, SNS23]. issued [BRM20]. Issues [BKW^{+}24, AR22, CHB^{+}20, CLL22a, GTMP20, HCW^{+}23, MMNC22, Min24, ORdS25, Pan25a, PLHK^{+}24, PGA^{+}20, Tan25, WZ25, Lv20]. Isurus [DSL20, Kai21]. Italian [ABC^{+}24, CSRG25, GBM^{+}25, MMF^{+}25, VMD^{+}24]. Italy [CDC^{+}25, Mar21a, Mor20, RBB^{+}24a, VMC^{+}21]. ITLOS [QSC24]. ITQ [GV22]. ITQ-managed [GV22]. IUCN [ABM21, Kai21, PNS^{+}23]. IUU [AHT25, BOBMG25, Faj22, GÁPSAP^{+}23, GY23, HM22c, HMY^{+}23, JS23, KL21a, KA21, MFAH25, Oan22, PHP25, Tei23, YW24]. IWC [Gal22, SOS25]. ix [Smi23b].

Jaeckel [SSSL25, Tan23b]. Jakobsen [Niu22]. Jamaica [LZRC23]. James [Tan23a]. Jane [NJW25]. Janeiro [SBP^{+}25, SdACdSLdAS24]. January [Ano20p, Ano21q, Ano22q, Ano23q, Ano24q, Ano25o]. Japan [Kim21, AKFY22, BI21, CL24, FRDK24, FL24, FMS^{+}22, HWU^{+}23, HWT24, HT23, HWMO20, Hor25, IMA22, JY20, KS23, Kay25, Kim20, Kob22, KWY^{+}22, Lia25, LSC25, MM21b, OKH^{+}24, OGM24, OI23, PL24a, QL24, SII22, SOS25, Sas25a, TYS24, TM23a, Tan25, TTA^{+}20, Tru23, TG25b, TYS22, WM21, WNM22, XTY22, YXDL24, ZTA23]. Japanese [GM25b, HWT23, HWT24, KSÖB25, KYMW21, KB20, KN21, ONA21, SRH^{+}22, TZM^{+}22, TSWC23, WS21]. Jariego [TGM22, SCD21]. Java [WFK^{+}24]. JCG [Zha25]. Jelly [MSZ^{+}23]. Jelly-FAD [MSZ^{+}23]. Jellyfish [KSV^{+}23, BCMCM21]. Jersey [SLS25]. Jiangsu [XWH^{+}22]. Jiaozhou [SL20, WLX21]. Job [PSPR20, ZWXL21]. Johansen [Niu22]. John [De 20]. Joint [Lin24, CHB^{+}20, Kim20, CHB^{+}20]. Jolla [KL21b]. Jor [AAKI21]. Joseph [Lv20]. joy [JBGC25]. Jr. [Fan24]. judgement [Pen22]. Judicial [CCX23, Li23b]. July [Ano20q, Ano21r, Ano22r, Ano23r, Ano24r, Ano25p]. jumbo [CEMRMB^{+}24, GWAPRPRU22]. June [Mar21a, Ano20r, Ano21s, Ano22s, Ano23s, Ano24s, Ano25q]. juridical [Gul23]. jurisdiction [BLL22, CWR20, CM25, DMNT20, DBZ24, FMS^{+}22, GNJR24, Gar20, HJ21, HM22a, HGL^{+}20, MDJ^{+}22, MH24, MCR22, OKR^{+}20, Pat22, Per24, Pri22, TD21, TMDN23, VHD20, VBC^{+}20, WF25, WGJ^{+}21, YG20, YW24, ZJ22, dO23, Has23, MTN23, Teo23, TM23b, TCD^{+}20]. jurisdictional [dO23]. jurisdictions [SHP^{+}21, YT20]. jus [Kuł22]. Just [ARP25, AC23, GABA^{+}23, KL21a, Lot23, SS21b, WLW^{+}22, dlPC24, SCF^{+}22]. Justice [Pen22, BC20a, BCN23, BBWC21, BAF^{+}23, Dau23, DUYEA21, PCHD^{+}25, YCF24, GFL^{+}21, SCJ22].

Ka’an [MMSB^{+}20]. Kadavu [RGK^{+}20]. Kao [LK24b]. Kaohsiung [TCT21]. KAP [KCCC^{+}21]. Kate [AFHM23]. kayaking [AH23].

Kazakhstan [AB20]. keeping [vLPBK21]. Kei [THT^{+}23]. kelp [ÁTGCE^{+}25, CTC^{+}23, GRPRVG23, GTTM21, HGG^{+}22b, HKKY22, QTBL23, TDC21]. Kenya [AOO^{+}23b, KWS20, KKAO^{+}25, KWMM20a, KWMM20b, LSB^{+}21, OWS20, OKNB22, OBNK24, OGH24, OKFN24, OOM^{+}25]. Kenyan [CMWT24, KOW^{+}25]. Kepulauan [HF20]. Kerala [DT23, NRF21, RBA23]. Kerguelen [TH22]. Key [HCW^{+}23, KKDK20, KSS23, MF20, MNHS25, NEF23, PDD^{+}25, PGA^{+}20, AHENNA^{+}24, AFHHD22, ABC^{+}24, BN20, CCCC21, GCP^{+}20, KCT^{+}24, KJS^{+}22, LSD^{+}22, MN23, SdACdSLdAS24, WFG22a, ZSLY20, ZCC^{+}24, Sin21]. Kil [Tan23a]. Killer [PTS^{+}25, Dam24, FMCD20, Shi22]. Killing [HM22c]. Kin [DCC^{+}20]. king [BRR^{+}22, DD22, VKN20, BRR^{+}22]. Kingdom [Cro21, FS21, KSB^{+}23b, SSA^{+}23, SPH^{+}20, VSD^{+}23]. Kiribati [Mal21]. Klang [RGH^{+}22]. Klaten [vLS24]. Klein [SSSL25, Tan23b]. Knowledge [CDC^{+}25, DVRK25, HBP^{+}21, MN23, MWM^{+}23b, NJW25, SBG^{+}25, AAKI21, Alm21a, AELR^{+}20, AHN^{+}23, ALLR^{+}24, BC20a, BB21, BMK^{+}22, BVT^{+}21, CPV^{+}21, CSV25, CAVJ^{+}20, CSVC23, DMd^{+}20, DSQ^{+}22, DG24b, DA23, FMTA^{+}25, FBC^{+}23, GNJR24, GSAF21, GCP^{+}20, GSM20, GAN^{+}25, HTP^{+}23, JK22, dSJNS20, KCCC^{+}21, KCH^{+}25, Kar22c, KDH^{+}21, LPMS24, Lat25, MZVC25, MMB^{+}23, MFG^{+}23, MWM^{+}23a, MMS24, CSdMR^{+}24, MFH^{+}20, NGT^{+}23, OCZ^{+}22, OOM^{+}25, PPvD^{+}24, PMM^{+}21, SBFHD^{+}23, SCCW21, SVP^{+}23, Sta23a, WF25, ZYTT21a, ZYTT21b, dJdMCL^{+}24, BÅ25a]. knowledge-base [GSM20]. knowledge-based [CPV^{+}21]. Known [HSGS^{+}23]. Komodo [SSL25]. Kong [KCT^{+}24, PSG25, Zhu22]. König [WQ25]. Korea [Kim21, CKY20, CLC21, CKY22, CJLJ24, HS22, JD20, KPN20, KNY20, Kim20, KNY22, Kim23, KSS23, KLK23, KGP^{+}22, LrS22, Lee24, PKR^{+}22, PKC23, PL24a, YK20]. Korea-Radionuclides [PL24a]. Korean [hC24, KLS22, KHH^{+}23, YXDL24]. Kraska [Tan23a]. krill [CMC22, Tra23b, TFJR25]. Kristoffer [Cha20b]. Kuang [LL24a]. Kuang-hao [LL24a]. Kum [LK24b]. Kuwait [AGS20].

label [AH21, BdLAdC^{+}20, CMR21]. labeling [GVL22, Kıl24a, WEF^{+}21]. labelling [CMNH21, JAK^{+}23, SFN^{+}21, WWC23]. Labels [PMK^{+}21]. labor [LXSX21, dSTHL^{+}24, WKBS20]. laboratory [HNB^{+}25]. laborers [TFP25]. Labour [CK23, KBC25, KCB24, PGSS22, SMCW22, SRY24, TOL^{+}22, WS23, YS21, YL21, YCF24, KL21a]. Labrador [NGRM^{+}22, MGNP25]. Lacepède [dSJNS20]. lack [TMDN23, VKN20]. lagoon [AC21, BASM22, Han21, RB22]. Lagos [BRR^{+}22, CLU^{+}24]. laissez [KK22, YL21]. laissez-faire [KK22, YL21]. Lake [WCN^{+}25]. Lancaster [KDH^{+}21]. Land [SVHV20, Bre25, CACJR^{+}21, Gug24, JKV^{+}22, LSH21, MX21, OPL20, PMAVM20, SVF20, SL20, SCF^{+}22, SRS^{+}25, BMM^{+}24, LPMS24]. land- [CACJR^{+}21, SVF20]. land-based [Gug24]. Land-sea [SVHV20, OPL20, SRS^{+}25]. Land-Sea-Interactions [BMM^{+}24]. landing [CLB^{+}20, FGPPGG23, GMPT^{+}20, MZVC25, OSV21, PMAVM20, PV23, SSZ^{+}24, SOLN21, YU20]. landing-ban [CLB^{+}20]. landings [AFMdSdTdCC22, IGO21, SOLN20, TDP20]. Landsat [XWH^{+}22]. landscape [DBH^{+}22, HHL^{+}24, KDHL24, Pan25b, Ton24]. landscapes [CSV^{+}24, HM21, PFM20]. lanes [VKS23]. language [ASKI20]. Lanka [AEL^{+}23, CLB^{+}20, CLB^{+}23, GRS^{+}23, JAH^{+}24, Kul23, NEH^{+}23, PDJ22]. Lankan [Pra20]. Lao [KJE21b]. Laoang [PCA21]. Large [ASGZ22, SLP^{+}22, WWG^{+}23, ABM21, BBB25, DLJ^{+}20, HYN^{+}23, HLW^{+}23, HSH^{+}21, HLO^{+}21, KNY20, KLK23, MZT^{+}23, PBG23, Sol23a, TBAR^{+}25, WD23, dO23, WWG^{+}25]. Large-scale [ASGZ22, BBB25, DLJ^{+}20, HYN^{+}23, HLW^{+}23, KNY20, MZT^{+}23, PBG23]. largest [CLP21, EMH^{+}23, HS24, NBB^{+}23, SBG^{+}24, SBG^{+}25, VLG^{+}24]. last [KGW20, LCMBGP22, VMC^{+}21, WEF^{+}21]. lasting [NSPRSS20]. late [GMVM23]. later [Ege23]. latest [LC23]. Latin [AFR25, AE23, GGMCMM^{+}24, MBC^{+}24, San25]. latissima [HKKY22]. Latvia [RCM24]. launches [LP21]. Law [Kuł22, LK24b, MY24, Ngu23, Niu22, Set23, Smi23b, SSSL25, Tan23a, TKC23, WQ25, Zha25, Álv23, Ber24, BMR23, CLS20, CS20, Cha21, CZ24, DSB^{+}25, DC23, Ege23, Gue23, HZC^{+}24, Hon20, Hor25, KS22, KDHL24, Koj23, LC20a, Li23a, Lia25, Lin24, NH20, Pan25a, QSC24, SK20b, Shi23a, Shi24a, Shi24b, SCZ^{+}24, SH25, Tan25, TZC23, Tse21, Wan23b, Wan23a, WZX24, Wan24c, Wan24b, WM22a, War20, WSDB20, YW24, ZT23b, Gul23, JH24, Koj23, Lia25, LH24, Mar21a, TH22, YCL25, ZY23, SSSL25, Tan23b]. law-making [Wan23b, Wan23a]. lawmakers [ADWP24]. Lawrence [TMBM^{+}25]. Laws [RS24, CMNH21, CSHO^{+}20, DZZC24, DRS^{+}25, Kle21, KGP^{+}22, LHW21, PK20a, Wil20, WEF^{+}21]. Lawscape [O’D21]. Lawships [GC22]. laying [KS22]. layout [LLHG23]. lead [HC25, MMK^{+}23, YHX24]. leader [HWT23]. leaders [SSB^{+}22]. leadership [GFD21, HGG22c, IMB^{+}24, OAdJC^{+}25]. leading [BA25b, DHR23, RE21]. learn [RMBR22]. learned [ADR^{+}21, BDMW22, BBJ^{+}22, BPJ^{+}24, CHC^{+}21, CNG^{+}23, DSL^{+}21, GLAVL21, GABA^{+}23, MF22, MCW^{+}22a, MCR22, NJBM^{+}23, SOP^{+}22, Squ21, TRL^{+}25, VWFW22, VVR^{+}22, ZGR^{+}25a, ZGR^{+}25b]. Learning [BMH^{+}22, KMTvT25, VBH^{+}23, CRJ22, OVA22, RB20, TRL^{+}25, TCT21, YHU22, ZYTT21a, JF20]. learning-experiences [RB20]. learnt [ALF^{+}22, BAVV24, KSB^{+}23b, MZT^{+}23, PKB^{+}21, UISM^{+}23]. leasing [EP20]. Leaving [RL24, TCF^{+}25]. led [DBS^{+}22b, HWT24, Lin24, MGDJ22, RBA23, VLG^{+}24, WdSRY23]. left [DMNT20]. Legal [Bov22, Cha20b, CWKW20, CLL22a, FSF^{+}21, LPO25, LCL24, LF25, MSSS22, Pan25a, Pie21, TN21, WL24b, dS22a, ATM^{+}24, BN20, BUR^{+}22, CC23, CZW20, CDY24, CCP25, CDRS22, CMD^{+}20, Fri24, Gal21, Gar22, GMVM23, JMB^{+}20, Kar22b, Kar23, KBC25, KL22, Kuł22, LPMG^{+}24, LC20a, Ma20, Min24, NCE21, OT23, PY20, PL23, RFPS20, SRvNN21, TFJR25, TK23, Urb22, Wan24a, WZ25, XZ23, Zha25, Cra20]. legalism [NW20]. legally [CWR20, LR20, MB22, Pri22]. Legislation [Lin24, TCZ21, ADWP24, GBC^{+}20, Kak20, KAN^{+}23, ÓHCÓ24, PMK^{+}21, Qaf24, SS23, Wil20, YT20, YD22]. Legislative [Li23b, DPB^{+}23, FFD^{+}23, KF23, Lov23, WCM^{+}24]. legitimacy [SvLGK24]. legitimate [Dau23]. legitimizing [RHG^{+}25]. leisure [BAJR24]. length [ASM21a]. length-based [ASM21a]. lens [DDGdS^{+}21, FFE20, SK20a]. lenses [PDJ22]. Leone [OOGS20a, OOGS20b, OOSG21]. Lepidorhombus [APV^{+}20]. less [SMC^{+}21, SMCW22]. Lessonia [ÁTGCE^{+}25, CBO^{+}21]. Lessons [ABDRV^{+}23, ALF^{+}22, BDMW22, BBB^{+}24, BPJ^{+}24, CNG^{+}23, DSL^{+}21, FOKM23, GABA^{+}23, HNMSG21, KMTvT25, KSB^{+}23b, MG25, MF22, MCW^{+}22a, MCR22, MAA^{+}25, MZT^{+}23, NJBM^{+}23, QTBL23, RCM^{+}22, SOP^{+}22, TRL^{+}25, THF^{+}25, UISM^{+}23, VVR^{+}22, WBB^{+}25, WCN^{+}25, ZX24, BAVV24, BBJ^{+}22, CMGF20, DPP^{+}20, FSKJ23, KHdSM^{+}22, Liu24, TBCB25, VWFW22, VBH^{+}23, ZGR^{+}25a, ZGR^{+}25b, ZT23a, ADR^{+}21, BC20a, CHC^{+}21, GLAVL21, KGW20, Kle21, OA20, PKB^{+}21, PGSFRMHH24, Squ21, SVKP21, SB20]. Lesvos [Sak21]. Let [CMWT24]. lethal [AMLC^{+}23, HSH^{+}21, SON21]. level [BADO20, CHEC21, Cha24a, CDC24, DSC^{+}23, HJLN24, IAB^{+}20, JY20, JLY23, Kim23, Mal21, Mba22, MCKZ21, NSPRSS20, Ora23, QF20, RXX24, SPG20, SPMN23]. levelled [HvdBH23]. levels [AFR25, AFMdSdTdCC22, AdSMS21, CLU^{+}24, GBDC24, HLCM21, MMW^{+}23, ZWJvL20]. leverage [GKO^{+}25]. Leveraging [Ber25, TMAS24, OGH24, DKW20]. lewini [ALLR^{+}24]. liability [FY24, Fei20, MH23, TN21]. Liberia [AAA^{+}24, JKKT20, LA23, NAS^{+}22, WM22b]. lice [GVR20]. licence [BHK^{+}25, KFM^{+}20, LMMT23, DN20]. license [Ber23, SS22]. licenses [OOGA22]. licensing [Her22]. lien [WL24a]. Life [GCC^{+}21a, ADB^{+}22, BBBS22, EAA^{+}21, Fro21, Gul20, LLLF22, MN23, MS21, WWC23, AP23]. light [Mar21a, NH20, Pri22, Sar23, TLT^{+}22]. Like [Ree20]. likely [Kim24, MHT^{+}24, SPN^{+}21, VPRR^{+}21]. Lima [VRBBM^{+}24]. limit [GHK20, TYS22]. Limited [NC25, CNB^{+}23, CDB20, Fuj21, GCS21, LLLF22, PWL^{+}23b]. limiting [lTS25]. limits [Aqu23, AMR21b, AS23a, Bau20, HNA25, MBDG23, Sam22, SRY24]. limu [MMAP25]. Lindi [YMG23]. line [DAA21, LCAX21, MVR21b]. linear [MEG^{+}21]. liner [BHEE22]. lines [KKDK20]. link [CCK^{+}21, HHH^{+}23, JBGC25, RMBR22, TRN^{+}24]. linkages [Chr20, CSL21, EB23]. linked [CT24, DVRK25, SRSGMGG21b]. Linking [CDS^{+}25, GABA^{+}22, dOJCSLdSB21, MWM^{+}23a, MFM21, OMC^{+}22, SCA^{+}22, WGR^{+}22, WWS^{+}25, dlPC24, SC24, SC25]. Lion [VCC^{+}21, EECE^{+}24]. lionfish [YTBM^{+}22]. lions [PBH^{+}23, dOPM^{+}20]. Lisbon [SCA^{+}22]. listed [HS23]. listening [WPvO24]. listing [Kai21, Sel20]. lists [CMNH21]. literacy [CJLJ24, CFS^{+}24, CSH23, LZL23, MF25, MMG^{+}25, PDS24, TSWC23]. literature [ABM21, CBH^{+}24b, CBH^{+}24a, CBH^{+}25, DGMF^{+}20, GLSVMB25, HLLK22, Jen23, KMSN25, LMT21, LRC24, ÖB22, SM23, SCM^{+}25]. Litter [MB22, AAA^{+}24, AVC24, ADR^{+}21, Che21a, DPB^{+}23, HPC^{+}23, KLK23, MTC^{+}23, Mor20, ONN^{+}20, PKB^{+}23, PBG^{+}22, WMS^{+}25]. littleneck [BRR^{+}22]. Liu [Sos21a]. live [CRJ22, KCT^{+}24, MFAH25, SAS^{+}25b, WGBS25]. lived [PHD^{+}24, UISM^{+}23]. livelihood [HP23, IAS25, LRN^{+}22, Pha20, WLW^{+}22]. livelihoods [BOW^{+}22, CTH^{+}23, FRLMW25, Han21, HBP^{+}21, KOBM24, LFLF22, MF22, Mba22, NAS^{+}22, OEA25, PTCC24, SF22, SGL^{+}21, SNK^{+}20, SAB^{+}24, YMG23, dlPdlLLC^{+}22, AFHM23]. livestock [TOHK20]. Living [BAAH24b, GB21, Sos21a, Liu20, PCA^{+}23, SCCW21, ZZMG^{+}24]. LMMA [LTJ^{+}21]. LNG [HPS23, MF23]. Lobos [dPJ21]. Lobster [KKO^{+}22, Bot20, BKG^{+}25, CMS^{+}25, DMS^{+}24, Doe21, EJC24, FRLMW25, Gor21, HS20, KdBK25, MM20, OAM20, OAM21, PLOG24, SOP^{+}22, TPG^{+}23]. Local [ADI^{+}25, Alm21a, ALLR^{+}24, FTM^{+}22, FSZ25, FS22, GPI^{+}24, MZVC25, MMB^{+}23, MPSF20, OCZ^{+}22, OOM^{+}25, WJB^{+}20, YML^{+}21, YD22, ALML23, AELR^{+}20, BC20a, BCR25, BMK^{+}22, CAVJ^{+}20, CS21, CSVC23, DB20, Dun21b, EVB21, GAN^{+}25, HRR^{+}20, HTP^{+}23, JL21b, dSJNS20, KYMW21, Kar22c, Kat20, KHMM20, Kim23, LTW23, LRN^{+}22, LLCH21, Los23, CSdMR^{+}24, NHD23, NAN^{+}24, NGT^{+}23, PCA21, Pha20, PJN24, SBG^{+}25, SCCW21, SVP^{+}23, SYY24, TCT21, VRBC^{+}21, VHS^{+}20, QUK21]. local-level [Kim23]. local-scale [PJN24]. localism [CSM^{+}25]. Locally [LTJ^{+}21, MGDJ22, RGK^{+}20, MB22, PLLJ24]. locals [ZNA^{+}23]. located [Lin20]. location [BLY^{+}24, COJZ^{+}21, KGW24]. locations [RNGRC22]. lock [MBG^{+}24b]. lock-in [MBG^{+}24b]. lockdown [GS21, MCBN22]. logbook [SIW^{+}21]. logbooks [CLS^{+}23]. logic [Fan24]. logistic [GSB20]. Logistical [BSY^{+}20]. logistics [CSV^{+}24]. logo [GVL22]. Lombok [AAKI21]. London [Ngu23]. Long [BSBD21, AHN25, AFMdSdTdCC22, BCBS22, BMH^{+}22, CVA22b, Jon21, McC21, MM20, NSPRSS20, SOP^{+}22, SGFM^{+}21, Wat22, SOP^{+}22, Fla21]. long-lasting [NSPRSS20]. Long-term [BSBD21, AHN25, AFMdSdTdCC22, CVA22b, Jon21, McC21, MM20, SOP^{+}22, SGFM^{+}21, Wat22, Fla21]. longer [Bra22]. longirostris [BDCJ22]. longitudinal [PLLJ24, WFG22c]. Longline [OKFN24, BdSL25, BDC^{+}21, CD23, Cha20a, CHC^{+}21, DSL20, HCS21, KB20, LLL^{+}24, dSRKTC24, SCSR^{+}22, SCSR^{+}23, SBW^{+}21]. Looking [TTA^{+}20]. losers [BCMCM21, PKMT23]. loss [FIB25, GMFD25, RWVH21, Tel23b]. losses [ABC^{+}25, GAA^{+}20, LASG^{+}25, YSWW21]. Lost [SG21, DAHS^{+}23, GRS^{+}23, GHW^{+}22, GAD^{+}23, HGG^{+}22a, Lov23, MDG^{+}23, NMN^{+}25, RCM24]. Low [ÁFFSC20, Wan23c, ZY21, ANC^{+}25, AH21, CSV^{+}24, DG24b, RF25, dlPdlLLC^{+}22, TCF^{+}25]. low-carbon [DG24b]. low-environmental-impact [AH21]. Low-impact [Wan23c, dlPdlLLC^{+}22]. Low-quality [ÁFFSC20]. Low-sulfur [ZY21]. low-trophic [ANC^{+}25, RF25]. low-tropic [CSV^{+}24]. Lower [EECE^{+}24]. LTEM [ANC^{+}25]. Lubchenco [NJW25]. Luce [LTJ^{+}21]. lucens [LSY20]. Lucia [TJ21a]. lucrative [FA24]. Lyme [SJ21].

M [NJW25, Sos21a, Sta23a]. Māori [RR20]. Macaronesia [GSGOP^{+}21, ANS^{+}20]. Macaronesian [GMGLV21]. mackerel [BDDE20, Gra21, MNS^{+}24, MKM^{+}20, SLL^{+}24, SGR21, RHST21]. Macmillan [AFHM23, LPO25]. Macro [BMC20, AVC24]. Macro-plastic [BMC20]. macroalgae [KWY^{+}22]. macroalgal [KWY^{+}22]. Madagascar [LTJ^{+}21, LJRHH21, PLLJ24, RRR^{+}23, TRL^{+}25]. made [RBNE20, ZCA^{+}23]. Madeira [DDGdS^{+}21]. Madura [HAsR24]. magnitude [GAD^{+}23, VDOE^{+}24]. Mahi [LSOP24]. Mahi-Mahi [LSOP24]. main [BN20, CM23, GSR21]. Maine [BLS21, BKG^{+}25, CCP25, CJH24, FSZ25, MM22]. mainland [BYL25]. mainstream [HM22c]. mainstreaming [TCF^{+}25]. maintaining [AH23, CCP25, McC21, ZD20b]. maintenance [KS22]. Major [SAA^{+}21, SFGEPC22, HCSBM^{+}23, LHPJ23, PL20a]. make [PK20b, RM21b]. makers [RB20]. Making [Men23, NGRM^{+}22, AMR21b, AH21, BL23a, BBV^{+}25, BBBS22, CBC24, DJC^{+}23, DBBL21, FAMW^{+}25, FRM22, GCP^{+}20, GSM20, GSG20, HDH21, HDCH21, JKCY25, KL22, KVL^{+}23, Mab24, MWM^{+}23b, MF23, MKB21, MMGAA^{+}21, Sep25, Sta25, SWW^{+}21, TDC21, Wan23b, Wan23a, YG20]. mako [DSL20, Kai21, Sel20]. Malacca [Poo23]. Malawi [MMK^{+}22, WCN^{+}25]. Malaysia [HMM^{+}24, KCCC^{+}21, MSBA22, PESO24, Pra24]. Malaysian [ALM^{+}21b, Lin20]. Maldives [YFD^{+}21]. Mali [AANM^{+}22, NSZ^{+}21]. Maluku [LLLF22, FKNG25]. Malvinas [AM22]. Mammal [BGK^{+}22, PEM^{+}25, BC20b, BC21, DOBE^{+}25, MHS^{+}25a, MHS^{+}25b, SMDM20]. mammals [BEM^{+}20, BDG^{+}25, BGK^{+}22, DMHS25, GDB^{+}24, KDH^{+}21, MGBS24, NBB^{+}25, RTMCVMDE22, VDOE^{+}24]. man [Wil21a, ZCA^{+}23]. man-made [ZCA^{+}23]. Manabique [GBC21a]. manage [BMGDS^{+}20, KLS22, VAG^{+}24]. Managed [LTJ^{+}21, OKS^{+}22, CdLN^{+}21, DTBP21, Dun21b, GV22, LJRHH21, MN23, PLLJ24, RGK^{+}20, RDLA^{+}20, ZRR^{+}25, AGK22]. Management [AS23a, APH24, BMR23, COJE24, DB20, EHM^{+}20, EHM^{+}22, FTW^{+}23, GM25a, GML^{+}24, KYMW21, KA24, MCW^{+}22a, NNM^{+}21, OC24, PCA21, Per24, Rub24, TJ21a, TK22, VLG^{+}24, APV^{+}20, AKFY22, Abo23, ABH25a, ABH25b, ÁFFSC20, AHAB^{+}22, AGM^{+}22, ABVPM20, AML^{+}21, ALF^{+}22, AGSI22, Art20, BCW^{+}25, BRSC^{+}24, BE20a, BD20, BMB^{+}24, BDM20, BEM^{+}20, BISSBdlMM22, Bea22, BSSL^{+}20, BFSP25, BMM^{+}20, BM21, BLM24, BDD^{+}21, BBT^{+}25, BBJ^{+}22, BAJR24, BDL^{+}24, BGMCG24, BWV25, BBGG22, BYL25, CSV25, CPR21, CBO^{+}21, CHH^{+}25, CMT^{+}23, CCK^{+}21, CCB^{+}21, CGB^{+}25, CCGM21, CD25, CTH^{+}21, CLRW22, CAM^{+}23, CHB^{+}20, CDP22, CLU^{+}24, CCP25, hC24, CAHR^{+}24, CWKE20, CES20, Cra20, CMD^{+}20, CDB20, CSL21, DL21, DMHS25, DT23, DBS22a, DH22, Dbo22, DSL^{+}21, DRY22, DSL20, DSV^{+}25, DAAP21, EdLP^{+}21, ETB25, ETSKA25, EAGV20]. management [EB23, EV22, EAA^{+}21, EBB^{+}25, FDGM24, FTGD^{+}20, FR24, FLB^{+}21, FCG25, FS21, FMTA^{+}25, FMLMM23, FSF^{+}21, Fuj21, GR25, GLPM24, GSAF21, GLS^{+}21, GN22, GCP^{+}20, GCC^{+}21a, GAD^{+}23, GLB^{+}24, GAN^{+}25, Gon25, GMMMRC^{+}25, GRPRVG23, GWAPRPRU22, GABA^{+}23, GTTM21, GDY25, GW21, GICM25, HMFH20, HDH21, HDCH21, HWU^{+}23, HLW^{+}20, HYN^{+}23, HGQJD24, HGG^{+}22b, HWT23, HWT24, Har22, HS23, HC25, HW22, HHBC20, HMG^{+}21b, HAsR24, HKZ25, HHH^{+}23, Hor25, Hor23, HvPPV20b, How20, HCS21, INW20, Isl25, JS23, JSW^{+}22a, JAH^{+}24, JPM^{+}23, JL21d, KWF^{+}21, KS23, Kai21, KFB^{+}20, Kar22c, KMM24, Kaw22, KSB^{+}23a, KSB^{+}23b, KKAO^{+}25, Kim23, KKO^{+}22, KGL22, KH24, KN21, KTC^{+}22, Kul23, KJS^{+}22, LGLL24, LDAC20, LLT^{+}20, LMC^{+}23a, LWH^{+}24, LBWS23, LDC^{+}20, LW22, LYHL23, LLCH21, LDL^{+}23, LH25, Liu23, LA23, LSGG^{+}20, LGW^{+}21]. management [LT24, Los24, LSY20, LLL23, LCMBGP22, MUR^{+}21a, MCRPC20, ML21, MHC^{+}25, MDLA21, MMB^{+}23, MVR21a, MN21, MFG^{+}23, MSBA22, MRD^{+}20, MAK22, MWM^{+}23b, MPH^{+}23, MNS^{+}24, MMW^{+}23, MGNP25, MSGAR20, MFF^{+}23, Nal24, NAP^{+}21, NUD^{+}23, NBB^{+}25, OEV23, ORSB22, OT23, OWCFDC21, OVSVPP^{+}22, Öİ24, PCD^{+}23, PHD^{+}24, PKB^{+}23, PVH^{+}21, PL20b, PCCR^{+}25, PK23, PMAVM20, PJN24, PMR23, PEM^{+}25, PLOG24, PMFR^{+}22, PFWM23, PKMT23, RMF^{+}21, RLS20, RR20, RYS^{+}21, RAFLRP21, RM21b, Ros22, RB22, RKS^{+}23, RSH^{+}25, SRvNN21, SK20a, SWI^{+}22, Sch24a, STIS20, STS^{+}23, SSZ^{+}24, SPHO23, SSOP^{+}21, SON21, SLW^{+}23, SCRCI24, SMC^{+}21, SCCW21, Ske24, SSMG21, SH23, Stø21, SLH22, SCBdA20, STKC22, SSL25, SLK^{+}23, SLS^{+}21, TSH22b, TGF^{+}25, TPAC24, TK23, TYS22, UISM^{+}23, VWFW22, VCC^{+}20, VDP^{+}22, VMC^{+}21, VGA25]. management [VLL^{+}24, WFK^{+}24, WJB^{+}20, WHPF20, WF25, Wat22, Wen20, WMRB22, WHAD^{+}25, WMS^{+}25, WLCG23, WCN^{+}25, WH20, WNBRSG22, WWC23, XHZ^{+}21, YWY20, YHU22, YW20, YCBW21, ZZMG^{+}24, ZJ22, ZWJvL20, ZPX^{+}22, ZHL^{+}23, ZCC^{+}24, dSdLSGB22, dMLB^{+}20, dSVT^{+}25, dlPC24, vNdB25, CAGBM22, CTH^{+}23, SH24a, De 20]. managers [Dun21b, GABA^{+}23, OCT^{+}22]. Managing [AVPAFM^{+}22, BDC23, CSLGFV20, ESD24, FRM22, Lov23, MT20, RRS25, RKS^{+}24, VTRS20, DMHS25, HIP^{+}21, HPPH22, MKAF21, SG20c, WWD24, YBW20, BMF20, Cha20b]. Manchón [GBC21b]. mandatory [CMR21, HBHJ20, MPL^{+}23]. maneuver [Lou22]. manganese [HKMS20]. Mangrove [RGH^{+}22, BNF21, dFVCGdS21, HAsR24, Isl25, MKAF21, MTC^{+}23, NHD23, QLU^{+}23, SLW^{+}23, TMS^{+}22]. mangroves [AS23b, HAsR24, IRK^{+}21, NEH^{+}23, RRR^{+}23]. mangroves-based [IRK^{+}21]. Manila [dZ20, CCG^{+}25]. mankind [dS21]. manning [Ste23]. manta [GRPE^{+}21, HS24, SEG^{+}22, DRCdMCJ21]. mantis [BMGDS^{+}20]. manufactured [FG25]. manufacturing [CQ20, FZYF23, HPH22]. Map [APF22, WP23]. MaPP [SSB^{+}23, DBS^{+}22b]. Mapping [AANM^{+}22, AN25, BE20b, BS22, FFE20, GBFMT23, HP20, HHL^{+}24, Ngu24, OTK^{+}21, Qi24, SSV^{+}23, AC21, BCJ21, CRP^{+}20, CMSH25, EBB^{+}25, Kar22c, MFRQ23, QRK^{+}23, SMC^{+}25, VGS^{+}20, WPWM^{+}22, WD20, CSCGC20]. maps [TZM^{+}22]. Mar [CBH^{+}25, DHU25, EHM^{+}22, FSRC25a, Mal21, MHS^{+}25a, Mot22, RP23a, SCSR^{+}23, SCM^{+}25, SC25, WWG^{+}25, ZGR^{+}25a]. Maranhão [dJdMCL^{+}24]. March [Ano20s, Ano21t, Ano22t, Ano23t, Ano24t, Ano25r]. Marcos [Fan24]. margin [CGB^{+}25]. marginalisation [OCD^{+}23]. marginalise [SMCW22]. marginalization [APGWTM^{+}24, Mba22]. Mariculture [RYL25, ANC^{+}25, BHB^{+}24, CHN^{+}24, DMS^{+}24, FAC^{+}23, HFD^{+}20, LHTR23, LDL^{+}23, MMW^{+}23, OWS20, OZF^{+}25, PBD^{+}23, RZ21b, Ren21a, RL24, Tel23a, XHY22, YYL20, YH20, YZG23]. marina [VWJMAJ20]. Marinas [BCS^{+}24]. Marine [BPE^{+}25, BJ25, BAAH24b, SONB^{+}22, CD23, CD25, Cha20a, CZW24, COSS21, CJ21, DBS^{+}22b, EAA^{+}22, EABA^{+}22, ESD24, FRM22, FMLMM23, FL24, GGP^{+}22, GBM^{+}25, GSGOP^{+}21, GSR21, Gar20, GLPM24, GB21, GMVM23, HSGNMGC21, Isr22, JMB^{+}20, Jon21, KJE21b, KAC22, Lia25, LK24b, LTJ^{+}21, Mah22, MFRQ23, McC21, MBB^{+}22, MTC^{+}23, NJBM^{+}23, OC24, ONN^{+}20, PCG^{+}23, PKB^{+}21, PDB^{+}23, PKB^{+}23, QTBL23, RSFvH^{+}22, RNGRC22, SMW^{+}21, SSB^{+}23, SPCCvT21, Smi23b, SV22, SSSL25, Sos21a, SLP^{+}22, Tan23b, TBCB25, TJ21a, TGM22, VFOA25, VBd^{+}21, VMC^{+}21, WGJ^{+}21, ZGY21, AHBAKH24, AAA^{+}24, AWBC20, AN25, ADK^{+}25, ACA^{+}21, ABH25b, AVC24, AED20, AHAB^{+}22, ASC25, AELR^{+}20, AE23, AP23, ADR^{+}21, AVPAFM^{+}22, ACG21, AHM22, ABVPM20, ASM^{+}21b, AFS^{+}23, AR23a, ADB^{+}22, BG25, BMR20, BPA^{+}22, BC20a, BBB25, BSM23]. marine [BBP25, BFdSdFL^{+}25, BDM20, BEM^{+}20, BB21, BJA^{+}24, BBB^{+}24, BUR^{+}22, BDG^{+}25, BGK^{+}22, Ber23, BCR25, BSI^{+}21, BCJ21, BC20b, BC21, BAAH24a, BBGG22, BMB^{+}21, CMS20, CH25, CHH^{+}25, CMGF20, CCS23, CBCA25, CMI^{+}22, CCBB^{+}25, CLRW22, CTCD21, CS20, CVH^{+}20, CQ20, CDHY22, CJL^{+}23, CL24, CDS^{+}25, CCKK22, CJ22, DDAB^{+}23, CES20, CBH^{+}24b, CBH^{+}24a, CBH^{+}25, Col20, CQAS^{+}25, CWR20, CNG^{+}23, CHB23, CM25, CLCV21, CMS^{+}21, DLJ^{+}20, DKW20, DMHS25, DFC^{+}20, DOBE^{+}25, DBS22a, DPB^{+}23, DHC22, DBBL21, DBH^{+}21, DLX21, DMS^{+}24, DSC^{+}23, DMG^{+}23, DG20, DL22, Dua22, DGMF^{+}20, EdLP^{+}21, Ehl21a, ETB25, ETSKA25, EV22, FTGD^{+}20, FFE^{+}24, FZYF23, FWGM24, FWSS^{+}25, FLB^{+}21, FR23, FSZ25, FWL^{+}22, FMS^{+}22, GFD21, GPA^{+}22, GNJR24, GPG^{+}21, GM25a, dlFGFFR^{+}20, GSG^{+}24, VKL^{+}25, GCC^{+}21a, GCMT20, GPM20]. marine [GCL21, GBC21a, GBC21b, GMMMRC^{+}25, GDB^{+}24, GZW^{+}20, GABA^{+}23, Gro20, GEB^{+}21, Gug24, GW21, GTMP20, Gul23, Guo20, HWM22, HS20, HJ21, HLCM21, HM22a, HvdBH23, HPH22, HZ23, HBR21, HPPH22, HMG^{+}21b, HPC^{+}23, HNB^{+}25, Hor23, How20, HLM^{+}20, HHL^{+}24, HGL^{+}20, HCT^{+}21, HWY22, HH25, HBHJ20, HCDN24, IOHC22, IMB^{+}24, JDF^{+}23, JHHG23, JLY23, JCMD20, JL21b, JF22, JH24, JPM^{+}23, JL21c, JSW22b, JL21d, KWF^{+}21, KAA^{+}24, Kar22c, KMM24, KFM^{+}20, KSB^{+}23a, KSB^{+}23b, KDHL24, KLK23, KWMM20a, KWMM20b, KvTNL20, KDM^{+}20, KGP^{+}22, KDH^{+}21, KGW24, Kra22, KTC^{+}22, KHdSM^{+}22, KVL^{+}23, Kuł22, Kul23, KHLS23, LR20, LHV^{+}20, LTSL22, LGF24, LJ20, LXSX21, LW22, LJ23, LWG23, LMC^{+}23b, LZX24, LYL^{+}22, LGP25, LWX^{+}20, Liu20, LXW^{+}21, LC22, LHPJ23, Liu24, LYLZ25, LASG^{+}25, LP21, LZRC23, LRC24, Lot22, MM21a, MFA^{+}24, MUR^{+}21a, MFG^{+}22]. marine [MCRPC20, MFT^{+}22, MG25, MBDG23, MSSS22, MN21, MSO23, MSBA22, MHS^{+}25a, MHS^{+}25b, MRD^{+}20, MFM21, MWM^{+}23b, MMG^{+}25, Men23, MSB^{+}25, MGBS24, MMO^{+}24, Mor20, MBL24, NNM^{+}21, NBB^{+}25, NW20, NO25, OSBA24, OGS25, OYKBA22, OOSG21, OKFN24, OCT^{+}22, OVSVPP^{+}22, PKC23, PLLJ24, PKH22, PBG23, PT23, PK23, PLWO24, Pha20, PS21, PEM^{+}25, PCCOS24, PC20, PdBA^{+}25, Pri22, PDJ22, PBH^{+}23, PMFR^{+}22, Puh23, PKMT23, QZZ20, QYD20, QS21, QTPL21, QYX21, RBA23, RSK^{+}25, RKC^{+}24, RSAP22, RLS20, RSS^{+}20, RDT^{+}23, RP23a, RJ21, Ren21b, RXN23, RXX24, RHC20, RGK^{+}20, RTMCVMDE22, RPY^{+}24, RKS^{+}23, Rus20, SAO^{+}21, SMDM20, SGB21, SAA^{+}21, San25, SC21, SvLGK24, SBM^{+}25a, SBP24, SPHO23, Sep25, SGLF^{+}23, SL20, SSDQ20, Sha20, SGK21, SLY21, SON21, SCM^{+}25, Shi24b, SBJF^{+}21, SBDC24, SS23, SJ21, SDG^{+}21]. marine [SDG^{+}22, ST24, SCC25, Sol23a, SZBA20, SPH^{+}20, SG20b, SRH23, SO20, SRS23, SWXM21, SYY24, SH25, SLY25, SSL25, SB20, TUBA22, TLYC20, TZZ^{+}21, TMAS24, THP22a, THT^{+}23, TBH^{+}21, TD21, Tho22, TV21, Ton22b, TEA^{+}22, Tra23a, THF^{+}25, TJ21b, Tru23, TCT24, TDP20, TBPC20, UPBB25, DR22, VWFW22, VDOE^{+}24, VAS^{+}25, VHD20, VAG^{+}24, VBC^{+}20, VWJMAJ20, VLG^{+}24, WD23, WWD24, WYH21, WXLD21, WLQD21, WSD21, WLY21, WCZS21, WW22, WHL22, WRL23, WNDZ23, WA24, WKK^{+}25, War23, WD20, WKN^{+}25, WBB^{+}25, WCL25, WMRB22, WHAD^{+}25, WFG^{+}22b, WMS^{+}25, WBPG20, WSDB20, WCM^{+}24, WL22b, Wu22, XHZ^{+}21, XXMZ23, XXLW24, YJG24, YCL25, YT20, YZH22, YW20, YML^{+}21, YCL22b, YD22, Yu22, YSL^{+}25, YC21b, ZRR^{+}25, ZZMG^{+}24, ZK21, ZJ22, ZLG^{+}24, ZW21, ZLPT23, ZPX^{+}22, ZHL^{+}23, ZWXL21, ZWC^{+}23, ZLH21, Zhu22, Zhu23, ZCA^{+}23, ZBF^{+}22]. marine [dSOW^{+}21, dS22a, dO23, dV21, vTHS21, ANS^{+}20, AIG^{+}20, ÁFFSC20, AP23, AVF23b, BI21, BvBN^{+}20, BEH^{+}21, BTN^{+}24, BPJ^{+}24, CN24, CR23, CS21, CLW^{+}21, DBH^{+}21, Ehl21b, FGK^{+}20, Gel25, GDB^{+}24, GKP^{+}22, GABA^{+}22, GDY25, JH24, JAK^{+}23, KWK^{+}25, KMT^{+}21, LGLL24, LZL25, LJ21, MCRPC20, MPD^{+}23, MDLA21, MPH^{+}23, Mor20, NRF21, ORSB22, QTD21, QTBL23, RHG^{+}25, RA20, RP23b, RCM^{+}22, Rub24, SK20a, SSB^{+}23, SRSGMGG21a, SRSGMGG21b, TOL^{+}22, VRBC^{+}21, VBH^{+}23, WWD23, WJLSH21, WB24, ZY21, De 20, GAD^{+}23, LPO25]. marine-derived [VAS^{+}25]. marine-related [QZZ20]. marine/maritime [ACG21, Ehl21a, ZK21]. Maritime [AB20, AHR24, BG20a, BRKS25, BMRM23, CLZ23, CZX24, DBH^{+}22, DZZC24, GJY23, HH21, HRMPCRRA22, Kar22b, KCB24, Kle21, LP20, Lv20, MK24, NM23, Øst20, PDB^{+}23, RBTW24, RJAT23, Smi23b, Ste25, SW22, TKC23, Tse21, VDP^{+}22, Wan24b, YC21a, ZYTT21b, Ada22, ATM^{+}24, ASKI20, ACG21, BG25, BN20, BRP^{+}23, Boy21, BL23b, CT24, CPV^{+}21, CVCV^{+}22, CSQ^{+}24, CCNGS25, CZW20, Cha21, CN23, CGP23, CZ24, Day24, Dom23, Dun23, DGMF^{+}20, Ehl21a, EVPMKP22, FFE20, Fra23, FGC21, GAGLS25, GEGRdL24, GSR21, GG22, GAD^{+}24, GCL21, GCC^{+}21b, GMC24, GW21, HPS23, HMM^{+}24, HS22, IGS^{+}23, JF24, JD20, Kar22a, KBC25, KMSN25, KSS23, KF23, LPMS24, LDG^{+}25, LKB^{+}23, LC20a, LLW22, Lin25, LF25, LEVET21, MC25, MLJ^{+}22, MCL22, Mar21a, MPL^{+}23, MBORGN24, MNHS25, OYECdOP24, PB20, PGP22, PGM^{+}24, PGA^{+}20]. maritime [Poo23, PK20b, PKS20, Qaf24, QF20, RKKM24, RLRS21, Shi24b, SVG24, SF22, SPNP24, SSZ^{+}22, SRS^{+}25, SPT^{+}20, SHZC21, SCZ^{+}24, SADP25, Tod20, Wan23a, Wan24a, WZX24, Wan24c, WOG20, WFG22c, WL24a, XT23, ZK21, ZD20b, ZMZ23, ZT23a, ZSLY20, ZYTT21a, Zou23, vLPBK21, vdBSvdV^{+}23, ANS^{+}20, AMK^{+}21, BAP21, BRP^{+}23, CK23, COJZ^{+}21, DBZ24, FTW^{+}23, GBCNCL^{+}21, KMTvT25, KBC25, RBB^{+}24a, SWGZ22, SGRW22, WN25, ZWG^{+}21]. markers [Los24]. Market [PGGA22, YWY20, AMIM21, APGWTM^{+}24, ABC^{+}24, CMI^{+}22, Cat20, CMNH21, Cro21, FGPPPV21, HCSBM^{+}23, Lin25, MMM^{+}22, MHT^{+}24, NTH^{+}21, PFM20, PLOG24, SGZH^{+}21, Sta25, dSTHL^{+}24, VJÓK25, VKN20, YZH22, YLSC23]. Market-based [YWY20, VJÓK25]. marketing [AOS^{+}24, CGAF25, GM21, GPM23]. markets [GAD^{+}20, Hor23, KKA^{+}21, OGH24, PSG25, SGA^{+}21, SOS20, SFN^{+}21]. Markov [AS23b]. MARPOL [VSV^{+}23]. Marshall [CTIH21]. MARTEC18 [ABVPM20]. Mascarene [PEM^{+}25]. MASS [JD20, MNHS25]. massive [HZC^{+}24, JG20, Kuł22]. master [CTC^{+}23]. match [RHC20]. matched [ZTM^{+}23]. Matching [GHW^{+}22]. materialities [Teo23]. Mathematical [vLPBK21]. matter [CSV^{+}24, HS23, JL21b, Øst20, TJZ^{+}21, WZZC21]. matters [BAP21, HRR^{+}20, LP20, OTK^{+}21, Sin21]. maturity [TPG^{+}23]. Mau [NHD23]. Mauritania [BGMCG24, FA24, TGHF^{+}22]. Mauritius [KMS20]. maw [BHdMCM^{+}21]. Maxent [AS23b]. maximisation [ASGS23]. Maximizing [PT23]. May [Mot22, MM20, Ano20t, Ano21u, Ano22u, Ano23u, Ano24u, Ano25s]. Maya [TGM22, SCD21, GÁSVPS^{+}25]. Maya-Jariego [TGM22, SCD21]. Mbenji [WCN^{+}25]. Mbour [CDNS23]. MCE [Day24]. MCH [BTN^{+}24]. MCS [FOKM23]. meadows [BAJR24, CMMVdPV25, KWY^{+}22, SML^{+}25]. mean [Bal22, CMR21, Kim20, QZZ20, SCF^{+}22]. means [MZ22]. measurable [KTC^{+}22]. measure [KYMW21, MR24, SPMN23]. measured [GS21]. Measurement [CDHY22]. Measurements [WHAD^{+}25, BdSL25, VSV^{+}23]. Measures [BJH21, SBM^{+}25b, APV^{+}20, Ada22, ADK^{+}25, CPR21, CHB^{+}20, CMWN^{+}23, DH22, DG23, DER^{+}21, EAA^{+}22, FTGD^{+}20, HRS^{+}24, Ma20, MBDG23, NNR^{+}21, PL24a, Ros22, SMLRB^{+}21, Sch24a, TAK^{+}22, WJB^{+}20, ZCC^{+}24, GKP^{+}22]. Measuring [ADR^{+}21, Che21a, CLW^{+}21, Gul20, POM24, Run24, SX25, SBW^{+}21, TMM^{+}23, WW21b, HCDN24, PWL^{+}23a, RBNE20]. meat [BBHN20, OAdJC^{+}25, SCBB^{+}22]. mechanism [BDD^{+}21, JCMD20, Lin24, LC20b, MFA^{+}24, SWHT22, TT22, XL25, XHZ^{+}21]. mechanisms [GBCNCL^{+}21, OPL20, SLO^{+}24, TOL^{+}22, WHAD^{+}25, YSL^{+}25]. MEDDZs [LZL25]. Media [HPG20, PBG23, CL24, CVA22b, ES22, LDL21, Los23, RM21a, SCD21, SMW^{+}21, TMM^{+}23, WYH21, WXLD21, ZCG^{+}23, ZLWY22]. mediate [RF25]. mediating [LM24, TNY23]. medicine [GLT23]. Medinipur [CDP22]. Mediterranean [BRSC^{+}24, CGB^{+}25, CMC^{+}23, GDY25, HRMPCRRA22, VCC^{+}21, ATF^{+}21, BSBD21, BSSL^{+}20, Bjø23, BMM^{+}24, CCS23, CCB^{+}21, CGMH24, CDD^{+}23, CCBB^{+}25, DL21, DFC^{+}20, DDF^{+}20, DSD^{+}24, DGIS^{+}22, EECE^{+}24, ES22, GM21, GCL21, GKP^{+}22, GMC24, HNMSG21, Kle21, KGL22, LSGG^{+}20, MMB^{+}23, MHF^{+}25, May22, NGT^{+}23, OCCSC23, OSZ^{+}25, OVSVPP^{+}22, SAS24, SPCCvT21, Tes20, VPB^{+}24, VGK23, WML25, dJRFVOA24]. medium [Bre25, HYN^{+}23]. medium-scale [HYN^{+}23]. medium-sized [Bre25]. meet [GDLC20, RSFvH^{+}22, VDP^{+}22]. Meeting [HAD24, GPH^{+}21, YK21]. megafauna [BUR^{+}22]. megalopolis [LDG^{+}25]. Megalops [GCM^{+}23]. Megaptera [MCW^{+}22b]. Mekong [NHD23]. Melanesia [ML21]. Member [ZGR^{+}25a, ZGR^{+}25b, Puh23]. membership [TSS^{+}24]. Memorializing [TCD^{+}20]. men [MMK^{+}22]. MENA [BEH^{+}21]. Mental [FAGS21, MEG^{+}21, PGK22, TAK^{+}22]. mer [PSG25]. Merchant [Fel22, HVN21, PZPB21, XZ23]. Merluccius [BDCJ22, RQB^{+}20]. mesh [GMM^{+}20]. mesopelagic [SG20a, SG21]. Mesopelagics [FTG^{+}23]. meta [BCB20]. meta-organization [BCB20]. metabarcoding [CCS23]. meteorology [FRM22]. Method [SX25, BMGDS^{+}20, BTV^{+}22, CA21, DXS^{+}21, LrS22, Okş23, PMC24, WLQD21, OTK^{+}21]. Methodological [IO23]. methodology [CGGA20, GKP^{+}22, POM24]. Methods [OSBA24, AMLC^{+}23, GAD^{+}23]. metrics [KTC^{+}22, SS21a]. metropolitan [PdBA^{+}25, SdCMN^{+}22]. Mexican [CSCGC20, LEERRM^{+}21, LCMBGP22, MVSP^{+}21, PGSFRMHH24, PL20b, RTMCVMDE22, SMLRB^{+}21, VSPMS^{+}22]. Mexico [SLP^{+}22, AAFK^{+}21, AFRFGP25, BCMCM21, CLL20, CEMRMB^{+}24, GÁPSAP^{+}23, GÁSVPS^{+}25, GBAECCRM25, GRSRSN^{+}24, GMMMRC^{+}25, GWCD22, HDAPG^{+}24, MMSB^{+}20, MAMCHP^{+}20, RSLR24, SMHPMO^{+}21, TBCB25, TRL^{+}25, VHCJdYC^{+}24, CSMMC^{+}25]. MEY [PHD^{+}24]. MFAD [WLRH^{+}20]. Miami [DMK24]. Michael [APH24, AFHM23]. Michel [Gir22]. micro [OVSVPP^{+}22]. micro-reserves [OVSVPP^{+}22]. microbeads [MLM23]. microbiological [PL23]. Micronesian [WGP^{+}21]. Microplastic [UPBB25, BR22, DVRK25, DDD22, Pra24, APH24]. microplastics [lTS25, Hos24, SZ24]. Mid [CSGC22]. Mid-Atlantic [CSGC22]. middle [DMNT20, JKCY25, BTN^{+}24, TCD^{+}20]. migrant [CJ23, YL21]. migrants [ALM^{+}21b]. migrating [MCW^{+}22b]. migration [DPG^{+}22, Kle21, MSSS22]. migratory [ABM21, MWM^{+}23b]. Mi’kmaq [DF24]. mile [WEF^{+}21]. militarized [HGLR22]. Military [CLD^{+}24, ABI24, Dun24, LF25, Rog24, Shi24a]. milkfish [MGS21, SGPF21]. Mind [HSSM24]. Mindanao [GM25a]. minds [CMWT24]. mine [ATF^{+}21]. mineral [KGW20, WRV20]. Ming [LK24b]. minimising [HRC^{+}24]. minimize [Dam24, SCSR^{+}22, SCSR^{+}23]. minimum [FMCD20, YU20]. Mining [Sin21, WRV20, AGM^{+}22, AR22, BHKJT23, BRM20, CGHV20, CDC20a, CDC20b, CL20, Cra20, DMd^{+}20, Fei20, GG20, HKMS20, HSSM24, HSK^{+}23, HZC^{+}24, Jae20, JACD20, Kak20, KGW20, MCR22, MM21b, PCG^{+}23, PL24b, PLHK^{+}24, RFPS20, Squ21, TML^{+}20, VMV24b, WSZ^{+}25, WABR^{+}22, WLCC23, Wil20, Wil21a, Wil22, Yad24, ZTA23, vdGD21]. Ministerial [ALLS23]. Miramar [HM21]. Mis [BMF20]. Misalignment [MS21]. miscellaneous [BBB23]. misinform [SMW^{+}21]. mislabelling [GAD^{+}20]. mismatch [BXV20]. Mismatches [PdBA^{+}25]. misperceptions [RWVH21]. missed [RL24]. misuse [HS23]. mitigate [BKW^{+}24, BRP^{+}23, GLMBMP25, JGWM^{+}23, MZT^{+}23, SBM^{+}25b, TOL^{+}22, VÖB20a, VÖB20b]. Mitigating [CHC^{+}21, AHN25, WS23]. mitigation [AFR25, ABVPM20, BEM^{+}20, CJLJ24, CLD^{+}24, DSB21, DBB^{+}22, GCB^{+}23, GAD^{+}23, HSGS^{+}23, MDLA21, MSGAR20, NHD23, SNP21, SAB^{+}20, UPBB25, WLW^{+}22, ZWJvL20, ZCSK23, Rub24]. mixed [CPR21, COJE24, GSG20, MH20a, NGP^{+}21, NSK20]. Miyakojima [Tru23]. MLS [YU20]. mm [GMM^{+}20]. mobile [FARBV^{+}24]. Mobilising [HPG20]. mobility [ATF^{+}21]. Mobilizing [WGP^{+}21]. modalities [TGP^{+}24]. model [BCB20, BCR25, DLYM22, HLW^{+}23, HRR^{+}20, JWX^{+}20, Kaw22, KDH^{+}21, MDG^{+}23, MLM23, MPN^{+}24, Obu20, PHF20, PDD^{+}25, PGSS22, PMT^{+}20, SGK21, VPRR^{+}21, WHPF20, WCZS21, WL24b, Zha21, ZHL^{+}23, ZYTT21b, GMG^{+}23]. Modeling [OR21, SA21, Cat20, SGFM^{+}21, WGR^{+}22]. Modelling [ZCG^{+}23, BVT^{+}21, CAM^{+}23, CFA^{+}24, MDLA21, PB20, THT^{+}23, vLPBK21]. models [AS23b, BDR20, DKW20, FAGS21, Gal21, GCL21, HC25, KPN20, MEG^{+}21, PKH22, PBT^{+}20, Pra20, WL22b]. moderating [Lin25, NRYG20]. Modern [LL24a, Ege23, WKBS20]. modernisation [KBC25]. modernist [Alf20]. modernization [TWK25]. modes [VHSV20, KA21]. moja [CMWT24]. Molidae [PNS^{+}23]. Molidea [PNS^{+}23]. mollusc [AOO^{+}23b, PL23]. momentum [LXSX21]. Mondego [SCBdA20]. money [MJMAA21, SCD21, TGM22]. monitor [GMG^{+}23, GLMBMP25, TLT^{+}22]. Monitoring [BLH^{+}20, LWSS21, NJBM^{+}23, PKB^{+}21, Rub24, SIW^{+}21, SG20c, WCF^{+}21, WCM^{+}24, XWH^{+}22, vdBSvdV^{+}23, AIG^{+}20, BCJ21, BDC^{+}21, CCS23, CWR20, EAA^{+}21, EHM^{+}20, EHM^{+}22, GCC^{+}21a, GDLC20, GAD^{+}23, KMS20, KDM^{+}20, LGLL24, MGBS24, OWS20, PSTW21, SCA^{+}20, SAS24, VAG^{+}24, WPWM^{+}22, WD20, ZLD^{+}21]. Mont [Gir22]. Mont-Saint-Michel [Gir22]. Monterrico [GBC21b]. Monument [Cha20a]. Moore [Ngu23]. Moorea [SO20]. moored [WLRH^{+}20]. moratorium [DLX21, ZMT^{+}25]. morio [HDAPG^{+}24, RSLR24]. Mortality [HS23, OKS^{+}22, SLMG^{+}21]. mosaic [THP^{+}22b]. most [Men23, OVSVPP^{+}22, SZ24, VPRR^{+}21]. mother [SMKS22]. mother-of-pearl [SMKS22]. motivate [WLQD21]. motivation [EVPMKP22, JBGC25]. Motivations [GRPE^{+}21, MFM21]. MoU [XWL^{+}20]. Mount [LA23]. move [JPM^{+}23]. movement [Lot22]. movements [CMSH25, GCV^{+}20]. Moving [GBAECCRM25, Gue24, RDM^{+}20, DUYEA21, VBH^{+}23, Cla20]. Mozambique [RDM^{+}20]. MPA [BTN^{+}24, FWSS^{+}25, GDY25, SSAAB^{+}24, VMC^{+}21, ABM21, BBK^{+}21b, CLS^{+}23, GLPM24, dOJCSLdSB21, MGNP25, NBB^{+}23, PSTW21, SRR^{+}21, SSAAB^{+}24, WD20]. MPAG [JL21c]. MPAs [CWB^{+}20, CSGCD25, DSD^{+}24, EAA^{+}22, GM25a, HS24, HNMSG21, Kat20, Mur21b, RCH22, TCZ21, VvSvH^{+}22, dPJ21]. MSC [BI21, GMG^{+}24, JAK^{+}23, SK20a, DQJ23, GBDC24, GML^{+}24, JC23, KMJ24, LLM^{+}23, MLA^{+}24, PCK22, SPG20, WS21]. MSC-certified [SPG20]. MSE [COJE24]. MSP [ZGR^{+}25a, ZGR^{+}25b, AMK^{+}21, BMM^{+}24, CMOJ^{+}24, HvT22, QRK^{+}23]. MSY [TM23a]. MSY-based [TM23a]. Mtwara [MYGM25]. Much [OSV21, DMNT20, SMHPMO^{+}21]. mud [DMS^{+}24, PLLJ24, RKR20, HCT^{+}21]. mullet [MKM^{+}20]. Multi [LRD^{+}22, NUD^{+}23, ZLPT23, ASV^{+}21, BCB20, COJE24, CPBO20, CvHF^{+}24, CDB20, hBDDF22, EAGV20, GRSRSN^{+}24, VKL^{+}25, HvdBH23, HW22, HBR^{+}23, KLSG21, Mba22, PLWO24, RBB^{+}24a, Sen25, SMB^{+}25, SCB25, SSZ^{+}22, TMS^{+}22, VTRS20, VAG^{+}24, VDN^{+}24, WOR^{+}25]. multi-criteria [ASV^{+}21, CPBO20, EAGV20, HW22, TMS^{+}22]. multi-decadal [PLWO24]. Multi-dimensional [LRD^{+}22, Sen25]. multi-level [Mba22]. multi-levelled [HvdBH23]. multi-national [VDN^{+}24]. multi-risk [HBR^{+}23]. multi-scalar [RBB^{+}24a]. Multi-scale [ZLPT23, VTRS20]. multi-specific [COJE24, GRSRSN^{+}24]. multi-stakeholder [BCB20, CDB20]. multi-usage [hBDDF22]. multi-use [CvHF^{+}24, VKL^{+}25, SMB^{+}25, SCB25, SSZ^{+}22, VAG^{+}24, WOR^{+}25]. multi-vessel [KLSG21]. Multi-year [NUD^{+}23]. multiagent [Qi22]. multidimensional [GGZRAHVS25, HIP^{+}21, RMBR22]. multifaceted [ACG21]. Multilateral [Poo23]. multilayered [ORSB22]. multilevel [SOLN21]. multilingual [Pau21]. multimodal [CQW24]. multinational [CRP^{+}20]. Multiple [DČFW23, DB20, PFWM23, BDM21, hBDF24, GAD^{+}24, GAD^{+}20, HvPPV20b, MM21a, MGF20, OCT^{+}22, PCD^{+}23, SHP^{+}21, WGR^{+}22, GBC21b, Has20]. Multiple-goal [PFWM23]. multiple-use [MM21a]. multispecies [BSLRR24, CTH^{+}23, GMPT^{+}20, GVH^{+}23]. Multivariate [MDG^{+}23]. municipality [IAB^{+}20, vLS24]. Murcia [ASV^{+}21]. murky [Wil22]. mussel [SPCCvT21]. mussels [Gir22]. mutual [JF20, LK24a, MPN^{+}24]. mutualism [CMI^{+}22]. Myanmar [HBJ21]. Myeik [HBJ21]. myths [Øst24].

Nam [Han21]. name [LSC25]. Namibia [FGK^{+}20]. naming [CMNH21]. Napoleon [AGSI22]. Nariño [PCCOS24]. narrative [CMOB^{+}20, MCL22, MSO23, Obu20]. narratives [ARP25, BDMW22, Dyr25, LLT^{+}20, TTA^{+}20]. narrow [GCV^{+}20]. Nash [Han20]. Natalie [SSSL25, Tan23b]. Nation [BMB^{+}21, HHH^{+}23]. National [Cha20a, COSS21, CJ21, GASLCP20, Has23, HNMSG21, KJE21a, MTN23, Teo23, TM23b, TCD^{+}20, YCL22b, BLL22, CWR20, CM25, DMNT20, Fox22, FMS^{+}22, GNJR24, Gar20, HJ21, HM22a, Hon20, HGL^{+}20, KAN^{+}23, LrS22, MYLV24, Mar21a, MDJ^{+}22, MH24, MCR22, ONA21, OKR^{+}20, PBD^{+}23, Pri22, QZZX20, Qi22, Qi24, SL23, TD21, TMDN23, VHD20, VAG^{+}24, VBC^{+}20, VDN^{+}24, WF25, Wil20, WGJ^{+}21, YG20, YML^{+}21, ZJ22, ZD20a, ZT21, dKMK21, ASGS23, HMM^{+}24, QTD21, SLP^{+}22]. nationality [HVN21]. nationally [Moh23]. nations [BDDE20, BA25b, TFH^{+}23, Gar20, KS22, MH24]. native [FSZ25, GTTM21]. Natura [PdBA^{+}25]. natural [AML^{+}21, BDDE20, BEH^{+}21, BYL25, EMN^{+}24, GMFD25, GWÁV^{+}25, KMM24, MYGM25, OGS24, RXX24, RMBR22, SMLRB^{+}21, SOP^{+}22, JL21d]. Nature [Sta23a, vNdB25, CWLD25, DMK24, DRY22, MR24, PCA^{+}23, RSFvH^{+}22, SVKP21, SRH^{+}22, Zha25, HDHM^{+}20, LGP25, OWC^{+}24]. nature-based [DMK24, MR24, RSFvH^{+}22, OWC^{+}24, LGP25]. nautical [GBFMT23]. naval [Cha24a, Dun23, Dun24]. navigate [WCC^{+}20]. Navigating [BBP25, DBZ24, GAD^{+}24, HCG^{+}25, KCB24, KDHL24, LF25, MRD^{+}20, OC24, SAC25, SYY24, SLS^{+}21, Wan24b, WKK^{+}25, CDY24, SS21b]. Navigation [Ngu23]. navigator [KP24]. Nazca [CMWN^{+}23, DBH^{+}22, GWÁV^{+}25, WvdMG^{+}21]. near [WPWM^{+}22]. nearly [LJ20]. nearshore [DLJ^{+}20, SCCW21]. necessary [BTE21, GR20a, VvSvH^{+}22]. necessity [Dau23, KCB24]. need [AGK22, BBB23, BB20, Cho22, FDF^{+}21, GCG^{+}21, HMM^{+}24, MHC^{+}25, MH24, Öİ24, Sin21, ZZMG^{+}24]. needed [MJMAA22, McC20, SOO20, SCD21, Tra23b, VCC^{+}20]. needs [ASKI20, GPG^{+}21, GWÁV^{+}25, HGG22c, MAMCHP^{+}20, MKSG20, UISM^{+}23, VDP^{+}22, WM22b]. negative [AHENNA^{+}24, Cao25, GLMBMP25, HACCM22, SCSR^{+}22, SCSR^{+}23]. negligible [QF20]. negotiable [CQW24]. Negotiating [Alf20, dlC24, DAB^{+}25]. negotiation [Sas25a]. Negotiations [TMDN23, BLL22, De 22, Dua22, HW22, HHD20, MTN23, SCGC20]. neighbours [PK20b]. Nemo [HM22c]. Nengye [Sos21a]. Neoliberal [Ree20]. Neostructural [BBGHR23]. neotropical [dMVJdAdSM25]. Nephrops [BMGDS^{+}20, SSW^{+}25]. Nerlovian [OAM20]. nesting [LTRA20]. net [ASPT24, BFSP25, SKL^{+}20, WLW^{+}22]. Netherlands [dV21]. Netnography [SCD21, TGM22, MJMAA21]. nets [DAHS^{+}23]. Netting [TFP25]. Network [ANR^{+}22, MFG^{+}22, SWH^{+}23, WJLSH21, vLS24, BBB^{+}24, CCS23, CCCC21, CSGCD25, FFD^{+}23, GLPM24, GPM23, GMRSP23, GGL22, KABT21, LZRC23, MF20, ORSB22, PTS^{+}25, PdBA^{+}25, Smy24, SLS^{+}20, VWFW22, VMD^{+}24, WW21a, WBB^{+}25, WGR^{+}22, WMRB22, XHZ^{+}21, YWYL22, ZLGX20, ZWG^{+}21, ZHL^{+}23]. networks [AOO^{+}23a, ABM21, ACV^{+}21, AHB^{+}23, BPE^{+}25, CS20, DGI^{+}24, DTBP21, HRC^{+}24, KBB^{+}25, MIMS^{+}20, PN20, RKS^{+}24, SCC25, TBCB25, TFP25, WMS21, KA21]. neural [KBB^{+}25]. neutral [GCB^{+}23]. neutrality [HZ23]. never [Gal22]. new-type [Zha22]. Newfoundland [DC20, EVB21, Fro21, NGRM^{+}22]. news [CL24, CVA22b, LDL21, RM21a, TMM^{+}23, YXDL24]. Next [YK21, BÖ24, Set23]. nexus [AHR24, DRS^{+}25, KSB^{+}23a, SPNP24]. NFDS [ASGS23]. Nha [KJE21b]. Nicaragua [LZRC23]. niche [TCF^{+}25]. Nicole [Lv20]. Niger [NSZ^{+}21, UOO20]. Nigeria [BOOO23, CLU^{+}24, KTI^{+}24, MAA^{+}24, NW20, NO25, OZF^{+}25, UOO20]. nighttime [TLT^{+}22]. Níjar [HSGNMGC21]. Ningaloo [Jon21]. Niño [HGLR22]. NIS [Kul23]. NL [MF20]. No [GVL22, BBB^{+}20, Bra22, Che21b, GV22, HLCM21, HS24, LC20a, QTPL21, QTBL23, TPAC24, WBPG20]. no-take [BBB^{+}20, HS24, QTPL21, QTBL23, WBPG20]. noa [MAR^{+}20b]. noa/we [MAR^{+}20b]. nocturnal [dSRKTC24]. nodes [HRC^{+}24, dlPC24]. nodule [HKMS20, JACD20, WABR^{+}22]. noise [AdSMS21, BBV^{+}25, BRP^{+}23, CLD^{+}24, DOBE^{+}25, KDH^{+}21, MG22, Tod25, VÖB20a, VÖB20b, VRBC^{+}21]. Non [BCS^{+}24, Cat20, SPN^{+}21, BPA^{+}22, CMS20, CC23, DR25, EMH^{+}23, HvT22, HWY22, KL22, Lin23, MCS^{+}25a, MJ23, MEG^{+}21, Ngu24, OAM21, Pet22, PCCOS24, QF20, RROP23, SCSR^{+}22, SCSR^{+}23, SON21, WMRB22, WCM^{+}24, ZRR^{+}25, ZT21]. non-binding [HvT22]. non-climate [PCCOS24]. non-competitive [ZT21]. non-compliance [Lin23, MJ23]. non-discretionary [OAM21]. non-disputant [Ngu24]. non-entangling [EMH^{+}23]. non-governmental [CMS20, HWY22, KL22, Pet22, WMRB22, ZRR^{+}25]. Non-indigenous [BCS^{+}24, SPN^{+}21, MCS^{+}25a, WCM^{+}24]. non-lethal [SON21]. non-linear [MEG^{+}21]. Non-market [Cat20]. non-negligible [QF20]. non-state [BPA^{+}22, CC23]. non-target [SCSR^{+}22, SCSR^{+}23]. non-use [DR25]. noncompliance [ÁTGCE^{+}25]. nondegradable [VJL^{+}21]. nonindigenous [Kul23]. Nonlinear [Sha20]. norm [AH21]. Normative [Fro21, Gra21]. norms [BPE^{+}25, MJ23]. North [KABT21, NEH^{+}23, NC25, SZBA20, vHSSK20, DXS^{+}21, HNMSG21, Mag23, MFRQ23, OSZ^{+}25, AFR25, ALML23, ÁFFSC20, BRSC^{+}24, BG20b, BvdBvdFK20, BDCJ22, BM21, BB20, BLY^{+}24, BGMCG24, DBS^{+}22b, DMS^{+}24, FSKJ23, FKNG25, FFE^{+}24, GLADGGVL23, Gra21, HS22, HvdBH23, Kai21, KMTvT25, MCRR23, MBB^{+}22, MM20, PS21, PKMT23, RHL20, SOT^{+}22, SNS23, SV22, SR22, TCS^{+}20, TYS22, VBH^{+}23, VGK23, WB24, YK20, dV21]. North-East [ÁFFSC20, FFE^{+}24, VBH^{+}23]. North-Eastern [DMS^{+}24]. north-west [DXS^{+}21]. North-western [NEH^{+}23, HNMSG21, MFRQ23, OSZ^{+}25, GLADGGVL23, BRSC^{+}24]. Northeast [CCGM21, CDRS22, dFVCGdS21, GICM25, JHL21, MKM23, MDLA21, WD23, Bjø23, BUM^{+}24, CPV^{+}21, EB24, FTSG21, HC21, MKSG20, RHST21]. Northeastern [WW21b]. Northern [PCA21, ÁTGCE^{+}25, BLM^{+}25, CBO^{+}21, CC21, CDHY22, EVPMKP22, GPTASM23, GRPRVG23, GNMGWM21, HDAPG^{+}24, JGWM^{+}23, dSJNS20, OKH^{+}24, WLW^{+}22, dPJ21, BISSBdlMM22, EMN^{+}24, EMS^{+}23, Gun21, HLLK22, KvTNL20, ONN^{+}20, RZ21a, RKKM24, SVG24, Tod22a, VMV24b, VJ24, WDG^{+}20, Wan23c]. NorthSEE [KMTvT25]. Northwest [dlFGFFR^{+}20, DC20, MSGAR20, HLW^{+}23, SBW^{+}21, ZNA^{+}23]. Northwestern [WCF^{+}21, BMR20]. norvegicus [BMGDS^{+}20]. Norway [ASGZ22, DNR21, Dyr25, EMS^{+}23, GBLBV20, GVL22, HJLN24, KKO^{+}22, LDC^{+}24, LAF^{+}24, MTS^{+}24, ONN^{+}20, SDI^{+}20, SLL^{+}24, SAS^{+}25b, SGZH^{+}21, SGH22, SGL20, SG20a, SH23, SA25, TSH22b, VMJ25]. Norwegian [Åm21, AOO^{+}23a, AHT25, BE20a, BE20b, BDR20, BDDE20, Fla21, GVR20, HO25, Her22, Knu25, RAB22, SZ20, SLSV25, SH24a, STG^{+}20, SAO^{+}24, TFF22, YHU22]. note [AEJ24, Tes20, ZT23b]. Notes [Smi22]. nothing [OSV21, TPAC24]. notification [For22]. notification-driven [For22]. notion [Urb22]. Nova [CA20, FRLMW25, KdBK25, TFF22]. novaeangliae [MCW^{+}22b]. novel [CAC23, CLCV21, DQJ23, IKB21, LJRHH21, PVH^{+}21, WWD24]. November [Ano20u, Ano21v, Ano22v, Ano23v, Ano24v, Ano25t]. nuclear [CL24, FL24, LWZ23, Lia25, LCL24, Mab24, QL24, WL22a, WZL25, XTY22, YXDL24]. nullification [Sch24a]. number [KPN20]. numbers [LLM^{+}23, WD23]. Nunavut [KDH^{+}21, BOW^{+}22]. nursery [MCW^{+}22b]. Nusa [RYS^{+}21, YC21b]. nutrient [RMF^{+}21]. nutrition [OGH24, TCC^{+}22]. nuts [SG20a]. Nuxalk [BMM^{+}20]. NW [WML25, CGB^{+}25, Sol23a, VCC^{+}21].

object [NM23, SHPW21, dZ20]. objection [Sch24a]. objectives [GDLC20, HGH^{+}20, HvPPV20b, LBL21, Mur21b, OGS24, PCD^{+}23, PFWM23, YW20, TML^{+}20]. objects [PKH22]. obligation [FGPPGG23, GMPT^{+}20, Guo20, KS22, OSV21, PMAVM20, PV23, SOLN21]. obligations [Har22, KL22, WL24b]. observation [MAMCHP^{+}20, TZM^{+}22, VHSV20, WBB^{+}25]. Observations [BMR20, CMS^{+}25, SBM^{+}25a]. observer [EHM^{+}20, EHM^{+}22, HTH25]. observers [FSKJ23, Gar24, HTH25]. obstacles [AMLC^{+}23, VJO23]. obtained [KMG^{+}20]. obvious [NBB^{+}23]. occupation [HVN21]. occupational [BBV^{+}25, PK20a, SHZ^{+}21, Sha22]. occur [CMS^{+}21]. Occurrence [Kul23, DVRK25, PY20]. Ocean [GMG^{+}23, HDHM^{+}20, HGG22c, Kar23, LL24a, MZ24, Mol24, MLAS23, NJW25, OAdJC^{+}25, Pol23, Shi24b, SBS21, Smi20, SWGZ22, Teo23, TSWC23, VDN^{+}24, VYB20, YK21, vNdB25, vSHHT20, AAA^{+}24, AAA^{+}21, AFHHD22, AHN^{+}23, BLH^{+}20, BH23b, BBWC21, BAF^{+}23, BCB20, BDM21, CSV25, CC23, CSV^{+}24, Cha24b, CMWZ24, CJLJ24, CFS^{+}24, CSH23, CNG^{+}23, COJS^{+}22, EFFP23, EBB^{+}25, FDGM24, FZL25, GPA^{+}22, GGP^{+}22, GBC^{+}20, GAGLS25, GWÁV^{+}25, HJPL23, HLO^{+}21, HCDN24, Joh25, JV20, Kim24, KvTNL20, KN21, LLT^{+}20, LZL23, Liu24, MLE^{+}24, MCL22, MZZ24, MSP23, MFG^{+}23, MGDJ22, MF25, MMNC22, MMG^{+}25, MSB^{+}25, Moh23, MM21b, NMCL24, OKNB22, OBNK24, PDS24, PBC^{+}24, PMN^{+}22, PNS^{+}23, PCHD^{+}25, PMC24, PSA^{+}25, Pyć23, SdACdSLdAS24, SBM^{+}25a, SX25, Shi20b, SWHT22, SMC^{+}21, SAS^{+}25b, SDG^{+}21, SDG^{+}22, SCA^{+}22, SZ24, TZM^{+}22, TSA^{+}22]. ocean [Tho22, TGP^{+}24, VHS^{+}20, WNJ25, WPWM^{+}22, WWS^{+}25, WNBHG24, XT23, YG20, YHX24, YWYL22, YC21a, ZJ22, ZCZ20, ZT21, ZZD^{+}23, ZTA23, Alf20, BdSL25, BUM^{+}24, CTH^{+}21, DFP^{+}25, DSL20, FIB25, GSGOP^{+}21, Gar20, VKL^{+}25, GB21, HLW^{+}23, KWF^{+}21, Kai21, KLS^{+}25, KY20, LCMBGP22, MFT^{+}22, MDG^{+}23, McC24, MAG^{+}22, MSGAR20, NOL^{+}23, NWS^{+}24, NT20, PH24, Pri22, SGWL21, Sta23a, TBCB25, TGF^{+}25, Tra23b, TMTC20, TYS22, VYB20, WS24, WCF^{+}21, Wen20, YJG24, YW24, ZZMG^{+}24, ZL24, Smi23b]. ocean-based [HLO^{+}21, Moh23]. oceangoing [ASGZ22]. oceanic [CFS^{+}24, LP20]. oceanica [CMMVdPV25]. Oceanographic [TK23, DFP^{+}25]. oceanography [Yu22]. Oceans [BCN23, WPvO24, BR22, BDM21, CSLGFV20, CWKW20, DKKS21, KFS^{+}22, RSFvH^{+}22, STF21, AIHR^{+}25, Smi22]. October [LK24b, Ano20v, Ano21w, Ano22w, Ano23w, Ano24w, Ano25u]. octopus [CSCGC20, DHR23, FA24, GVB^{+}25, LMC^{+}23a, SPG20, GÁSVPS^{+}25]. octopuses [BBB23]. Ocypode [YSEK22]. odd [Wil21a]. Odisha [KHMM20]. OECD [LCC20, SLGAG24]. OECMs [EAA^{+}22, MBDG23]. Off [TMDN23, AML^{+}21, BH23a, CHEC21, Dun24, HDAPG^{+}24, PdBA^{+}25, SOP^{+}22, Tod20, TBAR^{+}25, WWG^{+}23, WWG^{+}25]. off-the-shelf [Dun24]. offer [HSH^{+}21, NCF25]. officers [KZR22, Tan22]. official [MZVC25, PKB^{+}21, Wat25]. offs [BCN23, GASLCP20, MB22, PKMT23, ZGY21]. offset [KWY^{+}22]. offsetting [MRKM23]. Offshore [Cha20b, Laj21, MCRR23, SLS25, BvdBvdFK20, BST23, BLY^{+}24, BWV25, CSLGFV20, CLL22a, CCW21, DLJ^{+}20, DWC^{+}20, DBHT22, DPSF25, FDVP24, FBC^{+}23, HRR^{+}20, HJ25, HBJ21, JFB^{+}24, JF22, JvH24, KNY20, KSO^{+}25, KABT21, LJ21, MBORGN24, OC20, ONA21, OKH^{+}24, PRM^{+}20, PKMT23, SMB^{+}25, SBT21, Smy24, SZBA20, SVKP21, TYS24, Wan24b, WOR^{+}25, WNMS^{+}24, YK20, vRB24, WML25]. often [MS21]. OHS [Sha22]. Oil [Cha20b, JHHG23, MFT^{+}22, dOSTB^{+}20, LZX24, SRvNN21, SBT^{+}23, SGFM^{+}21, dSTHL^{+}24, WDG^{+}20, PKH22]. Okinawa [Tru23]. old [Gal22, OEA25, VBd^{+}21, ASM^{+}21b]. Olympics [HWMO20]. OMAF [VKL^{+}25]. Oman [ASGS23, MDLA21]. on-board [ASKI20, MRD^{+}20]. on-land [PMAVM20]. onboard [DG23, PGK22, VAMR^{+}20, dOPM^{+}20]. One [Mag23, PBG^{+}22, SCA^{+}22, TRL^{+}25, dO23, CMWT24, HS24, Liu21, OVSVPP^{+}22, PT23, APF22, Her22]. One-Size-Fits-All [Mag23]. Ong [LK24b]. ongoing [SEG^{+}22]. online [MGS21, RPY^{+}24, STIS20]. only [WFR23]. oNova [DF24]. onshore [KSS23]. ontologies [Teo23]. Open [SBM^{+}25a, WOG20, AFHHD22, AR22, Art20, OA20, SOS20, VGA25]. open-air [SOS20]. open-ocean [AFHHD22]. Opening [GEB^{+}21, BSY^{+}20, CC22, CNB^{+}23, GGL22, Tod20]. openness [Ton24]. operate [LMMT23]. Operating [BBBS22, BCJ21, DSL20, HBJ21, KZR22]. operation [DG24a, HKMS20, PWL^{+}23b, SC22]. Operational [KABT21, Yu22, BTV^{+}22, DG23, Guo20, Sas25a, TBV^{+}23, WD20, WKN^{+}25]. Operationalisation [GPG^{+}21]. operationalise [GMG^{+}24]. operationalising [GG20]. operationalization [MZ22]. operations [CZ24, LF25, MBL24, Stø21]. operator [SCB25]. operators [BHEE22, EP20, GGZRAHVS25, WW21a]. opinion [GLT23, KSB^{+}23a, QSC24, WFG22a]. opinions [CPR21, DDD22, KHLS23]. Opportunities [FLLML25, Hsi20, LWH^{+}24, LGW^{+}21, MTS^{+}24, OYECdOP24, SMF^{+}23, SSZ^{+}24, Tho25, AMLC^{+}23, BI21, SONB^{+}22, BAD21, CWB^{+}20, CMC24, CJH24, CWR20, Dav25, DF24, DS22b, GGMCMM^{+}24, JF24, KWF^{+}21, KFS^{+}22, LGP25, MG25, OGS24, QWBY21, SPFA^{+}20, SLB21, SCGC20, SMB^{+}25, SF22, SGL^{+}21, TNE^{+}24, TBPC20, WvdMG^{+}21, WGR^{+}22, WRV20]. opportunity [Dbo22, ETB25, KFP20, KMM21, MZ24, dSRKTC24, RL24, SC24, SC25, SSZ^{+}22, Sta23b]. Oppositions [GR25]. Optimal [EHLT21, XTY22, PHF20, SCSR^{+}22, SCSR^{+}23, SKE22]. optimise [CRP^{+}20]. Optimistic [SBT21]. Optimization [GSB20, BMC20, Mot20, Mot22, TBV^{+}23, YSL^{+}25]. Options [RJAT23, Tod22a, BKW^{+}24, CSRG25, DBHT22, FGPPPV21, IBC^{+}20, KGW20, OWC^{+}24, Shi20a]. orange [Squ20]. orange-spotted [Squ20]. orca [Dam24]. Orchid [KWG23]. Orcinus [Dam24]. order [BN20, MZZ24, SHZC21]. ordering [TK23]. Oregon [HCWC21, SCCW21]. Organic [SOK^{+}22]. organisation [Ros22, Per24]. organisational [SKL23]. Organisations [GML^{+}24, DMT25, GLAVL21, HWY22, KL22]. organised [BE20c]. organisms [SLO^{+}24]. Organization [EHM^{+}22, BCB20, GMRSP23, dOJCSLdSB21, SLB21, SCBdA20, BRKS25, CMSN^{+}22, EHM^{+}20, Gug22, RJAT23]. Organizational [BCB20, HSdT21, KLSG21, PTS^{+}25, RKS^{+}24, ZYTT21a, SOS^{+}21b]. organizations [CMS20, CJN25, Gon25, HMFH20, LSY20, PCCR^{+}25, Pet22, WMRB22, ZRR^{+}25]. orientation [LDAC20, VKN20]. oriented [MLVT21, PCCOS24, WNJ25]. origin [RKC^{+}24]. Origins [UPBB25]. ornamental [CMI^{+}22, PDJ22]. orphaned [PPvD^{+}24]. Oslo [Qaf24]. OSPAR [TCZ21, RCH22]. Osthagen [Smi23b]. Otaria [PBH^{+}23]. other [BDG^{+}25, EAA^{+}22, GCL21, MBDG23, SGA^{+}21, SZBA20, YG20, dS21]. otherwise [GRS^{+}23, Lov23, RCM24]. otter [NSK20]. our [CBCA25, CMWT24, JV20, LMSLMQ21, RSFvH^{+}22, Cha24b]. outbreak [ABC^{+}24, NK21]. outbreaks [HCW^{+}23]. outcome [JDF^{+}23, MGNP25, PCHD^{+}25]. outcomes [BBB25, BDD^{+}21, CT24, GABA^{+}22, GABA^{+}23, Has23, MFM21, PLWO24, SSAAB^{+}24, VPRR^{+}21]. outdated [CMNH21]. outer [dS21, dS22a]. outermost [CVCV^{+}22]. outlets [CL24]. Outlook [SK20b, FGC21, WBB^{+}25]. Output [dlFGFFR^{+}20, BG20a, BSSL^{+}20, SRSGMGG21b, TM23a, WW22, ZCCQ22a, ZT21, ZZD^{+}23]. outstanding [PLHK^{+}24, Sin21]. Ouvea [LDAC20]. Over-fishing [dMLB^{+}20]. overall [AGK22]. Overcoming [JAH^{+}24, MG25, MBG^{+}24b]. overexploitation [BRSC^{+}24]. overfishing [FTSG21, ST24]. Overlap [AVC24, JKCY25, SSMG21]. overlapping [SWH^{+}23]. Overlaps [DS22b]. overlooked [Gar24]. Overseas [DHC22, Min24]. Overview [LZRC23, BBP25, CDRS22, CL20, DG24a, Mab24, WA24, NSZ^{+}21]. own [FGK^{+}20, Ste23]. owned [AIHR^{+}25]. owner [EP20, PYVF24]. owner-operators [EP20]. ownership [EP20, HvPPV20a, KLS^{+}25, SBG^{+}25]. Oxford [Smi22, Ste25]. Oxfordshire [LL24a]. Oxon [WQ25]. oxyrinchus [DSL20, Kai21]. oyster [BABC20, CHN^{+}24, LCL20, PZZ^{+}21, SN23, WM21].

P [BRR^{+}22, LK24b]. P. [BRR^{+}22]. Pêche [Ken20]. PA [Hor25]. Pacific [Abo23, AFHM23, CDG^{+}21, DBS^{+}22b, HAD24, HLW^{+}23, Kai21, KLS^{+}25, LCMBGP22, MBB^{+}22, Per24, SOT^{+}22, TYS22, WLS^{+}25, War23, WB24, Abo23, AL20, BBC24, BBS^{+}21, Ber23, BDC^{+}21, Can25, CMWN^{+}23, CZX24, CLD^{+}24, DPG^{+}22, FSKJ23, FTM^{+}22, FIB25, FOKM23, GWÁV^{+}25, GBC21b, Gov25, GC22, HAD24, HWT23, JHHG23, JF24, Kak20, KFB^{+}20, Lou22, MZ24, MGPR^{+}25, NWS^{+}24, NNR^{+}21, OWCFDC21, OAAMS24, Pan25b, RBB^{+}24b, RROP23, RTMCVMDE22, SMYH22, SMKS22, SGRW22, SW22, STF21, WKH22, Wat25, Yad24, YS21]. packages [PDD^{+}25]. PADDD [ACA^{+}21]. Pagellus [SFGEPC22]. pair [LXW^{+}21]. Pairing [MLJ^{+}22]. Pakistan [CJK24, KCB24, KBC25]. Palau [DLJ^{+}20]. Palawan [MFG^{+}22, MFAH25]. Palestine [Qaf24]. Palgrave [AFHM23, LPO25]. Palos [HSGNMGC21]. Palos-Islas [HSGNMGC21]. Pan [DB20]. panacea [OOSG21]. Panama [GHK20]. pandemic [AEL^{+}23, ASGZ22, ASG^{+}24, BSI^{+}21, CDNS23, CRCP^{+}23, DME23, Ero22, FTM^{+}22, GPTASM23, HDH21, HCW^{+}23, IKB21, Kiz24, LKE^{+}22, MÇ23, MK24, PGK22, PAK22, RSK^{+}25, SVV24, Set23, SSB^{+}22, SRH^{+}22, Tan22, TAK^{+}22, VMM21, VDP^{+}22, VMP^{+}24, WHL22, ZTZ^{+}21, ZTW23, ZTM^{+}23, GMCMJA^{+}21]. Pandora [Tod20]. panel [PNS^{+}23, SGK21, XSCL21, Sha20]. panels [SKL^{+}20]. Pangkajene [HF20]. Papahānaumokuākea [Cha20a]. Paper [RP23a, RP23b]. paperback [Ste25]. Papua [HFD^{+}20, HKZ25, SWI^{+}22, SMKS22]. Papuan [SEG^{+}22]. paradigm [CCB^{+}21, MSZ^{+}23, SLW^{+}23]. paradise [Tea20]. paradox [DPB^{+}23, LGLL24, SAS^{+}25b]. Paralympics [HWMO20]. Parapenaeus [BDCJ22]. Paris [Moh23]. Park [COSS21, CJ21, ESD24, KJE21a, QTBL23, RP23a, RP23b, Tan23a, CLW^{+}21, GASLCP20, HNMSG21, MCRPC20, YCL22b]. Parks [Jon21, ALML23, HWM22, VAG^{+}24]. Parnaíba [dFVCGdS21]. Part [APH24, SSSL25, RB22]. Partelow [Sta23a]. partially [dSRKTC24]. participants [HACCM22, WNMS^{+}24]. participate [NMN^{+}25]. participated [SCBdA20]. participating [Cha22]. participation [BDC23, CR23, DA24, Dua22, GTMP20, HvdBH23, Kak20, MMC^{+}22, MFM21, MCR22, MMGAA^{+}21, NEF23, OSBA24, OBNK24, ORSB22, RYS^{+}21, RPY^{+}24, SAO^{+}21, SBD^{+}23, SCA^{+}20, SN23, Szy20, TWK25, WXLD21, WLQD21, dMVJdAdSM25]. participative [JvH24]. Participatory [AC21, CCG^{+}25, DHR23, Joh25, PMC24, WFK^{+}24, ASV^{+}21, BJGG^{+}24, GSGOP^{+}21, HC25, Kar22c, MGS21, MM23, OSBA24, PMT^{+}20, RA20, dSOW^{+}21]. Particularly [Cho22, RJAT23]. particulate [WZZC21]. parties [Bov22, FY24, GR25]. Partnership [CZX24, DBS^{+}22b, MBB^{+}22, SSB^{+}23, HC22, Liu22, DBH^{+}21]. partnerships [LTJ^{+}23, LLA^{+}25, LGW^{+}21, Pan25b, WJLSH21]. party [Kim21]. passage [GCV^{+}20, CDRS22, FFE^{+}24, JHL21, TCD^{+}20, ZNA^{+}23]. Passive [MGBS24, KDM^{+}20]. Past [AS23b, MCRR23, OSZ^{+}25, DL21, MGNP25, NWS^{+}24, WCN^{+}25, Liu21]. Patagonia [AHB^{+}23]. Patagonian [CHC^{+}21, SVG24]. patchwork [Wil20]. Patella [OVSVPP^{+}22]. Patellidae [OVSVPP^{+}22]. Path [TUBA22, Ber22, PMM^{+}21]. paths [DČFW23, GY23, Tod25]. pathway [ETSKA25, RMBS25, ZWXL21]. Pathways [DTBP21, AJDA24, ADB^{+}22, CMWN^{+}23, DP23, HSGS^{+}23, LWZ23, NNR^{+}21, RBB^{+}24b, RL24, TGF^{+}25]. Patrol [Laj21]. Patron [RMU^{+}22]. Patron-client [RMU^{+}22]. Pattern [APH24, TC23]. Patterns [BSK^{+}20, SLY25, BMK^{+}22, CT24, CHEC21, CLS^{+}23, FLC^{+}23, GS21, LCC20, MNS^{+}24, MKM^{+}20]. Paul [Smi20]. Paulo [PMC24]. pay [AFRFGP25, BCS^{+}24, BR22, CLL20, CLC21, CKY22, LC20a, XLLS21, XXLW24]. Payment [RRR^{+}23]. payments [JZW21, WL24a]. pays [Zhu23]. Paz [GBAECCRM25]. peace [hC24, CSHO^{+}20]. pearl [SMKS22, WHPF20]. peer [BSK^{+}20, GR20a, PGK22]. peer-review [GR20a]. pelagic [AGN^{+}20, AONF21, BCBS22, BdSL25, CC21, COJE24, CDC20a, hBDDF22, FZ25, KOBM24, MPD^{+}23, OKFN24, OSZ^{+}25, PFWM23, SBW^{+}21, TSA^{+}22]. Pelagics [hBDF24, ABM21, BGMCG24, MGS21, TGHF^{+}22]. Penghu [LTJ^{+}23, LLA^{+}25]. penguin [TD21]. Penida [YC21b]. Peninsula [HDAPG^{+}24, hC24, TDC21]. People [Hos24, AHB^{+}23, BEM^{+}20, BS22, DBBL21, EVB21, WNBHG24, YXDL24, TKC23]. peoples [BCR25, Mol24, VHS^{+}20, WF25]. Perceived [HGQJD24, LHTR23, MHC^{+}25, SMB^{+}25, CD24, NGP^{+}21, SBDC24, SKS21, TAK^{+}22, TCS^{+}20, ZWJvL20]. Perception [CDP22, FTGD^{+}20, NGA23, OCT^{+}22, ASG^{+}24, BMMF22, CCW21, FKNG25, HDAPG^{+}24, Hon21, LDL21, RPK23, SBG^{+}24, SBG^{+}25, SVP^{+}23, WFR23, dMVJdAdSM25, dOPM^{+}20]. Perceptions [BCS^{+}24, GLADGGVL23, Hos24, McC24, MTC^{+}23, OGM24, Tru23, WNBRSG22, AHENNA^{+}24, ADI^{+}25, AGK22, AMLC^{+}23, BAAH24b, Cao25, CMS^{+}25, CGGA20, CLU^{+}24, CBH^{+}24b, CBH^{+}24a, CBH^{+}25, CLB^{+}20, CSRG25, CHN^{+}24, DGI^{+}24, EVPMKP22, FSZ25, GTA^{+}20, GPM20, GRPRVG23, GCM^{+}23, HWT24, HPC^{+}23, HACCM22, HHEJB20, HvPPV20a, KWG23, KK22, KMG^{+}20, KNY22, KWMM20a, NMN^{+}25, OSZ^{+}25, OMC^{+}22, PVH^{+}21, PBS^{+}23, PLOG24, PSA^{+}25, RP23a, RP23b, RHC20, RKS^{+}24, SMLRB^{+}21, SDI^{+}20, SML^{+}25, SON21, SCM^{+}25, SCH^{+}24b, SPJB22, SAB^{+}20, SLS^{+}21, TDC21, VBd^{+}21, VDN^{+}24, VWJMAJ20, WLCG23]. perceptual [QRK^{+}23]. perfect [Wil20]. Performance [BDDE20, DAA21, DDGdS^{+}21, EAAG21, RSLR24, AE21, AACdlT23, AAW^{+}23, AGA^{+}21, BJGG^{+}24, CDG^{+}21, DMS^{+}24, DMG^{+}23, EP20, GLS^{+}21, GABA^{+}22, GBDC24, HMFH20, HMY^{+}23, HvPPV20a, Kar22a, KB20, KWMM20a, KWMM20b, LZX24, LFPBD^{+}20, MVR21a, MGK^{+}25, NSZ^{+}21, OMC^{+}22, PKR^{+}22, QTD21, SGZH^{+}21, SZBA20, SSL25, WD23, WMS^{+}25, YYM25, ZYTT21a, YW20]. performances [ZZW^{+}21]. period [HSHS24]. peripheral [HM21]. permanent [vHSSK20]. permits [BBJ^{+}22]. perpetuated [Mba22]. Persian [HGG^{+}22a, SONB^{+}22, DSB^{+}25]. persist [NAP^{+}21]. Persistence [SLGAG24, KMM24]. personal [Gel25]. perspective [ADR^{+}21, BM21, BMM^{+}24, SONB^{+}22, CSM^{+}25, CWKW20, CZX24, CKY20, CBB^{+}24, CZ24, CLCV21, Doe21, EVB21, FRM22, FMTA^{+}25, FWL^{+}22, FL24, Gel25, Guo20, HWT23, Hon20, Kak20, KKDK20, KLS22, Kow24, LLW22, LSH21, MSO23, CSdMR^{+}24, Min24, MTS^{+}24, MB21, OEV23, OC20, OZF^{+}25, OAAMS24, PYVF24, Qi22, QYD20, QUK21, RR20, Ren21a, RTMCVMDE22, SVHV20, SBM^{+}25a, SBGR24, Set23, Shi23a, SCGC20, SH25, TTA^{+}20, TD21, TK23, WDG^{+}20, War23, WD20, YZL24, YXDL24, ZTZ^{+}21, ZW21, ZWG^{+}21, ZYTT21a, dRA25, LPO25]. Perspectives [BL23a, CSGCD25, DAAP21, Pra24, AHT25, BCB20, BLS21, CWMM^{+}21, FMLMM23, GNJR24, Has23, HO25, HC21, Kas21, MGBS24, MHEW20, PZPB21, PMN^{+}22, QWBY21, STS^{+}23, Sol23b, SKW^{+}22, VJO23, VJÓK25, Wat25, WOR^{+}25, WFG22c, WL24b, HDHM^{+}20]. pertaining [DDD22]. Peru [GWAPRPRU22, GNMGWM21, JGWM^{+}23, JTR^{+}20, LASG^{+}25, dPJ21]. Peruvian [APGWTM^{+}24, CC21, DGH25, DVRK25, GPTASM23, GWAPT21, GRPE^{+}21, LSOP24, ODG25]. perverse [hBDF22]. PES [RRR^{+}23, SGA^{+}21]. Pescatourism [GDY25]. PESTLE [THS^{+}24]. Peter [NJW25]. PGIS [dSOW^{+}21]. phase [HCS21]. philanthropy [GEB^{+}21]. Philippine [SLS^{+}20, SLS^{+}21]. Philippines [ADI^{+}25, AR23a, BMB^{+}24, DRS^{+}25, GM25a, HJPL23, MLMD21, MMC^{+}22, MMK^{+}23, MNGM24, MFAH25, MH20b, MIH^{+}24, OT23, PCA21, PCSS^{+}21, WKO^{+}22, dZ20]. Phu [KK22]. Physiological [BMGDS^{+}20]. phytoplankton [SCA^{+}20]. pickers [vLS24]. Pier [QCM^{+}20]. Pierucci [LLM^{+}23]. pilchardus [EEIER25]. Pilot [RL24, GRS^{+}23, GDLC20, RKKM24, ZLD^{+}21]. Piloting [DSV^{+}25]. pilots [ZT23a]. Pingtan [SWW^{+}21]. pink [FCG25]. pinnipeds [WCT^{+}20]. pioneering [Gar22]. pipelines [KS22]. piracy [FLC^{+}23, GJY23, PGM^{+}24, Poo23]. Place [RF25, CD24, KM25, SCF^{+}22, SWW^{+}21, VJL^{+}21]. Place-based [RF25].

Placentia [MF20]. Plague [HO25]. plaice [NSK20]. Plan [DBH^{+}21, DBS^{+}22b, MBB^{+}22, SSB^{+}23, GAD^{+}24, PKB^{+}23, DL21]. planet [HJ21]. Planned [DLYM22, DPSF25, LYQM21, MLM23, TRL^{+}25, XXLW24]. Planning [DKW20, LXH^{+}22, PDB^{+}23, SV22, TBH^{+}21, WJLSH21, AHBAKH24, ACG21, AHM22, AML^{+}21, BG25, BN20, BMMF22, BJA^{+}24, BBB^{+}24, Bir20, BAAH24b, Bre25, BMB^{+}21, CPV^{+}21, CVCV^{+}22, CSQ^{+}24, CCNGS25, CHH^{+}25, CSV^{+}24, CTCD21, CHB23, DBH^{+}21, DBS^{+}22b, DSV^{+}25, DGMF^{+}20, Ehl21a, EVPMKP22, FGC21, GPA^{+}22, GGP^{+}22, GPG^{+}21, GSR21, GOH^{+}20, GCC^{+}21b, HLW^{+}23, HLLK22, JDF^{+}23, Jae20, JHHG23, JMB^{+}20, KMSN25, KvTNL20, KVL^{+}23, LPMS24, LKB^{+}23, LJ20, LJ23, LEVET21, MFA^{+}24, MG25, Mah22, MBORGN24, MSB^{+}25, OEV23, OWS20, Öİ24, PBS^{+}23, PGA^{+}20, PS21, PMC24, PKS20, RSAP22, RLS20, RDT^{+}23, RYS^{+}21, RKKM24, RHO24, SGB21, SAA^{+}21, San25, SBP24, SSOP^{+}21, SCRCI24, SMB^{+}25, SVG24, SZBA20, SSZ^{+}22, SRS^{+}25, SG20b, SSF23, TJ21b, WLS^{+}25, WMRB22, WGJ^{+}21, ZK21, ZJ22, ZBF^{+}22, dKMK21, dV21, vdBSvdV^{+}23]. planning [vTHS21, ANS^{+}20, AMK^{+}21, BRP^{+}23, COJZ^{+}21, Ehl21b, FGK^{+}20, FTW^{+}23, GSGOP^{+}21, GBCNCL^{+}21, KMTvT25, KMT^{+}21, RBB^{+}24a, SRSGMGG21a, SRSGMGG21b, WJLSH21]. Plans [SGLF^{+}23, BD20, BRP^{+}23, CBO^{+}21, IKB21, NM23]. plant [FGPPPV21, Mab24]. Plastic [PCN^{+}24, WWD23, AAA^{+}24, AWBC20, AVF^{+}23a, AFS^{+}23, BMC20, SONB^{+}22, Dau23, DBS22a, DDD22, Gug24, HLCM21, JAH^{+}24, KK22, KLK23, LASG^{+}25, MFT^{+}22, MGDJ22, MTC^{+}23, Men23, MLM23, NW20, Osm23, VFOA25, VRBBM^{+}24, WYH21, WHL22, Wan23b, WFG^{+}22b, WMS^{+}25, WNBRSG22, Wu22, XHZ^{+}21, ZMR20, ZHL^{+}23, vLS24, NW20]. Plastics [lTS25, AVF^{+}23a, DAB^{+}25, Hos24, RU20]. Plateau [TH22]. platessa [NSK20]. platform [GS21, GVH^{+}23, HCT^{+}21, WLQD21, WCC^{+}20]. platforms [Cha24b, TZM^{+}22, Wan24b]. play [Lin25]. players [MF20]. Plenty [TGM22, MJMAA21, SCD21]. Pleuronectes [NSK20]. plight [JS23]. plucked [SSH^{+}23]. plume [WABR^{+}22]. plundered [LHW21]. plural [Mba22]. PMS [VCC^{+}20]. poachers [GRA^{+}22]. poaching [BSLRR24, MJ23, TCT24]. point [MZVC25]. points [BHB^{+}24, GKO^{+}25, SAO^{+}21]. poisoning [Nal24]. Poland [MCTZ24]. Polar [Sos21a, WPvO24, Geh23, KOBS21, Liu20, FFE20, KOBS21]. POLARIS [FFE20]. polarisation [CVA22a]. Policies [AWBC20, ARP25, ABH25b, AS23a, BDVR25, CBC24, CBH^{+}24b, CBH^{+}24a, CBH^{+}25, DRCdMCJ21, hBDF22, DRS^{+}25, FRDK24, GBC^{+}20, GSB20, HTH25, HT23, HSHS24, KKAO^{+}25, KGP^{+}22, KM25, Kon20, LBWS23, LGP25, MDLA21, MVSP^{+}21, McC20, NSM25, OKNB22, PK23, PK20a, QS21, RBEA^{+}22, RXX24, San25, SCM^{+}25, Shi23b, TDC21, VDN^{+}24, WYH21, WYD^{+}22, XXMZ23, XLW24, YW20, YYL20, YYD23, YYM25, ZV20, dSdLSGB22]. Policy [ABH25b, BJ25, CBH^{+}24b, CBH^{+}25, DBHT22, DHU25, ESD24, EHM^{+}22, Fan24, FZYF23, FGC21, GZW^{+}20, HPS23, KSS23, KL22, Mal21, MHS^{+}25a, Mot22, PCCOS24, PMFR^{+}22, QTD21, RP23a, RJAT23, SCSR^{+}23, SCM^{+}25, SBJF^{+}21, SC25, TGM22, VMV^{+}24a, VAS^{+}25, WM22a, WWG^{+}25, YWYL22, ZGR^{+}25a, ZWC^{+}23, Åm21, ASPT24, AOO^{+}23a, ALML23, AMR21a, ALF^{+}22, AFS^{+}23, ADB^{+}22, AC23, BH23a, BDMW22, BBFA^{+}24, BB21, BISSBdlMM22, Bau24, BUR^{+}22, BOW^{+}22, BRKS25, BA25b, Bre25, BDM21, CA20, CMWN^{+}23, CGAF25, CBL24, CNG^{+}23, CLCV21, CMS^{+}21, DL21, DC20, DKKS21, Dbo22, De 20, DFP^{+}25, DMS^{+}24, DLL25, Ero22, ETB25, FXC^{+}22, FAMW^{+}25, FAC^{+}23, FSRC25a, FSRC25b, FFD^{+}23, GPI^{+}24, GSA23, GR20a, GJY23, Gor21, GVR20, GABA^{+}23, GMC24, Guo20, HWT24, HLCM21, HS22, HCG^{+}25, HLLK22, Hon21, Hos24, HSHS24, HHD20, HWY22, Isl25, Isr22]. policy [JC23, JKV^{+}22, JAH^{+}24, JSW22b, KL21a, KA21, KCCC^{+}21, Kao23, KFP20, KSB^{+}23a, KSB^{+}23b, KdBK25, KvTNL20, LTW23, LDC^{+}24, Lat25, LDAC20, LL25, LDC^{+}20, LXSX21, LGP25, LSC25, Lov23, LSD^{+}22, LFVCF24, Mar20a, MPH^{+}23, MAA^{+}25, MCS25b, MF23, MBG^{+}24b, Mor20, NSZ^{+}21, NH20, NCE21, NAP^{+}21, OBNK24, ÓHCÓ24, OGH24, OOGA22, PKR^{+}22, PKC23, PKH22, Pra24, QYD20, RAMEG24, RRR^{+}23, RA20, RHC20, RS24, SK20b, SDI^{+}20, Sep25, SSDQ20, SCRCI24, Smi20, Smy24, ST24, SS21b, Sta23a, SRH23, SWXM21, SCY24, SYY24, SLY25, SSL25, TYS24, TM23a, TKP23, TMM^{+}23, TMTC20, THS^{+}24, TAP^{+}23, VÖB20a, VÖB20b, VMD^{+}24, VSPMS^{+}22, VHCJdYC^{+}24, VLL^{+}24, WXLD21, WZX24, War23, WML25, WGR^{+}22, WtDD^{+}21, WWS^{+}25, WCM^{+}24, WFG22c, XL25, XXMZ23, YU20, YH20, YML^{+}21, YZL24, YCF24, YCBW21, ZLPT23, ZSZ20]. policy [ZLZ21, ZY21, ZLH21, ZTA23, APF22, BTE21, CdJK24, DMG^{+}23, Gun21, HBP^{+}21, KGW20, PV23, SPFA^{+}20, WH20, YBW20, ZSLY20, Smi23a, GAD^{+}23]. policy-making [Sep25]. Policy-oriented [PCCOS24]. policy-relevant [WML25]. policymakers [ABC^{+}24]. policymaking [DSQ^{+}22, QLU^{+}23]. polish [BN20, Bre25, CMOJ^{+}24, PKS20, RB22]. political [JL21b, Lou22, Mal21, SDL^{+}22, Tse21]. politics [Abo23, BÅ25a, DAB^{+}25, KGP^{+}20, Mar21a, SBS21, Smi20]. pollutant [VGK23]. pollutants [YZG23]. polluter [Zhu23]. Pollution [Rub24, AWBC20, ALML23, AFS^{+}23, BG20a, BMC20, CAC23, CZW24, CLD^{+}24, Dau23, DBS22a, DPB^{+}23, DDD22, FL24, Gug24, HLCM21, HKMS20, Hos24, JL21b, KK22, KSS23, Kuł22, LTJ^{+}23, LCL24, LASG^{+}25, MGDJ22, MTC^{+}23, Men23, NW20, Osm23, PCN^{+}24, Pra24, RMF^{+}21, SGB21, Sha20, Tru23, UPBB25, VÖB20a, VÖB20b, VMV^{+}24a, VFOA25, WYH21, WHL22, WRL23, Wan23b, WFG^{+}22b, WNBRSG22]. polymetallic [JACD20, VCC^{+}20, WABR^{+}22]. Polynesia [FLB^{+}21, SO20]. Pool [KL21b]. pooling [SII22]. Poor [RPY^{+}24, Alm21a, AML^{+}21, BSLRR24, PBS^{+}23, SCBdA20, YTG^{+}23]. population [CNB^{+}23, IAB^{+}20, KYMW21, KTC^{+}22, MGBS24, dMVJdAdSM25]. populations [BMF20, PEM^{+}25, PMFR^{+}22, SSH^{+}23, SO20, SLK^{+}23, YNZ22]. porcelain [LHW21]. Port [AE23, BHEE22, COJZ^{+}21, Hon20, OEV23, Osm23, ZZH21, ZSWJ23, BDMW22, Bre25, CMOJ^{+}24, Dun21a, FXC^{+}22, GAMR20, GEGRdL24, GLSVMB25, HPS23, HLLK22, HMY^{+}23, HFY^{+}24, KKDK20, KSS23, KABT21, Okş23, PGSS22, RC22, Şan20, SPMN23, WP23, WMM24, WdSRY23, ZWG^{+}21, BJH21, BCABG24, BHCL23, HNMSG21, KA24, Kar22a, LMSLMQ21, XWL^{+}20]. Port-based [ZZH21, ZSWJ23, PGSS22]. Port-Cros [HNMSG21]. port-level [SPMN23]. portion [DAHS^{+}23]. Portofino [VMC^{+}21]. portrayal [CL24]. portrayals [LDL21]. Ports [ABME22, GLMBMP25, BGLTC21, CDNS23, CZW24, CCOP25, CM23, GGL22, HMY^{+}23, LMSLMQ21, Öİ24, PYVF24, SBGR24, SOK^{+}22, SADP25, VÖB20b, VGK23, ZZH21]. Portugal [MCRPC20, ACV^{+}21, ADM^{+}25, BASM22, CSQ^{+}24, CACJR^{+}21, CLCV21, MFRQ23, RFPS20, SVP^{+}23, SCBdA20]. Portuguese [CHEC21, ZLG^{+}24]. posed [SBG^{+}24, Wen20]. Poseidon [Smi22]. poses [SBT^{+}23]. Posidonia [CMMVdPV25]. positional [dSOW^{+}21]. Positive [BHB^{+}24, YBW20]. positively [YNZ22]. possibilities [ASM^{+}21b, DMT25, SOK^{+}22]. Possibility [KLK23, IGS^{+}23]. Possible [Kao23, Bre25, CJ22, LZ24, ONA21, MCPS20]. Post [ABC^{+}25, BM21, DGS^{+}24, FFGJ22, HWC^{+}25, KFP20, LLHG23, YCBW21, ZX24, Sta25]. post-2020 [FFGJ22]. post-Brexit [DGS^{+}24, KFP20, ZX24]. post-COVID-19 [HWC^{+}25]. post-epidemic [LLHG23]. Post-harvest [ABC^{+}25, Sta25]. postharvest [GAA^{+}20]. posthuman [CTH^{+}23]. pot [HGG^{+}22a]. Potential [CDC20a, DD22, ETB25, FOKM23, GGL22, KWS20, LWZ23, WKO^{+}22, WWG^{+}23, WWG^{+}25, YK20, ZZD^{+}23, vdGD21, AS23b, BSM23, CVH^{+}20, Chr20, GVB^{+}25, GMGLV21, GMGLV22, HSH^{+}21, HKMS20, KCB24, LZRC23, MMF^{+}25, MdCLL^{+}23, ONA21, OKH^{+}24, ORdS25, PCG^{+}23, PL20b, Pra20, RSFvH^{+}22, SGB21, SHP^{+}21, Squ20, Tes20, TG25a, Tod22a, VRBBM^{+}24, VCC^{+}21, WML25, WBB^{+}25, YH20, dSOW^{+}21, EV22]. potentially [CAHR^{+}24]. potentials [WPvO24, HIP^{+}21]. poverty [LSOP24, OWS20, RMBR22, Sol23b, ZZW^{+}21]. Power [AH21, OSBA24, Sen25, APGWTM^{+}24, BHCL23, DAB^{+}25, DBHT22, Har21, HJ25, KNY20, KSS23, Mab24, MCL22, MMK^{+}22, Sta23b, Wan24a, ZRR^{+}25, vLS24]. powered [CTR24]. powerhouse [Wil21a]. Pp [Smi23b, LPO25, LL24a, Niu22, Rub24, Smi22, Sta23a]. pp232 [HH21]. PPP [YXX^{+}24]. Practical [HHD20, BDG^{+}25, CSL21, MB22, NOB22, AAA^{+}21, SG21]. Practical- [SG21]. Practice [BKW^{+}24, BPJ^{+}24, ALF^{+}22, BAVV24, CLL22a, CJ22, EFCU25, FGF^{+}24, FRCR24, GG20, GLB^{+}24, HBP^{+}21, JSW22b, Lin24, MJMAA22, Oan22, SWI^{+}22, Shi24b, SVKP21, TCZ21, TAP^{+}23, WM22a, YD22, YCF24]. practices [AIHR^{+}25, AAKI21, ABM21, BLY^{+}24, BGQL24, BSK^{+}20, CN24, CRP^{+}20, CGB^{+}25, CWKW20, DT23, DPB^{+}23, DRY22, FZL25, FTM^{+}22, FGK^{+}20, GMPT^{+}20, GMG^{+}24, HT23, KCCC^{+}21, KSO^{+}25, KMT^{+}21, MF22, MWM^{+}23a, MMGAA^{+}21, ONN^{+}20, OEA25, Pra20, SBFHD^{+}23, SPHO23, Squ21, TOAPJ21, Tse21, WH20, WSDB20, dlPdlLLC^{+}22, vdBSvdV^{+}23]. Practitioner [OWC^{+}24, GLB^{+}24]. practitioners [HZ23, SBP24]. pre [MÇ23, MLB^{+}25, BM21]. pre-pandemic [MÇ23]. pre-requisite [MLB^{+}25]. preamble [Lot23]. precarious [YS23]. Precarity [CJ23]. precaution [KGL22, Pen22]. precautionary [Hor25, MXZ23, PL24b, WH20]. precedents [hC24]. Preconditions [COJZ^{+}21]. precursor [GR20a]. predators [MLAS23]. predict [GABA^{+}22, HHEJB20, VPRR^{+}21]. predictable [BdSL25]. Predicting [FIB25]. prediction [KBB^{+}25]. predictions [BDL^{+}24, SR22]. Prefecture [Tru23]. preference [CMR21, MMW^{+}23, WLX21, WYD^{+}22]. Preferences [DSB21, CDD^{+}23, DWC^{+}20, DMK24, FRDK24, GTTM21, MM21b, POM24, Pra20, Sol23a, SSL25, WMS21, XXMZ23, YFD^{+}21, vRB24, BR22]. Preferential [MMO^{+}24]. preferred [OZF^{+}25]. Preliminary [GMCMJA^{+}21, MH20b, DDD22]. premium [SPG20]. preparation [Set23]. prepare [IGS^{+}23]. preparedness [McC24, dKMK21]. Preparing [MN23]. presence [PCD^{+}23, Hon21]. present [BMMF22, DL21, Fra23, Liu21, MCRR23, OSZ^{+}25]. present-day [Fra23]. presentist [TDP20]. presents [OGS24]. preservation [JACD20]. preserve [GPM20, KGW20]. preserved [SCZ^{+}24]. preserving [FWGM24]. presidential [TLYC20]. Press [Niu22, Smi20, Smi22, Ste25, Tan23a]. Pressing [GAGLS25]. pressure [DXS^{+}21, GABA^{+}22, Mil22, SH22]. pressures [AH23, MLJ^{+}22, PEM^{+}25]. Prevalence [HHH^{+}23]. Prevent [BBV^{+}25, MM20]. Preventing [TCT24]. prevention [CLC21, LDAC20, LC20b, LFVCF24]. previous [hC24]. Previously [TJ21a]. Price [Gor21, JG23, Cat20, CDS^{+}25, FR23, JTdH25, LT24, MMM^{+}22, May22, QL24, SPG20, SVV24, SKE22, VJ24]. Priced [EP20]. prices [GICM25, NTH^{+}21, dSTHL^{+}24]. pricing [RLRS21]. primary [RKC^{+}24]. primer [HSK^{+}23]. principal [MF23]. principal-agent [MF23]. principle [Che21b, LC20a, MXZ23, Zhu23]. principled [Dbo22]. Principles [YCF24, Dbo22, DC23, HNA25, JSW^{+}22a, Koj23, Shi23a, War20]. Prionace [DSL20, SSA^{+}23]. prior [DLJ^{+}20, Lev20]. priorities [BDM21, CLRW22, DVRK25, FWGM24, GAD^{+}23, HPPH22, LJ20, Mur21b]. prioritization [MCTN23]. prioritizing [OCCSC23]. Priority [BAAH24a, GWAPT21, SZ24, VBC^{+}20]. prisoner [Zha21]. Private [WKBS20, CRJ22, DSWB20, GEB^{+}21, RSN25, GBC21b]. private-boat [CRJ22]. privatization [CAHR^{+}24]. Privy [RROP23]. pro [CCK^{+}21, WFG^{+}22b]. pro-active [CCK^{+}21]. pro-environmental [WFG^{+}22b]. proactive [HM22a]. problem [Art20, Fen23, FDVP24, GVR20, MF23, OOGS20b, Pie21, PTCC24, TCT21]. problem/project [TCT21]. Problematizing [Lot22]. Problems [JPM^{+}23, Liu25, Gro20, HIP^{+}21, Pau21, SOLN21, AGS20, ZL20]. procedure [HFY^{+}24]. procedures [Kak20, Sch24a, BBBS22]. process [ANS^{+}20, ANR^{+}22, BD20, Bau24, BAAH24b, hC24, DBS^{+}22b, HF20, JDF^{+}23, JC23, ML22, Mab24, MF25, MPH^{+}23, MKB21, RYL25, Sas25a, SS21b, SB20, Wan23b, Wan23a, WMM24, ZK21, WJLSH21, APH24]. processed [GSG^{+}24]. processes [CdJK24, GM21, RA20, vHSSK20]. processing [CBD^{+}23, DSWB20, HRC^{+}24, PCSS^{+}21]. processors [IOHC22]. procurement [Cha24a, Kow24]. produced [BAJR24, GEB^{+}21, WXLD21]. Producer [CGGA20, DMT25, HLCM21]. product [PESO24, QZZX20, QZZ20, SLY25, TOHK20]. production [AOO^{+}23a, AED20, BABC20, CSM^{+}25, CSCGC20, EMS^{+}23, FXC^{+}22, FTM^{+}22, HCSBM^{+}23, IGO21, JY20, KCH^{+}25, LGF24, LLW22, LJ22, LSY20, MTS^{+}24, PL23, SCJ22, SHP^{+}21, SLGAG24, SOS^{+}21b, SB20, Tar25, dSTHL^{+}24, TFJR25, TTL^{+}25, VMV24b, YCL22a]. production-based [LLW22]. productive [FZL25]. Productivity [dSVT^{+}25, FWGM24, LXSX21, NHA23, TJZ^{+}21, WLY21, WW21b]. products [CMR21, CLCV21, DVRK25, LHV^{+}20, LAF^{+}24, MHT^{+}24, MLM23, PMK^{+}21, QZZ20, QL24, SPG20, Sol24, SFN^{+}21, WM21, WBB^{+}25]. professional [BGQL24]. professionals [MLE^{+}24, MMM^{+}25]. Profit [OAM20]. Profitability [CCGM21, JHL21, TCS^{+}20, BV21, EECE^{+}24, RAFLRP21]. profits [MM20, dSRKTC24]. program [AACdlT23, AGK22, IMA22, MLA^{+}24, dSTHL^{+}24, ZABI25, TCT21, WZ25]. Programme [FFGJ22]. Programmes [GKP^{+}22, OKNB22]. programming [PFWM23]. programs [AACdlT23, CTC^{+}23, EHM^{+}20, EHM^{+}22, HTH25, STIS20, VAG^{+}24, ZLD^{+}21]. Progress [FTSG21, MG22, MDJ^{+}22, RHO24, CTCD21, DAB^{+}25, EdLP^{+}21, Ehl21b, FWSS^{+}25, GMG^{+}23, GKP^{+}22, Mah22, QLZ^{+}24, RZ21b, Ren21a, Shi20b, VSV^{+}23, WSZ^{+}25, ZL20]. prohibited [CMI^{+}22]. prohibiting [VJL^{+}21]. Prohibition [Kuł22, NW20]. Project [Smi22, BHCL23, CTIH21, GBCNCL^{+}21, KMTvT25, KNY20, LJ20, QTD21, TCT21, WPvO24, BMM^{+}20, MF20]. projections [Gor21]. Projects [SSW^{+}25, BDMW22, GRSRSN^{+}24, JB25, LLSY22, MPSF20, OWS20, RBEA^{+}22, TT22, WS23, GGMCMM^{+}24]. prolifera [XLLS21]. proliferation [Rog24, SMCW22]. prominent [WNBRSG22]. promote [CSHO^{+}20, GCV^{+}21, Gir22, GMRSP23, HPH22, Liu22, NWZ^{+}21, TLYC20, TMBM^{+}25]. Promoting [ASC25, HC22, cL22, WFG^{+}22b, CN24, FRDK24, HWMO20, OCD^{+}23, Sas25a, WKO^{+}22, WP23]. promotion [JL21b, Wan24a]. proofing [CHH^{+}25]. propagation [ZXGL21]. propensity [PDD^{+}25]. Property [Run24, AELR^{+}20, RDLA^{+}20, VAS^{+}25]. Proposal [SLP^{+}22]. Proposals [MPD^{+}23, LH24]. proposed [Dun21a, HGL^{+}20, LR20, MK24, SWHT22, TSH^{+}22a, TBAR^{+}25]. propositions [OCD^{+}23]. propulsion [KDS22, THS^{+}24]. prosecution [Álv23, Pat22]. prospect [Wu22]. prospective [CHEC21, DC20, KGW24]. Prospects [PH24, TCC^{+}22, CCX23, DLL25, GCC^{+}21b, HLM^{+}20, Liu25, PDS24, SHS^{+}25, Wan23b, WSZ^{+}25, YW20, YYM25, ZL20, ZLZ21, ZMM^{+}23]. prosperity [LP20]. Protect [RJAT23, ALML23, ADB^{+}22, BAJR24, DMG^{+}23, Mag23, MH24, RMBS25, SMF^{+}23, SMCW22]. Protected [AIG^{+}20, ACA^{+}21, ÁFFSC20, AVF23b, BAAH24b, BvBN^{+}20, BEH^{+}21, BTN^{+}24, CCS23, FMLMM23, Gel25, GMG^{+}24, GABA^{+}22, GDY25, HSGNMGC21, IMB^{+}24, KJE21b, LGLL24, NJBM^{+}23, ORSB22, QTBL23, RHG^{+}25, RA20, RP23b, RCM^{+}22, VBd^{+}21, VMC^{+}21, WJLSH21, WB24, AHAB^{+}22, ASC25, ASM^{+}21b, BBB25, BDM20, BPE^{+}25, BBB^{+}24, BDG^{+}25, BD23, BCR25, BCJ21, BUM^{+}24, BWV25, BBK^{+}21b, CHH^{+}25, CMGF20, CCBB^{+}25, CLRW22, CJ22, DDAB^{+}23, Col20, DLJ^{+}20, DKW20, De 20, DB20, DSC^{+}23, Dua22, EV22, FTGD^{+}20, FWSS^{+}25, GFD21, GM25a, Gar20, GCMT20, GCL21, GBC21a, GBC21b, GMVM23, HJ21, HPPH22, HLM^{+}20, HCT^{+}21, JPM^{+}23, JL21c, KFM^{+}20, KDM^{+}20, KOCL23, Liu24, LZRC23, MM21a, MCRPC20, MG25, MSBA22, MMO^{+}24, NNM^{+}21, OSBA24, PBG23, Pha20, PdBA^{+}25, PBH^{+}23, PKMT23, RKC^{+}24, RP23a, RPY^{+}24, Rus20, SAO^{+}21]. protected [SMLRB^{+}21, SPHO23, SJ21, SDG^{+}21, SDG^{+}22, SBT^{+}23, SCC25, Sol23a, SO20, SB20, TLYC20, THT^{+}23, TBH^{+}21, TBCB25, TEA^{+}22, Tra23a, TJ21b, TCT24, VWFW22, VAG^{+}24, VLG^{+}24, WD20, WMRB22, WHAD^{+}25, YT20, YCL22b, YD22, YC21b, ZRR^{+}25, ZPX^{+}22, dS22a]. Protecting [AH23, CMWN^{+}23, HJ21, SB20, TD21, Tod22b, dS22a, AMLC^{+}23, Li23b, Pyć23, WvdMG^{+}21]. Protection [BGK^{+}22, CR23, JH24, Lia25, MSP23, RCM^{+}22, WM22a, YCL25, BNF21, BC20b, BC21, BAAH24a, BAJR24, BL23b, BS22, CJ23, CLP21, COJS^{+}22, Dav25, DSC^{+}23, FFYE21, GWÁV^{+}25, GCG^{+}21, GW21, Gul23, Guo20, GMVM23, HVS22, HMM^{+}24, IMB^{+}24, KWG23, KDHL24, KN21, LHW21, LWG23, LZL25, Lin20, Liu25, MGPR^{+}25, PT23, PC20, RHC20, RCH22, SMDM20, SAS24, SDG^{+}21, SDG^{+}22, ST24, SOS^{+}21b, SRLLG23, Tra23a, TFJR25, VTRS20, DR22, VBC^{+}20, VvSvH^{+}22, WYD^{+}22, WBPG20, YL21, YSL^{+}25, LC23]. protective [SML^{+}25]. protein [Tel23a]. provide [GCM^{+}23, WHAD^{+}25]. provided [CMMVdPV25]. provides [KFP20, dSRKTC24]. Province [HFD^{+}20, LYL^{+}22, RYS^{+}21, JTR^{+}20, NHD23, QS21, XWH^{+}22]. provinces [CQ20, RZ21b, RJ21]. Provincial [Zho23]. provision [CWKE20]. provisioning [SWH^{+}23]. Provisions [BGK^{+}22, Sch24a, ML22, WW25]. proviso [Tan24]. prowess [PRM^{+}20]. proxies [DR22]. proxy [DA24]. PSMA [HMY^{+}23]. PSSA [Cho22]. psychological [WNJ25]. psychology [YBW20, YCBW21]. Public [CKY20, CCW21, DMK24, hBDF22, FRDK24, GPM20, HPC^{+}23, HHEJB20, KNY20, KNY22, MCR22, ORdS25, VDP^{+}22, VHCJdYC^{+}24, WLX21, AACdlT23, BBV^{+}25, Bea22, CKY22, DSWB20, EV22, Fen23, GBB^{+}21, GTTM21, Hon21, Kak20, KSB^{+}23a, LDL21, MVSP^{+}21, MBB^{+}22, OKNB22, Pra24, QL24, RHC20, RS24, RPY^{+}24, STIS20, VSPMS^{+}22, WWD23, WFG22a, YHX24, vRB24]. publications [AMR21b]. Publishing [LL24a, Sos21a, Sta23a, Tan23b]. Puerto [LPMG^{+}24, MTC^{+}23, PSPR20, VGAH^{+}21]. puerulus [ASPT24]. Puget [CTC^{+}23]. pulse [KGP^{+}20, Pen22]. punishment [GM25b]. Punta [GBC21a, HNMSG21]. Punta-Coda [HNMSG21]. Purba [CDP22]. purchase [DLYM22, MLM23]. purpose [GNJR24, Laj21, WS23]. purposes [SSMG21]. Purse [RROP23, AIHR^{+}25, BLH^{+}20, Cao25, EEIER25, EMH^{+}23, HWT24, MNS^{+}24, SAS^{+}25b, TYS22]. purse-seine [MNS^{+}24]. Pursuing [OI23]. pursuit [BLL22]. purview [CJK24]. Putting [Jon21]. puzzle [RE23].

Q [CGGA20]. Q-methodology [CGGA20]. Qatar [Hon20]. Qatar-Gulf [Hon20]. Qin [Sos21a]. Qingdao [XLLS21]. QRE [ZLGX20]. Quad [Can25]. qualify [WS21]. qualitative [KDS22, MKB21, NEF23, QTD21, QUK21]. qualitatively [QLU^{+}23]. Quality [Run24, SGH22, ASV^{+}21, ÁFFSC20, BDR20, Cat20, CSGCD25, FZL25, FWSS^{+}25, Kon20, LXW^{+}21, MKAF21, RYL25, RB22, SGZH^{+}21, WLY21]. Quality-enhancing [SGH22, BDR20]. Quantification [BAP21, NSM25, OVA22]. quantify [MEG^{+}21, VAMR^{+}20]. Quantifying [ADK^{+}25, HKMS20, SRS23]. Quantitative [HSHS24, KCCC^{+}21, KDS22, MKB21, OMC^{+}22, PMT^{+}20]. quasi [RXX24]. queen [MVR21b]. Queensland [CMSH25]. quest [Per24]. question [SGA^{+}21, AM22]. questionnaire [Zhu22]. questionnaires [Öİ24]. questions [HMG^{+}21a, WLCC23]. quiet [Kar23]. quo [Zhu22, Wan23b]. Quoc [KK22]. Quota [FDF^{+}21, AGN^{+}20, BFMS21, BADO20, Cro21, HWT23, HWT24, HvPPV20a, KV20, SS22, EP20]. quotas [BXV20, HvPPV20b, SSAP24].

Race [Lat25, BCMCM21]. Racing [AR22]. radio [BMRM23]. radioactive [MXZ23, WM21]. Radionuclides [PL24a]. Radovich [LPO25]. rahui [FLB^{+}21]. Raising [FAMW^{+}25]. Raja [SWI^{+}22, WMRB22]. Rallying [Fel22]. ranching [DG20, DL22, HS23, QYD20, QS21, WXLD21, WLQD21, YW20, YWYL22]. ranchings [QTD21]. range [CMS^{+}25, CMS^{+}21]. range-expansion [CMS^{+}25]. rankings [FMGV20]. Rapa [DUYEA21]. rapid [AHAB^{+}22, GTMP20, HF20, HAsR24, MIH^{+}24, SNK^{+}20]. rapidly [MCW^{+}22b]. RAPMangroves [HAsR24]. raster [CA21]. Rate [HGG^{+}22a, PHD^{+}24, TBAR^{+}25]. Rates [MLA^{+}24, Tar25]. ratification [KCB24]. rating [IMA22]. Ray [MGPR^{+}25, BdLAdC^{+}20, BBT^{+}25, GSB24, GRPE^{+}21, PMM^{+}21, SEG^{+}22]. Rayfuse [SSSL25, Tan23b]. rays [HS24, HMG^{+}21a, LWH^{+}24, RPK23, TJS22]. RBM [ALF^{+}22]. RCEP [Cha22]. Re [BVvKL21, OGS24, Ren21a, LMMT23, Öİ24]. re-centre [LMMT23]. Re-examine [Ren21a]. Re-framing [BVvKL21]. re-planning [Öİ24]. Re-thinking [OGS24]. reached [MTN23]. reaching [SPCCvT21]. reacted [SRH^{+}22]. Ready [FAB^{+}23, GTA^{+}20, Gar20, LGW^{+}21, MWM^{+}23a, RDT^{+}23, RRS25, SSAP24]. Real [FGPPPV21, TZM^{+}22, WPWM^{+}22]. real-time [TZM^{+}22, WPWM^{+}22]. Realising [SDI^{+}20, WBB^{+}25, NSN22]. realities [AVF^{+}23a, Sas25a, SPHO23]. reality [BOBMG25, HBL^{+}21]. realize [Ren21a]. Really [Gue23]. realm [BSM23]. reasons [GWAPRPRU22]. Reassessing [SLK^{+}23]. rebuilding [AMR21a, Ber22, DSL^{+}21, DGS^{+}24]. Rebuttal [TGM22]. ReCAAP [Poo23]. recall [GVH^{+}23]. reception [Osm23]. reciprocal [KTI^{+}24, TRL^{+}25]. Reclamation [Squ21, LDC^{+}20, LYHL23, MX21, SL20, XWH^{+}22]. recognise [LX24]. Recognising [PJN24]. recognition [BMMF22, EAA^{+}22, OVA22]. recognized [GBC21a, VvSvH^{+}22]. Recommendations [GPH^{+}21, PNS^{+}23, TCF^{+}25, ATM^{+}24, ALF^{+}22, BAJR24, SONB^{+}22, CGAF25, DBBL21, Ero22, FMLMM23, GR25, GPG^{+}21, GSR21, HBP^{+}21, Isr22, LGP25, NCD23, NBB^{+}25, OBNK24, PKB^{+}23, SAS24, TYS24, VWFW22, VMV^{+}24a, WHAD^{+}25, LHV^{+}20, MT20, SBM^{+}25a]. Reconciling [Che21b, LMC^{+}23b]. Reconfiguration [VKS23]. Reconsidering [CD23]. Reconstructing [AHR24]. reconstruction [GÁPSAP^{+}23, GMMMRC^{+}25, VKS23]. reconstructions [dlPC24]. recordings [BBB23]. records [BBK21a, OWCFDC21, SSZ^{+}24, WWCS21]. recover [KYMW21]. Recovering [PPvD^{+}24]. recovery [BMGDS^{+}20, CNB^{+}23, DS22b, HWC^{+}25, IKB21, Isr22, WH20]. recreation [WFR23]. Recreational [DWC^{+}20, RSH^{+}25, SVP^{+}23, VMC^{+}21, AVPAFM^{+}22, CACJR^{+}21, CAM^{+}23, CRJ22, ES22, dlFGFFR^{+}20, GSG20, GCL21, GMCMJA^{+}21, GVH^{+}23, JBGC25, KWS20, KWF^{+}21, MRKM23, MYLV24, MJMAA21, MJMAA22, MMO^{+}24, OGS24, PHF20, QWBY21, RDD^{+}21, SCD21, SSA^{+}23, SLS25, SBT21, SMC^{+}25, TGM22, ZCCQ22a, ZCCQ22b, ZCC^{+}24]. Recreationist [SON21]. recreationists [PC20]. recruitment [CC21, CCK^{+}21, NSM25, QTPL21, UISM^{+}23]. Recycling [DN20, HP20]. Red [EMN^{+}24, HDAPG^{+}24, BMK^{+}22, DD22, GdCC^{+}25, May22, MKM^{+}20, RSLR24, SPCCvT21, VKN20, GMFD25, XL25, YSEK22, YNZ22]. Redefining [BFSP25]. Redistribution [MRKM23]. reduce [AWBC20, APGWTM^{+}24, BDG^{+}25, BAD21, CBR21, CPR21, FMWW21, HLCM21, HSH^{+}21, KAA^{+}24, KSS23, NSPRSS20, NW20, PBG^{+}22, SDG^{+}23, SMLRB^{+}21, SNP21, SSB^{+}22, SLMG^{+}21, TBAR^{+}25, WYH21]. reduced [CSM^{+}25, SWF^{+}24]. reduces [CFA^{+}24, Shi22]. Reducing [DPB^{+}23, GMM^{+}20, MGDJ22, MM20, SADP25]. Reduction [SGLF^{+}23, AAA^{+}24, AFS^{+}23, HLCM21, HRS^{+}24, MPL^{+}23, MM23, OOGK^{+}24, RQB^{+}20, Sol23b, ZZW^{+}21]. Reef [DR25, MHC^{+}25, VBd^{+}21, AKFY22, ADWP24, BBC24, CWB^{+}20, CLRW22, CWKE20, FR24, HVS22, HF20, KCT^{+}24, LRD^{+}22, Lev20, McC20, McC21, MBG^{+}24b, NTH^{+}21, OOM^{+}25, PDJ22, RBEA^{+}22, SKE22, SO20, WB24, YFD^{+}21, dSF24, DR22, Wat22]. reef-building [AKFY22]. reefs [AHN^{+}23, BDMW22, BS22, Dun21a, Dun21b, GA20, GN22, MDG^{+}23, SVF20, WWD24]. Reeling [JBGC25]. reference [JACD20, TCZ21]. refining [ZGY21]. reflect [ÓHCÓ24]. reflection [MY24]. Reflections [EFFP23, YCL25, ZX24, AAA^{+}21, XT23, ZY23, BG20b, DPP^{+}20]. reform [CMOB^{+}20, SK20b, ZPX^{+}22]. reforms [MVR21a, Sak21]. Refuge [GBC21a, GMMMRC^{+}25]. refugia [CWB^{+}20]. regard [Lin24, MS20]. regarding [CLL22a, CMS^{+}25, EV22, EVPMKP22, GLADGGVL23, Hos24, Pen22, SCGC20, SS23, Tan24]. Regency [HF20]. regime [Ber24, CC21, Fri24, Gar22, KBC25, KY20, LV23, LPMG^{+}24, OA20, Tel23b, TBH^{+}21, TN21, YK21, XWL^{+}20]. regimes [GSAF21, GRPRVG23, Kar23]. Region [BRR^{+}22, FMTA^{+}25, GBC21b, OA20, YMG23, Als20, ABRA20, ADB^{+}22, BEH^{+}21, CVCV^{+}22, CBO^{+}21, CRCP^{+}23, FKNG25, GGMCMM^{+}24, Kas21, LEVET21, PKB^{+}23, TVN23, Wat25, ZFGL20, EMN^{+}24, GAA^{+}20, KMTvT25, Kiz24, MÇ23, Şan20, SR22, Tod22b]. Regional [EHM^{+}22, FFGJ22, HDH21, HDCH21, IOHC22, LXH^{+}22, LSH21, LFLF22, Per24, RCH22, Ros22, War23, BDVR25, BLL22, BJ25, Bet25, BDD^{+}21, BMM^{+}24, BLM^{+}25, CBC24, CS21, CCCC21, Cla20, COM20, Dav25, hBDF24, EVPMKP22, dlFGFFR^{+}20, Gon25, HvT22, HMFH20, HJPL23, JLY23, JK22, LJ22, Mar21a, MCTZ24, Osm23, PCCR^{+}25, PWL^{+}23a, PWL^{+}23b, QYX21, SOLN20, SLO^{+}24, SLH22, SHFP21, SHZC21, ZLG^{+}24, EHM^{+}20, GML^{+}24]. regionalisation [VJ24]. regions [Alm21a, FMGV20, HM22c, KOBS21, Liu20, NWZ^{+}21, Sha20, SGK21, Sos21a]. register [HVN21]. register-based [HVN21]. registration [AHT25, PKR^{+}22, XL25]. registries [WOG20]. regression [FSRC25a, FSRC25b, JHL21]. regulate [ABI24, ZL24]. regulated [SKL23]. Regulating [Doe21, HSSM24, KM25, WMKG21, Che21b, LCL24, NC25]. Regulation [AKL^{+}22, GdCC^{+}25, WL20, AKFY22, CQ20, DG24b, DER^{+}21, GG20, KMM21, Kon20, LC23, LC22, NWZ^{+}21, OOGA22, RJ21, Sam22, SKE22, SDG^{+}21, SDG^{+}22, TBV^{+}23, TML^{+}20, WZX24, YL21, YZG23, ZW21, Zhu22, KA21]. regulation-centered [SDG^{+}21, SDG^{+}22]. Regulations [MK24, Álv23, AP23, BJ25, Bet25, BBT^{+}25, Bre25, CLL20, Dam24, FMCD20, HO25, JKV^{+}22, MHT^{+}24, NSPRSS20, PLHK^{+}24, SFN^{+}21, VMV^{+}24a, YU20, ML22]. regulators [MCR22]. Regulatory [ATM^{+}24, STS^{+}23, AR22, BFdSdFL^{+}25, BTV^{+}22, DBZ24, DSC^{+}23, FAC^{+}24, Lov23, OC20, Squ21, WZL25, Wu22]. Reichelt [Rub24]. Reichelt-Brushett [Rub24]. reinvented [BCR25]. Rejoinder [SCM^{+}25, CBH^{+}25]. Related [HFY^{+}24, AH23, AGM^{+}22, CDRS22, FTGD^{+}20, MN23, Pau21, PEM^{+}25, QZZ20, SAO^{+}21, SWH^{+}23, TM23a, VMJ25, WMS21]. relation [CMR21, Dev22, IOHC22]. relational [HC22, Lat25, MUR^{+}21a, PPK23]. relations [Gra21, KA21, LJ23, MMK^{+}22, Pau21, Sen25]. relationship [CLW^{+}21, DC23, PGM^{+}24, SSA^{+}23, Shi23a, Tse21, VRBBM^{+}24, WSD21, WB24, ZY23]. relationships [CCK^{+}21, MMG^{+}25, MEG^{+}21, PCA^{+}23, RMU^{+}22, Tar25, TJ21b, WP23]. relative [HGH^{+}20, IAB^{+}20]. release [CL24, LCL24, WL22a, YXDL24]. releases [Mab24]. releasing [Lia25, MXZ23]. relevance [GSB24, TMTC20]. relevant [BCBS22, Gal22, GHW^{+}22, KVL^{+}23, War20, WML25]. Reliability [KMG^{+}20]. Remaking [LLT^{+}20]. Remediation [APH24]. remedies [FBC^{+}23]. remote [ASM^{+}21b, HMG^{+}21a, HAsR24, JB25, KZR22, LWSS21, VSV^{+}23, YCL22b]. remote-controlled [KZR22]. remotely [Stø21]. removal [ADR^{+}21, WL24b]. removing [VSDR^{+}22]. Remuneration [BGQ25, Mat20]. renaissance [FLB^{+}21]. renewable [GSR21, WNMS^{+}24, vRB24]. Renewal [DD22]. Rent [AMIM21, Åm21]. Reorganising [Bot20]. Reorganization [LCL20]. Reorienting [MC25, CFM^{+}21]. reply [SCD21]. report [AJDA24, BUM^{+}24, CLS20, CS20, DRY22, DČFW23, DC23, DZZC24, GMG^{+}23, LLHG23, LSH21, MdCLL^{+}23, MH20b, QYX21, RFPS20, RBB^{+}24b, SWGZ22, SCA^{+}22, SHZC21, TZC23, Ton22a, Ton22b, TC23, Ton24, WFP25, YC21a, ZCZ20]. reporting [HBHJ20, ZLD^{+}21]. Reports [SBP24, PKB^{+}21, THF^{+}25, TMBM^{+}25, YXDL24]. representations [HPG20, NGP^{+}21]. representative [CES20]. representativeness [MPD^{+}23]. reproductive [LLLF22, MWB^{+}21]. Republic [TKC23, CJLJ24, KPN20, Kim20, Kim21, KLK23, PSPR20]. require [THF^{+}25]. requirements [BRM20, GMG^{+}24, KMJ24, MPH^{+}23, RSFvH^{+}22]. requires [AHAB^{+}22, WLW^{+}22]. requisite [MLB^{+}25]. rescue [BMRM23, CDRS22, KMSN25, Sar23, CGP23]. Research [CMC22, CRIMLS24, GWÁV^{+}25, HSHS24, Ngu23, Ren21b, RXX24, SSSL25, VCC^{+}20, WSZ^{+}25, WQ25, XHZ^{+}21, XHY22, ZHL^{+}23, ZWXL21, AN25, BBP25, BMM^{+}20, CKY22, CCG^{+}25, DVRK25, DLX24, GEB^{+}21, HZ23, HHL^{+}24, Lat25, LMT21, LSD^{+}22, MIMS^{+}20, MMG^{+}25, MH24, MM23, MZT^{+}23, OWCFDC21, PSD23, RKC^{+}24, RBB^{+}24b, Shi24b, SCA^{+}22, SRL25, STF21, TV21, TMM^{+}23, VAS^{+}25, WHL22, WRL23, WKN^{+}25, WSDB20, Yu22, ZLY23, SSSL25, Tan23b]. researchers [PBC^{+}24]. researches [AFS^{+}23]. reservations [SBT21]. Reserve [FSRC25a, GBC21b, ADB^{+}22, McC21, QTPL21, ZCC^{+}24, CD23, CD25, FSRC25b, MMSB^{+}20]. Reserves [RNGRC22, DTBP21, KAC22, KKO^{+}22, OVSVPP^{+}22, WBPG20]. Reshaping [AOO^{+}23a]. residency [NUD^{+}23]. Resident [PTS^{+}25, WMS^{+}25, XLLS21, XXMZ23, XXLW24]. residents [HP23, KWG23, NNRS24, LYQM21]. residual [TN21]. Resilience [ABME22, BSI^{+}21, FGPVPP^{+}24, HCWC21, LMSLMQ21, AJDA24, BDVR25, BKG^{+}25, CDC24, CSVC23, CBD^{+}23, FTM^{+}22, GABSD21, Gov25, Jon21, KAC22, LHPJ23, LSOP24, MKSG20, MYGM25, PGSS22, PALB20, Pyć23, Ros25, SCLM^{+}21, SPMN23, TCS^{+}20, VSPMS^{+}22, VLS^{+}23, WCL25, WHAD^{+}25, YG20, YFD^{+}21, ZLH21]. resiliency [CSGC22]. resilient [TC23, VLL^{+}24, YG20, dlPC24]. Resolution [Yu22, CMSH25, Kar23]. resolved [Sin21]. Resolving [COSS21, GVR20]. Resort [HM21]. Resounding [SEG^{+}22]. Resource [MM21b, Åm21, AML^{+}21, BDDE20, hBDDF22, DCC^{+}20, GSAF21, GRPRVG23, GABA^{+}23, Hor23, HGL^{+}20, Isr22, KMM24, Kaw22, KGW20, LR20, MJ20, MWM^{+}23b, PK23, SOP^{+}22, SCCW21, VJÓK25, VKN20, WWG^{+}23, WWG^{+}25, YMG23, ZRR^{+}25]. resource-based [VJÓK25]. Resources [BAAH24b, Smi23b, Sos21a, BH23b, BM21, CCK^{+}21, CHB^{+}20, FR23, JKV^{+}22, KN21, KTI^{+}24, LDL^{+}23, Liu20, MN21, Ren21b, SCJ22, Sep25, ST24, TNE^{+}24, THT^{+}23, WRV20, ZZMG^{+}24, ZSWJ23, GB21]. respect [Ros22]. responding [Dun21b, LC20b]. Response [CBH^{+}25, MJMAA22, MJ23, DOBE^{+}25, FL24, HGQJD24, JHHG23, Kar22b, LCL24, MZZ24, Muk21, PL24a, PKH22, PV25, SGK21, SPJB22, SHFP21]. Responses [DSWB20, BC21, BvdH23, CLB^{+}20, LEERRM^{+}21, MAK22, MBG^{+}24b, Zou23]. responsibilities [Che21b]. Responsibility [TCT21, HLCM21, LZZ22, Liu22, RFG24, SMCW22, SH25, UOO20]. Responsible [MJMAA22, SCD21, FBC^{+}23, LLJZ22, TK22]. responsiveness [DKKS21]. rest [BBK21a, BCABG24]. Restoration [GN22, DRY22, EFCU25, FSZ25, GTTM21, LRD^{+}22, QTBL23, RBEA^{+}22, SBJF^{+}21, SPCCvT21, TMBM^{+}25, WLX21, XXLW24]. restoring [Wat22]. restraint [KMM21]. Restricted [VCC^{+}21, OCCSC23]. restriction [McC21, WYD^{+}22]. restrictions [CZW24, KL21b, SLL^{+}24, SNK^{+}20, ZLG^{+}24]. restructuring [Kıl24b, Ton22a]. resulting [CWKE20, Fei20, JF22]. Results [ALF^{+}22, DAA21, RZ21b, CLL20]. Results-based [ALF^{+}22]. retention [JAK^{+}23, WFG22c]. Rethinking [AVF^{+}23a, APGWTM^{+}24, CTH^{+}23, MM21a, RU20, DCC^{+}20]. retirements [Ber23]. retrieval [ABDRV^{+}23, NMN^{+}25]. Reunion [Los23, Los24]. Reuniting [PPvD^{+}24]. reveal [AGK22, AMLC^{+}23, BC20b, PVH^{+}21, WWCS21]. revealed [GPTASM23, TSLK23]. Revealing [LMC^{+}23a]. reveals [ALLR^{+}24, EdLP^{+}21, OCZ^{+}22, SWH^{+}23]. revenue [Gor21, MGK^{+}25]. revenues [Dom23]. Reverse [BB21]. reversing [BRSC^{+}24]. Review [APH24, AFHM23, Cha20b, De 20, DSJ^{+}23, HH21, LPO25, LL24a, LK24b, Lv20, Ngu23, Niu22, NJW25, Rub24, Smi20, Smi22, Smi23a, Smi23b, SSSL25, Sos21a, Sta23a, Ste25, SHS^{+}25, Tan23a, Tan23b, WQ25, YYM25, BDM20, BOW^{+}22, BRP^{+}23, BABC20, BSK^{+}20, CMWN^{+}23, CJL^{+}23, CTDY25, CNG^{+}23, DHU24, DHU25, DGMF^{+}20, DRS^{+}25, EdLP^{+}21, GPA^{+}22, GR20a, GLSVMB25, GMVM23, HHBC20, HLLK22, Jen23, KBB^{+}25, KMSN25, LRC24, MLB^{+}25, OSBA24, OT23, ÖB22, PNS^{+}23, PCK22, PV23, QLU^{+}23, QK24, RKC^{+}24, RBEA^{+}22, SMDM20, SM23, Sen25, SRH23, SRL25, TSH^{+}22a, WvdMG^{+}21, WNDZ23, YL21, ZLY23, vdBSvdV^{+}23, BPJ^{+}24, JC23]. Reviewing [DR22]. Revillagigedo [CWB^{+}20]. revised [DMG^{+}23, KDHL24]. Revision [YCL25, HMM^{+}24]. revisited [Zha21]. Revisiting [RSN25, VSD^{+}23, Sel20]. revitalisation [CS21]. revitalization [CCCC21, HC22, LLSY22, TCT21]. revitalize [JXWL25]. rewards [PY20]. RFMO [Sch24a]. RFMOs [GML^{+}24, And24, Fis22, FOKM23, Kao23, TG25a]. Rhode [SCB25]. Ria [BASM22]. Rica [RJAT23, Teo23, PSPR20, WB24]. Rican [UPBB25].

Richard [HH21]. Richest [OYKBA22]. Rico [LPMG^{+}24, PSPR20, VGAH^{+}21]. Ridge [BMR20, KY20]. ridges [DBH^{+}22, GWÁV^{+}25, WvdMG^{+}21, CMWN^{+}23]. right [Bea22, BMK^{+}25, Gal21, MM20, NM23, SPFA^{+}20, SHAC23, TDC21, RCM^{+}22, RM21b]. Rights [FMCT^{+}21, HDHM^{+}20, Run24, AR23a, AR23b, AEJ24, BBS^{+}21, BOW^{+}22, CdLN^{+}21, CK23, DF24, FFYE21, FGF^{+}24, GE23, HLW^{+}20, Jen23, Li23b, MSP23, Mrd23, NOB22, NHA23, OAAMS24, Pyć23, Rog24, RDLA^{+}20, SCGC20, SRY24, SRLLG23, WYD^{+}22, Wan23c, Who23, WKBS20, WS23, YL21, YLSC23, vNdB25]. rights-based [CdLN^{+}21]. rigid [RQB^{+}20]. Rio [SBP^{+}25, SdACdSLdAS24, dS21]. rise [Fla21, GABSD21, Liu20, Mal21, Ora23, TG25b, dS21]. Rising [WQ25]. Risk [Cha20b, CLL^{+}22b, CL20, JF20, AVC24, AMR21a, BB21, CH25, CAC23, DDAB^{+}23, CMSH25, CGP23, DBS22a, GJY23, GDB^{+}24, GWAPRPRU22, GHK20, HHBC20, HBR21, HBR^{+}23, HVN21, HSdT21, HMY^{+}23, JL24, JSW22b, LDL21, LMT21, LTRA20, Los23, Los24, LLL23, MDLA21, MCW^{+}22b, MHN^{+}20, OKH^{+}24, PB20, RKS^{+}24, Ros25, SWF^{+}24, SLS^{+}21, TBAR^{+}25, VGS^{+}20, VDP^{+}22, VRBC^{+}21, VBH^{+}23, VHCJdYC^{+}24, WMM24, WCT^{+}20, ZCSK23, dMVJdAdSM25, SLS^{+}21]. risk-based [GDB^{+}24, VRBC^{+}21, VBH^{+}23]. Risk-sharing [JF20]. Risks [Gov25, BBWC21, CBAPM20, DS22b, FKNG25, FBC^{+}23, GCS21, LDAC20, NWS^{+}24, PL24b, SCH^{+}24b, SBT^{+}23, SLS^{+}21, TOL^{+}22, WWG^{+}23, WWG^{+}25, PY20]. River [JZW21, NHD23, QYX21, SCBdA20, WHPF20, CCP25, SCA^{+}22, DDGdS^{+}21]. RMSI [Poo23]. Road [HH21, BMH^{+}22, CMC24, SF22, TBPC20, ZWG^{+}21]. roadmap [ABH25a, DN20]. Robert [De 20]. Roberts [Lv20]. rock [EJC24, OAM20, OAM21]. rockfish [GBSC22]. rocks [Gul23]. Role [BBJ^{+}22, GA20, dlC24, ABRA20, And24, ALM^{+}21b, Bea22, Ber22, BGMCG24, CMI^{+}22, CBB^{+}24, CBL24, FFGJ22, HRMPCRRA22, HACCM22, IGO21, Kat20, KL22, LMC^{+}23a, LGF24, Lin25, LJ21, MS20, MBL24, NSN22, OOGS20a, OAM21, Pet22, PK20a, RU20, RFG24, RSN25, RC22, SvLGK24, SGL^{+}23, SML^{+}25, SCC25, SFB24, SOLN20, SSW^{+}25, SPJB22, TN21, TNY23, THS^{+}24, TAP^{+}23, UE25, VÖB20b, WMRB22, WN25, ZLH21]. Roles [DG23, LNRM23, CCP25, LLSY22, OBNK24, SWH^{+}23, ZZH21]. Romania [VTRS20]. Romanian [Tea20]. rookery [KL21b]. roots [BV21]. rose [BDCJ22]. Rosemary [SSSL25, Tan23b]. Roses [WML25]. Ross [ADB^{+}22, BBK^{+}21b]. rough [LL24b]. Route [TBV^{+}23, HCG^{+}25, RZ21a, ZLGX20, ZXGL21, BSY^{+}20, Gun21, HLLK22, RZ21a, Tod22a, WDG^{+}20]. routes [Boy21, GSB20, GGL22]. Routledge [LL24a, Ngu23, Smi23a, WQ25]. Rule [CS20, LK24b, KP24, BGK^{+}22]. rules [CC21, CLL^{+}22b, GML^{+}24, LZ24, Mot20, Mot22, NT20, Pra20]. run [BCBS22, MLVT21]. Rural [DA24, CSV^{+}24, DCC^{+}20, FTM^{+}22, Fro21, RC22]. rush [FTG^{+}23]. Russia [WFP25, AB20, EB24, Tod20, WFP25]. Russian [DD22, FFE20, KV20, Tod22a]. RVSPM [GBC21a].

S [APH24, Fla21, LPO25, Lv20, BSLRR24]. S-fisheries [BSLRR24]. S. [Sta23a]. Saccharina [HKKY22]. Safe [AFHHD22, HMY^{+}23, Ste23]. safely [NGRM^{+}22]. Safety [AKL^{+}22, KA24, Sha22, Stø21, VJL^{+}21, YBW20, BvdBvdFK20, BISSBdlMM22, CMSH25, FRM22, JG23, Kar22a, NSPRSS20, PK20a, SNS23, TSH22b, WM21, WW21a, WZX24, BAP21]. Saharan [Sen25, UOO20]. Saiko [OEA25]. sailed [KOBS21]. Sailing [LL24b, KP24]. Saint [Gir22, TJ21a]. Sainte [LTJ^{+}21]. Sakuraebi [LSY20]. Salaam [SML^{+}25]. salar [FSZ25]. Salas [CMWN^{+}23, DBH^{+}22, GWÁV^{+}25, WvdMG^{+}21]. Sale [May22, Lev20, NTH^{+}21, PSMR20]. Sales [CCX23]. salinity [RKR20]. Salish [DAHS^{+}23, FMCD20]. Salmo [FSZ25]. Salmon [Ber23, CBC24, MF20, Åm21, BVvKL21, BE20a, BBGHR23, BMH^{+}22, BHK^{+}25, CAHR^{+}24, DA24, EHLT21, FCG25, FSZ25, GVL22, HO25, KS23, Knu25, KM25, LJ22, LKE^{+}22, MTS^{+}24, Mrd23, POM24, RFG24, SZ20, SCLM^{+}21, SLMG^{+}21, STG^{+}20, TFF22, WFG22a, YK20]. salmon-farming [Mrd23]. saltfish [Ken20]. salutogenic [CHB23]. salvage [FY24, LC20a, WL24a]. salvaged [PPvD^{+}24]. salvors [FY24]. Sam [GAA^{+}20]. Samar [PCA21]. same [MAR^{+}20b, OEA25, dOPM^{+}20]. sample [HACCM22]. samples [ZTM^{+}23]. Sampling [GCC^{+}21a]. San [PB20, PCCOS24, THGR^{+}20]. sanctions [Fel22, Oan22, Ros22, Tei23, WOG20]. sanctuaries [INW20]. Sanctuary [KHMM20, MLJ^{+}22, SBG^{+}24, SBG^{+}25, Mar20a]. Sandwich [Tra23a]. Santa [QCM^{+}20]. Santay [MTC^{+}23]. Santos [PCSS^{+}21]. São [PMC24]. sapidus [ABC^{+}24, BLM^{+}25]. saratan [YSEK22]. Sardina [EEIER25]. sardine [JTdH25, MIH^{+}24, WWG^{+}23, WWG^{+}25]. sargassum [LPMG^{+}24, HGQJD24]. satisfaction [CLW^{+}21, JBGC25, LSY20, PSPR20, QWBY21]. satisfactory [HWT23]. Satisfying [UISM^{+}23]. Saving [PT23, JV20]. say [BBBS22]. Sbragaglia [MJMAA22]. Scalar [Abo23, RBB^{+}24a]. Scale [CMC^{+}23, AOS^{+}24, AP24, ÁCT20, ACV^{+}21, AFMdSdTdCC22, Art20, ASGZ22, AH21, BPA^{+}22, BBB25, BSLRR24, BMB^{+}24, BDM20, BJGG^{+}24, BBS^{+}21, BCG^{+}24, BSI^{+}21, BASM22, CJN25, CPBO20, CRIMLS24, CSM^{+}25, CAVJ^{+}20, CSV^{+}24, CCKK22, DLJ^{+}20, DAN^{+}23, DDGdS^{+}21, EBB^{+}25, FKNG25, FRCR24, FS22, GLADGGVL23, GRSRSN^{+}24, GSG^{+}24, GGZRAHVS25, GSB24, GPTASM23, GMRSP23, GNMGWM21, GMCMJA^{+}21, GFL^{+}21, HLW^{+}20, HYN^{+}23, Han21, HLW^{+}23, HMG^{+}21a, HTP^{+}23, HBP^{+}21, IAS25, JAdS^{+}20, JKKT20, KMS20, KB20, KNY20, KWMM20a, KWMM20b, KJS^{+}22, LRN^{+}22, LBWS23, LJRHH21, LEERRM^{+}21, Lov23, LLS^{+}20, MMK^{+}23, ML21, MB22, MF22, MMB^{+}23, MM23, Mil22, MIH^{+}24, MMGAA^{+}21, MZT^{+}23, NH20, NAS^{+}22, OCZ^{+}22, OYKBA22, OOGS20b, OCD^{+}23, OMC^{+}22, OA20, OEA25, PLLJ24, PSPR20, PBG23, PGSFRMHH24, PL20b, PJN24, PALB20, PMC24, PKS20]. scale [RTMCVMDE22, RDLA^{+}20, RKS^{+}23, SPFA^{+}20, SIW^{+}21, SLB21, SCJ22, SBM^{+}25b, SSZ^{+}24, Sen25, SdCMN^{+}22, SS21b, SSP^{+}21a, SLO^{+}24, SPA^{+}21, SGL^{+}21, SRH23, SSP^{+}21b, TTA^{+}20, TPG^{+}23, UOO20, VTRS20, VGC^{+}22, VGAH^{+}21, WKO^{+}22, WKN^{+}25, WGBS25, WTO24, YCL22a, ZZMG^{+}24, ZLG^{+}24, ZLPT23, ZMT^{+}25, ZJ20, dSF24, dJdMCL^{+}24, dJRFVOA24, dSVT^{+}25, dlPdlLLC^{+}22, NCE21]. scaled [EFCU25]. scales [KBD^{+}20, KVL^{+}23, OPL20, SHPW21, VRBC^{+}21]. Scaling [Obu20, KMM24, LRD^{+}22, MHT^{+}24, OGM24]. Scallop [BBFA^{+}24, BDD^{+}21, YML^{+}21, YTG^{+}23]. scalloped [ALLR^{+}24, Mag23, NC25]. scandals [VM21]. SCAR [BL23a, CBB^{+}24]. scarse [CPBO20]. Scenario [IGS^{+}23, YHU22, dKMK21, GOH^{+}20, HM22b, PMC24, SMC^{+}21, TMG24]. scenarios [BCBS22, CPV^{+}21, CVCV^{+}22, CEMRMB^{+}24, EJC24, KDS22, PJN24, VPRR^{+}21]. scenery [CDP22]. schedules [vLPBK21]. scheme [BE20a, BLM^{+}25, GHK20, HCS21, dlÁPAFGO20, RRR^{+}23, WLXY22]. schemes [AO20, BMB^{+}24, BI21, Sah22, WWC23]. Scholarly [XT23, ABM21]. school [TSWC23]. schoolchildren [WFG^{+}22b]. schools [CFS^{+}24, cL22, PDS24]. sci [SLY21]. sci-tech [SLY21]. Science [CBB^{+}24, De 20, KGP^{+}20, MZ24, Pen22, SvLGK24, TBCB25, AFHHD22, AMR21b, CSV25, CMT^{+}23, CD25, CWLD25, CMS^{+}21, FLB^{+}21, GNJR24, GSA23, Geh23, GS21, GVH^{+}23, HWU^{+}23, HLCM21, HMG^{+}21b, HGG22c, Joh25, KDH^{+}21, Kra22, LNRM23, LGW^{+}21, MFM21, MH24, MBG^{+}24b, OSBA24, OKNB22, OBNK24, PKH22, PZPB21, PCHD^{+}25, Ren21b, SBM^{+}25a, SBM^{+}25b, Shi24b, STKC22, SB20, TSLK23, TMM^{+}23, VLL^{+}24, WSD21, WP23, WGR^{+}22, WPvO24, WFG^{+}22b, vSHHT20, Pol23, PCHD^{+}25]. science-based [CMT^{+}23, HGG22c, STKC22]. science-driven [CD25]. science-policy [GSA23, MBG^{+}24b, PKH22]. Science. [LK24b]. Scientific [DFP^{+}25, MIMS^{+}20, RFPS20, SLMG^{+}21, AGM^{+}22, BD20, CBR21, DGMF^{+}20, GDLC20, GCG^{+}21, HTP^{+}23, Los24, MH24, MKSG20, PCCR^{+}25, SCCW21, SG21, WRL23, WSDB20, Yu22, CGHV20, CMD^{+}20, GNJR24]. scientists [DMG^{+}23, PDD^{+}25]. scientometrics [SRL25]. scope [AOS^{+}24]. scores [KMJ24]. Scoring [Fox22]. Scotia [CA20, DF24, FRLMW25]. Scotian [TFF22, KdBK25]. Scotland [Cro21, HWM^{+}25]. Scottish [OC20]. scrappage [WWCS21]. script [PHP25, TCT24]. Scrutinizing [LMR25]. Scuba [AHN^{+}23, LM24, MMO^{+}24, NSN22, FR24]. SDG [CMOB^{+}20, CCOP25, Gul20, Liu23, Tho25]. SDG14b [SPFA^{+}20]. SDGs [LL25, Obu20]. SEA [HM22a, ASM^{+}21b, BSY^{+}20, BMM^{+}24, BK23, BBK^{+}21b, CMS^{+}25, Cho22, DMS^{+}24, EMN^{+}24, GÁPSAP^{+}23, Gul23, Gun21, HvdBH23, HLLK22, KMTvT25, Kiz24, LPMS24, LYHL23, LXH^{+}22, LNRM23, MÇ23, Ngu23, Niu22, RZ21a, RJAT23, RCH22, Şan20, SR22, TBH^{+}21, Tod22a, WDG^{+}20, ATM^{+}24, AB20, ADM^{+}25, AR22, BEM^{+}20, Bau20, BHKJT23, Boy21, BE20c, BMK^{+}25, CSMMC^{+}25, CHEC21, CDD^{+}23, CBCA25, CAM^{+}23, CLS20, CJ23, CKY20, CvHF^{+}24, CSHO^{+}20, CDC20b, CL20, De 21, Ege23, EAA^{+}21, EHM^{+}20, EHM^{+}22, FMWW21, GDB^{+}24, GVR20, Gue24, HFD^{+}20, Har21, HSSM24, HVN21, HZC^{+}24, Hon20, HSHS24, Jen21, JACD20, KAA^{+}24, Kak20, KS22, LMSLMQ21, LDC^{+}20, Li23a, Lia25, Lin24, LSH21, MXZ23, Mal21, MMB^{+}23, Mat20, MK24, MSGAR20, OPL20, Ora23, OKR^{+}20, ÖB22, Pan25a, PL24b, PLHK^{+}24, PMR23, Pra20, PBH^{+}23]. sea [PSG25, QF20, RE21, RE23, RFPS20, Rog24, Sar23, SVF20, SVHV20, SSA^{+}23, SCBB^{+}22, Set23, Shi23a, SAS24, SBW^{+}21, SC24, SC25, SPNP24, SZBA20, SLO^{+}24, SSZ^{+}22, SRS^{+}25, SRY24, SRLLG23, TZC23, VJL^{+}21, VPB^{+}24, WNJ25, WSZ^{+}25, WLCC23, Wil20, Wil21a, Wil22, WDC^{+}20, WSDB20, YWY20, YNZ22, YLSC23, YBW20, ZL20, ZTM^{+}23, dOPM^{+}20, vdGD21, KGW20, AP24, ABH25b, ATF^{+}21, ADB^{+}22, BSBD21, BRSC^{+}24, BDC23, BvdBvdFK20, BFMS21, BM21, BMM^{+}24, BLM^{+}25, BLY^{+}24, CRP^{+}20, CGB^{+}25, CCBB^{+}25, CMC^{+}23, CHB^{+}20, CDY24, COYS24, CZX24, CWKE20, DGG22, DFC^{+}20, DDF^{+}20, DUYEA21, DSD^{+}24, DAHS^{+}23, DD22, EEIER25, EECE^{+}24, EVPMKP22, ES22, Fan24, FGPPGG22, FGPPGG23, FMCD20, GMFD25, GC22, Guo20, HJPL23, HWT24, HJ25, Hsi20, IMB^{+}24, KGL22, Koj23, LSH21, LEVET21, MCS^{+}25a, MCRR23]. Sea [MHF^{+}25, Ngu24, NGT^{+}23, OSZ^{+}25, PKH22, PL20a, Pie21, PS21, PWL^{+}23b, PKMT23, RCM24, RKS^{+}23, SDL^{+}22, SPN^{+}21, SV22, SZBA20, SSZ^{+}22, SRS^{+}25, SW22, SHZC21, Tes20, TH22, VJ24, VCC^{+}21, VKS23, VPB^{+}24, VGK23, XL25, YSEK22, YNZ22, Zha21, ZFGL20, ZLGX20, Zou23, dV21, vHSSK20, Sas25a, HH21, Smi23b, Tan23a]. sea-based [SVF20]. sea-level [Mal21]. sea-use [LDC^{+}20]. sea-uses [SZBA20]. seabass [LFPBD^{+}20]. Seabed [AR22, BHKJT23, CM25, GSM20, Sin21, AGM^{+}22, BRM20, CGHV20, CDC20a, Cra20, Fei20, GG20, HSK^{+}23, Jae20, Mal21, MCR22, PSTW21, SDG^{+}23, SSMG21, Squ21, TML^{+}20, TCD^{+}20, VvSvH^{+}22, WRV20]. seabird [AVC24, BKW^{+}24, PdBA^{+}25]. seabirds [CHC^{+}21, GDB^{+}24, LTRA20, OKH^{+}24]. seabream [LFPBD^{+}20, LT24]. Seafarer [BBK21a, Dev22, TAK^{+}22, DG24a, DG24b, Min24, NEF23, PGK22, Tan22, ZSWJ23]. Seafarers [BCABG24, VMM21, ASKI20, DG23, Gue23, HVN21, LL24b, PK20a, Sam22, SHZ^{+}21, TNY23, YTÇ22, YBW20, ZWJvL20, ZZH21, ZTW23, ZTM^{+}23]. seafloor [JG20]. Seafood [BI21, EMS^{+}23, FRLMW25, IMA22, MFF^{+}23, AOS^{+}24, AKF^{+}25, CN24, CSM^{+}25, CMNH21, CGAF25, CMWT24, CBD^{+}23, CSH23, CM23, DVRK25, DLYM22, EB23, FGC^{+}24, FFYE21, FGL^{+}21, FG25, GM21, GPM23, HPG20, HT23, HRC^{+}24, HWMO20, KSÖB25, LGF24, LJ22, MAA^{+}24, NOB22, NGA23, PMK^{+}21, RL24, SFB24, SMCW22, SVKP21, SAO^{+}24, TOHK20, VMJ25, VMP^{+}24, VVR^{+}22, WKBS20, WEF^{+}21, WMS21]. seagrass [BMMF22, BAJR24, CMMVdPV25, EFCU25, GABA^{+}23, KWY^{+}22, SML^{+}25, SGA^{+}21, SWH^{+}23, SLH22]. seagrass-data [SWH^{+}23]. seagrasses [HTP^{+}23]. Seal [EVB21, KL21b, SLK^{+}23]. seals [PBH^{+}23, WJB^{+}20]. Seamount [BMR20, DGIS^{+}22]. seamounts [ALLR^{+}24, BMR20]. Seaport [SKDK23, OCD^{+}20]. seaports [CMOJ^{+}24]. search [BMRM23, CDRS22, KMSN25, CGP23]. Searching [FS21]. Seas [BAAH24a, FFGJ22, HCG^{+}25, TCZ21, AFHHD22, BMR20, CMD^{+}20, DBH^{+}22, GBM^{+}25, GR20b, Gul23, Lam21, LX24, LL24b, MCPS20, MLAS23, MSGAR20, PCA^{+}23, Ros22, WvdMG^{+}21, WCT^{+}20, vdGD21, HGLR22, LTRA20, PKB^{+}23, RKKM24, SSMG21, WQ25, Smi20]. Seascape [EABA^{+}22, AL20, CACJR^{+}21, ETB25, LFLF22, SEG^{+}22]. season [MLMD21, MMM^{+}22, MMC^{+}22, MAA^{+}25, OA20, SGPF21]. Seasonal [JKKT20, DLX21]. Seasonality [SGPF21]. seatrout [SKS21]. seawater [SCA^{+}20]. Seaweed [Chr20, FAB^{+}23, ADI^{+}25, ARP25, ALM^{+}21b, BvdBvdFK20, Ber25, CSV^{+}24, KCCC^{+}21, LWSS21, LTW23, MM22, MBC^{+}24, MPN^{+}24, MSHK25, RMF^{+}21, SLS^{+}20, SLS^{+}21, TCS^{+}20]. Seaweeds [HSN^{+}21, MMAP25, OR21, VMV24b]. seaworthiness [SM23]. Second [SSSL25, CLP21]. Secrets [BMK^{+}25]. section [EAAG21]. sector [AHENNA^{+}24, AOS^{+}24, AEL^{+}23, AVPAFM^{+}22, ASG^{+}24, BADO20, CBD^{+}23, CM23, DT23, DRCdMCJ21, DRS^{+}25, Fen23, dHJTGO21, HS20, HvdBH23, KMM21, LEVET21, MCR22, MGK^{+}25, MMGAA^{+}21, NVP^{+}23, OCD^{+}20, PCSS^{+}21, RSK^{+}25, SYA25, SN23, TTL^{+}25, VMP^{+}24, VWJMAJ20, WLXY22, dV21, vLM22, vRB24]. sectorial [FGC^{+}24]. sectors [BGQ25, WM22b]. secure [GWAPRPRU22, dlPdlLLC^{+}22]. Securing [CCX23, FGL^{+}21, NH20]. security [Als20, AHR24, Can25, CN23, CTIH21, Dav25, hBDF24, DG20, DL22, GPI^{+}24, GSB24, Gov25, HS22, Hon20, HFY^{+}24, Jen21, JF24, MF22, MM21b, OYECdOP24, OOGK^{+}24, OGH24, RBA23, SNK^{+}20, SHZC21, Tel23a, TFH^{+}23, TRN^{+}24, TGHF^{+}22, TCC^{+}22, WCZS21, YH20, ZWC^{+}23, Zou23, Lv20, Ste25]. Sediment [RB22, DPSF25]. see [BHEE22]. seed [ZWC^{+}23]. Seeing [Ree20]. Seeking [Ske24, SRLLG23]. segmentation [PFM20]. segments [HRS^{+}24]. Segura [TPAC24]. seine [BLH^{+}20, Cao25, EMH^{+}23, HWT24, MNS^{+}24, NSK20, SII22, SAS^{+}25b, TYS22]. Seiner [RROP23]. seiners [AIHR^{+}25, EEIER25]. Seizing [SC24, SC25]. seizures [GÁPSAP^{+}23]. selectable [SKE22]. Selected [Smi22, NNR^{+}21, dKMK21]. selecting [KGW24, MKB21]. selection [CA21, CSLGFV20, CMWZ24, HPS23, JKV^{+}22, MMO^{+}24, vTHS21]. selections [WL24b]. selective [MKM^{+}20]. Selectivity [BRSC^{+}24, GMM^{+}20]. Selectivity-based [BRSC^{+}24]. self [ASC25, LSY20, MMAP25, SLB21]. self-assessment [ASC25]. self-determination [MMAP25]. self-governance [LSY20, SLB21]. self-organization [SLB21]. selfishness [AAA^{+}21]. Senegal [CDNS23, hBDF22, hBDDF22, KA21, NSZ^{+}21, SPHO23]. sensation [SRLLG23]. sense [CD24]. sensing [ASM^{+}21b, HAsR24, LWSS21]. Sensitive [Cho22, RJAT23]. sensitivity [KMJ24]. sensors [TZM^{+}22]. sentiment [QL24, ZLWY22]. sentinels [WtDD^{+}21]. separation [GHK20, RQB^{+}20]. September [Ano20w, Ano21x, Ano22x, Ano23x, Ano24x, Ano25v]. sequence [LR20]. sequential [CVA22a, GCB^{+}23, NGP^{+}21]. sequestration [TMTC20]. Sergia [LSY20]. series [APH24, SFGEPC22, SSSL25]. serious [Kuł22]. service [HPPH22, SL20, YCBW21]. Services [RRR^{+}23, CMMVdPV25, CDD^{+}23, CWKE20, DJC^{+}23, DA23, DLX24, FRVB22, GPG^{+}21, HPC^{+}23, JZW21, LW22, LRC24, MKAF21, MFC^{+}24, MWWS21, POM24, QK24, RFG24, RGH^{+}22, SAO^{+}21, SAC25, SMC^{+}25, ZZH21, ZCG^{+}23, dSOW^{+}21, vSHHT20, vTHS21]. serving [IGS^{+}23]. session [MTN23]. set [LXW^{+}21]. setback [vRB24]. setbacks [DAB^{+}25]. sets [dSRKTC24]. Setting [TML^{+}20]. settings [DJC^{+}23]. Settlement [Shi20a, HC22, Li23b, Pan25a, ZX24]. Settlements [PK20b]. seven [ZV20]. severe [DDAB^{+}23, SZ24]. severity [BMK^{+}22]. Sexual [PK20a]. Seychelles [BC20a, BCN23, BAVV24, COSS21, CJ21, DFTK20, HH25, Kıl24b]. SFPA [KA21]. SFPAs [THP^{+}22b]. shad [INW20]. shadow [JL21c]. Shake [TMDN23]. Shandong [QS21]. Shanghai [ZCSK23]. Shanshui [LLA^{+}25]. Shanyuan [HM21]. shape [BDMW22, RMU^{+}22]. shaped [SWF^{+}24]. shapes [SLB21]. shaping [UE25]. share [MCKZ21, STIS20, SLSV25]. Shared [MZ24, BM21, BLM24, CDG^{+}21, Fox22, Han20, Wan23a, XT23, ZCZ20]. Sharing [BL22, CMS20, SII22, BFMS21, BTE21, GSB20, JF20, SLSV25, SCZ^{+}24, WLCC23, dlC24]. Shark [Jon21, Los23, MCW^{+}22a, MGPR^{+}25, PMM^{+}21, BDTD25, BMF20, BdLAdC^{+}20, BBT^{+}25, BL22, BBHN20, CDC^{+}25, CLB^{+}20, CRJ22, CMSH25, DSL20, GBAECCRM25, GMGLV21, GMGLV22, GKO^{+}25, GBDC24, LDL21, Los24, MAMCHP^{+}20, MT20, NUD^{+}23, OT23, OAdJC^{+}25, PGGA22, SMHPMO^{+}21, SBG^{+}24, SBG^{+}25, SBDC24, SMC^{+}21, SAB^{+}20, TMM^{+}23, HACCM22]. shark-bite [SAB^{+}20]. shark-diving [GMGLV21, GMGLV22]. shark-human [SMC^{+}21]. shark-marine [SBDC24]. Sharks [LM24, PMFR^{+}22, RM21a, Alm21a, AMLC^{+}23, ALLR^{+}24, BDTD25, BBHN20, CLB^{+}23, GBB^{+}21, HMG^{+}21a, HACCM22, KL22, LDL21, LWH^{+}24, MH20b, NGP^{+}21, NC25, SSA^{+}23, SBG^{+}24, Sel20, SON21, TJS22]. shelf [CSGCD25, Dun24, GdCC^{+}25, KS22, MFRQ23, WCF^{+}21, dS21, dS22a, AML^{+}21, LFLF22]. shelled [AOO^{+}23b]. shellfish [BSLRR24, BHB^{+}24, CCP25, CWKE20, Gir22, JXWL25, LDL^{+}23, MR24, MM22, MWWS21, MFF^{+}23]. Shellfisherie [WWS^{+}25]. shift [MCKZ21, MSZ^{+}23, VDOE^{+}24]. shift-share [MCKZ21]. shifted [Tel23b]. Shifting [TOAPJ21, YFD^{+}21, CCB^{+}21, CD24, GTMP20, KSB^{+}23a, CSdMR^{+}24, VLS^{+}23]. shifts [CC21, CSGC22, NGA23, VGC^{+}22]. Shih [LK24b]. Shih-Ming [LK24b]. shimmering [Gul23]. Ship [BBK21a, COJS^{+}22, DN20, MCW^{+}22b, WWCS21, BVT^{+}21, CLL^{+}22b, Dun24, Gun21, HSH^{+}21, HP20, KDH^{+}21, KDS22, KMM21, LMT21, LCL24, MÇ23, MTN23, NK21, NNR^{+}21, Öİ24, PKR^{+}22, PYVF24, RZ21a, SNP21, Ste23, TBAR^{+}25, XL25, GHK20]. ship-derived [Öİ24]. ship-source [KDH^{+}21]. shipbuilding [Lee24]. shipowner [XLW24]. Shipping [Bau24, Tod22b, Wan23c, ZD20b, AHN25, ATM^{+}24, BH23a, BHEE22, BTV^{+}22, Cha22, Che21b, CDS^{+}25, DG24b, FLLML25, Gun21, GGL22, GCV^{+}20, HRS^{+}24, JHL21, JD20, KKDK20, KLS22, Kiz24, Kon20, KABT21, LDC^{+}24, LrS22, MF23, MCTN23, NCD23, NNR^{+}21, PZPB21, QLZ^{+}24, RZ21a, RSN25, SGB21, Sar23, SSK23, SPN^{+}21, Tod25, Ton22a, TC23, Ton24, THS^{+}24, VGK23, WDG^{+}20, WA24, WCT^{+}20, WLXY22, XSCL21, ZWJvL20, ZXGL21, ZLWY22, ZLY23, ZNA^{+}23, vLM22]. Shiprider [WZ25]. Ships [Laj21, ATM^{+}24, ASKI20, AKL^{+}22, CCX23, CAC23, DG23, DG24a, HVN21, KA24, MNHS25, TMG24, VMV^{+}24a, WZX24, XZ23, CCX23, DBZ24]. Shipyard [CA21]. Shoal [FMLMM23]. Shocks [SMKS22, CDC24, Gor21, SPMN23, SAB^{+}24]. shopping [NT20]. shore [BHCL23, FHBS23, MTN23, Wan24a]. shore-based [FHBS23]. Shoreline [BD20, BS22, CTC^{+}23]. shores [TMAS24]. shorescapes [BvdH23]. Short [AHN25, ATM^{+}24, LK24a, MNS^{+}24, PHD^{+}24, UISM^{+}23]. short-lived [UISM^{+}23]. short-term [LK24a]. shortcomings [SS23]. Shortfalls [MGPR^{+}25]. shortfin [DSL20, Kai21]. Should [ABI24, CMOJ^{+}24]. Shouting [MF25]. shows [HLCM21]. Shrimp [AFMdSdTdCC22, BDCJ22, CPBO20, DME23, GABSD21, HCT^{+}21, KTI^{+}24, LFLF22, MVR21a, May22, PAK22, TRN^{+}24, TCS^{+}20]. shrimp-tilapia-seaweed [TCS^{+}20]. Sian [MMSB^{+}20]. Sicilian [AVC24, DPB^{+}23]. side [GLT23]. side-effects [GLT23]. sides [dOPM^{+}20]. sideways [SH24a]. SIDS [CTIH21]. Sierra [OOGS20a, OOGS20b, OOSG21]. signals [WKO^{+}22]. Signe [Niu22]. significance [Sas25a, Wan23a]. Significant [MFA^{+}24, BMR20, DSV^{+}25]. Silk [HH21, ZWG^{+}21, SF22]. similarities [BMB^{+}21]. simmering [TMDN23]. SIMNORAT [GBCNCL^{+}21]. Simulating [RHST21]. simulation [CDHY22, JHL21, PBT^{+}20, ZLGX20]. simulation-based [PBT^{+}20]. since [GMVM23, PL20a, SWXM21]. sing [WPvO24]. Singapore [CJCS^{+}22, ZSLY20]. single [AAA^{+}24, AWBC20, Fra23]. single-use [AAA^{+}24, AWBC20]. sink [HZ23]. sinks [WL22b]. Sino [Yad24]. site [CA21, JKV^{+}22, vTHS21, CSLGFV20]. siting [JB25]. situation [QLZ^{+}24, CLS20]. situational [DAN^{+}23]. situations [CPBO20, RE21]. Size [Mag23, OCZ^{+}22, SKE22, TPG^{+}23, YU20, GMM^{+}20]. sized [Bre25]. Skagerrak [NSK20]. skate [GPH^{+}21]. Skepticism [Pol23]. skill [GCP^{+}20]. skills [ASKI20]. skipjack [PWL^{+}23b]. sky [FTG^{+}23]. slavery [Gue24, VM21, WKBS20]. sleep [vLPBK21]. sleepiness [vLPBK21]. Sleuthing [KDM^{+}20]. Slow [LL25, cL22, SVF20]. Sluggish [Kar22b]. Small [ACV^{+}21, APF22, AJDA24, Art20, BMB^{+}24, CJN25, CPBO20, CMC^{+}23, CSV^{+}24, hBDDF22, hBDF24, FKNG25, FRCR24, GSG^{+}24, MF22, Moh23, SPFA^{+}20, SCB25, TTA^{+}20, TFH^{+}23, Tho25, YCL22a, AGN^{+}20, AP24, ÁCT20, AFMdSdTdCC22, AONF21, AH21, BCBS22, BPA^{+}22, BSLRR24, BDM20, BJGG^{+}24, BBS^{+}21, BCG^{+}24, BSI^{+}21, BCJ21, BASM22, BGMCG24, Bre25, CC21, COJE24, CRIMLS24, CAVJ^{+}20, CCKK22, COSS21, DAN^{+}23, DBB^{+}22, DHU24, DHU25, DDGdS^{+}21, FZ25, FGPPPV21, FMCD20, FS22, GLADGGVL23, GRSRSN^{+}24, GGZRAHVS25, GSB24, GPTASM23, GMRSP23, GNMGWM21, GMCMJA^{+}21, GFL^{+}21, HLW^{+}20, Han21, HTP^{+}23, HBP^{+}21, IAS25, JAdS^{+}20, JKKT20, KMS20, KOBM24, KB20, KWMM20a, KWMM20b, KJS^{+}22, LRN^{+}22, LBWS23, LJRHH21, LEERRM^{+}21, Lov23, LLS^{+}20, MGS21, MMK^{+}23, ML21, MMB^{+}23, MM23, Mil22, MIH^{+}24, MMGAA^{+}21, NRYG20, NSN22, NAS^{+}22]. small [OCZ^{+}22, OYKBA22, OOGS20b, OCD^{+}23, OSZ^{+}25, OMC^{+}22, OA20, OEA25, PLLJ24, PSPR20, PGSFRMHH24, PL20b, PALB20, PMC24, PKS20, PFWM23, RTMCVMDE22, RDLA^{+}20, SIW^{+}21, SLB21, SCJ22, SBM^{+}25b, SSZ^{+}24, Sen25, SdCMN^{+}22, SS21b, SSP^{+}21a, SPA^{+}21, SGL^{+}21, SRH23, SSP^{+}21b, THP22a, THT^{+}23, TGHF^{+}22, TPG^{+}23, UOO20, VGC^{+}22, VGAH^{+}21, WKO^{+}22, WKN^{+}25, WGBS25, WTO24, ZLG^{+}24, ZMT^{+}25, ZJ20, dSF24, dJdMCL^{+}24, dJRFVOA24, dSVT^{+}25, dlPdlLLC^{+}22, BAVV24, CTIH21, NH20, NCE21, WQ25]. small-island [NRYG20]. small-pelagic [OSZ^{+}25]. Small-Scale [CMC^{+}23, ACV^{+}21, Art20, BMB^{+}24, CJN25, CPBO20, CSV^{+}24, FKNG25, FRCR24, GSG^{+}24, MF22, SPFA^{+}20, TTA^{+}20, YCL22a, AP24, ÁCT20, AFMdSdTdCC22, AH21, BPA^{+}22, BDM20, BJGG^{+}24, BBS^{+}21, BCG^{+}24, BSI^{+}21, BASM22, CRIMLS24, CAVJ^{+}20, CCKK22, DAN^{+}23, DDGdS^{+}21, FS22, GLADGGVL23, GRSRSN^{+}24, GGZRAHVS25, GSB24, GPTASM23, GMRSP23, GMCMJA^{+}21, GFL^{+}21, HLW^{+}20, Han21, HTP^{+}23, HBP^{+}21, IAS25, JAdS^{+}20, JKKT20, KMS20, KWMM20a, KWMM20b, KJS^{+}22, LRN^{+}22, LBWS23, LJRHH21, LEERRM^{+}21, Lov23, LLS^{+}20, MMK^{+}23, ML21, MMB^{+}23, MM23, Mil22, MIH^{+}24, MMGAA^{+}21, NAS^{+}22, OCZ^{+}22, OYKBA22, OOGS20b, OCD^{+}23, OMC^{+}22, OA20, OEA25, PLLJ24, PSPR20, PGSFRMHH24, PL20b, PALB20, PMC24, PKS20, RTMCVMDE22, RDLA^{+}20, SIW^{+}21, SLB21, SCJ22, SBM^{+}25b, SSZ^{+}24, SdCMN^{+}22, SS21b, SSP^{+}21a, SPA^{+}21, SGL^{+}21, SRH23, SSP^{+}21b, TPG^{+}23]. small-scale [UOO20, VGC^{+}22, VGAH^{+}21, WKO^{+}22, WKN^{+}25, WGBS25, WTO24, ZLG^{+}24, ZMT^{+}25, ZJ20, dSF24, dJdMCL^{+}24, dJRFVOA24, dSVT^{+}25, dlPdlLLC^{+}22, NCE21, NH20]. smart [CCNGS25, OOGK^{+}24, THP22a]. Smokescreen [LGLL24]. smuggling [ASPT24]. snakes [GDB^{+}24]. snapper [GdCC^{+}25, QTPL21, RYS^{+}21]. snow [Gor21, SH24a]. Social [CMBB^{+}25, CFA^{+}24, CHN^{+}24, FHBS23, HFD^{+}20, KFM^{+}20, LL24a, MVSP^{+}21, NGP^{+}21, PN20, RDLA^{+}20, SBP^{+}25, Sol23a, TCT21, VGC^{+}22, AE21, BMMF22, BPE^{+}25, BHK^{+}25, CH25, CJ23, CCCC21, CWLD25, CLB^{+}23, DUYEA21, Doe21, ES22, FFYE21, FRCR24, FAGS21, GPM23, GMRSP23, GABA^{+}23, GTMP20, HRR^{+}20, HHEJB20, JDF^{+}23, JMB^{+}20, dOJCSLdSB21, KBD^{+}20, KBP24, LIK^{+}24, LZZ22, Liu22, LMMT23, MR24, MCW^{+}22a, MHF^{+}25, MF20, MTC^{+}23, MYGM25, ONA21, ORSB22, PFG^{+}21, PWL^{+}23a, PLWO24, PV23, RBA23, RFG24, RF25, RC22, SOO20, SCD21, SCF^{+}22, SOLN20, SMCW22, SSW^{+}25, Squ20, SLS^{+}20, STF21, TMM^{+}23, UOO20, XXLW24, ZCG^{+}23, ZLWY22, ZCA^{+}23, ZCC^{+}24, dlPdlLLC^{+}22, PZPB21, Sol23b, WGP^{+}21]. Social-ecological [RDLA^{+}20, VGC^{+}22, FAGS21, GABA^{+}23, HRR^{+}20, MYGM25, PWL^{+}23a, PLWO24, SOO20]. societal [MLVT21, RSFvH^{+}22]. society [HRR^{+}20]. Socio [BDCJ22, PDJ22, SSA^{+}23, SCBB^{+}22, ZNA^{+}23, CVA22b, DGH25, FMLMM23, JTR^{+}20, LPMS24, PBS^{+}23, RGK^{+}20, SRR^{+}21, SDL^{+}22, TMS^{+}22, VBd^{+}21, dRA25]. Socio-cultural [SSA^{+}23]. socio-ecological [CVA22b, DGH25, JTR^{+}20, LPMS24, PBS^{+}23, SRR^{+}21, dRA25]. Socio-economic [BDCJ22, PDJ22, SCBB^{+}22, ZNA^{+}23, FMLMM23, RGK^{+}20, TMS^{+}22]. socio-environmental [VBd^{+}21]. socio-political [SDL^{+}22]. sociocultural [GG22, WNMS^{+}24]. socioecological [MHS^{+}25a, MHS^{+}25b, SVG24, VMV24b]. Socioecology [CAHR^{+}24]. Socioeconomic [BKG^{+}25, CDNS23, EEIER25, NO25, dMVJdAdSM25, BSY^{+}20, CDG^{+}21, CM23, EJC24, HWT23, JAdS^{+}20, MLVT21, MFC^{+}24, WNMS^{+}24, ZCCQ22a, dJdMCL^{+}24]. Softcover [Sta23a]. soil [SOK^{+}22]. sole [DMT25, MKM^{+}20]. solidarity [FRCR24]. Solution [hC24, GOH^{+}20, TPAC24]. Solutions [LGP25, Niu22, AHN25, AHENNA^{+}24, Ada22, AHT25, AGS20, ADR^{+}21, BN20, BBWC21, CSV25, CJ22, DMK24, DQJ23, DBZ24, EFCU25, Gug24, HHD20, KMSN25, LZ24, ORdS25, PZZ^{+}21, RSFvH^{+}22, SNP21, SBJF^{+}21, SH25, WLW^{+}22, OWC^{+}24]. solve [Fen23, SAS^{+}25b]. solved [Pie21]. solving [PVH^{+}21]. Some [Zhu23, ZY23, VMC^{+}21]. Songhees [BMB^{+}21]. Soufriere [TJ21a]. Sound [KDH^{+}21, KDM^{+}20, CTC^{+}23, SOP^{+}22]. Sounding [WPvO24]. Sounds} [WPvO24]. source [AAA^{+}24, BHdMCM^{+}21, KDH^{+}21, SGLF^{+}23]. sourced [KA24]. sources [GAD^{+}20, Gug24, KTI^{+}24, MCTN23, Pan25a, SSK23]. South [BBGG22, FMLMM23, HGLR22, MCW^{+}22a, MGPR^{+}25, SHAC23, YMG23, ADR^{+}21, CACJR^{+}21, CSHO^{+}20, Mag23, MLB^{+}25, WWG^{+}23, WWG^{+}25, WDC^{+}20, ZL20, Abo23, AN25, AP24, AKF^{+}25, AM22, BBHN20, CHB^{+}20, CDY24, CDG^{+}21, COYS24, CZX24, CKY20, CKY22, CES20, CHC^{+}21, CHN^{+}24, DFP^{+}25, DPG^{+}22, DSV^{+}25, EJC24, Fan24, FWSS^{+}25, GC22, Guo20, HWM22, HJPL23, Han21, Hsi20, KFB^{+}20, KNY20, KNY22, Kim23, KSS23, Lee24, MFT^{+}22, Mar20a, MGDJ22, Ngu24, NC25, OAM20, OAM21, PKC23, PL24a, PKB^{+}23, PBG23, PL20a, Per24, PGP22, PK20b, PWL^{+}23b, PBH^{+}23, RBTW24, RBB^{+}24b, SRvNN21, SRR^{+}21, SON21, SBDC24, dOSTB^{+}20, SBT^{+}23, SS21b, SSP^{+}21a, SW22, SHZC21, Tra23a, WLCG23, YXDL24, Zha21, ZFGL20, ZLGX20, Zou23, dOPM^{+}20]. south-east [ADR^{+}21]. South-Eastern [YMG23, WWG^{+}23, WWG^{+}25]. Southeast [MGDJ22, RSLR24, LTJ^{+}21, CMWN^{+}23, Dav25, PPvD^{+}24]. southeastern [GWÁV^{+}25, GM25a]. Southern [GÁSVPS^{+}25, Isr22, MH20b, RCM^{+}22, BMK^{+}25, CC21, DSQ^{+}22, FMCT^{+}21, cL22, Liu24, Mar21b, MJMAA21, MIH^{+}24, PBH^{+}23, PKMT23, dSRKTC24, SCD21, SMB^{+}25, SVP^{+}23, SHAC23, TDC21, TGM22, AFMdSdTdCC22, BdLAdC^{+}20, BMK^{+}25, CTH^{+}21, EEIER25, Gar20, GB21, KTI^{+}24, MMS24, NEH^{+}23, PTS^{+}25, SBFHD^{+}23, Smy24, SPA^{+}21, Tra23b, TMTC20, WS24, Wen20, YW24, ZZMG^{+}24, ZL24]. southwest [DFP^{+}25, BOOO23, BdSL25, OZF^{+}25]. Southwestern [CMI^{+}22, GDY25, ALLR^{+}24, RRR^{+}23, PCG^{+}23]. Sovereignty [Gug24, BMF20, Kow24, Mar21b, MMAP25, SW22, WD22]. Sowing [MdCLL^{+}23]. Space [LP21, APF22, Bre25, QZZX20, QZZ20, Qi24]. spaces [SVG24]. Spain [BSBD21, CRCP^{+}23, dlFGFFR^{+}20, HPC^{+}23, dlÁPAFGO20, SPG20, SBM^{+}25b, TPAC24, TGM22, ASV^{+}21, BDLO23, BISSBdlMM22, FGPPPV21, GLADGGVL23, GSR21, Gar22, GLMBMP25, HSGNMGC21, JTdH25, MJMAA21, RS24, Sol23a, Tei23, SCD21]. span [Dun23]. Spanish [AIHR^{+}25, Álv23, BLH^{+}20, CCBB^{+}25, DL21, DSL20, FGPPGG22, FGPPGG23, FR23, dHJTGO21, HRMPCRRA22, JG23, Oan22, Pat22, RAMEG24, RAFLRP21, VMP^{+}24, WOR^{+}25]. Spanish-owned [AIHR^{+}25]. Spatial [ANS^{+}20, AMK^{+}21, AGN^{+}20, AVC24, BRP^{+}23, CLS^{+}23, COJZ^{+}21, DGIS^{+}22, Ehl21b, FLC^{+}23, FGK^{+}20, GBCNCL^{+}21, GASLCP20, KMTvT25, KMT^{+}21, LDG^{+}25, LYL^{+}22, PDB^{+}23, RBB^{+}24a, SLL^{+}24, SV22, SSMG21, SRSGMGG21a, SRSGMGG21b, TZM^{+}22, TMS^{+}22, WLS^{+}25, APV^{+}20, AFMdSdTdCC22, ACG21, AHM22, BG25, BN20, BAJR24, CPV^{+}21, CVCV^{+}22, CSQ^{+}24, CCNGS25, CSLGFV20, Cat20, CTCD21, CMD^{+}20, CHB23, DBH^{+}21, DSV^{+}25, DGMF^{+}20, Ehl21a, EVPMKP22, EBB^{+}25, FTW^{+}23, FSRC25a, FSRC25b, FGC21, GPA^{+}22, GGP^{+}22, GPG^{+}21, GSR21, GAD^{+}24, GABA^{+}22, GCC^{+}21b, GGL22, HLW^{+}23, JDF^{+}23, JKV^{+}22, JMB^{+}20, KvTNL20, KVL^{+}23, LPMS24, LKB^{+}23, LJ20, LJ23, LEVET21, MFA^{+}24, MFG^{+}22, Mah22, MBORGN24, MSB^{+}25, NM23, PHF20, PBT^{+}20, PGA^{+}20, PS21, PKS20, PKMT23, RSAP22, RLS20, RDT^{+}23, RYL25, RKKM24, RDLA^{+}20, SGB21, SAA^{+}21, San25, SBP24, SZBA20, SSZ^{+}22]. spatial [SRS^{+}25, SPT^{+}20, TJ21b, VPRR^{+}21, WNM22, WGJ^{+}21, ZK21, ZJ22, ZBF^{+}22, dV21, vdBSvdV^{+}23, vdGD21, CHEC21, GSGOP^{+}21]. Spatial-temporal [LYL^{+}22]. Spatio [FARBV^{+}24, RHST21, WL22b, CGB^{+}25]. Spatio-temporal [FARBV^{+}24, RHST21, WL22b, CGB^{+}25]. Spatiotemporal [GVH^{+}23, JK22, BDL^{+}24, MNS^{+}24]. Spatiotemporally [JL24]. spawning [SGPF21]. spearfishermen [MJ23]. spearfishing [MJMAA22]. Special [VFOA25, EAAG21, Ehl21a, GAD^{+}23, LBWS23, PZPB21, SNS23]. specialization [ÁCT20]. Species [BADO20, AC21, BFdSdFL^{+}25, BdSL25, BBB23, BDG^{+}25, BdLAdC^{+}20, BCS^{+}24, BLM^{+}25, BUM^{+}24, CHEC21, COJE24, CMI^{+}22, CAHR^{+}24, DOBE^{+}25, DUYEA21, EBB^{+}25, FGC^{+}24, FCG25, FRDK24, GÁPSAP^{+}23, GMM^{+}20, GMG^{+}24, GTTM21, JSW22b, KCT^{+}24, KKAO^{+}25, Kul23, MCS^{+}25a, MN23, MWM^{+}23b, OGS25, OTK^{+}21, OVA22, PHD^{+}24, SCSR^{+}22, SCSR^{+}23, SSDQ20, SMW^{+}21, SPN^{+}21, TBPC20, WML25, WCM^{+}24]. species-at-risk [JSW22b]. Species-level [BADO20]. species-specific [BBB23, FGC^{+}24]. specific [BBB23, COJE24, FGC^{+}24, GRSRSN^{+}24, GHW^{+}22]. spectrum [PFG^{+}21]. speed [GHK20, MPL^{+}23]. Spermonde [HF20]. Sphyrna [ALLR^{+}24]. spill [JHHG23, LZX24, MFT^{+}22, PKH22, SRvNN21, dOSTB^{+}20, SBT^{+}23]. spillover [JK22]. spillovers [CZW24]. spills [SGFM^{+}21]. spin [MB22]. spin-offs [MB22]. spiny [DMS^{+}24, Doe21, PLOG24]. Spiraling [RC22]. Spitsbergen [Tod20]. Splitting [Mat20]. sponges [MSGAR20]. sponsorship [GR25]. sport [KWF^{+}21]. spot [CBR21, CSGCD25]. spotted [SKS21, Squ20]. spp [APV^{+}20, FRDK24, MH20b, YNZ22]. Spratlys [Mar21b]. spring [GS21]. Springer [APH24, NJW25, Rub24, Sta23a]. Sputc [BMM^{+}20]. square [GMM^{+}20]. squid [CEMRMB^{+}24, GWAPRPRU22, MFAH25, PLOG24, TM23a]. squid-related [TM23a]. Squids [MLAS23]. Squilla [BMGDS^{+}20]. Sri [JAH^{+}24, NEH^{+}23, AEL^{+}23, CLB^{+}20, CLB^{+}23, GRS^{+}23, Kul23, Pra20, PDJ22]. SSF [BMB^{+}24]. St [CJ21]. St. [GMRSP23, TMBM^{+}25]. Stability [JZW21, GC22, LGF24]. stable [Gue24]. stage [HHD20]. stages [Pol23, YC21b]. stagnate [LG24]. stake [Gon25]. Stakeholder [AMLC^{+}23, BOOO23, BAAH24b, BYL25, BPJ^{+}24, CdLN^{+}21, CR23, DGI^{+}24, FOH^{+}20, FMLMM23, GNJR24, HWT24, HDAPG^{+}24, JB25, KGW24, MHF^{+}25, MKSG20, OCD^{+}20, OZF^{+}25, RYS^{+}21, SBG^{+}24, SLS25, WLCG23, ASG^{+}24, BDC23, BCB20, BDM21, CLU^{+}24, CDB20, DSL^{+}21, Doe21, EVPMKP22, FTW^{+}23, GRA^{+}22, HvdBH23, HPPH22, dOJCSLdSB21, KMG^{+}20, MPN^{+}24, MHEW20, OMC^{+}22, Öİ24, POM24, RA20, RP23a, RP23b, STS^{+}23, VDN^{+}24, ZLWY22, HO25, PVH^{+}21]. Stakeholder-defined [MKSG20]. Stakeholder-government [CdLN^{+}21]. stakeholders [AHN25, AHENNA^{+}24, ABC^{+}24, CSM^{+}25, CLW^{+}21, CHN^{+}24, DSL20, FDGM24, HS20, HC21, JCMD20, MBB^{+}22, ÓHCÓ24, PMFR^{+}22, SBG^{+}25, SWH^{+}23, VJ24, XTY22, ZK21, ZXGL21, GTA^{+}20]. stalked [GRA^{+}22]. Standard [BPJ^{+}24, GMG^{+}24, JC23, BJ25, SK20a, BBBS22, CN24, KMJ24]. standards [Bet25, BGMCG24, GLS^{+}21, HPG20, RSN25, Sha22, SGA^{+}21]. start [MMF^{+}25]. State [AP23, BJH21, BCABG24, HMY^{+}23, KA24, Kar22a, RBA23, SBP^{+}25, Sam22, WL22a, XWL^{+}20, BPA^{+}22, CT24, CC23, CLL20, CLU^{+}24, COSS21, DBZ24, GBC^{+}20, Har22, JAdS^{+}20, KS22, MS21, RCT20, Şan20, SFB24, SRY24, Tod25, WdSRY23, YT20, LWZ23, Shi22, Tho25, THF^{+}25, Wil21b]. state-led [WdSRY23]. statement [Dbo22]. States [CTIH21, Ngu23, WQ25, ZGR^{+}25a, Hon20, HLO^{+}21, KMS20, Kar22a, NSN22, Puh23, TN21, WdSRY23, dS21, AOS^{+}24, BAVV24, Cha24a, CGAF25, CNB^{+}23, Col20, FG25, FBC^{+}23, HWU^{+}23, JFB^{+}24, JB25, LAF^{+}24, Moh23, MS20, NGA23, PWL^{+}23a, SBT21, SRH23, TG25b, WW21b, Wil21a, ZGR^{+}25b]. static [CMD^{+}20, DFF23, FARBV^{+}24, FWL^{+}22]. station [KGW24]. statistical [RZ21a]. Status [KSV^{+}23, MPD^{+}23, PLHK^{+}24, Tho22, WLRH^{+}20, ASM21a, BGK^{+}22, BHKJT23, BRM20, BRR^{+}22, CZW20, CES20, DFC^{+}20, DUYEA21, DLX24, FZYF23, Fox22, FMCT^{+}21, GCM^{+}23, HF20, HGH^{+}20, Kas21, Kul23, LC20a, LYHL23, MBG^{+}24a, MIH^{+}24, NGT^{+}23, OZF^{+}25, OKR^{+}20, PCG^{+}23, Pie21, Puh23, TYS24, WNDZ23, WM22b, ZGR^{+}25a, ZGR^{+}25b, Zhu22, ZMM^{+}23, GKP^{+}22, Wan23b]. statute [Shi23a]. steelhead [JKCY25]. Step [GCMT20, APF22, CHB^{+}20, COJS^{+}22, GJY23, Kay25, LC23, NCE21, SPFA^{+}20, SCA^{+}22]. steps [BDL^{+}24, LPMG^{+}24, PCCR^{+}25, YK21]. Stewards [Har21]. Stewardship [BI21, BPJ^{+}24, CN24, GDB^{+}24, JAK^{+}23, LJ21, MPH^{+}23, NRF21, SK20a, TOL^{+}22, CWLD25, GBSC22, HBCA^{+}24, KSÖB25, MMAP25, HDHM^{+}20]. still [BCBS22, HWT23, Sin21]. Stock [CCK^{+}21, BBB23, CDG^{+}21, CD24, GRSRSN^{+}24, HHH^{+}23, LT24, MZVC25, PDD^{+}25, WFK^{+}24]. Stock-recruitment-environment [CCK^{+}21]. stocks [BLM24, DPSF25, EECE^{+}24, Fox22, Han20, HGH^{+}20, Hor25, OOSG21]. stop [BDTD25]. Stopping [SGLF^{+}23]. storage [QTBL23, SAS^{+}25b]. Storm [HBR21, YSWW21]. story [MB22, NW20, OEA25]. Strait [HAsR24, Tod22b, BRKS25, COYS24, STS^{+}23]. straits [Lot22, Tod22a, Poo23]. stranded [BBBS22]. stranding [OWCFDC21, SPJB22]. Strategic [BXV20, CVCV^{+}22, Jae20, KV20, PS21, TML^{+}20, AHBAKH24, ABH25a, CFA^{+}24, FZL25, MZZ24, OEV23, PV25, WWS^{+}25, YZH22, HM22a]. Strategies [ANC^{+}25, ABME22, COJE24, DDF^{+}20, LL25, WYH21, AFMdSdTdCC22, BB21, CMS20, CDNS23, CCK^{+}21, Cha24a, CGAF25, CBAPM20, DSB21, DSL20, FAC^{+}23, GLMBMP25, GML^{+}24, HSH^{+}21, HM22b, MM21b, Nal24, Okş23, PCCOS24, Rus20, RHL20, SIW^{+}21, SADP25, SLS^{+}21, UPBB25, Wan24a, Wan24c, WCL25, WGBS25, WWS^{+}25, WNBRSG22, XTY22, ZTZ^{+}21, ZCSK23, dMVJdAdSM25]. Strategy [ANS^{+}20, ASGS23, Ber23, GKP^{+}22, KWK^{+}25, MPD^{+}23, MFRQ23, Mor20, PKB^{+}21, PDB^{+}23, VRBC^{+}21, VBH^{+}23, VJO23, BÖ24, CMWT24, CBL24, DFC^{+}20, GBM^{+}25, GW21, HJPL23, Pre21, Puh23, RKR20, SC24, SC25, TCT21, VSV^{+}23, VPRR^{+}21, WLQD21, YZG23, ZD20a, ZMZ23, ZWC^{+}23, MCW^{+}22a, OGS25]. stream [CA20, CKY22, JB25]. strengthen [dlPC24]. Strengthening [Bet25, BMM^{+}20, CWR20, GSM20, GOH^{+}20, VSDR^{+}22, HJ21, BJ25]. stressor [WGR^{+}22]. stressors [ASG^{+}24, Dun21b, MNGM24, OCT^{+}22, PCCOS24, SCLM^{+}21]. stricter [BH23a]. strike [MCW^{+}22b, TBAR^{+}25]. strikes [FMWW21]. stringency [WN25]. striped [BSBD21]. Strong [AEJ24, AR23a, AR23b, Jen23, NHA23]. Structural [NHA23, CEMRMB^{+}24, ZZD^{+}23]. structure [CQ20, SZ20, SAO^{+}24, SKL23, TMTC20, YWYL22]. structured [ÖB22]. structures [CCP25, DMK24, Gal21, MCRR23, SVF20, VTRS20, ZCA^{+}23]. struggle [Smi22, SCF^{+}22]. Stuck [DMNT20]. students [DDD22, TSWC23]. Studies [HH21, AANM^{+}22, BMMF22, CJ23, CES20, GPA^{+}22, GCC^{+}21b, JL21c, LLLF22, LAF^{+}24, SSAAB^{+}24, SRSGMGG21b, ZWJvL20]. Study [WB24, ZLGX20, Åm21, APV^{+}20, ASV^{+}21, AKF^{+}25, AED20, BvdBvdFK20, BMGDS^{+}20, BJGG^{+}24, Cao25, CDC^{+}25, CWB^{+}20, CRIMLS24, CMC^{+}23, CJK24, COYS24, CZW24, CDC24, CCCC21, CSHO^{+}20, CVA22b, dFVCGdS21, Cro21, COJZ^{+}21, CMOJ^{+}24, DRCdMCJ21, DSQ^{+}22, DXS^{+}21, DMS^{+}24, DG20, FGPPPV21, Fro21, FMWW21, FMS^{+}22, GRS^{+}23, GMFD25, GBAECCRM25, GLADGGVL23, GBDC24, GAA^{+}20, Han21, HWT24, HLW^{+}23, HRR^{+}20, HHB^{+}21, HS23, HZ23, HRS^{+}24, HVN21, Hor23, HCT^{+}21, IBC^{+}20, JAH^{+}24, JAdS^{+}20, JMB^{+}20, KMG^{+}20, KGP^{+}22, KABT21, LDC^{+}24, LDAC20, cL22, LL25, LWG23, LTJ^{+}23, LLSY22, LMR25, LJ21, LSY20, MLJ^{+}22, MBDG23, MFAH25, MLB^{+}25, MMSB^{+}20, MGF20, MBL24, NSZ^{+}21, NCD23, NHD23, NNRS24, NEH^{+}23, OCD^{+}20, Okş23, Øst24, PCA21, PKC23, PESO24, PGM^{+}24, PC20, PPK23, Qi24, RBB^{+}24a, RP23a, RP23b, RYL25, RCM24, RKKM24, RDM^{+}20, RNGRC22]. study [SMHPMO^{+}21, SVV24, SSZ^{+}24, Shi24b, SMKS22, SMC^{+}21, SKL^{+}20, SVG24, SSW^{+}25, SRH^{+}22, SSL25, TM23a, THGR^{+}20, TSH22b, TLT^{+}22, TCT21, TMT20, TPG^{+}23, UISM^{+}23, WD23, WJB^{+}20, WLQD21, WSD21, WML25, WCL25, WFG^{+}22b, WDC^{+}20, WFG22c, WLW^{+}22, Wu22, WWC23, YFD^{+}21, YS21, YTÇ22, YXX^{+}24, YC21b, ZJ20, ZCCQ22b, ZTM^{+}23, ZCSK23, ZCS^{+}23, Zhu22, dJdMCL^{+}24, vLS24, vRB24]. sturgeon [BUM^{+}24]. sub [KKA^{+}21, UOO20, WCT^{+}20, Sen25]. sub-Arctic [KKA^{+}21, WCT^{+}20]. sub-Saharan [UOO20, Sen25]. Subarea [CHC^{+}21]. Subdistrict [SSL25]. subgroups [WNJ25]. subject [YZL24]. subjected [CBAPM20]. submarine [Dav25, KS22]. submissions [BL23a]. subregional [VRBC^{+}21]. subsidies [BCBS22, BHKK23, CMOB^{+}20, CMSN^{+}22, DLL25, FR23, For22, GICM25, KGP^{+}20, LBL21, MVSP^{+}21, SC22, SS21a, SSV^{+}23, ST24, VSD^{+}23, VSDR^{+}22, ALLS23]. subsidy [DGG22, SS21a, dSTHL^{+}24, WLQD21]. subsistence [Gal21, QCM^{+}20]. substitution [SKS21]. subsystems [LJL22]. success [DKW20, MB22, SEG^{+}22, ZSLY20]. successes} [CMS^{+}21]. Successful [GMC24, CCP25, EdLP^{+}21, Geh23, Rus20, VvSvH^{+}22]. sufficiency [ADK^{+}25]. sufficient [CLD^{+}24, KYMW21]. sugar [HKKY22]. suggestions [KOBS21, LL25]. suitability [WLS^{+}25]. suitable [OKH^{+}24, RZ21b]. Sulawesi [ETB25]. sulfide [JG20, VCC^{+}20]. sulfides [HZC^{+}24]. Sulfur [CGMH24, WZZC21, ZY21]. sulphur [Kon20, Tes20]. Sulu [ETB25, MIH^{+}24]. Summary [MGNP25]. summer [ZMT^{+}25]. Sunda [EABA^{+}22]. Sundarbans [Isl25]. sunfishes [PNS^{+}23]. sunken [Lin20]. Sunscreen [Lev20, ALML23]. sunscreens [Lev20]. super [MT20]. super-shark [MT20]. supervision [WXLD21]. supply [CTDY25, FLLML25, GKO^{+}25, HCSBM^{+}23, HRC^{+}24, How20, HP20, KSS23, KOW^{+}25, Lin25, MMM^{+}22, Tan22, TCC^{+}22, VMJ25, VVR^{+}22, WKBS20, WEF^{+}21, YZH22, ZCSK23, dlPC24]. support [ADWP24, AGK22, BST23, CWB^{+}20, CTH^{+}21, CVH^{+}20, CDB20, DKW20, hBDF22, DSD^{+}24, DB20, EBB^{+}25, FDGM24, Fla21, GPG^{+}21, GLS^{+}21, HS20, JKV^{+}22, KWMM20a, KVL^{+}23, LTW23, LDL21, LMC^{+}23a, MLMD21, MMC^{+}22, MF23, MM20, NC25, PGK22, PEM^{+}25, RLS20, RSS^{+}20, RBNE20, RKS^{+}23, SCZ^{+}24, THP22a, TEA^{+}22, TFMT^{+}23, VLG^{+}24, YYM25, ZLZ21]. Supporting [LHV^{+}20, PWL^{+}23b, SSZ^{+}22, BGMCG24, BYL25, CCG^{+}25, KGP^{+}22, Koe22, MH24, VDN^{+}24]. supportive [TAK^{+}22]. surf [MMS24]. Surface [DBZ24, Squ21, ATM^{+}24, BK23, GAN^{+}25, MNHS25]. surfer [BL22]. surfing [MYLV24]. surge [YSWW21]. Surigao [MMK^{+}23, MNGM24]. Suriname [DB20]. surrounding [MSO23]. surveillance [AIG^{+}20, CWR20, DGMF^{+}20, MNS^{+}24, SAB^{+}20]. survey [CC21, CSRG25, DDD22, KCCC^{+}21, KOBS21, KP24, MM21b, MSGAR20, NK21, DNR21]. survey-based [CC21]. surveys [CdLN^{+}21, GVH^{+}23, PBD^{+}23, SSOP^{+}21]. Survival [NSK20, OYKBA22]. suspects [TV21]. suspended [TMDN23]. Sustain [HCS21, MJ23, SWHT22]. Sustainability [ASM21a, CLU^{+}24, GAGLS25, GV22, HRMPCRRA22, HAsR24, LCAX21, NOL^{+}23, STS^{+}23, VMJ25, WKH22, AANM^{+}22, AVPAFM^{+}22, AMR21a, AH21, BMF20, BGLTC21, BDR20, BBO21, Cha24b, CCOP25, Doe21, FOH^{+}20, FMCT^{+}21, FG25, GdACT20, GCV^{+}21, GKO^{+}25, HPG20, HHH^{+}23, KBH^{+}24, KgJqY^{+}24, Kob22, KBP24, LSC25, MMK^{+}23, MMF^{+}25, MDG^{+}23, MAMCHP^{+}20, MPN^{+}24, MM20, NO25, NRF21, OR21, OI23, OGS24, PFG^{+}21, PBC^{+}24, PMM^{+}21, RC22, dSRKTC24, Sah22, Sep25, TUBA22, TOHK20, UISM^{+}23, VCC^{+}21, VAG^{+}24, WTO24, ZLWY22, Als20]. Sustainable [BH23b, BCN23, CTIH21, KA21, LHTR23, LBL21, LJ22, MZ24, NH20, NJW25, Ton22b, YHX24, ZLY23, dJRFVOA24, vSHHT20, AAA^{+}24, AAA^{+}21, ABH25a, AO20, Aqu23, BE20a, BAVV24, BCB20, BÅ25a, BB20, BDM21, BHB^{+}24, BYL25, BRR^{+}22, CN24, CSV25, CBO^{+}21, CMC^{+}23, CJL^{+}23, CdJK24, DGH25, Dbo22, DFF23, DZZC24, EMN^{+}24, ETSKA25, EAA^{+}21, ESD24, FR24, FGL^{+}21, FMS^{+}22, GNJR24, GBAECCRM25, GLAVL21, Gir22, GMG^{+}23, GLSVMB25, GMMMRC^{+}25, GMRSP23, Gug22, HS20, HPH22, HT23, HWMO20, HSN^{+}21, HC21, HP23, HM21, HBL^{+}21, IKB21, IMA22, JV20, KWF^{+}21, KS23, Kay25, KFP20, Koe22, LWH^{+}24, LZZ22, LZL25, LLL^{+}24, LLA^{+}25, MZ22, MSBA22, McC21, MMNC22, MBC^{+}24, MCTN23, NSN22, OGM24, OCD^{+}23, PSD23, PN20, PKB^{+}23, Pen22, Pri22, PŚ22, RJ21, Ren21a, RBB^{+}24b, RDM^{+}20, SdACdSLdAS24, Sas25a]. sustainable [SHP^{+}21, SHZ^{+}21, SGL^{+}21, TNE^{+}24, TOAPJ21, Tel23a, THP22a, Tho22, TGF^{+}25, Tra23b, WP23, WFP25, WN25, YCL25, ZJ20, ZWG^{+}21, ZXGL21, Zha22, dlPdlLLC^{+}22, dlPC24]. sustaining [FWGM24, NSPRSS20]. Svalbard [Tod20, Liu21, NT20, Øst24]. Svendsen [Cha20b]. Sweden [LDC^{+}24]. Swedish [DME23, PYVF24, WJB^{+}20]. swimmer [OCZ^{+}22]. Swimming [LSOP24]. Switzerland [Sta23a]. sword [Gug24]. swordfish [NOL^{+}23]. SWOT [AHBAKH24]. SWOT-AHP [AHBAKH24]. symbiosis [WCZS21]. symbolism [BST23]. Symes [Smi23a]. Symposium [SWGZ22]. sync [WEF^{+}21]. syndrome [CSdMR^{+}24]. Synergetic [GW21]. Synergies [EHÓ25, ANS^{+}20, BPA^{+}22, PT23, WOR^{+}25]. Synergistic [GGP^{+}22, ZLH21]. synergy [CCB^{+}21, LJL22]. synthesis [BB21, HMG^{+}21a, TFMT^{+}23, WKN^{+}25, WNMS^{+}24]. synthesizes [DBS22a]. system [ADB^{+}22, CHH^{+}25, CMC24, CLL22a, CDHY22, CJL^{+}23, DLX21, DSJ^{+}23, EMN^{+}24, FLC^{+}23, FFE20, FDF^{+}21, FMTA^{+}25, GDLC20, HYN^{+}23, HCW^{+}23, HJPL23, HLM^{+}20, JKV^{+}22, JF22, KL21a, KL24, KABT21, LG24, LDC^{+}20, LZX24, LXH^{+}22, Lin25, PVH^{+}21, PBS^{+}23, Qi22, Qi24, RMU^{+}22, SII22, SK20a, SIW^{+}21, SSZ^{+}24, SdCMN^{+}22, SVG24, SLO^{+}24, SCA^{+}22, SH25, SSP^{+}21b, THP22a, YLSC23, AM22, CBB^{+}24, lTS25, Gar20, TMBM^{+}25, WS24]. Systematic [WMM24, CJL^{+}23, CTDY25, EdLP^{+}21, HLLK22, KBB^{+}25, LRC24, OSBA24, QK24, RKC^{+}24, TSH^{+}22a]. Systems [dSOW^{+}21, BISSBdlMM22, BCJ21, BSK^{+}20, CSM^{+}25, FTM^{+}22, FAGS21, GM21, GPM23, GTMP20, HWT23, HWT24, HRR^{+}20, Her22, JL21d, LPMS24, LLS^{+}20, Mba22, MTS^{+}24, MdCLL^{+}23, MYGM25, MEG^{+}21, PWL^{+}23a, PLWO24, RDLA^{+}20, SOO20, SB25, Sta23a, STKC22, SN23, STF21, VGS^{+}20, VGA25, dSVT^{+}25, dRA25, GMG^{+}23].

T [HH21]. table [RL24, TM23a]. TAC [DSJ^{+}23, ZLD^{+}21, MH20a]. tacit [Zou23]. tackle [BR22]. Tackling [CDC^{+}25, GE23, MG22]. tactical [CPR21, CBR21]. tag [MFG^{+}23]. Tagus [SCA^{+}22]. Taiwan [CS21, COYS24, CCCC21, CJ22, HC21, HP23, HFY^{+}24, HM21, HCS21, HCT^{+}21, HWY22, Kao23, KWG23, KOCL23, LHW21, LTSL22, cL22, LTJ^{+}23, Liu22, LLL^{+}24, LLA^{+}25, LSY20, Sas25a, Shi20b, Shi24b, Tse21, TK22, TK23, TWK25, VM21, WYH21, WWC23, YL21, LL24a]. Taiwanese [LLCH21, LLSY22, SCGC20, TSWC23, YCL22a]. take [BBB^{+}20, HS24, NOB22, QTPL21, QTBL23, WBPG20]. Taking [CLD^{+}24, MZ24, TCZ21, VKN20, Wan23a, XL25, LC20b]. tale [PL24b]. Tallurutiup [KDH^{+}21]. Tam [Han21]. Tangerang [MKAF21]. tangled [OPL20]. Tanzania [BYL25, KOBM24, KMM24, MFC^{+}24, MYGM25, SML^{+}25, YMG23]. Tanzanian [IOHC22]. Target [GA20, BdSL25, CHEC21, DUYEA21, HGH^{+}20, SCSR^{+}22, SCSR^{+}23, WNJ25]. targeting [BLH^{+}20, BDCJ22, GBC^{+}20]. targets [EHÓ25, SPCCvT21, VvSvH^{+}22, Wan24a, Mor20]. tarpon [GCM^{+}23]. task [MBL24, RE21]. tasks [Dun23]. Tasmania [BHK^{+}25, POM24]. Tasting [CSH23]. Taurua [MRD^{+}20]. Tavolara [HNMSG21]. Tawi [ADI^{+}25, MH20b]. Tawi-Tawi [ADI^{+}25, MH20b]. tax [HYN^{+}23, RAB22, VSD^{+}23]. taxation [Åm21]. taxes [GICM25, MVSP^{+}21]. taxonomy [FFE^{+}24]. Taylor [Ngu23, WQ25]. Taza [GDY25]. teaching [PDS24]. Tears [SWW^{+}21]. tech [SLY21]. Technical [CPR21, GNJR24, DER^{+}21, GAMR20, RFPS20]. technique [KCCC^{+}21]. techniques [DBB^{+}22, Shi24b, SRSGMGG21b]. Technological [SDG^{+}23, VMD^{+}24, GLADGGVL23, HZC^{+}24, KSB^{+}23a, RZ21b, RJ21, Ren21a, Sha20]. technologies [ABVPM20, CTDY25, JD20, SBP24, SG20c, XLW24]. Technology [KZR22, OYECdOP24, Tan23a, GLS^{+}21, HPH22, Lin25, MBL24, NNR^{+}21, Ren21b, RBNE20, Shi24b, SA25, VHD20, WSD21, ZGY21]. technology-compatible [GLS^{+}21]. technology-enabled [Lin25]. Tel [PC20]. Telecouplings [SGL^{+}23]. temperate [QTPL21]. temperature [BK23]. Temporal [KCT^{+}24, dSF24, CGB^{+}25, FARBV^{+}24, GASLCP20, LYL^{+}22, RHST21, WL22b]. temporary [BBB^{+}20, DSB^{+}25, LBL21]. ten [Her22, BBWC21]. tendencies [NGP^{+}21]. Tenggara [RYS^{+}21]. tensions [CSM^{+}25, COSS21]. Tenualosa [DAAP21]. tenure [GWAPRPRU22, Li23b].

Teouma [SGPF21]. term [AHN25, AFMdSdTdCC22, BSBD21, CVA22b, Jon21, Kim20, LK24a, McC21, MM20, SOP^{+}22, SGFM^{+}21, Wat22, Fla21]. terminal [BHEE22, COJZ^{+}21, WW21a]. terminating [BNF21]. terminologies [HGSD23]. terminology [LIK^{+}24]. Terms [ABME22, CMC^{+}23]. Terrace [YXX^{+}24]. terrestrial [MCR22]. territorial [HLW^{+}20, Lin20, MPSF20, ONA21, FMCT^{+}21, RM21b]. territorial-based [HLW^{+}20]. territorialisation [SW22]. territorialization [Lam21]. territory [YT20]. test [LM24, SCLM^{+}21]. tested [HVS22]. Testing [KMJ24, RKS^{+}23, ALML23]. Texas [SKS21]. Text [BDLO23, ZTA23]. texts [HSHS24]. textual [YYD23]. Thai [CCKK22, MNS^{+}24]. Thailand [CJ23, CCKK22, Isr22, KL21a, KA21, VM21, WKBS20]. Thames [BMC20]. Their [CBH^{+}24b, CBH^{+}24a, CBH^{+}25, Laj21, SCM^{+}25, AANM^{+}22, BMB^{+}24, BGQL24, CJN25, CCK^{+}21, CDC20b, CNG^{+}23, DGI^{+}24, DA23, GBB^{+}21, Has20, HSK^{+}23, Kim20, NGP^{+}21, NMN^{+}25, PMK^{+}21, PDD^{+}25, PBG^{+}22, PMFR^{+}22, RLRS21, TAK^{+}22, VRBBM^{+}24, WMS21, WL22b, FMCT^{+}21, ZZH21]. them [Gal21, GLMBMP25, NT20]. thematic [Sta23a]. themes [RSS^{+}20]. theoretic [BM21, ZGY21]. Theoretical [MZZ24, YLSC23, CSL21, MWB^{+}21, YCF24]. theories [BSM23, EFFP23]. theorise [SG20b]. Theory [ABME22, Alf20, GLB^{+}24, Gra21, GSB20, LYQM21, MLM23, PHF20, PFG^{+}21, TMAS24, XXLW24, ZY21, DLYM22, HVS22]. theory-based [GSB20]. there [NT20, OOGS20b]. Thermal [RJAT23, Teo23]. thin [Dyc24]. thing [BDTD25]. thinking [JL21d, OGS24, STKC22, Wan24c]. Third [Kim21, DMNT20, FY24]. thirty [Szy20]. Thoothoor [PRM^{+}20]. those [AO20, KOBS21]. thought [CGHV20, DMG^{+}23]. thoughts [Zhu23]. threat [Kar22b, SSOP^{+}21, SZ24]. threaten [dMLB^{+}20]. Threatened [GMG^{+}24, BdLAdC^{+}20, BUM^{+}24, FRDK24, HS24, Jon21, OGS25, Smi20, TBPC20, dJRFVOA24]. threatens [CMI^{+}22, KgJqY^{+}24]. threats [CCP25, FTGD^{+}20, GCM^{+}23, KHLS23, LTRA20, MIMS^{+}20, McC24, OTK^{+}21, OOM^{+}25, PEM^{+}25]. Three [Dyr25, HM22c, SAO^{+}21, TUBA22, YZL24, ZRR^{+}25, BI21, Cha24a, CLU^{+}24, FOKM23, GPM20, HNMSG21, LXSX21, Mur21b, SRSGMGG21b, BH23a, STG^{+}20]. Three-dimensional [YZL24]. three-level [Cha24a]. thresher [CLB^{+}20, MH20b]. threshold [Sha20, VDOE^{+}24]. Thresholds [HSK^{+}23, CHN^{+}24, DOBE^{+}25, PWL^{+}23a, PLWO24, TPG^{+}23]. thriving [ANC^{+}25]. throughout [FGC^{+}24, LGP25, VKN20]. thwarted [OSV21]. Tianbao [Sos21a]. tidal [AH23, BMC20, CA20, CKY22, HHEJB20, JB25, ZCG^{+}23]. tide [BMK^{+}22, RCH22]. tides [SYY24]. tiered [HGL^{+}20, YLSC23]. Tierra [dPJ21]. tilapia [TCS^{+}20]. timber [VVR^{+}22]. Time [vLM22, Bau20, Bov22, BASM22, DMNT20, Dev22, Fox22, HRS^{+}24, MNS^{+}24, NCF25, SFGEPC22, TZM^{+}22, WPWM^{+}22, WL22b]. time-dependent [WL22b]. time-frequency [NCF25]. Timeline [BLH^{+}20]. timely [ZLD^{+}21]. times [GABSD21, LBL21, LEERRM^{+}21]. Timothy [Ste25]. tipping [BHB^{+}24]. title [CCX23]. Toda [dHJTGO21]. Togo [NSZ^{+}21]. Tokyo [HWMO20, XWL^{+}20]. tolerable [AS23a]. too [SMHPMO^{+}21]. tool [ASC25, BCJ21, BRP^{+}23, CRP^{+}20, CDB20, GMMMRC^{+}25, JHHG23, KKO^{+}22, OVSVPP^{+}22, PHD^{+}24, PS21, Sol23b, SAB^{+}20, WS23, YZL24]. tool-direction-subject [YZL24]. toolkit [DMHS25, MFC^{+}24]. tools [DMHS25, GGP^{+}22, QYD20, SMCW22, TBCB25, WF25]. toothfish [ADB^{+}22, CHC^{+}21]. top [SLB21]. top-down [SLB21]. topography [MLJ^{+}22]. TOPSIS [DG20, Kar22a]. Torre [Rus20]. Total [Sta23b, EJC24, DSJ^{+}23, ZLD^{+}21]. tour [Dev22]. Tourism [Pha20, SWW^{+}21, AVPAFM^{+}22, AAFK^{+}21, BBC24, DHC22, ESD24, GMFD25, GGZRAHVS25, GMGLV21, GMGLV22, HHBC20, HCDN24, Jon21, JK22, LLJZ22, LM24, MAMCHP^{+}20, Mil22, Mur21b, NUD^{+}23, NAN^{+}24, PFM20, PN20, RDM^{+}20, SBDC24, SRL25, SCY24, SAB^{+}24, Ton22b, TEA^{+}22]. tourist [AFRFGP25, LLJZ22, NNRS24, QWBY21]. tours [HACCM22]. toxic [Lev20]. toxicity [Lev20]. TPB [LYQM21]. trabeculata [CBO^{+}21]. traceability [CTDY25, DSWB20, HRC^{+}24, LFVCF24, THP^{+}22b]. traceable [DLYM22]. track [MKB21]. Tracking [CSGC22, TMBM^{+}25]. Trade [BCN23, BBT^{+}25, CMSN^{+}22, EB23, FXC^{+}22, PKMT23, AED20, BDTD25, BdLAdC^{+}20, CJCS^{+}22, DC20, GAD^{+}20, GASLCP20, IOHC22, IO23, KCT^{+}24, KL24, LP20, LCC20, LWG23, Ma20, MAA^{+}24, MMK^{+}22, NC25, OAdJC^{+}25, PMM^{+}21, SMF^{+}23, SGL^{+}23, SSH^{+}23, SX25, SK24, SFB24, SFN^{+}21, SAO^{+}24, SLY25, TJS22, XSCL21, ZTZ^{+}21, ZLGX20, ZT21, ZGY21, vRB24]. Trade-offs [BCN23, PKMT23, GASLCP20, ZGY21]. traders [IOHC22]. trades [CMI^{+}22]. Trading [WLXY22, WXLD21, WLQD21]. Traditional [DMd^{+}20, MFH^{+}20, WF25, dJdMCL^{+}24, BBP25, CSV25, DSQ^{+}22, FLB^{+}21, HLW^{+}20, HTP^{+}23, HKZ25, Li23b, SBP^{+}25]. Traffic [MCW^{+}22b, TKC23, BRKS25, Boy21, Gun21, GHK20, HWM^{+}25, MLJ^{+}22, PEM^{+}25, SMF^{+}23, SSOP^{+}21]. trafficking [MGF20]. tragedy [Zha21]. training [CLD^{+}24, DG24a, KZR22, RKR20, WN25]. traits [WWC23]. trammel [WLW^{+}22]. trammel-net [WLW^{+}22]. Trang [KJE21b]. Trans [hC24, RB22]. Trans-boundary [hC24, RB22]. transaction [YLSC23]. Transboundary [GBCNCL^{+}21, LJ20, Abo23, CC21, CZW24, DAAP21, ETB25, FDGM24, HNA25, HSdT21, KMTvT25, KMT^{+}21, LJ23, MFA^{+}24, PTS^{+}25, PGA^{+}20, RKS^{+}24, SMYH22, Tel23b, VLS^{+}23]. Transdisciplinary [GCP^{+}20, STF21, SONB^{+}22, FMCT^{+}21, GTMP20, RBB^{+}24b]. transfer [VHD20, ZWXL21]. transferable [HvPPV20b, SSAP24]. transference [JPM^{+}23]. transfers [NRYG20]. Transformation [SHS^{+}25, BJGG^{+}24, ETSKA25, EFFP23, LC22, MZZ24, YHX24, HCS21]. Transformational [IRK^{+}21]. transformations [ARP25, BSM23, Kow24]. transformative [GM21, HCG^{+}25, MdCLL^{+}23, VGC^{+}22]. transformed [MMM^{+}25, PCA^{+}23]. Transforming [BBGG22, CCS23, HLW^{+}20, SA25, MMM^{+}25, PMC24, SLB21]. Transition [BGQL24, ZLD^{+}21, AC23, CSRG25, Dom23, EMH^{+}23, GdACT20, GWAPRPRU22, KZR22, MMF^{+}25, MF23, NNR^{+}21, SS21b, THS^{+}24, WP23, WZX24, ZTA23, vHSSK20]. Transitions [KDS22, MM22, CMOB^{+}20]. transits [RZ21a]. Translating [SWI^{+}22]. translation [Fla21, SG21]. transmission [JTdH25, JG23]. transnational [BE20c, CAGBM22, TK23, Wu22]. Transparency [DH22, DFTK20, Gug22, GJS22, Pet22, TFF22, WKH22, Wil22, AMR21b, CMNH21, CTDY25, For22, Har22, Koe22, Ske24, VVR^{+}22]. transparent [BD20, Fis22, HGG22c, SMYH22]. Transpolar [BSY^{+}20]. transport [Ada22, AB20, ADB^{+}22, BG20a, CQW24, Dom23, GAGLS25, LLW22, RBTW24, RLRS21, VGK23]. transportation [IGS^{+}23]. Transposition [BN20]. Transshipment [SOT^{+}22, BJ25, Bet25, HLLK22, SMYH22]. Transversal [ST24]. traps [ADR^{+}23, DGH25, VGC^{+}22]. trawl [AE21, ADM^{+}25, BDCJ22, DDF^{+}20, GMPT^{+}20, JGWM^{+}23, KGP^{+}20, MKM^{+}20, NSK20, RQB^{+}20, SC22, SCCW21, ZV20]. trawlers [DXS^{+}21]. Trawling [DAA21, ATF^{+}21, BMGDS^{+}20, CHEC21, CGB^{+}25, DL21, FARBV^{+}24]. treat [BBV^{+}25]. Treated [Mab24]. treatment [Che21b]. Treaty [AM22, CBB^{+}24, lTS25, Gar20, WS24, Dau23, DAB^{+}25, DF24, HHD20, Kra22, MH24, MS20, MK24, TM23b, Liu21, TH22]. trend [AIHR^{+}25, RXN23, SWXM21, PZZ^{+}21]. Trends [KFB^{+}20, LBWS23, Mah22, QK24, Wu22, AN25, BSBD21, BMC20, BRR^{+}22, CLRW22, DDF^{+}20, DLX24, DAAP21, FARBV^{+}24, GRSRSN^{+}24, GZW^{+}20, Gun21, HBJ21, KCT^{+}24, LSD^{+}22, MAG^{+}22, MIH^{+}24, PL20a, SKDK23, Tho22, TDP20, VAS^{+}25, WLRH^{+}20, LJ20, SGL^{+}21]. Tri [SLP^{+}22]. Tri-National [SLP^{+}22]. Trial [CBR21]. trials [CBR21, CC22]. triggers [PHD^{+}24]. Trindade [dS22a]. tripartite [HZ23]. Triple [MVR21b, DAA21, LCAX21]. trips [DWC^{+}20]. Trophic [TCF^{+}25, ANC^{+}25, RF25]. tropic [CSV^{+}24]. tropical [BLH^{+}20, GCC^{+}21b, LLLF22, MDLA21, NOL^{+}23, RTMCVMDE22, RHO24, SdCMN^{+}22, DMS^{+}24]. tropics [MFG^{+}22]. Trouble [Tea20, BL23b]. Troubled [Smi23a, FS21]. true [NCF25]. trunculus [SVP^{+}23]. trust [DGS^{+}24, EVPMKP22, FOH^{+}20, FS21, HSdT21, How20, KJS^{+}22, RKS^{+}24]. trusts [KBH^{+}24, HBCA^{+}24]. Tudela [MJ23]. Tumaco [PCCOS24]. Tuna [Kar23, SBS21, Abo23, AIHR^{+}25, And24, BLH^{+}20, Bjø23, BDC^{+}21, DSWB20, EMH^{+}23, dHJTGO21, GSG20, GSB20, HAD24, HWT23, HCS21, JG23, Kao23, KB20, Kaw22, KMG^{+}20, KgJqY^{+}24, KLS^{+}25, Koe22, LLL^{+}24, PCSS^{+}21, PWL^{+}23b, RHO24, SMYH22, SAS^{+}25b, STF21, TOAPJ21, THP^{+}22b, TGF^{+}25, WTO24]. tuna-RFMOs [And24]. tunas [NOL^{+}23]. Tunisia [BDCJ22]. turbot [FGPPPV21]. turbulence [Lin25]. TURF [FSRC25a, FMCT^{+}21, FSRC25b]. TURF-Reserve [FSRC25a, FSRC25b]. TURFs [AS23a]. Turkey [CA21, YU20]. Turkish [ASG^{+}24, Knu25]. Turkiye [Okş23, Ero22]. Turkmenistan [AB20]. Turning [NBB^{+}23, RCH22]. turtle [HBHJ20, SBW^{+}21, VPB^{+}24]. turtles [FMWW21, GDB^{+}24, SAS24]. Tuscan [DSD^{+}24]. Twenty [GLSVMB25]. Twilight [YJG24]. Twitter [BDLO23]. Two [Ehl21b, KKO^{+}22, VMV24b, AFMdSdTdCC22, BBB^{+}20, CES20, DKW20, DBB^{+}22, Fro21, GJY23, GBC21b, GRPRVG23, JDF^{+}23, LJ20, LLSY22, MKM^{+}20, Pau21, PT23, RNGRC22, SON21, Sin21, TRL^{+}25, WWD24, ZWJvL20, ZTM^{+}23, dOPM^{+}20, dPJ21]. two-step [GJY23]. two-year [Sin21]. Tying [ADB^{+}22]. type [MB22, Zha22]. types [MÇ23, NNR^{+}21]. typology [BDC23, KLSG21, MPSF20]. Tyrrhenian [PBG^{+}22].

U.S. [ADWP24, AP23, AL20, BCW^{+}25, BGK^{+}22, Ber25, BBJ^{+}22, DAHS^{+}23, FGC^{+}24, GLS^{+}21, HKKY22, LGF24, MKSG20, MM20, PH24, RMF^{+}21, RBNE20, RL24, SMB^{+}25, Squ21]. Ucides [dFVCGdS21]. UK [EB23, LL24a, BWV25, CSM^{+}25, Che21a, Dbo22, DPSF25, DFF23, EFCU25, FDF^{+}21, Fox22, GVL22, KFP20, KSB^{+}23a, LP21, OGM24, RSS^{+}20, SCC25, SSW^{+}25, WtDD^{+}21, WCM^{+}24, WL24a, ZX24]. Ulrikke [Niu22]. Ulva [XLLS21]. umbilicals [VJL^{+}21]. unable [Han21]. Unambiguously [VvSvH^{+}22]. unavoidable [IBC^{+}20]. Unbound [Wil21a]. Uncertain [Ber24, FG25, BvdH23, DC20, Dyc24]. uncertainties [GSM20, SLS^{+}21]. uncertainty [BA25b, CBL24, NMCL24, SPCCvT21, UISM^{+}23]. uncharted [DBZ24, LF25]. UNCLOS [DBZ24, DC23, Ege23, Fra23, Guo20, HGL^{+}20, LR20, Men23, Ora23, Shi24a, Tel23b, Who23, Yu22, ZY23, ZT23b]. Uncovering [ABC^{+}24, WFR23]. Uncrewed [Dun23, Ros25, Ste23]. undelimited [Kim21]. under-management [dMLB^{+}20]. undercut [LWG23]. underdeveloped [GCP^{+}20, NWZ^{+}21]. Underlying [RE21, PLHK^{+}24]. Undermanaged [SSH^{+}23]. undermine [CMNH21, MGPR^{+}25]. undermining [BLL22, LV23, Tan24]. underneath [JXWL25]. underperforming [MVR21b]. underprotected [RCH22]. underrepresents [HWM^{+}25]. Undersea [Can25, Gov25, TG25b, Wat25, Pan25b, Ros25]. undersized [NSK20]. understand [BMK^{+}22, CSRG25, GMRSP23, MFM21]. Understanding [BSM23, CGB^{+}25, CHN^{+}24, CMS^{+}21, DAB^{+}25, DLYM22, DDGdS^{+}21, GBB^{+}21, HPS23, HP23, KP24, LZX24, LLJZ22, NEH^{+}23, OBNK24, SBDC24, SKL23, TRN^{+}24, XXMZ23, Zha21, ZCCQ22b, vRB24, AVF^{+}23a, FHBS23, IO23, LLM^{+}23, MPN^{+}24, NHD23, O’D21, SOO20, SCCW21, STF21, WMM24, WNBHG24, Zou23, KDM^{+}20, Ste25]. understandings [Has20]. Underwater [LC23, War23, AIG^{+}20, BRP^{+}23, EHÓ25, HMG^{+}21a, KOCL23, LHW21, Lin24, LF25, ML22, MLJ^{+}22, MG22, MJMAA21, PT23, PPvD^{+}24, SCD21, Tod25, TGM22, VÖB20a, VÖB20b, VRBC^{+}21, WM22a]. UNDRIP [Wan23c]. undulatus [HS23]. UNECA [MFC^{+}24]. UNEP [FFGJ22]. unequal [Ber24]. UNFCCC [Tel23b]. unfeasible [OGS25]. unified [ZLY23]. Unilateral [Li23a, BGK^{+}22]. Union [Ada22, ARP25, Álv23, Aqu23, FGC21, KL24, KGP^{+}20, LFPBD^{+}20, NVP^{+}23, OAdJC^{+}25, PŚ22, SPFA^{+}20, SOLN21, VSDR^{+}22]. unique [Tse21]. United [Cro21, WW21b, AOS^{+}24, CGAF25, CNB^{+}23, Col20, FS21, FG25, FBC^{+}23, Gar20, Gul20, HWU^{+}23, JFB^{+}24, JB25, KS22, KSB^{+}23b, LAF^{+}24, MH24, NGA23, PWL^{+}23a, SSA^{+}23, SBT21, Squ20, SPH^{+}20, SRH23, TG25b, VSD^{+}23, Wil21a]. universally [FRDK24]. universities [LLSY22, OKNB22]. University [LTJ^{+}23, Liu22, Niu22, Smi20, Smi22, Ste25, Tan23a, TCT21]. University-driven [LTJ^{+}23]. unjust [SSV^{+}23]. unknown [DSC^{+}23]. unlikely [HSGS^{+}23]. Unlocking [WOR^{+}25]. unmanaged [OKS^{+}22]. unmanned [CZW20, LF25, XZ23]. Unmasking [SSP^{+}21a]. unmet [ABC^{+}24]. unofficial [SSZ^{+}24]. Unpacking [HvdBH23, NCE21, THP^{+}22b, HGSD23, LV23]. Unprecedented [AM22]. Unraveling [BMK^{+}25, CT24, CJCS^{+}22, MAMCHP^{+}20, TFP25]. Unravelling [ASPT24, ADM^{+}25, CBO^{+}21, Mba22, SLS^{+}20]. unregulated [AGS20, BJH21, BJ25, Bet25, DGG22, KL21a, Ma20, PHP25, TSH^{+}22a]. Unreported [SOLN20, AGS20, BJH21, BJ25, Bet25, DGG22, KL21a, Ma20, PHP25, SSH^{+}23, TSH^{+}22a]. unsolved [RE23]. unstable [Ber24]. Untangling [BFSP25, CMGF20, EFFP23, JBGC25, LPMG^{+}24]. Untimely [AMR21b]. unveil [PHP25]. unveiled [BdLAdC^{+}20]. Unveiling [DG24b, KLS^{+}25]. unwanted [CPR21, IBC^{+}20]. upcoming [BRP^{+}23]. update [HAD24]. Updated [CN24]. Updating [GMG^{+}24, CMC24]. uphold [NOB22]. upon [CCCC21]. Upper [ABDRV^{+}23]. ups [CSQ^{+}24]. upsurge [SK24]. upward [TDP20]. urban [Bre25, JB25, KL21b, LHPJ23, MCW^{+}22b, NWZ^{+}21, QCM^{+}20]. urbanization [SGK21, Zha22]. urchin [CSMMC^{+}25]. urgency [SHZ^{+}21]. Uruguay [BJGG^{+}24]. USA [SLS25]. usage [hBDDF22, MGK^{+}25, RZ21a, YCBW21]. Use [BCW^{+}25, BMK^{+}22, DR25, DB20, FMCT^{+}21, GBC21b, LXH^{+}22, Liu23, Ngu23, PCD^{+}23, RM21b, VAMR^{+}20, ADI^{+}25, AKFY22, AAA^{+}24, AWBC20, ABH25a, AAW^{+}23, BH23b, BMRM23, Cha21, CLU^{+}24, CvHF^{+}24, ESD24, FMS^{+}22, GNJR24, GSAF21, VKL^{+}25, HS23, Her22, HAsR24, Kle21, KH24, Lev20, LDC^{+}20, LJRHH21, MM21a, MM23, OVA22, Pri22, PBH^{+}23, SCD21, SS21a, SMB^{+}25, SCB25, SSZ^{+}22, TLT^{+}22, TJZ^{+}21, DR22, VPRR^{+}21, VAG^{+}24, WYH21, WOR^{+}25, XZ23, YMG23, YWY20, YLSC23, Zha25]. used [DQJ23, GCV^{+}20, HGSD23, LW22]. useful [PMC24]. User [Mrd23, Run24, AR23a, AR23b, AEJ24, EBB^{+}25, Jen23, NHA23, SAB^{+}20, dRA25]. users [SMC^{+}21, SMC^{+}25, ZRR^{+}25]. Uses [Los24, ASM^{+}21b, DGMF^{+}20, MSB^{+}25, SV22, SZBA20]. Using [CCS23, CAVJ^{+}20, CLB^{+}20, EAGV20, ES22, HLW^{+}23, HLCM21, KDH^{+}21, MNS^{+}24, MEG^{+}21, PHP25, PMT^{+}20, STIS20, SNP21, SSOP^{+}21, SKL^{+}20, STKC22, ASV^{+}21, BVT^{+}21, CRP^{+}20, CSH23, DLX21, DLYM22, EAAG21, EV22, FLC^{+}23, GMM^{+}20, GCS21, HF20, HTP^{+}23, JKV^{+}22, KCCC^{+}21, Kar22a, Kaw22, KPN20, KDM^{+}20, LWSS21, LrS22, LYQM21, MVR21a, MFC^{+}24, MLM23, NGT^{+}23, OGM24, Okş23, RQB^{+}20, SGK21, SGWL21, SSAAB^{+}24, TM23a, TMAS24, TPG^{+}23, WD23, WSD21, XSCL21, XWH^{+}22, ZTA23, Cat20, FGK^{+}20]. Ustica [HNMSG21]. usual [TV21]. Utilising [CTH^{+}21]. Utilization [TJS22, Hor23, SOK^{+}22]. Utilizing [CSHO^{+}20].

UUVs [Ros25].

v3.0 [GMG^{+}24]. Vague [CMNH21]. Valdés [TDC21]. validation [LMSLMQ21]. valorisation [IBC^{+}20, PMAVM20]. Valuable [CLB^{+}23, dMLB^{+}20, KVL^{+}23, VKN20]. Valuation [MFC^{+}24, CMMVdPV25, Cat20, CKY20, LrS22, LW22, NNRS24, QTPL21, SMC^{+}25, ZMR20]. Value [dFVCGdS21, JAdS^{+}20, LLL^{+}24, YTG^{+}23, AANM^{+}22, ALM^{+}21b, BRM^{+}21, BBJ^{+}22, Bot20, CSMMC^{+}25, CJN25, CLW^{+}21, CLB^{+}23, CSCGC20, CEMRMB^{+}24, DLJ^{+}20, FGC^{+}24, FFYE21, GSG^{+}24, GNMGWM21, HPG20, HKKY22, HCDN24, JTdH25, JKKT20, KWMM20a, KWMM20b, Knu25, LTW23, LCL20, LWH^{+}24, LDL^{+}23, LRC24, LSOP24, MGS21, MBDG23, MFAH25, MWB^{+}21, MMW^{+}23, NAS^{+}22, PSTW21, PCSS^{+}21, PSG25, QTBL23, RMU^{+}22, Seu22, SdCMN^{+}22, THT^{+}23, THP^{+}22b, TGHF^{+}22, TOHK20, TPG^{+}23, DR22, VAS^{+}25, VKN20, WNM22, WCC^{+}20, WMS21, ZW21]. value-added [WCC^{+}20]. value-adding [LTW23]. value-chain [ALM^{+}21b]. value-satisfaction-demand [CLW^{+}21]. values [BJA^{+}24, CH25, CL24, FRVB22, GCL21, MSB^{+}25, NSM25, RF25, TW22, ZWC^{+}23, ZCA^{+}23]. Valuing [DJC^{+}23, MKAF21, QWBY21, RNGRC22, SL20, WFR23]. Vancouver [IGS^{+}23]. Vanuatu [SGPF21, SNK^{+}20]. variability [CGB^{+}25, CAVJ^{+}20, KOBM24]. variables [Kao23, OAM21]. variation [EMS^{+}23, GABA^{+}22]. Variations [HPPH22, Zho23]. Varied [Pan25b]. various [LWX^{+}20, vLPBK21]. vary [CLL20]. varying [CDG^{+}21]. vector [SdACdSLdAS24, SGK21]. vehicles [CZW20, Kle21, LF25, Ros25, VJL^{+}21, WABR^{+}22]. Veierud [Niu22]. Veiled [APGWTM^{+}24]. venue [SC24, SC25]. venus [BSBD21]. Verde [Aqu23]. versa [AFMdSdTdCC22]. versus [BSSL^{+}20, CCKK22, COM20, DMd^{+}20, dOPM^{+}20]. Vessel [BCJ21, AMIM21, AdSMS21, BRKS25, CSGC22, Dam24, FMWW21, HWM^{+}25, KLS22, KLSG21, KDM^{+}20, MNS^{+}24, NSPRSS20, PEM^{+}25, SSOP^{+}21, Shi22, Stø21, TMBM^{+}25]. vessels [BV21, CDY24, CTR24, Dun23, FMCD20, KZR22, KLS^{+}25, LDG^{+}25, MCW^{+}22b, OVA22, PSMR20, Set23, Ste23, SRY24, dSTHL^{+}24, TBV^{+}23, Tse21, VÖB20a, VÖB20b, VAMR^{+}20, YS23]. VI [VSV^{+}23]. via [BDLO23, JHL21, MWB^{+}21, VSPMS^{+}22, ZLWY22]. viability [DAN^{+}23, YTBM^{+}22]. viable [RZ21a]. vice [AFMdSdTdCC22]. vice-versa [AFMdSdTdCC22]. vicinity [Shi22]. vicious [YS21]. victories [CMS20]. video [AIG^{+}20, HMG^{+}21a, SKL^{+}20]. Viet [Han21]. Vietnam [DBHT22, HS20, KJE21a, KJE21b, LAF^{+}24, NHD23, TCS^{+}20, TVN23]. Vietnamese [Cao25, NMN^{+}25]. view [KGW24, Kuł22]. viewpoints [RKKM24]. views [EFCU25, JB25, VJÓK25, ZLWY22]. VIII [WL20]. violence [Bra22]. Violeta [LPO25]. Virgin [GMRSP23, YTBM^{+}22]. Virginia [WGBS25]. Virtual [HDH21]. Vis [KBC25]. Vis-à-Vis [KBC25]. visceral [CSH23]. vision [RFPS20, VAMR^{+}20, XT23, ZMZ23]. visitor [DHC22]. Vistula [RB22]. visualization [STIS20]. Visualizing [KBD^{+}20, ZTA23, WCF^{+}21]. Vitoria [dS22a]. VMS [CLS^{+}23]. voice [LEERRM^{+}21]. voices [FAMW^{+}25]. void [MF25]. Vol [Mot22]. Volatility [VSPMS^{+}22, JTdH25]. volcanoes [SNK^{+}20]. volume [SVV24, SSH^{+}23]. Voluntary [Cha24b, HWT24, MM23, RSN25, SMCW22, WS23, WTO24, NH20]. volunteers [MFM21]. voting [Sta23b]. voyage [Bov22]. vs [BCABG24, GABSD21, KLSG21, Rog24]. vu [SBT^{+}23]. Vulnerabilities [GPTASM23, HSM^{+}23]. Vulnerability [CBAPM20, KKAO^{+}25, KHH^{+}23, MGS21, APGWTM^{+}24, CD24, CSGC22, DAN^{+}23, DGIS^{+}22, FMGV20, Gar24, HTH25, HP23, IAS25, JTR^{+}20, KOBM24, KDVR20, RDLA^{+}20, SMLRB^{+}21, SWF^{+}24, SCLM^{+}21, SPA^{+}21, VSPMS^{+}22]. vulnerable [BMR20, HRC^{+}24, JHHG23, KAA^{+}24, dMLB^{+}20].

W [De 20]. Wadden [CWKE20]. Wades [RROP23]. wages [YTÇ22]. waka [MAR^{+}20b, MRD^{+}20]. Waka-Taurua [MRD^{+}20]. wake [FGL^{+}21]. wakes [FAC^{+}24]. Walbaum [EEIER25]. Wales [BBGG22, MCW^{+}22a, BD20, BHB^{+}24, SSDQ20]. Walking [SH24a]. wall [dZ20]. war [Gra21, Rog24]. warming [CTH^{+}21, KS23, LDAC20]. warning [SK24, SLO^{+}24, WCZS21]. warnings [SLMG^{+}21]. Warships [Ngu23, GC22, Lin20]. Washington [Shi22]. waste [DMd^{+}20, DBS22a, FDVP24, JAH^{+}24, MB22, Öİ24, PCN^{+}24, XTY22, XHZ^{+}21, YXDL24, ZMR20, ZHL^{+}23, vLS24]. wastewater [FL24, Lia25, LCL24, PL24a, QL24, WZL25]. watch [AdSMS21, vLPBK21]. watching [AdSMS21, AFRFGP25, STS^{+}23, Shi22, SHAC23, SRL25]. Water [Zho23, AP24, ABH25b, BDCJ22, Bre25, CL24, hC24, HCS21, JKV^{+}22, KLS22, LWZ23, Lia25, LYQM21, LCL24, MXZ23, Mab24, MSGAR20, Shi23b, TFP25, WL22a, XTY22, YS23, YL21, YYM25, YXDL24, NSN22, Gul20]. watercraft [SMF^{+}23]. Waters [Smi23a, BBK^{+}21b, CMWN^{+}23, CMOJ^{+}24, DWC^{+}20, Dbo22, DER^{+}21, DBZ24, DD22, FS21, GBC^{+}20, JL24, KKO^{+}22, LDG^{+}25, Lin20, LF25, MLJ^{+}22, MNS^{+}24, MGBS24, ONA21, OOGS20a, ÓHCÓ24, OKFN24, RMBS25, RCM24, SCGC20, TNE^{+}24, TZM^{+}22, Tod22b, TYS22, Wil22, WtDD^{+}21, Zha25, vdGD21]. watershed [JZW21]. wave [MMS24]. wavelet [Tar25]. waves [BL22, KDHL24]. way [CQW24, DSJ^{+}23, DHR23, FGK^{+}20, JDF^{+}23, MS21, VGC^{+}22, Wil21b, YHX24]. ways [JPM^{+}23]. weakfish [dSJNS20]. Wealth [ASGS23, ODG25, hBDF24, GV22]. weather [FRM22, GAMR20, SPA^{+}21]. Weaving [CSV25]. web [OPL20, PBT^{+}20, SLS^{+}20, THP22a, TMTC20]. web-based [THP22a]. Weddell [TBH^{+}21]. wedge [SVP^{+}23]. wedgefishes [CJCS^{+}22]. Welfare [Gro20, ASGS23, CAM^{+}23, DBB^{+}22, HTH25, STG^{+}20, SOS^{+}21b, TSS^{+}24, ZZH21, ZSWJ23]. well [AVF^{+}23a, BBB25, BCG^{+}24, GCF^{+}23, GSA23, GGZRAHVS25, MNGM24, MTC^{+}23, MH20a, QTPL21, RGH^{+}22, SCF^{+}22, SKL23, TRL^{+}25]. well-being [AVF^{+}23a, BBB25, BCG^{+}24, GCF^{+}23, GGZRAHVS25, MNGM24, MTC^{+}23, RGH^{+}22, SCF^{+}22]. well-established [QTPL21]. well-planned [TRL^{+}25]. well-regulated [SKL23]. wellbeing [CFM^{+}21, MYLV24, dlPdlLLC^{+}22]. West [BG20b, BGMCG24, GSG^{+}24, NRF21, RHL20, SWI^{+}22, SN23, DXS^{+}21, GSA23, AAA^{+}24, AWBC20, BCBS22, BDVR25, hBDF24, DP23, EJC24, FTGD^{+}20, HKZ25, KABT21, MZVC25, RYS^{+}21, TNE^{+}24, WFK^{+}24]. Western [HAD24, Jon21, ABRA20, BSBD21, BDC^{+}21, HNMSG21, KDH^{+}21, LSGG^{+}20, MFRQ23, McC24, NWZ^{+}21, NEH^{+}23, OCCSC23, OSZ^{+}25, Abo23, AVC24, CCK^{+}21, GLADGGVL23, GAA^{+}20, KWF^{+}21, MAG^{+}22, OA20, RDD^{+}21, BRSC^{+}24]. wetland [PLLJ24]. wetlands [WLX21, ZCG^{+}23]. whacking [YML^{+}21]. Whale [AdSMS21, AFRFGP25, Mar20a, RCM^{+}22, GBAECCRM25, GCV^{+}20, MM20, OKS^{+}22, STS^{+}23, SNP21, Shi22, SHAC23, SRL25, TDC21]. whale-ship [SNP21]. Whale-watch [AdSMS21]. whale-watching [AdSMS21, Shi22, SHAC23, SRL25]. whales [BMK^{+}25, Dam24, FMCD20, GHK20, HSH^{+}21, MCW^{+}22b, Shi22, SHAC23, TBAR^{+}25, PTS^{+}25]. Whaling [WL20, LSC25, WNM22]. what-if [PLHK^{+}24]. what-now [PLHK^{+}24]. wheelchair [SMC^{+}25]. whelk [DUYEA21]. Where [And24, BÖ24, CFS^{+}24, RSFvH^{+}22, SZ24]. whether [HRR^{+}20]. which [RE21]. white [AS23b, BBHN20, HACCM22, MAMCHP^{+}20, SMHPMO^{+}21, HACCM22]. Who [CSCGC20, MHT^{+}24, MMF^{+}25, RKC^{+}24, KOBS21, LV23, Pre21, MK24]. Whose [MB21]. Wicked [PTCC24, Gro20, OOGS20b]. wide [HMG^{+}21a]. widespread [RZ21a]. Wild [Geh23, JKCY25, MM22]. wild-caught [MM22]. Wilderness [McC20]. Wildlife [KHMM20, Lot22, NUD^{+}23, GBC21a]. Will [BGK^{+}22, MGNP25, PK23, Bal22, Ber23, JD20, KZR22]. Willingness [CLC21, AFRFGP25, BCS^{+}24, BR22, BUM^{+}24, CLL20, CKY22, NMN^{+}25, SLSV25, XLLS21, XXLW24]. willingness-to-pay [CLL20]. win [Ton22b]. Wind [WML25, BvdBvdFK20, BST23, BLY^{+}24, CSLGFV20, CTR24, CCW21, DWC^{+}20, DBHT22, DPSF25, FAC^{+}24, HRR^{+}20, HJ25, JvH24, KNY20, MBORGN24, OC20, ONA21, OKH^{+}24, PKMT23, SLS25, SMB^{+}25, SBT21, Smy24, SZBA20, SVKP21, TYS24, WOR^{+}25, vRB24, FAC^{+}24]. wind-powered [CTR24]. Windfarms [DS22b]. winding [BMH^{+}22]. window [Lot23]. windows [ETB25]. Winners [PKMT23, BCMCM21]. wish [HNA25]. withdraw [PCK22]. withdrawal [MLA^{+}24, SOS25]. withdrawals [LLM^{+}23]. within [BDTD25, BCJ21, BDL^{+}24, Bre25, CCBB^{+}25, CTH^{+}21, lTS25, Gar24, GG22, MPD^{+}23, MMAP25, NBB^{+}23, PTS^{+}25, RAMEG24, SMKS22, TOL^{+}22, VHD20, dS21]. without [MLM23]. WMO [Yu22]. Women [FS22, MMC^{+}22, MMK^{+}23, ABRA20, FFYE21, GE23, LL24b, MLB^{+}25, MMK^{+}22, NEF23, OBNK24, PK20a, SBP^{+}25, Szy20, UOO20]. words [HBL^{+}21]. Work [BCABG24, TNY23, ASKI20, BBK21a, JSW^{+}22a, Kuł22, LJ20, LFLF22, LSD^{+}22, MMM^{+}25, MWWS21, Pau21, SGA^{+}21, TJS^{+}25, YS23, YTÇ22, NSN22]. Work-family [TNY23]. work-related [Pau21]. Work/rest [BCABG24]. Worker [SMCW22]. Worker-less [SMCW22]. workers [EEIER25, SMCW22, VM21, WS23]. workforce [JD20, TFP25]. Working [Bau20, AFHHD22, GPH^{+}21, HTH25, JSW^{+}22a, PBC^{+}24, RAFLRP21, VM21]. Workplace [ÖB22, NEF23, WW21a]. workplaces [NOB22]. works [CHB^{+}20]. Workshop [MGBS24, GRA^{+}22, HMG^{+}21a, WFP25]. world [BLM24, BMR23, BS22, CLP21, EMH^{+}23, HS24, HM22c, KFP20, SBG^{+}24, SBG^{+}25, Seu22, Smi22, ZJ22, CMSN^{+}22]. worldwide [CMGF20]. worst [MHT^{+}24, OGM24]. worth [Tod20]. Would [BBHN20, ALLS23, EECE^{+}24]. wrasse [HS23, AGSI22]. wreck [PT23]. wrecks [WL24b]. writing [Gel25]. WTO [ALLS23, PL24a]. WWII [Lin20].

x [Smi22]. Xávega [ACV^{+}21]. xiii [Smi20]. Xin’an [JZW21]. xv [Rub24]. XX [WL20, LL24a, Sta23a].

yacht [VGK23]. Yaldad [DSQ^{+}22]. Yamamoto [dHJTGO21]. Yangtze [QYX21]. Yantai [DG20]. year [Bau20, MFT^{+}22, NUD^{+}23, PBG^{+}22, Sin21, VMM21, VBd^{+}21]. Years [FSKJ23, BMR23, Ege23, FFE^{+}24, Gal22, GdCC^{+}25, GLSVMB25, Gul23, KSB^{+}23b, Liu21, NAP^{+}21, SCRCI24, SAS24, Smi23a, SRL25, Szy20, ZABI25]. yellow [Pan25a, TOAPJ21, Cho22, LSH21]. yellow-fin [TOAPJ21]. Yen [LK24b]. Yen-Chiang [LK24b]. yields [McC21]. Yong [TCT21]. Yong-An [TCT21]. York [Ngu23, Smi23a, WQ25]. Young [MMM^{+}25, Tan23a, DBBL21, WNBHG24]. youth [SN23]. Yucatan [HDAPG^{+}24, RSLR24, VHCJdYC^{+}24]. Yuen [LK24b].

Zambia [MMK^{+}22]. Zanzibar [MSHK25]. Zealand [PCN^{+}24, HNB^{+}25, LLT^{+}20, MAR^{+}20b, QTBL23, RR20, RM21a, RNGRC22, RRM^{+}24, SPJB22, WJLSH21]. zero [GCMT20]. zetu [CMWT24]. Zhejiang [LYL^{+}22]. Zhikong [YML^{+}21]. Zhoushan [YD22]. ziwe [CMWT24]. Zone [BMR23, CAGBM22, BCW^{+}25, BGMCG24, CDHY22, COYS24, CLP21, CZ24, FRVB22, OKFN24, RXX24, SCRCI24, dSdLSGB22, FXC^{+}22, WLS^{+}25, YJG24]. Zones [LZL25, BBB^{+}20, CZ24, JACD20, PMR23, QF20, dO23, GMMMRC^{+}25, SOLN20]. zoning [CMC24, MFG^{+}22, ONA21, SRR^{+}21, SSB^{+}23, TZZ^{+}21, NAP^{+}21].

References

Addo:2021:ISP

[AAA^{+}21] Kwasi Appeaning Addo, Olanike Adeyemo, Ibukun Jacob Adewumi, Robert Blasiak, S. Kim Juniper, Dwight Owens, Tessa Owens, and U. Rashid Sumaila. Informed selfishness — Practical reflections on building a sustainable ocean economy. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003468.

Abrokwah:2024:ASS

[AAA^{+}24] Sika Abrokwah, Ivy Serwaa Gyimah Akuoko, Margaret Fafa Awushie Akwetey, Mike Izava Olendo, Peter Kershaw, and Denis Worlanyo Aheto. Achieving sustainable source reduction of marine litter for ocean conservation in West Africa: Insights from single-use plastic consumers in Liberia. Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004700.

Alvarez:2023:CPC

[AACdlT23] Lourdes Alvarez, Gregory Arcaya, Francisco Cueto, and Renato de la Torre. Can public credit programs improve artisanal fisher performance? The case of FONDEPES a credit program. Marine Policy, 156(??):??, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003251.

Arcos-Aguilar:2021:DTM

[AAFK^{+}21] Ramiro Arcos-Aguilar, Fabio Favoretto, Joy A. Kumagai, Victoria Jiménez-Esquivel, Adán L. Martínez-Cruz, and Octavio Aburto-Oropeza. Diving tourism in Mexico — Economic and conservation importance. Marine Policy, 126 (??):??, April 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000208.

AlAmin:2021:CKA

[AAKI21] M. A. Al Amin, L. Adrianto, T. Kusumastanto, and Z. Imran. Community knowledge, attitudes and practices towards environmental conservation: Assessing influencing factors in Jor Bay Lombok Indonesia. Marine Policy, 129 (??):??, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001329.

Acosta-Alba:2022:MFV

[AANM^{+}22] Ivonne Acosta-Alba, Gian Nicolay, Adama Mbaye, Moustapha Dème, Ludovic Andres, Marc Oswald, Henri Zerbo, Joseph Ndenn, and Angel Avadí. Mapping fisheries value chains to facilitate their sustainability assessment: Case studies in The Gambia and Mali. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004656.

Ankrah:2023:CUI

[AAW^{+}23] Martinson Twumasi Ankrah, Dennis Asante, Pengcheng Wang, Evans Brako Ntiamoah, and Yuansheng Jiang. Can the use of the Internet improve fish farmers’ financial performance? Evidence from Ghana. Marine Policy, 149(??): ??, March 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000210.

Akbulaev:2020:MTE

[AB20] Nurkhodzha Akbulaev and Gadir Bayramli. Maritime transport and economic growth: Interconnection and influence (an example of the countries in the Caspian sea coast; Russia, Azerbaijan, Turkmenistan, Kazakhstan and Iran). Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302726.

Azzurro:2024:UUD

[ABC^{+}24] Ernesto Azzurro, Sara Bonanomi, Marina Chiappi, Rocco De Marco, Gian Marco Luna, Matilde Cella, Stefano Guicciardi, Francesco Tiralongo, Andrea Bonifazi, and Pierluigi Strafella. Uncovering unmet demand and key insights for the invasive blue crab (Callinectes sapidus) market before and after the Italian outbreak: Implications for policymakers and industry stakeholders. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24002938.

Acharjee:2025:PHF

[ABC^{+}25] Debasish Chandra Acharjee, Shakila Begum, Sumon Roy Chowdhury, Md Ruhul Amin, Kamal Gosh, and G. M. Monirul Alam. Post-harvest fish losses by the fish farmers in Haor areas of Bangladesh. Marine Policy, 176(??):??, June 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000533.

Aceves-Bueno:2023:DGI

[ABDRV^{+}23] E. Aceves-Bueno, L. Davids, J. A. Rodriguez-Valencia, A. M. Jaramillo-Legorreta, E. Nieto-Garcia, G. Cárdenas-Hinojosa, E. Hidalgo-Pla, A. Bonilla-Garzón, A. J. Diaz de Leon, L. Rojas-Bracho, and M. A. Cisneros-Mata. Derelict gear from an illegal fishery: Lessons from gear retrieval efforts in the Upper Gulf of California. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004341.

Alam:2025:EDF

[ABH25a] Md. Mehedi Alam, Md. Saifullah Bin Aziz, and Mohammad Mahfujul Haque. The extent of destructive fishing gear use in Bangladesh: Ecological impacts and strategic roadmap for sustainable fisheries management. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25002337.

Albrecht:2025:PCE

[ABH25b] Eerika Albrecht, Antti Belinskij, and Elina Heikkilä. Policy coherence for ecosystem-based management: Implementing EU water and marine policies in the Archipelago Sea. Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004275.

Afriansyah:2024:SGI

[ABI24] Arie Afriansyah, Leonardo Bernard, and Christou Imanuel. Should Indonesia regulate foreign military activities in its EEZ? Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004645.

Allan:2021:ECD

[ABM21] Jamie C. Allan, Karen F. Beazley, and Anna Metaxas. Ecological criteria for designing effective MPA networks for large migratory pelagics: Assessing the consistency between IUCN best practices and scholarly literature. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20308654.

Ayaz:2022:RSP

[ABME22] İlke Sezin Ayaz, Umur Bucak, Mahmut Mollaoğlu, and Soner Esmer. Resilience strategies of ports against Covid-19 in terms of chaos theory. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003700.

Abolhassani:2023:SPT

[Abo23] Angela Abolhassani. Scalar politics in transboundary fisheries management: the Western and Central Pacific Fisheries Commission as an eco-scalar fix for South Pacific albacore tuna management. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001100.

Ameyaw:2020:FCA

[ABRA20] Anita B. Ameyaw, Annette Breckwoldt, Hauke Reuter, and Denis W. Aheto. From fish to cash: Analyzing the role of women in fisheries in the western region of Ghana. Marine Policy, 113(??):??, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306541.

Antelo:2020:IMT

[ABVPM20] Luis T. Antelo, Jose M. Bellido, Julio Valeiras, and Ricardo I. Pérez-Martín. Innovative marine technologies applied to discard mitigation and management: the MARTEC18 conference. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20302013.

Azzurro:2021:PMI

[AC21] Ernesto Azzurro and Jacopo Cerri. Participatory mapping of invasive species: a demonstration in a coastal lagoon. Marine Policy, 126(??):??, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000221.

Axon:2023:BBE

[AC23] Stephen Axon and Sammy Collier. Breaking blue: Establishing comprehensive policy for a just and inclusive transition for the Blue Economy. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003906.

Albrecht:2021:PAD

[ACA^{+}21] Renee Albrecht, Carly N. Cook, Olive Andrews, Kelsey E. Roberts, Martin F. J. Taylor, Michael B. Mascia, and Rachel E. Golden Kroner. Protected area downgrading, downsizing, and degazettement (PADDD) in marine protected areas. Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000476.

Ansong:2021:MAB

[ACG21] J. Ansong, H. Calado, and P. M. Gilliland. A multifaceted approach to building capacity for marine/maritime spatial planning based on European experience. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18304056.

Alvarez:2020:DSA

[ÁCT20] Antonio Álvarez, Lorena Couce, and Lourdes Trujillo. Does specialization affect the efficiency of small-scale fishing boats? Marine Policy, 113(??):??, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301721.

Alves:2021:SSA

[ACV^{+}21] Fátima Alves, Paula Casaleiro, Cristina Sá Valentim, Paula Castro, and Miguel Pardal. Small-scale artisanal fisheries and its networks: the case of Arte Xávega in Portugal. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000956.

Adamowicz:2022:DMT

[Ada22] Magdalena Adamowicz. Decarbonisation of maritime transport — European Union measures as an inspiration for global solutions? Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001324.

Ashford:2022:TPS

[ADB^{+}22] Julian Ashford, Michael Dinniman, Cassandra Brooks, Lian Wei, and Guoping Zhu. Tying policy to system: Does the Ross Sea region marine reserve protect transport pathways connecting the life history of Antarctic toothfish? Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005145.

Abduraup:2025:LPU

[ADI^{+}25] Yennyriza T. Abduraup, Araniza M. Diansuy, Roger A. Irilis, Argamar A. Habibuddin, Akkil S. Injani, Mohammad Amilussin T. Ammang, Fauriza J. Saddari, Adzmer A. Juaini, Elsie S. Omandan, Mary Joyce Z. Guinto-Sali, and Richard N. Muallil. Local perceptions on the use of inorganic commercial fertilizers in eucheumatoid seaweed farming: Insights from Tawi-Tawi islands, Philippines. Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400486X.

Ahtiainen:2025:QES

[ADK^{+}25] Heini Ahtiainen, Luke F. Dodd, Samuli Korpinen, Kristīne Pakalniete, and Liisa Saikkonen. Quantifying effectiveness and sufficiency of measures – an application of the DPSIR framework for the marine environment. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004809.

Aranha:2025:UDA

[ADM^{+}25] Sofia Graça Aranha, Ester Dias, Tiago Marsili, Amos Barkai, Nuno Queiroz, Pedro Pires da Rocha, and Alexandra Teodósio. Unravelling the deep: Assessing the bycatch of deep-sea elasmobranchs in crustacean bottom trawl fisheries in Portugal. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005554.

Andres:2021:MCS

[ADR^{+}21] Marga Andrés, Matthias Delpey, Irene Ruiz, Amandine Declerck, Caroline Sarrade, Peggy Bergeron, and Oihane C. Basurko. Measuring and comparing solutions for floating marine litter removal: Lessons learned in the southeast coast of the Bay of Biscay from an economic perspective. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000609.

Antonelis:2023:DED

[ADR^{+}23] Kyle Antonelis, Joan Drinkwin, Paul Rudell, Jason Morgan, James Selleck, Donald Velasquez, and Don P. Rothaus. Determining effectiveness of Dungeness crab escapement in derelict traps. Marine Policy, 149(??):??, March 2023.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300026X.

Arranz:2021:WWV

[AdSMS21] Patricia Arranz, Natacha Aguilar de Soto, Peter T. Madsen, and Kate R. Sprogis. Whale-watch vessel noise levels with applications to whale-watching guidelines and conservation. Marine Policy, 134(??):??, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003870.

Alva:2024:WCB

[ADWP24] Amanda Alva, Kelly Dunning, Ryan Williamson, and Kampol Pannoi. What compels bipartisan lawmakers to support coral reef legislation in the U.S. Congress. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000927.

Akpalu:2021:ESE

[AE21] Wisdom Akpalu and Håkan Eggert. The economic, social and ecological performance of the industrial trawl fishery in Ghana: Application of the FPIs. Marine Policy, 125(??): ??, March 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20308873.

Anbleyth-Evans:2023:PDM

[AE23] Jeremy Anbleyth-Evans. Port developments and marine democracy in Latin America. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003007.

Alshubiri:2020:IFP

[AED20] Faris Alshubiri, Mohamed Elheddad, and Nadia Doytch. The impact of fish production on marine trade balance and foreign direct investment: an empirical study of the GCC economies. Marine Policy, 116(??):??, June 2020.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301320.

Arnason:2024:SUR

[AEJ24] Ragnar Arnason, Håkan Eggert, and Frank Jensen. Strong user rights in fisheries: an editorial note. Marine Policy, 161 (??):??, March 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000071.

Amaralal:2023:ICP

[AEL^{+}23] K. H. M. L. Amaralal, E. M. A. P. Edirimanna, W. A. S. W. Lakmini, K. K. D. Chamodi, A. U. Kuragodage, R. G. Sanuja, E. G. K. Y. C. Bandara, M. M. A. S. Maheepala, M. N. D. F. Abeykoon, K. P. G. L. Sadaruwan, S. Kuganathan, and K. H. M. Ashoka Deepananda. Impacts of COVID-19 pandemic on the fisheries sector of Sri Lanka. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003864.

Anbleyth-Evans:2020:TMD

[AELR^{+}20] Jeremy Anbleyth-Evans, Francisco Araos Leiva, Francisco Ther Rios, Ricardo Segovia Cortés, Vreni Häussermann, and Carolina Aguirre-Munoz. Toward marine democracy in Chile: Examining aquaculture ecological impacts through common property local ecological knowledge. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19303513.

Alvarez-Fernandez:2020:LQM

[ÁFFSC20] Inmaculada Álvarez-Fernández, Juan Freire, and Noela Sánchez-Carnero. Low-quality management of Marine Protected Areas in the North-East Atlantic. Marine Policy, 117 (??):??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19303847.

Amon:2022:SWE

[AFHHD22] Diva J. Amon, Zoleka Filander, Lindsey Harris, and Harriet Harden-Davies. Safe working environments are key to improving inclusion in open-ocean, deep-ocean, and highseas science. Marine Policy, 137(??):??, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005583.

Azzahra:2023:BRA

[AFHM23] Tami Azzahra, Muh. Farhan, Muh Aldhy Hatmar, and Mamnunah. Book review: Asia–Pacific Fishing Livelihoods, Michael Fabinyi, Kate Barclay. Palgrave Macmillan, Cham (2021). Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002373.

Andriguetto-Filho:2022:SFV

[AFMdSdTdCC22] José Milton Andriguetto-Filho, Rodrigo Pereira Medeiros, André Martins Vaz dos Santos, and Paulo de Tarso da Cunha Chaves. Shrimp up, fish down, and vice-versa: Fishers’ strategies and long-term changes in small-scale fisheries landings at two spatial levels in Southern Brazil. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002317.

Afzal:2025:ABE

[AFR25] Ayesha Afzal, Saba Fazal Firdousi, and Myra Imran Rafiq. Assessing the blue economy’s impact on carbon mitigation growth levels: Evidence from North and Latin America. Marine Policy, 180(??):??, October 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002180.

Avila-Foucat:2025:WWC

[AFRFGP25] Véronique Sophie Avila-Foucat, Daniel Revollo-Fernández, Diane Gendron, and Ernesto Israel Popoca. Whale watching choice experiment to assess boat crowding and whale abundance on tourist willingness to pay in Mexico. Marine Policy, 172(??):??, February 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005104.

Arifin:2023:IPR

[AFS^{+}23] Zainal Arifin, Dede Falahudin, Hiroaki Saito, Tuti Hendrawati Mintarsih, Muhammad Hafizt, and Yulianto Suteja. Indonesian policy and researches toward 70% reduction of marine plastic pollution by 2025. Marine Policy, 155(??): ??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002191.

Asche:2021:FPA

[AGA^{+}21] Frank Asche, Taryn M. Garlock, Wisdom Akpalu, Ebele Chinelo Amaechina, Robert Botta, Nnaemeka Andegbe Chukwuone, Håkan Eggert, Ken Hutchings, Razack Lokina, Byela Tibesigwa, and Jane K. Turpie. Fisheries performance in Africa: an analysis based on data from 14 countries. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2030909X.

Alves:2022:FPB

[AGK22] Catherine L. Alves, Olga Denise Garcia, and Randall A. Kramer. Fisher perceptions of Belize’s Managed Access program reveal overall support but need for improved enforcement. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002391.

Amon:2022:ASG

[AGM^{+}22] Diva J. Amon, Sabine Gollner, Telmo Morato, Craig R. Smith, Chong Chen, Sabine Christiansen, Bronwen Currie, Jeffrey C. Drazen, Tomohiko Fukushima, Matthew Gianni, Kristina M. Gjerde, Andrew J. Gooday, Georgina Guillen Grillo, Matthias Haeckel, Thembile Joyini, Se-Jong Ju, Lisa A. Levin, Anna Metaxas, Kamila Mianowicz, Tina N. Molodtsova, Ingo Narberhaus, Beth N. Orcutt, Alison Swaddling, Joshua Tuhumwire, Patricio Urueña Palacio, Michelle Walker, Phil Weaver, Xue-Wei Xu, Clement Yow Mulalap, Peter E. T. Edwards, and Chris Pickens. Assessment of scientific gaps related to the effective environmental management of deep-seabed mining. Marine Policy, 138 (??):??, April 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000537.

Aedo:2020:SDS

[AGN^{+}20] Gustavo Aedo, Cristóbal Garcés, Edwin Niklitschek, Selim Musleh, Luis A. Cubillos, and Renato A. Quiñones. Spatial distribution of small pelagic fishes: Implications for fishing quota allocations. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304956.

Alqattan:2020:IUU

[AGS20] Mohammad E. A. Alqattan, Tim S. Gray, and Selina M. Stead. The illegal, unreported and unregulated fishing in Kuwait: Problems and solutions. Marine Policy, 116(??): ??, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19302672.

Arieta:2022:CIA

[AGSI22] Siti Arieta, Budhi Gunawan, R. Widya Setiabudi Sumadinata, and Yudi Nurul Ihsan. Calling for institutional arrangements in Napoleon Wrasse conservation and management in Indonesia: the Anambas Islands case. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003050.

Autzen:2021:WSB

[AH21] Mathilde Højrup Autzen and Troels Jacob Hegland. When ‘sustainability’ becomes the norm: Power dynamics in the making of a new eco-label for low-environmental-impact, small-scale fisheries. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003535.

Alducci:2023:PCI

[AH23] Beatrice Alducci and Stephen Hynes. Protecting coastal infrastructure in the face of increasing climate related pressures: a cost-benefit analysis of maintaining a tidal kayaking feature in Ireland. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000775.

Amkieltiela:2022:REI

[AHAB^{+}22] Amkieltiela, Christian N. Handayani, Dominic A. Andradi-Brown, Estradivari, Amanda K. Ford, Maria Beger, Amehr Hakim, Dominic K. Muenzel, Eleanor Carter, Firdaus Agung, Laura Veverka, Mohamad Iqbal, Muhammad E. Lazuardi, Muhammad N. Fauzi, Sylvie N. Tranter, and Gabby N. Ahmadia. The rapid expansion of Indonesia’s marine protected area requires improvement in management effectiveness. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003049.

Araos:2023:FBA

[AHB^{+}23] Francisco Araos, Carlos Hidalgo, Francisco Brañas, Jeremy Anbleyth-Evans, Florencia Diestre, and Allan Yu Iwama. Facing the blue Anthropocene in Patagonia by empowering indigenous peoples’ action networks. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004444.

Abdel-Hady:2024:CAS

[AHBAKH24] Mahmoud M. Abdel-Hady, Ahmed A. Barrania, Zeinab M. Abdel-Khalek, and Shaimaa M. Haggag. A comprehensive approach to strategic planning for marine aquaculture in Egypt: SWOT-AHP analysis. Marine Policy, 162(??):??, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000551.

Abdel-Hady:2024:CNP

[AHENNA^{+}24] Mahmoud M. Abdel-Hady, Thanaa El-Noby, Ahmed M. Nasr-Allah, Seham A. Hashem, Zeinab M. Abdel-Khalek, Shaimaa M. Haggag, and Abdel-Fattah M. El-Sayed. Challenging the negative perceptions of key stakeholders towards aquaculture sector in Egypt: Evidence-based solutions. Marine Policy, 169(??):??, November 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003294.

Ansong:2022:IBI

[AHM22] Joseph Onwona Ansong, Ritchie Heather, and Linda McElduff. Institutional barriers to integrated marine spatial planning on the island of Ireland. Marine Policy, 141(??): ??, July 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001294.

Apps:2023:SDC

[AHN^{+}23] Kirin Apps, Elizabeth Heagney, Quach Thi Khanh Ngoc, Kay Dimmock, and Kirsten Benkendorff. Scuba divers, coral reefs, and knowledge of ocean acidification. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003123.

Aalto:2025:SLT

[AHN25] Pami Aalto, Minna Hanhijärvi, and Daiju Narita. Short and long-term solutions for mitigating black carbon emissions in Arctic shipping: Implications for stakeholders. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005657.

Arifin:2024:MBF

[AHR24] Ridwan Arifin, Margaretha Hanita, and Arthur Josias Simon Runturambi. Maritime border formalities, facilitation and security nexus: Reconstructing immigration clearance in Indonesia. Marine Policy, 163(??):??, May 2024.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400099X.

Ahlquist:2025:FSN

[AHT25] Ina Helene Ahlquist, Hanne Hjelle Hatlebrekke, and Rachel Tiller. Fishing for solutions: Norwegian fishers’ perspectives on the implementation of automatic catch registration for combating IUU fishing. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25001654.

Aguzzi:2020:EVM

[AIG^{+}20] J. Aguzzi, N. Iveša, M. Gelli, C. Costa, A. Gavrilovic, N. Cukrov, M. Cukrov, Nu Cukrov, D. Omanovic, M. Štifanić, S. Marini, M. Piria, E. Azzurro, E. Fanelli, and R. Danovaro. Ecological video monitoring of Marine Protected Areas by underwater cabled surveillance cameras. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20302669.

Acevedo-Iglesias:2025:BTF

[AIHR^{+}25] Sergio Acevedo-Iglesias, Miguel Herrera, María Lourdes Ramos, José Carlos Báez, Jon Ruiz, Gonzalo Rodríguez-Rodríguez, Vanessa Rojo, Pedro J. Pascual-Alayón, and Francisco J. Abascal. Bycatch trend and its fate of the Spanish-owned tuna purse seiners fleet from the Atlantic and Indian Oceans: Impacts of the implementation of good practices. Marine Policy, 177(??):??, July 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001095.

Arifin:2024:SIF

[AJDA24] Zainal Arifin, Joelianingsih Joelianingsih, Yuncheng Deng, and Noor Cahyo Dwi Aryanto. Small island futures: a conference report on the pathways to resilience and development. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002641.

Almada:2025:IEA

[AKF^{+}25] Amalia Aruda Almada, A. J. Kim, Heather Fenney, Diane Y. Kim, Nathan Churches, Linda Chilton, Maria Madrigal, Maira Quintanilla, and Victoria Campbell-Árvai. Improving equitable access to seafood: a case study of a community-driven effort in South Central Los Angeles, California. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005621.

Abe:2022:DUM

[AKFY22] Hiroya Abe, Yuko F. Kitano, Tomohiro Fujita, and Hiroya Yamano. Distribution, use, management, regulation, and future concerns of reef-building corals based on administrative documents in Japan. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001373.

Arguello:2022:RSA

[AKL^{+}22] Gabriela Argüello, Niels Krabbe, David Langlet, Ida-Maja Hassellöv, Claes Martinson, and Astrid Helmstad. Regulation of ships at anchor: Safety and environmental implications. Marine Policy, 140(??):??, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000999.

Ayers:2020:ESC

[AL20] Adam L. Ayers and Kirsten Leong. Examining the seascape of compliance in U.S. Pacific island fisheries. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19305573.

Alff:2020:NCI

[Alf20] Henryk Alff. Negotiating coastal infrastructures: an evolutionary governance theory (EGT) approach to Chinese high-modernist development along the Indian Ocean. Marine Policy, 112(??):??, February 2020. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18303348.

Arias:2022:RBM

[ALF^{+}22] Juliana Arias, Unn Laksá, Benvido D’Oliveira Fonseca, Mamadou Diallo, Khallahi Brahim, Alexandre Rodríguez, Sonia Martin Doblado, Juliana Antunes Galvão, Jónas R. Vidarsson, and Ragnhildur Fridriksdóttir. Results-based management in practice: Lessons learnt and policy recommendations from the implementation of RBM in European fisheries outside Europe. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000859.

Ayres:2024:LDK

[ALLR^{+}24] Kathryn A. Ayres, Frida Lara-Lizardi, Callum M. Roberts, Walter Pisco-Limones, Peter Klimley, Salvador J. Jorgensen, Felipe Galván-Magaña, Mauricio Hoyos-Padilla, and James T. Ketchum. Local diver knowledge reveals decline in scalloped hammerhead sharks (Sphyrna lewini) at seamounts in the southwestern Gulf of California. Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004487.

Alger:2023:WWA

[ALLS23] Justin Alger, Philippe Le Billon, Eric Leinberger, and U. Rashid Sumaila. What would Article 5.1 of the 2022 WTO Ministerial Agreement on Fisheries Subsidies accomplish? Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001689.

Almojil:2021:LEK

[Alm21a] Dareen Almojil. Local ecological knowledge of fisheries charts decline of sharks in data-poor regions. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002499.

Asri:2021:RMM

[ALM^{+}21b] Azam Asri, Virginie Le Masson, Valéria Montalescot, Phaik Eem Lim, Adibi M. Nor, Hanafi Hussin, and Louise Shaxson. The role of migrants in the Malaysian seaweed value-chain. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004231.

Akerlof:2023:WLP

[ALML23] K. L. Akerlof, Jacqueline Loevenich, Sara Melena, and Christine A. Lipsky. What’s local policy got to do with it? Testing the efficacy of federal interventions to protect coastal parks from sunscreen chemical pollution in Hawaii and North Carolina. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003822.

Alshubiri:2020:EBE

[Als20] Faris Alshubiri. Editorial: the blue economy of the GCC region: Sustainability and security. Marine Policy, 116(??): ??, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20300816.

Alvarez:2023:PFC

[Álv23] Inma Valeije Álvarez. The prosecution of fisheries crime in Spanish criminal law: the impact of European Union regulations. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003748.

Aam:2021:CPS

[Åm21] Heidrun Åm. A critical policy study on why introducing resource rent taxation in Norwegian salmon aquaculture failed. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003031.

Arpi:2022:FAS

[AM22] Bruno Arpi and Jeffrey McGee. Fishing around the South Georgia Islands and the “Question of the Falklands/ Malvinas”: Unprecedented challenges for the Antarctic Treaty System. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002482.

Abe:2021:RDE

[AMIM21] Keita Abe, Yusuke Matsuki, Gakushi Ishimura, and Shunsuke Managi. Rent distribution in an ex-vessel auction market of fisheries. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003821.

Abspoel:2021:CMS

[AMK^{+}21] Lodewijk Abspoel, Igor Mayer, Xander Keijser, Harald Warmelink, Rhona Fairgrieve, Malena Ripken, Andrej Abramic, Andreas Kannen, Roland Cormier, and Sue Kidd. Communicating Maritime Spatial Planning: the MSP Challenge approach. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304044.

Araujo:2021:GIP

[AML^{+}21] Laís S. Araujo, Ulises Rodrigo Magdalena, Tayana S. Louzada, Paulo S. Salomon, Fernando C. Moraes, Beatrice P. Ferreira, Eduardo T. C. Paes, Alex C. Bastos, Renato C. Pereira, Leonardo T. Salgado, Maria Lucia Lorini, Patricia Yager, and Rodrigo L. Moura. Growing industrialization and poor conservation planning challenge natural resources’ management in the Amazon Shelf off Brazil. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000762.

Atkins:2023:SPR

[AMLC^{+}23] Shanan Atkins, Judy Mann-Lang, Geremy Cliff, Neville Pillay, and Mauricio Cantor. Stakeholder perceptions reveal obstacles and opportunities to change lethal methods of protecting bathers from sharks. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002956.

Archibald:2021:IGC

[AMR21a] D. W. Archibald, R. McIver, and R. Rangeley. The implementation gap in Canadian fishery policy: Fisheries rebuilding and sustainability at risk. Marine Policy, 129(??): ??, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001019.

Archibald:2021:UPD

[AMR21b] D. W. Archibald, R. McIver, and R. Rangeley. Untimely publications: Delayed Canadian fisheries science advice limits transparency of decision-making. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003018.

Afinowi:2025:MCM

[AN25] Olubunmi A. Afinowi and Godwell Nhamo. Mapping coastal and marine research in South Africa: a bibliometric analysis highlighting trends and future directions. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005669.

Asaduzzaman:2025:ALT

[ANC^{+}25] Md Asaduzzaman, Md Nahiduzzaman, Muhammad Tanvir Hossain Chowdhury, Md Moshiur Rahman, Abdullah-Al Mamun, and Mohammad Mokarrom Hossain. Advancing low-trophic extractive mariculture (LTEM): Strategies for a thriving blue economy in Bangladesh. Marine Policy, 173 (??):??, March 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24005578.

Andreassen:2024:RTR

[And24] Ingrid Solstad Andreassen. The role of tuna-RFMOs in combating ‘ghost fishing’: Where is the catch? Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002379.

Anonymous:2020:A

[Ano20a] Anonymous. August 2020. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:D

[Ano20b] Anonymous. December 2020. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:EBa

[Ano20c] Anonymous. Editorial Board. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19308681.

Anonymous:2020:EBb

[Ano20d] Anonymous. Editorial Board. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20300063.

Anonymous:2020:EBc

[Ano20e] Anonymous. Editorial Board. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20300865.

Anonymous:2020:EBd

[Ano20f] Anonymous. Editorial Board. Marine Policy, 114:??, April 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20301937.

Anonymous:2020:EBe

[Ano20g] Anonymous. Editorial Board. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20302384.

Anonymous:2020:EBf

[Ano20h] Anonymous. Editorial Board. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20303687.

Anonymous:2020:EBg

[Ano20i] Anonymous. Editorial Board. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2030470X.

Anonymous:2020:EBh

[Ano20j] Anonymous. Editorial Board. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20305388.

Anonymous:2020:EBi

[Ano20k] Anonymous. Editorial Board. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20307831.

Anonymous:2020:EBj

[Ano20l] Anonymous. Editorial Board. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20308496.

Anonymous:2020:EBk

[Ano20m] Anonymous. Editorial Board. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20308800.

Anonymous:2020:EBl

[Ano20n] Anonymous. Editorial Board. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309544.

Anonymous:2020:F

[Ano20o] Anonymous. February 2020. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:Ja

[Ano20p] Anonymous. January 2020. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:Jc

[Ano20q] Anonymous. July 2020. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:Jb

[Ano20r] Anonymous. June 2020. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:Ma

[Ano20s] Anonymous. March 2020. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:Mb

[Ano20t] Anonymous. May 2020. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:N

[Ano20u] Anonymous. November 2020. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:O

[Ano20v] Anonymous. October 2020. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2020:S

[Ano20w] Anonymous. September 2020. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Aa

[Ano21a] Anonymous. April 2021. Marine Policy, 126(??):??, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Ab

[Ano21b] Anonymous. August 2021. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:D

[Ano21c] Anonymous. December 2021. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:EBb

[Ano21d] Anonymous. Editorial Board. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309957.

Anonymous:2021:EBc

[Ano21e] Anonymous. Editorial Board. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000051.

Anonymous:2021:EBd

[Ano21f] Anonymous. Editorial Board. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000270.

Anonymous:2021:EBe

[Ano21g] Anonymous. Editorial Board. Marine Policy, 126(??):??, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000683.

Anonymous:2021:EBf

[Ano21h] Anonymous. Editorial Board. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001068.

Anonymous:2021:EBg

[Ano21i] Anonymous. Editorial Board. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001494.

Anonymous:2021:EBh

[Ano21j] Anonymous. Editorial Board. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002086.

Anonymous:2021:EBi

[Ano21k] Anonymous. Editorial Board. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002396.

Anonymous:2021:EBj

[Ano21l] Anonymous. Editorial Board. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003298.

Anonymous:2021:EBk

[Ano21m] Anonymous. Editorial Board. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003699.

Anonymous:2021:EBl

[Ano21n] Anonymous. Editorial Board. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004292.

Anonymous:2021:EB

[Ano21o] Anonymous. Editorial Board. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100470X.

Anonymous:2021:F

[Ano21p] Anonymous. February 2021. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Ja

[Ano21q] Anonymous. January 2021. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Jc

[Ano21r] Anonymous. July 2021. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Jb

[Ano21s] Anonymous. June 2021. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Ma

[Ano21t] Anonymous. March 2021. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:Mb

[Ano21u] Anonymous. May 2021. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:N

[Ano21v] Anonymous. November 2021. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:O

[Ano21w] Anonymous. October 2021. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2021:S

[Ano21x] Anonymous. September 2021. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Aa

[Ano22a] Anonymous. April 2022. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Ab

[Ano22b] Anonymous. August 2022. Marine Policy, 142(??):??, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:D

[Ano22c] Anonymous. December 2022. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:EBa

[Ano22d] Anonymous. Editorial Board. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005066.

Anonymous:2022:EBb

[Ano22e] Anonymous. Editorial Board. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000124.

Anonymous:2022:EBc

[Ano22f] Anonymous. Editorial Board. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000483.

Anonymous:2022:EBd

[Ano22g] Anonymous. Editorial Board. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000756.

Anonymous:2022:EBe

[Ano22h] Anonymous. Editorial Board. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001105.

Anonymous:2022:EBf

[Ano22i] Anonymous. Editorial Board. Marine Policy, 140(??):??, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001610.

Anonymous:2022:EBg

[Ano22j] Anonymous. Editorial Board. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200197X.

Anonymous:2022:EBh

[Ano22k] Anonymous. Editorial Board. Marine Policy, 142(??):??, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002135.

Anonymous:2022:EBi

[Ano22l] Anonymous. Editorial Board. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002937.

Anonymous:2022:EBj

[Ano22m] Anonymous. Editorial Board. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003177.

Anonymous:2022:EBk

[Ano22n] Anonymous. Editorial Board. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003645.

Anonymous:2022:EB

[Ano22o] Anonymous. Editorial Board. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2200416X.

Anonymous:2022:F

[Ano22p] Anonymous. February 2022. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Ja

[Ano22q] Anonymous. January 2022. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Jc

[Ano22r] Anonymous. July 2022. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Jb

[Ano22s] Anonymous. June 2022. Marine Policy, 140(??):??, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Ma

[Ano22t] Anonymous. March 2022. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:Mb

[Ano22u] Anonymous. May 2022. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:N

[Ano22v] Anonymous. November 2022. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:O

[Ano22w] Anonymous. October 2022. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2022:S

[Ano22x] Anonymous. September 2022. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Aa

[Ano23a] Anonymous. April 2023. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Ab

[Ano23b] Anonymous. August 2023. Marine Policy, 154(??):??, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:D

[Ano23c] Anonymous. December 2023. Marine Policy, 158(??):??, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:EBa

[Ano23d] Anonymous. Editorial Board. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004638.

Anonymous:2023:EBb

[Ano23e] Anonymous. Editorial Board. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000143.

Anonymous:2023:EBc

[Ano23f] Anonymous. Editorial Board. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000684.

Anonymous:2023:EBd

[Ano23g] Anonymous. Editorial Board. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000921.

Anonymous:2023:EBe

[Ano23h] Anonymous. Editorial Board. Marine Policy, 151(??):??, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001446.

Anonymous:2023:EBf

[Ano23i] Anonymous. Editorial Board. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001781.

Anonymous:2023:EBg

[Ano23j] Anonymous. Editorial Board. Marine Policy, 153(??):??, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2300204X.

Anonymous:2023:EBih

[Ano23k] Anonymous. Editorial Board. Marine Policy, 154(??):??, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002506.

Anonymous:2023:EBj

[Ano23l] Anonymous. Editorial Board. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003445.

Anonymous:2023:EBk

[Ano23m] Anonymous. Editorial Board. Marine Policy, 156(??):??, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003561.

Anonymous:2023:EBl

[Ano23n] Anonymous. Editorial Board. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004190.

Anonymous:2023:EBm

[Ano23o] Anonymous. Editorial Board. Marine Policy, 158(??):??, December 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004360.

Anonymous:2023:F

[Ano23p] Anonymous. February 2023. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Ja

[Ano23q] Anonymous. January 2023. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Jc

[Ano23r] Anonymous. July 2023. Marine Policy, 153(??):??, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Jb

[Ano23s] Anonymous. June 2023. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Ma

[Ano23t] Anonymous. March 2023. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:Mb

[Ano23u] Anonymous. May 2023. Marine Policy, 151(??):??, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:N

[Ano23v] Anonymous. November 2023. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:O

[Ano23w] Anonymous. October 2023. Marine Policy, 156(??):??, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2023:S

[Ano23x] Anonymous. September 2023. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. April 2024. Marine Policy, 162(??):??, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:Ab

[Ano24b] Anonymous. August 2024. Marine Policy, 166(??):??, August 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:D

[Ano24c] Anonymous. December 2024. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:EBa

[Ano24d] Anonymous. Editorial Board. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23005043.

Anonymous:2024:EBb

[Ano24e] Anonymous. Editorial Board. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000204.

Anonymous:2024:EBc

[Ano24f] Anonymous. Editorial Board. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000654.

Anonymous:2024:EBd

[Ano24g] Anonymous. Editorial Board. Marine Policy, 162(??):??, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000897.

Anonymous:2024:EBe

[Ano24h] Anonymous. Editorial Board. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24001696.

Anonymous:2024:EBf

[Ano24i] Anonymous. Editorial Board. Marine Policy, 164(??):??, June 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24001945.

Anonymous:2024:EBg

[Ano24j] Anonymous. Editorial Board. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24002550.

Anonymous:2024:EBh

[Ano24k] Anonymous. Editorial Board. Marine Policy, 166(??):??, August 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24002690.

Anonymous:2024:EBi

[Ano24l] Anonymous. Editorial Board. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003075.

Anonymous:2024:EBj

[Ano24m] Anonymous. Editorial Board. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2400349X.

Anonymous:2024:EBk

[Ano24n] Anonymous. Editorial Board. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003890.

Anonymous:2024:EB

[Ano24o] Anonymous. Editorial Board. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24004640.

Anonymous:2024:F

[Ano24p] Anonymous. February 2024. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:Ja

[Ano24q] Anonymous. January 2024. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:Jc

[Ano24r] Anonymous. July 2024. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:Jb

[Ano24s] Anonymous. June 2024. Marine Policy, 164(??):??, June 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:Ma

[Ano24t] Anonymous. March 2024. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:Mb

[Ano24u] Anonymous. May 2024. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:N

[Ano24v] Anonymous. November 2024. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:O

[Ano24w] Anonymous. October 2024. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2024:S

[Ano24x] Anonymous. September 2024. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2025:Aa

[Ano25a] Anonymous. April 2025. Marine Policy, 174(??):??, April 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2025:A

[Ano25b] Anonymous. August 2025. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2025:EBa

[Ano25c] Anonymous. Editorial Board. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24005232.

Anonymous:2025:EBb

[Ano25d] Anonymous. Editorial Board. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24005499.

Anonymous:2025:EBc

[Ano25e] Anonymous. Editorial Board. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25000259.

Anonymous:2025:EBd

[Ano25f] Anonymous. Editorial Board. Marine Policy, 174(??): Article 106643, April 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X25000582.

Anonymous:2025:EBe

[Ano25g] Anonymous. Editorial Board. Marine Policy, 175(??):??, May 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25000806.

Anonymous:2025:EBf

[Ano25h] Anonymous. Editorial Board. Marine Policy, 176(??):??, June 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25000909.

Anonymous:2025:EBg

[Ano25i] Anonymous. Editorial Board. Marine Policy, 177(??):??, July 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2500140X.

Anonymous:2025:EBh

[Ano25j] Anonymous. Editorial Board. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2500171X.

Anonymous:2025:EBi

[Ano25k] Anonymous. Editorial Board. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25002076.

Anonymous:2025:EBj

[Ano25l] Anonymous. Editorial Board. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25002520.

Anonymous:2025:EBk

[Ano25m] Anonymous. Editorial Board. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2500288X.

Anonymous:2025:F

[Ano25n] Anonymous. February 2025. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2025:Ja

[Ano25o] Anonymous. January 2025. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2025:Jc

[Ano25p] Anonymous. July 2025. Marine Policy, 177(??):??, July 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2025:Jb

[Ano25q] Anonymous. June 2025. Marine Policy, 176(??):??, June 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2025:Ma

[Ano25r] Anonymous. March 2025. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2025:Mb

[Ano25s] Anonymous. May 2025. Marine Policy, 175(??):??, May 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2025:N

[Ano25t] Anonymous. November 2025. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2025:O

[Ano25u] Anonymous. October 2025. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Anonymous:2025:S

[Ano25v] Anonymous. September 2025. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic).

Ayostina:2022:NAB

[ANR^{+}22] Ines Ayostina, Lucentezza Napitupulu, Barakalla Robyn, Cynthia Maharani, and Daniel Murdiyarso. Network analysis of blue carbon governance process in Indonesia. Marine Policy, 137(??):??, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000021.

Abramic:2020:IMS

[ANS^{+}20] A. Abramic, N. Nogueira, P. Sepulveda, M. Cavallo, Y. Fernández-Palacios, C. Andrade, S. Kaushik, and R. Haroun. Implementation of the Marine Strategy Framework Directive in Macaronesia and synergies with the Maritime Spatial Planning process. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309192.

Amundsen:2020:BCB

[AO20] Vilde Steiro Amundsen and Tonje Cecilie Osmundsen. Becoming certified, becoming sustainable? Improvements from aquaculture certification schemes as experienced by those certified. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030169X.

Asiedu:2021:FDA

[AONF21] Berchie Asiedu, Paulina Okpei, Francis Kofi Ewusie Nunoo, and Pierre Failler. A fishery in distress: an analysis of the small pelagic fishery of Ghana. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001111.

Afewerki:2023:IPN

[AOO^{+}23a] Samson Afewerki, Tonje Osmundsen, Marit Schei Olsen, Kristine Vedal Størkersen, Andreas Misund, and Trine Thorvaldsen. Innovation policy in the Norwegian aquaculture industry: Reshaping aquaculture production innovation networks. Marine Policy, 152(??):??, June 2023.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001513.

Alati:2023:GAF

[AOO^{+}23b] Victor Mwakha Alati, Kennedy Osuka, Levy Michael Otwoma, Paul Tuda, and Lina Mtwana Nordlund. Gender analysis in fisheries: the case of the shelled mollusc fisheries in Kenya. Marine Policy, 158(??):??, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003962.

Advani:2024:ESS

[AOS^{+}24] Sahir Advani, Jeffrey K. O’Hara, Sarah M. Shoffler, Patricia Pinto da Silva, Juan Agar, JuliAnna Arnett, Lilian Brislen, Matthew Cutler, Abigail Harley, Justin Hospital, Karma Norman, Edward Ragland, Dale Squires, Brent Stoffle, Marysia Szymkowiak, Americo J. Vega-Labiosa, and Joshua S. Stoll. Estimating the scope, scale, and contribution of direct seafood marketing to the United States seafood sector. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001866.

Anderson:2023:CAU

[AP23] Paul A. Anderson and Robert S. Pomeroy. Comparative assessment of the U.S. State of Florida’s Marine Life regulations for its commercial marine aquarium fish fishery. Marine Policy, 149(??):??, March 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000507.

Ahn:2024:EEI

[AP24] Soojung Ahn and Robert Pomeroy. Evaluating the economic impact of distant water fleet activities on small-scale fishers in the South China Sea. Marine Policy, 168(??): ??, October 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003221.

Andrefouet:2022:IIN

[APF22] Serge Andréfouët, Mégane Paul, and A. Riza Farhan. Indonesia’s 13558 islands: a new census from space and a first step towards a One Map for Small Islands Policy. Marine Policy, 135(??):??, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004590.

Alonso-Poblacion:2024:VCC

[APGWTM^{+}24] Enrique Alonso-Población, Renato Gozzer-Wuest, Josymar Torrejón-Magallanes, Iván Gómez, Percy Bayona, and Rolando Labraña-Cornejo. Veiled costs, concealed incentives: Rethinking market approaches to address institutional marginalization and reduce vulnerability of artisanal fishers to power abuses in the Peruvian dolphinfish fishery. Marine Policy, 170(??):??, December 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400410X.

Asranudin:2024:BRM

[APH24] Asranudin, Agus Wedi Pratama, and Holilah. Book review: Microplastic in the Environment: Pattern and Process, Michael S. Bank, in: Part of the book series: Environmental Contamination Remediation and Management. Springer (2022). Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24002239.

Abad:2020:ISM

[APV^{+}20] Esther Abad, Maria Grazia Pennino, Julio Valeiras, Raul Vilela, José María Bellido, Antonio Punzón, and Francisco Velasco. Integrating spatial management measures into fisheries: the Lepidorhombus spp. case study. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18308509.

Aquino:2023:LEU

[Aqu23] Mateus Lopes Aquino. The limits of the European Union’s fisheries agreements as sustainable development instruments: the case of Cape Verde. Marine Policy, 148(??): ??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22005024.

Ardito:2022:RCR

[AR22] Giovanni Ardito and Marzia Rovere. Racing the clock: Recent developments and open environmental regulatory issues at the International Seabed Authority on the eve of deep-sea mining. Marine Policy, 140(??):??, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200121X.

Arnason:2023:AIS

[AR23a] Ragnar Arnason and Birgir Th. Runolfsson. Assessing the impacts of strong user rights in fisheries: the case of the Philippines marine fisheries. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2300129X.

Arnason:2023:DIS

[AR23b] Ragnar Arnason and Birgir Th. Runolfsson. Delineating the impacts of strong user rights in fisheries. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23001938.

Albrecht:2025:IBT

[ARP25] Moritz Albrecht, Rijal Ramdani, and Kevin Arya Pranaja. Imaginaries of blue transformations: Just seaweed narratives in European Union and Indonesian seaweed farming policies. Marine Policy, 178(??):??, August 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001435.

Arthur:2020:SSF

[Art20] Robert Ian Arthur. Small-scale fisheries management and the problem of open access. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19307043.

Arias:2023:MPI

[AS23a] Natalia M. Arias and Wolfgang B. Stotz. Management policies impact fishers’ income beyond tolerable limits: the case of the Chilean TURFs in Central Chile. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003676.

Asgarian:2023:PPF

[AS23b] Ali Asgarian and Alireza Soffianian. Past and potential future distribution of white mangroves in an arid estuarine environment: Integration of Maxent and CA-Markov models. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200392X.

Anariba:2025:PGE

[ASC25] Sara E. Bonilla Anariba, Arie Sanders, and Steven W. J. Canty. Promoting gender equity in marine protected areas: a self-assessment tool. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24005268.

Atalay:2024:SPG

[ASG^{+}24] M. A. Atalay, H. Sevgili, D. Giannetto, G. Sarà, M. Kanyılmaz, S. Erkan, and M. C. Mangano. A stakeholders’ perception: Turkish aquaculture sector under COVID-19 pandemic effect with consideration of anthropogenic stressors. Marine Policy, 164(??):??, June 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001519.

Al-Subhi:2023:EON

[ASGS23] Khalid Al-Subhi, Tim Gray, and Selina M. Stead. An evaluation of Oman’s National Fisheries Development Strategy (NFDS) 2013–2020: Wealth or welfare maximisation? Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002142.

Asche:2022:LSF

[ASGZ22] Frank Asche, Geir Sogn-Grundvåg, and Dengjun Zhang. Large-scale fisheries during the COVID-19 pandemic: the case of the oceangoing groundfish fleet in Norway. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002706.

Ahmmed:2020:NAM

[ASKI20] Raju Ahmmed, Begum Shahnaz Sinha, Rubina Khan, and Dewan Mazharul Islam. A needs analysis of maritime English language skills for Bangladeshi seafarers to work onboard ships. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301044.

Amorim:2021:SSG

[ASM21a] Patrícia Amorim, Pedro Sousa, and Gui M. Menezes. Sustainability status of the grouper fishery in the Azores archipelago: a length-based approach. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21001731.

Appleby:2021:SPO

[ASM^{+}21b] Thomas Appleby, Matthew Studley, Brendon Moorhouse, Judith Brown, Chad Staddon, and Emma Bean. Sea of possibilities: Old and new uses of remote sensing data for the enforcement of the Ascension Island marine protected area. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18300289.

Adiputra:2024:CNU

[ASPT24] Yudha Trinoegraha Adiputra, Suadi, Ayu Pratiwi, and Andi Sandi Antonius Tabusassa Tonralipu. Caught in the net: Unravelling policy challenges and smuggling dynamics in Indonesia’s puerulus exploitation. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003348.

Aguado:2021:QFG

[ASV^{+}21] S. H. Aguado, I. S. Segado, M. E. S. Vidal, T. J. Pitcher, and M. E. Lam. The quality of fisheries governance assessed using a participatory, multi-criteria framework: a case study from Murcia, Spain. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2030926X.

Armelloni:2021:ADM

[ATF^{+}21] Enrico Nicola Armelloni, Anna Nora Tassetti, Carmen Ferrà, Alessandro Galdelli, Martina Scanu, Adriano Mancini, Gianna Fabi, and Giuseppe Scarcella. AIS data, a mine of information on trawling fleet mobility in the Mediterranean Sea. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001822.

Avila-Thieme:2025:ENI

[ÁTGCE^{+}25] M. Isidora Ávila-Thieme, Catalina A. Galaz-Catalán, Rodrigo A. Estévez, Pablo Romero, Josh C. Donlan, and Stefan Gelcich. Exploring noncompliance in the intertidal kelp fishery Lessonia berteroana of northern Chile. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005748.

Ahmed:2024:RLF

[ATM^{+}24] Yaseen Adnan Ahmed, Gerasimos Theotokatos, Ilia Maslov, Lars Andreas Lien Wennersberg, and Dag Atle Nesheim. Regulatory and legal frameworks recommendations for short sea shipping maritime autonomous surface ships. Marine Policy, 166(??):??, August 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002240.

Alessi:2024:SRA

[AVC24] Jessica Alessi, Alessandra Vanacore, and Elena Cicoria. Spatial risk analysis of floating marine macro litter along the Western Sicilian coast: Overlap with bottlenose dolphin and seabird hotspots. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24002720.

Akuoko:2023:RPR

[AVF^{+}23a] Ivy Serwaa Gyimah Akuoko, Jessica Vandenberg, Jill C. Falman, Karin Otsuka, Gabriel K. Lau, Marlena Skrobe, Suzy An, Elaine M. Faustman, and Yoshitaka Ota. Rethinking plastic realities in Ghana: a call for a well-being approach to understanding human–plastics entanglements for more equitable plastics governance. Marine Policy, 158(??): ??, December 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003895.

Arcas:2023:EEC

[AVF23b] Esther Arcas, Carlos Valle, and Aitor Forcada. Evaluation of the economic crisis on the conservation of the ichthyofauna in Marine Protected Areas. Marine Policy, 147(??): ??, January 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003943.

Angulo-Valdes:2022:MMR

[AVPAFM^{+}22] Jorge Angulo-Valdes, Fabián Pina-Amargos, Tamara Figueredo-Martin, Rod Fujita, Sepp Haukebo, Valerie Miller, Eduardo Boné-Morón, and Daniel Whittle. Managing marine recreational fisheries in Cuba for sustainability and economic development with emphasis on the tourism sector. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003013.

Adam:2020:PRS

[AWBC20] Issahaku Adam, Tony R. Walker, Joana Carlos Bezerra, and Andrea Clayton. Policies to reduce single-use plastic marine pollution in West Africa. Marine Policy, 116(??): ??, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19304865.

Abbott:2021:FDE

[AWX21] Joshua K. Abbott, Daniel Willard, and Jintao Xu. Feeding the dragon: the evolution of China’s fishery imports. Marine Policy, 133(??):??, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003444.

Bogen:2025:CFC

[BÅ25a] Ellinor Bogen and Heidrun Åm. Closing facilities or closing debates? Knowledge politics in sustainable aquaculture governance. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001897.

Brahmana:2025:ICP

[BA25b] Rayenda Khresna Brahmana and Adnan Aslam. The impact of climate policy uncertainty on the dynamic connectedness among leading fish-exporting nations. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001757.

Boothroyd:2024:PAM

[BAAH24a] Anne Boothroyd, Vanessa Adams, Karen Alexander, and Nicole Hill. Priority areas for marine protection in the Amundsen and Bellingshausen Seas, Antarctica. Marine Policy, 167(??):??, September 2024. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002306.

Boothroyd:2024:SPC

[BAAH24b] Anne Boothroyd, Vanessa Adams, Karen Alexander, and Nicole Hill. Stakeholder perceptions of the Commission for the Conservation of Antarctic Marine Living Resources Marine Protected Area planning process. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003798.

Botta:2020:RGO

[BABC20] Robert Botta, Frank Asche, J. Scott Borsum, and Edward V. Camp. A review of global oyster aquaculture production and consumption. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305433.

Byrne:2021:FII

[BAD21] Conor Byrne, Sveinn Agnarsson, and Brynhildur Davidsdottir. Fuel Intensity in Icelandic fisheries and opportunities to reduce emissions. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000580.

Byrne:2020:SLQ

[BADO20] Conor Byrne, Sveinn Agnarsson, Brynhildur Davidsdottir, and Maartje Oostdijk. Species-level quota concentration in the Icelandic harvesting sector. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20303110.

Bennett:2023:EJA

[BAF^{+}23] Nathan J. Bennett, Juan José Alava, Caroline E. Ferguson, Jessica Blythe, Elisa Morgera, David Boyd, and Isabelle M. Côté. Environmental (in)justice in the Anthropocene ocean. Marine Policy, 147(??):??, January 2023.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004304.

Boucek:2024:WFA

[BAJR24] Ross E. Boucek, Kristin A. Anderson, Benjamin L. Jones, and Jennifer S. Rehage. When fishers ask for more protection: Co-produced spatial management recommendations to protect seagrass meadows from leisure boating. Marine Policy, 167(??):??, September 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002252.

Balton:2022:WWB

[Bal22] David Balton. What will the BBNJ agreement mean for the Arctic fisheries agreement? Marine Policy, 142(??): ??, August 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X1930449X.

Baumler:2021:QII

[BAP21] Raphael Baumler, Maria Carrera Arce, and Anne Pazaver. Quantification of influence and interest at IMO in Maritime Safety and Human Element matters. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003572.

Braga:2022:CCT

[BASM22] Heitor O. Braga, Ulisses M. Azeiteiro, Alexandre Schiavetti, and Luísa Magalhães. Checking the changes over time and the impacts of COVID-19 on cockle (Cerastoderma edule) small-scale fisheries in Ria de Aveiro coastal lagoon, Portugal. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004541.

Baumler:2020:WTL

[Bau20] Raphael Baumler. Working time limits at sea, a hundredyear construction. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20301998.

Baumann:2024:GHF

[Bau24] Jennifer Baumann. From global to human factors: Shipping emissions policy process in the IMO. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002896.

Benzaken:2024:CPF

[BAVV24] Dominique Benzaken, Jean Paul Adam, John Virdin, and Michelle Voyer. From concept to practice: financing sustainable blue economy in Small Island Developing States, lessons learnt from the Seychelles experience. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000708.

Bonnett:2020:CCC

[BB20] Nicole Bonnett and S. Jeff Birchall. Coastal communities in the Circumpolar North and the need for sustainable climate adaptation approaches. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2030302X.

Bastardie:2021:RDC

[BB21] Francois Bastardie and Elliot John Brown. Reverse the declining course: a risk assessment for marine and fisheries policy strategies in Europe from current knowledge synthesis. Marine Policy, 126(??):??, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000191.

Bostedt:2020:BCT

[BBB^{+}20] Göran Bostedt, Charlotte Berkström, Runar Brännlund, Ola Carlén, Ann-Britt Florin, Lars Persson, and Ulf Bergström. Benefits and costs of two temporary notake zones. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308108.

Barrett:2023:CCM

[BBB23] C. J. Barrett, K. Bradley, and A. Brazier. Common, curled or miscellaneous: the need for species-specific recordings of octopuses to inform stock assessments. Marine Policy, 153 (??):??, July 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001598.

Beaty:2024:DIL

[BBB^{+}24] Fiona Beaty, Kil Hltaanuwaay Tayler Brown, Julien Braun, Steve Diggon, Emery Hartley, Aaron Heidt, Heather Maddin, Avery Maloney, Rebecca Martone, Chris Mc-Dougall, Mike Reid, Carrie Robb, Emily Rubidge, Charles Short, and Kristin Worsley. From design to implementation: Lessons from planning the first marine protected area network in Canada. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003580.

Baker:2025:HWB

[BBB25] Dana Baker, Nathan J. Bennett, and Natalie C. Ban. Human well-being outcomes of large-scale marine protected areas. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005797.

Boys:2022:WHS

[BBBS22] Rebecca M. Boys, Ngaio J. Beausoleil, Emma L. Betty, and Karen A. Stockin. When and how to say goodbye: an analysis of Standard Operating Procedures that guide end-of-life decision-making for stranded cetaceans in Australasia. Marine Policy, 138(??):??, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005601.

Bartelet:2024:ECD

[BBC24] Henry A. Bartelet, Michele L. Barnes, and Graeme S. Cumming. Estimating and comparing the direct economic contributions of reef fisheries and tourism in the Asia– Pacific. Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004724.

Bakit:2024:SAG

[BBFA^{+}24] José Bakit, Valeria Burgos-Fuster, Alejandro Abarca, Ivonne Etchepare, Juan Enrique Illanes, Sebastian Villasante, Enzo Bonilla, Rodrigo Rojas, Benjamin Dudouet, and Niris Cortés. Scallop aquaculture growth: Four decades of economic policy in Chile. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24001374.

Brooks:2022:TCM

[BBGG22] Kate Brooks, Kate Barclay, R. Quentin Grafton, and Natalie Gollan. Transforming coastal and marine management: Deliberative democracy and integrated management in New South Wales, Australia. Marine Policy, 139(??): ??, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20301470.

Barton:2023:NID

[BBGHR23] Jonathan R. Barton, Sebastián Baeza-González, Joaquín Zenteno Hopp, and Álvaro Román. Neostructural innovation and directionality in Chilean salmon aquaculture. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000453.

Braccini:2020:WES

[BBHN20] Matias Braccini, Nick Blay, Alastair Harry, and Stephen J. Newman. Would ending shark meat consumption in Australia contribute to the conservation of white sharks in South Africa? Marine Policy, 120(??):??, October 2020.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20306370.

Bonito:2022:UEF

[BBJ^{+}22] Lindsay Bonito, Lyall Bellquist, Alexis M. Jackson, Kate Kauer, Mary G. Gleason, Jono Wilson, and Stuart Sandin. U.S. exempted fishing permits: Role, value, and lessons learned for adaptive fisheries management. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000392.

Baumler:2021:SFS

[BBK21a] Raphael Baumler, Bikram Singh Bhatia, and Momoko Kitada. Ship first: Seafarers’ adjustment of records on work and rest hours. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309388.

Brooks:2021:RSA

[BBK^{+}21b] Cassandra M. Brooks, Evan Bloom, Andrea Kavanagh, Emily S. Nocito, George M. Watters, and John Weller. The Ross Sea, Antarctica: a highly protected MPA in international waters. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004061.

Bottema:2021:AAS

[BBO21] Mariska J. M. Bottema, Simon R. Bush, and Peter Oosterveer. Assuring aquaculture sustainability beyond the farm. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002694.

Baral:2025:NDB

[BBP25] Kabita Baral, Ramakrushna Baral, and Bagmi Priyadarshani. Navigating the discourse: a bibliometric overview of research on fisherwomen in traditional marine fisheries. Marine Policy, 180(??):??, October 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2500212X.

Bennett:2021:ARC

[BBS^{+}21] Nathan J. Bennett, Natalie C. Ban, Anna Schuhbauer, Dacotah-Victoria Splichalova, Megan Eadie, Kiera Vandeborne, Jim McIsaac, Eric Angel, Joshua Charleson, Erika R. Gavenus, Sarah Harper, Terre Satterfield, Tasha Sutcliffe, and Rashid Sumaila. Access rights, capacities and benefits in small-scale fisheries: Insights from the Pacific Coast of Canada. Marine Policy, 130(??):??, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001925.

Bond:2025:TRD

[BBT^{+}25] Mark E. Bond, Hollie Booth, Akshay Tanna, Sarah L. Fowler, Carlos J. Polo-Silva, K. H. Stanley Shea, Floriane Cardiec, Elisabeth F. Mansur, and Rima W. Jabado. Trade regulations drive improved global shark and ray management. Marine Policy, 180(??):??, October 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001484.

Bakit:2025:PTO

[BBV^{+}25] José Bakit, Enzo Bonilla, Sebastian Villasante, Niris Cortés, and Begoña Peceño. Prevent or treat occupational noise exposure in artisanal fishers? Decision-making based on public expenditure. Marine Policy, 174(??):??, April 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000338.

Bennett:2021:BGB

[BBWC21] Nathan James Bennett, Jessica Blythe, Carole Sandrine White, and Cecilia Campero. Blue growth and blue justice: Ten risks and solutions for the ocean economy. Marine Policy, 125(??):??, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310381.

Baker:2020:EJI

[BC20a] Susan Baker and Natasha Louise Constant. Epistemic justice and the integration of local ecological knowledge for marine conservation: Lessons from the Seychelles. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304920.

Bisack:2020:CMM

[BC20b] Kathryn Bisack and Patricia M. Clay. Compliance with marine mammal protection: Focus groups reveal factors in commercial fishermen’s decisions. Marine Policy, 115(??): ??, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19301794.

Bisack:2021:BRC

[BC21] Kathryn D. Bisack and Patricia M. Clay. Behavioral responses to competing incentives and disincentives: Compliance with marine mammal protection. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002852.

Bhatia:2024:SVP

[BCABG24] Bikram Singh Bhatia, Maria Carrera-Arce, Raphael Baumler, and Michelle Rita Grech. Seafarers vs. Port State Control: Decoding work/rest compliance data disparity. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001039.

Berkowitz:2020:OPS

[BCB20] Heloise Berkowitz, Larry B. Crowder, and Cassandra M. Brooks. Organizational perspectives on sustainable ocean governance: a multi-stakeholder, meta-organization model of collective action. Marine Policy, 118(??):??, August 2020.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305159.

Ba:2022:SSR

[BCBS22] Aliou Ba, Christian Chaboud, Patrice Brehmer, and Jörn O. Schmidt. Are subsidies still relevant in West African artisanal small pelagic fishery? Insights from long run bioeconomic scenarios. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003414.

Bennett:2024:AWB

[BCG^{+}24] Nathan J. Bennett, Antonio Calò, Paolo Guidetti, Marco Milazzo, Giulia Prato, Emna Ben Lamine, Claudia Scianna, and Antonio Di Franco. Access and well-being in smallscale fisheries. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003269.

Birchenough:2021:VMS

[BCJ21] Sarah Elizabeth Birchenough, Patrick Arthur Cooper, and Antony Charles Jensen. Vessel monitoring systems as a tool for mapping fishing effort for a small inshore fishery operating within a marine protected area. Marine Policy, 124(??): ??, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309726.

Brotz:2021:RJW

[BCMCM21] Lucas Brotz, Andrés M. Cisneros-Montemayor, and Miguel Ángel Cisneros-Mata. The race for jellyfish: Winners and losers in Mexico’s Gulf of California. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003869.

Baker:2023:OJT

[BCN23] Susan Baker, Natasha Constant, and Poppy Nicol. Oceans justice: Trade-offs between sustainable development goals in the Seychelles. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004043.

Beuret:2025:GMP

[BCR25] Jean-Eudes Beuret, Anne Cadoret, and Newton José Rodrigues Da Silva. A generic marine protected area model, challenged by indigenous peoples and local communities: How can the model be adapted or reinvented? Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25002684.

Blayac:2024:NIS

[BCS^{+}24] Thierry Blayac, Pierre Courtois, Lucille Sevaux, Hélène Rey-Valette, Anais Page, Nicole Lautrédou-Audouy, Jean-Michel Salles, and Frédérique Viard. Non-indigenous species and ecological degradations in marinas: Perceptions and willingness to pay for improvements. Marine Policy, 166(??):??, August 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002227.

Bachman:2025:UCA

[BCW^{+}25] Michelle Bachman, Jessica Coakley, David Witherell, Deirdre Boelke, Mark Fitchett, John Froeschke, Kerry Griffin, Brett Holycross, Roger Pugliese, Eric Reid, and Liajay Rivera. Use of conservation areas for fisheries management and ecosystem conservation in the U.S. exclusive economic zone. Marine Policy, 175(??):??, May 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2500048X.

Ballinger:2020:SMP

[BD20] Rhoda C. Ballinger and Wendy Dodds. Shoreline management plans in England and Wales: a scientific and transparent process? Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17301422.

Bergseth:2023:CLA

[BD23] Brock Bergseth and Jon C. Day. Compliance — the ‘Achilles heel’ of protected areas. Marine Policy, 155(??): ??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002610.

Brown:2021:EMI

[BDC^{+}21] Christopher J. Brown, Amelia Desbiens, Max D. Campbell, Edward T. Game, Eric Gilman, Richard J. Hamilton, Craig Heberer, David Itano, and Kydd Pollock. Electronic monitoring for improved accountability in western Pacific tuna longline fisheries. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100275X.

Ballesteros:2023:MPA

[BDC23] Marta Ballesteros and Mark Dickey-Collas. Managing participation across boundaries: a typology for stakeholder engagement in the International Council for the Exploration of the Sea. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004365.

BenArfa:2022:SEA

[BDCJ22] Yessine Ben Arfa, Antonio Di Cintio, Luca Ceriola, and Othmen Jarboui. Socio-economic analysis of the trawl fleet targeting deep-water rose shrimp (Parapenaeus longirostris) and European hake (Merluccius merluccius) in North Tunisia (2015–2017). Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005637.

Bloor:2021:BBC

[BDD^{+}21] Isobel S. M. Bloor, Peter F. Duncan, Samuel P. Dignan, Jack Emmerson, David Beard, Fiona R. Gell, Karen McHarg, and Michel J. Kaiser. Boom not bust: Cooperative management as a mechanism for improving the commercial efficiency and environmental outcomes of regional scallop fisheries. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002608.

Bertheussen:2020:PDB

[BDDE20] Bernt Arne Bertheussen, Bent Magne Dreyer, Sofie Dreyer, and Susann Evenseth. Performance differences between nations exploiting a common natural resource: the Icelandic– Norwegian mackerel case. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309155.

Bellanger:2025:PFE

[BDG^{+}25] Manuel Bellanger, Benjamin Dudouet, Sophie Gourguet, Olivier Thébaud, Lisa T. Ballance, Nicolas Becu, Kathryn D. Bisack, Annie Cudennec, Fabienne Daurès, Sigrid Lehuta, Rebecca Lent, C. Tara Marshall, David Reid, Vincent Ridoux, Dale Squires, and Clara Ulrich. A practical framework to evaluate the feasibility of incentive-based approaches to reduce bycatch of marine mammals and other protected species. Marine Policy, 177(??):??, July 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000764.

Bertheussen:2021:IFE

[BDHI21] Bernt Arne Bertheussen, Bent Magne Dreyer, Øystein Hermansen, and John Roald Isaksen. Institutional and financial entry barriers in a fishery. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309489.

Bowlby:2024:GHP

[BDL^{+}24] Heather D. Bowlby, Jean-Noël Druon, Jon Lopez, Maria José Juan-Jordá, María Teresa Carreón-Zapiain, Frederic Vandeperre, Agostino Leone, Brittany Finucci, Philippe S. Sabarros, Barbara A. Block, Haritz Arrizabalaga, Pedro Afonso, Michael K. Musyl, Rui Coelho, Enric Cortés, Luis Gustavo Cardoso, Bruno Mourato, Nuno Queiroz, Jorge Fontes, Francisco J. Abascal, Antonella Zanzi, Humberto Gomes Hazin, Pascal Bach, David W. Sims, and Paulo Travassos. Global habitat predictions to inform spatiotemporal fisheries management: Initial steps within the framework. Marine Policy, 164(??):??, June 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001532.

Bernardo:2020:LCA

[BdLAdC^{+}20] Cristina Bernardo, Aisni Mayumi Corrêa de Lima Adachi, Vanessa Paes da Cruz, Fausto Foresti, Robin H. Loose, and Hugo Bornatowski. The label “Cação” is a shark or a ray and can be a threatened species! Elasmobranch trade in Southern Brazil unveiled by DNA barcoding. Marine Policy, 116(??):??, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309261.

Baraibar-Diez:2023:TEA

[BDLO23] Elisa Baraibar-Diez, Ignacio Llorente, and María D. Odriozola. Text emotion analysis in aquaculture communication via Twitter: the case of Spain. Marine Policy, 152(??): ??, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2300132X.

Barreto:2020:HDM

[BDM20] Giovanna C. Barreto, Maikon Di Domenico, and Rodrigo Pereira Medeiros. Human dimensions of marine protected areas and small-scale fisheries management: a review of the interpretations. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301381.

Britton:2021:ASO

[BDM21] Easkey Britton, Christine Domegan, and Patricia McHugh. Accelerating sustainable ocean policy: the dynamics of multiple stakeholder priorities and actions for oceans and human health. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309829.

Bailey:2022:HNS

[BDMW22] Sabine Bailey, Kelly Heber Dunning, Daniel Morris, and Ryan D. Williamson. How narratives shape policy: Lessons learned from port projects adjacent to coral reefs in Florida and the Cayman Islands. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002809.

Bertheussen:2020:ESQ

[BDR20] Bernt Arne Bertheussen, Bent Dreyer, and Audun Reiertsen. Economic sustainability of quality-enhancing business models in the Norwegian cod industry. Marine Policy, 113 (??):??, March 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19307778.

Barcellos:2025:HCC

[BdSL25] Laís R. Barcellos, Thaísy E. F. da Silva, and Rosângela P. T. Lessa. Historical changes in conservation measurements on target, predictable bycatch and bycatch species caught by pelagic longline fisheries in Southwest Atlantic Ocean. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005670.

Baez:2025:IST

[BDTD25] José C. Báez, Andrés Domingo, Davinia Torreblanca, and Ignacio Doadrio. Initiative to stop trade of shark fins within the EU: Is it a good thing for sharks? Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001319.

Backler:2025:ECR

[BDVR25] Scott Backler, K. Sian Davies-Vollum, and Debadayita Raha. An evaluation of coastal resilience policies in Anglophone West Africa through a regional comparative analysis. Marine Policy, 176(??):??, June 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000636.

Bailey:2020:ISS

[BE20a] Jennifer L. Bailey and Sigrid Sandve Eggereide. Indicating sustainable salmon farming: the case of the new Norwegian aquaculture management scheme. Marine Policy, 117(??): ??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X1930452X.

Bailey:2020:MAA

[BE20b] Jennifer L. Bailey and Sigrid Sandve Eggereide. Mapping actors and arguments in the Norwegian aquaculture debate. Marine Policy, 115(??):??, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302702.

Bueger:2020:BCC

[BE20c] Christian Bueger and Timothy Edmunds. Blue crime: Conceptualising transnational organised crime at sea. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20300270.

Bean:2022:HHR

[Bea22] Emma Bean. Helping or hindering? The role of the public right to fish in the management of coastal fisheries. Marine Policy, 138(??):??, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000380.

Breen:2021:ICN

[BEH^{+}21] Colin Breen, Crystal El Safadi, Harmen Huigens, Sophie Tews, Kieran Westley, Georgia Andreou, Rodrigo Ortiz Vazquez, Julia Nikolaus, and Lucy Blue. Integrating cultural and natural heritage approaches to Marine Protected Areas in the MENA region. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002876.

Barz:2020:BDF

[BEM^{+}20] Fanny Barz, Josefa Eckardt, Steffi Meyer, Sarah B. M. Kraak, and Harry V. Strehlow. ‘Boats don’t fish, people do’ — how fishers’ agency can inform fisheriesmanagement on bycatch mitigation of marine mammals and sea birds. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309143.

Bertheussen:2022:RPD

[Ber22] Bernt Arne Bertheussen. The role of path–dependent institutions during the collapse and rebuilding of a fishery. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005558.

Bertram:2023:PSS

[Ber23] Douglas F. Bertram. Pacific Salmon Strategy Initiative will conserve marine birds with fishery closures and gillnet license retirements. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000787.

Beringen:2024:UUU

[Ber24] Ethan Beringen. Uncertain, unstable and unequal: Can regime interaction help international fisheries law address Anthropocene challenges? Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24001568.

Berman:2025:LEG

[Ber25] Martha Berman. Leveraging EO 14801 to grow the U.S. seaweed industry. Marine Policy, 177(??):??, July 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000983.

Bethel:2025:SFG

[Bet25] Lonna Bethel. Strengthening fisheries governance in the Caribbean through transshipment regulations to combat illegal, unreported and unregulated fishing: the intersection of Caribbean regional fishery bodies and global standards. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2500226X. See corrigen-

dum [BJ25].

Barbosa-Filho:2025:FAC

[BFdSdFL^{+}25] Márcio L. V. Barbosa-Filho, Gabriel Barros Gonçalves de Souza, Sérgio de Faria Lopes, Salvatore Siciliano, Rachel Ann Hauser-Davis, and José da Silva Mourão. Fisher awareness is crucial for compliance with regulatory frameworks concerning endangered marine species. Marine Policy, 177(??):??, July 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001034.

Bjorndal:2021:BCQ

[BFMS21] Trond Bjørndal, Torben Foss, Gordon R. Munro, and Mogens Schou. Brexit and consequences for quota sharing in the Barents Sea cod fishery. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002335.

Bertelsen:2025:UNR

[BFSP25] Sissel Kolls Bertelsen, Jordan P. Feekings, and Marie Storr-Paulsen. Untangling the net: Redefining fisheries management with fully documented fisheries. Marine Policy, 177 (??):??, July 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25000995.

Bagoulla:2020:MTF

[BG20a] Corinne Bagoulla and Patrice Guillotreau. Maritime transport in the French economy and its impact on air pollution: an input-output analysis. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305408.

Ballinger:2020:ERC

[BG20b] Rhoda C. Ballinger and Jeremy Gault. Editorial: Reflections on coastal climate change adaptation in North West Europe. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309679.

Baameur:2025:EMB

[BG25] Wahiba Tiziri Baameur and Samir Grimes. Enhancing marine biodiversity conservation through maritime spatial planning. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002404.

Bering:2022:WUA

[BGK^{+}22] Janet Bering, Henry Gargan, Jess Kuesel, Maria Morrison, Claire Mullaney, Andrew J. Read, Stephen E. Roady, and Addyson Rowe. Will unilateral action improve the global conservation status of marine mammals? A first analysis of the U.S. Marine Mammal Protection Act’s Import Provisions Rule. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004437.

Barreiro-Gen:2021:GES

[BGLTC21] María Barreiro-Gen, Rodrigo Lozano, Melis Temel, and Angela Carpenter. Gender equality for sustainability in ports: Developing a framework. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002049.

Braham:2024:RCS

[BGMCG24] Cheikh-Baye Braham, Jo Gascoigne, Carlos Montero-Castaño, and Christophe Godard. The role of certification standards in supporting improved fisheries management for small pelagics in Mauritania: an example to follow in the North West African zone? Marine Policy, 161(??): ??, March 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000526.

Borriello:2025:RAE

[BGQ25] Antonio Borriello, Jordi Guillen, and Simone Quatrini. Remuneration and attractiveness of EU blue economy sectors. Marine Policy, 178(??):??, August 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001356.

Bras:2024:TAA

[BGQL24] Quentin Le Bras, Didier Gascuel, Florian Quemper, and Harold Levrel. Transition and adaptation: an analysis of how professional fishermen change their practices. Marine Policy, 164(??):??, June 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001520.

Bach:2023:ICD

[BH23a] Hanna Bach and Teis Hansen. IMO off course for decarbonisation of shipping? Three challenges for stricter policy. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004262.

Bailey:2023:SUO

[BH23b] David M. Bailey and Charlotte R. Hopkins. Sustainable use of ocean resources. Marine Policy, 154(??):??, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001999.

Brown:2024:PTP

[BHB^{+}24] A. Ross Brown, Tim Huntington, Lisa K. Bickley, Tom Powell, and Charles R. Tyler. Positive tipping points for the sustainable growth of bivalve shellfish mariculture in England and Wales. Marine Policy, 162(??):??, April 2024.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000563.

Bullock:2023:ISP

[BHCL23] Simon Bullock, Eliott Higgins, James Crossan, and Alice Larkin. Improving shore power project economics at the Port of Aberdeen. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001525.

Ben-Hasan:2021:CFM

[BHdMCM^{+}21] Abdulrahman Ben-Hasan, Yvonne Sadovy de Mitcheson, Miguel A. Cisneros-Mata, Érica Antunes Jimenez, Moslem Daliri, Andrés M. Cisneros-Montemayor, Rekha J. Nair, Sangeetha A. Thankappan, Carl J. Walters, and Villy Christensen. China’s fish maw demand and its implications for fisheries in source countries. Marine Policy, 132(??): ??, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003079.

Bastug:2022:PCD

[BHEE22] Sedat Baştuğ, Hercules Haralambides, Soner Esmer, and Enes Eminoğlu. Port competitiveness: Do container terminal operators and liner shipping companies see eye to eye? Marine Policy, 135(??):??, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004772.

Buckwell:2025:SLS

[BHK^{+}25] Andrew Buckwell, Syezlin Hasan, Alana Knight, Christopher Fleming, Jeremy Harte, and James C. R. Smart. The social licence of salmon aquaculture in Tasmania: Fulfilling the social bargain. Marine Policy, 176(??):??, June 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000685.

Blanchard:2023:CSD

[BHKJT23] Catherine Blanchard, Ellycia Harrould-Kolieb, Emily Jones, and Michelle L. Taylor. The current status of deepsea mining governance at the International Seabed Authority. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004432.

Benediktsdottir:2023:AFS

[BHKK23] Brynhildur Benediktsdóttir, Gunnar Haraldsson, Helga Kristjánsdóttir, and Ögmundur Knútsson. Assessing fisheries subsidies. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003688.

Blandon:2021:SCS

[BI21] Abigayil Blandon and Hiroe Ishihara. Seafood certification schemes in Japan: Examples of challenges and opportunities from three Marine Stewardship Council (MSC) applicants. Marine Policy, 123(??):??, January 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309258.

Birchall:2020:CCA

[Bir20] S. Jeff Birchall. Coastal climate adaptation planning and evolutionary governance: Insights from Homer, Alaska. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18303919.

Basterretxea-Iribar:2022:BMP

[BISSBdlMM22] Imanol Basterretxea-Iribar, Iranzu Sotés, Javier Sanchez-Beaskoetxea, and María de las Mercedes Maruri. Beach management policy analysis concerning safety flag systems in Northern Spain. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002731.

Bethel:2025:CLS

[BJ25] Lonna Bethel and Henning Jessen. Corrigendum to ‘Strengthening fisheries governance in the Caribbean through transshipment regulations to combat illegal, unreported and unregulated fishing: the intersection of Caribbean regional fishery bodies and global standards’ [Marine Policy 180 (2025) 106811]. Marine Policy, 181(??): ??, November 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25002568.  See

[Bet25].

Beaty:2024:CCV

[BJA^{+}24] Fiona Beaty, Bridget John, Myia Antone, Jonathan Williams, Nathan Bennett, Nikita Wallia, and Christopher Harley. Centering community values in marine planning. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003610.

Bausero-Jorcin:2024:APP

[BJGG^{+}24] Sofía Bausero-Jorcin, Stefan Gelcich, Ignacio Gianelli, Gabriela Jorge-Romero, Cecilia Lezama, and Omar Defeo. Assessing the performance of a participatory governance transformation in small-scale fisheries: a case study from Uruguay. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004979.

Bethel:2021:IPS

[BJH21] Lonna Bethel, Henning Jessen, and Johan Hollander. Implementing the Port State Measures Agreement to combat illegal, unreported and unregulated fishing in the Caribbean. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002542.

Bjorndal:2023:NAM

[Bjø23] Trond Bjørndal. The Northeast Atlantic and Mediterranean bluefin tuna fishery: Back from the brink. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003810.

Brahmana:2023:SST

[BK23] Rayenda Khresna Brahmana and Maria Kontesa. Sea surface temperature anomalies and cash holdings: Evidence from fisheries companies. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004997.

Burnham:2025:SIR

[BKG^{+}25] Theresa L. U. Burnham, Joelle Kilchenmann, Carla Guenther, Maggie O’Shea, Kathleen Reardon, and Joshua S. Stoll. Socioeconomic indicators of resilience in Maine’s American lobster fishery. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24005438.

Baker:2024:IAB

[BKW^{+}24] G. Barry Baker, Valeriya Komyakova, Alexia Wellbelove, Nicola Beynon, and Marcus Haward. The implementation of ACAP Best Practice Advice to mitigate seabird bycatch in fisheries: Issues and options. Marine Policy, 160(??): ??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004128.

Boyle:2022:SWE

[BL22] Apryl Boyle and Brianna Le Busque. Sharing the waves: an exploration of surfer and shark interactions. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003074.

Bai:2023:HSI

[BL23a] Jiayu Bai and Xiaoyu Li. How SCAR informs decisionmaking for Antarctic governance: Perspectives from SCAR submissions. Marine Policy, 155(??):??, September 2023.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002907.

Bueger:2023:CMI

[BL23b] Christian Bueger and Tobias Liebetrau. Critical maritime infrastructure protection: What’s the trouble? Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003056.

Baez:2020:MSF

[BLH^{+}20] José Carlos Báez, María Lourdes Ramos, Miguel Herrera, Hilario Murua, José Luis Cort, Santiago Déniz, Vanessa Rojo, Jon Ruiz, Pedro J. Pascual-Alayón, Anertz Muniategi, Alejandro Pérez San Juan, Javier Ariz, Felipe Fernández, and Francisco Abascal. Monitoring of Spanish flagged purse seine fishery targeting tropical tuna in the Indian ocean: Timeline and history. Marine Policy, 119(??): ??, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19307559.

Beringen:2022:APU

[BLL22] Ethan Beringen, Nengye Liu, and Michelle Lim. Australia and the pursuit of “not undermining” regional bodies at the biodiversity beyond national jurisdiction negotiations. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005406.

Bjorndal:2024:CMI

[BLM24] Trond Bjørndal, Marko Lindroos, and Gordon Munro. The cooperative management of internationally shared fish stocks in a world of instability. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004943.

Bonanomi:2025:ARC

[BLM^{+}25] Sara Bonanomi, Simone Libralato, Loretta Malvarosa, Karolina Maria Czechowska, Giulia Sandalli, Alessia Cariani, Alice Ferrari, Ernesto Azzurro, and Giuseppe Scarcella. Applying a regional certification scheme for aquatic invasive species fishery: the case of blue crab (Callinectes sapidus) in the northern Adriatic Sea. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2500274X.

Britsch:2021:DPA

[BLS21] Melissa L. Britsch, Heather M. Leslie, and Joshua S. Stoll. Diverse perspectives on aquaculture development in Maine. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003080.

Bonsu:2024:CLF

[BLY^{+}24] Prince Owusu Bonsu, Jonas Letschert, Katherine L. Yates, Jon C. Svendsen, Jörg Berkenhagen, Marcel J. C. Rozemeijer, Thomas R. H. Kerkhove, Jennifer Rehren, and Vanessa Stelzenmüller. Co-location of fisheries and offshore wind farms: Current practices and enabling conditions in the North Sea. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004748.

Bjorndal:2021:GTP

[BM21] Trond Bjørndal and Gordon R. Munro. A game theoretic perspective on the management of shared North Sea fishery resources: Pre and post Brexit. Marine Policy, 132(??): ??, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002803.

Buscher:2021:DSB

[BMB^{+}21] Elena Buscher, Darcy L. Mathews, Cheryl Bryce, Kathleen Bryce, Darlene Joseph, and Natalie C. Ban. Differences and similarities between Indigenous and conventional marine conservation planning: the case of the Songhees Nation, Canada. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001317.

Barboza:2024:SSF

[BMB^{+}24] Anna Barboza, Edison D. Macusi, Elaine Q. Borazon, Mudjekeewis D. Santos, Richard N. Muallil, and Ivy M. Nallos. Small-scale fisheries (SSF) management and conservation schemes and their application in the Philippines. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000162.

Bernardini:2020:MPP

[BMC20] Giulia Bernardini, A. J. McConville, and Arturo Castillo Castillo. Macro-plastic pollution in the tidal Thames: an analysis of composition and trends for the optimization of data collection. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300397.

Baker-Medard:2020:FMF

[BMF20] Merrill Baker-Médard and Jake Faber. Fins and (Mis)fortunes: Managing shark populations for sustainability and food sovereignty. Marine Policy, 113(??):??, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301617.

Barragan-Mendez:2020:PRA

[BMGDS^{+}20] C. Barragán-Méndez, M. M. González-Duarte, I. Sobrino, Y. Vila, J. M. Mancera, and I. Ruiz-Jarabo. Physiological recovery after bottom trawling as a method to manage discards: the case study of Nephrops norvegicus and Squilla mantis. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308157.

Billi:2022:LCL

[BMH^{+}22] Marco Billi, Aldo Mascareño, Pablo A. Henríquez, Ignacia Rodríguez, Felipe Padilla, and Gonzalo A. Ruz. Learning from crises? The long and winding road of the salmon industry in Chiloé Island, Chile. Marine Policy, 140(??):??, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001166.

Blake:2022:UFL

[BMK^{+}22] Suzana Dumitrita Blake, Matthew McPherson, Mandy Karnauskas, Skyler Rose Sagarese, Adyan Rios, Amanda Daria Stoltz, Anthony Mastitski, and Michael Jepson. Use of fishermen’s local ecological knowledge to understand historic red tide severity patterns. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003001.

Bueno:2025:SSU

[BMK^{+}25] Patrícia Falcão Bueno, Rodrigo Machado, Sarah Longhi Kunzler, Renato Azevedo Matias Silvano, Karina Rejane Groch, and Alexandre Schiavetti. Secrets of the sea: Unraveling the interactions, between artisanal fishing and southern right whales, in Southern Brazil. Marine Policy, 174 (??):??, April 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25000181.

Beveridge:2020:NSE

[BMM^{+}20] Rachelle Beveridge, Megan Moody, Grant Murray, Chris Darimont, and Bernie Pauly. The Nuxalk Sputc (Eulachon) Project: Strengthening indigenous management authority through community-driven research. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19304506.

Bocci:2024:LSI

[BMM^{+}24] Martina Bocci, Marina Markovic, Aleš Mlakar, Margarita Stancheva, Michelle Borg, Fabio Carella, Andrea Barbanti, and Emiliano Ramieri. Land-Sea-Interactions in MSP and ICZM: a regional perspective from the Mediterranean and the Black Sea. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004578.

Baranano:2022:ERS

[BMMF22] Carlota Barañano, Gonzalo Méndez, María Isabel Míguez, and Emilio Fernández. Enhancing recognition of seagrass ecosystems through communication planning based on historical and present social perception studies. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003438.

Baco:2020:OVM

[BMR20] Amy R. Baco, Nicole B. Morgan, and E. Brendan Roark. Observations of vulnerable marine ecosystems and significant adverse impacts on high seas seamounts of the northwestern Hawaiian Ridge and Emperor Seamount Chain. Marine Policy, 115(??):??, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302611.

Botero:2023:YCZ

[BMR23] Camilo M. Botero, Celene B. Milanes, and Sebastian Robledo. 50 years of the Coastal Zone Management Act: the bibliometric influence of the first coastal management law on the world. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000751.

Brown:2023:MSR

[BMRM23] Robert Brown, David Molyneux, Joshua Ryan, and Alexandria Major. Maritime search and rescue in Canada and the use of emergency radio beacons. Marine Policy, 147(??): ??, January 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003086.

Bakowski:2020:TDE

[BN20] Tomasz Bakowski and Justyna Nawrot. Transposition of the directive establishing a framework for maritime spatial planning into the polish legal order (main aspects and key solutions). Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303276.

Bernardino:2021:CTM

[BNF21] Angelo F. Bernardino, Gabriel N. Nóbrega, and Tiago O. Ferreira. Consequences of terminating mangrove’s protection in Brazil. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2031040X.

Bilgili:2024:ISW

[BÖ24] Levent Bilgili and Aykut I. Ölçer. IMO 2023 strategy — where are we and what’s next? Marine Policy, 160(??): ??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004864.

Bartlett:2025:BCR

[BOBMG25] Brittany Bartlett, Rodrigo Oyanedel, Hollie Booth, and E. J. Milner-Gulland. Bringing contextual reality to IUU fishing. Marine Policy, 180(??):??, October 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2500199X.

Badru:2023:SEC

[BOOO23] Gbolahan Badru, Shakirudeen Odunuga, Ademola Omojola, and Emmanuel Oladipo. Stakeholder engagement in coastal adaptation interventions: Insights from the Southwest coast of Nigeria. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002178.

Botha:2020:RLF

[Bot20] Mark Botha. Reorganising the lobster fishery value chain. Marine Policy, 120(??):??, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303781.

Boviatsis:2022:LAB

[Bov22] M. Boviatsis. Legal assessment of BIMCO’s infectious or contagious diseases (IOCD) clauses for voyage and time charter parties. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002536.

Bishop:2022:HIG

[BOW^{+}22] Breanna Bishop, Jade Owen, Lisette Wilson, Tagalik Eccles, Aldo Chircop, and Lucia Fanning. How icebreaking governance interacts with Inuit rights and livelihoods in Nunavut: a policy review. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000045.

Boylan:2021:IMT

[Boy21] Brandon M. Boylan. Increased maritime traffic in the Arctic: Implications for governance of Arctic sea routes. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001779.

Baiju:2022:AGS

[BPA^{+}22] K. K. Baiju, Shinoj Parappurathu, S. Abhilash, C. Ramachandran, P. S. Swathi Lekshmi, Praveen Padmajan, Shelton Padua, and Thomson Kaleekal. Achieving governance synergies through institutional interactions among non-state and state actors in small-scale marine fisheries in India. Marine Policy, 138(??):??, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000379.

Battaglia:2025:MPA

[BPE^{+}25] Maria B. Battaglia, Jeremy Pittman, Graham B. Epstein, Giulia Bernardi, and Derek Armitage. Marine protected areas governance, social norms, and social networks. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005761.

Burns:2024:DBP

[BPJ^{+}24] Polly Burns, Beth Polidoro, Ernesto Jardim, David McElroy, Emily McGregor, and Nikki Wood. Defining best practice in global stakeholder consultations: Lessons learned from the Marine Stewardship Council’s Fishery Standard Review. Marine Policy, 167(??):??, September 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002367.

Borriello:2022:IMO

[BR22] Antonio Borriello and John M. Rose. The issue of microplastic in the oceans: Preferences and willingness to pay to tackle the issue in Australia. Marine Policy, 135(??): ??, January 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004863.

Braithwaite:2022:NLB

[Bra22] Jeremy Braithwaite. This is no longer a Bristol Bay Fishery”: Fisheries dispossession and colonial violence in Bristol Bay, Alaska. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002196.

Bres:2025:CPR

[Bre25] Justyna Breś. Coordination of planning regulations at the edge of water and land. Conflicts within planning policy for urban blue space in small and medium-sized Polish port cities and possible directions of change. Marine Policy, 173 (??):??, March 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24005682.

Block:2025:AEI

[BRKS25] Izabella R. Block, Michele F. Remer, Kelly E. Kapsar, and Jennifer I. Schmidt. Analysis on the effectiveness of International Maritime Organization’s ‘Areas to be Avoided’ policy on vessel traffic in the Bering Strait (2015–2022). Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002659.

Brager:2020:CSE

[BRM20] Stefan Bräger, Gabriela Q. Romero Rodriguez, and Sandor Mulsow. The current status of environmental requirements for deep seabed mining issued by the International Seabed Authority. Marine Policy, 114(??):??, April 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304329.

Belton:2021:CIA

[BRM^{+}21] Ben Belton, Leah Rosen, Lucinda Middleton, Saadiah Ghazali, Abdullah-Al Mamun, Jacqueline Shieh, Hamia S. Noronha, Goutam Dhar, Mohammod Ilyas, Christopher Price, Ahmed Nasr-Allah, Ibrahim Elsira, Bikram K. Baliarsingh, Arun Padiyar, Suresh Rajendran, A. B. C. Mohan, Ravi Babu, Michael Joseph Akester, Ei Ei Phyo, Khin Maung Soe, Ajibola Olaniyi, Sunil N. Siriwardena, John Bostock, David C. Little, Michael Phillips, and Shakuntala H. Thilsted. COVID-19 impacts and adaptations in Asia and Africa’s aquatic food value chains. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001342.

Bosi:2023:MSP

[BRP^{+}23] Sofia Bosi, Emiliano Ramieri, Marta Picciulin, Stefano Menegon, Michol Ghezzo, Antonio Petrizzo, Thomas Folegot, Fantina Madricardo, and Andrea Barbanti. Is Maritime Spatial Planning a tool to mitigate the impacts of underwater noise? A review of adopted and upcoming maritime spatial plans in Europe. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002580.

Bustos:2022:IDS

[BRR^{+}22] Eduardo Bustos, Hugo Robotham, Guillermo Rodríguez, Francisco Ther-Ríos, Jorge Muñoz, Lorenzo Andrade, Nancy. Barahona, and Alejandro Clement. Indicators for the determination of the status and trends of king’s littleneck clam fishery (Ameghinomya antiqua, P. P. King, 1832) in Los Lagos Region, Chile: an integral sustainable approach. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001750.

Bahamon:2024:SBM

[BRSC^{+}24] N. Bahamon, L. Recasens, J. Sala-Coromina, B. Calero, J. A. Garcia, G. Rotllant, A. Maurer, A. Rojas, L. Muth, J. Quevedo, M. Vigo, J. Ribera-Altimir, and J. B. Company. Selectivity-based management for reversing overexploitation of demersal fisheries in North-western Mediterranean Sea. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001830.

Burke:2022:SPW

[BS22] Lauretta Burke and Mark Spalding. Shoreline protection by the world’s coral reefs: Mapping the benefits to people, assets, and infrastructure. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2200358X.

Baeta:2021:LTT

[BSBD21] Marc Baeta, Marco Antonio Solís, Manuel Ballesteros, and Omar Defeo. Long-term trends in striped venus clam (Chamelea gallina) fisheries in the western Mediterranean Sea: the case of Ebro Delta (NE Spain). Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004097.

Bhowmik:2021:RSS

[BSI^{+}21] Joy Bhowmik, Samiya Ahmed Selim, Haseeb Md. Irfanullah, Jannat Shancharika Shuchi, Rumana Sultana, and Shaikh Giasuddin Ahmed. Resilience of small-scale marine fishers of Bangladesh against the COVID-19 pandemic and the 65-day fishing ban. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2100405X.

Brown:2020:PPF

[BSK^{+}20] Stephen K. Brown, Manoj Shivlani, Roberto F. Koeneke, David Agnew, Julia Byrd, Martin Cryer, Cathy Dichmont, David Die, William Michaels, Keith Reid, Jake Rice, Henrik Sparholt, and James Weinberg. Patterns and practices in fisheries assessment peer review systems. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301241.

Ballesteros:2024:EIC

[BSLRR24] Hugo. M. Ballesteros, Eduardo Sánchez-Llamas, and Gonzalo Rodríguez-Rodríguez. Estimating illegal catches in data-poor S-fisheries: Insights from multispecies shellfish poaching in Galician small scale fisheries. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000824.

Banikoi:2023:UTM

[BSM23] Hudu Banikoi, Achim Schlüter, and Aisa O. Manlosa. Understanding transformations in the marine coastal realm: the explanatory potential of theories of institutional change. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2300324X.

Bellido:2020:IVO

[BSSL^{+}20] Jose M. Bellido, U. Rashid Sumaila, Jose Luis Sánchez-Lizaso, Maria Lourdes Palomares, and Daniel Pauly.

Input versus output controls as instruments for fisheries management with a focus on Mediterranean fisheries. Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303720.

Bidwell:2023:ASO

[BST23] David Bidwell, Tiffany Smythe, and Grant Tyler. Anglers’ support for an offshore wind farm: Fishing effects or clean energy symbolism. Marine Policy, 151(??):??, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000957.

Bennett:2020:OTS

[BSY^{+}20] Mia M. Bennett, Scott R. Stephenson, Kang Yang, Michael T. Bravo, and Bert De Jonghe. The opening of the Transpolar Sea Route: Logistical, geopolitical, environmental, and socioeconomic impacts. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2030453X.

Brugere:2021:MTF

[BTE21] Cecile Brugere, Max Troell, and Hampus Eriksson. More than fish: Policy coherence and benefit sharing as necessary conditions for equitable aquaculture development. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309179.

Breen:2024:DIF

[BTN^{+}24] Colin Breen, Sophie Tews, Julia Nikolaus, Nick Ray, Georgia Holly, Georgia Andreou, Chris McGonigle, Kieran Westley, and Lucy Blue. Developing an integrated framework for Marine Cultural Heritage (MCH) and Marine Protected Areas (MPA) across Africa and the Middle East. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002161.

Browne:2022:MEO

[BTV^{+}22] Thomas Browne, Trung Tien Tran, Brian Veitch, Doug Smith, Faisal Khan, and Rocky Taylor. A method for evaluating operational implications of regulatory constraints on Arctic shipping. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004504.

Breve:2024:FWR

[BUM^{+}24] Niels W. P. Brevé, Kateryna Urbanovych, AlberTinka J. Murk, Paul A. M. van Zwieten, Leopold A. J. Nagelkerke, and Marloes Kraan. Fishers’ willingness to report incidental bycatches of endangered, threatened and protected fish species: the case of European sturgeon in the Northeast Atlantic Ocean. Marine Policy, 162(??):??, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400054X.

Begum:2022:APL

[BUR^{+}22] Amany Begum, Md Kutub Uddin, Md Mizanur Rahman, Md Mostafa Shamsuzzaman, and Mohammad Mahmudul Islam. Assessing policy, legal and institutional frameworks of marine megafauna conservation in Bangladesh. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002342.

Bertheussen:2021:IBR

[BV21] Bernt Arne Bertheussen and Terje Vassdal. Institutionbased roots to fishing vessels profitability. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X20309325.

Brander:2020:GCB

[BvBN^{+}20] Luke M. Brander, Pieter van Beukering, Lynn Nijsten, Alistair McVittie, Corinne Baulcomb, Florian V. Eppink, and Jorge Amrit Cado van der Lelij. The global costs and benefits of expanding Marine Protected Areas. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19302386.

Banach:2020:FSD

[BvdBvdFK20] J. L. Banach, S. W. K. van den Burg, and H. J. van der Fels-Klerx. Food safety during seaweed cultivation at offshore wind farms: an exploratory study in the North Sea. Marine Policy, 120(??):??, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301500.

Brugnach:2023:EAC

[BvdH23] Marcela Brugnach and Ronald van den Hoek. Embracing ambiguity in climate change adaptation for more effective responses to new uncertain shorescapes conditions. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001537.

Browne:2021:CMA

[BVT^{+}21] Thomas Browne, Brian Veitch, Rocky Taylor, Jennifer Smith, Doug Smith, and Faisal Khan. Consequence modelling for Arctic ship evacuations using expert knowledge. Marine Policy, 130(??):??, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001937.

Bachmann-Vargas:2021:RFS

[BVvKL21] Pamela Bachmann-Vargas, C. S. A. (Kris) van Koppen, and Machiel Lamers. Re-framing salmon aquaculture in the aftermath of the ISAV crisis in Chile. Marine Policy, 124(??): ??, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20310095.

Bridier:2025:BHE

[BWV25] T. Bridier, H. Wauchope, and J. Vad. Beyond the horizon: Evidence of management deficiencies in UK offshore protected areas. Marine Policy, 179(??):??, September 2025.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001563.

Bertheussen:2020:SIC

[BXV20] Bernt Arne Bertheussen, Jinghua Xie, and Terje Vassdal. Strategic investments in catch capacity and quotas: How costly is a mismatch for a firm? Marine Policy, 117(??): ??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19307092.

Bungwa:2025:SAS

[BYL25] Zainabu Shabani Bungwa, Pius Zebhe Yanda, and James Lyimo. Stakeholders’ analysis for supporting sustainable natural gas development in mainland Tanzania: Implications for coastal forest ecosystem management. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001186.

Carlson:2020:ACS

[CA20] Jordan Thomas Carlson and Michelle Adams. Assessing the consistency of in-stream tidal energy development policy in Nova Scotia, Canada. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1930048X.

Caner:2021:SSS

[CA21] Halil Ibrahim Caner and Cevdet Coşkun Aydin. Shipyard site selection by raster calculation method and AHP in GIS environment, İskenderun, Turkey. Marine Policy, 127(??): ??, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100049X.

Ceylan:2023:NFA

[CAC23] Bulut Ozan Ceylan, Demir Ali Akyar, and Mehmet Serdar Celik. A novel FMEA approach for risk assessment of air pollution from ships. Marine Policy, 150(??):??, April 2023.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000635.

Cardoso-Andrade:2021:ALS

[CACJR^{+}21] Mariana Cardoso-Andrade, Frederico Cruz-Jesus, Francisco Castro Rego, Mafalda Rangel, and Henrique Queiroga. Assessing the land- and seascape determinants of recreational diving: Evidence for Portugal’s south coast. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309313.

Caviedes:2022:ICZ

[CAGBM22] Verónica Caviedes, Pedro Arenas-Granados, and Juan Manuel Barragán-Muñoz. Integrated Coastal Zone Management on a transnational area: the Gulf of Honduras. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100542X.

Cid-Aguayo:2024:SFS

[CAHR^{+}24] Beatriz Eugenia Cid-Aguayo, Alfonso Henríquez, Loreto Arias Ramírez, Marcelo Durán Pérez, Chris Harrod, and Daniel Gomez-Uchida. Socioecology of farmed salmon escapes: ’commons’ and de-privatization of escaped salmon for better management of a potentially invasive species. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004682.

Cevenini:2023:AWI

[CAM^{+}23] Fabio Cevenini, Barnaby Andrews, Angela Muench, Philip Lamb, Silvia Ferrini, and Kieran Hyder. Assessing the welfare impacts of changes in recreational fisheries management: a modelling approach for European sea bass. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004559.

Cannon:2025:UCS

[Can25] Brendon J. Cannon. Undersea cable security in the Indo-Pacific: Enhancing the Quad’s collaborative approach. Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004159.

Cao:2025:FPN

[Cao25] Nga Thi Hong Cao. Fishermen’s perceptions of negative events affecting fishing activities. a case study of a Vietnamese purse seine fishery. Marine Policy, 175(??):??, May 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25000430.

Catma:2020:NMV

[Cat20] Serkan Catma. Non-market valuation of beach quality: Using spatial hedonic price modeling in Hilton Head Island, SC. Marine Policy, 115(??):??, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303100.

Cavole:2020:ULE

[CAVJ^{+}20] Leticia Maria Cavole, Solange Andrade-Vera, José R. Marin Jarrin, Daniela Faggiani Dias, Octavio Aburto-Oropeza, and María José Barrágan-Paladines. Using local ecological knowledge of fishers to infer the impact of climate variability in Galápagos’ small-scale fisheries. Marine Policy, 121(??): ??, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20300439.

Creach:2020:VCA

[CBAPM20] Axel Creach, Emilio Bastidas-Arteaga, Sophie Pardo, and Denis Mercier. Vulnerability and costs of adaptation strategies for housing subjected to flood risks: Application to La Guérinière France. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306303.

Chown:2024:SAI

[CBB^{+}24] Steven L. Chown, Kees Bastmeijer, Cassandra M. Brooks, Neil Gilbert, Laura M. Phillips, and Akiho Shibata. Science advice for international governance — an evidencebased perspective on the role of SCAR in the Antarctic Treaty System. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001416.

Carrasco-Bahamonde:2024:EBD

[CBC24] Daniel Carrasco-Bahamonde and Antònia Casellas. Evolving blue development discourses and policies: Salmon farming industry and regional making in Chile. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2400109X.

Carrasco-Bahamonde:2025:GOS

[CBCA25] Daniel Carrasco-Bahamonde, Antònia Casellas, and Francisco Araos. Getting our sea back: Indigenous governance and biocultural conservation of coastal and marine commons. Marine Policy, 178(??):??, August 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001204.

Cramer:2023:ISP

[CBD^{+}23] Lori A. Cramer, Jennifer Beaullieu, Jamie Doyle, Marta Maldonado, Hillary Egna, Maria Johnson, and Flaxen D. L. Conway. The importance of the seafood processing sector to coastal community resilience. Marine Policy, 156(??): ??, October 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003305.

Cochrane:2024:EBM

[CBH^{+}24a] K. L. Cochrane, D. S. Butterworth, R. Hilborn, A. M. Parma, É. E. Plagányi, and M. P. Sissenwine. Errors and bias in marine conservation and fisheries literature: Their impact on policies and perceptions. Marine Policy, 168(??):

??, October 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003270. See re-

joinder [SCM^{+}25], response [CBH^{+}25], and corrigendum [CBH^{+}24b].

Cochrane:2024:CEB

[CBH^{+}24b] K. L. Cochrane, D. S. Butterworth, R. Hilborn, A. M. Parma, É.E. Plagányi, and M. P. Sissenwine. Corrigendum to “Errors and bias in marine conservation and fisheries literature: Their impact on policies and perceptions” [Marine Policy 168 (2024) 10639]. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003749. See [CBH^{+}24a].

Cochrane:2025:RRC

[CBH^{+}25] K. L. Cochrane, D. S. Butterworth, R. Hilborn, A. M. Parma, É.E. Plagányi, and M. P. Sissenwine. Response to rejoinder to: Cochrane et al., Errors and bias in marine conservation and fisheries literature: Their impact on policies and perceptions [Mar. Policy 168 (2024) 106329]. Marine Policy, 172(??):??, February 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005153. See [SCM^{+}25].

Cret:2024:HII

[CBL24] Laurence Cret, Marc Baudry, and Frédéric Lantz. How to implement the 2023 IMO GHG strategy? Insights on the importance of combining policy instruments and on the role of uncertainty. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003300.

Campos:2021:ULT

[CBO^{+}21] Leonardo Campos, Fernando Berrios, Rómulo Oses, Jorge E. González, and Estefanía Bonnail. Unravelling Lessonia trabeculata management in coastal areas of the Atacama region of northern Chile through a DPSIR approach: Insights for sustainable plans. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003481.

Calderwood:2021:TET

[CBR21] Julia Calderwood, Guillaume Bal, and David G. Reid. Trial and error: Tactical changes in fishing behaviour can help reduce discards and exposure to chokes, but scientific trials can fail to spot this. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310162.

Canales:2021:ESB

[CC21] Cristian M. Canales and Luis A. Cubillos. Empirical survey-based harvest control rules in a transboundary small pelagic fishery under recruitment regime shifts: the case of the northern Chilean–southern Peruvian anchovy. Marine Policy, 134(??):??, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100395X.

Cao:2022:OCC

[CC22] Xingguo Cao and Yen-Chiang Chang. The opening of cabotage: China’s trials and challenges. Marine Policy, 143(??): ??, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002214.

Cao:2023:DLB

[CC23] Xingguo Cao and Yen-Chiang Chang. Developing the legal basis for non-state actors in China’s ocean governance. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002609.

Cardinale:2021:MFM

[CCB^{+}21] Massimiliano Cardinale, Francesco Colloca, Angelo Bonanno, Giuseppe Scarcella, Enrico Arneri, Angélique Jadaud, Claire Saraux, Salvatore Aronica, Simona Genovese, Marco Barra, Gualtiero Basilone, Silvia Angelini, Fabio Falsone, Vita Gancitano, Alberto Santojanni, Fabio Fiorentino, Giacomo Milisenda, Matteo Murenu, Tommaso Russo, Piera Carpi, Beatriz Guijarro, José Luis Pérez Gil, María González, Pedro Torres, Ana Giráldez, Cristina García, Antonio Esteban, Encarnación García, Miguel Vivas, Enric Massutí, Francesc Ordines, Antoni Quetglas, and Juan Gil Herrera. The Mediterranean fishery management: a call for shifting the current paradigm from duplication to synergy. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002232.

Castro-Cadenas:2025:FAW

[CCBB^{+}25] María D. Castro-Cadenas, Martín Barreiros, Maria Bas, Miquel Ortega, Joachim Claudet, Marta Coll, and Valerio Sbragaglia. Fishing activities within Spanish marine protected areas in the Mediterranean Sea. Marine Policy, 177 (??):??, July 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25001010.

Chuang:2021:EKD

[CCCC21] Min-Ta Chuang, Wen-Hsuan Chou, Chia-Hsuan Chang, and Wen-Lun Chou. Examining the key drivers for regional revitalization based upon social network analysis: a case study of Badouzi in Taiwan. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003651.

Correia:2025:PAR

[CCG^{+}25] Maria João Correia, Paula Chainho, Thomas Goulding, Frederico Carvalho, Sara Cabral, Filipa Gomes Ferreira, and Lia Vasconcelos. Participatory action research supporting adaptive governance of Manila clam fisheries. Marine Policy, 174(??):??, April 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2500020X.

Carvalho:2021:PMC

[CCGM21] Natacha Carvalho, John Casey, Jordi Guillen, and Jann Th. Martinsohn. Profitability and management costs in the EU Northeast Atlantic fisheries. Marine Policy, 123(??): ??, January 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309271.

Caputi:2021:SRE

[CCK^{+}21] Nick Caputi, Arani Chandrapavan, Mervi Kangas, Simon de Lestang, Anthony Hart, Danielle Johnston, and James Penn. Stock-recruitment-environment relationships of invertebrate resources in Western Australia and their link to pro-active management harvest strategies. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003390.

Chumchuen:2022:CTM

[CCKK22] Watcharapong Chumchuen, Shiela Villamor Chumchuen, Krisada Kajonrit, and Kraison Krueajun. COVID-19 and Thai marine capture fishery in the Gulf of Thailand: a case of small-scale fishery versus industrial fishery. Marine Policy, 144(??):??, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002871.

Calado:2025:FAC

[CCNGS25] H. Calado, C. Cervera-Núñez, D. Gutierrez, and Stojanovic. Forging ahead: Climate-smart maritime spatial planning for the future. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005037.

Cunha:2025:SAS

[CCOP25] Darliane Ribeiro Cunha, Markus Dannyllo Carneiro Costa, Clóvis Bôsco Mendonça Oliveira, and Newton Narciso Pereira. SDG attributes: a sustainability assessment framework for Brazilian ports. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2500257X.

Chivers:2025:REL

[CCP25] Sophia A. Chivers, Kenneth S. Cline, and Christopher W. Petersen. The roles of ecology, legal structures, and threats in maintaining successful shellfish and river herring comanagement in Maine. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25001666.

Capurso:2023:TMM

[CCS23] Ginevra Capurso, Brendan Carroll, and Kathryn A. Stewart. Transforming marine monitoring: Using eDNA metabarcoding to improve the monitoring of the Mediterranean Marine Protected Areas network. Marine Policy, 156(??):??, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X23003408.

Cronin:2021:PPO

[CCW21] Yvonne Cronin, Valerie Cummins, and Eric Wolsztynski. Public perception of offshore wind farms in Ireland. Marine Policy, 134(??):??, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004255.

Cao:2023:SCT

[CCX23] Xingguo Cao, Leyi Chang, and Houqun Xing. Securing the clean title of ships: the prospects and achievements of the Convention on the International Effects of Judicial Sales of Ships. Marine Policy, 154(??):??, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002129.

Castrejon:2023:RLB

[CD23] Mauricio Castrejón and Omar Defeo. Reconsidering the longline ban in the Galapagos Marine Reserve. Marine Policy, 151(??):??, May 2023. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23001161.

Conley:2024:ASP

[CD24] Elizabeth D. Conley and Emily P. Diamond. Angler’s sense of place as an indicator for perceived vulnerability to shifting stock distributions. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000058.

Castrejon:2025:TEA

[CD25] Mauricio Castrejón and Omar Defeo. Towards an ecosystem approach to fisheries in the Galapagos Marine Reserve: a science-driven and adaptive co-management framework. Marine Policy, 181(??):??, November 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002738.

Crosman:2020:EFM

[CDB20] Katherine M. Crosman, Natalie A. Dowling, and Ann Bostrom. The effects of Fishpath, a multi-stakeholder decision-support tool, on stakeholder buy-in to management in data-limited fisheries. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20308617.

Christiansen:2020:PED

[CDC20a] Bernd Christiansen, Anneke Denda, and Sabine Christiansen. Potential effects of deep seabed mining on pelagic and benthopelagic biota. Marine Policy, 114:??, April 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306407.

Clark:2020:EIA

[CDC20b] Malcolm R. Clark, Jennifer M. Durden, and Sabine Christiansen. Environmental Impact Assessments for deep-sea mining: Can we improve their future effectiveness? Marine Policy, 114:??, April 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307309.

Chen:2024:RIE

[CDC24] Keqi Chen, Qianbin Di, and Xiaolong Chen. The resilience of island economy in the face of adverse shocks and its influencing factors: a case study from county-level islands in China. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001763.

Carabelli:2025:KDC

[CDC^{+}25] A. Carabelli, M. Dal Maso, M. V. Carola, V. Ba, M. Carraturo, I. Milani, R. Bettinetti, and G. Boldrocchi. Knowledge drives conservation: Tackling the shark consumption issue in Italy — a case study. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25002210.

Carlesi:2023:EPM

[CDD^{+}23] Lorenzo Carlesi, Emilia Cubero Dudinskaya, Roberto Danovaro, Gianfranco D’Onghia, Serena Mandolesi, Simona Naspetti, and Raffaele Zanoli. Estimating preferences for Mediterranean deep-sea ecosystem services: a discrete choice experiment. Marine Policy, 151(??):??, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001203.

Chavez:2021:PSF

[CDG^{+}21] Carlos Chávez, Jorge Dresdner, Nuria González, Mauricio Leiva, and Miguel Quiroga. The performance of shared fish stock fisheries under varying institutional and socioeconomic conditions: Evidence from the South Eastern Pacific Anchoveta Fishery. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309088.

Chen:2022:MCE

[CDHY22] Xiaolong Chen, Qianbin Di, Zhiwen Hou, and Zhe Yu. Measurement of carbon emissions from marine fisheries and system dynamics simulation analysis: China’s northern marine economic zone case. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003268.

Christy:2024:CSF

[CdJK24] Des Christy, Edwin B. P. de Jong, and Luuk Knippenberg. The conundrum of sustainable fishery: fishers’ experiences of the exclusionary processes of the EU Common Fisheries Policy. Marine Policy, 164(??):??, June 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001659.

Caputi:2021:SGC

[CdLN^{+}21] Nick Caputi, Simon de Lestang, Stephen J. Newman, Gary Jackson, and Kim Smith. Stakeholder-government collaboration in developing cost-effective fishery-independent surveys in rights-based and co-managed fisheries. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001214.

Camara:2023:SEC

[CDNS23] Fatoumata Camara, Idrissa Diedhiou, Waly Ndiaye, and Alassane Sarr. Socioeconomic effects of the COVID-19 pandemic on artisanal fishing in Senegal: Adaptation strategies of actors in the fishing ports of Dakar and Mbour. Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001604.

Chatterjee:2022:BBB

[CDP22] Poushali Chatterjee, Rajarshi Dasgupta, and Ashis Kumar Paul. Beach beauty in Bengal: Perception of scenery and its implications for coastal management in Purba Medinipur district, eastern India. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000811.

Christodoulou:2022:OLS

[CDRS22] Anastasia Christodoulou, Dimitrios Dalaklis, Peter Raneri, and Rebecca Sheehan. An overview of the legal search and rescue framework and related infrastructure along the Arctic Northeast Passage. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200032X.

Christodoulou:2025:LEA

[CDS^{+}25] Anastasia Christodoulou, Tuan Dong, Alessandro Schönborn, Aykut I. Ölçer, and Dimitrios Dalaklis. Linking the employment of alternative marine fuels to a carbon price for shipping. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004524.

Chang:2024:LBG

[CDY24] Yen-Chiang Chang, Xingyi Duan, and Ling Yang. The legal basis for Chinese vessels navigating in the South China Sea. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002276.

Cruz-Escalona:2024:VCJ

[CEMRMB^{+}24] Víctor Hugo Cruz-Escalona, Yuri Astrid Moncaleano-Rubio, Enrique Morales-Bojórquez, Dana Isela Arizmendi-Rodríguez, Xchel Aurora Pérez-Palafox, Arelly Ornelas-Vargas, Juan Ángel Payán-Alcacio, and Francisco Javier Urcádiz-Cazares. The value chain of the jumbo squid (Dosidicus gigas) in Mexico: Diagnostic, structural analysis, and scenarios. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000976.

Cochrane:2020:DSE

[CES20] Kevern L. Cochrane, Jessica Eggers, and Warwick H. H. Sauer. A diagnosis of the status and effectiveness of marine fisheries management in South Africa based on two representative case studies. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19307006.

Condie:2024:SIM

[CFA^{+}24] C. M. Condie, E. A. Fulton, K. A. Alexander, J. Vince, M. Haward, and S. A. Condie. Social influence modelling demonstrates that strategic communication and depoliticization reduces conflict in aquaculture. Marine Policy, 165 (??):??, July 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24002094.

Campbell:2021:BEB

[CFM^{+}21] Lisa M. Campbell, Luke Fairbanks, Grant Murray, Joshua S. Stoll, Linda D’Anna, and Julia Bingham. From Blue Economy to Blue Communities: reorienting aquaculture expansion for community wellbeing. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20310125.

Costa:2024:WOL

[CFS^{+}24] Ana C. Costa, Diana Freitas, Ana I. Santos, Andrea Z. Botelho, Manuela I. Parente, André Behr, Ana M. Rodrigues, Hélia Guerra, José Cascalho, and Armando Mendes. Where is ocean literacy in oceanic islands’ schools? The Azores case. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000605.

Chuah:2025:USC

[CGAF25] Sharon Xin Ying Chuah, Zhifeng Gao, Nicole L. Arnold, and Razieh Farzad. United States consumers’ attitudes towards cell-based seafood: Effective marketing strategies and policy recommendations. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24005542.

Carreton:2025:UST

[CGB^{+}25] Marta Carreton, Eve Galimany, Marta Blanco, Mariona Garriga-Panisello, Ana I. Colmenero, Marc Balcells, Ferran Bustos, Mireia G. Mingote, Cristina López-Pérez, David Nos, Xènia Puigcerver-Segarra, Marta Pujol-Baucells, John G. Ramírez, Jordi Ribera-Altimir, Alberto J. Rico, Alba Rojas, Joan Sala-Coromina, Ricardo Santos-Bethencourt, Mireia Silvestre, Antoni Lombarte, Laura Recasens, and Joan B. Company. Understanding the spatiotemporal variability of fisheries data for better bottom trawling management practices in the Catalan margin (NW Mediterranean Sea). Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24005128.

Chikudza:2020:PPI

[CGGA20] Lynnlee Chikudza, Claire Gauzente, Patrice Guillotreau, and Karen A. Alexander. Producer perceptions of the incentives and challenges of adopting ecolabels in the European finfish aquaculture industry: a Q-methodology approach. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20303146.

Christiansen:2020:EGD

[CGHV20] S. Christiansen, H. Ginzky, K. Houghton, and A. Vink. Environmental governance of deep seabed mining — Scientific insights and food for thought. Marine Policy, 114:??, April 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20300646.

Cariou:2024:EIM

[CGMH24] Pierre Cariou, Alice Thébault Guët, Jason Monios, and Ronald A. Halim. The economic impact of the Mediterranean Sulfur Emission Control Area. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002987.

Cucinelli:2023:ERG

[CGP23] Jessica Cucinelli, Floris Goerlandt, and Ronald Pelot. Exploring risk governance deficits of maritime Search and Rescue in Canada. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000386.

Campbell:2025:HMB

[CH25] Marnie L. Campbell and Chad L. Hewitt. A holistic marine biosecurity risk framework that is inclusive of social, cultural, economic and ecological values. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24005116.

Chan:2020:EIP

[Cha20a] Hing Ling Chan. Economic impacts of Papahānaumokuākea Marine National Monument expansion on the Hawaii longline fishery. Marine Policy, 115(??):??, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300090.

Chang:2020:BRM

[Cha20b] Yen-Chiang Chang. Book review: Managing the Risk of Offshore Oil and Gas Accidents — The International Legal Dimension, Gunther Handl, Kristoffer Svendsen (Eds.). Edward Elgar (2019), 482, ISBN 978-1-78643-673-3. Marine Policy, 111(??):??, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307122.

Chang:2021:UFD

[Cha21] Yen-Chiang Chang. The use of force during law enforcement in disputed maritime areas. Marine Policy, 124(??):

??, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20310332.

Chang:2022:ADI

[Cha22] Yen-Chiang Chang. Accelerating domestic and international dual circulation in the Chinese shipping industry — an assessment from the aspect of participating RCEP. Marine Policy, 135(??):??, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004735.

Chaker:2024:AGS

[Cha24a] Rachid Chaker. Arab Gulf States’ naval procurement strategies: a three-level approach. Marine Policy, 166(??):??, August 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24002355.

Chan:2024:VCO

[Cha24b] Nicholas Chan. Voluntary commitments for ocean sustainability: Comparing the UN Ocean Conference and Our Ocean Conference platforms. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2300489X.

Chang:2020:SFJ

[CHB^{+}20] Shui-Kai Chang, Nien-Tsu Alfred Hu, Samsudin Basir, Hai Vu Duyen, Praulai Nootmorn, Mudjekeewis D. Santos, and Fayakun Satria. A step forward to the joint management of the South China Sea fisheries resources: Joint works on catches, management measures and conservation issues. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305500.

Curran:2023:HMS

[CHB23] Kaitlyn Curran, Kayla M. Hamelin, and Megan Bailey. Humanizing marine spatial planning: a salutogenic approach. Marine Policy, 154(??):??, August 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001872.

Collins:2021:MIL

[CHC^{+}21] M. A. Collins, P. R. Hollyman, J. Clark, M. Soeffker, O. Yates, and R. A. Phillips. Mitigating the impact of longline fisheries on seabirds: Lessons learned from the South Georgia Patagonian toothfish fishery (CCAMLR Subarea 48.3). Marine Policy, 131(??):??, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002293.

Chen:2021:MLD

[Che21a] Yong Chen. Measuring litter distribution on UK beaches. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002037.

Chen:2021:RCD

[Che21b] Yuli Chen. Reconciling common but differentiated responsibilities principle and no more favourable treatment principle in regulating greenhouse gas emissions from international shipping. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309647.

Campos:2021:DST

[CHEC21] Aida Campos, Victor Henriques, Karim Erzini, and Margarida Castro. Deep-sea trawling off the Portuguese continental coast — Spatial patterns, target species and impact of a prospective EU-level ban. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000774.

Cannizzo:2025:FPG

[CHH^{+}25] Zachary J. Cannizzo, Karen L. Hunter, Sara Hutto, Jennifer C. Selgrath, and Lauren Wenzel. Future-proofing the global system of marine protected areas: Integrating climate change into planning and management. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004202.

Cribbs:2024:SCC

[CHN^{+}24] Tyler W. Cribbs, Jeffrey C. Hallo, William C. Norman, Matthew T. J. Brownlee, and Elizabeth A. Forys. Social carrying capacity of oyster mariculture on coast of South Carolina: Understanding stakeholders perceptions and thresholds for oyster mariculture development. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003026.

Choi:2022:AND

[Cho22] Junghwan Choi. Assessing the need for the designation of the Yellow Sea Particularly Sensitive Sea Area (PSSA). Marine Policy, 137(??):??, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000185.

Christensen:2020:SCF

[Chr20] Lotte Dalgaard Christensen. Seaweed cultivation in the Faroe Islands: Analyzing the potential for forward and fiscal linkages. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301342.

Cockerell:2021:GAS

[CJ21] Lucy M. Cockerell and Peter J. S. Jones. Governance analysis of St Anne Marine National Park, Seychelles. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306426.

Chung:2022:IMP

[CJ22] Huey-Shian Elly Chung and Juei-Cheng Jao. Improving marine protected area governance: Concerns and possible solutions from Taiwan’s practice. Marine Policy, 140(??): ??, June 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001257.

Chantavanich:2023:PSP

[CJ23] Supang Chantavanich and Waranya Jitpong. Precarity and social protection of migrant fishers in Thailand: Case studies of death and disappearance at sea. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002762.

Choy:2022:UTW

[CJCS^{+}22] Christina Pei Pei Choy, Rima W. Jabado, Naomi Clark-Shen, Danwei Huang, Min Yi Choo, and Madhu Rao. Unraveling the trade in wedgefishes and giant guitarfishes in Singapore. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100525X.

Cleaver:2024:COI

[CJH24] Caitlin Cleaver, Teresa R. Johnson, and Samuel P. Hanes. Challenges and opportunities for integrating aquaculture with commercial fishing in Maine. Marine Policy, 170(??): ??, December 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24004056.

Chang:2024:BAP

[CJK24] Yen-Chiang Chang, Muneeba Javid, and Mehran Idris Khan. BBNJ agreement in the purview of developing countries: a case study of Pakistan. Marine Policy, 165(??): ??, July 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24001994.

Chen:2023:FSM

[CJL^{+}23] Fenggui Chen, Yuhuan Jiang, Zhenghua Liu, Ruijuan Lin, and Wei Yang. Framework system of marine sustainable development assessment based on systematic review. Marine Policy, 154(??):??, August 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002166.

Choi:2024:ABO

[CJLJ24] Yongjin Choi, Duckhee Jang, Moon-Suk Lee, and Se-Jun Jin. Association between ocean literacy and climate change mitigation efforts in the Republic of Korea. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001556.

Caldas:2025:SSF

[CJN25] Juan Pablo Caldas, Laura Jaramillo, and Andrés Felipe Navia. Small-scale fisherfolk organizations: Do current conditions facilitate their effective contribution to the value chain? Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004238.

Chang:2023:MLC

[CK23] Yen-Chiang Chang and Muneeb Khan. The Maritime Labour Convention 2006 in human rights context: an appraisal. Marine Policy, 154(??):??, August 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002154.

Choi:2020:PPC

[CKY20] Kyung-Ran Choi, Ju-Hee Kim, and Seung-Hoon Yoo. Public perspective on constructing sea forests as a public good: a contingent valuation experiment in South Korea. Marine Policy, 120(??):??, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302979.

Choi:2022:PWP

[CKY22] Kyung-Ran Choi, Ju-Hee Kim, and Seung-Hoon Yoo. The public willingness to pay for the research and demonstration of tidal stream energy in South Korea. Marine Policy, 138 (??):??, April 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000288.

Cormier:2020:RAD

[CL20] Roland Cormier and Jemma Londsdale. Risk assessment for deep sea mining: an overview of risk. Marine Policy, 114:??, April 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X18306109.

Lee:2022:PSF

[cL22] Hsiao chien Lee. Promoting slow fish education in southern Taiwan coastal areas: an empirical case study of five elementary schools. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000422.

Chen:2024:DPG

[CL24] Cheng Chen and Renping Liu. Discrepancies in the portrayal of Japan’s nuclear water release in Chinese and US media outlets: a corpus-based news values analysis of marine disputes. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002100.

Clark:2020:IAN

[Cla20] Nichola A. Clark. Institutional arrangements for the new BBNJ agreement: Moving beyond global, regional, and hybrid. Marine Policy, 122(??):??, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308796.

Collins:2020:UPE

[CLB^{+}20] Claire Collins, Tom Bech Letessier, Annette Broderick, Isuru Wijesundara, and Ana Nuno. Using perceptions to examine human responses to blanket bans: the case of the thresher shark landing-ban in Sri Lanka. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20303973.

Collins:2023:VBE

[CLB^{+}23] Claire Collins, Tom B. Letessier, Aloka Benaragama, Annette Broderick, Isuru Wijesundara, Dilhara Wijetunge, and Ana Nuno. Valuable bycatch: Eliciting social importance of sharks in Sri Lanka through value chain analysis. Marine Policy, 157(??):??, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003652.

Choi:2021:WPP

[CLC21] Eun Chul Choi, Joo Suk Lee, and Jeong-In Chang. Willingness to pay for the prevention of beach erosion in Korea: the case of Haeundae beach. Marine Policy, 132(??): ??, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002785.

Custodio:2021:HNM

[CLCV21] Marco Custódio, Ana I. Lillebø, Ricardo Calado, and Sebastián Villasante. Halophytes as novel marine products — a consumers’ perspective in Portugal and policy implications. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003420.

Cox:2024:MTG

[CLD^{+}24] Kieran D. Cox, Audrey Looby, Hailey L. Davies, Kelsie A. Murchy, Brittnie Spriel, Aaron N. Rice, Francis Juanes, and Isabelle M. Côté. Military training in the Canadian Pacific: Taking aim at critical habitat or sufficient mitigation of noise pollution impacts? Marine Policy, 160(??): ??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004785.

Carter:2020:DAW

[CLL20] David W. Carter, Sabrina J. Lovell, and Christopher Liese. Does angler willingness-to-pay for changes in harvest regulations vary by state? Results from a choice experiment in the Gulf of Mexico. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20302487.

Chang:2022:LIR

[CLL22a] Yen-Chiang Chang, Xinyi Liu, and Shuang Liu. Legal issues regarding the establishment of an offshore data collection system — a practice from China. Marine Policy, 140(??): ??, June 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001245.

Chen:2022:RAS

[CLL^{+}22b] Yuli Chen, Naiyuan Lou, Guangli Liu, Yang Luan, and Han Jiang. Risk analysis of ship detention defects based on association rules. Marine Policy, 142(??):??, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001701.

Claudet:2021:CGP

[CLP21] Joachim Claudet, Charles Loiseau, and Antoine Pebayle. Critical gaps in the protection of the second largest exclusive economic zone in the world. Marine Policy, 124(??): ??, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20310307.

Ceccarelli:2022:EMP

[CLRW22] Daniela M. Ceccarelli, Asri P. Lestari, Rudyanto, and Alan T. White. Emerging marine protected areas of eastern Indonesia: Coral reef trends and priorities for management. Marine Policy, 141(??):??, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001385.

Chang:2020:CSC

[CLS20] Yen-Chiang Chang, Hefei Liu, and Yue Sun. Current Situation and Challenges in the law of armed conflict at sea: Conference report. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308814.

Campos:2023:SPF

[CLS^{+}23] Aida Campos, Pedro Leitão, Luís Sousa, Patrícia Gaspar, and Victor Henriques. Spatial patterns of fishing activity inside the Gorringe bank MPA based on VMS, AIS and elogbooks data. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004031.

Chinwendu:2024:CTC

[CLU^{+}24] Igbodiegwu Gloria Chinwendu, Markus Leibrecht, Okoli Moses Ugochukwu, Obasi Ephraim Uzoma, and Lingfeng Huang. Croakers in three coastal areas of Lagos state, Nigeria: Sustainability levels and stakeholders’ perceptions on the use of ecosystem-based fisheries management. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002951.

Chen:2021:MRA

[CLW^{+}21] Fenggui Chen, Jinwen Liu, Jianwei Wu, Nurul Dhewani Mirah Sjafrie, Puji Rahmadi, and Respaty Yudha Putranto. Measuring the relationship among stakeholders from valuesatisfaction-demand in the development of ecotourism of Marine Park. Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001305.

Chang:2023:CGM

[CLZ23] Yen-Chiang Chang, Shuang Liu, and Xu Zhang. The construction of Global Maritime Capital — current development in China. Marine Policy, 151(??):??, May 2023.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001033.

Curtin:2023:SIS

[CM23] Richard Curtin and Neil McCullough. The socioeconomic impact of the seafood sector at Ireland’s main ports. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001549.

Currie:2025:GMB

[CM25] Duncan Eric James Currie and Johannes-Alexander Müller. The governance of marine biodiversity beyond national jurisdiction: the BBNJ agreement and the International Seabed Authority. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001046.

Cisneros-Montemayor:2025:SEI

[CMBB^{+}25] A. M. Cisneros-Montemayor, H. Breakey, S. Bossier, F. Croft, I. Issifu, J. Keefer, G. G. Singh, M. Voyer, and Y. Ota. Social equity indicators for a blue economy: Guidelines for application. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25002301.

Cappell:2022:RFE

[CMC22] Rod Cappell, Graeme MacFadyen, and Andrew Constable. Research funding and economic aspects of the Antarctic krill fishery. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002470.

Cavraro:2023:SSF

[CMC^{+}23] Francesco Cavraro, Marco Anelli Monti, Alberto Caccin, Fabio Fiori, Fabio Grati, Elisabetta Russo, Giuseppe Scarcella, Dario Vrdoljak, Sanja Matić-Skoko, and Fabio Pranovi. Is the small-scale fishery more sustainable in terms of GHG emissions? A case study analysis from the Central Mediterranean Sea. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000015.

Castrejon:2024:BRC

[CMC24] Mauricio Castrejón, Nicolas Moity, and Anthony Charles. The bumpy road to conservation: Challenges and opportunities in updating the Galapagos zoning system. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001441.

Crespo:2020:BSS

[CMD^{+}20] Guillermo Ortuño Crespo, Joanna Mossop, Daniel Dunn, Kristina Gjerde, Elliott Hazen, Gabriel Reygondeau, Robin Warner, Derek Tittensor, and Patrick Halpin. Beyond static spatial management: Scientific and legal considerations for dynamic management in the high seas. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309248.

Canovas-Molina:2020:UWC

[CMGF20] Almudena Cánovas-Molina and Eduardo García-Frapolli. Untangling worldwide conflicts in marine protected areas: Five lessons from the five continents. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20304097.

Carvalho:2022:MOM

[CMI^{+}22] Laila M. Carvalho, Miguel Mies, Kelly Y. Inagaki, Eduardo G. Sanches, Marcelo R. Souza, Guilherme O. Longo, and Acácio R. G. Tomás. The marine ornamental market in Brazil (Southwestern Atlantic) frequently trades prohibited and endangered species, and threatens the ecosystem role of cleaning mutualism. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003529.

Caparros-Martinez:2025:FCC

[CMMVdPV25] Jose Luis Caparrós-Martínez, Rosa María Martínez-Vázquez, and Jaime de Pablo-Valenciano. Future challenges in the conservation and valuation of environmental services provided by the seagrass meadows of “Posidonia oceanica”. Marine Policy, 175(??):??, May 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25000454.

Cawthorn:2021:VLL

[CMNH21] Donna-Mareè Cawthorn, Torie E. Murphy, Amanda M. Naaum, and Robert H. Hanner. Vague labelling laws and outdated fish naming lists undermine seafood market transparency in Canada. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309842.

Cisneros-Montemayor:2020:CNF

[CMOB^{+}20] Andrés M. Cisneros-Montemayor, Yoshitaka Ota, Megan Bailey, Christina C. Hicks, Ahmed S. Khan, Anthony Rogers, U. Rashid Sumaila, John Virdin, and Kevin K. He. Changing the narrative on fisheries subsidies reform: Enabling transitions to achieve SDG 14.6 and beyond. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302529.

Czermanski:2024:MPA

[CMOJ^{+}24] Ernest Czermański, Magdalena Matczak, Aneta Oniszczuk-Jastrzabek, Karolina Krośnicka, and Joanna Witkowska. MSP for port areas — to what extent should we interfere with governance of ports’ waters? Case study of Polish seaports. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005183.

Cantillo:2021:ALM

[CMR21] Javier Cantillo, Juan Carlos Martín, and Concepción Román. Assessing the label’s mandatory information for fishery and aquaculture products in the EU28. A consumer approach based on a consistent fuzzy preference relation with geometric Bonferroni mean. Marine Policy, 128(??): ??, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001263.

Cadman:2020:SVC

[CMS20] Rachael Cadman, Bertrum H. MacDonald, and Suzuette S. Soomai. Sharing victories: Characteristics of collaborative strategies of environmental non-governmental organizations in Canadian marine conservation. Marine Policy, 115(??): ??, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19304038.

Cvitanovic:2021:UEB

[CMS^{+}21] C. Cvitanovic, M. Mackay, R. J. Shellock, E. I. van Putten, D. B. Karcher, and M. Dickey-Collas. Understanding and evidencing a broader range of ‘successes’ that can occur at the interface of marine science and policy. Marine Policy, 134(??):??, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004139.

Cheng:2025:OPC

[CMS^{+}25] Helen Cheng, Marissa D. McMahan, Steven B. Scyphers, Loren McClenachan, and Jonathan H. Grabowski. Observations, perceptions and concerns of the American lobster industry regarding the range-expansion of Black Sea Bass. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005177.

Cross:2025:HRM

[CMSH25] Maddison Cross, Jonathan Mitchell, and Tracey Scott-Holland. High-resolution mapping of shark movements from drone footage at Queensland beaches improves risk assessment for beach safety. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25001526.

Cisneros-Montemayor:2022:CCW

[CMSN^{+}22] Andrés M. Cisneros-Montemayor, Hussain Sinan, Tu Nguyen, José María Da Rocha, U. Rashid Sumaila, Daniel J. Skerritt, Anna Schuhbauer, Enrique Sanjurjo, and Megan Bailey. A constructive critique of the World Trade Organization draft agreement on harmful fisheries subsidies. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004838.

Capello:2023:DSB

[CMT^{+}23] Manuela Capello, Gorka Merino, Mariana Tolotti, Hilario Murua, and Laurent Dagorn. Developing a science-based framework for the management of drifting Fish Aggregating Devices. Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001847.

Chavez-Molina:2023:PSG

[CMWN^{+}23] Vasco Chavez-Molina, Daniel Wagner, Emily S. Nocito, Michelle Benedum, Carlos F. Gaymer, Duncan Currie, Emily Golden Beam, and Cassandra M. Brooks. Protecting the Salas y Gomez and Nazca Ridges: a review of policy pathways for creating conservation measures in the international waters of the Southeast Pacific. Marine Policy, 152 (??):??, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001215.

Cohn:2024:AZZ

[CMWT24] Rachel M. Cohn, Ruth Mbeyu, Andrew Wamukota, and Melva Treviño. “Akili zetu ziwe moja / Let our minds be one”: Emergent strategy for adapting to seafood insecurity in coastal Kenyan fishing communities. Marine Policy, 170(??):??, December 2024. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003403.

Chen:2024:BDN

[CMWZ24] Dongxing Chen, Guizhen Mu, Xutao Wang, and Yafeng Zhang. Brief discussion on the new approach of selection and designation of ocean dumping areas in China. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004827.

Chapsos:2023:MSG

[CN23] Ioannis Chapsos and Elizabeth Ann Norman. Is maritime security gender-blind? Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004468.

Cadrin:2024:MSC

[CN24] Steven X. Cadrin and Shuya Nakatsuka. Marine Stewardship Council’s Fisheries Standard: Updated best practices for promoting sustainable seafood. Marine Policy, 163(??): ??, May 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000812.

Coleman:2023:COL

[CNB^{+}23] Felicia C. Coleman, José A. C. C. Nunes, Áthila A. Bertoncini, Leonardo S. Bueno, Matheus O. Freitas, Maíra Borgonha, Jonas R. Leite, Márcio J. C. A. Lima-Júnior, Beatrice Ferreira, Bianca Bentes, Christopher C. Koenig, Christopher R. Malinowski, Eduardo G. Sanches, Maurício Hostim-Silva, and Cláudio L. S. Sampaio. Controversial opening of a limited fishery for Atlantic goliath grouper in the United States: Implications for population recovery. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002853.

Cummins:2023:GRO

[CNG^{+}23] Gabrielle H. Cummins, Matthew L. Navarro, Kingsley Griffin, Julian Partridge, and Tim J. Langlois. A global review of ocean ecosystem accounts and their data: Lessons learned and implications for marine policy. Marine Policy, 153(??): ??, July 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2300163X.

Canales:2024:MSE

[COJE24] Cristian M. Canales, Gonzalo Olea, Viviana Jurado, and Miguel Espíondola. Management strategies evaluation (MSE) in a mixed and multi-specific fishery based on indicator species: an example of small pelagic fish in Ecuador. Marine Policy, 162(??):??, April 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000423.

Czermanski:2022:IEE

[COJS^{+}22] Ernest Czermański, Aneta Oniszczuk-Jastrzabek, Eugen F. Spangenberg, Łukasz Kozłowski, Magdalena Adamowicz, Jakub Jankiewicz, and Giuseppe T. Cirella. Implementation of the Energy Efficiency Existing Ship Index: an important but costly step towards ocean protection. Marine Policy, 145(??):??, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003062.

Czermanski:2021:PNC

[COJZ^{+}21] Ernest Czermański, Aneta Oniszczuk-Jastrzabek, Jacek Zaucha, Barbara Pawłowska, Magdalena Matczak, and Łukasz Szydłowski. Preconditions of new container terminal location in the Maritime Spatial Planning framework. A case study for the Central Port Concept in Gdańsk. Marine Policy, 130(??):??, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001962.

Collier:2020:ECC

[Col20] Chantal E. Collier. Enabling conditions for communitybased comanagement of marine protected areas in the United States. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308903.

Cooper:2020:CDV

[COM20] J. A. G. Cooper, M. C. O’Connor, and S. McIvor. Coastal defences versus coastal ecosystems: a regional appraisal. Marine Policy, 111(??):??, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16300331.

Clifton:2021:RCD

[COSS21] Julian Clifton, Eslam O. Osman, David J. Suggett, and David J. Smith. Resolving conservation and development tensions in a small island state: a governance analysis of Curieuse Marine National Park, Seychelles. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304737.

Chen:2024:CGZ

[COYS24] Wei-Chung Chen, Ching-Hsiewn Ou, Ming-Hao Yang, and Yi-Che Shih. China’s gray zone actions in the East China Sea, Taiwan Strait, and South China Sea: a comparative study and impact on fisheries. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24002446.

Carvalho:2020:SSS

[CPBO20] Adriana Rosa Carvalho, Maria Grazia Pennino, Jose Maria Bellido, and George Olavo. Small-scale shrimp fisheries bycatch: a multi-criteria approach for data-scarse situations. Marine Policy, 116(??):??, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307164.

Calderwood:2021:TTM

[CPR21] Julia Calderwood, Debbi Pedreschi, and David G. Reid. Technical and tactical measures to reduce unwanted catches in mixed fisheries: Do the opinions of Irish fishers align with management advice? Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309362.

Calado:2021:EKB

[CPV^{+}21] H. Calado, C. Pegorelli, M. Vergílio, C. Hipólito, A. Campos, F. Moniz, A. C. Costa, C. Pereira da Silva, C. Fonseca, C. Frazão Santos, D. Gabriel, J. Guerreiro, A. J. F. Gil, D. Johnson, K. Ng, M. M. Monwar, M. A. Ventura, J. L. Suárez de Vivero, M. Pinho, P. Borges, M. Caña-Varona, and E. A. Papaioannou. Expert knowledge-based co-development of scenarios for maritime spatial planning in the Northeast Atlantic. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003523.

Chen:2020:EME

[CQ20] Xuan Chen and Weiwen Qian. Effect of marine environmental regulation on the industrial structure adjustment of manufacturing industry: an empirical analysis of China’s eleven coastal provinces. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19303744.

Comte:2025:EME

[CQAS^{+}25] Adrien Comte, Frédéric Quemmerais-Amice, Pierre Scemama, Rémi Mongruel, Clément Surun, and Harold Levrel. Experimenting marine extent and condition accounts in France. Marine Policy, 172(??):??, February 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005323.

Cao:2024:NCN

[CQW24] Xingguo Cao, Jingyi Qiu, and Mengyan Wang. A new convention for negotiable multimodal transport documents? Challenges ahead and the way forward. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003701.

Cariou:2023:SPI

[CR23] Pierre Cariou and Laingo M. Randrianarisoa. Stakeholders’ participation at the IMO Marine Environmental Protection Committee. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000337.

Craik:2020:IAM

[Cra20] Neil Craik. Implementing adaptive management in deep seabed mining: Legal and institutional challenges. Marine Policy, 114:??, April 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18303178.

Cousido-Rocha:2023:DEF

[CRCP^{+}23] Marta Cousido-Rocha, Marta González Carballo, Maria Grazia Pennino, Marta Coll, and José C. Báez. Differential effect of fisheries to the COVID-19 pandemic in the region of Andalusia (Spain). Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004894.

Casal-Ribeiro:2024:RFI

[CRIMLS24] Morgan Casal-Ribeiro, Ualerson Iran Peixoto, Wendell Medeiros-Leal, and Régis Santos. Research fatigue’s impact on small-scale fishers’ engagement: a case-study from Azores fisheries. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004020.

Coulson:2022:CPB

[CRJ22] Peter G. Coulson, Karina L. Ryan, and Gary Jackson. Are charter and private-boat recreational fishers learning to live with shark depredation? Marine Policy, 141(??): ??, July 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001439.

Cross:2021:CQM

[Cro21] Helen Cross. Challenging quota market efficiencies: a casestudy of Scotland (United Kingdom). Marine Policy, 126 (??):??, April 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20310356.

Calderwood:2020:HMC

[CRP^{+}20] Julia Calderwood, Marianne Robert, Lionel Pawlowski, Youen Vermard, Zachary Radford, Thomas L. Catchpole, and David G. Reid. Hotspot mapping in the Celtic Sea: an interactive tool using multinational data to optimise fishing practices. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307693.

Chang:2020:RLM

[CS20] Yen-Chiang Chang and Xinxiang Shi. Rule of law and the marine environmental networks: Conference report. Marine Policy, 113(??):??, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309315.

Chen:2021:BEB

[CS21] Tai-An Peter Chen and Yi-Che Shih. Blue economy based on local DNA in Taiwan: Marine regional revitalisation under the collaboration of the local and central government. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002797.

Coronado:2020:WWV

[CSCGC20] Eva Coronado, Silvia Salas, María Fernanda Cepeda-González, and Ratana Chuenpagdee. Who’s who in the value chain for the Mexican octopus fishery: Mapping the production chain. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308784.

Medeiros:2024:CAF

[CSdMR^{+}24] Macelly Correia Medeiros, Anderson Silva Pinto, Erimágna de Morais Rodrigues, Tiago Almeida de Oliveira, José Valberto de Oliveira, Sérgio de Faria Lopes, and José da Silva Mourão. Changes in artisanal fishing as a driver of the shifting baseline syndrome: from the perspective of local ecological knowledge. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002021.

Cutler:2022:TSV

[CSGC22] Matthew Cutler, Angela Silva, Lauren Gentile, and Lisa Colburn. Tracking shifts in the vulnerability and resiliency of commercial fishing vessel crew and hired captains in New England and the Mid-Atlantic. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000276.

Coelho-Souza:2025:BSC

[CSGCD25] Sergio A. Coelho-Souza, Daniel Gorman, Aurea M. Ciotti, and Gustavo M. Dias. A bright spot of conservation in a Brazilian hotspot of cumulative impacts considering ecosystem quality in a cross-shelf gradient: Perspectives to MPAs network. Marine Policy, 172(??):??, February 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004810.

Cretella:2023:TOH

[CSH23] Agnese Cretella, Cordula Scherer, and Poul Holm. Tasting the ocean: How to increase ocean literacy using seafood heritage with a visceral approach. Marine Policy, 149(??): ??, March 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000039.

Clark-Shen:2020:CSF

[CSHO^{+}20] Naomi Clark-Shen, Amanda Hsiao, Leon Oei, Madhu Rao, Ning-Yi Sun, and Alexander Douglas. A comparative study of fishery and environment laws in the south China sea: Utilizing existing laws to promote peace. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303756.

Cucuzza:2021:ETP

[CSL21] Marina Cucuzza, Joshua S. Stoll, and Heather M. Leslie. Evaluating the theoretical and practical linkages between ecosystem-based fisheries management and fisheries comanagement. Marine Policy, 126(??):??, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310411.

Castro-Santos:2020:MOS

[CSLGFV20] Laura Castro-Santos, María Isabel Lamas-Galdo, and Almudena Filgueira-Vizoso. Managing the oceans: Site selection of a floating offshore wind farm based on GIS spatial analysis. Marine Policy, 113(??):??, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303707.

Caveen:2025:GLB

[CSM^{+}25] Alex J. Caveen, Bryce D. Stewart, Cameron Moffat, Daniel Skerritt, Estelle Jones, Huw Thomas, Lara Funk, Magnus Johnson, Mike Cohen, Michael Roach, Neil A. Auchterlonie, Samuel Peacock, Sarah Hussey, Tim Gray, Tom Pickerell, Umi Muawanah, Vasiliki Kioupi, Yvonne Sadovy de Mitcheson, Neil M. Burns, and Charlotte R. Hopkins. Globalism, localism and blue food systems — how can cross-scale tensions be reduced? A production perspective from UK seafood stakeholders. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25002349.

Cabrera-Sanchez:2025:AVC

[CSMMC^{+}25] Claudia T. Cabrera-Sánchez, Gabriela Montaño-Moctezuma, Eva Coronado, Claudia E. Delgado-Ramírez, and M. Cristina Garza-Lagler. Analysis of value chain of the artisanal sea urchin fishery of Baja California, México. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24005098.

Calado:2024:UDM

[CSQ^{+}24] H. Calado, C. Frazão Santos, A. Quintela, C. Fonseca, and D. Gutierrez. The ups and downs of maritime spatial planning in Portugal. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005171.

Cozzolino:2025:CEA

[CSRG25] Maria Cozzolino, Roberta Salomone, Bruno Ricca, and Teresa Maria Gulotta. Circular economy in aquaculture: an Italian survey to understand perceptions, challenges and options for transition. Marine Policy, 176(??):??, June 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000752.

Cerca:2024:SSL

[CSV^{+}24] Mariana Cerca, Amanda Sosa, Charlene Vance, Priya Pollard, Julie Maguire, and Fionnuala Murphy. Small-scale low-tropic ocean farming and coastal rural landscapes: Why the logistics of seaweed matter? Insights from Ireland for collaborative planning. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24001386.

Caldeira:2025:WST

[CSV25] Mariana Caldeira, Alumita Talei Sekinairai, and Marjo Vierros. Weaving science and traditional knowledge: Toward sustainable solutions for ocean management. Marine Policy, 174(??):??, April 2025. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25000065.

Costa:2023:FAR

[CSVC23] H. S. Carvalho Costa, R. Scachetti, M. Vivacqua, and R. Choueri. A framework to assess resilience attributes based on local knowledge. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002993.

Caesar:2024:UPC

[CT24] Livingstone Divine Caesar and Reuben Tamakloe. Unraveling the patterns of critical contributory factors and flag-state affiliations linked with maritime incident outcomes. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001337.

Conroy:2023:SMP

[CTC^{+}23] Katherine Conroy, Daniel Tonnes, Max Calloway, Nicole Naar, Kalloway Page, and Betsy Peabody. Shoreline master programs: a hindrance or a help to kelp conservation in Puget Sound? Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003044.

Chalastani:2021:BAP

[CTCD21] Vasiliki I. Chalastani, Vasiliki K. Tsoukala, Harry Coccossis, and Carlos M. Duarte. A bibliometric assessment of progress in marine spatial planning. Marine Policy, 127 (??):??, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309763.

Cromwell:2025:DTT

[CTDY25] Jonas Cromwell, Charles Turkson, Manoj Dora, and Fred Amofa Yamoah. Digital technologies for traceability and transparency in the global fish supply chains: a systematic review and future directions. Marine Policy, 178(??):

??, August 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25001150.

Cavanagh:2021:UIA

[CTH^{+}21] Rachel D. Cavanagh, Philip N. Trathan, Simeon L. Hill, Jess Melbourne-Thomas, Michael P. Meredith, Philip Hollyman, Björn A. Krafft, Monica M. C. Muelbert, Eugene J. Murphy, Martin Sommerkorn, John Turner, and Susie M. Grant. Utilising IPCC assessments to support the ecosystem approach to fisheries management within a warming Southern Ocean. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002001.

Copeland:2023:RCF

[CTH^{+}23] Kellen Copeland, Bastian Thomsen, Michael Harte, Shaozeng Zhang, David Fennell, and David Lewis. Rethinking 21st Century Fisheries Management: posthuman multispecies livelihoods approach. Marine Policy, 152(??): ??, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001501.

Cheng:2021:SFS

[CTIH21] Yan-Tzong Cheng, Yun-Ching Tseng, Yoko Iwaki, and Michael C. Huang. Sustainable food security in Small Island Developing States (SIDS): a case of Horticulture project in Marshall Islands. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310290.

Corruble:2024:FDT

[CTR24] Philippe Corruble, Youssef Tliche, and Hamdi Radhoui. Financing decarbonization through wind-powered vessels. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001611.

Condie:2022:IPA

[CVA22a] Corrine M. Condie, Joanna Vince, and Karen A. Alexander. Increasing polarisation in attitudes to aquaculture: Evidence from sequential government inquiries. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004784.

Condie:2022:LTE

[CVA22b] Corrine M. Condie, Joanna Vince, and Karen A. Alexander. The long-term evolution of news media in defining socio-ecological conflict: a case study of expanding aquaculture. Marine Policy, 138(??):??, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000355.

Calado:2022:SSM

[CVCV^{+}22] Helena Calado, Marta Vergílio, Mario Caña-Varona, Camila Pegorelli, Cláudia Hipólito, Aida Silva, Gilberto Carreira, Maria Paramio, and Eva Papaioannou. Strategic scenarios for maritime spatial planning in an European outermost region — the case of the Azores. Marine Policy, 145(??): ??, November 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003025.

Chen:2020:EAP

[CVH^{+}20] Wenting Chen, Kristof Albert M. Van Assche, Stephen Hynes, Trine Bekkby, Hartvig C. Christie, and Hege Gundersen. Ecosystem accounting’s potential to support coastal and marine governance. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19302933.

Ciravegna:2024:HCM

[CvHF^{+}24] E. Ciravegna, L. van Hoof, C. Frier, F. Maes, H. B. Rasmussen, A. Soete, and S. W. K. van den Burg. The hidden costs of multi-use at sea. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000150.

Carter:2020:AOS

[CWB^{+}20] Amber L. Carter, A. Meriwether W. Wilson, Maximiliano Bello, E. Mauricio Hoyos-Padilla, Mark E. Inall, James T. Ketchum, Andrew Schurer, and Alexander W. Tudhope. Assessing opportunities to support coral reef climate change refugia in MPAs: a case study at the Revillagigedo Archipelago. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301654.

Cobacho:2020:ISR

[CWKE20] Sara P. Cobacho, Sonja Wanke, Zoi Konstantinou, and Ghada El Serafy. Impacts of shellfish reef management on the provision of ecosystem services resulting from climate change in the Dutch Wadden Sea. Marine Policy, 119(??): ??, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19306876.

Chang:2020:LPC

[CWKW20] Yen-Chiang Chang, Chuanliang Wang, Mehran Idris Khan, and Nannan Wang. Legal practices and challenges in addressing climate change and its impact on the oceans — a Chinese perspective. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18304925.

Church:2025:CSE

[CWLD25] Emma K. Church, Kerrie A. Wilson, Jennifer Loder, and Angela J. Dean. From citizen science experiences to stewardship action: the importance of both nature and social experiences. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005372.

Conejo-Watt:2021:FPB

[CWMM^{+}21] Heather Conejo-Watt, Angela Muench, Stephen C. Mangi, Keith Jeffery, and Kieran Hyder. Fishers perspectives on the barriers for the English inshore fleet to diversify into aquaculture. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002219.

Cremers:2020:SMC

[CWR20] Klaudija Cremers, Glen Wright, and Julien Rochette. Strengthening monitoring, control and surveillance of human activities in marine areas beyond national jurisdiction: Challenges and opportunities for an international legally binding instrument. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19309534.

Cui:2024:ELE

[CZ24] Hong-Ming Cui and Khadija Zulfiqar. Effective law enforcement in gray zone operations: a Chinese perspective of coastguard law implementation in delimited maritime zones. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003646.

Chang:2020:ILS

[CZW20] Yen-Chiang Chang, Chao Zhang, and Nannan Wang. The international legal status of the unmanned maritime vehicles. Marine Policy, 113(??):??, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304774.

Chen:2024:MFR

[CZW24] Xiaohong Chen, Junjun Zhou, and Yangjie Wang. Marine fuel restrictions and air pollution: a study on Chinese ports considering transboundary spillovers. Marine Policy, 163 (??):??, May 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24001349.

Chen:2024:IIP

[CZX24] Xidi Chen, Xudong Zhang, and Qi Xu. The implications of the Indo-Pacific Partnership for Maritime Domain Awareness (IPMDA) in the South China Sea: a Chinese perspective. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002458.

Do:2023:GFG

[DA23] Huu-Luat Do and Claire W. Armstrong. Ghost fishing gear and their effect on ecosystem services — identification and knowledge gaps. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000556.

Donkersloot:2024:RPA

[DA24] Rachel Donkersloot and Andria Agli. Rural is not a proxy: Accounting for Indigenous participation in Bristol Bay salmon fisheries. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400321X.

Danielsen:2021:TTB

[DAA21] Rannvá Danielsen, Christopher M. Anderson, and Sveinn Agnarsson. Trawling for triple bottom line results: Applying the Fishery Performance Indicators in the Faroe Islands. Marine Policy, 125(??):??, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308964.

Dutta:2021:BET

[DAAP21] Sachinandan Dutta, Ibtisam Al-Abri, and Sourav Paul. Bioeconomic trends of Hilsa (Tenualosa ilisha) fishery: Perspectives of transboundary management between India and Bangladesh. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000944.

Dauvergne:2025:CAP

[DAB^{+}25] Peter Dauvergne, Jen Iris Allan, Simon Beaudoin, Bethanie Carney Almroth, Jennifer Clapp, Emily Cowan, Babet de Groot, Trisia Farrelly, Natalia de Miranda Grilli, Alice Mah, Elizabeth Mendenhall, Rosetta Paik, Rob Ralston, Peter Stoett, Aleke Stöfen-O’Brien, Jack Taggart, Rachel Tiller, Patricia Villarrubia-Gómez, and Joanna Vince. Competing axes of power in the global plastics treaty: Understanding the politics of progress and setbacks in negotiating a high-ambition agreement. Marine Policy, 181(??): ??, November 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25002350.

Drinkwin:2023:ILF

[DAHS^{+}23] Joan Drinkwin, Kyle Antonelis, Madison Heller-Shipley, Paul Rudell, Michael Etnier, Tom Good, Anna Elz, and Jason Morgan. Impacts of lost fishing nets in the U.S. portion of the Salish Sea. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004778.

Damato:2024:EFI

[Dam24] Nikol Damato. Evaluating factors influencing compliance with vessel regulations to minimize disturbances to endangered killer whales (Orcinus orca). Marine Policy, 161(??): ??, March 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000344.

Dias:2023:VVS

[DAN^{+}23] Ana Carolina Esteves Dias, Derek Armitage, Prateep Kumar Nayak, Shehu Latunji Akintola, Evans Kwasi Arizi, Ratana Chuenpagdee, Basanta Kumar Das, Sabiha Ahmed Diba, Raktima Ghosh, Moenieba Isaacs, Gazi Md Nurul Islam, Alioune Kane, Yinji Li, Moffat Mzama Manase, Ahmadou Aly Mbaye, Paul Onyango, Sarmistha Pattanaik, Aliou Sall, Indah Susilowati, Sika Abrokwah, Denis Worlanyo Aheto, Olajumoke Akiode, Afful Chineboaba Araba, Foluke Omotayo Areola, Frederick Ato Armah, Esinam Attipoe, Aliyu Hamza Danagalan, Piyashi DebRoy, Alyne Delaney, Mariama Dieng, Kafayat A. Fakoya, Aini Nur Furoida, Cheikh Amadou Bamba Gaye, Adama Gueye, Fama Gueye, Mafaniso Hara, Haseeb Md. Irfanullah, Taofeek Iyanda, Kungwan Juntarashote, Ayorinde Kappo, Ishmael B. M. Kosamu, Hapsari Ayu Kusumawardhani, Tracey Lee, Fatou Gueye, Jenia Mukherjee, Tamano Namikawa, Sachiko Harada, Friday J. Njaya, Pranaya Kumar Parida, Betty Ngui Chiew Pieng, Made Ika Prasetyadewi, Erwin Prayogi, Daniel Quilter, Syed Tauheed Raihan, Aparna Roy, Alberta Ama Sagoe, Madalo Esther Samati, Clement Sambou, Revarunan Sammogam, Alassane Sarr, Khady Yama Sarr, Suvaluck Satumanatpan, Izumi Seki, Samiya A. Selim, Amrita Sen, Oumar Sow, I. Gusti Lanang Tantra, Suphakarn Traesupap, Vanessa Warren, Batuli Mohammed Yahya, Thamasak Yeemin, Sevil Berenji, Jessica Blythe, Alida Bundy, and Simron Singh. From vulnerability to viability: a situational analysis of small-scale fisheries in Asia and Africa. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002646.

Dauvergne:2023:NJF

[Dau23] Peter Dauvergne. The necessity of justice for a fair, legitimate, and effective treaty on plastic pollution. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003184.

Davenport:2025:PSC

[Dav25] Tara Maria Davenport. The protection of submarine cables in Southeast Asia: the security gap and challenges and opportunities for regional cooperation. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004354.

Dayan:2024:CMC

[Day24] Eleanor Dayan. Conceptualizing maritime constabulary effectiveness (MCE). Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000459.

Djosetro:2020:BLS

[DB20] Marijem Djosetro and Jelle Hendrik Behagel. Building local support for a coastal protected area: Collaborative governance in the Bigi Pan Multiple Use Management Area of Suriname. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306286.

Dolman:2022:IWC

[DBB^{+}22] Sarah J. Dolman, Charlotte N. Breen, Philippa Brakes, Andy Butterworth, and Simon J. Allen. The individual welfare concerns for small cetaceans from two bycatch mitigation techniques. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001737.

Devenport:2021:IRI

[DBBL21] Emilie Devenport, Esther Brooker, Adam Brooker, and Chris Leakey. Insights and recommendations for involving young people in decision making for the marine environment. Marine Policy, 124(??):??, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309593.

Diggon:2021:MPP

[DBH^{+}21] Steve Diggon, Caroline Butler, Aaron Heidt, John Bones, Russ Jones, and Craig Outhet. The Marine Plan Partnership: Indigenous community-based marine spatial planning. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305268.

Delgado:2022:HLM

[DBH^{+}22] James P. Delgado, Michael L. Brennan, Sergio A. Rapu Haoa, Julianna H. Rapu Leong, Carlos F. Gaymer, Diego Carabias, Emily Stokes, and Daniel Wagner. The hidden landscape: Maritime cultural heritage of the Salas y Gómez and Nazca ridges with implications for conservation on the high seas. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004887.

Do:2022:POO

[DBHT22] Thang Nam Do, Paul J. Burke, Llewelyn Hughes, and Ta Dinh Thi. Policy options for offshore wind power in Vietnam. Marine Policy, 141(??):??, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001270.

Dbouk:2022:EBP

[Dbo22] Wassim Dbouk. The Environment Bill’s policy statement on environmental principles: a golden opportunity for a principled framework for the sustainable management of UK waters. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005315.

Davis:2022:FRA

[DBS22a] Kaleigh E. Davis, Fiona Beaty, and Carolina Sánchez. A flexible risk assessment framework for marine plastic pollution that synthesizes waste management and ecological impact data. Marine Policy, 137(??):??, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004449.

Diggon:2022:MPP

[DBS^{+}22b] Steve Diggon, John Bones, Charles J. Short, Joanna L. Smith, Megan Dickinson, Kelly Wozniak, Karen Topelko, and Kylee A. Pawluk. The marine plan partnership for the North Pacific Coast — MaPP: a collaborative and co-led marine planning process in British Columbia. Marine Policy, 142(??):??, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19306657.

Dong:2024:NUW

[DBZ24] Bingying Dong, Lowell Bautista, and Ling Zhu. Navigating uncharted waters: Challenges and regulatory solutions for flag state jurisdiction of Maritime Autonomous Surface Ships under UNCLOS. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2400037X.

Daly:2020:FUF

[DC20] Jack Daly and Ratana Chuenpagdee. Facing an uncertain future: a prospective analysis of trade policy on coastal communities in northwest Newfoundland. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19306475.

Duan:2023:RBG

[DC23] Xingyi Duan and Yen-Chiang Chang. The relationship between the general principles of international law and UNCLOS: Conference report. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000799.

Donkersloot:2020:KCD

[DCC^{+}20] Rachel Donkersloot, Jesse Coleman, Courtney Carothers, Danielle Ringer, and Paula Cullenberg. Kin, community, and diverse rural economies: Rethinking resource governance for Alaska rural fisheries. Marine Policy, 117(??): ??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19300910.

Deng:2023:MPT

[DČFW23] Yuncheng Deng, Anica Čuka, Yanyan Fu, and Jiangyue Wu. Multiple paths towards Eco islands and blue development: Conference report. Marine Policy, 149(??):??, March 2023.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000532.

Dvoretsky:2022:RAR

[DD22] Alexander G. Dvoretsky and Vladimir G. Dvoretsky. Renewal of the amateur red king crab fishery in Russian waters of the Barents Sea: Potential benefits and costs. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005273.

Clippele:2023:EAS

[DDAB^{+}23] Laurence H. De Clippele, Laura Alonso Díaz, Dominic A. Andradi-Brown, Muhammad Erdi Lazuardi, Mohamad Iqbal, Imam Musthofa Zainudin, Derta Prabuning, Ruben van Hooidonk, Amehr Hakim, Firdaus Agung, Agus Dermawan, and Sebastian J. Hennige. Evaluating annual severe coral bleaching risk for marine protected areas across Indonesia. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004754.

Dowarah:2022:PSA

[DDD22] Kaushik Dowarah, Hemashree Duarah, and Suja P. Devipriya. A preliminary survey to assess the awareness, attitudes/behaviours, and opinions pertaining to plastic and microplastic pollution among students in India. Marine Policy, 144(??):??, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002676.

DeAngelis:2020:STB

[DDF^{+}20] Paola De Angelis, Lorenzo D’Andrea, Simone Franceschini, Stefano Cataudella, and Tommaso Russo. Strategies and trends of bottom trawl fisheries in the Mediterranean Sea. Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301263.

Doria:2021:UID

[DDGdS^{+}21] Carolina Rodriguesda Costa Doria, Jynessa Dutka-Gianelli, Suelen Taciane Brasil de Sousa, Jingjie Chu, and Taryn Marie Garlock. Understanding impacts of dams on the small-scale fisheries of the Madeira River through the lens of the Fisheries Performance Indicators. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309076.

DeSanto:2020:BRM

[De 20] Elizabeth De Santo. Book review: Marine protected areas: Science, policy and management, John Humphreys, Robert W. E. Clark (Eds.). Elsevier, Amsterdam (2020). Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20307181.

DeBeukelaer:2021:CBC

[De 21] Christiaan De Beukelaer. COVID-19 border closures cause humanitarian crew change crisis at sea. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002724.

DeLucia:2022:BNE

[De 22] Vito De Lucia. The BBNJ negotiations and ecosystem governance in the Arctic. Marine Policy, 142(??):??, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19306025.

Dolman:2021:INT

[DER^{+}21] S. J. Dolman, P. G. H. Evans, F. Ritter, M. P. Simmonds, and J. Swabe. Implications of new technical measures regulation for cetacean bycatch in European waters. Marine Policy, 124(??):??, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309672.

Devereux:2022:SIR

[Dev22] Helen Devereux. Seafarer injuries in relation to time into tour of duty. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004760.

Denny:2024:COM

[DF24] Shelley K. Denny and Lucia M. Fanning. Challenges and opportunities for Mi’kmaq Aboriginal and treaty rights to fish in oNova Scotia, Canada. Marine Policy, 166(??):??, August 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24002215.

Danovaro:2020:TMS

[DFC^{+}20] R. Danovaro, E. Fanelli, M. Canals, T. Ciuffardi, M.-C. Fabri, M. Taviani, M. Argyrou, E. Azzurro, S. Bianchelli, A. Cantafaro, L. Carugati, C. Corinaldesi, W. P. de Haan, A. Dell’Anno, J. Evans, F. Foglini, B. Galil, M. Gianni, M. Goren, S. Greco, J. Grimalt, Q. Güell-Bujons, A. Jadaud, L. Knittweis, J. L. Lopez, A. Sanchez-Vidal, P. J. Schembri, P. Snelgrove, S. Vaz, L. Angeletti, M. Barsanti, J. A. Borg, M. Bosso, A. Brind’Amour, G. Castellan, F. Conte, I. Delbono, F. Galgani, G. Morgana, S. Prato, A. Schirone, and E. Soldevila. Towards a marine strategy for the deep Mediterranean Sea: Analysis of current ecological status. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19305524.

Drakeford:2023:BSF

[DFF23] Benjamin M. Drakeford, Andy Forse, and Pierre Failler. Biodegradability and sustainable fisheries: the case for static gear in the UK Channel fishery. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2300307X.

Defeo:2025:FOF

[DFP^{+}25] Omar Defeo, Bárbara C. Franco, Alberto R. Piola, Claudio C. Buratti, Luis Gustavo Cardoso, Federico Cortés, Ignacio Gianelli, Micaela Giorgini, Gabriela Jorge-Romero, Diego Lercari, Erika Meerhoff, Leonardo Ortega, Jose Angel Alvarez Perez, Nicolás Prandoni, and Rodrigo Sant’Ana. Facing oceanographic, fisheries, and governance hotspots: Scientific evidence and policy implications from the southwest South Atlantic Ocean. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24005840.

Drakeford:2020:IFT

[DFTK20] Benjamin M. Drakeford, Pierre Failler, Ben Toorabally, and Elyes Kooli. Implementing the Fisheries Transparency Initiative: Experience from the Seychelles. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309285.

Crespo:2021:VCM

[dFVCGdS21] Maria de Fátima Vieira Crespo, Jaíra Maria Alcobaça Gomes, and Regina Oliveira da Silva. Value chain of the mangrove crab (Ucides cordatus): a case study of the Parnaíba Delta, Northeast Brazil. Marine Policy, 131(??): ??, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002530.

Du:2020:ESE

[DG20] Yuan-Wei Du and Kun Gao. Ecological security evaluation of marine ranching with AHP-entropy-based TOPSIS: a case study of Yantai, China. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20308691.

Dewan:2023:RCS

[DG23] Mohammud Hanif Dewan and Radu Godina. Roles and challenges of seafarers for implementation of energy efficiency operational measures onboard ships. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002798.

Dewan:2024:OSE

[DG24a] Mohammud Hanif Dewan and Radu Godina. An overview of seafarers’ engagement and training on energy efficient operation of ships. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005134.

Dewan:2024:USA

[DG24b] Mohammud Hanif Dewan and Radu Godina. Unveiling seafarers’ awareness and knowledge on energy-efficient and low-carbon shipping: a decade of IMO regulation enforcement. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000356.

Dagtekin:2022:CAI

[DGG22] Murat Dağtekin, Ali Cemal Gücü, and Yaşar Genç. Concerns about illegal, unreported and unregulated fishing, carbon footprint, and the impact of fuel subsidy — an economic analysis of the Black Sea anchovy fishery. Marine Policy, 140(??):??, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001142.

Damonte:2025:ESG

[DGH25] Gerardo Damonte, Isabel Gonzales, and Susana Higueras. The elusive sustainable growth: the formation of institutional socio-ecological traps in Peruvian anchoveta fisheries. Marine Policy, 176(??):??, June 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000727.

Das:2024:SPB

[DGI^{+}24] Jewel Das, Maheshwaran Govender, Haseeb Md. Irfanullah, Samiya Ahmed Selim, and Marion Glaser. Stakeholder perceptions of blue economy governance networks and their equity implications in Bangladesh. Marine Policy, 170(??): ??, December 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003579.

Diez:2022:SAB

[DGIS^{+}22] A. De la Torriente Diez, J. M. González-Irusta, A. Serrano, R. Aguilar, F. Sánchez, M. Blanco, and A. Punzón. Spatial assessment of benthic habitats vulnerability to bottom fishing in a Mediterranean seamount. Marine Policy, 135(??): ??, January 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004619.

Dupont:2020:EUM

[DGMF^{+}20] Clément Dupont, Françoise Gourmelon, Catherine Meur-Ferec, Frédérick Herpers, and Christophe Le Visage. Exploring uses of maritime surveillance data for marine spatial planning: a review of scientific literature. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19308930.

Dixon:2024:IRT

[DGS^{+}24] Maximilian Dixon, Gaetano Grilli, Bryce D. Stewart, Rosalind H. Bark, and Silvia Ferrini. The importance of rebuilding trust in fisheries governance in post-Brexit England. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000320.

Davis:2022:TFC

[DH22] Ruth A. Davis and Quentin Hanich. Transparency in fisheries conservation and management measures. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20301172.

Deely:2022:OVD

[DHC22] John Deely, Stephen Hynes, and Mary Cawley. Overseas visitor demand for marine and coastal tourism. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002238.

Garcia-del-Hoyo:2021:GCB

[dHJTGO21] Juan José García del Hoyo, Ramón Jiménez-Toribio, and Félix García-Ordaz. Granger causality between the canning sector and the Spanish tuna fleet: Evidence from the Toda–Yamamoto approach. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003122.

Dudayev:2023:PFG

[DHR23] Rayhan Dudayev, Lugas Lukmanul Hakim, and Indah Rufiati. Participatory fisheries governance in Indonesia: Are octopus fisheries leading the way? Marine Policy, 147(??): ??, January 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003852.

Dolman:2024:RSC

[DHU24] Sarah J. Dolman, Nicola K. Hodgins, and Fernando Ugarte. A review of small cetacean hunts in Greenland. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0308597X24003993. See corrigendum [DHU25].

Dolman:2025:CRS

[DHU25] Sarah J. Dolman, Nicola K. Hodgins, and Fernando Ugarte. Corrigendum to: “A review of small cetacean hunts in Greenland” [Mar. Policy 170 (2024) 106401]. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004147. See [DHU24].

DeValck:2023:VES

[DJC^{+}23] Jeremy De Valck, Diane Jarvis, Anthea Coggan, Ella Schirru, Petina Pert, Victoria Graham, and Maxine Newlands. Valuing ecosystem services in complex coastal settings: an extended ecosystem accounting framework for improved decision-making. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002944.

deJesus:2024:TKS

[dJdMCL^{+}24] Paulo Protasio de Jesus, Ana Melissa de Moraes Câmara, Moisés Meireles Leal, Regiane da Silva Almeida, Caio Brito Lourenço, and Izabel Cristina da Silva Almeida Funo. Traditional knowledge and socioeconomic aspects of smallscale bivalve fishing on the Amazon coast: a case study of Iphigenia brasiliensis on the Island of Maranhão. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000745.

deJuan:2024:THS

[dJRFVOA24] Silvia de Juan, Ana Ruiz-Frau, Sebastián Villasante, and Andres Ospina-Alvarez. A threatened heritage: Sustainable alternatives for Mediterranean small-scale fisheries. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001465.

Daly:2021:CCB

[DKKS21] Jack Daly, Christine Knott, Poppy Keogh, and Gerald G. Singh. Changing climates in a blue economy: Assessing the climate-responsiveness of Canadian fisheries and oceans policy. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002347.

deKlerk:2021:SPA

[dKMK21] Yvette de Klerk, Michael Ekow Manuel, and Momoko Kitada. Scenario planning for an autonomous future: a comparative analysis of national preparedness of selected countries. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000385.

Dahood:2020:PSL

[DKW20] Adrian Dahood, Emily S. Klein, and George M. Watters. Planning for success: Leveraging two ecosystem models to support development of an Antarctic marine protected area. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309145.

Dahlet:2021:FCM

[DL21] Lol Iana Dahlet and José Luis Sánchez Lizaso. Fisheries co-management, past and present: From the Plan Castellón (1961–1966) for Spanish Mediterranean trawling fisheries to the current EU fisheries policy. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000919.

Du:2022:CFI

[DL22] Yuan-Wei Du and Xiao-Xue Li. Critical factor identification of marine ranching ecological security with hierarchical DEMATEL. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200029X.

Pineiro-Antelo:2020:FSG

[dlÁPAFGO20] María de los Ángeles Piñeiro-Antelo, Jesús Felicidades-García, and Brendan O’Keeffe. The FLAG scheme in the governance of EU coastal areas. The cases of Ireland and Galicia (Spain). Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18302045.

delaConcepcion:2024:NFE

[dlC24] Richard Tur de la Concepción. Negotiating fair and equitable sharing of benefits in the BBNJ agreement: Role of the Group of 77 and China. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000836.

Garcia-de-la-Fuente:2020:CCC

[dlFGFFR^{+}20] Laura García de-la Fuente, Lucía García-Flórez, Ma. Pino Fernández-Rueda, Jorge Alcázar-Álvarez, Arturo Colina-Vuelta, Esteban Fernández-Vázquez, and Carmen Ramos-Carvajal. Comparing the contribution of commercial and recreational marine fishing to regional economies in Europe. An input-output approach applied to Asturias (Northwest Spain). Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306980.

Dacks:2020:DBV

[DLJ^{+}20] Rachel Dacks, Staci A. Lewis, Philip A. S. James, Lincy L. Marino, and Kirsten L. L. Oleson. Documenting baseline value chains of Palau’s nearshore and offshore fisheries prior to implementing a large-scale marine protected area. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301782.

Dong:2025:CPF

[DLL25] Yichong Dong, Wei Liu, and Jingmei Li. Chinese policy on fisheries subsidies: Evolution, challenges, and new prospects. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002416.

delaPuente:2024:LCR

[dlPC24] Santiago de la Puente and Villy Christensen. Linking catch reconstructions with downstream supply chain nodes can help strengthen management actions in favour of just, sustainable and resilient futures. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003853.

delaPuente:2022:ASL

[dlPdlLLC^{+}22] Santiago de la Puente, Rocío López de la Lama, Camila Llerena-Cayo, Benny R. Martínez, Gonzalo Rey-Cama, Villy Christensen, María Rivera-Ch, and Armando Valdés-Velasquez. Adoption of sustainable low-impact fishing practices is not enough to secure sustainable livelihoods and social wellbeing in small-scale fishing communities. Marine Policy, 146(??):??, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003682.

Ding:2021:DFI

[DLX21] Lili Ding, Mengtong Lu, and Yuemei Xue. Driving factors on implementation of seasonal marine fishing moratorium system in China using evolutionary game. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003183.

Du:2024:CST

[DLX24] Rui Du, Baiqiong Liu, and Min Xu. Current status and trends in island ecosystem services research: a bibliometric analysis from 2000 to 2023. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000964.

Ding:2022:UCC

[DLYM22] Lili Ding, Mengxi Liu, Ying Yang, and Wen Ma. Understanding Chinese consumers’ purchase intention towards traceable seafood using an extended Theory of Planned Behavior model. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000203.

DaCostaOliveira:2020:TKF

[DMd^{+}20] Pablo Da Costa Oliveira, Ana Paula Madeira Di Beneditto, Valéria da Silva Quaresma, Alex Cardoso Bastos, and Camilah Antunes Zappes. Traditional knowledge of fishers versus an environmental disaster from mining waste in Central Brazil. Marine Policy, 120(??):??, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301585.

Dutto:2023:AEE

[DME23] Davide Dutto, Krister Mars, and Håkan Eggert. Assessing the economic effects of the COVID-19 pandemic on Swedish shrimp fishers. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000234.

Drouineau:2023:FTF

[DMG^{+}23] Hilaire Drouineau, Fabien Moullec, Didier Gascuel, Francis Laloë, Sterenn Lucas, Nicolas Bez, Patrice Guillotreau, Jérôme Guitton, Pierre-Yves Hernvann, Martin Huret, Sigrid Lehuta, Marc Léopold, Stéphanie Mahévas, Marianne Robert, Mathieu Woillez, and Youen Vermard. Food for thought from French scientists for a revised EU Common Fisheries Policy to protect marine ecosystems and enhance fisheries performance. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22005073.

Dallison:2025:EMA

[DMHS25] Thomas Dallison, Sarah Manuel, Benjamin Haskell, and Francis Staub. Effectively managing area-based management tools: a new toolkit for the conservation of marine mammals. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004925.

Dario:2024:PPC

[DMK24] Carlie Dario, Renato Molina, and David L. Kelly. Public preferences for coastal adaptation: Economic evidence from a discrete choice experiment for hard structures and naturebased solutions in Miami, Florida. Marine Policy, 165(??): ??, July 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2400215X.

deMitcheson:2020:VVF

[dMLB^{+}20] Yvonne J. Sadovy de Mitcheson, Christi Linardich, João Pedro Barreiros, Gina M. Ralph, Alfonso Aguilar-Perera, Pedro Afonso, Brad E. Erisman, David A. Pollard, Sean T. Fennessy, Athila A. Bertoncini, Rekha J. Nair, Kevin L. Rhodes, Patrice Francour, Thierry Brulé, Melita A. Samoilys, Beatrice P. Ferreira, and Matthew T. Craig. Valuable but vulnerable: Over-fishing and undermanagement continue to threaten groupers so what now? Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19306116.

DeSanto:2020:SMY

[DMNT20] Elizabeth M. De Santo, Elizabeth Mendenhall, Elizabeth Nyman, and Rachel Tiller. Stuck in the middle with you (and not much time left): the third intergovernmental conference on biodiversity beyond national jurisdiction. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308334.

Divu:2024:EPM

[DMS^{+}24] Damodaran Nair Divu, Suresh Kumar Mojjada, Pratheesh Omana Sudhakaran, Swathi Lekshmi Perumal Sundaram, Muktha Menon, Ramesh Kumar Mojjada, Mayur Shivdas Tade, Vinuja Syamala Vishwambharan, Jai Shree, Aarsha Subramanian, Boby Ignatius, and Achamveetil Gopalakrishnan. Economic performance and marine policy implications of mud spiny lobster mariculture in Tropical Sea Cages, North-Eastern Arabian Sea, India: an empirical study in marine economics. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000393.

Dudouet:2025:AIF

[DMT25] Benjamin Dudouet, Claire Macher, and Olivier Thébaud. Allocating individual fishing possibilities through producer organisations: the case of the Bay of Biscay common sole fishery in France. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25001794.

deMelo:2025:ERP

[dMVJdAdSM25] Paulo Wanderley de Melo, Arleu Barbosa Viana-Junior, Maria Elisabeth de Araújo, and José da Silva Mourão. Environmental risk perception and adaptative strategies in a neotropical fishing population: Socioeconomic aspects and community participation. Marine Policy, 175(??):??, May 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25000387.

Devaux:2020:DES

[DN20] Caroline Devaux and Jean-Philippe Nicolaï. Designing an EU Ship Recycling Licence: a Roadmap. Marine Policy, 117 (??):??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19301757.

Diekert:2021:CAC

[DNR21] Florian Diekert, Linda Nøstbakken, and Andries Richter. Control activities and compliance behavior — Survey evidence from Norway. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310320.

deOliveira:2023:OJH

[dO23] César Soares de Oliveira. One jurisdiction away from a healthier ecosystem? The impacts of jurisdictional zones on the health of large marine ecosystems. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002257.

Darias-OHara:2025:BRT

[DOBE^{+}25] A. K. Darias-O’Hara, C. G. Booth, C. Erbe, S. Isojunno, V. M. Janik, K. Lucke, B. Southall, J. Tougaard, A. M. von Benda-Beckmann, and U. K. Verfuss. Behavioural response thresholds for the assessment of noise impact on Antarctic marine mammal species. Marine Policy, 179(??): ??, September 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25001538.

Doerr:2021:RSE

[Doe21] Angee N. Doerr. Regulating for social and environmental sustainability: a stakeholder perspective from the Bahamian spiny lobster fishery. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20310174.

Junior:2021:LSO

[dOJCSLdSB21] José Gilmar Cavalcante de Oliveira Júnior, João V. Campos-Silva, Richard J. Ladle, and Vandick da Silva Batista. Linking social organization, attitudes, and stakeholder empowerment in MPA governance. Marine Policy, 130(??): ??, August 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001548.

Dominioni:2023:TET

[Dom23] Goran Dominioni. Towards an equitable transition in the decarbonization of international maritime transport: Exemptions or carbon revenues? Marine Policy, 154(??):??, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001963.

deOliveira:2020:AEI

[dOPM^{+}20] Larissa Rosa de Oliveira, Ana Carolina Pont, Rodrigo Machado, Mônica Tais Engel, Paulo Henrique Ott, Enrique A. Crespo, and Silvio Marchini. Assessing the economic impact caused by South American sea lions based on onboard check versus fishermen’s perception: the two sides of the same coin. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19306517.

Soares:2020:OSS

[dOSTB^{+}20] Marcelo de Oliveira Soares, Carlos Eduardo Peres Teixeira, Luís Ernesto Arruda Bezerra, Sandra Vieira Paiva, Tallita Cruz Lopes Tavares, Tatiane Martins Garcia, Jorge Thé de Araújo, Carolina Coelho Campos, Sarah Maria Cavalcante Ferreira, Helena Matthews-Cascon, Alice Frota, Tarin Cristino Frota Mont’Alverne, Solange Teles Silva, Emanuelle Fontenele Rabelo, Cristiane Xerez Barroso, João Eduardo Pereira de Freitas, Mauro de Melo Júnior, Renata Polyana de Santana Campelo, Claudeilton Severino de Santana, Pedro Bastos de Macedo Carneiro, Antônio Jeovah Meirelles, Bráulio Almeida Santos, André Henrique Barbosa de Oliveira, Paulo Horta, and Rivelino Martins Cavalcante. Oil spill in South Atlantic (Brazil): Environmental and governmental disaster. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19308346.

Drakopulos:2023:FCI

[DP23] Lauren Drakopulos and Melissa Poe. Facing change: Individual and institutional adaptation pathways in West Coast fishing communities. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004109.

DeDomenico:2023:RML

[DPB^{+}23] Francesca De Domenico, Cristina Pedà, Claudio Berti, Pierpaolo Consoli, Francesco Longo, Silvestro Greco, and Teresa Romeo. Reducing marine litter pollution from anchored fish aggregating devices: Sicilian good practices and the legislative paradox. Marine Policy, 154(??):??, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002221.

Delord:2022:FEM

[DPG^{+}22] Karine Delord, Timothée Poupart, Nicolas Gasco, Henri Weimerskirch, and Christophe Barbraud. First evidence of migration across the South Pacific in endangered Amsterdam albatross and conservation implications. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005327.

delPozo:2021:GAT

[dPJ21] Daniela Laínez del Pozo and Peter J. S. Jones. Governance analysis of two historical MPAs in northern Peru: Isla Lobos de Tierra and Isla Lobos de Afuera. Marine Policy, 127 (??):??, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19309492.

Daniell:2020:EEG

[DPP^{+}20] Katherine A. Daniell, Roel Plant, Victoria Pilbeam, Catherine Sabinot, Nicolas Paget, Karen Astles, Ruben Steffens, Olivier Barreteau, Severine Bouard, Peter Coad, Anna Gordon, Nils Ferrand, Pierre-Yves Le Meur, Caroline Lejars, Pierre Maurel, Ana Rubio, Jean-Emmanuel Rougier, and Ian White. Evolutions in estuary governance? Reflections and lessons from Australia, France and New Caledonia. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18303336.

Donald:2025:DIP

[DPSF25] Daniel Donald, Rachael Priest, Christopher Sweeting, and Clare Fitzsimmons. Disturbance impacts of planned offshore wind expansion on UK sediment blue carbon stocks. Marine Policy, 181(??):??, November 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002635.

Davies:2023:NFE

[DQJ23] Tim K. Davies, Elise Quinn, and Ernesto Jardim. A novel framework to evaluate the accuracy of information used in MSC fisheries assessments: Development challenges and solutions. Marine Policy, 158(??):??, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004025.

Valck:2022:RUP

[DR22] Jeremy De Valck and John Rolfe. Reviewing the use of proxies to value coastal and marine biodiversity protection: the Great Barrier Reef in Australia. Marine Policy, 136(??): ??, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21005017.

DeValck:2025:UNU

[DR25] Jeremy De Valck and John Rolfe. Use or non-use? A deep dive into the amenity and ecological benefits of the great barrier reef. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002192.

delCarmenPena-Puch:2025:CGG

[dRA25] Angelina del Carmen Peña-Puch and Evelia Rivera-Arriaga. Challenges for good governance in coastal socio-ecological systems of Campeche: Analysis from the expert users’ perspective. Marine Policy, 176(??):??, June 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000491.

Delgado:2021:AFE

[DRCdMCJ21] Erick Rodolfo Menéndez Delgado, Elvira Bernardita Rodríguez Ríos, Darío Bernal Casasola, and María del Mar Cerbán Jiménez. Artisanal fishery in Ecuador. A case study of Manta city and its economic policies to improve competitiveness of the sector. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2030960X.

Durano:2025:RIL

[DRS^{+}25] Francis Levi Abdala Durano, Glory Dee Antero Romo, Jon Marx Paredes Sarmiento, Rodgessa Apas Lopez, Jonna Belle Draug Hollite, and Princess Anne Badilla Padios. A review of the impacts of laws and policies on the nexus of the fisheries sector and gender in the Philippines. Marine Policy, 180(??):??, October 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002258.

Deng:2022:IER

[DRY22] Yuncheng Deng, James Randall, and Fang Ye. Island ecological restoration and management practices based on nature: Conference report. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002354.

daSilva:2021:BAA

[dS21] Alexandre Pereira da Silva. Brazil advances over the Area: the inclusion of the Rio Grande Rise within the Brazilian outer continental shelf and its consequences for other states and for the common heritage of mankind. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000099.

daSilva:2022:PVT

[dS22a] Alexandre Pereira da Silva. Protecting the Vitoria– Trindade Chain: Legal clarifications and alternatives toward a marine protected area in the Brazilian outer continental shelf. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005376.

Dunkley:2022:WFB

[DS22b] Frith Dunkley and Jean-Luc Solandt. Windfarms, fishing and benthic recovery: Overlaps, risks and opportunities. Marine Policy, 145(??):??, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003098.

DelBene:2021:PDG

[DSB21] James A. DelBene, Andrew M. Scheld, and Donna M. Bilkovic. Preferences for derelict gear mitigation strategies by commercial fishers. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002736.

Daliri:2025:TET

[DSB^{+}25] Moslem Daliri, Ali Salarpouri, Siamak Behzadi, Saud M. Al Jufaili, Saeid Gorgin, Elyas Soleimani, Danial Ghaderi, and Mohammad Momeni. Toward an effective temporary fishing ban law enforcement in the Persian Gulf. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25002283.

Driedger:2023:GMP

[DSC^{+}23] Alex Driedger, Jennifer Sletten, Claire Colegrove, Timothé Vincent, Virgil Zetterlind, Joachim Claudet, and Barbara Horta e Costa. Guidance on marine protected area protection level assignments when faced with unknown regulatory information. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004882.

DiCintio:2024:IAF

[DSD^{+}24] Antonio Di Cintio, Erik Sulanke, Selene Di Genio, Federico Niccolini, Valerio Sbragaglia, Francesca Visintin, and Fabio Bulleri. Investigating artisanal fishers’ support for MPAs: Evidence from the Tuscan Archipelago (Mediterranean Sea). Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002586.

deLima:2022:ECC

[dSdLSGB22] Andre de Souza de Lima, Marinez Eymael Garcia Scherer, Tiago Borges Ribeiro Gandra, and Jarbas Bonetti. Exploring the contribution of climate change policies to integrated coastal zone management in Brazil. Marine Policy, 143(??): ??, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002275.

daSilveira:2024:TCS

[dSF24] Mariana Ferreira da Silveira and Beatrice Padovani Ferreira. Temporal changes in a small-scale artisanal reef fishery in Brazil: Coastal development and its impacts. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001842.

Ding:2023:ECT

[DSJ^{+}23] Qi Ding, Xiujuan Shan, Xianshi Jin, Harry Gorfine, Lisha Guan, and Tao Yang. Evolution of China’s Total Allowable Catch (TAC) system: Review and way forward. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004377.

Junior:2020:BEB

[dSJNS20] Oswaldo Gomes de Souza Junior, José Leocyvan Gomes Nunes, and Renato Azevedo Matias Silvano. Biology, ecology and behavior of the acoupa weakfish Cynoscion acoupa (Lacepède, 1801) according to the local knowledge of fishermen in the northern coast of Brazil. Marine Policy, 115 (??):??, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19302593.

Dinkel:2020:ISE

[DSL20] T. M. Dinkel and J. L. Sánchez-Lizaso. Involving stakeholders in the evaluation of management strategies for shortfin mako (Isurus oxyrinchus) and blue shark (Prionace glauca) in the Spanish longline fisheries operating in the Atlantic Ocean. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301883.

Deith:2021:LLC

[DSL^{+}21] Mairin C. M. Deith, Daniel J. Skerritt, Roberto Licandeo, Daniel E. Duplisea, Caroline Senay, Divya A. Varkey, and Murdoch K. McAllister. Lessons learned for collaborative approaches to management when faced with diverse stakeholder groups in a rebuilding fishery. Marine Policy, 130(??):??, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001664.

daSilva:2021:PGI

[dSOW^{+}21] Carina Vieira da Silva, Margarida Ortigão, Tom Willaert, Renato Rosa, Luís Catela Nunes, and Maria A. Cunha e Sá. Participatory Geographic Information Systems (PGIS): Alternative approaches to identify potential conflicts and positional accuracy in marine and coastal ecosystem services. Marine Policy, 131(??):??, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100261X.

Delgado:2022:ITK

[DSQ^{+}22] Luisa E. Delgado, Camila Sandoval, Paola Quintanilla, David Quiñones-Guerrero, Ignacio A. Marín, and Víctor H. Marín. Including traditional knowledge in coastal policymaking: Yaldad Bay (Chiloé, southern Chile) as a case study. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002287.

Rodrigues:2024:EDF

[dSRKTC24] Lucas dos S. Rodrigues, Paul Gerhard Kinas, Vinni S. Thykjaer, and Luis Gustavo Cardoso. Enjoy the darkness: Forsake partially nocturnal sets provides a good opportunity to improve profits and sustainability in the southern Brazil longline fishery. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24001477.

Teixeira:2024:SPD

[dSTHL^{+}24] Gibran da Silva Teixeira, Vinícius Halmenschlager, Pedro Henrique Soares Leivas, Patrízia Raggi Abdallah, Ussif Rashid Sumaila, and Larissa Carvalho Solino Silva. The subsidy program for diesel oil prices for fishing vessels and its impacts on the labor market and fish production in Brazil. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001003.

Dunga:2025:PCS

[DSV^{+}25] Loyiso V. Dunga, Kerry J. Sink, Megan G. Van der Bank, Linda R. Harris, Lara Van Niekerk, Myriam Perschke, Ruan Brand, Judy Mann-Lang, Mike Bruton, Mia Strand, Nina Rivers, Stephen Lamberth, Harriet Clift, Bernadette Snow, Edward Knott, and Kira Gee. Piloting a culturally significant areas framework for spatial planning and management in the coastal environment of South Africa. Marine Policy, 180(??):??, October 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002222.

deSantiago:2025:PCM

[dSVT^{+}25] Jose Alberto de Santiago, Sebastián Villasante, Ana Tubío, Manel Antelo, Pablo Pita, and Jose María Da Rocha. Productivity changes and management systems in small-scale fisheries. Marine Policy, 174(??):??, April 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000144.

Doddema:2020:RIT

[DSWB20] Mandy Doddema, Gert Spaargaren, Budy Wiryawan, and Simon R. Bush. Responses of Indonesian tuna processing companies to enhanced public and private traceability. Marine Policy, 119(??):??, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302207.

Daniel:2023:DFG

[DT23] Damaris Benny Daniel and Saly N. Thomas. Derelict fishing gear abundance, its causes and debris management practices — insights from the fishing sector of Kerala, India. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004766.

Domondon:2021:PEM

[DTBP21] Paolo Roberto Domondon, Raquel Sanchez Tirona, Steve Box, and Robert Pomeroy. Pathways to establishing managed access and networks of reserves. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001913.

Duan:2022:CPD

[Dua22] Wen Duan. China’s participation in the discussion on marine protected areas in the BBNJ negotiations and its implications. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200313X.

Dunning:2021:AGP

[Dun21a] Kelly Heber Dunning. Adaptive Governance: the proposed port expansion in the Cayman Islands and its impacts to coral reefs. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308940.

Dunning:2021:HMR

[Dun21b] Kelly Heber Dunning. How are managers responding to local and global ecological stressors? The case of Indonesian co-managed coral reefs in the Anthropocene. Marine Policy, 131(??):??, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001718.

Dunley:2023:UNV

[Dun23] Richard Dunley. Uncrewed naval vessels and the span of maritime tasks. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000088.

Dunley:2024:AMS

[Dun24] Richard Dunley. Australianised military off-the-shelf: Australia’s naval ship design ethos and the Anzac Class frigates. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000010.

Demirel:2021:MTA

[DUYEA21] Nazli Demirel, Aylin Ulman, Taner Yıldız, and Pinar Ertör-Akyazi. A moving target: Achieving good environmental status and social justice in the case of an alien species, Rapa whelk in the Black Sea. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002980.

deVrees:2021:AMS

[dV21] Leo de Vrees. Adaptive marine spatial planning in the Netherlands sector of the North Sea. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X18304597.

Deville:2025:KGF

[DVRK25] Alejandro Deville, Ian Vázquez-Rowe, and Ramzy Kahhat. Knowledge gaps and future research priorities linked to microplastic abundance and occurrence in Peruvian fisheries and seafood products. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2400544X.

Dalton:2020:RBP

[DWC^{+}20] Tracey Dalton, Michael Weir, Aislyne Calianos, Nelle D’Aversa, and Julia Livermore. Recreational boaters’ preferences for boating trips associated with offshore wind farms in US waters. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308629.

Devi:2021:EPA

[DXS^{+}21] Manoharmayum Shaya Devi, K. A. Martin Xavier, Asem Sanjit Singh, Leela Edwin, Veerendra Veer Singh, and Latha Shenoy. Environmental pressure of active fishing method: a study on carbon emission by trawlers from northwest Indian coast. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000646.

Dyck:2024:TIA

[Dyc24] Carol Dyck. On thin ice: the Arctic Council’s uncertain future. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000587.

Dyrset:2025:BGN

[Dyr25] Guri Dyrset. Blue growth in Norway — three emerging narratives. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001332.

Zoysa:2020:BWD

[dZ20] Rapti Siriwardane de Zoysa. Beyond the wall: Dyking as an object of everyday governance in the Bay of Manila, Philippines. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303567.

Duan:2024:MLS

[DZZC24] Xingyi Duan, Xiaonan Zhao, Marlene Zou, and Yen-Chiang Chang. Maritime laws and sustainable development of Blue Economy: Conference report. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003373.

Estes:2021:EML

[EAA^{+}21] Maurice Estes, Clarissa Anderson, Ward Appeltans, Nic Bax, Nina Bednaršek, Gabrielle Canonico, Samy Djavidnia, Elva Escobar, Peer Fietzek, Marilaure Gregoire, Elliott Hazen, Maria Kavanaugh, Franck Lejzerowicz, Fabien Lombard, Patricia Miloslavich, Klas O. Möller, Jacquomo Monk, Enrique Montes, Hassan Moustahfid, Monica M. C. Muelbert, Frank Muller-Karger, Lindsey E. Peavey Reeves, Erin V. Satterthwaite, Jörn O. Schmidt, Ana M. M. Sequeira, Woody Turner, and Lauren V. Weatherdon. Enhanced monitoring of life in the sea is a critical component of conservation management and sustainable economic growth. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003109.

Estradivari:2022:MCB

[EAA^{+}22] Estradivari, Muh. Firdaus Agung, Dedi Supriadi Adhuri, Sebastian C. A. Ferse, Ita Sualia, Dominic A. Andradi-Brown, Stuart J. Campbell, Mohamad Iqbal, Harry D. Jonas, Muhammad Erdi Lazuardi, Hellen Nanlohy, Fitryanti Pakiding, Ni Kadek Sri Pusparini, Hikmah C. Ramadhana, Toni Ruchimat, I Wayan Veda Santiadji, Natelda R. Timisela, Laura Veverka, and Gabby N. Ahmadia. Marine conservation beyond MPAs: Towards the recognition of other effective area-based conservation measures (OECMs) in Indonesia. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005509.

Eggert:2021:AGF

[EAAG21] Håkan Eggert, Christopher M. Anderson, James L. Anderson, and Taryn M. Garlock. Assessing global fisheries using Fisheries Performance Indicators: Introduction to special section. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030899X.

Estradivari:2022:MCS

[EABA^{+}22] Estradivari, Dominic A. Andradi-Brown, Amkieltiela, Christian N. Handayani, Fikri F. Sjahruddin, Muh. Firdaus Agung, Stuart J. Campbell, Kelly Claborn, Matheus De Nardo, Helen E. Fox, Louise Glew, Amehr Hakim, Muhammad Erdi Lazuardi, Hellen Nanlohy, Wira Sanjaya, Edy Setyawan, Natelda R. Timisela, Laura Veverka, Nara Wisesa Wiwardhana, Marthen Welly, Imam Musthofa Zainudin, and Gabby N. Ahmadia. Marine conservation in the Sunda Banda Seascape, Indonesia. Marine Policy, 138(??): ??, April 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000410.

Elvarsson:2020:UMC

[EAGV20] Bjarki Tór Elvarsson, Sveinn Agnarsson, Saga Gudmundsdóttir, and Jónas Vidarsson. Using multi-criteria analysis the assess impacts of change in ecosystem-based fisheries management: the case of the Icelandic cod. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18308303.

Elvestad:2023:EUT

[EB23] Christel Elvestad and Trond Bjørndal. The EU–UK Trade and Cooperation Agreement: Issue linkages and its implications for fisheries management and trade in seafood. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004274.

Ellefsen:2024:GFI

[EB24] Hans Ellefsen and Daniel W. Bromley. Faroe Islands fisheries in the Northeast Atlantic: Brexit, Russia, and economic incoherence. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004281.

Exeter:2025:FSM

[EBB^{+}25] Owen M. Exeter, Julian Branscombe, Annette C. Broderick, Tom Hooper, Jan Maclennan, Trudy Russell, Kate Sugar, Alice Trevail, Julie Webber, and Kristian Metcalfe. Fine-scale mapping of ocean user groups to support species and habitat spatial management. Marine Policy, 174(??): ??, April 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2500003X.

Eger:2021:SRI

[EdLP^{+}21] S. L. Eger, R. C. de Loë, J. Pittman, G. Epstein, and S. C. Courtenay. A systematic review of integrated coastal and marine management progress reveals core governance characteristics for successful implementation. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002992.

Erauskin-Extramiana:2024:LFE

[EECE^{+}24] Maite Erauskin-Extramiana, Adrien Chevallier, Bruno Ernande, Yunne-Jai Shin, and Gorka Merino. Lower fishing effort would benefit fish stocks and fisheries profitability in the Gulf of Lion in the Mediterranean Sea. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001210.

ElOumrassi:2025:SCS

[EEIER25] Hajar El Oumrassi, Sana El Arraf, Mohammed Malouli Idrissi, and Hasan Er-Raioui. Socioeconomic consequences of Sardina pilchardus (Walbaum, 1792) decline on the purse seiners activity and workers along the Southern Alboran Sea. Marine Policy, 172(??):??, February 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005025.

Evans:2025:CGP

[EFCU25] Ally J. Evans, Ellie Ford, and Leanne Cullen-Unsworth. Connecting governance and practice views on the barriers and solutions to scaled seagrass restoration in the UK. Marine Policy, 180(??):??, October 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001964.

Evans:2023:UTT

[EFFP23] Tegan Evans, Stephen Fletcher, Pierre Failler, and Jonathan Potts. Untangling theories of transformation: Reflections for ocean governance. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002439.

Egede:2023:UAC

[Ege23] Edwin Egede. UNCLOS 82: Africa’s contributions to the development of modern law of the sea 40 years later. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005103.

Ehler:2021:ISI

[Ehl21a] Charles Ehler. Introduction to a special issue on marine/ maritime spatial planning. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20306424.

Ehler:2021:TDP

[Ehl21b] Charles N. Ehler. Two decades of progress in Marine Spatial Planning. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303816.

Engehagen:2021:OHF

[EHLT21] Sondre Engehagen, Marius Hornslien, Maria Lavrutich, and Stian Tønnessen. Optimal harvesting of farmed salmon during harmful algal blooms. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001391.

Ewell:2020:ERF

[EHM^{+}20] Christopher Ewell, John Hocevar, Elizabeth Mitchell, Samantha Snowden, and Jennifer Jacquet. An evaluation of Regional Fisheries Management Organization at-sea compliance monitoring and observer programs. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0308597X19305913. See corrigendum [EHM^{+}22].

Ewell:2022:CER

[EHM^{+}22] Christopher Ewell, John Hocevar, Elizabeth Mitchell, Samantha Snowden, and Jennifer Jacquet. Corrigendum to “An evaluation of Regional Fisheries Management Organization at-sea compliance monitoring and observer programs” [Mar. Policy 115 (2020) 103842]. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000707. See [EHM^{+}20].

Edvardsson:2025:SBU

[EHÓ25] Ragnar Edvardsson, Einar Hjörleifsson, and Gudbjörg Ásta Ólafsdóttir. Synergies between underwater cultural heritage and biodiversity conservation in Iceland: Insights for achieving 30 × 30 targets. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25001976.

Eggers:2024:ESI

[EJC24] Jessica M. Eggers, Susan J. Johnston, and Kevern L. Cochrane. The ecological and socioeconomic implications of future total allowable catch decisions and illegal fishing scenarios for South Africa’s West Coast rock lobster fishery. Marine Policy, 170(??):??, December 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400397X.

Escalle:2023:TNE

[EMH^{+}23] Lauriane Escalle, Jennyfer Mourot, Paul Hamer, Steven R. Hare, Naiten Bradley Phillip, and Graham M. Pilling. Towards non-entangling and biodegradable drifting fish aggregating devices — baselines and transition in the world’s largest tuna purse seine fishery. Marine Policy, 149(??): ??, March 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000271.

Eladawy:2024:TCH

[EMN^{+}24] Ahmed Eladawy, Neil C. Mitchell, Takashi Nakamura, Momen El-Husseiny, Yuta A. Takagi, Nabil Elhady, Brook Muller, Sara Abdel-Hamid, Asmaa Mohammed, Kazuo Nadaoka, and J. P. Walsh. Towards a combined human– natural system approach in the Northern Red Sea Region: Ecological challenges, sustainable development, and community engagement. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004505.

Espinasse:2023:SPN

[EMS^{+}23] Marina Espinasse, Eirik Mikkelsen, Sigrunn Holbek Sørbye, Mette Skern-Mauritzen, Jannike Falk-Andersson, and Per Fauchald. Seafood production in Northern Norway: Analyzing variation and co-development in aquaculture and coastal fisheries. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300310X.

Edwards:2020:POQ

[EP20] D. N. Edwards and E. Pinkerton. Priced out of ownership: Quota leasing impacts on the financial performance of owner-operators. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304105.

EROL:2022:FEI

[Ero22] Sercan Erol. Financial and economic impacts of the COVID-19 pandemic on aquaculture in Türkiye and financial policy recommendations. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003608.

Eryasar:2022:USM

[ES22] Ahmet Raif Eryaşar and İsmet Saygu. Using social media to identify recreational bluefish angling in the Mediterranean and Black Sea. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004450.

Ewart:2024:MCT

[ESD24] Marisha Ewart, Pascal Scherrer, and Kay Dimmock. Managing commercial tourism for conservation and sustainable use: Policy instrument interactions in Cape Byron Marine Park, Australia. Marine Policy, 166(??):??, August 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002318.

Ellett:2025:PBW

[ETB25] Lindsey G. Ellett, Jennifer M. Thomsen, and Charles Besançon. Potential barriers and windows of opportunity for transboundary marine management in the Sulu–Sulawesi seascape: a policy analysis. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24004731.

Elsira:2025:DTE

[ETSKA25] Ibrahim Elsira, Alexander Tilley, Mohamed Samy-Kamal, and Walid Aly. Digital transformation of Egyptian marine fisheries: a pathway to sustainable fisheries management. Marine Policy, 179(??):??, September 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001812.

Engel:2022:BPA

[EV22] Monica T. Engel and Jerry J. Vaske. Balancing public acceptability and consensus regarding marine protected areas management using the Potential for Conflict Index_{2}. Marine Policy, 139(??):??, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000896.

Engel:2021:SHN

[EVB21] Monica T. Engel, Jerry J. Vaske, and Alistair J. Bath. Seal hunting in Newfoundland from the perspective of local people. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001020.

Erkkila-Valimaki:2022:CFS

[EVPMKP22] Anne Erkkilä-Välimäki, Mari Pohja-Mykrä, Jenny Katila, and Riitta Pöntynen. Coastal fishery stakeholders’ perceptions, motivation, and trust regarding maritime spatial planning and regional development: the case in the Bothnian Sea of the northern Baltic Sea. Marine Policy, 144(??): ??, October 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002524.

Fall:2024:AOF

[FA24] Assane Dedah Fall and Berchie Asiedu. The artisanal octopus fishery in Mauritania: a lucrative fishery in decline. Marine Policy, 169(??):??, November 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003336.

Fujita:2023:SBC

[FAB^{+}23] Rod Fujita, Simona Augyte, Jennifer Bender, Poppy Brittingham, Alejandro H. Buschmann, Max Chalfin, Jamie Collins, Kristen A. Davis, John Barry Gallagher, Rebecca Gentry, Rebecca L. Gruby, Kristin Kleisner, Monica Moritsch, Nichole Price, Loretta Roberson, John Taylor, and Charles Yarish. Seaweed blue carbon: Ready?

Or not? Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002804.

Fisher:2023:ECC

[FAC^{+}23] Jeffrey Fisher, Dror Angel, Myriam Callier, Daniel Cheney, Ramon Filgueira, Bobbi Hudson, Christopher W. McKindsey, Lisa Milke, Heather Moore, Francis O’Beirn, Jack O’Carroll, Berit Rabe, Trevor Telfer, and Carrie J. Byron. Ecological carrying capacity in mariculture: Consideration and application in geographic strategies and policy. Marine Policy, 150(??):??, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300043X.

Finseraas:2024:GWW

[FAC^{+}24] Eirik Finserås, Ignacio Herrera Anchustegui, Etienne Cheynet, Cristian Guillermo Gebhardt, and Joachim Reuder. Gone with the wind? Wind farm-induced wakes and regulatory gaps. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2300430X.

Furman:2021:MMA

[FAGS21] Kelsi L. Furman, Payam Aminpour, Steven A. Gray, and Steven B. Scyphers. Mental models for assessing coastal social-ecological systems following disasters. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309830.

Fajardo:2022:CCI

[Faj22] Teresa Fajardo. To criminalise or not to criminalise IUU fishing: the EU’s choice. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002597.

Fassina:2025:CCA

[FAMW^{+}25] C. M. Fassina, L. R. Almeida, E. S. Morris-Webb, L. Y. Xavier, G. D. Blanco, L. M. Diele-Viegas, and L. R. Gonçalves. Climate change adaptation: Raising fisherwomen’s voices to policy making. Marine Policy, 181(??): ??, November 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2500243X.

Fan:2024:LMJ

[Fan24] Zhiying Fan. The logic of the Marcos Jr. government’s South China Sea policy. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2400246X.

Fernandez-Arcaya:2024:BFB

[FARBV^{+}24] Ulla Fernandez-Arcaya, Augusto Rodríguez-Basalo, Patricia Verísimo, José Rodriguez, Elvira Ceballos, Jose Manuel Gonzalez-Irusta, Ana García-Alegre, Maider Plaza-Morlote, Alberto Serrano, and Antonio Punzón. Bottom fishing beyond trawling. spatio-temporal trends of mobile and static bottom fisheries on benthic habitats. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2300338X.

Fujita:2023:TER

[FBC^{+}23] Rod Fujita, Poppy Brittingham, Ling Cao, Halley Froehlich, Matt Thompson, and Taylor Voorhees. Toward an environmentally responsible offshore aquaculture industry in the United States: Ecological risks, remedies, and knowledge gaps. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003980.

Finnsson:2025:ISM

[FCG25] Hjörleifur Finnsson, Catherine Chambers, and Gudni Gudbergsson. Invasive species management: the case of pink salmon in Iceland. Marine Policy, 173(??):??, March 2025.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005396.

Forse:2021:BBU

[FDF^{+}21] Andy Forse, Benjamin Drakeford, Pierre Failler, Jonathan Potts, and Negar Akbari. Beyond Brexit — Is the UK’s Fixed Quota Allocation (FQA) system in need of a fix? Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001743.

Fabbri:2024:GIS

[FDGM24] Federico Fabbri, Jean-Noël Druon, Brendan J. Godley, and David March. Geospatial identification of stakeholders to support dynamic ocean management in transboundary areas. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000332.

Filho:2024:HGP

[FDVP24] Walter Leal Filho, Maria Alzira Pimenta Dinis, Claudio R. P. Vasconcelos, and Arminda Paço. Handling the growing problem of offshore food waste. Marine Policy, 163(??): ??, May 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24001350.

Feichtner:2020:CLE

[Fei20] Isabel Feichtner. Contractor liability for environmental damage resulting from deep seabed mining activities in the area. Marine Policy, 114:??, April 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307115.

Feldman:2022:RAF

[Fel22] Nizan Feldman. Rallying around the “flag of convenience”: Merchant fleet flags and sanctions effectiveness. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001762.

Fennell:2023:CBC

[Fen23] Hannah L. Fennell. Changing behavior: Can intervention design from the public health sector help solve the problem of fishing gear conflict? Marine Policy, 151(??):??, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000544.

Frohlich:2023:NAA

[FFD^{+}23] Miguel Frohlich, Pedro Fidelman, Ian Dutton, Marcus Haward, Brian W. Head, Dianne Maynard, David Rissik, and Joanna Vince. A network approach to analyse Australia’s blue economy policy and legislative arrangements. Marine Policy, 151(??):??, May 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300115X.

Fedi:2020:MAM

[FFE20] Laurent Fedi, Olivier Faury, and Laurent Etienne. Mapping and analysis of maritime accidents in the Russian Arctic through the lens of the Polar Code and POLARIS system. Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030230X.

Fedi:2024:AIT

[FFE^{+}24] Laurent Fedi, Olivier Faury, Laurent Etienne, Ali Cheaitou, and Patrick Rigot-Muller. Application of the IMO taxonomy on casualty investigation: Analysis of 20 years of marine accidents along the North-East Passage. Marine Policy, 162(??):??, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000599.

Ferreira:2022:RUR

[FFGJ22] Maria Adelaide Ferreira, Christopher Barrio Froján, Vikki Gunn, and David E. Johnson. A role for UNEP’s Regional Seas Programme under the post-2020 global biodiversity framework. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005418.

Finkbeiner:2021:CPW

[FFYE21] Elena M. Finkbeiner, Juno Fitzpatrick, and Whitney Yadao-Evans. A call for protection of women’s rights and economic, social, cultural (ESC) rights in seafood value chains. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000932.

Froehlich:2025:UUS

[FG25] Halley E. Froehlich and Jessica A. Gephart. Uncertain United States seafood sustainability in a manufactured crisis. Marine Policy, 180(??):??, October 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002106.

Friess:2021:POR

[FGC21] Bernhard Friess and Marie Grémaud-Colombier. Policy outlook: Recent evolutions of maritime spatial planning in the European Union. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304652.

Ferreira:2024:ECU

[FGC^{+}24] João-Pedro Ferreira, Taryn Garlock, Christa D. Court, James L. Anderson, and Frank Asche. The economic contribution of U.S. seafood imports throughout the value chain: a sectorial and species-specific analysis. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003737.

Finkbeiner:2024:ICP

[FGF^{+}24] Elena M. Finkbeiner, Christopher Giordano, Juno Fitzpatrick, Ashley Apel, Cecilia Blasco, Kathryn H. Dalton, Juan Carlos Jerí, Ines Lopez-Ercilla, Gabrielle E. Lout, Chris Madden, Ivan Martinez-Tovar, Pablo Obregon, Jada Tullos Anderson, and John N. Kittinger. Insights from a community of practice: Integrating human rights in fisheries improvement. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000988.

Finke:2020:NWM

[FGK^{+}20] Gunnar Finke, Kira Gee, Anja Kreiner, Maria Amunyela, and Rodney Braby. Namibia’s way to Marine Spatial Planning — Using existing practices or instigating its own approach? Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030333X.

Froehlich:2021:SSF

[FGL^{+}21] Halley E. Froehlich, Rebecca R. Gentry, Sarah E. Lester, Richard S. Cottrell, Gavin Fay, Trevor A. Branch, Jessica A. Gephart, Easton R. White, and Julia K. Baum. Securing a sustainable future for US seafood in the wake of a global crisis. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309751.

Fernandez-Gonzalez:2022:CSC

[FGPPGG22] Raquel Fernández-González, Marcos I. Pérez-Pérez, and María Dolores Garza-Gil. COVID-19 and the Spanish Celtic Sea fishery: an economic analysis. Marine Policy, 143(??): ??, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002512.

Fernandez-Gonzalez:2023:AEE

[FGPPGG23] Raquel Fernández-González, Marcos I. Pérez-Pérez, and María Dolores Garza-Gil. Assessing the economic effects of the European landing obligation on the Spanish fishery in the Celtic Sea. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005000.

Fernandez-Gonzalez:2021:ROS

[FGPPPV21] Raquel Fernández-González, Marcos I. Pérez-Pérez, and Raisa Pérez-Vas. Real options for a small company in a context of market concentration: a case study of investment in a turbot farming plant in Spain. Marine Policy, 134(??): ??, December 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004395.

Fernandez-Gonzalez:2024:RAG

[FGPVPP^{+}24] Raquel Fernández-González, Raisa Pérez-Vas, Marcos Pérez-Pérez, Zita Graça Teixeira Pereira, Félix Puime-Guillén, and Rosa María Ricoy-Casas. Resilience and adaptation: Galician canning fish industry evolution. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2400085X.

Furman:2023:SES

[FHBS23] Kelsi L. Furman, Sharon L. Harlan, Luiz Barbieri, and Steven B. Scyphers. Social equity in shore-based fisheries: Identifying and understanding barriers to access. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200402X.

Frenkel:2025:PFF

[FIB25] Caitlin M. Frenkel, Josephine Iacarella, and Natalie C. Ban. Predicting factors of fishing gear loss and distribution across Canada’s Pacific Ocean. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25002775.

Fischer:2022:HTR

[Fis22] Johanne Fischer. How transparent are RFMOs? Achievements and challenges. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19307158.

Febrica:2025:SSF

[FKNG25] Senia Febrica, Herson Keradjaan, Laely Nurhidayah, and Michael Gustaf. Small-scale fishers’ perception of risks in Indonesia’s cross-border region of North Maluku. Marine Policy, 175(??):??, May 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25000399.

Fu:2024:MEG

[FL24] Benchao Fu and Heran Li. Marine environmental governance for nuclear pollution: From the perspective of China’s response to Japan’s Fukushima nuclear wastewater discharge. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002409.

Flaaten:2021:RDF

[Fla21] Ola Flaaten. The rise and decline of fishing industry support — with a translation from Norwegian of Bjørn S. Brochmann’s 1981 article ‘Long-term effects of government support to the fisheries’. Marine Policy, 126(??):??, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20304541.

Filous:2021:FSM

[FLB^{+}21] Alexander Filous, Robert J. Lennox, Jean Pierre Beaury, Hinano Bagnis, Mathew Mchugh, Alan M. Friedlander, Eric E. G. Clua, Steven J. Cooke, Todd K. Fuller, and Andy J. Danylchuk. Fisheries science and marine education catalyze the renaissance of traditional management (rahui) to improve an artisanal fishery in French Polynesia. Marine Policy, 123(??):??, January 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309374.

Fan:2023:SPC

[FLC^{+}23] Hanwen Fan, Jing Lyu, Zheng Chang, Xuzhuo He, and Shu Guo. Spatial patterns and characteristics of global piracy analyzed using a geographic information system. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003494.

Fullonton:2025:GAA

[FLLML25] Abhilasha Fullonton, Amanda R. Lea-Langton, Fatima Madugu, and Alice Larkin. Green ammonia adoption in shipping: Opportunities and challenges across the fuel supply chain. Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004445.

Fraser:2020:CSV

[FMCD20] Molly D. Fraser, Lauren H. McWhinnie, Rosaline R. Canessa, and Chris T. Darimont. Compliance of small vessels to minimum distance regulations for humpback and killer whales in the Salish Sea. Marine Policy, 121(??): ??, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20306709.

Franco-Melendez:2021:TUR

[FMCT^{+}21] Milagros Franco-Meléndez, Luis A. Cubillos, Jorge Tam, Simón Hernández Aguado, Renato A. Quiñones, and Aldo Hernández. Territorial Use Rights for Fisheries (TURF) in central-southern Chile: Their sustainability status from a transdisciplinary holistic approach. Marine Policy, 132(??): ??, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002554.

Fernandez-Macho:2020:AAV

[FMGV20] Javier Fernández-Macho, Pilar González, and Jorge Virto. Assessing anthropogenic vulnerability of coastal regions: DEA-based index and rankings for the European Atlantic Area. Marine Policy, 119(??):??, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302694.

Francolini:2023:SPS

[FMLMM23] E. M. Francolini, J. B. Mann-Lang, E. McKinley, and B. Q. Mann. Stakeholder perspectives on socio-economic challenges and recommendations for better management of the Aliwal Shoal Marine Protected Area in South Africa. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005218.

Fujii:2022:IEC

[FMS^{+}22] Iwao Fujii, Miko Maekawa, Nozomi Shimizu, Naohisa Kanda, Nariaki Mikuni, Kazunobu Suzuki, Izumi Tsurita, Miriam C. Balgos, and Marjo K. Vierros. Implications of existing capacity building efforts for the conservation and sustainable use of marine biological diversity of areas beyond national jurisdiction: a case study of Japan. Marine Policy, 138(??):??, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000513.

Franco-Melendez:2025:CCM

[FMTA^{+}25] Milagros Franco-Meléndez, Jorge Tam, Simón Hernández Aguado, Luis A. Cubillos, and Beatriz Cid-Aguayo. Characterising the co-management system in the Biobío Region, Chile, based on fishers’ and experts’ knowledge: a dual perspective. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001642.

Fuentes:2021:CIR

[FMWW21] Mariana M. P. B. Fuentes, Zoë A. Meletis, Natalie E. Wildermann, and Matthew Ware. Conservation interventions to reduce vessel strikes on sea turtles: a case study in Florida. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000828.

Fleming:2020:STH

[FOH^{+}20] Aysha Fleming, Emily Ogier, Alistair J. Hobday, Linda Thomas, Jason R. Hartog, and Bianca Haas. Stakeholder trust and holistic fishery sustainability assessments. Marine Policy, 111(??):??, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306335.

Fujii:2023:PCR

[FOKM23] Iwao Fujii, Yumi Okochi, Hajime Kawamura, and Mitsutaku Makino. Potential cooperation of RFMOs for the integrity of MCS: Lessons from the three RFMOs in the Asia–Pacific. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002816.

Formenti:2022:ATF

[For22] Lorenzo Formenti. Assessing transparency in fisheries subsidies: a notification-driven analysis. Marine Policy, 136(??): ??, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19307766.

Fox:2022:SSU

[Fox22] Clive J. Fox. Scoring the status of UK shared and national fish stocks around the time of Brexit. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004620.

Florido:2023:ECP

[FR23] Alberto Roca Florido and Emilio Padilla Rosa. Everything comes at a price: the impact of eliminating harmful subsidies in the Spanish marine resources industry. Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001884.

Fernandez:2024:SDB

[FR24] Pelayo Rico Fernández and Antonella Rivera. SCUBA diver behavior, diver carrying capacity and its implications for sustainable reef management. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24001970.

Franckx:2023:SMB

[Fra23] Erik Franckx. A single maritime boundary: From UNCLOS III to present-day developments. Marine Policy, 148(??): ??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004729.

Fontenele:2024:SSF

[FRCR24] Raimundo Eduardo Silveira Fontenele, Francisco Wellington Ribeiro, Maria Regiane Araújo Cavalcante, and Dimitri-Alexander Lenine Bivar Ribeiro. Small-scale fisheries as a common good and a social and solidarity economy practice. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003038.

Froehlicher:2024:PPP

[FRDK24] Hermione Froehlicher, Tina Rambonilaza, Françoise Daverat, and Kenzo Kaifu. Public preferences for policies promoting the conservation of a universally threatened species (Anguilla spp.): Insights from a choice experiment in Japan. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003233.

Friedman:2024:IBA

[Fri24] Shani Friedman. The interaction of the BBNJ agreement and the legal regime of the Area, and its influence on the implementation of the BBNJ agreement. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002331.

Flaherty:2025:SDL

[FRLMW25] Mark Flaherty, Gregor Reid, Leah Lewis-McCrea, and Therese Wilson. Seafood dependent livelihoods and climate change: Insights from the lobster fishery in Nova Scotia. Marine Policy, 178(??):??, August 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001149.

Finnis:2022:MWF

[FRM22] Joel Finnis and Emily Reid-Musson. Managing weather & fishing safety: Marine meteorology and fishing decisionmaking from a governance and safety perspective. Marine Policy, 142(??):??, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001671.

Frost:2021:NBE

[Fro21] Rebecca Frost. Normative beliefs and economic life: a case study of the fishing industry in two communities in rural Newfoundland. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000336.

Filho:2022:ESV

[FRVB22] Luiz Magalhães Filho, Peter Roebeling, Sebastián Villasante, and Maria Isabel Bastos. Ecosystem services values and changes across the Atlantic coastal zone: Considerations and implications. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003128.

Ford:2021:SBT

[FS21] Eleanor Ford and Bryce D. Stewart. Searching for a bridge over troubled waters: an exploratory analysis of trust in United Kingdom fisheries management. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002979.

Freeman:2022:WES

[FS22] Richard Freeman and Kristina Svels. Women’s empowerment in small-scale fisheries: the impact of Fisheries Local Action Groups. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005182.

Frohlich:2021:LBA

[FSF^{+}21] Miguel F. Frohlich, Timothy F. Smith, Pedro Fidelman, Claudia Baldwin, Chris Jacobson, and R. W. (Bill) Carter. Legal barriers to adaptive coastal management at a coastal erosion hotspot in Florianópolis, Brazil. Marine Policy, 127 (??):??, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000464.

Faunce:2023:FOE

[FSKJ23] Craig H. Faunce, Jaclyn Smith, Andy Kingham, and Dennis Jaszka. Fisheries observers as enforcement assets: 21 years of lessons from the North Pacific. Marine Policy, 158(??): ??, December 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004013.

Foliente:2025:CIC

[FSRC25a] Joville Andrea P. Foliente, Asa Jose U. Sajise, Paul Joseph B. Ramirez, and Maria Angeles O. Catelo. Corrigendum to “Impacts of a community-based TURF-Reserve policy on CPUE and fishers’ income: Evidence from a spatial fuzzy regression discontinuity design” [Mar. Policy 180 (2025) 106812]. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002428. See [FSRC25b].

Foliente:2025:ICB

[FSRC25b] Joville Andrea P. Foliente, Asa Jose U. Sajise, Paul Joseph B. Ramirez, and Maria Angeles O. Catelo. Impacts of a community-based TURF-Reserve policy on CPUE and fishers’ income: Evidence from a spatial fuzzy regression discontinuity design. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25002271. See corrigen-

dum [FSRC25a].

Flye:2025:LPM

[FSZ25] Melissa E. Flye, Carly C. Sponarski, and Joseph Zydlewski. Local perceptions of marine conservation aquaculture for the restoration of native Atlantic salmon (Salmo salar) in Downeast, Maine. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400424X.

Fjeld:2023:MNG

[FTG^{+}23] Kristian Fjeld, Rachel Tiller, Eduardo Grimaldo, Leif Grimsmo, and Inger-Beate Standal. Mesopelagics — new gold rush or castle in the sky? Marine Policy, 147(??): ??, January 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004067.

Failler:2020:PTR

[FTGD^{+}20] Pierre Failler, Grégoire Touron-Gardic, Benjamin Drakeford, Oumar Sadio, and Marie-Suzanne Traoré. Perception of threats and related management measures: the case of 32 marine protected areas in West Africa. Marine Policy, 117 (??):??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19304452.

Ferguson:2022:LPP

[FTM^{+}22] Caroline E. Ferguson, Teri Tuxson, Sangeeta Mangubhai, Stacy Jupiter, Hugh Govan, Victor Bonito, Semese Alefaio, Maxine Anjiga, Jonathan Booth, Tracey Boslogo, Delvene Boso, Ambroise Brenier, Akanisi Caginitoba, Ana Ciriyawa, Joeli Bili Fahai’ono, Margaret Fox, Andy George, Hampus Eriksson, Alec Hughes, Eugene Joseph, Sean Kadannged, Eferemo Kubunavanua, Sesimani Loni, Semisi Meo, Fiorenza Micheli, Elizah Nagombi, Rebecca Omaro, Anouk Ride, Annisah Sapul, Ann Singeo, Karen Stone, Margaret Tabunakawai-Vakalalabure, Marama Tuivuna, Caroline Vieux, Vutaieli B. Vitukawalu, and McKenzie Waide. Local practices and production confer resilience to rural Pacific food systems during the COVID-19 pandemic. Marine Policy, 137(??):??, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200001X.

Froese:2021:PTE

[FTSG21] Rainer Froese, Athanassios C. Tsikliras, Giuseppe Scarcella, and Didier Gascuel. Progress towards ending overfishing in the Northeast Atlantic. Marine Policy, 125(??): ??, March 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309283.

Flynn:2023:MSD

[FTW^{+}23] Sarah Flynn, Elizabeth Tray, Tom Woolley, Adam Leadbetter, Kellie Heney, Deirdre O’Driscoll, Caitriona Nic Aonghusa, and Andrew Conway. Management of spatial data integrity including stakeholder feedback in Maritime Spatial Planning. Marine Policy, 156(??):??, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003329.

Fujita:2021:AMD

[Fuj21] Rod Fujita. The assessment and management of data limited fisheries: Future directions. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003419.

Fermepin:2024:GMC

[FWGM24] Solange Fermepin, James E. M. Watson, Hedley S. Grantham, and Martin Mendez. Global marine conservation priorities for sustaining marine productivity, preserving biodiversity and addressing climate change. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000149.

Fu:2022:EIA

[FWL^{+}22] Xiu-Mei Fu, Li-Xia Wang, Chun-Yu Lin, Wan-Yu Wu, Hong-Li Ku, Shan-Shan Jiang, and Ying Liu. Evaluation of the innovation ability of China’s marine fisheries from the perspective of static and dynamic. Marine Policy, 139 (??):??, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000793.

Field:2025:ASA

[FWSS^{+}25] Laurel C. Field, Kendyl Wright, Jenna Sullivan-Stack, Jean M. Harris, Stephen P. Kirkman, Peter J. Fielding, Ané Oosthuizen, Siyabonga Dlulisa, Anastasiya Laznya, Emanuel J. Gonçalves, and Kirsten Grorud-Colvert. Assessing South Africa’s marine protected area quality and progress towards conservation goals: an application of The MPA Guide framework. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2400513X.

Fan:2022:CFT

[FXC^{+}22] Guohao Fan, Xiaoke Xie, Jihong Chen, Zheng Wan, Mingzhu Yu, and Jia Shi. Has China’s Free Trade Zone policy expedited port production and development? Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005625.

Faure:2024:LST

[FY24] Michael G. Faure and Haiyang Yu. The liability of salvors towards third parties in case of environmental salvage: Economic analysis. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004475.

Fang:2025:ICC

[FZ25] Xiaohan Fang and Ying Zhang. The impact of climate change and economic development on the catches of small pelagic fisheries. Marine Policy, 175(??):??, May 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000466.

Feng:2025:ENQ

[FZL25] Yuan Feng, Shiyang Zhang, and Chaomin Liu. Exploring new quality productive forces in ocean governance: Insights from China’s practices and strategic directions. Marine Policy, 176(??):??, June 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000715.

Feng:2023:CMM

[FZYF23] Fan Feng, Chenqi Zhai, Zhenguo Yang, and Kuo Feng. China’s marine manufacturing export growth: Policy, status, and driving components. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000222.

Gairin:2020:RHD

[GA20] Emma Gairin and Serge Andréfouët. Role of habitat definition on Aichi Target 11: Examples from New Caledonian coral reefs. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306505.

Gyan:2020:APF

[GAA^{+}20] Watson Ray Gyan, Elliot Haruna Alhassan, Amos Asase, Daniel Nsoh Akongyuure, and Yang Qi-Hui. Assessment of postharvest fish losses: the case study of Albert Bosomtwi-Sam fishing harbour, Western Region, Ghana. Marine Policy, 120(??):??, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301536.

Griffiths:2022:LHF

[GABA^{+}22] Laura L. Griffiths, Dominic A. Andradi-Brown, Gabby N. Ahmadia, Purwanto, Awaludinnoer, Dale Bryan-Brown, and Christopher J. Brown. Linking historical fishing pressure to biodiversity outcomes to predict spatial variation in Marine Protected Area performance. Marine Policy, 139 (??):??, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000719.

Grimm:2023:FFJ

[GABA^{+}23] Kerry E. Grimm, Jessica L. Archibald, Sara E. Bonilla-Anariba, Nadia Bood, and Steven W. J. Canty. Framework for fostering just and equitable seagrass policy, management, and social-ecological outcomes: Lessons learned from Belizean marine resource managers. Marine Policy, 152 (??):??, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001331.

Goti-Aralucea:2021:EVR

[GABSD21] Leyre Goti-Aralucea, Jörg Berkenhagen, Erik Sulanke, and Ralf Döring. Efficiency vs resilience: the rise and fall of the German brown shrimp fishery in times of COVID 19. Marine Policy, 133(??):??, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002864.

Giovos:2020:AMS

[GAD^{+}20] Ioannis Giovos, Marco Arculeo, Nikolaos Doumpas, Dimitra Katsada, Mary Maximiadi, Eleni Mitsou, Vangelis Paravas, Roxani Naasan Aga-Spyridopoulou, Vasilis-Orestis Stoilas, Francesco Tiralongo, Ioannis E. Tsamadias, Luca Vecchioni, and Dimitrios K. Moutopoulos. Assessing multiple sources of data to detect illegal fishing, trade and mislabelling of elasmobranchs in Greek markets. Marine Policy, 112(??): ??, February 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19302027.

Gilman:2023:IMP

[GAD^{+}23] Eric Gilman, Kyle Antonelis, Joan Drinkwin, Saeid Gorgin, Petri Suuronen, Saly N. Thomas, and Jono Wilson. Introduction to the Marine Policy special issue on abandoned, lost and discarded fishing gear: Causes, magnitude, impacts, mitigation methods and priorities for monitoring and evidence-informed management. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002713.

Gee:2024:NMS

[GAD^{+}24] Kira Gee, Philipp Arndt, Marie Dahmen, Nele-Kristin Geitel, Anna Hunke, Bettina Käppeler, Dominic Plug, Ulrich Scheffler, and Kai Trümpler. A new maritime spatial plan for the German EEZ: Navigating multiple aspirations and spatial developments. Marine Policy, 161(??): ??, March 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000046.

Garcia-Alonso:2025:SEO

[GAGLS25] Lorena Garcia-Alonso, Fernando González-Laxe, and Ricardo J. Sanchez. Sustainability, externalities and ocean grabbing: Pressing challenges for maritime transport. Marine Policy, 174(??):??, April 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25000132.

Galligan:2021:FER

[Gal21] Bryan P. Galligan. Fisheries extractivism and the right to subsistence: Conflicting governance models and the legal structures that enact them. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003407.

Gales:2022:YOI

[Gal22] Nick Gales. At 75 years old, the IWC has never been more globally relevant. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001361.

Garcia-Alonso:2020:EWC

[GAMR20] Lorena Garcia-Alonso, Ticiana Grecco Zanon Moura, and David Roibas. The effect of weather conditions on port technical efficiency. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305895.

Gomes:2025:FBS

[GAN^{+}25] Marisa A. Gomes, Catarina M. Alves, Regina Neto, Fábio Faria, Jesus Troncoso, and Pedro T. Gomes. Fishing beyond the surface: Harnessing local ecological knowledge to improve local fisheries management and conservation. Marine Policy, 179(??):??, September 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001782.

Gamboa-Alvarez:2023:SCC

[GÁPSAP^{+}23] Miguel Á. Gamboa-Álvarez, Alicia V. Poot-Salazar, Alfonso Aguilar-Perera, R. Isaac Rojas-González, and Cuauhtémoc Ruiz-Pineda. Sea cucumber catch reconstruction and species composition as determined by seizures analyses from the IUU fishing in the Campeche Bank, Mexico. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003384.

Gardiner:2020:MPA

[Gar20] Natasha B. Gardiner. Marine protected areas in the Southern Ocean: Is the Antarctic Treaty System ready to co-exist with a new United Nations instrument for areas beyond national jurisdiction? Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20308587.

Garcia:2022:SPC

[Gar22] Isabel Artime García. Spain: a pioneering country in the fight against the infringement of the international legal regime for fisheries. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002779.

Garcia:2024:FOO

[Gar24] Eleanor L. Garcia. Fisheries observers: an overlooked vulnerability for crime and corruption within the global fishing industry. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000277.

Gonzalez-Andres:2020:IFI

[GASLCP20] Cristina González-Andrés, José Luis Sánchez-Lizaso, Jorge Cortés, and Maria Grazia Pennino. Illegal fishing in Isla del Coco National Park: Spatial–temporal distribution and the economic trade-offs. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19302192.

Gamboa-Alvarez:2025:EIF

[GÁSVPS^{+}25] Miguel Ángel Gamboa-Álvarez, Jesús Miguel Soto-Vázquez, Alicia Virginia Poot-Salazar, Sebastián Caña-Hernández, José Francisco Chávez-Villegas, Saúl Pensamiento-Villarauz, Cynthia Gutiérrez-Pérez, R. Isaac Rojas-González, and Alfonso Aguilar-Perera. Evidence of illegal fishing of Octopus maya in the Campeche Bank, Southern Gulf of Mexico. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001824.

Goldsworthy:2021:CCS

[GB21] Lynda Goldsworthy and Eaven Brennan. Climate change in the Southern Ocean: Is the Commission for the Convention for the Conservation of Antarctic Marine Living Resources doing enough? Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001603.

Garcia-Baciero:2025:MTS

[GBAECCRM25] Alberto García-Baciero, Karen Mariana Acevedo-Escobedo, Maritza Cruz-Castillo, and Dení Ramírez-Macías. Moving towards sustainable whale shark–human interactions: a case study in Bahía de La Paz, Baja California Sur, Mexico. Marine Policy, 174(??):??, April 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000211.

Giovos:2021:UPA

[GBB^{+}21] Ioannis Giovos, Adi Barash, Monica Barone, Claudio Barría, Diego Borme, Cecile Brigaudeau, Anastasia Charitou, Cristina Brito, Jock Currie, Mareike Dornhege, Lara Endrizzi, Kerstin Forsberg, Armelle Jung, Periklis Kleitou, Alison MacDiarmid, Dimitrios K. Moutopoulos, Shotaro Nakagun, João Neves, Flávia L. D. Nunes, Dorien Schröder, Ruth H. Thurstan, Malcolm Tull, Sezginer Tuncer, and Carlotta Mazzoldi. Understanding the public attitude towards sharks for improving their conservation. Marine Policy, 134(??):??, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100422X.

Galdies:2020:EPL

[GBC^{+}20] Charles Galdies, Richard Bellerby, Donata Canu, Wenting Chen, Enrique Garcia-Luque, Blaženka Gašparović, Jelena Godrijan, Paul J. Lawlor, Frank Maes, Alenka Malej, Dionisios Panagiotaras, Beatriz Martinez Romera, Claire E. Reymond, Julien Rochette, Cosimo Solidoro, Robert Stojanov, Rachel Tiller, Isabel Torres de Noronha, Grzegorz Uścinowicz, Nataşa Vaidianu, Cormac Walsh, and Roberta Guerra. European policies and legislation targeting ocean acidification in European waters — Current state. Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309054.

Gonzalez-Bernat:2021:GAG

[GBC21a] Mariá José González-Bernat and Julian Clifton. A governance analysis of Guatemala’s first recognized marine protected area: the Wildlife Refuge of Punta de Manabique (RVSPM). Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304646.

Gonzalez-Bernat:2021:GAT

[GBC21b] María José González-Bernat and Julian Clifton. A governance analysis of two marine protected areas in the Pacific Region of Guatemala: the Multiple Use Area of Monterrico and the Private Reserve La Chorrera–Manchón Guamuchal. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304610.

Gomez-Ballesteros:2021:TCM

[GBCNCL^{+}21] M. Gómez-Ballesteros, C. Cervera-Núñez, M. Campillos-Llanos, A. Quintela, L. Sousa, M. Marques, F. L. Alves, C. Murciano, N. Alloncle, P. Sala, A. Lloret, A. P. Simão, A. C. Costa, D. Carval, D. Bailly, C. Nys, H. Sybill, and J. Dilasser. Transboundary cooperation and mechanisms for Maritime Spatial Planning implementation. SIM-NORAT project. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000440.

Gutteridge:2024:HAP

[GBDC24] A. N. Gutteridge, A. Bräutigam, K. Dewar, and R. J. C. Currey. How are appropriate performance levels developed for MSC certification? A case study assessing shark finning. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001179.

Gomez:2023:MNC

[GBFMT23] Aina G. Gómez, Pau Balaguer, Àngels Fernández-Mora, and Joaquín Tintoré. Mapping the nautical carrying capacity of anchoring areas of the Balearic Islands’ coast. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003081.

Gjertsen:2020:AAA

[GBLBV20] Arild Gjertsen, Ingrid Bay-Larsen, Hilde Bjørkhaug, and Bjørn Vidar Vangelsten. Access to areas for algae cultivation in Norway. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306050.

Gallegati:2025:MSF

[GBM^{+}25] S. Gallegati, S. Bianchelli, F. Marcellini, R. G. Boschetto, A. Capriolo, C. Corinaldesi, G. Giorgi, R. A. Mascolo, E. Nepote, A. Sagrati, and R. Danovaro. Marine strategy framework directive: a comprehensive cost assessment for the Italian seas. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25001307.

Gordon:2022:EFI

[GBSC22] Jesse Y. Gordon, Anne H. Beaudreau, Emma M. Saas, and Courtney Carothers. Engaging formal and informal institutions for stewardship of rockfish fisheries in the Gulf of Alaska. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002172.

Guilfoyle:2022:LWC

[GC22] Douglas Guilfoyle and Edward Sing Yue Chan. Lawships or warships? Coast guards as agents of (in)stability in the Pacific and South and East China Sea. Marine Policy, 140 (??):??, June 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000951.

Gilman:2023:BNF

[GCB^{+}23] Eric Gilman, Milani Chaloupka, Hollie Booth, Martin Hall, Hilario Murua, and Jono Wilson. Bycatchneutral fisheries through a sequential mitigation hierarchy. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000490.

Gilbey:2021:LDS

[GCC^{+}21a] John Gilbey, Gary Carvalho, Rita Castilho, Ilaria Coscia, Mark W. Coulson, Geir Dahle, Sofie Derycke, Sara M. Francisco, Sarah J. Helyar, Torild Johansen, Claudia Junge, Kara K. S. Layton, Jann Martinsohn, Iveta Matejusova, Joana I. Robalo, Naiara Rodríguez-Ezpeleta, Gonçalo Silva, Ilona Strammer, Anti Vasemägi, and Filip A. M. Volckaert. Life in a drop: Sampling environmental DNA for marine fishery management and ecosystem monitoring. Marine Policy, 124(??):??, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309787.

Guerreiro:2021:GPM

[GCC^{+}21b] José Guerreiro, Ana Carvalho, Daniela Casimiro, Marie Bonnin, Helena Calado, Hilde Toonen, Philippe Fotso, Ibrahima Ly, Osvaldina Silva, and Solange Teles da Silva. Governance prospects for maritime spatial planning in the tropical Atlantic compared to EU case studies. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309404.

Gamarra:2023:CFH

[GCF^{+}23] N. C. Gamarra, A. C. L. Costa, M. A. C. Ferreira, L. M. Diele-Viegas, A. P. O. Santos, R. J. Ladle, A. C. Malhado, and J. V. Campos-Silva. The contribution of fishing to human well-being in Brazilian coastal communities. Marine Policy, 150(??):??, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000489.

Gollner:2021:ASC

[GCG^{+}21] S. Gollner, A. Colaço, A. Gebruk, P. N. Halpin, N. Higgs, E. Menini, N. C. Mestre, P.-Y. Qian, J. Sarrazin, K. Szafranski, and C. L. Van Dover. Application of scientific criteria for identifying hydrothermal ecosystems in need of protection. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002529.

Gomez:2021:CHE

[GCL21] Sílvia Gómez, Arnau Carreño, and Josep Lloret. Cultural heritage and environmental ethical values in governance models: Conflicts between recreational fisheries and other maritime activities in Mediterranean marine protected areas. Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001408.

Griffin:2023:AGP

[GCM^{+}23] Lucas P. Griffin, Grace A. Casselberry, Ezra M. Markowitz, Jacob W. Brownscombe, Aaron J. Adams, Bill Horn, Steven J. Cooke, and Andy J. Danylchuk. Angler and guide perceptions provide insights into the status and threats of the Atlantic tarpon (Megalops atlanticus) fishery. Marine Policy, 151(??):??, May 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000969.

Giraldi-Costa:2020:SZM

[GCMT20] Ana Clara Giraldi-Costa, Rodrigo Pereira Medeiros, and Liliani Marília Tiepolo. Step zero of marine protected areas of Brazil. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306529.

Giebels:2020:TKM

[GCP^{+}20] Diana Giebels, Jana Carus, Maike Paul, Michael Kleyer, Bernd Siebenhüner, Arne Arns, Alexander Bartholomä, Vanessa Carlow, Jürgen Jensen, Britta Tietjen, Achim Wehrmann, and Boris Schröder. Transdisciplinary knowledge management: a key but underdeveloped skill in EBM decision-making. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1930644X.

Gilman:2021:ERD

[GCS21] Eric Gilman, Milani Chaloupka, and Chrissie Sieben. Ecological risks of a data-limited fishery using an ensemble of approaches. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003638.

Guzman:2020:HWM

[GCV^{+}20] Héctor M. Guzman, Juan J. Capella, Carlos Valladares, Jorge Gibbons, and Richard Condit. Humpback whale movements in a narrow and heavily-used shipping passage, Chile. Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302049.

Giacomarra:2021:AFE

[GCV^{+}21] Marcella Giacomarra, Maria Crescimanno, Demetris Vrontis, Lluís Miret Pastor, and Antonino Galati. The ability of fish ecolabels to promote a change in the sustainability awareness. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309386.

Gerhardinger:2020:CST

[GdACT20] Leopoldo Cavaleri Gerhardinger, Mariana Martins de Andrade, Marina Ribeiro Corrêa, and Alexander Turra. Crafting a sustainability transition experiment for the Brazilian blue economy. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20305509.

Good:2024:AMS

[GDB^{+}24] Stephanie D. Good, Kate Dewar, Polly Burns, Keith Sainsbury, Richard A. Phillips, Bryan P. Wallace, Caterina Fortuna, Vinay Udyawer, Bruce Robson, Edward F. Melvin, and Rohan J. C. Currey. Adapting the Marine Stewardship Council risk-based framework to estimate impacts on seabirds, marine mammals, marine turtles and sea snakes. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001167.

GarciaCavalleirodeMacedoKlautau:2025:RGC

[GdCC^{+}25] Alex Garcia Cavalleiro de Macedo Klautau, Rafael Anaisce das Chagas, Paulo Emílio Costa Santos, Bruno Barbosa Iespa, Israel Hidenburgo Aniceto Cintra, Ana Patricia Barros Cordeiro, Wagner Cesar Rosa dos Santos, Jose Augusto Negreiros Aragão, Carlos Eduardo Matos Carvalho Bastos, Flávio de Almeida Alves-Junior, Beatrice Padovani Ferreira, and Bianca Bentes da Silva. Regulation of the Caribbean red snapper fishery on the Amazon continental shelf: 60 years of conflict between conservation and exploitation. Marine Policy, 174(??):??, April 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000375.

Gilman:2020:CPF

[GDLC20] Eric Gilman, Vicente De Ramón Castejón, Eparama Loganimoce, and Milani Chaloupka. Capability of a pilot fisheries electronic monitoring system to meet scientific and compliance monitoring objectives. Marine Policy, 113(??): ??, March 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19306608.

Guedri:2025:PCM

[GDY25] Salah Eddine Guedri, Amina Hana Djabi, and Ibrahim Yahiaoui. Pescatourism’s contribution to the management effectiveness of the Marine Protected Area (MPA) of “Taza” (Algeria, Southwestern Mediterranean). Marine Policy, 173 (??):??, March 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24005815.

Golo:2023:TCC

[GE23] Harrison K. Golo and Bolanle Erinosho. Tackling the challenges confronting women in the Elmina fishing community of Ghana: a human rights framework. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003967.

Gruby:2021:OBB

[GEB^{+}21] Rebecca L. Gruby, Ashley Enrici, Michele Betsill, Elodie Le Cornu, and Xavier Basurto. Opening the black box of conservation philanthropy: a co-produced research agenda on private foundations in marine conservation. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002566.

Garcia-Echalar:2024:GEA

[GEGRdL24] Andrés García-Echalar, Rosa G. González-Ramírez, and Daniella de Luca. Gender equity analysis in the maritime and port industry in Chile. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24004093.

Gehrke:2023:WPB

[Geh23] Charlotte Gehrke. Wild polar bear conservation: a case of successful Arctic science diplomacy? Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003160.

Gell:2025:CCM

[Gel25] Fiona R. Gell. Communicating the complexity of Marine Protected Areas in creative writing: a personal perspective. Marine Policy, 178(??):??, August 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001277.

Gabela-Flores:2021:ELD

[GFD21] María Virginia Gabela-Flores and Amy Diedrich. Exploring leadership dynamics in community-based marine protected areas. Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001809.

Gustavsson:2021:GBJ

[GFL^{+}21] Madeleine Gustavsson, Katia Frangoudes, Lars Lindström, María Catalina Ávarez, and Maricela de la Torre Castro. Gender and Blue Justice in small-scale fisheries governance. Marine Policy, 133(??):??, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003547.

Gerber:2020:COG

[GG20] Leonardus J. Gerber and Renée L. Grogan. Challenges of operationalising good industry practice and best environmental practice in deep seabed mining regulation. Marine Policy, 114:??, April 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304639.

GarciaCh:2022:ESC

[GG22] María Catalina García Ch. and Joyeeta Gupta. Environmental and sociocultural claims within maritime boundary disputes. Marine Policy, 139(??):??, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000902.

Guo:2022:PSE

[GGL22] Jianke Guo, Shu Guo, and Jing Lv. Potential spatial effects of opening Arctic shipping routes on the shipping network of ports between China and Europe. Marine Policy, 136(??): ??, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004966.

Gomez-Gomez:2024:DCO

[GGMCMM^{+}24] A. Gomez-Gomez, L. Malpica-Cruz, C. G. Montaño-Moctezuma, A. M. Cisneros-Montemayor, A. K. Salomon, and G. Seingier. Diversity, challenges, and opportunities of Fishery Improvement Projects in the Latin America and the Caribbean region. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001143.

Gacutan:2022:MSP

[GGP^{+}22] Jordan Gacutan, Ibon Galparsoro, Kemal Pınarbaşı, Arantza Murillas, Ibukun J. Adewumi, Teerapong Praphotjanaporn, Emma L. Johnston, Ken P. Findlay, and Ben M. Milligan. Marine spatial planning and ocean accounting: Synergistic tools enhancing integration in ocean governance. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005479.

Gil-Gonzalez:2025:CAM

[GGZRAHVS25] Damaris A. Gil-González, Manuel J. Zetina-Rejon, Luis C. Almendarez-Hernandez, and Francisco Vergara-Solana. Comparative analysis of the multidimensional well-being of small-scale fishers and tourism charter operators. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001903.

Guzman:2020:SCT

[GHK20] Hector M. Guzman, Natasha Hinojosa, and Stefanie Kaiser. Ship’s compliance with a traffic separation scheme and speed limit in the Gulf of Panama and implications for the risk to humpback whales. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2030021X.

Gilman:2022:MFS

[GHW^{+}22] Eric Gilman, Jenn Humberstone, Jono R. Wilson, Emmanuel Chassot, Alexis Jackson, and Petri Suuronen. Matching fishery-specific drivers of abandoned, lost and discarded fishing gear to relevant interventions. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001440.

Guillen:2025:EFP

[GICM25] Jordi Guillen, James Innes, Natacha Carvalho, and Roger Martini. Effects of fuel prices, subsidies and taxes on CO_{2} emissions and management: an application to the EU Northeast Atlantic fisheries. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25002295.

Girard:2022:CGI

[Gir22] Sophie Girard. Can Geographical Indications promote sustainable shellfish farming? The example of Bay of Mont-Saint-Michel mussels. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004565.

Guggisberg:2022:TFG

[GJS22] Solène Guggisberg, Aline Jaeckel, and Tim Stephens. Transparency in fisheries governance: Achievements to date and challenges ahead. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002505.

Gong:2023:MPR

[GJY23] Xiaoxing Gong, Huilin Jiang, and Dong Yang. Maritime piracy risk assessment and policy implications: a two-step approach. Marine Policy, 150(??):??, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300074X.

Gupta:2025:ILP

[GKO^{+}25] Trisha Gupta, Divya Karnad, Rodrigo Oyanedel, Hollie Booth, Tejaswi Abhiram, Harsha Gaonkar, and E. J. Milner-Gulland. Identifying leverage points for sustainability in India’s shark supply chains. Marine Policy, 173(??): ??, March 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24005803.

Gorjanc:2022:EPA

[GKP^{+}22] Sašo Gorjanc, Katja Klančnik, Nadia K. Papadopoulou, Arantza Murillas-Maza, Klara Jarni, Theodora Paramana, Mišo Pavičić, Francesca Ronchi, Maria C. Uyarra, Špela Koren, Manos Dassenakis, Olja Vidjak, Christopher J. Smith, and Sanda Skejić. Evaluating the progress in achieving Good Environmental Status in the Mediterranean: a methodology to assess the effectiveness of Marine Strategy Framework Directive’s Programmes of Measures. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005005.

Garcia-Lorenzo:2023:PSS

[GLADGGVL23] Iria García-Lorenzo, Lucy Amigo-Dobaño, María Dolores Garza-Gil, and Manuel Varela-Lafuente. Perceptions in small-scale fisheries regarding institutional, economic, technological and environmental factors. Case study in North-Western Spain. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300057X.

Garcia-Lorenzo:2021:CBF

[GLAVL21] Iria García-Lorenzo, Dewan Ahsan, and Manuel Varela-Lafuente. Community-based fisheries organisations and sustainable development: Lessons learned from a comparison between European and Asian countries. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002839.

Golden:2024:EAC

[GLB^{+}24] Abigail S. Golden, Arielle Levine, Marissa L. Baskett, Dan Holland, Katherine E. Mills, Jacqueline M. Vogel, and Timothy Essington. Enhancing the adaptive capacity of fisheries to climate change: Bridging academic theory and management practice through practitioner interviews. Marine Policy, 168(??):??, October 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003191.

Gonzalez-Laxe:2025:PNE

[GLMBMP25] Fernando González-Laxe, Federico Martín-Bermúdez, and Federico Martín-Palmero. Ports and negative externalities: Institutional strategies to monitor and mitigate them in Spain. Marine Policy, 181(??):??, November 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002489.

Gardner:2024:MCC

[GLPM24] Jonathan P. A. Gardner, Barbara Lausche, Simon J. Pittman, and Anna Metaxas. Marine connectivity conservation: Guidance for MPA and MPA network design and management. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002483.

Garren:2021:HPS

[GLS^{+}21] Melissa Garren, Forrest Lewis, Laura Sanchez, Daniella Spina, and Annie Brett. How performance standards could support innovation and technology-compatible fisheries management frameworks in the U.S. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002426.

Gonzalez-Laxe:2025:TYS

[GLSVMB25] Fernando González-Laxe, Javier Seijo-Villamizar, and Federico Martín-Bermúdez. Twenty years of sustainable development and port authorities: a critical review of the literature. Marine Policy, 172(??):??, February 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004913.

Guo:2023:CAO

[GLT23] Jianping Guo, Wanqiang Li, and Haoyu Tian. The climate advisory opinion: a medicine with side-effects? Marine Policy, 156(??):??, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003500.

Gomez:2021:ASM

[GM21] Sílvia Gómez and Francesc Maynou. Alternative seafood marketing systems foster transformative processes in Mediterranean fisheries. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000427.

Galveia:2025:MEM

[GM25a] Maria C. Galveia and Edison D. Macusi. Management effectiveness of marine protected areas (MPAs) in the Southeastern Mindanao, Philippines. Marine Policy, 174(??):??, April 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25000119.

Goto:2025:CPC

[GM25b] Jun Goto and Takahiro Matsui. Counterproductive punishment in closed communities: Experimental evidence from Japanese fishery. Marine Policy, 177(??):??, July 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004937.

Guet:2024:SAM

[GMC24] Alice Thébault Guët, Jason Monios, and Pierre Cariou. Successful adoption of maritime environmental policy: the Mediterranean emission control area. Marine Policy, 166(??):??, August 2024. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002264.

Guerra-Marrero:2021:PAI

[GMCMJA^{+}21] Airam Guerra-Marrero, Lorena Couce-Montero, David Jiménez-Alvarado, Ana Espino-Ruano, Raibel Núñez-González, Airam Sarmiento-Lezcano, Ángelo Santana del Pino, and José J. Castro. Preliminary assessment of the impact of Covid-19 Pandemic in the small-scale and recreational fisheries of the Canary Islands. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003237.

Gallegati:2025:INC

[GMFD25] Silvia Gallegati, Paula Masiá, Emanuela Fanelli, and Roberto Danovaro. The impact of natural capital loss on blue-tourism economy: the Red Sea case study. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24005074.

GoncalvesLoureiro:2023:OAA

[GMG^{+}23] Tainã Gonçalves Loureiro, Ben Milligan, Jordan Gacutan, Ibukun J. Adewumi, and Ken Findlay. Ocean accounts as an approach to foster, monitor, and report progress towards sustainable development in a changing ocean — the Systems and Flows Model. Marine Policy, 154(??):??, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001951.

Good:2024:URE

[GMG^{+}24] Stephanie D. Good, Shaun McLennan, Matt Gummery, Rebecca Lent, Timothy E. Essington, Bryan P. Wallace, Richard A. Phillips, Tom Peatman, G. Barry Baker, Keith Reid, and Rohan J. C. Currey. Updating requirements for Endangered, Threatened and Protected species MSC Fisheries Standard v3.0 to operationalise best practices. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001155.

Gonzales-Mantilla:2021:CCP

[GMGLV21] Pedro G. Gonzáles-Mantilla, Austin J. Gallagher, Carmelo J. León, and Gabriel M. S. Vianna. Challenges and conservation potential of shark-diving tourism in the Macaronesian archipelagos. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002438.

Gonzales-Mantilla:2022:EIC

[GMGLV22] Pedro G. Gonzáles-Mantilla, Austin J. Gallagher, Carmelo J. León, and Gabriel M. S. Vianna. Economic impact and conservation potential of shark-diving tourism in the Azores Islands. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004802.

Gutteridge:2024:MAD

[GML^{+}24] A. N. Gutteridge, M. C. Melnychuk, C. Longo, R. J. C. Currey, and E. Jardim. The MSC approach for developing harvest strategies and harvest control rules for Regional Fisheries Management Organisations (RFMOs). Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000022.

Gatti:2020:RDD

[GMM^{+}20] Paul Gatti, Sonia Méhault, Fabien Morandeau, Marie Morfin, and Marianne Robert. Reducing discards of demersal species using a 100 mm square mesh cylinder: Size selectivity and catch comparison analysis. Marine Policy, 116 (??):??, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X18307358.

Gonzalez-Medina:2025:HRF

[GMMMRC^{+}25] Tránsito González-Medina, Gabriela Montaño-Moctezuma, Laura Rodríguez-Cardozo, Jorge Torre, Alejandro Espinoza-Tenorio, and Nuria Jiménez García. Historical reconstruction of Fishing Refuge Zones in Mexico: an innovative marine conservation and sustainable fisheries co-management tool. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24005050.

Gamaza-Marquez:2020:DPG

[GMPT^{+}20] Maria Angeles Gamaza-Márquez, Maria Grazia Pennino, Maria Angeles Torres, Juan Jose Acosta, Karim Erzini, and Ignacio Sobrino. Discard practices in the Gulf of Cadiz multispecies trawl fishery. Implications for the EU ‘landing obligation’. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302349.

Grace-McCaskey:2023:CSN

[GMRSP23] Cynthia A. Grace-McCaskey, Maria C. Ramos, Anja Sjostrom, and Sarah E. Page. Combining social network analysis and ethnography to better understand fishers’ organization and promote sustainable small-scale fisheries in St. Croix, US Virgin Islands. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001008.

Guzman:2023:MPA

[GMVM23] Diego Hernández Guzmán, Roberto Lastra Mier, Alba Vergara, and Celene B. Milanes. Marine protected areas in Colombia: a historical review of legal marine protection since the late 1960s to 2023. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002592.

Gibbs:2022:RHN

[GN22] Mark T. Gibbs and Maxine Newlands. Restoration heralds’ new management challenges for coral reefs. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005224.

Gaebel:2024:ISK

[GNJR24] Christine Gaebel, Paula Novo, David E. Johnson, and J. Murray Roberts. Institutionalising science and knowledge under the agreement for the conservation and sustainable use of marine biodiversity of areas beyond national jurisdiction (BBNJ): Stakeholder perspectives on a fit-for-purpose Scientific and Technical Body. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23005316.

Grillo-Nunez:2021:ICV

[GNMGWM21] J. Grillo-Núñez, T. Mendo, R. Gozzer-Wuest, and J. Mendo. Impacts of COVID-19 on the value chain of the hake small scale fishery in northern Peru. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100419X.

Gray:2020:SCA

[GOH^{+}20] S. Gray, C. O’Mahony, J. Hills, B. O’Dwyer, R. Devoy, and J. Gault. Strengthening coastal adaptation planning through scenario analysis: a beneficial but incomplete solution. Marine Policy, 111(??):??, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16302275.

Goncalves:2025:GBR

[Gon25] Leandra R. Gonçalves. Brazil and regional fisheries management organizations — what is at stake? Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004871.

Gordon:2021:PRP

[Gor21] Daniel V. Gordon. Price and revenue projections under alternative policy shocks due to the coronavirus: Canadian lobster and snow crab. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001676.

Govella:2025:UCG

[Gov25] Kristi Govella. Undersea cables, geoeconomics, and security in the Indo–Pacific: Risks and resilience. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25002246.

Gacutan:2022:EIB

[GPA^{+}22] Jordan Gacutan, Kemal Pınarbaşı, Messan Agbaglah, Crystal Bradley, Ibon Galparsoro, Arantza Murillas, Ibukun Adewumi, Teerapong Praphotjanaporn, Michael Bordt, Kenneth Findlay, Coulson Lantz, and Ben M. Milligan. The emerging intersection between marine spatial planning and ocean accounting: a global review and case studies. Marine Policy, 140(??):??, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001026.

Galparsoro:2021:OES

[GPG^{+}21] Ibon Galparsoro, Kemal Pınarbaşı, Elena Gissi, Fiona Culhane, Jordan Gacutan, Jonne Kotta, David Cabana, Sonja Wanke, Robert Aps, Daniel Bazzucchi, Giacomo Cozzolino, Marco Custodio, Mihhail Fetissov, Miguel Inácio, Susanna Jernberg, Alessandro Piazzi, Keshav Prasad Paudel, Alex Ziemba, and Daniel Depellegrin. Operationalisation of ecosystem services in support of ecosystem-based marine spatial planning: insights into needs and recommendations. Marine Policy, 131(??):??, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002207.

Garbett:2021:CEF

[GPH^{+}21] Amy Garbett, Natasha D. Phillips, Jonathan D. R. Houghton, Paulo Prodöhl, James Thorburn, Sophie L. Loca, Lawrence E. Eagling, Gary Hannon, Daniel Wise, Liz Pothanikat, Cat Gordon, Maurice Clarke, Peter Williams, Rebecca Hunter, Ronan McShane, Aafke Brader, Jane Dodd, Chris McGonigle, Heidi McIlvenny, Olivia Daly, Ronald Surgenor, Sarah Varian, Peter Verhoog, Gijs Van Zonneveld, Lylian R. Burke, Ian Davies, Terri A. Souster, Paul A. Mayo, Tanja N. Schwanck, Catherine S. Jones, and Patrick C. Collins. The critically endangered flapper skate (Dipturus intermedius): Recommendations from the first flapper skate working group meeting. Marine Policy, 124(??):??, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310186.

Galappaththi:2024:FSP

[GPI^{+}24] Eranga K. Galappaththi, Chrishma D. Perera, Gayanthi A. Illangarathna, Sithuni M. Jayasekara, and Hannah Garbutt. Food security policy and coastal climate adaptation among Indigenous and local communities. Marine Policy, 170(??): ??, December 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24004068.

Gkargkavouzi:2020:PPM

[GPM20] Anastasia Gkargkavouzi, Stefanos Paraskevopoulos, and Steriani Matsiori. Public perceptions of the marine environment and behavioral intentions to preserve it: the case of three coastal cities in Greece. Marine Policy, 111(??): ??, January 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19301290.

Gomez:2023:ISS

[GPM23] Sílvia Gómez, Beatriz Patraca, and José Luis Molina. Improving seafood systems with social network analysis: the case of cuttlefish marketing in Catalonia. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000441.

Gonzalez-Pestana:2023:VNP

[GPTASM23] Adriana Gonzalez-Pestana, Daniela C. S. Thorne, Joanna Alfaro-Shigueto, and Jeffrey C. Mangel. Vulnerabilities of northern Peruvian small-scale fishing communities revealed by the COVID-19 pandemic. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000301.

Gerwing:2020:CCP

[GR20a] Travis G. Gerwing and Joshua A. Rash. Constructive and collegial peer-review as a necessary precursor to datadriven environmental policy. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19303367.

Green:2020:CHS

[GR20b] Jessica F. Green and Bryce Rudyk. Closing the high seas to fishing: a club approach. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19303161.

Gaertner:2025:OAB

[GR25] Daniel Gaertner and Nastassia Reyes. Oppositions and alliances between ICCAT contracting parties through an analysis of co-sponsorship of management recommendations. Marine Policy, 181(??):??, November 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002465.

Gray:2021:NTI

[Gra21] Tim Gray. Normative theory of international relations and the ‘mackerel war’ in the North East Atlantic. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002311.

Geiger:2022:CPE

[GRA^{+}22] Katja J. Geiger, Antonella Rivera, Alba Aguión, Jorge Álvarez, Julio Arrontes, Yaisel Juan Borrell, Teresa Cruz, Dominique Davoult, Jesús Dubert, Marieke E. Feis, Joana N. Fernandes, Consolación Fernández, Lucía García-Flórez, David Jacinto, Didier Jollivet, Gonzalo Macho, Elena Mateo, David Mateus, Paloma Morán, Carlota Muñiz, Amandine Nicolle, Rita Nolasco, Marina Parrondo, Henrique Queiroga, José Rico, Alina Sousa, Salvador Román, Teresa Silva, Eric Thiébaut, Elsa Vázquez, and José Luis Acuña. Coping with poachers in European stalked barnacle fisheries: Insights from a stakeholder workshop. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004371.

Groeneveld:2020:WEW

[Gro20] Rolf A. Groeneveld. Welfare economics and wicked problems in coastal and marine governance. Marine Policy, 117 (??):??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19303951.

Guirkinger:2021:MCP

[GRPE^{+}21] Lucie Guirkinger, Stefany Rojas-Perea, Isabel Ender, Mark Ramsden, Charley Lenton-Lyons, and Jonas Geldmann. Motivations for compliance in Peruvian manta ray fisheries. Marine Policy, 124(??):??, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309623.

Gonzalez-Roca:2023:AFP

[GRPRVG23] Fernando González-Roca, Ángel Pérez-Ruzafa, Julio A. Vásquez, and Stefan Gelcich. Artisanal fishers’ perceptions of two co-existing kelp resource management regimes in northern Chile. Marine Policy, 153(??):??, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001379.

Gallagher:2023:ALO

[GRS^{+}23] Anthony Gallagher, Peter Randall, David Sivyer, Umberto Binetti, Gayathri Lokuge, and Mohamed Munas. Abandoned, lost or otherwise discarded fishing gear (ALDFG) in Sri Lanka — a pilot study collecting baseline data. Marine Policy, 148(??):??, February 2023. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200433X.

Garcia-Rodriguez:2024:BEF

[GRSRSN^{+}24] Emiliano García-Rodríguez, Luz Erandi Saldaña-Ruiz, Oscar Sosa-Nishizaki, Arturo Fajardo-Yamamoto, Frida Cisneros-Soberanis, and Alesa Flores-Guzmán. From bottom to up: Effects of fishery improvement projects on the stock trends of multi-specific small-scale fisheries from Mexico. Marine Policy, 168(??):??, October 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002902.

Goodman:2024:ICG

[GRvdM24] Camille Goodman, Mercedes Rosello, and Eva R. van der Marel. Innovating for change in global fisheries governance (again): an introduction. Marine Policy, 169(??): ??, November 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003543.

Gundelund:2021:CAD

[GS21] Casper Gundelund and Christian Skov. Changes in angler demography and angling patterns during the Covid-19 lockdown in spring 2020 measured through a citizen science platform. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100213X.

Gangal:2023:HWD

[GSA23] Mayuresh Gangal, Venkatachalam Suri, and Rohan Arthur. How well does Indian fisheries policy engage with fisheries biology? Exploring the science-policy interface of coastal capture fisheries along the west coast of India. Marine Policy, 156(??):??, October 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003299.

Garmendia:2021:UFK

[GSAF21] Vladimir Garmendia, Maria Dulce Subida, Ainara Aguilar, and Miriam Fernández. The use of fishers’ knowledge to assess benthic resource abundance across management regimes in Chilean artisanal fisheries. Marine Policy, 127 (??):??, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100035X.

Groba:2020:OTF

[GSB20] Carlos Groba, Antonio Sartal, and Gustavo Bergantiño. Optimization of tuna fishing logistic routes through information sharing policies: a game theory-based approach. Marine Policy, 113(??):??, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306699.

Glaus:2024:CGF

[GSB24] Kerstin Glaus, Rusila Savou, and Juerg M. Brunnschweiler. Characteristics of Fiji’s small-scale ray fishery and its relevance to food security. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000800.

Goldsmith:2020:DMM

[GSG20] William M. Goldsmith, Andrew M. Scheld, and John E. Graves. Decision-making in a mixed commercial-recreational fishery for Atlantic bluefin tuna. Marine Policy, 115(??): ??, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19303495.

Gbedomon:2024:SSM

[GSG^{+}24] Rodrigue Castro Gbedomon, Kolawolé Valère Salako, Setondé Constant Gnansounou, Kisito Gandji, Pierre Failler, Achille Ephrem Assogbadjo, and Romain Glèlè Kakaï. Small-scale marine fishing in Benin, West Africa: a comprehensive assessment of the processed fish value chain. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000307.

Garcia-Sanabria:2021:MSP

[GSGOP^{+}21] Javier García-Sanabria, Javier García-Onetti, Víctor Cordero Penín, María de Andrés, Claudia Millán Caravaca, Eleonora Verón, and Cristina Pallero-Flores. Marine Spatial Planning cross-border cooperation in the ‘European Macaronesia Ocean’: a participatory approach. Marine Policy, 132(??): ??, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002827.

Ginzky:2020:SIS

[GSM20] Harald Ginzky, Pradeep A. Singh, and Till Markus. Strengthening the International Seabed Authority’s knowledgebase: Addressing uncertainties to enhance decisionmaking. Marine Policy, 114:??, April 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19300661.

Garcia:2021:MRE

[GSR21] Pablo Quero García, Javier García Sanabria, and Juan Adolfo Chica Ruiz. Marine renewable energy and maritime spatial planning in Spain: Main challenges and recommendations. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000543.

Gamaza:2020:WRI

[GTA^{+}20] MariAngeles Gamaza, Maria Angeles Torres, Juan Jose Acosta, Karim Erzini, and Ignacio Sobrino. Are we ready to implement a discard ban in the Gulf of Cádiz? Stakeholders’ perceptions. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307528.

Guillotreau:2020:ATP

[GTMP20] Patrice Guillotreau, Brice Trouillet, Stéphanie Mahévas, and Sophie Pardo. Addressing transdisciplinary and participation issues to cope with rapid changes shifting marine social ecological systems. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20302463.

Grover:2021:IPP

[GTTM21] Isobella M. Grover, Mark S. Tocock, Dugald R. Tinch, and Darla Hatton MacDonald. Investigating public preferences for the management of native and invasive species in the context of kelp restoration. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002918.

Guelker:2023:FSI

[Gue23] Daphne Guelker. Fishers and seafarers in international law — really so different? Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22005243.

Guelker:2024:FSG

[Gue24] Daphne Guelker. Finding stable ground: Moving from ‘sea slavery’ to ‘continuum of exploitation’. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2400263X.

Guggisberg:2022:TAF

[Gug22] Solène Guggisberg. Transparency in the activities of the Food and Agriculture Organization for sustainable fisheries. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001093.

Guggisberg:2024:FES

[Gug24] Solène Guggisberg. Finding equitable solutions to the landbased sources of marine plastic pollution: Sovereignty as a double-edged sword. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004931.

Gulseven:2020:MAT

[Gul20] Osman Gulseven. Measuring achievements towards SDG 14, life below water, in the United Arab Emirates. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20300488.

Gullett:2023:CLS

[Gul23] Warwick Gullett. The contribution of the Law of the Sea Convention to marine environmental protection: 40 years of shimmering seas or high hopes dashed against juridical rocks? Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000258.

Gunnarsson:2021:RST

[Gun21] Björn Gunnarsson. Recent ship traffic and developing shipping trends on the Northern Sea Route — Policy implications for future Arctic shipping. Marine Policy, 124(??): ??, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20310204.

Guo:2020:DME

[Guo20] Jianping Guo. The developments of marine environmental protection obligation in article 192 of UNCLOS and the operational impact on China’s marine policy — a South China Sea fisheries perspective. Marine Policy, 120(??): ??, October 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19308917.

Gunnlaugsson:2022:SWC

[GV22] Stefan B. Gunnlaugsson and Hreidar Valtysson. Sustainability and wealth creation, but no consensus: Recent decades in Iceland’s ITQ-managed fisheries. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004474.

Gleadall:2025:BAP

[GVB^{+}25] Ian G. Gleadall, Roger Villanueva, Gregory J. Barord, Zoe Doubleday, Felipe Aguado-Giménez, Nobuhiko Akiyama, Eduardo Almansa, Cheryl L. Ames, Alexander Arkhipkin, Otilio Avendaño, Chris Barrett, Giambattista Bello, John R. Bower, Ramiro Braga, Felipe A. Briceño, Paco Bustamante, Claudia Caamal-Monsreal, Miguel Cabanellas-Reboredo, Sergio A. Carrasco, Sheila Castellanos-Martínez, Jesús Cerezo Valverde, Leo Ji-Ho Che, Wen-Sung Chung, Shigeki Dan, Mariana Díaz-Santana-Iturrios, Pedro Domingues, Erica Donlon Durante, Alejandro Escánez, Viviana Espinoza, Ana Farías, Fernando Ángel Fernández-Álvarez, Pedro Ferreiro-Velasco, Graziano Fiorito, Hidetaka Furuya, Pedro Gallardo, Kostas Ganias, Camino Gestal, Alexey V. Golikov, Ángel F. González, Roberto González-Gómez, Jessica Gordon, Angel Guerra, Juergen Guerrero-Kommritz, Karina Hall, Manuel Haimovici, Katsuyuki Hamasaki, Jorge Hernández-Urcera, Jorge Hernández-Velásquez, Noritaka Hirohashi, Kazuki Hirota, Neil Hutchinson, Pamela Imperadore, Yoko Iwata, Maria de Lourdes Jiménez-Badillo, Yoshiki Kato, Oleg N. Katugin, Ryosuke Kimbara, Zdenek Lajbner, Germaine Lau, Unai Markaida, Lorenzo Marquez, M. Virginia Martín, Maite Mascaro, Natalie Moltschaniwskyj, Óscar Monroig, Amalia E. Morales, Ana Moreno, Piedad S. Morillo-Velarde, Jaruwat Nabhitabhata, Manuel Nande, Juan C. Navarro, Goh Nishitani, Harini Nishshanka, Atsushi Ogura, Aurelio Ortega, Nicolás Ortiz, Jaime Otero, Rodrigo Oyanedel, Yumeng Pang, Cristina Pascual, Catalina Perales-Raya, João Manuel Figueiredo Pereira, Cristina Pita, Giovanna Ponte, Anne Marie Power, Delta Putra, Antoni Quetglas, Tiago Repolho, Jean-Paul Robin, Francisco Rocha, Alex Romero, Rui Rosa, Carlos Rosas, Rigoberto Rosas-Luis, Katina Roumbedakis, Álvaro Roura, Rushan M. Sabirov, João B. L. Sales, Sanchez Pilar Sánchez, Noriyosi Sato, Warwick H. H. Sauer, Paul W. Shaw, Shuichi Shigeno, Roxana De Silva-Dávila, Chikatoshi Sugimoto, Yasuo Tsukahara, Iker Uriarte, Maria Valls, Silvina Van der Molen, Inmaculada Varó, Iván Velázquez-Abunader, Diego G. Vilarnau, José C. Xavier, Masa aki Yoshida, Xiumei Zhang, Jian Zheng, Xiaodong Zheng, and Mehmet Arif Zoral. A balanced approach to the potential of octopus aquaculture. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25000971.

Gundelund:2023:SMC

[GVH^{+}23] Casper Gundelund, Paul Venturelli, Bruce W. Hartill, Kieran Hyder, Hans Jakob Olesen, and Christian Skov. Spatiotemporal and multispecies comparisons between a citizen science platform and recall surveys in recreational fisheries. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003135.

Gulbrandsen:2022:NLF

[GVL22] Lars H. Gulbrandsen, Irja Vormedal, and Mari Lie Larsen. No logo? The failure of ASC salmon labeling in Norway and the UK. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000343.

Greaker:2020:EPI

[GVR20] Mads Greaker, Irja Vormedal, and Kristin Rosendal. Environmental policy and innovation in Norwegian fish farming: Resolving the sea lice problem? Marine Policy, 117(??): ??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19309686.

Guilan:2021:SMS

[GW21] Lin Guilan and Dong Weiwei. Synergetic management strategy for maritime cultural heritage protection and marine development in China. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310344.

Gozzer-Wuest:2022:WRD

[GWAPRPRU22] Renato Gozzer-Wuest, Enrique Alonso-Población, Stefany Rojas-Perea, and Rubén H. Roa-Ureta. What is at risk due to informality? Economic reasons to transition to secure tenure and active co-management of the jumbo flying squid artisanal fishery in Peru. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004978.

Gozzer-Wuest:2021:IPA

[GWAPT21] Renato Gozzer-Wuest, Enrique Alonso-Población, and Geoffrey A. Tingley. Identifying priority areas for improvement in Peruvian Fisheries. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001561.

Gaymer:2025:RAC

[GWÁV^{+}25] C. F. Gaymer, D. Wagner, R. Álvarez-Varas, B. Boteler, L. Bravo, C. M. Brooks, V. Chavez-Molina, D. Currie, J. Delgado, B. Dewitte, E. E. Easton, A. M. Friedlander, M. A. Gallardo, M. Gianni, K. Gjerde, M. Gorny, S. Hormazábal, R. Hucke-Gaete, G. Luna-Jorquera, A. Mecho, N. Morales-Serrano, L. Morgan, P. Nuñez, M. Ramos, J. Rapu, C. Rodrigo, J. Sellanes, E. Soto, M. Thiel, L. Van der Meer, and D. Véliz. Research advances and conservation needs for the protection of the Salas y Gómez and Nazca ridges: a natural and cultural heritage hotspot in the southeastern Pacific ocean. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24004536.

Guillot-Wright:2022:IED

[GWCD22] Shannon Guillot-Wright, Ellie Cherryhomes, and Lacy Davis. The impact of economic deregulation for health disparities among Gulf of Mexico commercial fishermen. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001464.

Gou:2023:DPI

[GY23] Yiping Gou and Cuibai Yang. Dilemmas and paths of international cooperation in China’s fight against IUU fishing analysis. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003226.

Greer:2020:CGT

[GZW^{+}20] Krista Greer, Dirk Zeller, Jessika Woroniak, Angie Coulter, Maeve Winchester, M. L. Deng Palomares, and Daniel Pauly. Corrigendum to ‘Global trends in carbon dioxide (CO_{2}) emissions from fuel combustion in marine fisheries from 1950 to 2016’ [Marine Policy, 107 (2019), 103382]. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308149.

Hoenicka:2022:RCH

[HACCM22] Ma. Alejandra Koeneke Hoenicka, Sara Andreotti, Humberto Carvajal-Chitty, and Conrad A. Matthee. The role of controlled human-animal interactions in changing the negative perceptions towards white sharks, in a sample of White Shark cage diving tours participants. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001774.

Haas:2024:PTU

[HAD24] Bianca Haas, Kamal Azmi, and Ruth Davis. Pacific tuna update: 20 ^{th} Meeting of the Western and Central Pacific Fisheries Commission. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2400318X.

Hannesson:2020:NCA

[Han20] Rögnvaldur Hannesson. The Nash–Cournot approach to shared fish stocks: an empirical investigation. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309431.

Hanh:2021:WFA

[Han21] Tong Thi Hai Hanh. Why are fisheries agencies unable to facilitate the development of alternative livelihoods in small-scale fisheries and aquaculture in the global South? A case study of the Tam Giang lagoon, Viet Nam. Marine Policy, 133(??):??, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003894.

Hart:2021:SSG

[Har21] Paul J. B. Hart. Stewards of the sea. Giving power to fishers. Marine Policy, 126(??):??, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000312.

Harrison:2022:ITO

[Har22] James Harrison. International transparency obligations in fisheries conservation and management: Inter-state and intra-state dimensions. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19306153.

Hassanali:2020:CBE

[Has20] Kahlil Hassanali. CARICOM and the blue economy — Multiple understandings and their implications for global engagement. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301123.

Hassanali:2023:ABA

[Has23] Kahlil Hassanali. The Agreement on Biodiversity in Areas Beyond National Jurisdiction — Caribbean Community perspectives on interests, asks and outcomes. Marine Policy, 156(??):??, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003330.

Hidayah:2024:SAM

[HAsR24] Zainul Hidayah, Abd. Rahman As-syakur, and Herlambang Aulia Rachman. Sustainability assessment of mangrove management in Madura Strait, Indonesia: a combined use of the rapid appraisal for mangroves (RAPMangroves) and the remote sensing approach. Marine Policy, 163(??): ??, May 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2400126X.

Hernandez-Blanco:2024:CAT

[HBCA^{+}24] Marcello Hernández-Blanco, Robert Costanza, Tundi Agardy, Joshua Farley, Elizabeth Fulton, Ida Kubiszewski, and Rifaee Rasheed. Common Asset Trusts for blue commons stewardship. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004906.

Deme:2022:SPF

[hBDDF22] El hadj Bara Deme, Moustapha Deme, and Pierre Failler. Small pelagic fish in Senegal: a multi-usage resource. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001300.

Deme:2022:PPS

[hBDF22] El hadj Bara Dème and Pierre Failler. Public policies to support artisanal fishing in Senegal: Between inconsistency and perverse effects. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000598.

Deme:2024:SPW

[hBDF24] El hadj Bara Dème and Pierre Failler. Small pelagics in West Africa face the multiple challenges of food security, wealth creation and regional governance. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003725.

Hurtubise:2020:EMR

[HBHJ20] Jessica A. Hurtubise, Emily P. Bond, Kelly E. Hall, and Michael C. James. Evaluating mandatory reporting of marine turtle bycatch in Atlantic Canadian fisheries. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930942X.

Hosch:2021:CET

[HBJ21] Gilles Hosch, Ben Belton, and Gareth Johnstone. Catch and effort trends in Myanmar’s offshore fleets operating out of Myeik — 2009–2018. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2030943X.

Huse:2021:CSF

[HBL^{+}21] Geir Huse, Åsmund Bjordal, Harald Loeng, Kari Østervold Toft, and Reidar Toresen. Commitments to sustainable fisheries: Empty words or reality? Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20310113.

Hoque:2021:KPE

[HBP^{+}21] Md. Sazedul Hoque, Despena Andrioti Bygvraa, Kate Pike, Md. Mahmudul Hasan, Muhammad Arifur Rahman, Shaida Akter, Darren Mitchell, and Eric Holliday. Knowledge, practice, and economic impacts of COVID-19 on small-scale coastal fishing communities in Bangladesh: Policy recommendations for improved livelihoods. Marine Policy, 131(??):??, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100258X.

Heck:2021:SRM

[HBR21] Nadine Heck, Michael W. Beck, and Borja Reguero. Storm risk and marine fisheries: a global assessment. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003092.

Heck:2023:GCC

[HBR^{+}23] Nadine Heck, Michael W. Beck, Borja Reguero, Kerstin Pfliegner, Max Ricker, and Ruben Prütz. Global climate change risk to fisheries — a multi-risk assessment. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004511.

Hsiao:2021:DPS

[HC21] Yao-Jen Hsiao and Jyun-Long Chen. Different perspectives of stakeholders on the sustainable development of fishery-based communities in Northeast Taiwan. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001871.

Ho:2022:PRC

[HC22] Li-Te Ho and Kuei-Chao Chang. Promoting relational collaborative partnership to implement coastal settlement revitalization. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002895.

Choi:2024:TBW

[hC24] Jee hyun Choi. Trans-boundary water management during the peace process in Korean peninsula: Solution finding from previous precedents. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004992.

Hayashida-Carrillo:2025:DPG

[HC25] Ana Harumi Hayashida-Carrillo. Do participatory governance models in fisheries lead to ecosystem-based fisheries management? Marine Policy, 174(??):??, April 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000090.

Hynes:2024:AAM

[HCDN24] Stephen Hynes, Mary Cawley, John Deely, and Daniel Norton. Alternative approaches to measuring the value of tourism in marine and coastal areas in ocean economy accounting. Marine Policy, 168(??):??, October 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002975.

Henry:2025:NTP

[HCG^{+}25] Lea-Anne Henry, Jason Cleland, Anna Gebruk, Richard Emmerson, Janos Hennicke, Tammy Davies, and J. Murray Roberts. Navigating a transformative policy route for high seas conservation. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25002003.

Huang:2021:SPT

[HCS21] Chao-Chin Huang, Shui-Kai Chang, and Shiahn-Wern Shyue. Sustain or phase out: Transformation of Taiwan’s management scheme on distant water tuna longline fisheries. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309428.

Hernandez-Casas:2023:ASD

[HCSBM^{+}23] Sergio Hernández-Casas, Juan Carlos Seijo, Luis Felipe Beltrán-Morales, Álvaro Hernández-Flores, Francisco Arreguín-Sánchez, and Germán Ponce-Díaz. Analysis of supply and demand in the international market of major abalone fisheries and aquaculture production. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004523.

Hung:2021:IMP

[HCT^{+}21] Yi-Ping Hung, Yi Chang, Minh-Phuong N. Truong, Jen-Han Yang, Tung-Yao Hsu, and Shih-Chun Hsiao. Improving marine protected area with coordination platform:

Mud shrimp conservation in Taiwan case study. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002189.

Han:2023:ICC

[HCW^{+}23] Wenjie Han, Jihong Chen, Kui Wei, Jia Shi, and Guoling Jia. International crew changes amid global pandemic outbreaks: Key issues and system innovations. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200389X.

Haugen:2021:RAC

[HCWC21] Brianna I. Haugen, Lori A. Cramer, George G. Waldbusser, and Flaxen D. L. Conway. Resilience and adaptive capacity of Oregon’s fishing community: Cumulative impacts of climate change and the graying of the fleet. Marine Policy, 126 (??):??, April 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000348.

Hernandez-Delgado:2024:SPC

[HDAPG^{+}24] Fredy Hernández-Delgado, Alfonso Aguilar-Perera, Vinicius J. Giglio, Virginia Nóh-Quiñones, Jorge I. Euán-Ávila, Wilian de Jesús Aguilar-Cordero, and Celia I. Sélem-Salas. Stakeholders’ perception on consumption, fishing, and conservation of red grouper, Epinephelus morio, off the northern coast of the Yucatan Peninsula, Mexico. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005328.

Haas:2021:RFMb

[HDCH21] Bianca Haas, Ruth Davis, Brooke Campbell, and Quentin Hanich. Regional fisheries management: COVID-19 calendars and decision making. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000853.

Haas:2021:RFMa

[HDH21] Bianca Haas, Ruth Davis, and Quentin Hanich. Regional fisheries management: Virtual decision making in a pandemic. Marine Policy, 125(??):??, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309349.

Harden-Davies:2020:RNP

[HDHM^{+}20] Harriet Harden-Davies, Fran Humphries, Michelle Maloney, Glen Wright, Kristina Gjerde, and Marjo Vierros. Rights of Nature: Perspectives for Global Ocean Stewardship. Marine Policy, 122(??):??, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20303365.

Hersoug:2022:OCT

[Her22] Bjørn Hersoug. “One country, ten systems” — the use of different licensing systems in Norwegian aquaculture. Marine Policy, 137(??):??, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005133.

Haya:2020:ACR

[HF20] La Ode Muhammad Yasir Haya and Masahiko Fujii. Assessment of coral reef ecosystem status in the Pangkajene and Kepulauan Regency, Spermonde Archipelago, Indonesia, using the rapid appraisal for fisheries and the analytic hierarchy process. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930524X.

Hair:2020:SEC

[HFD^{+}20] Cathy Hair, Simon Foale, Nicholas Daniels, Peter Minimulu, John Aini, and Paul C. Southgate. Social and economic challenges to community-based sea cucumber mariculture development in New Ireland Province, Papua New Guinea. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307316.

Hsu:2024:EPS

[HFY^{+}24] Wen-Kai Hsu, Johonsan A. Fabilane, Hui-Lung Yu, Guan-Wei Wang, and Wen-Hong Liu. Enhancing port security in Taiwan by establishing security inspection procedure to combat fisheries-related crimes. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24002847.

Haghighatjou:2022:RAA

[HGG^{+}22a] N. Haghighatjou, S. Gorgin, R. Ghorbani, E. Gilman, R. A. Naderi, H. Raeisi, and Sh Farrukhbin. Rate and amount of abandoned, lost and discarded gear from the Iranian Persian Gulf Gargoor pot fishery. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001476.

Hamilton:2022:EBM

[HGG^{+}22b] Sara L. Hamilton, Mary G. Gleason, Natalio Godoy, Norah Eddy, and Kirsten Grorud-Colvert. Ecosystem-based management for kelp forest ecosystems. Marine Policy, 136(??): ??, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21005303.

HortaeCosta:2022:OCN

[HGG22c] Barbara Horta e Costa, Jorge M. S. Gonçalves, and Emanuel J. Gonçalves. UN Ocean Conference needs transparent and science-based leadership on ocean conservation. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002445.

Hill:2020:ASA

[HGH^{+}20] Nicholas J. Hill, Caleb Gardner, Malcolm Haddon, Klaas Hartmann, L. Richard Little, Jeremy M. Lyle, and Bradley R. Moore. Assessing the status of Australia’s fish stocks relative to target objectives. Marine Policy, 112(??):

??, February 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19301952.

Humphries:2020:TAM

[HGL^{+}20] Fran Humphries, Hiroko Muraki Gottlieb, Sarah Laird, Rachel Wynberg, Charles Lawson, Michelle Rourke, Morten Walløe Tvedt, Maria Julia Oliva, and Marcel Jaspars. A tiered approach to the marine genetic resource governance framework under the proposed UNCLOS agreement for biodiversity beyond national jurisdiction (BBNJ). Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19308218.

Hendrix:2022:GCN

[HGLR22] Cullen S. Hendrix, Sarah M. Glaser, Joshua E. Lambert, and Paige M. Roberts. Global climate, El Niño, and militarized fisheries disputes in the East and South China Seas. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001841.

Hamel:2024:PSE

[HGQJD24] Ken Hamel, Carlos Garcia-Quijano, Di Jin, and Tracey Dalton. Perceived Sargassum event incidence, impacts, and management response in the Caribbean Basin. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002124.

Haas:2023:FFU

[HGSD23] Bianca Haas, Camille Goodman, Hussain Sinan, and Ruth Davis. Fact or fiction? Unpacking the terminologies used in fisheries allocation discussions. Marine Policy, 152(??): ??, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001574.

Hebbar:2021:BRM

[HH21] Anish Arvind Hebbar and Shantanu Anish Hebbar. Book review: The Maritime Silk Road: China’s Belt and Road at Sea, Richard T. Griffiths. International Institute for Asian Studies (2020), pp232, ISBN: 978-90-823810-4-7(e). Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001159.

Hunt:2025:SBF

[HH25] Alister Hunt and Ray Hilborn. Seychelles’ blue finance: a blueprint for marine conservation? Marine Policy, 179(??): ??, September 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25001320.

Hasan:2021:ICF

[HHB^{+}21] Neaz A. Hasan, Richard D. Heal, Abul Bashar, Alif Layla Bablee, and Mohammad Mahfujul Haque. Impacts of COVID-19 on the finfish aquaculture industry of Bangladesh: a case study. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001883.

Healy:2020:GRE

[HHBC20] Teleah Joy Healy, Nicholas James Hill, Adam Barnett, and Andrew Chin. A global review of elasmobranch tourism activities, management and risk. Marine Policy, 118(??): ??, August 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19302143.

Humphries:2020:PPS

[HHD20] Fran Humphries and Harriet Harden-Davies. Practical policy solutions for the final stage of BBNJ treaty negotiations. Marine Policy, 122(??):??, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308605.

Hooper:2020:PPT

[HHEJB20] Tara Hooper, Caroline Hattam, Andrew Edwards-Jones, and Nicola Beaumont. Public perceptions of tidal energy: Can you predict social acceptability across coastal communities in England? Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20303791.

Hill:2023:PEF

[HHH^{+}23] Nicholas J. Hill, Malcolm Haddon, Klaas Hartmann, L. Richard Little, Jeremy M. Lyle, Bradley R. Moore, and Simon Nicol. Prevalence of effective fisheries management in a developed nation and its link to stock sustainability. Marine Policy, 151(??):??, May 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001239.

Huang:2024:MLM

[HHL^{+}24] Yong Huang, Xiangfeng He, Zhiguang Lian, Zhirong Yang, and Qingbo Jiang. Mapping the landscape of marine cultural heritage research from 2000 to 2023: a bibliometric analysis. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000848.

Haapasaari:2021:IGM

[HIP^{+}21] Päivi Haapasaari, Suvi Ignatius, Mia Pihlajamäki, Andreas Bryhn, Simo Sarkki, Jouni Tuomisto, Lauri Nevalainen, Annukka Lehikoinen, Timo Assmuth, Atso Romakkaniemi, Heikki Peltonen, and Sakari Kuikka. Integrated governance for managing multidimensional problems: Potentials, challenges, and arrangements. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309222.

Hammond:2021:PBH

[HJ21] Amy Hammond and Peter J. S. Jones. Protecting the ‘blue heart of the planet’: Strengthening the governance framework for marine protected areas beyond national jurisdiction. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309064.

Herronen:2025:FIE

[HJ25] Vilma Herronen and Jamie Jenkins. Feasibility of implementing ecological compensation for offshore wind power in the Finnish Baltic Sea area. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25002647.

Hagspiel:2024:CLB

[HJLN24] Verena Hagspiel, Markus Bjørkli Jansen, Maria Lavrutich, and Gaute Nepstad. Company level biodiversity impact assessment: an application to the aquaculture industry in Norway. Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004323.

Han:2023:PHS

[HJPL23] Zenglin Han, Kaiping Jiang, Fei Peng, and Shuqin Li. The Philippines’ hedging strategy against China in the South China Sea dispute: Based on the human-ocean regional system. Marine Policy, 151(??):??, May 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001057.

Heidkamp:2022:GCC

[HKKY22] C. Patrick Heidkamp, Louie Vincent Krak, Miriah M. Russo Kelly, and Charles Yarish. Geographical considerations for capturing value in the U.S. sugar kelp (Saccharina latissima) industry. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002688.

Heinrich:2020:QFC

[HKMS20] Luise Heinrich, Andrea Koschinsky, Till Markus, and Pradeep Singh. Quantifying the fuel consumption, greenhouse gas emissions and air pollution of a potential commercial manganese nodule mining operation. Marine Policy, 114:??, April 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19300818.

Hidayat:2025:BEE

[HKZ25] Nur Ismu Hidayat, Fery Kurniawan, and Zairion Zairion. Balancing economy and ecology: Ecosystem approach to enhanced traditional demersal fisheries management in West Papua, Indonesia. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25002593.

Harris:2021:UCS

[HLCM21] Lucas Harris, Max Liboiron, Louis Charron, and Charles Mather. Using citizen science to evaluate extended producer responsibility policy to reduce marine plastic debris shows no reduction in pollution levels. Marine Policy, 123(??): ??, January 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309660.

Hermann:2022:ATH

[HLLK22] Roberto Rivas Hermann, Ning Lin, Julien Lebel, and Alina Kovalenko. Arctic transshipment hub planning along the Northern Sea Route: a systematic literature review and policy implications of Arctic port infrastructure. Marine Policy, 145(??):??, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003220.

Hu:2020:CMP

[HLM^{+}20] Wenjia Hu, Jie Liu, Zhiyuan Ma, Yuyu Wang, Dian Zhang, Weiwei Yu, and Bin Chen. China’s marine protected area system: Evolution, challenges, and new prospects. Marine Policy, 115(??):??, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303987.

Hume:2021:TOB

[HLO^{+}21] Andrew Hume, Jim Leape, Kirsten L. L. Oleson, Emily Polk, Kevin Chand, and Robert Dunbar. Towards an ocean-based large ocean states country classification. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003778.

Halim:2020:TTM

[HLW^{+}20] Abdul Halim, Neil R. Loneragan, Budy Wiryawan, Rod Fujita, Dedi S. Adhuri, Adrian R. Hordyk, and M. Fedi A. Sondita. Transforming traditional management into contemporary territorial-based fisheries management rights for small-scale fisheries in Indonesia. Marine Policy, 116(??): ??, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19309297.

Hao:2023:UHC

[HLW^{+}23] Huang Hao, Cai Lei, Lei Wang, Weiwen Li, Ou Danyun, Wu Lijun, Ou Rimei, and Lin Lianghua. Using habitat classification and large-scale spatial planning model to identify environmental areas: a case study in the area of the Northwest Pacific Ocean. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004079.

Huang:2021:CLS

[HM21] Yi-Chen Huang and Leslie Mabon. Coastal landscapes, sustainable consumption and peripheral communities: Evaluating the Miramar Resort controversy in Shanyuan Bay, Taiwan. Marine Policy, 123(??):??, January 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309295.

Hassanali:2022:EPG

[HM22a] Kahlil Hassanali and Robin Mahon. Encouraging proactive governance of marine biological diversity of areas beyond national jurisdiction through Strategic Environmental Assessment (SEA). Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005431.

Holmes:2022:HSA

[HM22b] Glen Holmes and Shana Miller. Harvest strategies and allocation — a chicken and egg scenario? Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004826.

Hosch:2022:KNT

[HM22c] G. Hosch and G. Macfadyen. Killing Nemo: Three world regions fail to mainstream combatting of IUU fishing. Marine Policy, 140(??):??, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001208.

Haas:2020:FIP

[HMFH20] Bianca Haas, Jeffrey McGee, Aysha Fleming, and Marcus Haward. Factors influencing the performance of regional fisheries management organizations. Marine Policy, 113 (??):??, March 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19303653.

Harvey:2021:BWA

[HMG^{+}21a] Euan S. Harvey, Dianne L. McLean, Jordan S. Goetze, Benjamin J. Saunders, Tim J. Langlois, Jacquomo Monk, Neville Barrett, Shaun K. Wilson, Thomas H. Holmes, Daniel Ierodiaconou, Alan R. Jordan, Mark G. Meekan, Hamish A. Malcolm, Michelle R. Heupel, David Harasti, Charlie Huveneers, Nathan A. Knott, David V. Fairclough, Leanne M. Currey-Randall, Michael J. Travers, Ben T. Radford, Matthew J. Rees, Conrad W. Speed, Corey B. Wakefield, Mike Cappo, and Stephen J. Newman. The BRUVs workshop — an Australia-wide synthesis of baited remote underwater video data to answer broad-scale ecological questions about fish, sharks and rays. Marine Policy, 127 (??):??, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000403.

Hermoso:2021:EDE

[HMG^{+}21b] María Isabel Hermoso, Victoria Y. Martin, Stefan Gelcich, Wolfgang Stotz, and Martin Thiel. Exploring diversity and engagement of divers in citizen science: Insights for marine management and conservation. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309635.

Hariry:2024:IMH

[HMM^{+}24] Nik Nurhalida Nik Hariry, Fahirah Syaliza Mokhtar, Nor Aieni Mokhtar, Rudiah Md Hanafiah, Suzanna Razali Chan, and Izyan Munirah Mohd Zaideen. Intangible maritime heritage protection in Malaysia: the need for a revision of the National Heritage Act of 2005. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003129.

Hosch:2023:ISH

[HMY^{+}23] Gilles Hosch, Nathan A. Miller, Yann Yvergniaux, Elaine Young, and Tim Huntington. IUU safe havens or PSMA ports: a global assessment of port state performance and risk. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002841.

Haas:2025:CWY

[HNA25] Bianca Haas, Rachel Nichols, and Kamal Azmi. Be careful what you wish (fish) for: Actioning distributive equity principles when determining catch limits in transboundary fisheries. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001678.

Hikuroa:2025:ANZ

[HNB^{+}25] Dan Hikuroa, James Nikitine, Harry Burkhardt, Kura Moeahu, Katherine Short, Mereana Berger, Jeanine Clarkin, Schannel van Dijken, and Kate Brown. Aotearoa New Zealand: a laboratory for future marine conservation. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001265.

Hogg:2021:LTN

[HNMSG21] Katie Hogg, Pedro Noguera-Méndez, and María Semitiel-García. Lessons from three north-western Mediterranean MPAs: a governance analysis of Port-Cros National Park, Tavolara Punta-Coda Cavallo and Ustica. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X17307388.

Haugen:2025:PCS

[HO25] Juliana Figueira Haugen and Jon Olaf Olaussen. Plague or cholera? Stakeholder perspectives on Norwegian salmon farming regulations. Marine Policy, 177(??):??, July 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001009.

Honniball:2020:PSC

[Hon20] Arron N. Honniball. Port states, coastal states and national security: a law of the sea perspective on the 2017 Qatar-Gulf crisis. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304889.

Hong:2021:CAP

[Hon21] Nong Hong. China and the Antarctic: Presence, policy, perception, and public diplomacy. Marine Policy, 134(??): ??, December 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003900.

Hornborg:2023:FHC

[Hor23] Sara Hornborg. Follow the herring — a case study on the interplay between management and markets for marine resource utilization. Marine Policy, 158(??):??, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004074.

Horiguchi:2025:DPA

[Hor25] Takeo Horiguchi. The development of the precautionary approach (PA) in international law and its implications for the management of fish stocks in Japan’s EEZ. Marine Policy, 179(??):??, September 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001587.

Hossain:2024:PAR

[Hos24] Md. Shakhawat Hossain. People’s attitudes regarding plastics and microplastics pollution: Perceptions, behaviors, and policy implications. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24002173.

Howson:2020:BTE

[How20] Peter Howson. Building trust and equity in marine conservation and fisheries supply chain management with blockchain. Marine Policy, 115(??):??, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307067.

Hsuan:2020:MSC

[HP20] Juliana Hsuan and Cristiana Parisi. Mapping the supply chain of ship recycling. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305858.

Hsu:2023:UVS

[HP23] Kang Hsu and Li-Pei Peng. Understanding vulnerability and sustainable livelihood factors from coastal residents in Taiwan. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003263.

Herrera:2023:PPM

[HPC^{+}23] Mariana Herrera, Pablo Pita, Daniela Castelo, C. Marisa R. Almeida, Sandra Ramos, and Sebastián Villasante. Public perceptions of marine litter and impacts on coastal ecosystem services in Galicia (Spain). Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002750.

Haas:2020:MRS

[HPG20] Bianca Haas, Michelle Phillipov, and Fred Gale. Media representations of seafood certification in Australia: Mobilising sustainability standards to attack or defend the value of an industry. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20303341.

He:2022:DDT

[HPH22] Xinhua He, Qiuyun Ping, and Wenfa Hu. Does digital technology promote the sustainable development of the marine equipment manufacturing industry in China? Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004796.

Heck:2022:VSE

[HPPH22] Nadine Heck, Adina Paytan, Donald Potts, and Brent Haddad. Variations in stakeholders’ ecosystem service priorities for managing a marine protected area. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003773.

Ha:2023:UCD

[HPS23] Min-Ho Ha, Ho Park, and Young-Joon Seo. Understanding core determinants in LNG bunkering port selection: Policy implications for the maritime industry. Marine Policy, 152 (??):??, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001355.

Hopkins:2024:ITS

[HRC^{+}24] C. R. Hopkins, S. I. Roberts, A. J. Caveen, C. Graham, and N. M. Burns. Improved traceability in seafood supply chains is achievable by minimising vulnerable nodes in processing and distribution networks. Marine Policy, 159(??): ??, January 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004438.

Herrera-Racionero:2022:RSM

[HRMPCRRA22] Paloma Herrera-Racionero, Luís Miret-Pastor, Roberto Cervelló-Royo, and Miguel Rodilla-Alama. The role of the Spanish Mediterranean Fisher’s Guilds in maritime sustainability. Marine Policy, 140(??):??, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001051.

Haraldsson:2020:HMS

[HRR^{+}20] Matilda Haraldsson, Aurore Raoux, Fabien Riera, Julien Hay, Jeffrey M. Dambacher, and Nathalie Niquil. How to model social-ecological systems? — A case study on the effects of a future offshore wind farm on the local society and ecosystem, and whether social compensation matters. Marine Policy, 119(??):??, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306530.

Hellstrom:2024:GER

[HRS^{+}24] Magnus Hellström, Rodrigo Rabetino, Henry Schwartz, Anastasia Tsvetkova, and Syed Hammad Ul Haq. GHG emission reduction measures and alternative fuels in different shipping segments and time horizons — a Delphi study. Marine Policy, 160(??):??, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005304.

Hai:2020:ISS

[HS20] Au Ton Nu Hai and Stijn Speelman. Involving stakeholders to support sustainable development of the marine lobster aquaculture sector in Vietnam. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305883.

Hastings:2022:MSF

[HS22] Justin V. Hastings and Annie Young Song. The maritime security fallout from North Korea’s fisheries policy. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005261.

Hau:2023:MMM

[HS23] C. Y. Hau and Y. J. Sadovy de Mitcheson. Mortality and management matter: Case study on use and misuse of ‘ranching’ for a CITES Appendix II-listed fish, humphead wrasse (Cheilinus undulatus). Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000428.

Harris:2024:IET

[HS24] Joanna L. Harris and Guy M. W. Stevens. The illegal exploitation of threatened manta and devil rays in the Chagos Archipelago, one of the world’s largest notake MPAs. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001088.

Hickey:2021:IOT

[HSdT21] Gordon M. Hickey, Hunter T. Snyder, Jasper R. deVries, and Owen Temby. On inter-organizational trust, control and risk in transboundary fisheries governance. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003833.

Hogg:2021:GAC

[HSGNMGC21] Katie Hogg, María Semitiel-García, Pedro Noguera-Méndez, and Jose Antonio García-Charton. A governance analysis of Cabo de Palos-Islas Hormigas and Cabo de Gata–Níjar Marine Protected Areas, Spain. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X17307327.

Howard:2023:BCP

[HSGS^{+}23] Jennifer Howard, Ariana E. Sutton-Grier, Lindsey S. Smart, Christian C. Lopes, Jill Hamilton, Joan Kleypas, Stefanie Simpson, Jennifer McGowan, Albert Pessarrodona, Heidi K. Alleway, and Emily Landis. Blue carbon pathways for climate mitigation: Known, emerging and unlikely. Marine Policy, 156(??):??, October 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003214.

Hausner:2021:DSO

[HSH^{+}21] Arjun Hausner, Jameal F. Samhouri, Elliott L. Hazen, Delgerzaya Delgerjargal, and Briana Abrahms. Dynamic strategies offer potential to reduce lethal ship collisions with large whales under changing climate conditions. Marine Policy, 130(??):??, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001767.

Huan:2024:REE

[HSHS24] Xiaomeng Huan, Jingzhu Shan, Limin Han, and Haiming Song. Research on the efficacy and effect assessment of deep-sea aquaculture policies in China: Quantitative analysis of policy texts based on the period 2004– 2022. Marine Policy, 160(??):??, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004967.

Hsiao:2020:OFE

[Hsi20] Amanda Hsiao. Opportunities for fisheries enforcement cooperation in the South China Sea. Marine Policy, 121(??): ??, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19303689.

Hitchin:2023:TDS

[HSK^{+}23] B. Hitchin, S. Smith, K. Kröger, DOB Jones, A. Jaeckel, NC Mestre, J. Ardron, E. Escobar, J. van der Grient, and T. Amaro. Thresholds in deep-seabed mining: a primer for their development. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000325.

Horta:2023:BFF

[HSM^{+}23] Paulo Horta, Marina Sissini, Carolina Melissa Mueller, Fernanda M. M. Soares, Paulo Pagliosa, Leonardo Rörig, José Bonomi-Barufi, Flavio Berchez, Leticia Cotrim da Cunha, Rodrigo Kerr, Sergio Rossi, Marcelo O. Soares, Jorge L. Rodrigues-Filho, Hudson T. Pinheiro, Luciano Henning, Marcos Aurélio Espindola, Nicole Figueiredo de Oliveira, and Alessandra Larissa Fonseca. Brazil fosters fossil fuel exploitation despite climate crises and the environmental vulnerabilities. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004705.

Hossain:2021:SFS

[HSN^{+}21] M. Shahadat Hossain, S. M. Sharifuzzaman, M. Nur Nobi, M. Shah Nawaz Chowdhury, Subrata Sarker, M. Alamgir, Sheikh Aftab Uddin, Sayedur Rahman Chowdhury, M. Mizanur Rahman, M. Shajjadur Rahman, Faisal Sobhan, and Showmitra Chowdhury. Seaweeds farming for sustainable development goals and blue economy in Bangladesh. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000804.

Heinrich:2024:MGC

[HSSM24] Luise Heinrich, Pradeep Singh, Karen Smith Stegen, and Till Markus. Mind the gap and close it: Regulating greenhouse gas emissions from deep-sea mining in the area. Marine Policy, 160(??):??, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004621.

Hirokawa:2023:INS

[HT23] Toya Hirokawa and Benjamin S. Thompson. The influence of new sustainable fisheries policies on seafood company practices and consumer awareness in Japan. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003524.

Hidalgo:2025:VOE

[HTH25] Marcelo Hidalgo, Peter Trott, and Bianca Haas. The vulnerability of observers – an evaluation of observer programs welfare and working conditions policies. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24005402.

Herrera:2023:DIB

[HTP^{+}23] Mariana Herrera, Ana Tubío, Pablo Pita, Elsa Vázquez, Celia Olabarria, Andrés Simón, José Carlos Mariño Balsa, Liliana Solís, Ignacio Gianelli, and Sebastian Villasante. Disentangling interactions between seagrasses and smallscale fisheries using scientific and local traditional knowledge. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002749.

Hatenboer:2023:DFS

[HvdBH23] Constant Hatenboer, Caspar van den Berg, and Ronald Holzhacker. The Dutch fisheries sector and the North Sea Accord: Unpacking stakeholder participation in multilevelled marine governance. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004110.

Herttua:2021:AAO

[HVN21] Kimmo Herttua, Jan Vork, and Jesper Bo Nielsen. Associations of age, occupation and nationality of seafarers with risk of evacuations at sea: a register-based study on Danish-flagged merchant ships. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1930908X.

Hoshino:2020:DQO

[HvPPV20a] Eriko Hoshino, Ingrid van Putten, Sean Pascoe, and Simon Vieira. Does quota ownership affect perceptions of fishery performance? Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20304589.

Hoshino:2020:ITQ

[HvPPV20b] Eriko Hoshino, Ingrid van Putten, Sean Pascoe, and Simon Vieira. Individual transferable quotas in achieving multiple objectives of fisheries management. Marine Policy, 113(??): ??, March 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X1930346X.

Halik:2022:DCR

[HVS22] Abdul Halik, Marco Verweij, and Achim Schlüter. Deliberating coral reef protection — Cultural Theory tested. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000835.

Haapasaari:2022:FRN

[HvT22] Päivi Haapasaari and Jan P. M. van Tatenhove. A Finnish regional non-binding MSP approach: What are the consequences for integrating Blue Growth and GES?

Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001488.

Hayes:2022:CMC

[HW22] Christopher Hayes and Lisa Wainger. Comparing multicriteria decision analysis to group negotiations in fisheries co-management. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000446.

Hartmann:2025:CFP

[HWC^{+}25] Lisa Hartmann, Yvonne Walz, David Clelland, Catalina Esquivel-Rodriguez, Patrizia Riganti, Marta Vicarelli, and Fabrice G. Renaud. A conceptual framework for a post-COVID-19 green recovery of the Blue Economy. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004469.

Haensch:2022:CAT

[HWM22] Juliane Haensch, Sarah Ann Wheeler, and Stephanie McWhinnie. Community attitudes towards marine parks in South Australia. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301457.

Hague:2025:ADU

[HWM^{+}25] Emily Hague, Alice Walters, Anna Moscrop, Emma Steel, Katie Dyke, Lauren Hartny-Mills, Alison Lomax, Rebecca Dudley, Pippa Garrard, Jenny Hampson, Sadie Gorvett, Hannah Lightley, Craig Mackie, Juliane Lehmann, Sebastian Olias, Carsten Hilgenfeld, Debbie Cole, Sarah MacDonald-Taylor, Carole Davis, Bernard Siddle, Peter Gulliver, Julie Tozer, Wendy Kilroe, Áine Purcell Milton, Rebecca Olaleye, Kathryn Allan, Tim Stenton, Emma Neave-Webb, Russell Neave, David Lambie, Anthony J. Morrison, Isabella Van Damme, Phil Dickinson, and Lauren McWhinnie. AIS data underrepresents vessel traffic around coastal Scotland. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001344.

Hori:2020:CAS

[HWMO20] Juri Hori, Hiroki Wakamatsu, Tsutom Miyata, and Yoshioki Oozeki. Has the consumers awareness of sustainable seafood been growing in Japan? Implications for promoting sustainable consumerism at the Tokyo 2020 Olympics and Paralympics. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304816.

Hanzawa:2023:SSA

[HWT23] Yudai Hanzawa, Shingo Watari, and Shion Takemura. Not satisfactory, but still acceptable: Exploring socioeconomic incentives of individual quota systems in Japanese Pacific bluefin tuna fisheries from the management leaders’ perspective. Marine Policy, 150(??):??, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000283.

Hanzawa:2024:SPP

[HWT24] Yudai Hanzawa, Shingo Watari, and Shion Takemura. Stakeholder perceptions and policy implications of voluntary and government-led individual quota systems in Japanese fisheries co-management: a case study of purse seine fisheries in the Sea of Japan. Marine Policy, 165(??): ??, July 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2400160X.

Hakala:2023:GSI

[HWU^{+}23] Siri Hakala, Shingo Watari, Shinji Uehara, Yujiro Akatsuka, Richard Methot, and Yoshi Oozeki. Governance and science implementation in fisheries management in Japan as it compares to the United States. Marine Policy, 155(??): ??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001975.

Hung:2022:CBG

[HWY22] Ling-Ya Hung, Shun-Mei Wang, and Ting-Kuang Yeh. Collaboration between the government and environmental nongovernmental organisations for marine debris policy development: the Taiwan experience. Marine Policy, 135(??): ??, January 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004607.

Halimatussadiah:2023:CML

[HYN^{+}23] Alin Halimatussadiah, Satoshi Yamazaki, Rachel Nichols, Umi Muawanah, Fachry Abdul Razak Afifi, and Muhammad Adriansyah. Co-management of large-scale and medium-scale fisheries: an assessment of the fishery tax system in Indonesia. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200505X.

He:2023:GSI

[HZ23] Yixiong He and Fengxuan Zhang. A game study on the implementation of marine carbon sink fisheries in the context of carbon neutrality — analysis of the tripartite behavior of fishery practitioners, research institutions, and the government. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004122.

Holzheid:2024:DSM

[HZC^{+}24] Astrid Holzheid, Hongbo Zhao, Tony Cabus, Lei Fan, Thomas Kuhn, Linlin Sun, Chunhui Tao, Matthias Haeckel, Duong Hoang, Norman Kelly, Terue Kihara, Bing Li, Jiangtao Li, Ju Ma, Nele Matz-Lück, Klaas Meyn, Massimiliano Molari, Sven Petersen, Katrin Pollmann, Martin Rudolph, Xiangxin Xu, and Yijia Zhang. Deep-sea mining of massive sulfides: Balancing impacts on biodiversity and ecosystem, technological challenges and law of the sea. Marine Policy, 167(??):??, September 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002872.

Iversen:2020:GDF

[IAB^{+}20] Audun Iversen, Frank Asche, Marcus Buck, Edgar Henriksen, Jonas Stein, and Sigbjørn Svalestuen. The growth and decline of fisheries communities: Explaining relative population growth at municipality level. Marine Policy, 112(??): ??, February 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19304555.

Imbwae:2025:ELS

[IAS25] Imikendu Imbwae, Shankar Aswani, and Warwick Sauer. Enhancing the livelihood of small-scale fishers in the context of vulnerability. Marine Policy, 177(??):??, July 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001058.

Inarra:2020:EUU

[IBC^{+}20] Bruno Iñarra, Carlos Bald, Marta Cebrián, Irene Peral, Raquel Llorente, and Jaime Zufía. Evaluation of unavoidable unwanted catches valorisation options: the Bay of Biscay case study. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308455.

Inguanzo:2021:RPF

[IGO21] Belén Inguanzo, María-José Gutiérrez, and Susan Orbe. The role of production factors on landings heterogeneity between EU countries. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002906.

Islam:2023:DDS

[IGS^{+}23] Samsul Islam, Floris Goerlandt, Quazi Mohammed Habibus Sakalayen, Yangyan Shi, and V. G. Venkatesh. Developing a “Disaster Scenario” to prepare for the possibility of disruptions to maritime transportation serving coastal communities of Vancouver Island. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000581.

Islam:2021:INC

[IKB21] Md Monirul Islam, Makidul Islam Khan, and Aparna Barman. Impact of novel coronavirus pandemic on aquaculture and fisheries in developing countries and sustainable recovery plans: Case of Bangladesh. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002220.

Iue:2022:DBS

[IMA22] Minako Iue, Mitsutaku Makino, and Misuzu Asari. The development of “Blue Seafood Guide,” a sustainable seafood rating program, and its implication in Japan. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100556X.

Isaac:2024:SCI

[IMB^{+}24] Janella Isaac, Rowena Mouda, David Barrett, Jackie Gould, Dermot Smyth, and Tanya Vernes. Sea Country Indigenous Protected Areas: Indigenous leadership in the protection of Australia’s marine environments. Marine Policy, 170(??): ??, December 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003567.

Islam:2020:FCM

[INW20] Mohammad Mahmudul Islam, Md. Nahiduzzaman, and Md. Abdul Wahab. Fisheries co-management in hilsa shad sanctuaries of Bangladesh: Early experiences and implementation challenges. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303380.

Ibengwe:2023:CIM

[IO23] Lilian J. Ibengwe and Paul O. Onyango. Capturing the invisible: Methodological approach in understanding informal cross border fish trade. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2300009X.

Ibengwe:2022:RTI

[IOHC22] Lilian J. Ibengwe, Paul O. Onyango, Aloyce S. Hepelwa, and Martin J. Chegere. Regional trade integration and its relation to income and inequalities among Tanzanian marine dagaa fishers, processors and traders. Marine Policy, 137(??):??, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000227.

Islam:2021:TAC

[IRK^{+}21] Md. Monirul Islam, Md. Asadur Rahman, Mohammad Shahneawz Khan, Gouri Mondal, and Makidul Islam Khan. Transformational adaptations to climatic hazards: Insights from mangroves-based coastal fisheries dependent communities of Bangladesh. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000865.

Islam:2025:CAP

[Isl25] Md. Ziaul Islam. A contemporary analysis of policy implementation for the coastal mangrove forest management in Bangladesh Sundarbans. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24004603.

Israngkura:2022:MRR

[Isr22] Adis Israngkura. Marine resource recovery in southern Thailand during COVID-19 and policy recommendations. Marine Policy, 137(??):??, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000197.

Jones:2020:ECI

[JACD20] Daniel O. B. Jones, Jeff A. Ardron, Ana Colaço, and Jennifer M. Durden. Environmental considerations for impact and preservation reference zones for deep-sea polymetallic nodule mining. Marine Policy, 118(??):??, August 2020.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301374.

Jimenez:2020:VCD

[JAdS^{+}20] Érica Antunes Jimenez, Marilu Teixeira Amaral, Pauliana Leão de Souza, Maria de Nazaré Ferreira Costa, Alex Souza Lira, and Flávia Lucena Frédou. Value chain dynamics and the socioeconomic drivers of small-scale fisheries on the Amazon coast: a case study in the state of Amapá, Brazil. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18301428.

Jaeckel:2020:SEP

[Jae20] Aline Jaeckel. Strategic environmental planning for deep seabed mining in the area. Marine Policy, 114:??, April 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1830722X.

Jayasinghe:2024:OPG

[JAH^{+}24] Rohantha Rukshan Jayasinghe, Wasudha Prabodhani Abeyrathna, Manuja Promodya Hendawitharana, Champika Liyanage, Karl Williams, and Rangika Umesh Halwatura. Overcoming policy gaps for effective plastic waste management: a case study of Dikkowita Fishery Harbour, Sri Lanka. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000265.

Jones:2023:ERI

[JAK^{+}23] Stuart T. Jones, Edward H. Allison, Kailin Kroetz, Yoshitaka Ota, and Sunny L. Jardine. Enrollment, retention, and inclusivity of Marine Stewardship Council (MSC) ecolabelling certifications. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002671.

Jenkins:2025:SVS

[JB25] Lekelia Danielle Jenkins and Ezra Beaver. Stakeholders’ views on siting in-stream tidal energy projects in urban and remote communities in the United States. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001174.

Jalali:2025:RJU

[JBGC25] Ali Jalali, Justin D. Bell, Harry K. Gorfine, and Simon D. Conron. Reeling in joy: Untangling the link between satisfaction and motivation in recreational fishing. Marine Policy, 179(??):??, September 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001575.

Jardim:2023:MFS

[JC23] Ernesto Jardim and Rohan J. C. Currey. The MSC Fisheries Standard Review policy development process. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003883.

Jiang:2020:GMS

[JCMD20] Daokui Jiang, Zhuo Chen, Levi McNeil, and Guilin Dai. The game mechanism of stakeholders in comprehensive marine environmental governance. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X18309485.

Jo:2020:DTS

[JD20] Sohyun Jo and Enrico D’agostini. Disrupting technologies in the shipping industry: How will MASS development affect the maritime workforce in Korea. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20304668.

Jacob:2023:TWP

[JDF^{+}23] Céline Jacob, Sereno DuPrey Diederichsen, Liam Fullbrook, Amanda T. Lombard, Siân E. Rees, Nina Rivers, Bernadette Snow, Mia Strand, Rachel Zuercher, and Holly J. Niner. A two way process — social capacity as a driver and outcome of equitable marine spatial planning. Marine Policy, 149(??):??, March 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000349.

Jensen:2021:FDC

[Jen21] Rikke Bjerg Jensen. Fragmented digital connectivity and security at sea. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309350.

Jensen:2023:ADI

[Jen23] Frank Jensen. Advantages and disadvantages of introducing strong user rights in fisheries. A literature review. Marine Policy, 157(??):??, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300372X.

Jiang:2020:RSC

[JF20] Minzhen Jiang and Michael Faure. Risk-sharing in the context of fishery mutual insurance: Learning from China. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030405X.

Jiang:2022:CSM

[JF22] Minzhen Jiang and Michael Faure. The compensation system for marine ecological damage resulting from offshore drilling in China. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001798.

Jin:2024:GAP

[JF24] Yongming Jin and Yuan Feng. Asia–Pacific maritime security cooperation: Challenges and opportunities. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2300475X.

Jacquet:2024:OAI

[JFB^{+}24] Jennifer Jacquet, Becca Franks, Heather Browning, Alfonso Donoso, Matthew Hayek, Kathy Hessler, and Laurie Sellars. Away from an offshore aquaculture industry in the United States. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005195.

Jamieson:2020:DAI

[JG20] J. W. Jamieson and A. Gartman. Defining active, inactive, and extinct seafloor massive sulfide deposits. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19307845.

Jin:2023:PTD

[JG23] Yan Jin and José María Gil. Price transmission during food safety incidents: the case of Spanish tuna fraud. Marine Policy, 156(??):??, October 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003275.

James:2023:IME

[JGWM^{+}23] M. A. James, R. Gozzer-Wuest, T. Mendo, I. Gomez, J. Grillo-Núñez, and J. Mendo. To ignore or mitigate — economic implications of an illegal artisanal trawl fishery in northern Peru. Marine Policy, 158(??):??, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003986.

Jiang:2024:EAC

[JH24] Xiaoyi Jiang and Zipeng Huang. Extraterritorial application of China’s Marine Environmental Protection Law: Towards effective marine environmental protection enforcement. Marine Policy, 167(??):??, September 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002392.

Jeffery:2023:OSR

[JHHG23] Sharon Jeffery, Lucie C. Hannah, Leif-Matthias Herborg, and Candice St Germain. Oil spill response planning in Pacific Canada: a tool for identifying vulnerable marine biota. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005176.

Jiang:2021:PCS

[JHL21] Miaomiao Jiang, Maixiu Hu, and Markus Leibrecht. Profitability of container shipping via the Arctic Northeast Passage: a simulation and regression analysis. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003493.

Juan:2022:FKS

[JK22] Wang Juan and Wang Kexin. Fishery knowledge spillover effects on tourism economic growth in China — spatiotemporal effects and regional heterogeneity. Marine Policy, 139 (??):??, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000665.

Jeanson:2025:CMI

[JKCY25] Amanda L. Jeanson, Andrew N. Kadykalo, Steven J. Cooke, and Nathan Young. Caught in the middle — inaction and overlap in governance and decision-making for Canada’s imperiled wild steelhead. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24005414.

Jueseah:2020:SFE

[JKKT20] Alvin Slewion Jueseah, Ogmundur Knutsson, Dadi Mar Kristofersson, and Tumi Tómasson. Seasonal flows of economic benefits in small-scale fisheries in Liberia: a value chain analysis. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305330.

Jayanthi:2022:ILW

[JKV^{+}22] M. Jayanthi, M. Kumaran, S. Vijayakumar, M. Duraisamy, P. R. Anand, M. Samynathan, S. Thirumurthy, S. Kabiraj, K. P. K. Vasagam, A. Panigrahi, and M. Muralidhar. Integration of land and water resources, environmental characteristics, and aquaculture policy regulations into site selection using GIS based spatial decision support system. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005571.

Ji:2021:DCF

[JL21a] Jianyue Ji and Yanming Li. The development of China’s fishery informatization and its impact on fishery economic efficiency. Marine Policy, 133(??):??, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003225.

Jiang:2021:DPP

[JL21b] Shan-Shan Jiang and Jing-Mei Li. Do political promotion incentive and fiscal incentive of local governments matter for the marine environmental pollution? Evidence from China’s coastal areas. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001160.

Jones:2021:ADR

[JL21c] Peter J. S. Jones and Stephen D. Long. Analysis and discussion of 28 recent marine protected area governance (MPAG) case studies: Challenges of decentralisation in the shadow of hierarchy. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310137.

Judd:2021:AST

[JL21d] Adrian Judd and Jemma-Anne Lonsdale. Applying systems thinking: the Ecosystem Approach and Natural Capital Approach — Convergent or divergent concepts in marine management? Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001287.

Jing:2024:SER

[JL24] Youfu Jing and Yang Liu. Spatiotemporally explicit risk assessment of fishing grounds in Asian waters. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003002.

Ji:2023:ERD

[JLY23] Jianyue Ji, Huimin Liu, and Xingmin Yin. Evaluation and regional differences analysis of the marine industry development level: the case of China. Marine Policy, 148(??): ??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004924.

Johnson:2020:MSP

[JMB^{+}20] Ayana Elizabeth Johnson, William James McClintock, Ogden Burton, Wayde Burton, Andrew Estep, Kathryn Mengerink, Read Porter, and Stephanie Tate. Marine spatial planning in Barbuda: a social, ecological, geographic, and legal case study. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309928.

Johannesen:2025:AGE

[Joh25] Ellen Johannesen. Advancing gender equality in international ocean science: Participatory approaches for institutional actions. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004214.

Jones:2021:GAN

[Jon21] Peter J. S. Jones. A governance analysis of Ningaloo and Shark Bay Marine Parks, Western Australia: Putting the ‘eco’ in tourism to build resilience but threatened in longterm by climate change? Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19303458.

Jompa:2023:TMP

[JPM^{+}23] Jamaluddin Jompa, Amanda Pricella Putri, Abigail Mary Moore, Hartati Tamti, and Siti Haerani. The transference of marine protected area management authority in Indonesia: Problems encountered, consequences and ways to move forward. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002890.

Jaleel:2023:PFM

[JS23] Azmath Jaleel and Hance D. Smith. The plight of the fishers and the management of IUU fishing. Marine Policy, 150 (??):??, April 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000842.

Jaya:2022:WPF

[JSW^{+}22a] Indra Jaya, Fayakun Satria, Wudianto, Duto Nugroho, Lilis Sadiyah, Eny A. Buchary, Alan T. White, Erik C. Franklin, Catherine A. Courtney, Gina Green, and Stuart J. Green. “Are the working principles of fisheries management at work in Indonesia?”. Marine Policy, 140(??):??, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200094X.

Jubinville:2022:PPA

[JSW22b] Isabelle Jubinville, Nancy L. Shackell, and Boris Worm. From policy to practice: Addressing bycatch for marine species-at-risk in Canada. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003475.

Jimenez-Toribio:2025:APV

[JTdH25] Ramón Jiménez-Toribio and Juan José García del Hoyo. Asymmetric price volatility transmission in the fresh sardine value chain: the case of Spain. Marine Policy, 175 (??):??, May 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25000429.

Jara:2020:CFS

[JTR^{+}20] Hans J. Jara, Jorge Tam, Borja G. Reguero, Francisco Ganoza, Gladis Castillo, Carlos Y. Romero, Manon Gévaudan, and Américo A. Sánchez. Current and future socio-ecological vulnerability and adaptation of artisanal fisheries communities in Peru, the case of the Huaura province. Marine Policy, 119(??):??, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305597.

Johansen:2020:CSO

[JV20] Despina F. Johansen and Rolf A. Vestvik. The cost of saving our ocean — estimating the funding gap of sustainable development goal 14. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305111.

Jurrius:2024:THP

[JvH24] Lobke H. Jurrius and Luc van Hoof. Towards holistic, participative and adaptable governance for offshore wind farm decommissioning. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24004111.

Jiang:2020:EGM

[JWX^{+}20] Bao Jiang, Xiaoqiong Wang, Hailiang Xue, Jian Li, and Yu Gong. An evolutionary game model analysis on emission control areas in China. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20301408.

Juteau:2025:BUC

[JXWL25] Christy Juteau, Harley Chappell Xwopokton, Sarah Marie Wiebe, and Robert Lapper. The beauty underneath: a critical coastal governance approach to revitalize indigenous shellfish harvest. Marine Policy, 174(??):??, April 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000053.

Jang:2020:CLA

[JY20] Ho Geun Jang and Satoshi Yamazaki. Community-level analysis of correlated fish production in fisheries and aquaculture: the case of Japan. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20308861.

Jiang:2021:SIF

[JZW21] Ke Jiang, Xinxin Zhang, and Yusheng Wang. Stability and influencing factors when designing incentive-compatible payments for watershed services: Insights from the Xin’an River Basin, China. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004358.

Kadfak:2021:SNM

[KA21] Alin Kadfak and Anna Antonova. Sustainable Networks: Modes of governance in the EU’s external fisheries policy relations under the IUU Regulation in Thailand and the SFPA with Senegal. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002670.

Kamal:2024:DDB

[KA24] Bunyamin Kamal and Abdullah Altunışık. A data-driven Bayes approach for investigating International Safety Management Code-sourced detention of ships in Port State Controls. Marine Policy, 169(??):??, November 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003440.

Kaikkonen:2024:IIA

[KAA^{+}24] Laura Kaikkonen, Teresa Amaro, Peter J. Auster, David M. Bailey, James B. Bell, Angelika Brandt, Malcolm R. Clark, Jeffrey C. Drazen, Cherisse Du Preez, Elva Escobar-Briones, Eva Giacomello, Matthew Gianni, Andrew F. Johnson, Lisa A. Levin, Rosanna J. Milligan, Stephen Oduware, Tabitha R. R. Pearman, Christopher K. Pham, Sofia P. Ramalho, Ashley A. Rowden, Tracey T. Sutton, Michelle L. Taylor, Les Watling, and Lissette Victorero. Improving impact assessments to reduce impacts of deep-sea fisheries on vulnerable marine ecosystems. Marine Policy, 167(??):??, September 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002793.

Kurt:2021:OCA

[KABT21] Ismail Kurt, Murat Aymelek, Evangelos Boulougouris, and Osman Turan. Operational cost analysis for a container shipping network integrated with offshore container port system: a case study on the West Coast of North America. Marine Policy, 126(??):??, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000105.

King:2022:MRR

[KAC22] Chloe King, Dedi S. Adhuri, and Julian Clifton. Marine reserves and resilience in the era of COVID-19. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001403.

Kai:2021:CIC

[Kai21] Mikihiko Kai. Are the current IUCN category and CITES listing appropriate for the conservation and management of shortfin mako, Isurus oxyrinchus, in the North Pacific Ocean? Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004012.

Kakee:2020:DSM

[Kak20] Tomoko Kakee. Deep-sea mining legislation in Pacific Island countries: From the perspective of public participation in approval procedures. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305718.

Kuemlangan:2023:EAA

[KAN^{+}23] Blaise Kuemlangan, Elizabeth-Rose Amidjogbe, Julia Nakamura, Alessandra Tomassi, Rudolph Hupperts, Buba Bojang, and Teresa Amador. Enforcement approaches against illegal fishing in national fisheries legislation. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000416.

Kao:2023:ATF

[Kao23] Shih-Ming Kao. Analysis of Taiwan’s foreign fisheries policy in tuna RFMOs: Possible variables and determinants. Marine Policy, 151(??):??, May 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001252.

Kara:2022:DMS

[Kar22a] Esma Gül Emecen Kara. Determination of maritime safety performance of flag states based on the Port State Control inspections using TOPSIS. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002032.

Karim:2022:MCI

[Kar22b] Md Saiful Karim. Maritime cybersecurity and the IMO legal instruments: Sluggish response to an escalating threat? Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001853.

Karnad:2022:ILE

[Kar22c] Divya Karnad. Incorporating local ecological knowledge aids participatory mapping for marine conservation and customary fishing management. Marine Policy, 135(??): ??, January 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004528.

Karim:2023:IOT

[Kar23] Md Saiful Karim. Indian Ocean Tuna Commission Climate Change Resolution: a quiet interaction of ocean and climate change legal regimes. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2200481X.

Kassouri:2021:EDF

[Kas21] Yacouba Kassouri. Exploring the dynamics of fishing footprints in the Gulf of Guinea and Congo Basin region: Current status and future perspectives. Marine Policy, 133(??): ??, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100350X.

Katikiro:2020:RLI

[Kat20] Robert Eliakim Katikiro. The role of local institutions in governance of MPAs. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303823.

Kawamoto:2022:CEG

[Kaw22] Taro Kawamoto. A challenge to estimate global canned tuna demand and its impact on future tuna resource management using the gamma model. Marine Policy, 139(??):

??, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2200063X.

Kayashima:2025:NST

[Kay25] Tomihiko Kayashima. A new step toward sustainable fisheries in Japan: the 2018 Amendments to the 1949 Fisheries Act and the Fisheries Cooperative Associations Act. Marine Policy, 178(??):??, August 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001253.

Kawamoto:2020:CFP

[KB20] Taro Kawamoto and Osamu Baba. Comparison of financial performance of Japanese and Australian small scale tuna longline fisheries. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930733X.

Khan:2025:HAN

[KBB^{+}25] Abdul Rehman Khan, Mohd Shahrizal Bin Ab Razak, Badronnisa Binti Yusuf, Helmi Zulhaidi Bin Mohd Shafri, and Noorasiah Binti Mohamad. Harnessing artificial neural networks for coastal erosion prediction: a systematic review. Marine Policy, 178(??):??, August 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001198.

Khan:2025:TLM

[KBC25] Muneeb Khan, Aiman Bibi, and Yen-Chiang Chang. Towards legal modernisation: Pakistan’s maritime legal regime vis-à-vis the Maritime Labour Convention 2006. Marine Policy, 179(??):??, September 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001605.

Krause:2020:VSA

[KBD^{+}20] Gesche Krause, Suzannah-Lynn Billing, John Dennis, Jon Grant, Lucia Fanning, Ramón Filgueira, Molly Miller, José Antonio Pérez Agúndez, Nardine Stybel, Selina M. Stead, and Wojciech Wawrzynski. Visualizing the social in aquaculture: How social dimension components illustrate the effects of aquaculture across geographic scales. Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307481.

Kauer:2024:AFS

[KBH^{+}24] Kate Kauer, Lyall Bellquist, Jenn Humberstone, Vienna Saccomanno, Dwayne Oberhoff, Sherry Flumerfelt, and Mary Gleason. Advancing fisheries sustainability and access through community fisheries trusts. Marine Policy, 165 (??):??, July 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24002082.

Kruk:2024:FLA

[KBP24] Sake R. L. Kruk, Simon R. Bush, and Michael Phillips. Federating ‘Aquaculture 4.0’ for data-driven social and environmental sustainability. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24003531.

Khan:2024:NPM

[KCB24] Muneeb Khan, Yen-Chiang Chang, and Aiman Bibi. Navigating Pakistan’s maritime industry potential in context of blue economy: an analysis of the necessity for ratification of maritime labour convention 2006. Marine Policy, 165 (??):??, July 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24001489.

Kambey:2021:EBP

[KCCC^{+}21] Cicilia S. B. Kambey, Iona Campbell, Elizabeth J. Cottier-Cook, Adibi R. M. Nor, Azhar Kassim, Ahemad Sade, and Phaik-Eem Lim. Evaluating biosecurity policy implementation in the seaweed aquaculture industry of Malaysia, using the quantitative knowledge, attitude, and practices (KAP) survey technique. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004115.

Karcher:2025:BBY

[KCH^{+}25] Denis B. Karcher, Christopher Cvitanovic, Alistair J. Hobday, Rebecca Shellock, Robert L. Stephenson, and Ingrid van Putten. Best bang for your buck: Considerations for cost-efficiency in knowledge co-production. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001848.

Kam:2024:TTK

[KCT^{+}24] Yan Chit Kam, Arthur Chung, Max Tin, Juan Diego Gaitan-Espitia, and Celia Schunter. Temporal trends of key commercial species under live reef food fish trade in Hong Kong. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001982.

Kim:2025:IEP

[KdBK25] Yulan Kim and Rebeca de Buen Kalman. Incorporating equity in policy analysis: the case of the Nova Scotian lobster fishery and COVID-19. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25001630.

Kochanowicz:2021:UWS

[KDH^{+}21] Zuzanna Kochanowicz, Jackie Dawson, William D. Halliday, Michael Sawada, Luke Copland, Natalie Ann Carter, Adrian Nicoll, Steven H. Ferguson, Mads Peter Heide-Jørgensen, Marianne Marcoux, Cortney Watt, and David J. Yurkowski. Using western science and Inuit knowledge to model ship-source noise exposure for cetaceans (marine mammals) in Tallurutiup Imanga (Lancaster Sound), Nunavut, Canada. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001688.

Khan:2024:NWC

[KDHL24] Mehran Idris Khan, BingYing Dong, Qihao He, and Moxin Liu. Navigating the waves of change — assessing the evolving landscape of China’s revised marine environmental protection law. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003695.

Kline:2020:SSU

[KDM^{+}20] Logan R. Kline, Annamaria I. DeAngelis, Candace McBride, Giverny G. Rodgers, Timothy J. Rowell, Jeremy Smith, Jenni A. Stanley, Andrew D. Read, and Sofie M. Van Parijs. Sleuthing with sound: Understanding vessel activity in marine protected areas using passive acoustic monitoring. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309005.

Kohler:2022:TSP

[KDS22] Jonathan Köhler, Ewa Dönitz, and Frank Schätter. Transitions for ship propulsion to 2050: the AHOY combined qualitative and quantitative scenarios. Marine Policy, 140 (??):??, June 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000963.

Koomson:2020:CVF

[KDVR20] Daniel Koomson, Katherine Sian Davies-Vollum, and Debadayita Raha. Characterising the vulnerability of fishing households to climate and environmental change: Insights from Ghana. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308590.

Kenchington:2020:GPD

[Ken20] Trevor J. Kenchington. La Grande Pêche: Drivers of change in the saltfish fisheries on Grand Bank through four centuries. Marine Policy, 115(??):??, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300132.

Knapp:2023:EAE

[KF23] Sabine Knapp and Philip Hans Franses. Empirical analysis of the effectiveness of the legislative framework in the maritime industry. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004481.

Karcher:2020:TSP

[KFB^{+}20] Denis B. Karcher, Elodie Fache, Annette Breckwoldt, Hugh Govan, Xochitl E. Elías Ilosvay, Juliette Kon Kam King, Léa Riera, and Catherine Sabinot. Trends in South Pacific fisheries management. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19306499.

Kelly:2020:SLM

[KFM^{+}20] Rachel Kelly, Aysha Fleming, Mary Mackay, Carolina García, and Gretta T. Pecl. Social licence for marine protected areas. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305196.

Kemp:2020:CPO

[KFP20] Paul S. Kemp, Rainer Froese, and Daniel Pauly. COVID-19 provides an opportunity to advance a sustainable UK fisheries policy in a post-Brexit brave new world. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20301032.

Kelly:2022:FFO

[KFS^{+}22] Rachel Kelly, Paul Foley, Robert L. Stephenson, Alistair J. Hobday, Gretta T. Pecl, Fabio Boschetti, Christopher Cvitanovic, Aysha Fleming, E. A. Fulton, Kirsty L. Nash, Barbara Neis, Gerald G. Singh, and E. Ingrid van Putten. Foresighting future oceans: Considerations and opportunities. Marine Policy, 140(??):??, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000689.

Khan:2024:IFT

[KgJqY^{+}24] Alexander M. A. Khan, Ming guo Jiang, Xiao qiang Yang, Izza Mahdiana Apriliani, Noir Primadona Purba, Budy Wiryawan, Am Azbas Taurusman, and Buntora Pasaribu. Illegal fishing threatens the sustainability of future tuna commodities in Indonesia. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004694.

Koehler:2022:APE

[KGL22] Lydia Koehler, Ioannis Giovos, and Jason Lowther. The application of precaution in elasmobranch conservation and management in the Mediterranean Sea. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004413.

Kraan:2020:SSP

[KGP^{+}20] Marloes Kraan, Rolf Groeneveld, Annet Pauwelussen, Tim Haasnoot, and Simon R. Bush. Science, subsidies and the politics of the pulse trawl ban in the European Union. Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309881.

Ko:2022:CSL

[KGP^{+}22] Dong-Hui Ko, Yunzheng Ge, Jin-Soon Park, Lei Liu, Changlei Ma, Fengyun Chen, Jingping Peng, Seung-Koo Kang, Jin-Hak Yi, and Weimin Liu. A comparative study of laws and policies on supporting marine energy development in China and Korea. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200104X.

Kirkham:2020:DSM

[KGW20] Nicholas R. Kirkham, Kristina M. Gjerde, and A. Meriwether W. Wilson. DEEP-SEA mining: Policy options to preserve the last frontier — Lessons from Antarctica’s mineral resource convention. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305615.

Kolakowski:2024:SVS

[KGW24] Paweł Kołakowski, Mateusz Gil, and Krzysztof Wróbel. Stakeholders’ view on selecting a prospective location for marine alternative fuels bunkering station. Marine Policy, 164(??):??, June 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001623.

Koehler:2024:UCA

[KH24] Holly Koehler and Bianca Haas. The use of corrective action frameworks in international fisheries management. Marine Policy, 168(??):??, October 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002884.

Kuhnen:2022:CLM

[KHdSM^{+}22] Vanessa V. Kuhnen, Kevin Hopkins, Lucas dos Santos Mota, Otavio M. Sousa, and Eduardo G. Sanches. Challenges and lessons from marine finfish farming in Brazil. Marine Policy, 138(??):??, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000264.

Kim:2023:VAK

[KHH^{+}23] Moo-Jin Kim, Jae-Beum Hong, In-Seong Han, Joon-Soo Lee, and Do-Hoon Kim. Vulnerability assessment of Korean fisheries to climate change. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002683.

Kvamsdal:2023:EOT

[KHLS23] Sturla Kvamsdal, Arnt O. Hopland, Yuanhao Li, and Simon Selle. Expert opinions on threats and impacts in the marine environment. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004298.

Khan:2020:HCC

[KHMM20] Wajiha Khan, Upamanyu Hore, Soham Mukherjee, and Gowri Mallapur. Human-crocodile conflict and attitude of local communities toward crocodile conservation in Bhitarkanika Wildlife Sanctuary, Odisha, India. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300739.

Kilic:2024:BBC

[Kıl24a] Arınç Onat Kılıç. Beyond bluewashing: a critical examination of labeling blue bonds. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24001507.

Kilic:2024:GSB

[Kıl24b] Arınç Onat Kılıç. Seychelles blue bond: Indebting ecological restructuring of fisheries. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24001428.

Kim:2020:WDR

[Kim20] Hyun Jung Kim. What did the Republic of Korea and Japan mean by the term “joint development” in their 1974 agreement? Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300154.

Kim:2021:TPF

[Kim21] Hyun Jung Kim. Third party fishing in an undelimited area under the Republic of Korea–Japan Fisheries Agreement. Marine Policy, 125(??):??, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000129.

Kim:2023:ACL

[Kim23] Sojeong Kim. Adaptive capacity and local-level fisheries co-management activities: a case of South Korea. Marine Policy, 154(??):??, August 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001926.

Kim:2024:LIB

[Kim24] Rakhyun E. Kim. The likely impact of the BBNJ Agreement on the architecture of ocean governance. Marine Policy, 165 (??):??, July 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2400188X.

Kizielewicz:2024:ECC

[Kiz24] Joanna Kizielewicz. Economic consequences of the COVID’19 pandemic for functioning of the cruise shipping industry in the Baltic Sea Region. Marine Policy, 159(??): ??, January 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2300458X.

Khuu:2021:GAD

[KJE21a] Duong T. Khuu, Peter J. S. Jones, and Paul Ekins. A Governance analysis of Con Dao National Park, Vietnam. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305421.

Khuu:2021:GAN

[KJE21b] Duong T. Khuu, Peter J. S. Jones, and Paul Ekins. Governance analysis of Nha Trang Bay and Cu Lao Cham Marine Protected Areas, Vietnam. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309775.

Kushardanto:2022:HFT

[KJS^{+}22] Hari Kushardanto, Raymond Jakub, Wahid Suherfian, Tarlan Subarno, Ahmad Isa Ansyori, La Sara, Naslina Alimina, Fajriah, La Ode Kardini, Emilio de la Rosa, Ade Yuliani, Eva Medianti, Imanda Pradana, Haris Setiawan, Yoni Muhammad, Lely Fajriah Djafar, Stephen Box, Courtney Cox, and Stuart J. Campbell. Household finances and trust are key determinants of benefits from small-scale fisheries co-management. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003311.

Kerber:2022:LFA

[KK22] Heide Kerber and Johanna Kramm. From laissez-faire to action? Exploring perceptions of plastic pollution and impetus for action. Insights from Phu Quoc Island. Marine Policy, 137(??):??, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005352.

Kaiser:2021:GEE

[KKA^{+}21] Brooks A. Kaiser, Melina Kourantidou, Dewan Ahsan, Sergey Bakanev, AnnDorte Burmeister, Ginny Eckert, Linda M. Fernandez, Hyun Pyo Hong, Albert Alexandre Monsalve, Darrell Mullowney, Bjørg Helen Nøstvold, Hojeong Park, Edward Poulsen, Lars Ravn-Jonsen, Chris Siddon, Jan H. Sundet, Kanae Tokunaga, and Masashi Yamamoto. Global ecological and economic connections in Arctic and sub-Arctic crab markets. Marine Policy, 127 (??):??, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100052X.

Kiilu:2025:VAE

[KKAO^{+}25] Benedict Kiilu, Boaz Kaunda-Arara, Remmy Oddenyo, Gladys Okemwa, Elizabeth Mueni, Peter Musembi, Bernerd Fulanda, Lameck Menya-Otieno, Maurine Okeri, Grace Nduku, Jonathan Musembei, Mohamed Omar, and Edward Kimani. Vulnerability assessment of elasmobranch species to fisheries in coastal Kenya: Implications for conservation and management policies. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24004597.

Kaliszewski:2020:KFC

[KKDK20] Adam Kaliszewski, Arkadiusz Kozłowski, Janusz Dabrowski, and Hanna Klimek. Key factors of container port competitiveness: a global shipping lines perspective. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19307146.

Knutsen:2022:LRM

[KKO^{+}22] Jan Atle Knutsen, Alf Ring Kleiven, Esben Moland Olsen, Halvor Knutsen, Sigurd Heiberg Espeland, Tonje Knutsen Sørdalen, Susanna Huneide Thorbjørnsen, Jeffrey A. Hutchings, Albert Fernández-Chacón, Mats Huserbråten, David Villegas-Ríos, Kim Tallaksen Halvorsen, Portia Joy Nillos Kleiven, Thomas Kiland Langeland, and Even Moland. Lobster reserves as a management tool in coastal waters: Two decades of experience in Norway. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005194.

Kadfak:2021:MTJ

[KL21a] Alin Kadfak and Sebastian Linke. More than just a carding system: Labour implications of the EU’s illegal, unreported and unregulated (IUU) fishing policy in Thailand. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000555.

Konrad:2021:CBA

[KL21b] Leilani Konrad and Arielle Levine. Controversy over beach access restrictions at an urban coastal seal rookery: Exploring the drivers of conflict escalation and endurance at Children’s Pool Beach in La Jolla, CA. Marine Policy, 132(??): ??, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002700.

Koehler:2022:PMS

[KL22] Lydia Koehler and Jason Lowther. Policy making for sharks and the role and contribution of non-governmental organisations in the fulfilment of international legal obligations. Marine Policy, 144(??):??, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002755.

Kim:2024:EIG

[KL24] Dae Eui Kim and Song Soo Lim. Economic impacts of the European Union carding system on global fish trade. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002069.

Klein:2021:MAV

[Kle21] Natalie Klein. Maritime autonomous vehicles and international laws on boat migration: Lessons from the use of drones in the Mediterranean. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000579.

Kim:2023:PAC

[KLK23] Boram Kim, Hyejin Lee, and Taeil Kim. Possibility analysis of complex marine accidents caused by large marine plastic litter in the Republic of Korea. Marine Policy, 152(??): ??, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001185.

Kim:2022:HCM

[KLS22] A-Rom Kim, Sung-Woo Lee, and Young-Joon Seo. How to control and manage vessels’ ballast water: the perspective of Korean shipping companies. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000549.

Kinds:2025:UHH

[KLS^{+}25] Arne Kinds, Natali Lazzari, Daniel J. Skerritt, Gillian B. Ainsworth, Adriana Rosa Carvalho, Katina Roumbedakis, Patricia Majluf, Maria Lourdes Palomares, Ussif Rashid Sumaila, and Sebastian Villasante. Unveiling the hidden hands: Analysis of corporate ownership of industrial tuna fishing vessels in the Eastern Pacific Ocean. Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004743.

Kinds:2021:CLA

[KLSG21] Arne Kinds, Pascal Le Floc’h, Stijn Speelman, and Olivier Guyader. Challenging the ‘artisanal vs. industrial’ dichotomy in French Atlantic fisheries: an organizational typology of multi-vessel fishing firms. Marine Policy, 134(??): ??, December 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100364X.

Kolavani:2025:RLF

[KM25] Narges Jalili Kolavani and Charles Mather. Regulating a ‘fish out of place’: a global assessment of farmed salmon escape policies and frameworks. Marine Policy, 173(??): ??, March 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24005724.

Khan:2020:RDT

[KMG^{+}20] Alexander M. A. Khan, Aileen C. Mill, Tim S. Gray, Mingguo Jiang, Hazmi Arief, Arthur Brown, Amirul Karman, and Nicholas V. C. Polunin. Reliability of the data on tuna catches obtained from the dockside in Indonesia: a study of stakeholders’ perceptions. Marine Policy, 122(??): ??, December 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20308885.

Koerner:2024:TSF

[KMJ24] Lauren M. Koerner, Michael C. Melnychuk, and Ernesto Jardim. Testing the sensitivity of fishery assessment scores to changes in the MSC Fisheries Standard requirements. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000034.

Kokosalakis:2021:ERE

[KMM21] George Kokosalakis, Anna Merika, and Xakousti-Afroditi Merika. Environmental regulation on the energy-intensive container ship sector: a restraint or opportunity? Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309246.

Katikiro:2024:EPS

[KMM24] Robert Katikiro, Shauna L. Mahajan, and Modesta Medard. The establishment, persistence, and scaling of marine community-based natural resource management in four coastal districts in Tanzania. Marine Policy, 169(??): ??, November 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003397.

Kalidin:2020:BEM

[KMS20] Bhavnah Komul Kalidin, Carlo Mattone, and Marcus Sheaves. Barriers to effective monitoring and evaluation of small-scale fisheries in small island developing states: an example from Mauritius. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1930541X.

Kilic:2025:CAD

[KMSN25] Kemal Ihsan Kilic, Samir Maity, Inkyung Sung, and Peter Nielsen. Challenges and AI-driven solutions in maritime search and rescue planning: a comprehensive literature review. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001071.

Kull:2021:IGP

[KMT^{+}21] M. Kull, J. R. Moodie, H. L. Thomas, S. Mendez-Roldan, A. Giacometti, A. Morf, and I. Isaksson. International good practices for facilitating transboundary collaboration in Marine Spatial Planning. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18304330.

Keijser:2025:LAB

[KMTvT25] Xander Keijser, Igor Mayer, Hilde Toonen, and Jan P. M. van Tatenhove. Learning across borders: Establishing transboundary coordination in Maritime Spatial Planning in the North Sea Region — lessons from the NorthSEE project. Marine Policy, 180(??):??, October 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2500185X.

Komatsu:2021:EMP

[KN21] Masayuki Komatsu and Tomoko Nakamura. Ecosystem management for protection of Japanese ocean and fishery resources. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002931.

Knudsen:2025:LTS

[Knu25] Ståle Knudsen. ‘Turkish salmon’, Norwegian entrepreneurs, and the global salmon value chain. Marine Policy, 171(??): ??, January 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24004706.

Kim:2020:PAL

[KNY20] Ju-Hee Kim, Jungho Nam, and Seung-Hoon Yoo. Public acceptance of a large-scale offshore wind power project in South Korea. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302499.

Kim:2022:PPB

[KNY22] Ju-Hee Kim, Jungho Nam, and Seung-Hoon Yoo. Public perceptions of blue carbon in South Korea: Findings from a choice experiment. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002834.

Kobayashi:2022:CIC

[Kob22] Masanori Kobayashi. The COVID-19 impacts and challenges to achieving sustainability in Japan’s fisheries and aquaculture. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002081.

Kapapa:2024:VFB

[KOBM24] Lyidia Kapapa, Paul Onyango, Philip Bwanthondi, and Prosper Mfilinge. The vulnerability of fisheries-based livelihoods to climate variability and change in coastal small pelagic fishing communities in Tanzania. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003427.

Karahalil:2021:EPC

[KOBS21] Meric Karahalil, Burcu Ozsoy, Ersan Basar, and Tanzer Satır. The evaluation of the Polar Code by the survey conducted with those who have sailed in polar regions, and suggestions for further improvement. Marine Policy, 128 (??):??, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001135.

Kuo:2023:CCD

[KOCL23] Ching-Ching Kuo, Ching-Hsiewn Ou, Wen-Yan Chiau, and Chao-Shing Lee. Challenges and conflicts in the designation of Taiwan’s underwater cultural heritage protected areas. Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001835.

Koehler:2022:CIT

[Koe22] Holly Koehler. The case for industry transparency in supporting sustainable tuna fisheries. Marine Policy, 136(??): ??, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19306384.

Kojima:2023:IGP

[Koj23] Chie Kojima. Integration of general principles of international environmental law into the Law of the Sea: Assessment and challenges. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000246.

Kontovas:2020:IAQ

[Kon20] Christos A. Kontovas. Integration of air quality and climate change policies in shipping: the case of sulphur emissions regulation. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304683.

Kowalski:2024:FPT

[Kow24] Jean-Marie Kowalski. The French procurement transformations: a historical perspective over sovereignty and innovation. Marine Policy, 167(??):??, September 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002343.

Kinyua:2025:SCA

[KOW^{+}25] Damaris Njeri Kinyua, Dalmas O. Oyugi, Andrew Wamukota, U. Rashid Sumaila, Sarah Ater, and Nelly Isigi Kadagi. A supply chain analysis of artisanal billfish fishery along the Kenyan coast. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25001770.

Kim:2024:USR

[KP24] Jong-Kwan Kim and Deuk-Jin Park. Understanding of sailing rule based on COLREGs: Comparison of navigator survey and automated collision–avoidance algorithm. Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300427X.

Kim:2020:DCN

[KPN20] Tae-Hyun Kim, Cheol-Hyung Park, and Jongoh Nam. The determinants of changes in the number of fishers employed by fisheries household in the Republic of Korea using count data models. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930569X.

Kraabel:2022:IAB

[Kra22] Kristine Dalaker Kraabel. Institutional arrangements in a BBNJ treaty: Implications for Arctic marine science. Marine Policy, 142(??):??, August 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305081.

Kaminski:2022:CSO

[KS22] Tomasz Kamiński and Rafał Szewczyk. The coastal state obligation not to impede the laying or maintenance of submarine pipelines on the continental shelf according to United Nations convention on the law of the sea. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001336.

Kaeriyama:2023:EBS

[KS23] Masahide Kaeriyama and Isao Sakaguchi. Ecosystem-based sustainable management of chum salmon in Japan’s warming climate. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003755.

Kemp:2023:FAU

[KSB^{+}23a] Paul S. Kemp, Gowshika Subbiah, Richard Barnes, Kristina Boerder, Bethan C. O’Leary, Bryce D. Stewart, and Chris Williams. Future advances in UK marine fisheries policy: Integrated nexus management, technological advance, and shifting public opinion. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003827.

Kemp:2023:FMF

[KSB^{+}23b] Paul S. Kemp, Gowshika Subbiah, Richard Barnes, Kristina Boerder, Bethan C. O’Leary, Bryce D. Stewart, and Chris Williams. The future of marine fisheries management and conservation in the United Kingdom: Lessons learnt from over 100 years of biased policy. Marine Policy, 147(??): ??, January 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001221.

Kuempel:2025:FBP

[KSO^{+}25] Caitlin D. Kuempel, Jackson Stockbridge, Tess O’Neill, Laura L. Griffiths, Christopher L. J. Frid, and Christopher J. Brown. Frameworks and best practices for cumulative impact accounting in offshore energy development. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005712.

Kageyama:2025:EEC

[KSÖB25] Shun Kageyama, Madlen Sobkowiak, Henrik Österblom, and Robert Blasiak. Exploring evidence of cascading change towards stewardship in the Japanese seafood industry. Marine Policy, 175(??):??, May 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000417.

Kim:2023:KBA

[KSS23] A-Rom Kim, Jungyong Seo, and Young-Joon Seo. Key barriers to adopting onshore power supply to reduce port air pollution: Policy implications for the maritime industry in South Korea. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003998.

Kumawat:2023:JFA

[KSV^{+}23] Tarachand Kumawat, Raju Saravanan, Kavungal Vinod, Ashok Kumar Jaiswar, Geetanjali Deshmukhe, Latha Shenoy, Damodaran Divu, Kunnummel Krishnan Joshi, and Achamveetil Gopalakrishnan. Jellyfish fisheries along the Gujarat coast, India: Status and challenges. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000817.

Krueck:2022:BMP

[KTC^{+}22] Nils C. Krueck, Colm Tong, Courtney Cox, Eric A. Treml, Kay Critchell, Iliana Chollett, Dedi S. Adhuri, Maria Beger, Dominic Muenzel, Daniel Holstein, Stuart J. Campbell, and Raymond Jakub. Benefits of measurable population connectivity metrics for area-based marine management. Marine Policy, 144(??):??, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002573.

Komi:2024:IRF

[KTI^{+}24] Gentle Wilson Komi, Nelson Turyahabwe, Prossy Isubikalu, Olufemi Martins Adesope, Stephen Okodudu, and Nkopuyo Udoekpo Sunday. Influence of reciprocal financial resources on shrimp fishers’ income sources in Southern Nigeria. Marine Policy, 162(??):??, April 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000484.

Kulaga:2022:PMS

[Kuł22] Łukasz Kułaga. Prohibition of massive and serious pollution of marine environment as a jus cogens. Identification, legal bases and consequences in view of the recent work of the International Law Commission. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002640.

Kularatne:2023:OMN

[Kul23] Ranil Kavindra Asela Kularatne. Occurrence of marine nonindigenous (NIS) species: Current status, management approaches and challenges in Sri Lanka. Marine Policy, 149 (??):??, March 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000040.

Kalinin:2020:SAR

[KV20] Nikita Kalinin and Mark Vershinin. Strategic analysis of the Russian crab quota auction in 2019. Marine Policy, 122(??): ??, December 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2030912X.

Kuismanen:2023:IEV

[KVL^{+}23] Lauri M. J. Kuismanen, Elina A. Virtanen, Juho Lappalainen, Lasse Kurvinen, Penina Blankett, and Markku Viitasalo. Identifying ecologically valuable marine areas to support conservation and spatial planning at scales relevant for decision making. Marine Policy, 158(??):??, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004232.

Kirkfeldt:2020:OAN

[KvTNL20] Trine Skovgaard Kirkfeldt, Jan P. M. van Tatenhove, Helle Nedergaard Nielsen, and Sanne Vammen Larsen. An ocean of ambiguity in Northern European marine spatial planning policy designs. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20303213.

Kadagi:2021:COS

[KWF^{+}21] Nelly Isigi Kadagi, Nina Wambiji, Sean Thomas Fennessy, Micheal S. Allen, and Robert N. M. Ahrens. Challenges and opportunities for sustainable development and management of marine recreational and sport fisheries in the Western Indian Ocean. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310010.

Kao:2023:CDP

[KWG23] Shih-Ming Kao, Wei-Ning Wu, and Meng-En Gu. Coastal disaster protection and hazard perceptions of the residents of Orchid Island, Taiwan. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22005061.

Koehler:2025:WMS

[KWK^{+}25] Lydia Koehler, Paddy Walker, Irene Kingma, Simone Niedermüller, and Jason Lowther. What the Marine Strategy Framework Directive did and did not change for Chondrichthyes. Marine Policy, 176(??):??, June 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000521.

Kimani:2020:APG

[KWMM20a] Patrick Kimani, Andrew Wamukota, Julius O. Manyala, and Chrisestom Mwatete Mlewa. Actors’ perceptions of government performance in support of value chain development in marine small-scale fisheries in Kenya. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20308678.

Kimani:2020:FIF

[KWMM20b] Patrick Kimani, Andrew Wamukota, Julius O. Manyala, and Chrisestom Mwatete Mlewa. Factors influencing financial performance in marine small-scale fisheries value chain in Kenya. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308642.

Kadagi:2020:PCR

[KWS20] N. I. Kadagi, N. Wambiji, and M. E. Swisher. Potential for conflicts in recreational and artisanal billfish fisheries on the coast of Kenya. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307341.

Kuwae:2022:IBC

[KWY^{+}22] Tomohiro Kuwae, Atsushi Watanabe, Satoru Yoshihara, Fujiyo Suehiro, and Yoshihisa Sugimura. Implementation of blue carbon offset crediting for seagrass meadows, macroalgal beds, and macroalgae farming in Japan. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000434.

Kullerud:2020:AGR

[KY20] Lars Kullerud and Oran R. Young. Adding a Gakkel Ridge regime to the evolving Arctic Ocean governance complex. Marine Policy, 122(??):??, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309167.

Kaifu:2021:MGE

[KYMW21] Kenzo Kaifu, Kazuki Yokouchi, Michael J. Miller, and Izumi Washitani. Management of glass eel fisheries is not a sufficient measure to recover a local Japanese eel population. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004176.

Kennard:2022:TTH

[KZR22] Alice Kennard, Pengfei Zhang, and Sriram Rajagopal. Technology and training: How will deck officers transition to operating autonomous and remote-controlled vessels? Marine Policy, 146(??):??, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003736.

Livingstone:2023:DFC

[LA23] Olivia Livingstone and Brandon P. Anthony. Documenting fisheries co-management in action: Case of the Grand Cape Mount County Co-Management Association, Liberia. Marine Policy, 157(??):??, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003779.

Love:2024:FAP

[LAF^{+}24] David C. Love, Frank Asche, Jillian Fry, Mark Brown, Ly Nguyen, Taryn M. Garlock, Elizabeth M. Nussbaumer, Gabriela L. Sarmiento, Sigbjørn Tveteraas, and Roni Neff. Fisheries and aquaculture by-products: Case studies in Norway, United States, and Vietnam. Marine Policy, 167(??): ??, September 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24002744.

Lajeunesse:2021:CAO

[Laj21] Adam Lajeunesse. Canada’s Arctic Offshore and Patrol Ships (AOPS): Their history and purpose. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309702.

Lambach:2021:FTH

[Lam21] Daniel Lambach. The functional territorialization of the high seas. Marine Policy, 130(??):??, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001901.

Llerena:2025:AEL

[LASG^{+}25] Pablo D. Llerena, Jorge J. Avila-Santamaría, María V. Gabela, Sara Purca, Carlos F. Mena, and Susana A. Cárdenas. Assessing economic losses in artisanal fisheries from marine plastic pollution in coastal Ecuador and Peru. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005530.

Latulippe:2025:RIK

[Lat25] Nicole Latulippe. Race, indigenous knowledge, and a relational alternative in fisheries policy research. Marine Policy, 175(??):??, May 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25000156.

LeBrenne:2021:SOC

[LBL21] Valérie Le Brenne, Laetitia Bisiaux, and Frédéric Le Manach. Sustainable objectives and commitments deceived by fisheries subsidies for ‘temporary cessations’ in times of COVID. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002815.

Lewin:2023:TEC

[LBWS23] Wolf-Christian Lewin, Fanny Barz, Marc Simon Weltersbach, and Harry V. Strehlow. Trends in a European coastal fishery with a special focus on small-scale fishers — implications for fisheries policies and management. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002075.

Li:2020:LSE

[LC20a] Xintong Li and Yen-Chiang Chang. The legal status of the “employment salvage contract” in the Chinese maritime salvage law: a challenge to the principle of ‘no cure, no pay’. Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300014.

Liu:2020:ERM

[LC20b] Xiaofei Liu and Yen-Chiang Chang. An emergency responding mechanism for cruise epidemic prevention — taking COVID-19 as an example. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20302451.

Liu:2022:ERF

[LC22] Xintian Liu and Suisui Chen. Has environmental regulation facilitated the green transformation of the marine industry? Marine Policy, 144(??):??, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002858.

Li:2023:SCC

[LC23] Xintong Li and Yen-Chiang Chang. A step closer to the convention on the Protection of Underwater Cultural Heritage 2001: China’s latest efforts in regulation. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003931.

Liu:2021:SCT

[LCAX21] Lijun Liu, Jingjie Chu, James L. Anderson, and Jintao Xu. Sustainability comparisons in the triple bottom line for Chinese fisheries. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309052.

Lee:2020:DDF

[LCC20] Tzu-Ming Lee, Pei-Yu Chi, and Kuo-I. Chang. Duration and determinants of fishery trade patterns: Evidence from OECD countries. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309583.

LeBihan:2020:RVC

[LCL20] Véronique Le Bihan, Marie Catalo, and Jeanine Le Bihan. Reorganization of the value chain activities of oyster companies on the Atlantic coast following health crises in France (2006–2013). Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16306868.

Liu:2024:LCR

[LCL24] Xiaofei Liu, Shuning Cao, and Yanan Liu. Legal challenges of regulating cross-border nuclear pollution from ship ballast water under the background of Fukushima wastewater release and China’s response. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24002860.

Luquin-Covarrubias:2022:LGE

[LCMBGP22] Marlene Anaid Luquin-Covarrubias, Enrique Morales-Bojórquez, and Sergio Scarry González-Peláez. The last geoduck: the experience of geoduck clam fishery management in the Mexican Pacific Ocean. Marine Policy, 143(??): ??, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001920.

LeDuff:2020:OCD

[LDAC20] Matthieu Le Duff, Pascal Dumas, Michel Allenbach, and Olivier Cohen. An orientation for coastal disaster risks management and prevention policy in a global warming context: Case study in Ouvea (New Caledonia). Marine Policy, 117 (??):??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X16307084.

Li:2020:CSR

[LDC^{+}20] Feixue Li, Dongdong Ding, Zhenjie Chen, Honghua Chen, Ting Shen, Qianlian Wu, and Chenye Zhang. Change of sea reclamation and the sea-use management policy system in China. Marine Policy, 115(??):??, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302738.

Latapi:2024:HFC

[LDC^{+}24] Mauricio Latapí, Brynhildur Davídsdóttir, David Cook, Lára Jóhannsdóttir, Andrea Marin Radoszynski, and Kenneth Karlsson. Hydrogen fuel cells in shipping: a policy case study of Denmark, Norway, and Sweden. Marine Policy, 163(??):??, May 2024. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001076.

Lang:2025:SFM

[LDG^{+}25] Stephen D. J. Lang, Philip D. Doherty, Brendan J. Godley, Shing Yip Lee, Felix Leung, Manmohan D. Sharma, and Kristian Metcalfe. Spatial footprint of maritime vessels in the waters of a fast-growing coastal megalopolis. Marine Policy, 179(??):??, September 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2500154X.

LeBusque:2021:INM

[LDL21] Brianna Le Busque, Jillian Dorrian, and Carla Litchfield. The impact of news media portrayals of sharks on public perception of risk and support for shark conservation. Marine Policy, 124(??):??, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030991X.

Lin:2023:EVM

[LDL^{+}23] Chun-Yu Lin, Gui-Lin Dai, Yang Liu, Meng-Qi Zhang, Ying Liu, Wen Jiang, Xiu-Mei Fu, and Han-Xue Chen. Ecological value of mariculture shellfish resources in China: Assessment and management. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004535.

Lee:2024:IBD

[Lee24] Dongkeun Lee. Influences behind the development of South Korea’s shipbuilding industry from the 1960s to the 2000s. Marine Policy, 167(??):??, September 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002495.

Lopez-Ercilla:2021:VMS

[LEERRM^{+}21] Ines Lopez-Ercilla, Maria Jose Espinosa-Romero, Francisco J. Fernandez Rivera-Melo, Stuart Fulton, Rebeca Fernández, Jorge Torre, Araceli Acevedo-Rosas, Arturo J. Hernández-Velasco, and Imelda Amador. The voice of Mexican small-scale fishers in times of COVID-19: Impacts, responses, and digital divide. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002177.

Levine:2020:SUA

[Lev20] Arielle Levine. Sunscreen use and awareness of chemical toxicity among beach goers in Hawaii prior to a ban on the sale of sunscreens containing ingredients found to be toxic to coral reef ecosystems. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19309029.

Luhtala:2021:BSI

[LEVET21] Hanna Luhtala, Anne Erkkilä-Välimäki, Søren Qvist Eliasen, and Harri Tolvanen. Business sector involvement in maritime spatial planning — Experiences from the Baltic Sea region. Marine Policy, 123(??):??, January 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309465.

Liu:2025:NUW

[LF25] Chaomin Liu and Yuan Feng. Navigating uncharted waters: Legal challenges and the future of unmanned underwater vehicles in maritime military cyber operations. Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004305.

Lout:2022:DWS

[LFLF22] Gabrielle Lout, Juno Fitzpatrick, Alejandro J. Garcia Lozano, and Elena Finkbeiner. Decent work in a seascape of livelihoods: Regional evaluation of the shrimp and groundfish fishery of the Guianas–Brazil Shelf. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002780.

Llorente:2020:AEP

[LFPBD^{+}20] Ignacio Llorente, José Fernández-Polanco, Elisa Baraibar-Diez, María D. Odriozola, Trond Bjørndal, Frank Asche, Jordi Guillen, Lamprakis Avdelas, Rasmus Nielsen, Maria Cozzolino, Manuel Luna, José L. Fernández-Sánchez, Ladislao Luna, Cristóbal Aguilera, and Bernardo Basurco. Assessment of the economic performance of the seabream and seabass aquaculture industry in the European Union. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307304.

Luna:2024:BBA

[LFVCF24] Manuel Luna, Simon Fernandez-Vazquez, Emilio Tereñes Castelao, and Álvaro Arias Fernández. A blockchain-based approach to the challenges of EU’s environmental policy compliance in aquaculture: From traceability to fraud prevention. Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004256.

Lasner:2024:ADS

[LG24] Tobias Lasner and Antje Gimpel. Aquaculture as a dysfunctional system of action; why does fish farming stagnate in Germany? Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004032.

Lester:2024:RMA

[LGF24] Sarah E. Lester, Rebecca R. Gentry, and Halley E. Froehlich. The role of marine aquaculture in contributing to the diversity and stability of U.S. seafood production. Marine Policy, 160(??):??, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005274.

Lauret:2024:SGP

[LGLL24] Valentin Lauret, Olivier Gimenez, Hélène Labach, and Nicolas Lescureux. Smokescreen or groundwork? The paradox of ecological monitoring in the management of Marine Protected Areas. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24003816.

Lise:2025:BNB

[LGP25] Frehen Lise, Ferraro Gianluca, and Failler Pierre. Blue Nature-based solutions in marine and coastal EU policies: Challenges, recommendations and policy opportunities throughout the policy cycle. Marine Policy, 177(??): ??, July 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25001022.

Lomonico:2021:OFP

[LGW^{+}21] Serena Lomonico, Mary G. Gleason, Jono R. Wilson, Darcy Bradley, Kate Kauer, Richard J. Bell, and Thomas Dempsey. Opportunities for fishery partnerships to advance climate-ready fisheries science and management. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20308988.

Liu:2024:CAP

[LH24] Ruiqiang Liu and Zhanfeng Hu. Controversies and amendment proposals on the China Coast Guard Law. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001131.

Link:2025:BCE

[LH25] Jason S. Link and Janne B. Haugen. The business case for ecosystem-based management. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24004858.

Liu:2023:ECC

[LHPJ23] Yang Liu, Limin Han, Zhaobin Pei, and Yiying Jiang. Evolution of the coupling coordination between the marine economy and urban resilience of major coastal cities in China. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005036.

Larson:2023:SID

[LHTR23] Silva Larson, Sreynov Hoy, Somony Thay, and Michael A. Rimmer. Sustainable and inclusive development of finfish mariculture in Cambodia: Perceived barriers to engagement and expansion. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004869.

Lear:2020:SER

[LHV^{+}20] D. Lear, P. Herman, G. Van Hoey, L. Schepers, N. Tonné, M. Lipizer, F. E. Muller-Karger, W. Appeltans, W. D. Kissling, N. Holdsworth, M. Edwards, E. Pecceu, H. Nygård, G. Canonico, S. Birchenough, G. Graham, K. Deneudt, S. Claus, and P. Oset. Supporting the essential — Recommendations for the development of accessible and interoperable marine biological data products. Marine Policy, 117 (??):??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19301915.

Leary:2021:GDP

[LHW21] David Leary, Nien-Tsu Alfred Hu, and Pei-Fu Wu. Gilt dragons, porcelain and plundered doubloons: a comparative analysis of Taiwan and Australian underwater cultural heritage protection laws. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003985.

Li:2023:UAD

[Li23a] Shuo Li. Unilateral actions in the development of the law of the sea. Marine Policy, 153(??):??, July 2023.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001859.

Li:2023:FTA

[Li23b] Tao Li. Fisheries tenure arrangement in China: Legislative ambiguity, judicial settlement conflicts, and the gap in protecting traditional fishers’ rights. Marine Policy, 152(??): ??, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001343.

Liao:2025:IAC

[Lia25] Boyu Liao. Impact of the amended Chinese Marine Environmental Protection Law on the law enforcement action of the Coast Guard — with a focus on the discharge of ballast water after Japan’s releasing of nuclear wastewater into the sea. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004317.

Leong:2024:AFT

[LIK^{+}24] Kirsten Mya Leong, Rebecca J. Ingram, Danika Kleiber, Sofya Hoshaiah Long, Anthony Mastitski, Karma Norman, Changhua Weng, and Sarah Wise. Aligning fisheries terminology with diverse social benefits. Marine Policy, 170(??): ??, December 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003750.

Lin:2020:PSW

[Lin20] Zhen Lin. The protection of sunken WWII warships located in Indonesian or Malaysian territorial waters. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19303240.

Lindley:2023:FNC

[Lin23] Jade Lindley. Fishing non-compliance and culture. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23001082.

Lin:2024:JLE

[Lin24] Zhen Lin. Joint law enforcement mechanism with regard to underwater cultural heritage at sea led by China Coast Guard: Legislation and practice. Marine Policy, 160(??): ??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23005122.

Lin:2025:EDB

[Lin25] Hsiu-Fen Lin. Examining the determinants of blockchain technology-enabled maritime supply chain system adoption intention: Does market turbulence play a moderating role? Marine Policy, 174(??):??, April 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000314.

Liu:2020:RCC

[Liu20] Nengye Liu. The rise of China and conservation of marine living resources in the polar regions. Marine Policy, 121(??): ??, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20303158.

Liu:2021:COH

[Liu21] Nengye Liu. China and one hundred years of the Svalbard Treaty: Past, present and future. Marine Policy, 124(??): ??, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20310046.

Liu:2022:USR

[Liu22] Wen-Hong Liu. University social responsibility to promote coastal partnership: Introducing Taiwan coastal communities. Marine Policy, 146(??):??, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003505.

Liu:2023:USI

[Liu23] Suyu Liu. Use of SDG Indicator 14.7.1 for ecosystembased fisheries management: Challenges and implications. Marine Policy, 151(??):??, May 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000878.

Liu:2024:EMP

[Liu24] Nengye Liu. Establishing marine protected areas in the southern ocean, lessons for the BBNJ agreement. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002148.

Liu:2025:ECA

[Liu25] Sizhu Liu. Ensuring compliance with Antarctic environmental protection through inspections: Problems and prospects. Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004330.

Li:2020:TMS

[LJ20] Shenghui Li and Stephen Jay. Transboundary marine spatial planning across Europe: Trends and priorities in nearly two decades of project work. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20301676.

Long:2021:GOG

[LJ21] Stephen Long and Peter J. S. Jones. Greenland’s offshore Greenland halibut fishery and role of the Marine Stewardship Council certification: a governance case study. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302736.

Lindfors:2022:SRI

[LJ22] Emil Tomson Lindfors and Stig-Erik Jakobsen. Sustainable regional industry development through co-evolution — the case of salmon farming and cell-based seafood production. Marine Policy, 135(??):??, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004668.

Li:2023:ATC

[LJ23] Shenghui Li and Stephen Jay. Addressing transboundary challenges: Exploring the interactive relations between collaborative governance and transboundary marine spatial planning in Europe. Marine Policy, 158(??):??, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2300413X.

Liu:2022:DIS

[LJL22] Chenxuan Liu, Qijun Jiang, and Markus Leibrecht. Does an increased synergy in an industry’s subsystems impact on industrial development? The case of the Fisheries industry in China. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004486.

Long:2021:GAC

[LJRHH21] Stephen Long, Peter J. S. Jones, Zoavina Randriana, and Jeneen Hadj-Hammou. Governance analysis of a community managed small-scale crab fishery in Madagascar: novel use of an empirical framework. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X17307522.

Li:2024:AIS

[LK24a] Wei Li and Mehran Idris Khan. Assessing the implications of short-term exemption clauses for insurers in Chinese fishery mutual insurance contracts. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2400112X.

Liu:2024:RRL

[LK24b] Xien Liu and Muneeb Khan. Review: “Rule of Law in the Governance of New Frontiers of the Marine Environment.”, Yen-Chiang Chang, Kum Fai Yuen, David Ong, Shih-Ming Kao, in: Frontiers in Marine Science.. (October 2023), p. 360, doi:10.3389/978 2-8325 3537-0. Marine Policy, 165 (??):??, July 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24001829.

Lees:2023:IMC

[LKB^{+}23] Liisi Lees, Krista Karro, Francisco R. Barboza, Ann Ideon, Jonne Kotta, Triin Lepland, Maili Roio, and Robert Aps. Integrating maritime cultural heritage into maritime spatial planning in Estonia. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003840.

Lorena:2022:ICC

[LKE^{+}22] Romero Lorena, Yévenes Karina, Pokrant Ekaterina, Baumberger Cecilia, Zavala Sebastián, Burgos José Miguel, Wacyk Jurij, and Cornejo Javiera. Impacts of COVID-19 on the Chilean salmon: a first approach to the effects of the pandemic in the industry. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002329.

Liu:2024:BRS

[LL24a] Wen-Hong Liu and Chih-Cheng Lin. Book review: The Social Construction of the Ocean and Modern Taiwan, Kuang-hao Hou. Routledge Publishing, Oxfordshire, England, UK (2024), xx +250 pp., ISBN 978-1-032-27099-9. Marine Policy, 167(??):??, September 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002598.

Lyu:2024:SRS

[LL24b] Wenyu Lyu and Lianbo Li. Sailing into rough seas: the difficulties and challenges Chinese women seafarers face. Marine Policy, 160(??):??, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004888.

Lee:2025:ISF

[LL25] Hsiao-Chien Lee and Wen-Hong Liu. Implementing the slow fish curriculum for SDGs: Strategies, challenges, and policy suggestions through a case study. Marine Policy, 173(??): ??, March 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24005384.

Liu:2025:ECP

[LLA^{+}25] Wen-Hong Liu, Hsiao-Chien Lee, Shingo Akaike, Li-Shu Chen, Chih-Cheng Lin, Hu Chun-Chieh, and Kuei-Chao Chang. Establishing coastal partnerships for sustainable development in Taiwan: a case of Shanshui area, Penghu. Marine Policy, 179(??):??, September 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001617.

Lin:2021:CAI

[LLCH21] Ta-Li Lin, Wen-Hong Liu, Yi Chang, and Shih-Chun Hsiao. Capacity assessment of integrated coastal management for Taiwanese local government. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003808.

Li:2023:DLC

[LLHG23] Hua Li, Jing Lu, Shanqi Hu, and Shengjun Gan. The development and layout of China’s cruise industry in the postepidemic era: Conference report. Marine Policy, 149(??): ??, March 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000374.

Liu:2022:UTE

[LLJZ22] Jia Liu, Jing Li, SooCheong (Shawn) Jang, and Yichen Zhao. Understanding tourists’ environmentally responsible behavior at coastal tourism destinations. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002251.

Luna:2023:CFR

[LLL23] Manuel Luna, Ignacio Llorente, and Ladislao Luna. A conceptual framework for risk management in aquaculture. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004249.

Liu:2024:TSD

[LLL^{+}24] Wen-Hong Liu, Chih-Cheng Lin, Cheng-Feng Lee, Chien-Shu Tsai, and Tsung-Yu Yang. Toward sustainable development of tuna longline fishery in Taiwan: Value chain analysis. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000083.

Limmon:2022:CDR

[LLLF22] Gino V. Limmon, Ken Longenecker, Ross Langston, and Erik C. Franklin. Capacity development in reproductive life history studies of tropical fishes in Ambon, Maluku, Indonesia for data limited fisheries. Marine Policy, 136(??): ??, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21005595.

Lees:2023:CNA

[LLM^{+}23] Samantha Lees, Catherine Longo, Michael C. Melnychuk, Polly Burns, and Beth Polidoro. Corrected numbers and assumptions for understanding MSC certified fishery withdrawals: Commentary on Pierucci et al. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002105.

Lowitt:2020:ESS

[LLS^{+}20] Kristen Lowitt, Charles Z. Levkoe, Andrew Spring, Colleen Turlo, Patricia L. Williams, Sheila Bird, Chief Dean Sayers, and Melaine Simba. Empowering small-scale, community-based fisheries through a food systems framework. Marine Policy, 120(??):??, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300373.

Liu:2022:RTU

[LLSY22] Wen-Hong Liu, Hsiao-Chien Lee, Wei-Ying Sung, and Tsung-Yu Yang. The roles of Taiwanese universities in coastal revitalization: a study of two case projects. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000975.

LeHeron:2020:ROG

[LLT^{+}20] Erena Le Heron, Richard Le Heron, Lara Taylor, Carolyn J. Lundquist, and Alison Greenaway. Remaking ocean governance in Aotearoa, New Zealand through boundary-crossing narratives about ecosystem-based management. Marine Policy, 122(??):??, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030868X.

Li:2022:EMT

[LLW22] Rongrong Li, Yi Liu, and Qiang Wang. Emissions in maritime transport: a decomposition analysis from the perspective of production-based and consumption-based emissions. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001725.

Lucrezi:2024:STC

[LM24] Serena Lucrezi and Tafadzwa Matiza. Sharks, tourism and conservation: a test of causative and mediating effects on scuba divers’ attitude. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23005298.

Leitao:2023:RRC

[LMC^{+}23a] Francisco Leitão, João N. Monteiro, Pedro Cabral, Maria A. Teodósio, and Ruben H. Roa-Ureta. Revealing the role of crab as bait in octopus fishery: an ecological and fishing approach to support management decisions. Marine Policy, 158(??):??, December 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004116.

Li:2023:RCD

[LMC^{+}23b] Yunzhou Li, Jian Ma, Abigail Costigan, Xiangyan Yang, Ellen Pikitch, and Yong Chen. Reconciling China’s domestic marine conservation agenda with the global 30 × 30 initiative. Marine Policy, 156(??):??, October 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003238.

Lyons:2023:EII

[LMMT23] Peci Lyons, Sara Mynott, and Jess Melbourne-Thomas. Enabling indigenous innovations to re-centre social licence to operate in the Blue Economy. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004316.

Liza:2025:SIB

[LMR25] Jannatul Islam Liza, Shapan Chandra Majumder, and Md. Hasanur Rahman. Scrutinizing the impact of blue economy factors on the economic growth in Bangladesh: an empirical study. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005426.

Leon-Mateos:2021:AOS

[LMSLMQ21] Fernando León-Mateos, Antonio Sartal, Lucas López-Manuel, and María A. Quintás. Adapting our sea ports to the challenges of climate change: Development and validation of a Port Resilience Index. Marine Policy, 130(??): ??, August 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001846.

Li:2021:HCC

[LMT21] Hua Li, Shuhan Meng, and Helong Tong. How to control cruise ship disease risk? Inspiration from the research literature. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002633.

Linke:2023:RAS

[LNRM23] Sebastian Linke, Kåre Nolde Nielsen, and Paulina Ramírez-Monsalve. Roles for advisory science in the international council for the exploration of the sea (ICES). Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22005206.

Losen:2023:SAR

[Los23] Barbara Losen. Shark attack risk on Reunion Island: Emphasis on local media construction. Marine Policy, 157(??): ??, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003846.

Losen:2024:USD

[Los24] Barbara Losen. Uses of scientific discourses and markers in shark attack risk management on Reunion Island. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000095.

Lott:2022:BWM

[Lot22] Alexander Lott. Barriers to wildlife movement in straits: Problematizing habitat connectivity across marine ecosystems. Marine Policy, 141(??):??, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001543.

Lothian:2023:BPM

[Lot23] Sarah Lothian. The BBNJ preamble: More than just window dressing. Marine Policy, 153(??):??, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001690.

Louey:2022:PBE

[Lou22] Philippa Louey. The Pacific blue economy: an instrument of political maneuver. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004917.

Lovell:2023:MAL

[Lov23] Tricia A. Lovell. Managing abandoned, lost and otherwise discarded fishing gear (derelict gear) in Eastern Caribbean small-scale fisheries: an assessment of legislative, regulatory and policy gaps. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004791.

Lane:2020:MDE

[LP20] Jesse M. Lane and Michael Pretes. Maritime dependency and economic prosperity: Why access to oceanic trade matters. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301743.

Lonsdale:2021:SLU

[LP21] Jemma-Anne Lonsdale and Claire Phillips. Space launches and the UK marine environment. Marine Policy, 129(??): ??, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000907.

Leon-Perez:2024:FST

[LPMG^{+}24] Mariana C. León-Pérez, Richard J. McLaughlin, James C. Gibeaut, Lisamarie Carrubba, Ricardo J. Colón-Rivera, and René Esteves. First steps towards untangling the sargassum legal regime in Puerto Rico. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24002008.

Lahde:2024:CCS

[LPMS24] Elisa Lähde, Mari Pohja-Mykrä, and Johanna Schreck. Cocreation of socio-ecological systems knowledge to adopt an Ecosystem-based Approach and Land–Sea Interactions in maritime spatial planning. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000770.

Lekatompessy:2025:MEE

[LPO25] Vanela Chatrin Lekatompessy, Pratiwi, and Lusya Sevyana Octavia. Book review: Marine Exploration and Exploitation of Hydrocarbons — an Environmental and Legal Perspective, Violeta S. Radovich (Ed.). Palgrave Macmillan, Cham (2024), 288, pp ISBN 978-3-031-75177-6. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001496.

Lawson:2020:DSI

[LR20] Charles Lawson and Michelle Rourke. Digital sequence information as a marine genetic resource under the proposed UNCLOS legally binding instrument. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19308061.

Lopez-Rivas:2024:WEV

[LRC24] Jose D. Lopez-Rivas and Juan-Camilo Cardenas. What is the economic value of coastal and marine ecosystem services? A systematic literature review. Marine Policy, 161 (??):??, March 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000319.

Lamont:2022:MDA

[LRD^{+}22] Timothy A. C. Lamont, Tries B. Razak, Rili Djohani, Noel Janetski, Saipul Rapi, Frank Mars, and David J. Smith. Multi-dimensional approaches to scaling up coral reef restoration. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002469.

Lekshmi:2022:GSS

[LRN^{+}22] P. S. Swathi Lekshmi, Kalidoss Radhakrishnan, R. Narayanakumar, V. P. Vipinkumar, Shinoj Parappurathu, Shyam S. Salim, B. Johnson, and Phalguni Pattnaik. Gender and small-scale fisheries: Contribution to livelihood and local economies. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005248.

Lee:2022:EVN

[LrS22] Taylor Tae-Hwee Lee and Ae rim So. Economic valuation of a national shipping company using a contingent valuation method (CVM): the case of Korea. Marine Policy, 135(??): ??, January 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004814.

Lau:2021:CIC

[LSB^{+}21] Jacqueline Lau, Sarah Sutcliffe, Michele Barnes, Emmanuel Mbaru, Innocent Muly, Nyawira Muthiga, Stephen Wanyonyi, and Joshua E. Cinner. COVID-19 impacts on coastal communities in Kenya. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004140.

Liu:2025:NCS

[LSC25] Ruisong Liu, Chengzhi Sun, and Dongshu Chen. In the name of culture and sustainability: a discourse-historical approach of Japan’s whaling policy. Marine Policy, 171(??): ??, January 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24004561.

Lozano:2022:DWF

[LSD^{+}22] Alejandro J. Garcia Lozano, Jessica L. Decker Sparks, Davina P. Durgana, Courtney M. Farthing, Juno Fitzpatrick, Birgitte Krough-Poulsen, Gavin McDonald, Sara McDonald, Yoshitaka Ota, Nicole Sarto, Andrés M. Cisneros-Montemayor, Gabrielle Lout, Elena Finkbeiner, and John N. Kittinger. Decent work in fisheries: Current trends and key considerations for future research and policy. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005339.

Lizaso:2020:NMF

[LSGG^{+}20] José Luis Sánchez Lizaso, Ivan Sola, Elena Guijarro-García, José Maria Bellido, and Ramón Franquesa. A new management framework for western Mediterranean demersal fisheries. Marine Policy, 112(??):??, February 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306104.

Liu:2021:REG

[LSH21] Chen-Hong Liu, Yue Sun, and Yang Han. Regional environmental governance of the Yellow Sea and Bohai Sea from the perspective of land and sea coordination: Conference report. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000567.

Lucas:2024:FPV

[LSOP24] Victor Lucas, Susanna Balaguer Serra, Marinka Ooteman, and Mariasole Petti. Fisher poverty, value chain equity, and resilience: the case of the Indonesian blue swimming crab and the Peruvian mahi-mahi. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2400407X.

Lu:2020:FAF

[LSY20] Yu-Heng Lu, Takahiro Sajiki, and Nobuyuki Yagi. Factors affecting fisherman satisfaction with fishermen’s selfgovernance organizations: a case study of the Taiwan Donggang Sakuraebi (Sergia lucens) production and management group. Marine Policy, 115(??):??, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304129.

Lorance:2024:ESC

[LT24] Pascal Lorance and Verena M. Trenkel. Effects of stock collapse and management on price dynamics of blackspot seabream. Marine Policy, 168(??):??, October 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003130.

Long:2021:CAG

[LTJ^{+}21] Stephen Long, Grace Thurlow, Peter J. S. Jones, Andrew Turner, Sylvestre M. Randrianantenaina, Tom Gammage, Jessica Savage, and Jeremie R. Ndriamanja. Critical analysis of the governance of the Sainte Luce Locally Managed Marine Area (LMMA), southeast Madagascar. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302635.

Lin:2023:UDC

[LTJ^{+}23] Chih-Cheng Lin, Kuo-Huan Ting, Hao-Tang Jhan, Chung-Lun Liu, Li-Shu Chen, and Wen-Hong Liu. Universitydriven coastal partnerships to address ghost gear pollution: a case study in Penghu Islands of Taiwan. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002658.

Lieske:2020:SRA

[LTRA20] David J. Lieske, Laura McFarlane Tranquilla, Robert A. Ronconi, and Sue Abbott. “Seas of risk”: Assessing the threats to colonial-nesting seabirds in Eastern Canada. Marine Policy, 115(??):??, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305263.

De-la-Torre:2025:PML

[lTS25] Gabriel Enrique De la Torre and Luis Santillán. Plastics and microplastics as limiting factors for conservation in Antarctica within the framework of the Antarctic Treaty System. Marine Policy, 181(??):??, November 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002787.

Lee:2022:EGG

[LTSL22] Hsiao-Chien Lee, Ying-Lee Tseng, Wei-Ying Sung, and Wen-Hong Liu. Exploring the gender gap in marine affairs in Taiwan. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005480.

Langford:2023:IIP

[LTW23] Alexandra Langford, Welem Turupadang, and Scott Waldron. Interventionist industry policy to support local valueadding: Evidence from the Eastern Indonesian seaweed industry. Marine Policy, 151(??):??, May 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300088X.

Lv:2020:BRI

[Lv20] Fang-Yuan Lv. Book review: Issues in Maritime Cyber Security, Joseph Direnzo, Nicole K. Drumhiller, Fried S. Roberts (Eds.) (2017), 568, ISBN: 978-1-78643-673-3. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20301639.

Langlet:2023:LUW

[LV23] Arne Langlet and Alice B. M. Vadrot. Not ‘undermining’ who? Unpacking the emerging BBNJ regime complex. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004195.

Li:2022:HWE

[LW22] Jing-Mei Li and Na Wang. How and to what extent is ecosystem services economic valuation used in coastal and marine management in China? Marine Policy, 138(??): ??, April 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000239.

Li:2023:DTP

[LWG23] Rongrong Li, Qiang Wang, and Yunfei Ge. Does trade protection undercut the green efficiency of the marine economy? A case study. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003974.

Leurs:2024:OCV

[LWH^{+}24] Guido Leurs, Matthew Bjerregaard Walsh, Alifa Bintha Haque, Claire Collins, Kerstin Glaus, Shoba Joe Kizhakudan, Muktha Menon, Efin Muttaqin, Benaya Simeon, Javier Tovar-Ávila, Laura L. Govers, Paddy A. Walker, and Kim J. Friedman. Opportunities and challenges in value chain analysis for sustainable management of sharks and rays. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001192.

Langford:2021:MCA

[LWSS21] Alexandra Langford, Scott Waldron, Sulfahri, and Hasnawati Saleh. Monitoring the COVID-19-affected Indonesian seaweed industry using remote sensing data. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000415.

Liu:2020:ECB

[LWX^{+}20] Baiqiong Liu, Jing Wang, Min Xu, Zaifeng Wang, and Lin Zhao. Evaluation of the comprehensive benefit of various marine exploitation activities in China. Marine Policy, 116 (??):??, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X1930315X.

Li:2023:FSD

[LWZ23] Fenghua Li, Leifan Wang, and Xiaoyu Zhou. Fulfiling the State duty to cooperate on the discharge of Fukushima nuclear contaminated water: Potential pathways. Marine Policy, 150(??):??, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000738.

Li:2024:CAB

[LX24] Jingchang Li and Wangwang Xing. A critical appraisal of the BBNJ agreement not to recognise the high seas decline as a common concern of humankind. Marine Policy, 163 (??):??, May 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24001295.

Lin:2022:IRS

[LXH^{+}22] Yanhong Lin, Xiongzhi Xue, Wenjia Hu, Qiuming Chen, Yang Zhang, Faming Huang, Jinliang Huang, and Chao Wang. Implementing regional sea use planning system in China: Evolution, drivers and experiences. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003451.

Li:2021:ECM

[LXSX21] Fuzhu Li, Wenxiu Xing, Meng Su, and Juan Xu. The evolution of China’s marine economic policy and the labor productivity growth momentum of marine economy and its three economic industries. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003882.

Liu:2021:ECM

[LXW^{+}21] Baiqiong Liu, Min Xu, Jing Wang, Zaifeng Wang, and Lin Zhao. Evaluation of China’s marine economic growth quality based on set pair analysis. Marine Policy, 126(??):??, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000154.

Li:2023:CSE

[LYHL23] Wei Li, Qi Yue, Heng Hu, and Liang Liu. Current status of ecological compensation for bohai sea reclamation and management countermeasures. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003111.

Lin:2022:STE

[LYL^{+}22] Ying Lin, Yiqun Yang, Peiliang Li, Cuicui Feng, Jieqiong Ding, Junyu Zhou, Qutu Jiang, and Guanqiong Ye. Spatialtemporal evaluation of marine ecological civilization of Zhejiang Province, China. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004462.

Liu:2025:EDM

[LYLZ25] Zifei Liu, Yuan Yuan, Fei Li, and Liangzhen Zang. An evaluation of degradation of marine capture fisheries in China. Marine Policy, 177(??):??, July 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000326.

Lili:2021:RAU

[LYQM21] Ding Lili, Yang Ying, Hu Qiuhui, and Liu Mengxi. Residents’ acceptance of using desalinated water in China based on the theory of planned behaviour (TPB). Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309398.

Li:2024:IER

[LZ24] Zhiwen Li and Bo Zhang. Internationalization of EIA rules in the BBNJ Agreement: Impediments and possible solutions. Marine Policy, 167(??):??, September 2024.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002434.

Liu:2023:TIF

[LZL23] Shiyu Liu, Wei Zeng, and Xiyao Li. Toward an integrated framework of ocean literacy: a Delphi-AHP approach. Marine Policy, 157(??):??, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003639.

Li:2025:CME

[LZL25] Fuzhu Li, Yuge Zhang, and Shuochen Luan. Can Marine Economic Development Demonstration Zones (MED-DZs) achieve a sustainable balance between economic development and ecological protection: Evidence from Chinese coastal cities. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001836.

Lopera:2023:OPI

[LZRC23] Luisa Lopera, Paula A. Zapata-Ramírez, and Yuley Cardona. Overview and potential implementation of a marine protected area network between Colombia, Nicaragua, and Jamaica. Marine Policy, 150(??):??, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004584.

Li:2024:UIM

[LZX24] Jingmei Li, Zhuqing Zhuang, and Zhihua Xu. Understanding the impact of marine oil spill on corporate performance for improving oil spill claims system. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002963.

Li:2022:CEG

[LZZ22] Hua Li, Li Zhu, and Mengyuan Zheng. Is the cruise enterprise giant also a giant of social responsibility for sustainable development? Marine Policy, 138(??):??, April 2022.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000562.

Ma:2020:ELA

[Ma20] Xuechan Ma. An economic and legal analysis of trade measures against illegal, unreported and unregulated fishing. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309455.

Marin:2024:DST

[MAA^{+}24] Charlotte Marin, Olugbenga M. Adewumi, Frank Asche, Taryn M. Garlock, Dadi M. Kristofersson, Kai Lorenzen, and Bixuan Yang. Does seafood trade enhance seafood availability in developing countries? The case of Nigeria. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000289.

Mensah:2025:CFS

[MAA^{+}25] Justice Mensah, John Oti Amoah, Edinam Bernice Amenumey, Precious A. D. Mattah, and Memuna Mawusi Mattah. The closed fishing season policy in Ghana: Lessons from the coastal fisherfolk. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400472X.

Mabon:2024:TWR

[Mab24] Leslie Mabon. Treated water releases from the Fukushima Dai’ichi nuclear power plant: an overview of the decisionmaking process and governing institutions. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X24001180.

Merten:2022:CIF

[MAG^{+}22] Wessley Merten, Richard Appeldoorn, Abby Grove, Alfonso Aguilar-Perera, Freddy Arocha, and Roberto Rivera. Condition of the international fisheries, catch and effort trends, and fishery data gaps for dolphinfish (Coryphaena hippurus) from 1950 to 2018 in the Western Central Atlantic Ocean. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002366.

Magnuson:2023:OSF

[Mag23] Briana Magnuson. One-size-fits-all? How north–south inequality challenges CITES’ ability to protect the scalloped hammerhead. Marine Policy, 158(??):??, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004001.

Mahadeo:2022:MSP

[Mah22] Sarah Mahadeo. Marine spatial planning in the Eastern Caribbean: Trends and progress. Marine Policy, 145(??): ??, November 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003244.

McClanahan:2022:FCR

[MAK22] T. R. McClanahan, M. K. Azali, and J. K. Kosgei. Fish catch responses to COVID-19 disease curfews dependent on compliance, fisheries management, and environmental contexts. Marine Policy, 144(??):??, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200286X.

Mallin:2021:CSL

[Mal21] Marc-Andrej Felix Mallin. Corrigendum to: From sealevel rise to seabed grabbing: the political economy of climate change in Kiribati [Mar. Policy 97 (2018) 244–252]. Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001202.

Meza-Arce:2020:UWS

[MAMCHP^{+}20] Mercedes I. Meza-Arce, Luis Malpica-Cruz, Mauricio E. Hoyos-Padilla, Francisco J. Mojica, María Concepción Arredondo-García, Claudia Leyva, Rebeca Zertuche-Chanes, and Omar Santana-Morales. Unraveling the white shark observation tourism at Guadalupe Island, Mexico: Actors, needs and sustainability. Marine Policy, 119(??): ??, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19307353.

Marcondes:2020:CGB

[Mar20a] Danilo Marcondes. Conservationist geopolitics: Brazilian foreign policy and the South Atlantic Whale Sanctuary. Marine Policy, 120(??):??, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301666.

Maxwell:2020:WEN

[MAR^{+}20b] Kimberley Maxwell, Shaun Awatere, Kelly Ratana, Kate Davies, and Caine Taiapa. He waka eke noa/we are all in the same boat: a framework for co-governance from Aotearoa, New Zealand. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308599.

Marghelis:2021:MDA

[Mar21a] Aris Marghelis. The maritime delimitation agreement between Greece and Italy of 9 June 2020: an analysis in the light of International Law, national interest and regional politics. Marine Policy, 126(??):??, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000130.

Martinson:2021:CSF

[Mar21b] Ryan Martinson. Catching sovereignty fish: Chinese fishers in the southern Spratlys. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2031023X.

Matthiasson:2020:SCR

[Mat20] Thorolfur Matthiasson. Splitting the catch, remuneration of Icelandic fishermen in a sea of change. Marine Policy, 117 (??):??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19307729.

Maynou:2022:SPF

[May22] Francesc Maynou. Sale price flexibilities of Mediterranean hake and red shrimp. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005157.

Moses:2021:WBC

[MB21] Jonathon W. Moses and Anne Margrethe Brigham. Whose benefit? A comparative perspective for the ISA. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001615.

Mannaart:2022:FLI

[MB22] Michael Mannaart and Arabelle Bentley. Fishing for litter: From the implementation of practical actions locally, to its spin-offs and the adoption of a new legally adopted waste type at continental scale, a success story. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003037.

Mbatha:2022:UPM

[Mba22] Philile Mbatha. Unravelling the perpetuated marginalization of customary livelihoods on the coast by plural and multi-level conservation governance systems. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001907.

McGee:2022:MPP

[MBB^{+}22] Gord McGee, Josie Byington, John Bones, Sally Cargill, Megan Dickinson, Kelly Wozniak, and Kylee A. Pawluk. Marine plan partnership for the North Pacific Coast: Engagement and communication with stakeholders and the public. Marine Policy, 142(??):??, August 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002244.

Mendez:2024:BBS

[MBC^{+}24] Cynthia Mendez, Danilo E. Bustamante, Martha S. Calderon, Cecilia Gauna, Leila Hayashi, Daniel Robledo, Claudia Tapia-Larios, Iona Campbell, Renato Westermeier, and Pedro Murúa. Biosecurity baseline for a sustainable development of seaweed aquaculture in Latin America. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004669.

Maini:2023:CVL

[MBDG23] Bani Maini, Jessica L. Blythe, Emily S. Darling, and Georgina G. Gurney. Charting the value and limits of other effective conservation measures (OECMs) for marine conservation: a Delphi study. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003979.

March:2024:SBE

[MBG^{+}24a] Antaya March, Michael Bennett, Matthew Germishuizen, Tegan Evans, and Pierre Failler. The status of Blue Economy development in Africa. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24002033.

Morrison:2024:OLS

[MBG^{+}24b] Tiffany H. Morrison, Jon Barnett, Georgina G. Gurney, Jacqueline Lau, Michele L. Barnes, Josh Cinner, Missaka Hettiarachchi, and Pip Cohen. Overcoming lock-in of science-policy responses to reef heating. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003786.

Mumtaz:2024:ARB

[MBL24] Ubaid Ullah Mumtaz, Paul Bergey, and Nicholas Letch. Assessing the role of blockchain technology for marine bunkering operations — a case study of task technology fit. Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004426.

Martin-Betancor:2024:EMS

[MBORGN24] M. Martín-Betancor, J. Osorio, A. Ruiz-García, and I. Nuez. Evaluation of maritime spatial planning for offshore wind energy in the Canary Islands: a comparative analysis. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000447.

Masalaci:2023:DDT

[MÇ23] Burcu Çelik Maşalacı and Erkan Çakır. Differences in deficiency types causing ship detentions at the Black Sea Region during the Covid-19 pandemic and pre-pandemic. Marine Policy, 151(??):??, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000805.

MacKay:2025:RAM

[MC25] Katharina MacKay and Richard Collins. Reorienting approaches to maritime boundary disputes: a case for hydrodiplomacy? Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004421.

Mallik:2022:ICL

[MCBN22] Abhijit Mallik, Puja Chakraborty, Shashi Bhushan, and Binaya Bhusan Nayak. Impact of COVID-19 lockdown on aquatic environment and fishing community: Boon or bane? Marine Policy, 141(??):??, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200135X.

McClanahan:2020:WCP

[McC20] Timothy Rice McClanahan. Wilderness and conservation policies needed to avoid a coral reef fisheries crisis. Marine Policy, 119(??):??, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307225.

McClanahan:2021:MRM

[McC21] Tim R. McClanahan. Marine reserve more sustainable than gear restriction in maintaining long-term coral reef fisheries yields. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000890.

McClanahan:2024:PPA

[McC24] Tim R. McClanahan. Perceptions of preparedness to address climate change threats in the western Indian Ocean. Marine Policy, 162(??):??, April 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000538.

Mogila:2021:BBG

[MCKZ21] Zbigniew Mogila, Dorota Ciolek, Jakub M. Kwiatkowski, and Jacek Zaucha. The Baltic blue growth — a countrylevel shift-share analysis. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004103.

Mallory:2022:COC

[MCL22] Tabitha Grace Mallory, Andrew Chubb, and Sallie Lau. China’s ocean culture and consciousness: Constructing a maritime great power narrative. Marine Policy, 144(??): ??, October 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002767.

Martinell:2020:CHS

[MCPS20] Daniel Peñalosa Martinell, Tim Cashion, Robert Parker, and U. Rashid Sumaila. Closing the high seas to fisheries: Possible impacts on aquaculture. Marine Policy, 115(??): ??, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19306931.

Menini:2022:PPS

[MCR22] Elisabetta Menini, Anindita Chakraborty, and Stephen E. Roady. Public participation in seabed mining in areas beyond national jurisdiction: Lessons learned from national regulators in the terrestrial mining sector. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003554.

Maestro:2020:AMP

[MCRPC20] María Maestro, Juan Adolfo Chica-Ruiz, and María Luisa Pérez-Cayeiro. Analysis of marine protected area management: the Marine Park of the Azores (Portugal). Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X1930819X.

Martins:2023:OES

[MCRR23] Maria Clara Iruzun Martins, Matt ID Carter, Sally Rouse, and Debbie JF Russell. Offshore energy structures in the North Sea: Past, present and future. Marine Policy, 152 (??):??, June 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001562.

Magliozzi:2025:DTA

[MCS^{+}25a] Chiara Magliozzi, Ana Cristina Cardoso, Mikhail O. Son, Yuriy Kvach, Bella Japoshvili, Kremena Stefanova, Güley Kurt, Elena Pantea, Cristina Tabarcea, Valentina Torodova, and Madona Varshanidze. Data towards assessing non-indigenous species introductions in the Black Sea. Marine Policy, 178(??):??, August 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001113.

Mombourquette:2025:IFP

[MCS25b] Daniel R. Mombourquette, Anthony Charles, and Robert L. Stephenson. Impacts of fishery policy on the distribution of access and community benefits. Marine Policy, 181(??): ??, November 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25002453.

Munim:2023:TPA

[MCTN23] Ziaul Haque Munim, Mohammed Mojahid Hossain Chowdhury, Hasan Mahbub Tusher, and Theo Notteboom. Towards a prioritization of alternative energy sources for sustainable shipping. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001069.

Mogila:2024:HIB

[MCTZ24] Zbigniew Mogila, Dorota Ciolek, Andrzej Toroj, and Jacek Zaucha. How important is the blue economy for regional development? — The case of Poland. Marine Policy, 168(??): ??, October 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003014.

Martin:2022:SDN

[MCW^{+}22a] Carol L. Martin, Belinda Curley, Kim Wolfenden, Marcel Green, and Natalie A. Moltschaniwskyj. The social dimension to the New South Wales Shark Management Strategy, 2015–2020, Australia: Lessons learned. Marine Policy, 141 (??):??, July 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001269.

Mayaud:2022:TNS

[MCW^{+}22b] Raphael Mayaud, Juliana Castrillon, Craig Wilson, David Peel, Joshua N. Smith, Greta Dalle Luche, Jenny Allen, and Susan Bengtson Nash. Traffic in a nursery: Ship strike risk from commercial vessels to migrating humpback whales (Megaptera novaeangliae) in a rapidly developing Australian urban embayment. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003797.

Moranta:2023:SPT

[MdCLL^{+}23] Joan Moranta, David Florido del Corral, Lucía López-López, Sílvia Gómez, Julio Agujetas, Marta Albo-Puigserver, Zoe W. Brent, Lamine Bathily, Rita Calvário, Marta Cavallé, Marta Coll, Annya Crane, Ericka C. L. de Oliveira, Irmak Ertör, Paloma Herrera-Racionero, Francisco Leitão, Lluís Miret-Pastor, Beatriz Patraca, Miquel Ortega, José J. Pascual-Fernández, Jaime Ramón-Bruquetas, Jorge Sáez-Jiménez, Abdoulaye Seck-Papalaye, and Sebastián Villasante. Sowing potential transformative changes in the fishing and agrifood systems — conference report. Marine Policy, 154(??):??, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002002.

McClanahan:2023:MEF

[MDG^{+}23] Timothy R. McClanahan, Stephanie D’Agata, N. A. J. Graham, Maxwell Azali Kodia, and Joseph M. Maina. Multivariate environment–fish biomass model informs sustainability and lost income in Indian Ocean coral reefs. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23001173.

Mendenhall:2022:DDP

[MDJ^{+}22] Elizabeth Mendenhall, Elizabeth De Santo, Mathias Jankila, Elizabeth Nyman, and Rachel Tiller. Direction, not detail: Progress towards consensus at the fourth intergovernmental conference on biodiversity beyond national jurisdiction. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003566.

Mansour:2021:GMT

[MDLA21] Shawky Mansour, Stephen Darby, Julian Leyland, and Peter M. Atkinson. Geospatial modelling of tropical cyclone risk along the northeast coast of Oman: Marine hazard mitigation and management policies. Marine Policy, 129(??): ??, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100155X.

Murphy:2021:UMM

[MEG^{+}21] Robert Murphy, Austin Estabrooks, John Gauvin, Steven Gray, Anita C. Kroska, Nathan Wolf, and Bradley P. Harris. Using mental models to quantify linear and nonlinear relationships in complex fishery systems. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003067.

Mendenhall:2023:MMW

[Men23] Elizabeth Mendenhall. Making the most of what we already have: Activating UNCLOS to combat marine plastic pollution. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003196.

Maxwell:2020:KPG

[MF20] R. J. Maxwell and R. Filgueira. Key players in the Grieg NL Placentia Bay Atlantic Salmon Aquaculture Project: a social network analysis. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19305184.

March:2022:SSF

[MF22] Antaya March and Pierre Failler. Small-scale fisheries development in Africa: Lessons learned and best practices for enhancing food security and livelihoods. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005364.

Monios:2023:PAP

[MF23] Jason Monios and Laurent Fedi. The principal-agent problem in hierarchical policy making: a failure of policy to support the transition to LNG as an alternative shipping fuel. Marine Policy, 157(??):??, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003792.

McKinley:2025:SVD

[MF25] Emma McKinley and Kathryn Fradera. Shouting into the void: Democratising ocean literacy through integrating process literacy. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001460.

Mackelworth:2024:EBS

[MFA^{+}24] P. Mackelworth, C. M. Fortuna, M. Antoninić, D. Holcer, D. Abdul Malak, K. Attia, M. Bricelj, F. Guerquin, M. Marković, E. Nunes, C. Perez-Valverde, E. Ramieri, I. Stojanović, L. Tunesi, and J. McGowan. Ecologically and Biologically Significant Areas (EBSAs) as an enabling mechanism for transboundary marine spatial planning. Marine Policy, 166(??):??, August 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400229X.

Matorres:2025:IFV

[MFAH25] Dane Erlo Matorres, Michael Fabinyi, and Rosalie Arcala-Hall. IUU fishing in the value chains of live grouper and squid exports: a case study of Palawan, Philippines. Marine Policy, 181(??):??, November 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002398.

Maskaeva:2024:ASE

[MFC^{+}24] Asiya Maskaeva, Pierre Failler, Honita Cowaloosur, Philippe Lallemand, and Jerry Mang’ena. Assessment of socioeconomic and ecosystem services of the blue economy in Tanzania using the UNECA’s Blue Economy Valuation Toolkit. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004530.

Murray:2023:SCM

[MFF^{+}23] Grant D. Murray, Robin Fail, Luke Fairbanks, Lisa M. Campbell, Linda D’Anna, and Joshua Stoll. Seafood consumption and the management of shellfish aquaculture. Marine Policy, 150(??):??, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000611.

Madarcos:2022:DMS

[MFG^{+}22] K. Madarcos, M. Fortnam, L. Gajardo, T. Chaigneau, RJ. Manucan, G. Cadigal, J. Matulac, L. Creencia, B. Gonzales, and L. Evans. Doing marine spatial zoning in coastal marine tropics: Palawan’s Environmental Critical Areas Network (ECAN). Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002548.

Marshall:2023:CET

[MFG^{+}23] Rachel C. Marshall, Jonathan A. D. Fisher, Paul Gatti, Erin H. Carruthers, Dominique Robert, and Arnault Le Bris. Combining electronic tag data and fish harvesters’ knowledge to identify ocean management concerns. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004407.

Mulalap:2020:TKB

[MFH^{+}20] Clement Yow Mulalap, Tekau Frere, Elise Huffer, Edvard Hviding, Kenneth Paul, Anita Smith, and Marjo K. Vierros. Traditional knowledge and the BBNJ instrument. Marine Policy, 122(??):??, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308620.

McAteer:2021:LMO

[MFM21] Benedict McAteer, Wesley Flannery, and Brendan Murtagh. Linking the motivations and outcomes of volunteers to understand participation in marine community science. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20310265.

Mamede:2023:HMP

[MFRQ23] Renato Mamede, Rosa Freitas, Ana Maria Rodrigues, and Victor Quintino. Habitat mapping of Portugal’s northwestern continental shelf and its application to implement the European Marine Strategy Framework Directive. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002981.

Magalhaes:2022:OPS

[MFT^{+}22] Karine Matos Magalhães, José Souto Rosa Filho, Carlos Eduardo Peres Teixeira, Clemente Coelho-Jr, Maria Cecilia Santana Lima, Ana Maria Costa Souza, and Marcelo O. Soares. Oil and plastic spill: 2021 as another challenging year for marine conservation in the South Atlantic Ocean. Marine Policy, 140(??):??, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001233.

Maruf:2022:TAU

[MG22] Maruf and Warwick Gullett. Tackling anthropogenic underwater noise through the Convention on Biological Diversity: Progress and future development. Marine Policy, 146(??): ??, December 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003402.

Magris:2025:OCI

[MG25] Rafael A. Magris and Leandra R. Gonçalves. Overcoming challenges and identifying opportunities for marine protected area planning: Lessons from Brazil. Marine Policy, 180(??):??, October 2025. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25002131.

Michel:2024:PAM

[MGBS24] Mathilde Michel, Benjamin Guichard, Julie Béesau, and Flore Samaran. Passive acoustic monitoring for assessing marine mammals population in European waters: Workshop conclusions and perspectives. Marine Policy, 160(??): ??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2300516X.

Mathis:2022:ROP

[MGDJ22] Jennifer Estelle Mathis, Marianne Carliez Gillet, Henri Disselkoen, and Jenna R. Jambeck. Reducing ocean plastic pollution: Locally led initiatives catalyzing change in South and Southeast Asia. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22001749.

Mileski:2020:HTC

[MGF20] Joan P. Mileski, Cassia Bomer Galvao, and Zaida Denise Forester. Human trafficking in the commercial fishing industry: a multiple case study analysis. Marine Policy, 116 (??):??, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19300533.

Miran:2025:ECI

[MGK^{+}25] Bülent Miran, Harun Güçlüsoy, Zafer Kızılkaya, Tolunay Tabak, and Kayhan Güçeli. The effects of COVID-19 on the input usage and revenue generation performance of the EU fishery sector. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400558X.

Morris:2025:WIC

[MGNP25] Corey J. Morris, John M. Green, Khanh Q. Nguyen, and Curtis Pennell. Will indigenous co-management improve the outcome of the Gilbert Bay, Labrador MPA: Summary of the past and hopes for the future. Marine Policy, 180(??): ??, October 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25001952.

Mouton:2025:SPI

[MGPR^{+}25] Théophile L. Mouton, Adriana Gonzalez-Pestana, Christoph A. Rohner, Ryan Charles, Emiliano García-Rodríguez, Peter M. Kyne, Amanda Batlle-Morera, Giuseppe Notarbartolo di Sciara, Asia O. Armstrong, Enzo Acuña, Joanna Alfaro-Shigueto, Randall Arauz, Cristopher G. Avalos-Castillo, Ely Augustinus, Sandra Bessudo, Enrique Barraza, Carlos Bustamante, Elpis J. Chávez, Eduardo Ramon Espinoza, Mario Espinoza, Ana Hacohen-Domené, Alex R. Hearn, Grettel M. Hernández, Felipe Galván-Magaña, José A. Gonzalez-Leiva, James T. Ketchum, Felipe Ladino, Frida Lara-Lizardi, Jorge Manuel Morales-Saldaña, Naití Morales Serrano, Jeffry Madrigal-Mesén, Paola A. Mejía-Falla, Andrés F. Navia, Gabriela M. Ochoa, Marta D. Palacios, César R. Peñaherrera-Palma, Francisco Polanco-Vásquez, Yehudi Rodríguez-Arriatti, Luz E. Saldaña-Ruiz, Oscar Sosa-Nishizaki, Javier Tovar-Ávila, Ángel J. Vega, Ximena Velez-Zuazo, Melany Villate-Moreno, Ilena Zanella, and Rima W. Jabado. Shortfalls in the protection of Important Shark and Ray Areas undermine shark conservation efforts in the Central and South American Pacific. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004482.

Macusi:2021:VDS

[MGS21] Edison D. Macusi, Rollan C. Geronimo, and Mudjekeewis D. Santos. Vulnerability drivers for small pelagics and milkfish aquaculture value chain determined through online participatory approach. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003213.

McQuaw:2020:WCM

[MH20a] Kristin McQuaw and Ray Hilborn. Why are catches in mixed fisheries well below TAC? Marine Policy, 117(??):

??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19300909.

Muallil:2020:PRA

[MH20b] Richard N. Muallil and Mohammad Gafor N. Hapid. Preliminary report on an artisanal fishery for thresher sharks (Alopias spp) in Tawi-Tawi, Southern Philippines. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304713.

Mendenhall:2023:BAL

[MH23] Elizabeth Mendenhall and Kahlil Hassanali. The BBNJ agreement and liability. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000763.

Mendenhall:2024:WBA

[MH24] Elizabeth Mendenhall and Rebecca R. Helm. When the “best available science” is not good enough: the need for supporting scientific research in the United Nations treaty to protect biodiversity beyond national jurisdiction. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004736.

Mankad:2025:PNM

[MHC^{+}25] Aditi Mankad, Elizabeth V. Hobman, Matt Curnock, Michelle Dyer, and Petina Pert. Perceived need for management interventions in the great barrier reef. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25002118.

Murphy:2020:CSP

[MHEW20] Robert Murphy, Bradley Harris, Austin Estabrooks, and Nathan Wolf. Capturing stakeholder perspectives through a collaboration with a commercial fishing cooperative.

Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19308425.

Massa:2025:ESA

[MHF^{+}25] Fabio Massa, Houssam Hamza, Davide Fezzardi, Linda Fourdain, Marianna Cavallo, Fabrizio Caruso, and John A. Theodorou. Enhancing social acceptability of aquaculture: Stakeholder engagement in guidelines development for the Mediterranean and Black Sea. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25001125.

Mendenhall:2020:CCI

[MHN^{+}20] Elizabeth Mendenhall, Cullen Hendrix, Elizabeth Nyman, Paige M. Roberts, John Robison Hoopes, James R. Watson, Vicky W. Y. Lam, and U. Rashid Sumaila. Climate change increases the risk of fisheries conflict. Marine Policy, 117 (??):??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19304622.

Matera:2025:CES

[MHS^{+}25a] Jaime Matera, Ellen Hines, Maritza Sepúlveda, María José Pérez-Alvarez, Fernanda Barilari, Stefan Gelcich, Rodrigo A. Estévez, Rebecca Lent, and Marcelo A. San Martín. Corrigendum to “Evaluating socioecological and governance characteristics associated with marine mammal bycatch in the artisanal fisheries of central Chile” [Mar. Policy 180 (2025) 106747]. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25002325.  See

[MHS^{+}25b].

Matera:2025:ESG

[MHS^{+}25b] Jaime Matera, Ellen Hines, Maritza Sepúlveda, María José Pérez-Alvarez, Fernanda Barilari, Stefan Gelcich, Rodrigo A. Estévez, Rebecca Lent, and Marcelo A. San Martín. Evaluating socioecological and governance characteristics associated with marine mammal bycatch in the artisanal fisheries of central Chile. Marine Policy, 180(??): ??, October 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25001629. See cor-

rigendum [MHS^{+}25a].

Mameno:2024:WLI

[MHT^{+}24] Kota Mameno, Chia-Hsuan Hsu, Takahiro Tsuge, Ayumi Onuma, and Takahiro Kubo. Who is likely to have illegal eel products after the market regulations? A best-worst scaling approach. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003713.

Muallil:2024:RAS

[MIH^{+}24] Richard N. Muallil, Roger A. Irilis, Argamar A. Habibuddin, Nur aisa S. Abdulmajid, Jaro O. Ajik, and Ronaldo A. Ancheta. A rapid assessment of the status, trends and challenges in small-scale commercial sardine fisheries in the Sulu Archipelago, southern Philippines. Marine Policy, 160(??): ??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004980.

Miller:2022:DTI

[Mil22] Karly Marie Miller. Disentangling tourism impacts on small-scale fishing pressure. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000070.

Marega-Imamura:2020:SCN

[MIMS^{+}20] Marcela Marega-Imamura, Fernanda Michalski, Kátia Silva, Alexandre Schiavetti, Yvonnick Le Pendu, and Leonardo de Carvalho Oliveira. Scientific collaboration networks in research on human threats to cetaceans in Brazil. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18305876.

Min:2024:COS

[Min24] Li Min. China’s overseas seafarer dispatching agencies: an analytical consideration of the legal issues in perspective. Marine Policy, 167(??):??, September 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002501.

Mazur:2020:EIF

[MJ20] Mackenzie Dale Mazur and Teresa R. Johnson. Effects of increases in fishery resource abundance on conservation compliance. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308630.

Maya-Jariego:2023:INS

[MJ23]  Isidro Maya-Jariego.  Informal norms sustain non-

compliance and poaching among spearfishermen. A response to Tudela et al. (2022). Marine Policy, 148(??): ??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22005231.  See

[TGM22].

Maya-Jariego:2021:PBM

[MJMAA21] Isidro Maya-Jariego, Inmaculada Martínez-Alba, and Deniza Alieva. “Plenty of black money”: Netnography of illegal recreational underwater fishing in southern Spain. Marine Policy, 126(??):??, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100021X. See reply [SCD21].

Maya-Jariego:2022:RPR

[MJMAA22] Maya-Jariego, Martínez-Alba, and D. Alieva. Responsible practice of recreational spearfishing is needed. A response to sbragaglia et al. (2021). Marine Policy, 139(??): ??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000690. See [SCD21].

Mulder:2024:WSM

[MK24] Thomas Mulder and Natalie Klein. WHO at sea: Maritime dimensions of the proposed pandemic treaty and changes to the International Health Regulations 2005. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004517.

Marlianingrum:2021:VHQ

[MKAF21] Peggy Ratna Marlianingrum, Tridoyo Kusumastanto, Luky Adrianto, and Achmad Fahrudin. Valuing habitat quality for managing mangrove ecosystem services in coastal Tangerang District, Indonesia. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003584.

Montenero:2021:QQD

[MKB21] Kelly Montenero, Chris Kelble, and Kathy Broughton. A quantitative and qualitative decision-making process for selecting indicators to track ecosystem condition. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001007.

Mehault:2020:EPR

[MKM^{+}20] Sonia Méhault, Dorothée Kopp, Marie Morfin, Jean Pierre Cuillandre, Pascal Larnaud, Fabien Morandeau, Jean Philippe Vacherot, and Thomas Rimaud. Escapement patterns of red mullet, sole, horse mackerel and hake facing two trawl selective devices. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18307930.

Maltby:2023:BEC

[MKM23] Katherine M. Maltby, Sabrina Kerin, and Katherine E. Mills. Barriers and enablers of climate adaptation in fisheries: Insights from northeast US fishing communities. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003785.

McShane:2021:AAF

[MKS21] Paul McShane, Ian Knuckey, and Sevaly Sen. Access and allocation in fisheries: the Australian experience. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003134.

Molino:2020:SDS

[MKSG20] Grace D. Molino, Melissa A. Kenney, and Ariana E. Sutton-Grier. Stakeholder-defined scientific needs for coastal resilience decisions in the Northeast U.S. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19306955.

Mangubhai:2021:EGI

[ML21] Sangeeta Mangubhai and Sarah Lawless. Exploring gender inclusion in small-scale fisheries management and development in Melanesia. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309337.

Ma:2022:APC

[ML22] Mingfei Ma and Zhe Li. Amendment process and comments of the provisions of underwater cultural heritage in the Draft Exploitation Regulations in the area. Marine Policy, 140(??):??, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001282.

Melnychuk:2024:RDF

[MLA^{+}24] Michael C. Melnychuk, Samantha Lees, Megan Atcheson, Polly Burns, Catherine Longo, Lauren Koerner, and Beth Polidoro. Rates and drivers of fishery certification and withdrawal from the MSC program. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003415.

Montecalvo:2023:OPS

[MLAS23] Isabella Montecalvo, Philippe Le Billon, Chris Arsenault, and Milko Schvartzman. Ocean predators: Squids, Chinese fleets and the geopolitics of high seas fishing. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23001112.

Matovu:2025:EBR

[MLB^{+}25] Baker Matovu, Isaac Lukambagire, Raimund Bleischwitz, Etta Linda A, Meltem Alkoyak-Yildiz, Ammu S. Suresh, and Anusree S. An evidence-based review of the prerequisite interventions for women’s inclusion in the blue economy in the global south: a case study of India. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004767.

Maia:2024:HBA

[MLE^{+}24] Michele Cristina Maia, Gabriela Lamego, Carla I. Elliff, Jana M. Del Favero, Juliana Leonel, and Catarina R. Marcolin. Harassment and bullying aboard: Impacts of gender inequality on ocean professionals. Marine Policy, 160(??): ??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004797.

Madon:2022:PAD

[MLJ^{+}22] Bénédicte Madon, Damien Le Guyader, Jean-Luc Jung, Benjamin De Montgolfier, Pascal Jean Lopez, Eric Foulquier, Laurent Bouveret, and Iwan Le Berre. Pairing AIS data and underwater topography to assess maritime traffic pressures on cetaceans: Case study in the Guadeloupean waters of the Agoa sanctuary. Marine Policy, 143(??): ??, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2200207X.

Mu:2023:ECC

[MLM23] Manning Mu, Mengxi Liu, and Wen Ma. Evaluating Chinese consumers’ purchase intentions for daily chemical products without plastic microbeads using an extended theory of planned behavior model. Marine Policy, 154(??):??, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001987.

Macusi:2021:FIC

[MLMD21] Edison D. Macusi, Arianne Kay O. Liguez, Erna S. Macusi, and Larry N. Digal. Factors influencing catch and support for the implementation of the closed fishing season in Davao Gulf, Philippines. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001895.

Mantziaris:2021:FRB

[MLVT21] Stamatis Mantziaris, Angelos Liontakis, George Valakas, and Irene Tzouramani. Family-run or business-oriented fisheries? Integrating socioeconomic and environmental aspects to assess the societal impact. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002025.

Myers:2020:REU

[MM20] Hannah J. Myers and Michael J. Moore. Reducing effort in the U.S. American lobster (Homarus americanus) fishery to prevent North Atlantic right whale (Eubalaena glacialis) entanglements may support higher profits and long-term sustainability. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309157.

Macedo:2021:RGB

[MM21a] Heitor Schulz Macedo and Rodrigo Pereira Medeiros. Rethinking governance in a Brazilian multiple-use marine protected area. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305529.

Motoori:2021:RSS

[MM21b] Ran Motoori and Benjamin C. McLellan. Resource security strategies and preferences for deep ocean mining from a community survey in Japan. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001226.

McClenachan:2022:TWC

[MM22] Loren McClenachan and Allie Moulton. Transitions from wild-caught fisheries to shellfish and seaweed aquaculture increase gender equity in Maine. Marine Policy, 146(??): ??, December 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003591.

Mesquita:2023:IRF

[MM23] Isabeli Cristina Gomes Mesquita and Rodrigo Pereira Medeiros. Integrating research and fishing extension approaches to engage small-scale fishers in the participatory evaluation and voluntary use of bycatch reduction devices. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001264.

McGuire:2025:SSG

[MMAP25] Gina McGuire, Alexander Mawyer, James J. Akau, and Noelani Puniwai. Seaweeds and sovereignty: Governance gaps and self-determination within limu stewardship in Hawai’i. Marine Policy, 181(??):??, November 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002441.

Marchessaux:2023:IBC

[MMB^{+}23] Guillaume Marchessaux, Maria Cristina Mangano, Sergio Bizzarri, Charaf M’Rabet, Elena Principato, Nicola Lago, Dimitri Veyssiere, Marie Garrido, Steven B. Scyphers, and Gianluca Sarà. Invasive blue crabs and small-scale fisheries in the Mediterranean sea: Local ecological knowledge, impacts and future management. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22005085.

Macusi:2022:WPS

[MMC^{+}22] Edison D. Macusi, Erna S. Macusi, Charlen Mae G. Canales, Anna Barboza, and Larry N. Digal. Women’s participation and support for the implementation of the closed fishing season in Davao Gulf, Philippines. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001804.

Masi:2025:ESI

[MMF^{+}25] M. Masi, E. S. Marrocco, G. Fusco, F. Adinolfi, and Y. Vecchio. The economic sustainability of Italian aquaculture farms: Who has the potential to start the blue transition? Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004780.

McRuer:2025:HOR

[MMG^{+}25] Jen McRuer, Emma McKinley, Diz Glithero, Ronaldo Christofoletti, and Diana Payne. Human-ocean relationships: Exploring alignment and collaboration between ocean literacy research and marine conservation. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004184.

Murillas-Maza:2021:EPS

[MMGAA^{+}21] Arantza Murillas-Maza, Antonio García-Allut, Martin Aranda, César Pazos, Ane Iriondo, and Miriam Montero. Enhancing the participation of small-scale fishing sector in decision-making: Good practices from Atlantic in EU. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001445.

Mudege:2022:ICG

[MMK^{+}22] Netsayi Noris Mudege, Catherine Mawia Mwema, Keagan Kakwasha, Andrew Chisopo, Chikondi Manyungwa-Pasani, Lisungu Banda, Emmanuel Kaunda, and Pamela Marinda. The impacts of Covid-19 on gender dynamics and power relations among men and women involved in cross border fish trade in Zambia and Malawi. Marine Policy, 146(??): ??, December 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003694.

Macusi:2023:WLS

[MMK^{+}23] Edison D. Macusi, Imee S. Maynawang, Robert E. Katikiro, Jee Grace Suyo, and Erna S. Macusi. Women lead smallscale fisheries to sustainability in Surigao del Sur, Philippines. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003019.

Macusi:2022:ICF

[MMM^{+}22] Edison D. Macusi, Ilah Dianne G. Morales, Erna S. Macusi, Aijelyn Pancho, and Larry N. Digal. Impact of closed fishing season on supply, catch, price and the fisheries market chain. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000550.

McNelly:2025:EGY

[MMM^{+}25] Carter McNelly, Élizabeth Melis, Hali Moreland, Noémie Roy, Nadia Dalili, Emily Wells, Caleigh Delle Palme, and Don Carruthers Den Hoed. Ecological grief: Young professionals transformed by and transforming conservation work. Marine Policy, 180(??):??, October 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003609.

McManus:2022:ICD

[MMNC22] Catherine McManus, Alistair McIlgorm, Rachel Nichols, and Andrew Cooper. An initial consideration of data availability issues in downscaling ocean accounting to inform sustainable aquaculture development: the example of Clew Bay, Ireland. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003335.

Morse:2024:PSM

[MMO^{+}24] Molly Morse, Douglas McCauley, Sara Orofino, Keenan Stears, Samantha Mladjov, Jenn Caselle, Tyler Clavelle, and Ryan Freedman. Preferential selection of marine protected areas by the recreational scuba diving industry. Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004414.

Matoso:2024:WKD

[MMS24] Thomaz Bosquetto Matoso, Gustavo Goulart Moreira Moura, and Thiago Zagonel Serafini. A wave knowledge description of a surf break in Southern Brazil. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2300461X.

Mendez-Medina:2020:ACD

[MMSB^{+}20] Crisol Méndez-Medina, Birgit Schmook, Xavier Basurto, Stuart Fulton, and Alejandro Espinoza-Tenorio. Achieving coordination of decentralized fisheries governance through collaborative arrangements: a case study of the Sian Ka’an Biosphere Reserve in Mexico. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18306377.

Mirera:2023:FPD

[MMW^{+}23] David Oersted Mirera, Esther Wairimu Magondu, Miriam Wambui Wainaina, Brendan Muli, Douglas Okemwa, Rose Angulu, Irene Heba, and Hellen Moyoni. Fish preference at different value chain levels and implications for management of mariculture. Marine Policy, 157(??):??, November 2023.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003780.

Marques:2021:ECM

[MN21] Pedro Henrique Dias Marques and Sebastião Cerqueira Neto. Environmental conflicts and management of marine resources in the Cassurubá Complex, Bahia, Brazil. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000452.

Mazloumi:2023:PCR

[MN23] Nastaran Mazloumi and Simon Nicol. Preparing for climate related impacts: Knowledge assessment of the life history of key exploited fish species managed in a climate hotspot. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004286.

Macusi:2024:FWB

[MNGM24] Edison D. Macusi, Ivy M. Nallos, Maria C. Galveia, and Erna S. Macusi. Fishers’ well-being as affected by anthropogenic stressors in Surigao del Sur, Philippines. Marine Policy, 162(??):??, April 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000472.

Munim:2025:KDC

[MNHS25] Ziaul Haque Munim, Theo Notteboom, Hercules Haralambides, and Halvor Schøyen. Key determinants for the commercial feasibility of maritime autonomous surface ships (MASS). Marine Policy, 172(??):??, February 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004822.

Meeanan:2024:UVS

[MNS^{+}24] Chonlada Meeanan, Pavarot Noranarttragoon, Piyachoke Sinanun, Wirat Sanitmajjaro, Yuki Takahashi, Methee Kaewnern, and Takashi Fritz Matsuishi. Using vessel surveillance data to estimate spatiotemporal patterns in the short mackerel purse-seine fishery: Implications for time–area closure management in Thai waters. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003804.

Mohan:2023:IND

[Moh23] Preeya S. Mohan. Implementing nationally determined contributions under the Paris Agreement: an assessment of ocean-based climate action in Caribbean Small Island Developing States. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003202.

Molenaar:2024:CAO

[Mol24] Erik J. Molenaar. The Central Arctic Ocean Fisheries Agreement and Arctic indigenous peoples. Marine Policy, 164(??):??, June 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001581.

Morseletto:2020:NFP

[Mor20] Piero Morseletto. A new framework for policy evaluation: Targets, marine litter, Italy and the Marine Strategy Framework Directive. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304002.

Mota:2020:FHC

[Mot20] Rui Pedro Mota. Fishery harvest control rules from bioeconomic optimization. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19308140. See corrigen-

dum [Mot22].

Mota:2022:CLF

[Mot22] Rui Pedro Mota. Corrigendum to ‘Fishery harvest control rules from bioeconomic optimization’ [Mar. Policy Vol. 115, May 2020, 103865]. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004498. See [Mot20].

Magliozzi:2023:SPH

[MPD^{+}23] Chiara Magliozzi, Marco Palma, Jean-Noël Druon, Andreas Palialexis, McQuatters-Gollop Abigail, Varkitzi Ioanna, González-Quirós Rafael, Gorokhova Elena, Heyden Birgit, Boicenco Laura, and Artigas Luis Felipe. Status of pelagic habitats within the EU-Marine Strategy Framework Directive: Proposals for improving consistency and representativeness of the assessment. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22005188.

McLennan:2023:NMS

[MPH^{+}23] Shaun McLennan, Beth Polidoro, Tim Huntington, Rod Cappell, Victor Restrepo, Karin Mundnich, and Ernesto Jardim. The new Marine Stewardship Council requirements to improve ghost gear management: Insights from the policy development process. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003731.

Marques:2023:CEI

[MPL^{+}23] Crístofer H. Marques, Paula C. Pereda, Andrea Lucchesi, Ramiro F. Ramos, Olav Fiksdahl, Luiz F. Assis, Newton N. Pereira, and Jean-David Caprace. Cost and environmental impact assessment of mandatory speed reduction of maritime fleets. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003815.

Muflikh:2024:BSM

[MPN^{+}24] Yanti N. Muflikh, Risti Permani, Nunung Nuryartono, Syamsul H. Pasaribu, Boedi S. Julianto, Fikri Sjahruddin, Nunung Kusnadi, and Ammar A. Aziz. Building stakeholders’ mutual understanding of seaweed sustainability in Indonesia: a group model building approach. Marine Policy, 168(??):??, October 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002811.

Miret-Pastor:2020:TTD

[MPSF20] Lluís Miret-Pastor, Kristina Svels, and Richard Freeman. Towards territorial development in fisheries areas: a typology of projects funded by Fisheries Local Action Groups. Marine Policy, 119(??):??, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20304759.

Maas:2024:SCB

[MR24] Pascale Maas and Sandra Rousseau. A social cost-benefit analysis of shellfish beds on the Dutch coast as a naturebased measure against climate change. Marine Policy, 163 (??):??, May 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000952.

Maxwell:2020:NTM

[MRD^{+}20] Kimberley H. Maxwell, Kelly Ratana, Kathryn K. Davies, Caine Taiapa, and Shaun Awatere. Navigating towards marine co-management with Indigenous communities onboard the Waka-Taurua. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301605.

Mrdalo:2023:URF

[Mrd23] Zvonimir P. D. Mrdalo. User rights in the fisheries and salmon-farming of the Faroe Islands. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003743.

Ma:2023:RFB

[MRKM23] Deqiang Ma, Jonathan Rhodes, Carissa J. Klein, and Martine Maron. Redistribution of fishery benefits among commercial and recreational fishers caused by offsetting. Marine Policy, 158(??):??, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004141.

Mossop:2020:ADR

[MS20] Joanna Mossop and Clive Schofield. Adjacency and due regard: the role of coastal States in the BBNJ treaty. Marine Policy, 122(??):??, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307511.

Miranda:2021:WDS

[MS21] Fabiola V. Miranda and Wolfgang B. Stotz. Why do state interventions in artisanal fisheries often fail in Chile? Misalignment to the way of life of fishers. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003043.

Meyer:2025:GBO

[MSB^{+}25] Abigail Meyer, Marinez Scherer, Jamani Balderamos, Nidia Chacon, Brooke Dixon, Andrew Estep, Shaistha Mohamed, Peter Menzies, Fathimath Nistharan, Matthew Paufve, Adriano Quintela, and William J. McClintock. Genderbased ocean uses and values: Implications for marine spatial planning. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2500106X.

Masud:2022:CMA

[MSBA22] Muhammad Mehedi Masud, Sharifah Muhairah Shahabudin, Angathevar Baskaran, and Rulia Akhtar. Comanagement approach to sustainable management of marine protected areas: the case of Malaysia. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000574.

Munoz:2020:CWC

[MSGAR20] Pablo Durán Muñoz, Mar Sacau, Ana García-Alegre, and Esther Román. Cold-water corals and deep-sea sponges by-catch mitigation: Dealing with groundfish survey data in the management of the northwest Atlantic Ocean high seas fisheries. Marine Policy, 116(??):??, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308066.

Mwanyoka:2025:AFS

[MSHK25] Iddi Ramadhani Mwanyoka, Mohamed Khamis Said, Kalista Peter Higini, and Abiud Lucas Kaswamila. Artisanal fishers and seaweed farmers’ engagement in Blue Economy in Zanzibar. Marine Policy, 174(??):??, April 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25000028.

Martin:2023:CNP

[MSO23] Angela Helen Martin, Maike Iris Esther Scheffold, and Bethan Christine O’Leary. Changing the narrative and perspective surrounding marine fish. Marine Policy, 156(??): ??, October 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003391.

Markiewicz-Stanny:2023:PGO

[MSP23] Joanna Markiewicz-Stanny and Dorota Pyć. Protection of the global ocean rights. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200519X.

Markiewicz-Stanny:2022:IMI

[MSSS22] Joanna Markiewicz-Stanny and Marta Szuniewicz-Stepień. Irregular migration and its impact on the marine environment: Legal aspects. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002639.

Moreno:2023:JFP

[MSZ^{+}23] Gala Moreno, Joaquín Salvador, Iker Zudaire, Jefferson Murua, Josep Lluís Pelegrí, Jon Uranga, Hilario Murua, Maitane Grande, Josu Santiago, and Victor Restrepo. The Jelly-FAD: a paradigm shift in the design of biodegradable Fish Aggregating Devices. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003992.

Molony:2020:MSS

[MT20] Brett W. Molony and Damian P. Thomson. Managing the super-shark: Recommendations to improve communication and debate about chondrichthyans. Marine Policy, 118(??): ??, August 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20301135.

McMullen:2023:MLS

[MTC^{+}23] Karly McMullen, Ana Tirapé, Paola Calle, Jessica Vandenberg, Omar Alvarado-Cadena, Yoshitaka Ota, Gustavo A. Domínguez, and Juan José Alava. Marine litter and social inequities entangle Ecuadorian mangrove communities: Perceptions of plastic pollution and well-being concerns in Puerto Hondo and Isla Santay, Ecuador. Marine Policy, 157(??):??, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003901.

Mendenhall:2023:SRS

[MTN23] Elizabeth Mendenhall, Rachel Tiller, and Elizabeth Nyman. The ship has reached the shore: the final session of the ‘Biodiversity Beyond National Jurisdiction’ negotiations. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002130.

Misund:2024:OCN

[MTS^{+}24] Andreas Misund, Trine Thorvaldsen, Andrea Viken Strand, Thea Lurås Oftebro, and Stine Wiborg Dahle. Opportunities and challenges in new production systems for salmon farming in Norway — industry perspective. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003920.

Mukherjee:2021:BRF

[Muk21] Zinnia Mukherjee. Behavioral response of fishers to hypoxia and the distributional impact on harvest. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003432.

Macpherson:2021:HAR

[MUR^{+}21a] Elizabeth Macpherson, Stephen C. Urlich, Hamish G. Rennie, Adrienne Paul, Karen Fisher, Laura Braid, Jill Banwell, Julia Torres Ventura, and Eric Jorgensen. ‘Hooks’ and ‘Anchors’ for relational ecosystem-based marine management. Marine Policy, 130(??):??, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100172X.

Murray:2021:GAT

[Mur21b] Ruth Murray. A governance analysis of three MPAs in Belize: Conservation objectives compromised by tourism development priorities? Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308897.

Marco:2021:EMR

[MVR21a] Jorge Marco, Diego Valderrama, and Mario Rueda. Evaluating management reforms in a Colombian shrimp fishery using fisheries performance indicators. Marine Policy, 125 (??):??, March 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309040.

Marco:2021:TBL

[MVR21b] Jorge Marco, Diego Valderrama, and Mario Rueda. Triple bottom line assessment for the historically underperforming Colombian queen conch fishery. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309039.

Martinell:2021:SEE

[MVSP^{+}21] Daniel Peñalosa Martinell, Francisco Javier Vergara-Solana, Marcelo Araneda Padilla, Germán Ponce Díaz, Annie Mejaes, Manuel María Varela Lafuente, and U. Rashid Sumaila. Social effects of energy subsidies and taxes on CO_{2} emissions: the case of Mexican aquaculture public policies. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000920.

Michael:2021:DEB

[MWB^{+}21] Pamela E. Michael, Chris Wilcox, Christophe Barbraud, Karine Delord, Michael Sumner, and Henri Weimerskirch. Dynamic enforcement of bycatch via reproductive value can increase theoretical efficiency. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002955.

Mason:2023:LKA

[MWM^{+}23a] Julia G. Mason, Sarah J. Weisberg, Janelle L. Morano, Richard J. Bell, Mark Fitchett, Roger B. Griffis, Elliott L. Hazen, William D. Heyman, Kirstin Holsman, Kristin M. Kleisner, Katie Westfall, Michele K. Conrad, Margaret Daly, Abigail S. Golden, Chris J. Harvey, Lisa A. Kerr, Gway Kirchner, Arielle Levine, Rebecca L. Lewison, Sean M. Lucey, Wendy Morrison, Brandon Muffley, Jameal F. Samhouri, Matthew Seeley, S. Kalei Shotwell, and Diana L. Stram. Linking knowledge and action for climate-ready fisheries: Emerging best practices across the US. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002919.

McLean:2023:CCC

[MWM^{+}23b] Mercedes McLean, Brian Warner, Robert Markham, Mibu Fischer, Jim Walker, Carissa Klein, Maia Hoeberechts, and Daniel C. Dunn. Connecting conservation & culture: the importance of indigenous knowledge in conservation decision-making and resource management of migratory marine species. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001094.

Michaelis:2021:CES

[MWWS21] Adriane K. Michaelis, William C. Walton, Donald W. Webster, and L. Jen Shaffer. Cultural ecosystem services enabled through work with shellfish. Marine Policy, 132(??): ??, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003006.

Miao:2021:CDI

[MX21] Dandan Miao and Zhihua Xue. The current developments and impact of land reclamation control in China. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003936.

Ma:2023:APP

[MXZ23] Deyi Ma, Qi Xu, and Jietao Zhou. The application of the precautionary principle in the decision of releasing Fukushima radioactive contaminated water into the sea. Marine Policy, 156(??):??, October 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003366.

Minyou:2024:LEI

[MY24] Yu Minyou and Ni Yao. Law enforcement in the implication of Blue Cooperation — a reflection of China. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000782.

Moshy:2025:ING

[MYGM25] Victoria H. Moshy, Pius Z. Yanda, Brown Gwambene, and Anselm Mwajombe. The impact of natural gas development on the resilience of coastal social-ecological systems amid climate change in Mtwara, Tanzania. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25002714.

Manero:2024:NAE

[MYLV24] Ana Manero, Asad Yusoff, Mark Lane, and Katja Verreydt. A national assessment of the economic and wellbeing impacts of recreational surfing in Australia. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002653.

Marentette:2022:MEI

[MZ22] Julie R. Marentette and Fan Zhang. From means to ends: Insights into the operationalization and evaluation of sustainable fisheries. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001348.

Mao:2024:TDO

[MZ24] Zhengkai Mao and Zhijun Zhang. Taking the “UN Decade of Ocean Science for Sustainable Development” as an opportunity to help build a “Community with a Shared Future between China and Pacific Island Countries”. Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004761.

Murua:2023:LLF

[MZT^{+}23] Hilario Murua, Iker Zudaire, Mariana Tolotti, Jefferson Murua, Manuela Capello, Oihane C. Basurko, Iñigo Krug, Maitane Grande, Igor Arregui, Jon Uranga, Jose Maria Ferarios, Philippe Sabarros, Jon Ruiz, Yannick Baidai, Maria Lourdes Ramos, Jose Carlos Báez, Francisco Abascal, Haritz Arrizabalaga, Gala Moreno, Laurent Dagorn, and Josu Santiago. Lessons learnt from the first large-scale biodegradable FAD research experiment to mitigate drifting FADs impacts on the ecosystem. Marine Policy, 148(??): ??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004419.

Macedo:2025:LKO

[MZVC25] Thais Peixoto Macedo, Patrizia Ziveri, Benalsy Varela, and André Carlo Colonese. Local knowledge and official landing data point to decades of fishery stock decline in West Africa. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004470.

Mao:2024:TEO

[MZZ24] Zhengkai Mao, Shiyang Zhang, and Zhijun Zhang. Theoretical exploration of the order transformation in global ocean governance and China’s strategic response. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002513.

Nalley:2024:HIC

[Nal24] Eileen M. Nalley. Highlighting the impacts of ciguatera poisoning on fishers and fish consumers in Hawai’i to inform future management strategies. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004086.

Nielsen:2024:DFA

[NAN^{+}24] Max Nielsen, Anna Andersson, Rasmus Nielsen, Johan Blomquist, and Staffan Waldo. Does fishery activity affect local tourism? Evidence from Denmark. Marine Policy, 161 (??):??, March 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000496.

Nicolodi:2021:CGI

[NAP^{+}21] João Luiz Nicolodi, Milton Lafourcade Asmus, Marcus Polette, Alexander Turra, Carlos Alberto Seifert, Fernanda Terra Stori, Deborah Campos Shinoda, Alexandre Mazzer, Vitor Alberto de Souza, and Rafael Kuster Gonçalves. Critical gaps in the implementation of Coastal Ecological and Economic Zoning persist after 30 years of the Brazilian coastal management policy. Marine Policy, 128 (??):??, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000816.

Nyiawung:2022:CSS

[NAS^{+}22] Richard A. Nyiawung, Raymond K. Ayilu, Neville N. Suh, Norbert N. Ngwang, Frederick Varnie, and Philip A. Loring. COVID-19 and small-scale fisheries in Africa: Impacts on livelihoods and the fish value chain in Cameroon and Liberia. Marine Policy, 141(??):??, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001518.

Nunes:2023:EIF

[NBB^{+}23] Diogo M. Nunes, Alan C. Bezerra, Wik M. S. Barros, Paulo V. N. Araújo, Ilka S. L. Branco-Nunes, Rafael A. Magris, Pedro H. C. Pereira, Iran C. Normande, Rafael S. L. Barboza, and Andrei T. C. Cardoso. Evidence of illegal fishing within the largest Brazilian coastal MPA: Turning a blind eye to the obvious. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003712.

Nunny:2025:HMM

[NBB^{+}25] Laetitia Nunny, Mike Bossley, Rebecca M. Boys, Philippa Brakes, Tilen Genov, E. C. M. Parsons, Katharina J. Peters, Naomi A. Rose, Claire A. Simeone, Karen A. Stockin, Courtney S. Vail, Krista van der Linde, Ingrid N. Visser, and Mark P. Simmonds. Out of habitat marine mammals — identification, causes, and management recommendations. Marine Policy, 177(??):??, July 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000673.

Nijman:2025:LSG

[NC25] Vincent Nijman and Jessica Chavez. Limited support for Global North–Global South inequality in regulating international trade in scalloped hammerhead sharks. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004433.

Nguyen:2023:BAC

[NCD23] Son Nguyen, Peggy Shu-Ling Chen, and Yuquan Du. Blockchain adoption in container shipping: an empirical study on barriers, approaches, and recommendations. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002579.

Nakamura:2021:ULP

[NCE21] Julia Nakamura, Ratana Chuenpagdee, and Mostafa El Halimi. Unpacking legal and policy frameworks: a step ahead for implementing the Small-Scale Fisheries Guidelines. Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001792.

Nefzi:2025:CBB

[NCF25] Nourhaine Nefzi, Chehir Chehibi, and Ella Farekh. Can blue bonds offer true diversification? Insights from timefrequency analysis. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001290.

Narayanan:2023:KFI

[NEF23] Saratkumar C. Narayanan, Gholam Reza Emad, and Jiangang Fei. Key factors impacting women seafarers’ participation in the evolving workplace: a qualitative exploration. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004547.

Nijamdeen:2023:UEI

[NEH^{+}23] T. W. G. F. Mafaziya Nijamdeen, Ngendahimana Ephrem, Jean Hugé, Kodikara Arachchilage Sunanda Kodikara, and Farid Dahdouh-Guebas. Understanding the ethnobiological importance of mangroves to coastal communities: a case study from Southern and North-western Sri Lanka. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004389.

Nguyen:2023:PSS

[NGA23] Ly Nguyen, Zhifeng Gao, and James L. Anderson. Perception shifts in seafood consumption in the United States. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004857.

Neves:2021:SRS

[NGP^{+}21] João Neves, Jean-Christophe Giger, Nuno Piçarra, Vasco Alves, and Joana Almeida. Social representations of sharks, perceived communality, and attitudinal and behavioral tendencies towards their conservation: an exploratory sequential mixed approach. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002712.

Neis:2022:FSD

[NGRM^{+}22] Barbara Neis, María Andrée López Gómez, Emily Reid-Musson, Brenda Greenslade, David Decker, Joel Finnis, and Christine Knott. Fishing safely during COVID-19 in Newfoundland and Labrador, Canada: Making it happen. Marine Policy, 145(??):??, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003281.

Nuez:2023:ACS

[NGT^{+}23] Ignasi Nuez, Ioannis Giovos, Francesco Tiralongo, Jaime Penadés-Suay, Ilija Cetkovic, Manfredi Di Lorenzo, Periklis Kleitou, Rigers Bakiu, Mohamed Nejmeddine Bradai, Sara A. A. Almabruk, Roxani Naasan Aga Spyridopoulou, Andréa Sabbio, and Manel Gazo. Assessing the current status of Hexanchus griseus in the Mediterranean Sea using local ecological knowledge. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004250.

Nguyen:2023:BRF

[Ngu23] Hong-Kong To Nguyen. Book review: Freedom of Navigation and the Law of the Sea: Warships, States and the Use of Force, Cameron Moore, in: Routledge Research on the Law of the Sea. Routledge Taylor & Francis, London and New York (2021), ISBN 978-1-138-38814-7. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001367.

Nguyen:2024:MDD

[Ngu24] Hong-Kong To Nguyen. Mapping a decade of disputant and non-disputant behaviors in the South China Sea dispute. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001878.

Nakamura:2020:ABF

[NH20] Julia Nakamura and Fábio Hazin. Assessing the Brazilian federal fisheries law and policy in light of the Voluntary Guidelines for Securing Sustainable Small-scale fisheries. Marine Policy, 113(??):??, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306359.

Nielsen:2023:SPC

[NHA23] Rasmus Nielsen, Ayoe Hoff, and Jesper L. Andersen. Structural and productivity changes from introducing strong user rights in the Danish demersal fisheries. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004328.

Nguyen:2023:ELC

[NHD23] Hai Nguyen, Richard. J. Harper, and Bernard Dell. Examining local community understanding of mangrove carbon mitigation: a case study from Ca Mau province, Mekong River Delta, Vietnam. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004456.

Niu:2022:BRL

[Niu22] Bingru Niu. Book review: The Law of the Sea and Climate Change — Solutions and Constraints, Elise Johansen, Signe Veierud Busch, Ingvild Ulrikke Jakobsen. Cambridge University Press (2021), 435 pp., ISBN 978-1-108-84226-6. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005522.

Noble-James:2023:MBH

[NJBM^{+}23] Tamsyn Noble-James, Ross Bullimore, Fionnuala McBreen, Joey O’Connor, James Highfield, Chris McCabe, Simeon Archer-Rand, Anna-Leena Downie, Jon Hawes, and Peter Mitchell. Monitoring benthic habitats in English Marine Protected Areas: Lessons learned, challenges and future directions. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003858.

Notanubun:2025:BRB

[NJW25] Christy Agathys Notanubun, Hamran Ulil Amrin Jamaludin, and Sakina Wally. Book review: The Blue Compendium: From Knowledge to Action for a Sustainable Ocean Economy, Jane Lubchenco, Peter M. Haugan (Eds.). Springer (2023), 915, ISBN: 978-3-031-16276-3 (hardback). Marine Policy, 171(??):??, January 2025.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004792.

Nam:2021:ICO

[NK21] Dong Nam and Mingyu Kim. Implication of COVID-19 outbreak on ship survey and certification. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002268.

Neimane:2023:MSP

[NM23] Leila Neimane and Séverine Michalak. Maritime spatial plans as an object of the right of access to information. Marine Policy, 158(??):??, December 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004037.

Navarro:2024:EUO

[NMCL24] Matt L. Navarro, Jacquomo Monk, Gabrielle H. Cummins, and Tim J. Langlois. Embracing uncertainty in ocean accounts. Marine Policy, 162(??):??, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000381.

Ngoc:2025:VFP

[NMN^{+}25] Quach Thi Khanh Ngoc, Alena Mychkova, Nguyen Thi Ngan, Pham Quoc Hung, Thomas Potempa, and Max Patrick Ehleben. Vietnamese fishers’ perceptions on the effects of abandoned, lost, and discarded fishing gear and their willingness to participate in retrieval efforts. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001800.

Nielsen:2021:MBF

[NNM^{+}21] Pernille Nielsen, Mette Møller Nielsen, Ciaran McLaverty, Kasper Kristensen, Kerstin Geitner, Jeppe Olsen, Camille Saurel, and Jens Kjerulf Petersen. Management of bivalve fisheries in marine protected areas. Marine Policy, 124(??): ??, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20310083.

Nuttall:2021:PID

[NNR^{+}21] P. Nuttall, A. Newell, I. Rojon, B. Milligan, and A. Irvin. Pacific island domestic shipping emissions abatement measures and technology transition pathways for selected ship types. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003158.

Nguyen:2024:CSR

[NNRS24] Lan Anh Nguyen, Manh-Hung Nguyen, Arnaud Reynaud, and Michel Simioni. A comparative study of residents and tourists’ valuation for a heterogeneous environmental good: the case of coastal erosion. Marine Policy, 161(??): ??, March 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000368.

Nwani:2025:IBE

[NO25] Chinazaekpere Nwani and Philip Chimobi Omoke. Investing in the blue economy: Socioeconomic dynamics in the impact of government expenditure on marine ecosystem sustainability in Nigeria. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25001162.

Nakamura:2022:PTD

[NOB22] Katrina Nakamura, Yoshitaka Ota, and Francisco Blaha. A practical take on the duty to uphold human rights in seafood workplaces. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004553.

Nobrega:2023:STS

[NOL^{+}23] Marcelo Francisco Nóbrega, Matheus Assis Oliveira, Marcelo Gomes Lira, Suelen de Souza Rocha, and Jorge Eduardo Lins Oliveira. Sustainability of tunas and swordfish exploitation in the equatorial tropical Atlantic Ocean. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002889.

Nyiawung:2021:MSC

[NRF21] Richard A. Nyiawung, Ajith Raj, and Paul Foley. Marine Stewardship Council sustainability certification in developing countries: Certifiability and beyond in Kerala, India and The Gambia, West Africa. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001378.

Nasrudin:2020:CCT

[NRYG20] Rus’an Nasrudin, Budy P. Resosudarmo, Satoshi Yamazaki, and Wardis Girsang. Contribution of cash transfers in moderating household food insecurity in smallisland communities: Experimental evidence from Indonesia. Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307705.

Noack:2020:SUP

[NSK20] Thomas Noack, Esther Savina, and Junita D. Karlsen. Survival of undersized plaice (Pleuronectes platessa) discarded in the bottom otter trawl and Danish seine mixed fisheries in Skagerrak. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301599.

Nahata:2025:QFC

[NSM25] Kimiharu Nahata, Nobuhiko Sato, and Kazushi Miyashita. Quantification of fishermen’s cognitive values to inform recruitment policies. Marine Policy, 174(??):??, April 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000089.

Nisa:2022:WBW

[NSN22] Zahidah Afrin Nisa, Clive Schofield, and Francis C. Neat. Work Below Water: the role of scuba industry in realising sustainable development goals in small island developing states. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005297.

Nunez-Sanchez:2020:GSL

[NSPRSS20] M. J. Nuñez-Sanchez, L. Perez-Rojas, L. Sciberras, and J. Ramos Silva. Grounds for a safety level approach in the development of long-lasting regulations based on costs to reduce fatalities for sustaining industrial fishing vessel fleets. Marine Policy, 113(??):??, March 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18309977.

NSouvi:2021:FAT

[NSZ^{+}21] Kodjo N’Souvi, Chen Sun, Haiqing Zhang, Debi Ahoefa Broohm, and Mawussé Komlagan Nézan Okey. Fisheries and aquaculture in Togo: Overview, performance, fisheries policy, challenges and comparative study with Ghana, Mali, Niger and Senegal fisheries and aquaculture. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100292X.

Nyman:2020:TCR

[NT20] Elizabeth Nyman and Rachel Tiller. ‘Is there a court that rules them all’? Ocean disputes, forum shopping and the future of Svalbard. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307401.

Nassiri:2021:HEC

[NTH^{+}21] A. Nassiri, O. Thébaud, S. J. Holbrook, M. Lauer, A. Rassweiler, R. J. Schmitt, and J. Claudet. Hedonic evaluation of coral reef fish prices on a direct sale market. Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001366.

Niella:2023:MYE

[NUD^{+}23] Yuri Niella, Vinay Udyawer, Michael Drew, Brett Simes, Hugh Pederson, and Charlie Huveneers. Multi-year effects of wildlife tourism on shark residency and implications for management. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004092.

Nielsen:2023:CIE

[NVP^{+}23] Rasmus Nielsen, Sebastian Villasante, Jose Manuel Fernandez Polanco, Jordi Guillen, Ignacio Llorente Garcia, and Frank Asche. The Covid-19 impacts on the European Union aquaculture sector. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004080.

Nwafor:2020:PBP

[NW20] Ndubuisi Nwafor and Tony R. Walker. Plastic Bags Prohibition Bill: a developing story of crass legalism aiming to reduce plastic marine pollution in Nigeria. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20303870.

Nomura:2024:EGR

[NWS^{+}24] Keiko J. Nomura, A. John Woodill, Jonathan Sweeney, Michael Harte, Jameal F. Samhouri, and James R. Watson. Emergent geopolitical risks from fishing activities and past conflicts in the Pacific Ocean. Marine Policy, 166(??): ??, August 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2400232X.

Nie:2021:CER

[NWZ^{+}21] Xin Nie, Jianxian Wu, Wei Zhang, Jun Zhang, Wenhuan Wang, Yinghui Wang, Yaping Luo, and Han Wang. Can environmental regulation promote urban innovation in the underdeveloped coastal regions of western China? Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003201.

Owusu:2020:OAR

[OA20] Victor Owusu and Edo Andriesse. From open access regime to closed fishing season: Lessons from small-scale coastal fisheries in the Western Region of Ghana. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309856.

Otumawu-Apreku:2024:FAG

[OAAMS24] Kofi Otumawu-Apreku, Transform Aqorao, Fiona Meke, and Kezyiah Lusa Saepioh. Fishing activities in Pacific Island countries: a human-rights perspective. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2400006X.

Ospina-Alvarez:2025:OLC

[OAdJC^{+}25] A. Ospina-Alvarez, S. de Juan, C. Cáceres, U. R. Sumaila, G. Sant, N. K. Dulvy, C. A. Simpfendorfer, R. García, S. Niedermüller, and S. Villasante. Ocean leadership: Can the European Union influence the global shark meat trade? Marine Policy, 177(??):??, July 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000740.

Otumawu-Apreku:2020:PES

[OAM20] Kofi Otumawu-Apreku and Stephanie F. McWhinnie. Profit Efficiency of the South Australian rock lobster fishery: a Nerlovian and directional distance function approach. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305950.

Otumawu-Apreku:2021:RFC

[OAM21] Kofi Otumawu-Apreku and Stephanie F. McWhinnie. The role of fixed cost and non-discretionary variables in fisheries: the case of South Australian rock lobster. Marine Policy, 133(??):??, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003560.

Oanta:2022:AAS

[Oan22] Gabriela A. Oanta. The application of administrative sanctions in the fight against IUU fishing: an assessment of Spanish practice. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002585.

Osterman:2022:WBH

[ÖB22] Cecilia Österman and Magnus Boström. Workplace bullying and harassment at sea: a structured literature review. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005212.

Ojwala:2024:UWR

[OBNK24] Renis Auma Ojwala, Susan Buckingham, Francis Neat, and Momoko Kitada. Understanding women’s roles, experiences and barriers to participation in ocean science education in Kenya: recommendations for better gender equality policy. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300533X.

Obura:2020:GSE

[Obu20] David O. Obura. Getting to 2030 — scaling effort to ambition through a narrative model of the SDGs. Marine Policy, 117(??):??, July 2020. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309340.

OHanlon:2020:CII

[OC20] Zoë O’Hanlon and Valerie Cummins. A comparative insight of Irish and Scottish regulatory frameworks for offshore wind energy — an expert perspective. Marine Policy, 117 (??):??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X1930418X.

OConnor:2024:CMM

[OC24] Helen Murray O’Connor and J. Andrew G. Cooper. Coastal and marine management — navigating islands of data. Marine Policy, 169(??):??, November 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400277X.

Ortega:2023:IPD

[OCCSC23] Miquel Ortega, María D. Castro-Cadenas, Jeroen Steenbeek, and Marta Coll. Identifying and prioritizing demersal fisheries restricted areas based on combined ecological and fisheries criteria: the western Mediterranean. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003834.

OKeeffe:2020:SAC

[OCD^{+}20] Jane M. O’Keeffe, Valerie Cummins, Robert J. N. Devoy, Donald Lyons, and Jeremy Gault. Stakeholder awareness of climate adaptation in the commercial seaport sector: a case study from Ireland. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16302408.

Okyere:2023:CDM

[OCD^{+}23] Isaac Okyere, Ernest Obeng Chuku, Selorm Awiah Dzantor, Violet Ahenkorah, and Richard Adade. Capacity deficit and marginalisation of artisanal fishers hamper effective fisheries governance in Ghana: Insights and propositions for promoting sustainable small-scale fisheries. Marine Policy, 153 (??):??, July 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001677.

Orobko:2022:PCC

[OCT^{+}22] Melissa Orobko, Gillian Curren, Thomas W. Therriault, Peter Lawton, Lucia Fanning, and Isabelle M. Côté. Perception and consideration of cumulative effects and multiple stressors by marine managers in Canada. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001695.

Obregon:2022:LFK

[OCZ^{+}22] Clara Obregón, Joseph Christensen, Dirk Zeller, Michael Hughes, James R. Tweedley, Andrea Gaynor, and Neil R. Loneragan. Local fisher knowledge reveals changes in size of blue swimmer crabs in small-scale fisheries. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001919.

ODonnell:2021:CLF

[O’D21] Tayanah O’Donnell. Coastal Lawscape: a framework for understanding the complexities of climate change adaptation. Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001433.

Orihuela:2025:WPF

[ODG25] Carlos Orihuela and Jose Davila-Garcia. Wealth of Peruvian fishery. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002313.

Owusu:2025:SOS

[OEA25] Victor Owusu, Rosina Sheburah Essien, and Moses Adjei. The same old story: ‘Saiko’ practices and coastal livelihoods in Ghana’s small-scale fisheries. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24005736.

OConnor:2023:PCP

[OEV23] Eamonn O’Connor, Natasha Evers, and Amaya Vega. Port capacity planning — a strategic management perspective. Marine Policy, 150(??):??, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000647.

OMahony:2020:IFC

[OGGC20] C. O’Mahony, S. Gray, J. Gault, and V. Cummins. ICZM as a framework for climate change adaptation action — Experience from Cork Harbour, Ireland. Marine Policy, 111(??): ??, January 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X15002924.

Omukoto:2024:FMF

[OGH24] Johnstone O. Omukoto, Nicholas A. J. Graham, and Christina C. Hicks. Fish markets facilitate nutrition security in coastal Kenya: Empirical evidence for policy leveraging. Marine Policy, 164(??):??, June 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001775.

Okazaki:2024:PSF

[OGM24] Ryotaro Okazaki, Klaus Glenk, and Mitsutaku Makino. Perceptions of “sustainable fisheries” in the UK and Japan using best–worst scaling. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24003981.

Orchard:2024:RTR

[OGS24] Shane Orchard, Shawn Gerrity, and David R. Schiel. Rethinking recreational fishing — how a natural disaster presents insights and opportunities for achieving sustainability and equity objectives. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004499.

Ojaveer:2025:CEB

[OGS25] Henn Ojaveer, Bella Galil, and Hanno Seebens. The commitment of the EU Biodiversity Strategy for 2030 is unfeasible for marine threatened species affected by invasive alien species. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005827.

Olafsdottir:2024:GLC

[ÓHCÓ24] Gudbjörg Ásta Ólafsdóttir, Theresa Henke, Catherine P. Chambers, and Steinunn Hilma Ólafsdóttir. Gaps in legislation and communication identified as stakeholders reflect on 30 × 30 policy in Icelandic waters. Marine Policy, 170(??): ??, December 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24004226.

Okubo:2023:PSE

[OI23] Ayako Okubo and Atsushi Ishii. Pursuing sustainability? Ecosystem considerations in Japan’s fisheries governance. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001306.

Ozkaynak:2024:BSQ

[Öİ24] Ömer Harun Özkaynak and Gönül Tuğrul İçemer. Based on stakeholder questionnaires of ship-derived waste, waste management re-planning in ports and determining the need for inspection. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004451.

Ontomwa:2024:LPF

[OKFN24] Mary Binsari Ontomwa, Edward Ndirui Kimani, Bernerd Mulwa Fulanda, and Chrisphine Sangara Nyamweya.

Longline pelagic fishery assemblage in Kenya’s exclusive economic zone marine waters. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23005225.

Obane:2024:AAS

[OKH^{+}24] Hideaki Obane, Kentaro Kazama, Hiroshi Hashimoto, Yu Nagai, and Kenji Asano. Assessing areas suitable for offshore wind energy considering potential risk to breeding seabirds in northern Japan. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23005158.

Ojwala:2022:EGP

[OKNB22] Renis Auma Ojwala, Momoko Kitada, Francis Neat, and Susan Buckingham. Effectiveness of gender policies in achieving gender equality in ocean science programmes in public universities in Kenya. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002846.

Orejas:2020:TCA

[OKR^{+}20] Covadonga Orejas, Ellen Kenchington, Jake Rice, Georgios Kazanidis, Andreas Palialexis, David Johnson, Matthew Gianni, Roberto Danovaro, and J. Murray Roberts. Towards a common approach to the assessment of the environmental status of deep-sea ecosystems in areas beyond national jurisdiction. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309236.

Oldach:2022:MUW

[OKS^{+}22] Eliza Oldach, Helen Killeen, Priya Shukla, Ellie Brauer, Nicholas Carter, Jennifer Fields, Alexandra Thomsen, Cassidy Cooper, Leah Mellinger, Kaiwen Wang, Carl Hendrickson, Anna Neumann, Pernille Sporon Bøving, and Nann Fangue. Managed and unmanaged whale mortality in the California Current Ecosystem. Marine Policy, 140(??):??, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000860.

Oksas:2023:CES

[Okş23] Olgay Okşaş. Carbon emission strategies for container handling equipment using the activity-based method: a case study of Ambarlı container port in Turkiye. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000076.

Ourens:2022:LSS

[OMC^{+}22] Rosana Ouréns, Michael C. Melnychuk, Larry B. Crowder, Nicolas L. Gutierrez, Ray Hilborn, Cristina Pita, and Omar Defeo. Linking small-scale fisheries performance to governance attributes: a quantitative assessment from stakeholders’ perceptions in the Americas and Europe. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004875.

Obane:2021:APA

[ONA21] Hideaki Obane, Yu Nagai, and Kenji Asano. Assessing the potential areas for developing offshore wind energy in Japanese territorial waters considering national zoning and possible social conflicts. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001251.

Olsen:2020:MLI

[ONN^{+}20] Julia Olsen, Leticia Antunes Nogueira, Anne Katrine Normann, Bjørn Vidar Vangelsten, and Ingrid Bay-Larsen. Marine litter: Institutionalization of attitudes and practices among fishers in Northern Norway. Marine Policy, 121(??): ??, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20308575.

Osmundsen:2022:APD

[OOGA22] Tonje C. Osmundsen, Marit Schei Olsen, Asle Gauteplass, and Frank Asche. Aquaculture policy: Designing licenses for environmental regulation. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000252.

Okeke-Ogbuafor:2024:CSF

[OOGK^{+}24] Nwamaka Okeke-Ogbuafor, Tim Gray, Sheku Kamara, Edward Sesay, Abdulai Dauda, Selina M. Stead, Danielle Robinson, and Kelechi Johnmary Ani. Climate-smart fisheries: CO_{2} emissions reduction and food security are complementary. Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004591.

Okeke-Ogbuafor:2020:CRF

[OOGS20a] Nwamaka Okeke-Ogbuafor, Tim Gray, and Selina Marguerite Stead. The controversial role of foreign fisheries consultants in Sierra Leone’s coastal waters. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18307036.

Okeke-Ogbuafor:2020:TWP

[OOGS20b] Nwamaka Okeke-Ogbuafor, Tim Gray, and Selina Marguerite Stead. Is there a ‘wicked problem’ of small-scale coastal fisheries in Sierra Leone? Marine Policy, 118(??): ??, August 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X18306341.

Otwoma:2025:LKC

[OOM^{+}25] Levy Michael Otwoma, Julia Obuya, Christopher Aura Mulanda, Victor Mwakha Alati, Iddy Munuphe, Jibril Olunga, Mathews Wafula, Jelvas Mwaura, and Christian Chimimba. Local knowledge corroborates threats of local extinctions in Kenya’s exploited reef fishes. Marine Policy, 177(??):??, July 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2500065X.

Okeke-Ogbuafor:2021:IAP

[OOSG21] Nwamaka Okeke-Ogbuafor, Selina Stead, and Tim Gray. Is inland aquaculture the panacea for Sierra Leone’s decline in marine fish stocks? Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002748.

OHagan:2020:ATW

[OPL20] Anne Marie O’Hagan, Shona Paterson, and Martin Le Tissier. Addressing the tangled web of governance mechanisms for land-sea interactions: Assessing implementation challenges across scales. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X18309163.

Ofori:2021:MIF

[OR21] Roland O. Ofori and Mark D. Rouleau. Modeling the impacts of floating seaweeds on fisheries sustainability in Ghana. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000373.

Oral:2023:USL

[Ora23] Nilufer Oral. UNCLOS and sea level rise. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22005012.

Orchard:2025:PCA

[ORdS25] Shane Orchard, Jesse Reiblich, and Maurício D. dos Santos. Public coastal access: Emerging issues and potential solutions. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001745.

Olaya-Restrepo:2022:MNA

[ORSB22] Julián Olaya-Restrepo, Alexandre Schiavetti, and Marcos S. Barbeitos. A multilayered network analysis of social participation in the management of Marine Protected Areas in Brazil. Marine Policy, 146(??):??, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003761.

OConnor:2024:PPS

[OSBA24] Ryan J. O’Connor, Ana K. Spalding, Alison W. Bowers, and Nicole M. Ardoin. Power and participation: a systematic review of marine protected area engagement through participatory science methods. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24001313.

Osmundsen:2023:PRF

[Osm23] Lori Osmundsen. Port reception facilities and a regional approach: a bridge for abating plastic pollution in the Arctic? Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004833.

Osthagen:2020:MBD

[Øst20] Andreas Østhagen. Maritime boundary disputes: What are they and why do they matter? Marine Policy, 120(??): ??, October 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20302426.

Osthagen:2024:MSG

[Øst24] Andreas Østhagen. The myths of Svalbard geopolitics: an Arctic case study. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24001817.

Ovegaard:2021:MAA

[OSV21] Mikael Ovegård, Andreas Sundelöf, and Daniel Valentinsson. Much ado about nothing: an example of how failed incentives thwarted the implementation of the EU landing obligation. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309507.

Ortega:2025:PPF

[OSZ^{+}25] Miquel Ortega, Sara Mohamed Santamaria, Christos Zografos, Yesmina Mascarell-Rocher, Marta Coll, and Valerio Sbragaglia. Past, present and future of small-pelagic fisheries in the north-western Mediterranean Sea through fishers’ perceptions. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004901.

Oposa:2023:RSC

[OT23] Anna R. Oposa and Erika J. Techera. A review of shark conservation and management legal frameworks in the Philippines. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002464.

Ostwald:2021:MTS

[OTK^{+}21] Ashlea Ostwald, Vivitskaia J. D. Tulloch, Peter M. Kyne, Nicholas J. Bax, Piers K. Dunstan, Luciana C. Ferreira, Michele Thums, Judy Upston, and Vanessa M. Adams. Mapping threats to species: Method matters. Marine Policy, 131(??):??, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002256.

Ovalle:2022:UDL

[OVA22] Juan Carlos Ovalle, Carlos Vilas, and Luís T. Antelo. On the use of deep learning for fish species recognition and quantification on board fishing vessels. Marine Policy, 139 (??):??, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000628.

Ostale-Valriberas:2022:AMM

[OVSVPP^{+}22] E. Ostalé-Valriberas, J. Sempere-Valverde, A. Pavón-Paneque, S. Coppa, F. Espinosa, and J. C. García-Gómez. Artificial marine micro-reserves as a new ecosystem-based management tool for marine conservation: the case of Patella ferruginea (Gastropoda, Patellidae), one of the most endangered marine invertebrates of the Mediterranean. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005285.

OLeary:2024:PIC

[OWC^{+}24] Bethan C. O’Leary, Louisa E. Wood, Cindy Cornet, Callum M. Roberts, and Catarina Fonseca. Practitioner insights on challenges and options for advancing blue Naturebased Solutions. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001027.

Ortiz-Wolford:2021:CSR

[OWCFDC21] Jenniffer Ortiz-Wolford, M. Fabiola Corona-Figueroa, Vanessa Dávila, and Andrea A. Cabrera. Cetacean stranding records along the Pacific coastline of Guatemala, 2007– 2021: Implications for management, conservation and research. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004383.

Odhiambo:2020:EME

[OWS20] Jacob Ochiewo Odhiambo, Joseph Wakibia, and Maurice M. Sakwa. Effects of monitoring and evaluation planning on implementation of poverty alleviation mariculture projects in the coast of Kenya. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19309595.

Okafor-Yarwood:2024:TMS

[OYECdOP24] Ifesinachi Okafor-Yarwood, Oliver Eastwood, Noleen Chikowore, and Lucas de Oliveira Paes. Technology and maritime security in Africa: Opportunities and challenges in Gulf of Guinea. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23005092.

Okafor-Yarwood:2022:SRF

[OYKBA22] Ifesinachi Okafor-Yarwood, Nelly I. Kadagi, Dyhia Belhabib, and Edward H. Allison. Survival of the Richest, not the Fittest: How attempts to improve governance impact African small-scale marine fisheries. Marine Policy, 135(??): ??, January 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004589.

Okomoda:2025:SPC

[OZF^{+}25] Victor Tosin Okomoda, Markose Chekol Zewdie, Idowu James Fasakin, Sunil Niranjan Siriwardena, Shehu Latunji Akintola, Abayomi Abdul-Azeez Jimoh, Matthew Adepoju Ajibade, Isa Olalekan Elegbede, Leila Basti, and Rodrigue Yossa. Stakeholders’ perspective on current status and preferred approaches to mariculture development in coastal areas of Southwest Nigeria. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25002751.

Pazir:2022:ECP

[PAK22] Mohammad Khalil Pazir, Amirhossein Ahmadi, and Parisa Hossein Khezri. The effect of COVID-19 pandemic on the shrimp industry of Iran. Marine Policy, 136(??): ??, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100511X.

Pomeroy:2020:FIB

[PALB20] Robert Pomeroy, Carlos Arango, Cristopher G. Lomboy, and Steve Box. Financial inclusion to build economic resilience in small-scale fisheries. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19306645.

Pan:2025:FLE

[Pan25a] Xiaolin Pan. Fisheries law enforcement in the yellow sea by Chinese coast guard agencies: Legal sources, core issues and dispute settlement. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24004329.

Panda:2025:IEU

[Pan25b] Jagannath Panda. India’s emerging undersea cable landscape: Varied Indo-Pacific partnerships to boost geopolitical ambitions? Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004172.

Patron:2022:CPI

[Pat22] Jose Manuel Sánchez Patrón. The criminal prosecution of illegal fishing and the jurisdiction of Spanish courts. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002561.

Pauksztat:2021:IRC

[Pau21] Birgit Pauksztat. Informal relations and communication about work-related problems in two multilingual crews. Marine Policy, 130(??):??, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305639.

Patterson:2020:ARM

[PB20] Douglas A. Patterson and Raj Bridgelall. Attack risk modelling for the San Diego maritime facilities. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20308563.

Penca:2024:BCR

[PBC^{+}24] Jerneja Penca, Andrea Barbanti, Christopher Cvitanovic, Amel Hamza-Chaffai, Ahmed Elshazly, Jean-Baptiste Jouffray, Nezha Mejjad, and Melita Mokos. Building competences for researchers working towards ocean sustainability. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001301.

Pecarevic:2023:ICC

[PBD^{+}23] Marijana Pećarević, Kruno Bonačić, Tatjana Dobroslavić, Leon Grubišić, Igor Čelić, Branko Glamuzina, Josip Maleš, Gianluca Sarà, and Maria Cristina Mangano. Impact of COVID-19 on Croatian mariculture: Findings from the first national surveys. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004420.

Poeta:2022:OYA

[PBG^{+}22] G. Poeta, M. Bazzichetto, L. Gallitelli, M. Garzia, F. Aprea, F. Bartoli, C. Battisti, S. Cascone, A. Corradi, D. D’Amelia, E. D’Amico, J. De Luca, F. Del Grosso, L. Iacobelli, S. Langone, C. Lembo Fazio, G. Locchi, M. Perrone, F. Petroni, D. Raimondi, F. Romiti, S. Secco, L. Sonet, A. Spinelli, S. Toscano, S. Vanadia, S. Vecchi, F. Zanon, and M. Malavasi. One year after on Tyrrhenian coasts: the ban of cotton buds does not reduce their dominance in beach litter composition. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002421.

Patrizzi:2023:MBC

[PBG23] Nayla S. Patrizzi, Francisco Barros, and Vinicius J. Giglio. Media bias on communicating conservation interventions: the case of large-scale marine protected areas in the South Atlantic. Marine Policy, 149(??):??, March 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000027.

Procksch:2023:HUS

[PBH^{+}23] Natália Procksch, Natalia Bragiola Berchieri, Rafael Kenji Horota, Vinícius Sales, Paulo Henrique Ott, Daniel Danilewicz, Taina Thomassim Guimaraes, Murilo Guimarães, Maurício Roberto Veronez, and Larissa Rosa de Oliveira. Habitat use by South American fur seals (Arctocephalus australis) and sea lions (Otaria flavescens) in a marine protected area in southern Brazil. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002208.

Pelage:2023:FPG

[PBS^{+}23] L. Pelage, A. Bertrand, S. C. W. Siqueira, A. C. A. P. Araújo, K. V. A. Avelino, C. Lopes da Silva, and T. Frédou. Fishers’ perceptions of global change to inform coastal planning in a data-poor socio-ecological system. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003172.

Pennino:2020:DBS

[PBT^{+}20] Maria Grazia Pennino, Ana Helena Bevilacqua, M. Angeles Torres, Jose M. Bellido, Jordi Sole, Jeroen Steenbeek, and Marta Coll. Discard ban: a simulation-based approach combining hierarchical Bayesian and food web spatial models. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307954.

Portman:2020:ABM

[PC20] Michelle E. Portman and André Zhulpa Camporesi. Attitudes and behaviours of marine recreationists towards conservation and environmental protection: a case study of Tel Aviv, Israel. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930884X.

Palm:2021:MLF

[PCA21] Karl E. Palm, Gary A. Campbell, and Jenny L. Apriesnig. Management of local fisheries: a case study of Laoang, Northern Samar, Philippines. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002682.

Pereira:2023:LIE

[PCA^{+}23] Laura M. Pereira, Guillermo Ortuño Crespo, Diva J. Amon, Renuka Badhe, Salomão Bandeira, Frida Bengtsson, Miranda Boettcher, Gabrielle Carmine, William W. L. Cheung, Bwalya Chibwe, Daniel Dunn, Maria A. Gasalla, Ghassen Halouani, David E. Johnson, Jean-Baptiste Jouffray, Silvana Juri, Patrick W. Keys, Hannah M. Lübker, Andrew S. Merrie, Farah Obaidullah, Juliano Palacios-Abrantes, Lynne J. Shannon, U. Rashid Sumaila, Edoardo Superchi, Naomi Terry, Colette C. C. Wabnitz, Moriaki Yasuhara, and Wei Zhou. The living infinite: Envisioning futures for transformed human-nature relationships on the high seas. Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001719.

Portilla-Cabrera:2024:POA

[PCCOS24] Cristiam Victoriano Portilla-Cabrera, Maria Alejandra Cifuentes-Ossa, and John Josephraj Selvaraj. Policyoriented adaptation strategies for artisanal marine fisheries under climate and non-climate stressors in San Andres de Tumaco, Nariño, Colombia. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23005262.

Pennino:2025:IFS

[PCCR^{+}25] Maria Grazia Pennino, Santiago Cerviño, Marta Cousido-Rocha, Ruth Fernandez, Diana Gonzalez-Troncoso, Fernando Gonzalez, Beatriz Guijarro, José Luis Pérez Gil, and Maria Soto. Identifying future steps to allow an efficient fisheries management advice among scientific committees in regional fisheries management organizations. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25002477.

Pascoe:2023:UDE

[PCD^{+}23] Sean Pascoe, Toni Cannard, Natalie A. Dowling, Catherine M. Dichmont, Frank Asche, and L. Richard Little. Use of Data Envelopment Analysis (DEA) to assess management alternatives in the presence of multiple objectives. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004912.

Paiva:2023:MCM

[PCG^{+}23] Sandra Vieira Paiva, Pedro Bastos Macedo Carneiro, Tatiane Martins Garcia, Tallita Cruz Lopes Tavares, Lidriana de Souza Pinheiro, Antonio Rodrigues Ximenes Neto, Tarin Cristino Montalverne, and Marcelo O. Soares. Marine carbonate mining in the Southwestern Atlantic: current status, potential impacts, and conservation actions. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004821.

Polejack:2025:HAO

[PCHD^{+}25] Andrei Polejack, Luciana Fernandes Coelho, Harriet Harden-Davies, Laura Elsler, Diva J. Amon, and Asha de Vos. Hope for an accessible ocean: Blue justice and ocean science diplomacy central to the outcome of the UN Decade of Ocean Science. Marine Policy, 176(??):??, June 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000545.

Pierucci:2022:GRM

[PCK22] Andrea Pierucci, Silvia Columbu, and Laurence T. Kell. A global review of MSC certification: Why fisheries withdraw? Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001713.

Peryman:2024:PPW

[PCN^{+}24] Matt Peryman, Romilly Cumming, Tina Ngata, Trisia A. Farrelly, Sascha Fuller, and Stephanie B. Borrelle. Plastic pollution as waste colonialism in Aotearoa (New Zealand). Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000769.

Prieto-Carolino:2021:GAP

[PCSS^{+}21] Alice Prieto-Carolino, Ida M. Siason, Marieta B. Sumagaysay, Rowena Paz L. Gelvezon, Harold M. Monteclaro, and Rosario H. Asong. A gender analysis of the processing sector of the tuna value chain in General Santos City, Philippines. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000889.

Paramana:2023:ACB

[PDB^{+}23] Th. Paramana, M. Dassenakis, N. Bassan, C. Dallangelo, P. Campostrini, S. Raicevich, F. Ronchi, G. Giorgi, A. Murillas-Maza, M. C. Uyarra, N. Papadopoulou, C. Smith, K. Jarni, Š. Koren Bačovnik, K. Klančnik, M. Pavičić, S. Skejić, O. Vidjak, J. F. Cadiou, L. López-López, I. Alvarez, L. Giannoudi, N. Streftaris, and P. Pagkou. Achieving coherence between the marine strategy framework directive and the maritime spatial planning directive. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300266X.

Poupart:2025:MBC

[PdBA^{+}25] Timothée Poupart, Thibaut de Bettignies, Matthieu Authier, Nicola Baccetti, Luca Börger, Iker Castège, Jacopo G. Cecere, Nicolas Courbin, Jamie Darby, Karine Delord, Ghislain Doremus, Hayley A. Douglas, Gilles Faggio, Giorgia Gaibani, Fabrice Gallien, Cécile Gicquel, Jacob González-Solís, David Grémillet, Simona Imperio, Jude V. Lane, Amélie Lescroël, Maite Louzao, Jeanne de Mazières, Noëmie Michez, Emilie Milon, Ellie Owen, Vitor H. Paiva, Federico De Pascalis, Clara Péron, Francesco Pezzo, Pascal Provost, Raül Ramos, Jaime A. Ramos, Solène Robert, Diego Rubolini, Olivier Scher, and Lorenzo Serra. Mismatches between the current marine Natura 2000 network and seabird distributions call for enhanced protected areas off metropolitan France. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25001940.

Pascoe:2025:KDM

[PDD^{+}25] Sean Pascoe, Natalie Dowling, Catherine M. Dichmont, Roy Deng, Andre E. Punt, and Ingrid van Putten. Key drivers of model choice by fisheries scientists and their propensity to adopt stock assessment packages. Marine Policy, 174 (??):??, April 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24005839.

Priyashadi:2022:SED

[PDJ22] M. S. V. H. Priyashadi, K. H. M. Ashoka Deepananda, and Asanka Jayasinghe. Socio-economic development of marine ornamental reef fish fishers in eastern Sri Lanka through the lenses of Human Development Index. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200183X.

Pazoto:2024:CPT

[PDS24] Carmen Edith Pazoto, Michelle Rezende Duarte, and Edson Pereira Silva. Challenges and prospects for teaching ocean literacy in Brazilian schools. Marine Policy, 166(??): ??, August 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24002185.

Plot:2025:AMT

[PEM^{+}25] Virginie Plot, Vanessa Estrade, Julie Martin, Thomas Rostaing, Tim Collins, and Violaine Dulau. Assessing marine traffic and related pressures and threats to cetacean populations to support vessel management in the Mascarene Important Marine Mammal Area. Marine Policy, 176(??):

??, June 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25000478.

Penca:2022:SPI

[Pen22] Jerneja Penca. Science, precaution and innovation for sustainable fisheries: the judgement by the Court of Justice of the EU regarding the electric pulse fishing ban. Marine Policy, 135(??):??, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004759.

Perraudeau:2024:IQC

[Per24] Maïa Perraudeau. Innovation in the quest for compatibility: Interplay of jurisdiction and allocation in the South Pacific Regional Fisheries Management Organisation. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400143X.

Pata:2024:EID

[PESO24] Ugur Korkut Pata, Sinan Erdogan, Sakiru Adebola Solarin, and Ilyas Okumus. Evaluating the influence of democracy, financial development, and fishery product consumption on fishing grounds: a case study for Malaysia. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002999.

Petersson:2022:TGF

[Pet22] Matilda T. Petersson. Transparency in global fisheries governance: the role of non-governmental organizations. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306682.

Parlee:2021:FSS

[PFG^{+}21] Courtenay E. Parlee, Paul Foley, María Andrée López Gómez, Md. Ruyel Miah, Charles Mather, and Robert L.

Stephenson. Full spectrum sustainability and a theory of access: Integrating social benefits into fisheries governance. Marine Policy, 134(??):??, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003754.

Pafi:2020:CTM

[PFM20] Maria Pafi, Wesley Flannery, and Brendan Murtagh. Coastal tourism, market segmentation and contested landscapes. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300774.

Purwanto:2023:MGB

[PFWM23] Purwanto, Erik C. Franklin, Alan White, and Ses Rini Mardiani. Multiple-goal bioeconomic programming to address conflicting management objectives in Indonesian small pelagic fisheries. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000465.

Pinarbasi:2020:KIT

[PGA^{+}20] Kemal Pınarbaşı, Ibon Galparsoro, Neil Alloncle, Frédéric Quemmerais, and Ángel Borja. Key issues for a transboundary and ecosystem-based maritime spatial planning in the Bay of Biscay. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20303298.

Pincinato:2022:MIS

[PGGA22] Ruth Beatriz Mezzalira Pincinato, Maria A. Gasalla, Taryn Garlock, and James L. Anderson. Market incentives for shark fisheries. Marine Policy, 139(??):??, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000781.

Pauksztat:2022:ICP

[PGK22] Birgit Pauksztat, Michelle R. Grech, and Momoko Kitada. The impact of the COVID-19 pandemic on seafarers’ mental health and chronic fatigue: Beneficial effects of onboard peer support, external support and Internet access. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005534.

Phayal:2024:CRB

[PGM^{+}24] Anup Phayal, Aaron Gold, Curie Maharani, Brandon Prins, and Sayed Fauzan Riyadi. The complementary relationship between illegal fishing and maritime piracy: a case study of the Gulf of Guinea. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002070.

Phayal:2022:IHM

[PGP22] Anup Phayal, Aaron Gold, and Brandon Prins. Interstate hostility and maritime crime: Evidence from South East Asia. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001816.

Pedroza-Gutierrez:2024:WFE

[PGSFRMHH24] Carmen Pedroza-Gutiérrez, Neyra Solano, Francisco Fernández-Rivera-Melo, and Isis Hernández-Herrera. What are the factors that enable or hinder gender equality? Lessons and experiences on gender dynamics in Mexican small-scale fisheries. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004773.

Phelan:2022:TMP

[PGSS22] Kieran Phelan, Alison Gardner, Elizabeth R. Selig, and Jessica L. Decker Sparks. Towards a model of portbased resilience against fisher labour exploitation. Marine Policy, 142(??):??, August 2022. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001555.

Pan:2024:CUC

[PH24] Min Pan and Henry P. Huntington. China–U.S. cooperation in the Arctic Ocean: Prospects for a new Arctic exceptionalism? Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002926.

Pham:2020:TMP

[Pha20] Thuy Thi Thanh Pham. Tourism in marine protected areas: Can it be considered as an alternative livelihood for local communities? Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303690.

Pascoe:2024:CRT

[PHD^{+}24] Sean Pascoe, Trevor Hutton, Roy Deng, Tom Kompas, and Annie Jarrett. Catch-rate triggers as a fisheries management tool for short lived species: Can they achieve MEY? Marine Policy, 164(??):??, June 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001635.

Palomo:2020:ISM

[PHF20] Leopoldo E. Palomo and Alvaro Hernández-Flores. Integrating a spatial model and decision theory towards optimal boating density and carrying capacity in a recreational fishery. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18305967.

Pons-Hernandez:2025:UCS

[PHP25] Monica Pons-Hernandez and Gohar A. Petrossian. Using crime script analysis to unveil illegal, unreported, and unregulated (IUU) fishing of European eels. Marine Policy, 173(??):??, March 2025. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24005694.

Pietkiewicz:2021:LSC

[Pie21] Michał Pietkiewicz. Legal status of Caspian Sea — problem solved? Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309684.

Pike:2024:IGF

[PJN24] Felicity Pike, Narriman S. Jiddawi, and Lina Mtwana Nordlund. Intertidal gleaning fisheries: Recognising localscale contributions and management scenarios. Marine Policy, 162(??):??, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000575.

Pineiro:2020:SHW

[PK20a] Laura Carballo Piñeiro and Momoko Kitada. Sexual harassment and women seafarers: the role of laws and policies to ensure occupational safety & health. Marine Policy, 117 (??):??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X1930106X.

Pratomo:2020:GAM

[PK20b] Eddy Pratomo and Jonathan Kwik. Good agreements make good neighbours: Settlements on maritime boundary disputes in South East Asia. Marine Policy, 117(??): ??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19308929.

Pentz:2023:WCC

[PK23] Brian Pentz and Nicole Klenk. Will climate change degrade the efficacy of marine resource management policies? Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005097.

Palialexis:2021:MBE

[PKB^{+}21] Andreas Palialexis, Vasiliki Kousteni, Laura Boicenco, Lisette Enserink, Kalliopi Pagou, Ulla Li Zweifel, Francesca Somma, Anna Cheilari, and David Connor. Monitoring biodiversity for the EU Marine Strategy Framework Directive: Lessons learnt from evaluating the official reports. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000841.

Patra:2023:MLM

[PKB^{+}23] Sivaji Patra, Muhammad Khurshid, Abas Basir, Pravakar Mishra, and M. V. Ramanamurthy. Marine litter management: a sustainable action plan and recommendations for the South Asian Seas region. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003871.

Park:2023:AMI

[PKC23] Sangwon Park, Chang Hoon Kim, and Minsoo Cheon. Analysis of marine incidents and policy implications: Case study of the South Korea. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002695.

Parviainen:2022:ESP

[PKH22] T. Parviainen, S. Kuikka, and P. Haapasaari. Enhancing science-policy interface in marine environmental governance: Oil spill response models as boundary objects in the Gulf of Finland, Baltic Sea. Marine Policy, 135(??): ??, January 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004747.

Puts:2023:TOB

[PKMT23] Miriam Püts, Alexander Kempf, Christian Möllmann, and Marc Taylor. Trade-offs between fisheries, offshore wind farms and marine protected areas in the southern North Sea — winners, losers and effective spatial management. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300101X.

Park:2022:EEP

[PKR^{+}22] Sunghwa Park, Taeil Kim, Heeyoung Ryu, Hanna Kim, and Janghan Kwon. The economic effect and policy performance of ship registration — a case of Korea. Marine Policy, 144(??):??, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002810.

Psuty:2020:ISS

[PKS20] Iwona Psuty, Tomasz Kulikowski, and Lena Szymanek. Integrating small-scale fisheries into Polish maritime spatial planning. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306979.

Pauly:2020:FSC

[PL20a] Daniel Pauly and Cui Liang. The fisheries of the South China Sea: Major trends since 1950. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19302015.

Pellowe:2020:IBF

[PL20b] Kara E. Pellowe and Heather M. Leslie. The interplay between formal and informal institutions and the potential for co-management in a Mexican small-scale fishery. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930781X.

Pinn:2023:IEL

[PL23] Eunice H. Pinn and Lewis Le Vay. Interpretation of the European legal framework for the microbiological classification of bivalve mollusc production areas. Marine Policy, 148(??): ??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000064.

Park:2024:DMG

[PL24a] Eon Kyung Park and Seokwoo Lee. Discharge measures for Fukushima wastewater and South Korea’s response and implications of the WTO Korea–Japan fishery dispute (Korearadionuclides) case. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004633.

Petrossian:2024:PTE

[PL24b] Gohar A. Petrossian and Jess Lettieri. A precautionary tale: Exploring the risks of deep-sea mining. Marine Policy, 162 (??):??, April 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2400071X.

Pickens:2024:WIW

[PLHK^{+}24] Chris Pickens, Hannah Lily, Ellycia Harrould-Kolieb, Catherine Blanchard, and Anindita Chakraborty. From what-if to what-now: Status of the deep-sea mining regulations and underlying drivers for outstanding issues. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23005006.

Parker:2024:LGA

[PLLJ24] Quinn M. Parker, Hoby Tsimijaly Longosoa, Stephen Long, and Peter J. S. Jones. A longitudinal governance analysis of a locally managed marine area: Ankobohobo wetland smallscale mud crab fishery, Madagascar. Marine Policy, 163 (??):??, May 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24001362.

Powell:2024:FPM

[PLOG24] Farrah Powell, Arielle Levine, and Lucia Ordonez-Gauger. Fishermen’s perceptions of management in the California spiny lobster and California market squid fisheries. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000137.

Perng:2024:MDA

[PLWO24] Lansing Y. Perng, Kirsten M. Leong, Mariska Weijerman, and Kirsten L. L. Oleson. A multi-decadal assessment of social thresholds and outcomes in marine social-ecological systems in Hawai’i. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001398.

Perez-Martin:2020:LMV

[PMAVM20] Ricardo I. Pérez-Martín, Luis T. Antelo, J. A. Vázquez, and Jesús Mirón. An on-land management and valorisation approach for biomass associated with landing obligation compliance. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308443.

Prado:2024:PSP

[PMC24] Deborah Santos Prado, Ivan Machado Martins, and Ronaldo Adriano Christofoletti. Participatory scenario planning as a useful method for transforming ocean conflicts: Insights from a small-scale fishing conflict in São Paulo coast, Brazil. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004554.

Puentes:2022:SMB

[PMFR^{+}22] Vladimir Puentes, Paola A. Mejía-Falla, John G. Ramirez, Luis M. Manjarrés-Martínez, Juan M. Rguez-Baron, Luis A. Zapata, José Tavera, Fabio Gómez-Delgado, Carlos G. Barreto, Emiliano Zambrano, Ángel A. Villa, and Andrés F. Navia. Sharks and marine batoids management in Colombia: Policy instruments, management duty and implications for their populations and stakeholders. Marine Policy, 145(??):??, November 2022. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003116.

Paolacci:2021:LSP

[PMK^{+}21] Simona Paolacci, Rogério Mendes, Regina Klapper, Amaya Velasco, Graciela Ramilo-Fernandez, Marta Muñoz-Colmenero, Tavis Potts, Sandra Martins, Solene Avignon, Julie Maguire, Enrique De Paz, Martin Johnson, Françoise Denis, Miguel A. Pardo, Dee McElligott, and Carmen Gonzalez Sotelo. Labels on seafood products in different European countries and their compliance to EU legislation. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004218.

Prasetyo:2021:SRT

[PMM^{+}21] Andhika P. Prasetyo, Allan D. McDevitt, Joanna M. Murray, Jon Barry, Firdaus Agung, Efin Muttaqin, and Stefano Mariani. Shark and ray trade in and out of Indonesia: Addressing knowledge gaps on the path to sustainability. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003250.

Perkiss:2022:CCA

[PMN^{+}22] Stephanie Perkiss, Alistair McIlgorm, Rachel Nichols, Anna R. Lewis, Kirti K. Lal, and Michelle Voyer. Can critical accounting perspectives contribute to the development of ocean accounting and ocean governance? Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005121.

Pinillos:2023:DCZ

[PMR23] Francisco José Vázquez Pinillos, Juan Manuel Barragán Muñoz, and Francisco Ther Ríos. Delimiting coastal zones for integrated management: the case of the island and the sea of Chiloé (Chile). Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000623.

Provot:2020:UQM

[PMT^{+}20] Z. Provot, S. Mahévas, L. Tissière, C. Michel, S. Lehuta, and B. Trouillet. Using a quantitative model for participatory geo-foresight: ISIS-Fish and fishing governance in the Bay of Biscay. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1630714X.

Partelow:2020:SNC

[PN20] Stefan Partelow and Katherine Nelson. Social networks, collective action and the evolution of governance for sustainable tourism on the Gili Islands, Indonesia. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X17308734.

Phillips:2023:OSF

[PNS^{+}23] Natasha Phillips, Marianne Nyegaard, Etsuro Sawai, Ching-Tsun Chang, Miguel Baptista, and Tierney Thys. The ocean sunfishes (family Molidea): Recommendations from the IUCN Molidae review panel. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002932.

Polejack:2023:DOS

[Pol23] A. Polejack. The UN Decade of Ocean Science stages of grief — skepticism, frustration, fear of failure, and hope. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001240.

Pethiyagoda:2024:MSP

[POM24] N. A. Pethiyagoda, E. M. Ogier, and D. Hatton MacDonald. Measuring stakeholder preferences for ecosystem services interacting with salmon aquaculture in Tasmania under expansion: Application of Q methodology. Marine Policy, 162 (??):??, April 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000514.

Poonnawatt:2023:MCA

[Poo23] Kornwika Poonnawatt. Multilateral cooperation against maritime piracy in the Straits of Malacca: From the RMSI to ReCAAP. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001550.

Prodhan:2023:FIF

[PPK23] Md. Masudul Haque Prodhan, Md. Salauddin Palash, and Md. Akhtaruzzaman Khan. Factors influencing the financial and relational governance in Bangladesh’s inland fishing industry: an empirical study. Marine Policy, 148(??): ??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004936.

Polkinghorne:2024:ROC

[PPvD^{+}24] Martin Polkinghorne, Natali Pearson, Wendy van Duivenvoorde, Widya Nayati, Zainab Tahir, Nia Naelul Hasanah Ridwan, Craig Forrest, Noel Hidalgo Tan, Rachel Popelka-Filcoff, Catherine Morton, Jarrad Kowlessar, and Mark Staniforth. Reuniting orphaned cargoes: Recovering cultural knowledge from salvaged and dispersed underwater cultural heritage in Southeast Asia. Marine Policy, 163 (??):??, May 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000721.

Paniagua:2024:GGF

[PR24] Pablo Paniagua and Veeshan Rayamajhee. Governing the global fisheries commons. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24001805.

Prasada:2020:APC

[Pra20] D. V. P. Prasada. Assessing the potential for closure rules in the Sri Lankan sea cucumber fishery: Empirical models of practices and preferences of fishers. Marine Policy, 120(??): ??, October 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20303006.

Praveena:2024:EPA

[Pra24] Sarva Mangala Praveena. Exploring public awareness, influencing factors and policy implications towards microplastic pollution: Perspectives from Malaysia. Marine Policy, 161 (??):??, March 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2400040X.

Prehn:2021:CSB

[Pre21] Michael Prehn. Climate strategy in the balance who decides? Marine Policy, 131(??):??, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002323.

Prip:2022:AOG

[Pri22] Christian Prip. Arctic Ocean governance in light of an of an international legally binding instrument on the conservation and sustainable use of marine biodiversity of areas beyond national jurisdiction. Marine Policy, 142(??):??, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19304762.

Parappurathu:2020:HAP

[PRM^{+}20] Shinoj Parappurathu, C. Ramachandran, Muktha Menon, K. K. Baiju, Prathibha Rohit, R. Narayana Kumar, Shelton Padua, and Shiv Kumar. Harnessing artisanal prowess in offshore fisheries: the case of Thoothoor fishers from India. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030049X.

Pinkau:2021:SEA

[PS21] Anne Pinkau and Kerstin S. Schiele. Strategic Environmental Assessment in marine spatial planning of the North Sea and the Baltic Sea — an implementation tool for an ecosystem-based approach? Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001585.

Puszkarski:2022:IIS

[PŚ22] Jakub Puszkarski and Olga Śniadach. Instruments to implement sustainable aquaculture in the European Union. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002627.

Purnomo:2025:DDO

[PSA^{+}25] Agus Heri Purnomo, Suhendar I. Sachoemar, Zainal Arifin, Joko Samiaji, Edwarsyah, Rosye H. R. Tanjung, Juliani, Atikah Nurhayati, Pudji Purwanti, Salnida Yuniarti Lumbessy, Nickson Kawung, Lestario Widodo, Haryanti, Terry Indrabudi, Widya Safitri, Mitsutaku Makino, Juri Hori, Olivier Thebaud, Daniel K. Lew, and Fabio Boschetti. Demographic dimension of ocean perceptions: Evidence from Indonesia. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001216.

Pace:2023:EFR

[PSD23] Lisa A. Pace, Ozcan Saritas, and Alan Deidun. Exploring future research and innovation directions for a sustainable blue economy. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004808.

Purcell:2025:DCV

[PSG25] Steven W. Purcell, Stanley K. H. Shea, and Benjamin C. T. Gray. Decadal changes in value of dried sea cucumbers (bêche-de-mer) in Hong Kong markets. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004500.

Petrossian:2020:FSE

[PSMR20] Gohar A. Petrossian, Monique Sosnowski, Dana Miller, and Diba Rouzbahani. Flags for sale: an empirical assessment of flag of convenience desirability to foreign vessels. Marine Policy, 116(??):??, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306372.

Partelow:2020:JSS

[PSPR20] Stefan Partelow, Tarsila Seara, Richard B. Pollnac, and Victor Ruiz. Job satisfaction in small-scale fisheries: Comparing differences between Costa Rica, Puerto Rico and the Dominican Republic. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930898X.

Pikesley:2021:BBF

[PSTW21] Stephen K. Pikesley, Jean-Luc Solandt, Colin Trundle, and Matthew J. Witt. Benefits beyond ‘features’: Cooperative monitoring highlights MPA value for enhanced seabed integrity. Marine Policy, 134(??):??, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004127.

Pearson:2023:STF

[PT23] Natali Pearson and Benjamin S. Thompson. Saving two fish with one wreck: Maximizing synergies in marine biodiversity conservation and underwater cultural heritage protection. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001409.

Purcell:2024:WPI

[PTCC24] Steven W. Purcell, Alejandro Tagliafico, Brian R. Cullis, and Nicole Cocks. Wicked problem of improving fishery livelihoods through capacity building. Marine Policy, 163 (??):??, May 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24001064.

Pedersen:2025:AIO

[PTS^{+}25] Dane Pedersen, Owen Temby, Antonia Sohns, Evelyn Roozee, and Gordon M. Hickey. Assessing interorganizational collaboration within the transboundary network governing the conservation of Southern Resident Killer Whales. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2400469X.

Puharinen:2023:AGM

[Puh23] Suvi-Tuuli Puharinen. Achieving good marine environmental status in the EU — implications of the marine strategy framework directive for member states and blue economic activities. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002452.

Prellezo:2023:ESI

[PV23] Raúl Prellezo and Sebastian Villasante. Economic and social impacts of the landing obligation of the European Common Fisheries Policy: a review. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004845.

Pulgar:2025:EAF

[PV25] Miroslav Pulgar and Constantino Villarroel. The emergence of artisanal fishing as a strategic response in the Chilean fishing field. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005141.

Pedreschi:2021:ITC

[PVH^{+}21] Debbi Pedreschi, Audric Vigier, Hannes Höffle, Sarah B. M. Kraak, and Dave G. Reid. Innovation through consultation: Stakeholder perceptions of a novel fisheries management system reveal flexible approach to solving fisheries challenges. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309878.

Perng:2023:IST

[PWL^{+}23a] Lansing Y. Perng, John Walden, Kirsten M. Leong, Geret S. DePiper, Cameron Speir, Suzana Blake, Karma Norman, Stephen Kasperski, Mariska Weijerman, and Kirsten L. L. Oleson. Identifying social thresholds and measuring social achievement in social-ecological systems: a cross-regional comparison of fisheries in the United States. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001227.

Prince:2023:SRC

[PWL^{+}23b] Jeremy Prince, Xuefeng Wang, Kun Lin, Ani Suryanti, Sallehudin Jamon, Mudjekeewis D. Santos, Francisco SB. Torres, Nicko Amor L. Flores, Vu Viet Ha, and Tran Thi Ngoc Anh. Supporting regional co-operation in the South China Sea with an initial assessment of the datalimited skipjack tuna fishery. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002440.

Partain:2020:HOE

[PY20] Roy Andrew Partain and Constantinos Yiallourides. Hydrate occurrence in Europe: Risks, rewards, and legal frameworks. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302062.

Pyc:2023:GOG

[Pyć23] Dorota Pyć. Global ocean governance: Towards protecting the ocean’s rights to health and resilience. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200375X.

Parsmo:2024:EDG

[PYVF24] Rasmus Parsmo, Erik Ytreberg, Maarten Verdaasdonk, and Erik Fridell. Environmental discounts for Swedish ports and fairways: a ship owner perspective. Marine Policy, 159(??): ??, January 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004839.

Pauksztat:2021:ISI

[PZPB21] Birgit Pauksztat, Minghua Zhao, Claire Pekcan, and Mike Barnett. Introduction to the special issue on “International merchant shipping in the 21st century: Social science perspectives”. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310241.

Peng:2021:OFC

[PZZ^{+}21] Daomin Peng, Shucheng Zhang, Hongzhi Zhang, Dazhong Pang, Qian Yang, Ruhao Jiang, Yuting Lin, Yongtong Mu, and Yugui Zhu. The oyster fishery in China: Trend, concerns and solutions. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001354.

Qafisheh:2024:GCG

[Qaf24] Mutaz M. Qafisheh. From Constantinople to Oslo to Gaza: Developments of Palestine’s maritime legislation under colonization. Marine Policy, 160(??):??, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004876.

Quimby:2020:IDE

[QCM^{+}20] Barbara Quimby, Stephen E. S. Crook, Karly Marie Miller, Jorge Ruiz, and David Lopez-Carr. Identifying, defining and exploring angling as urban subsistence: Pier fishing in Santa Barbara, California. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20304863.

Qiu:2020:NNI

[QF20] Wenxian Qiu and Jeremy Firestone. The non-negligible influence of global sea level change on the distribution of maritime zones. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309131.

Qi:2022:CFN

[Qi22] Xiaofei Qi. The conceptual framework of the national blue economic system: a multiagent perspective. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003347.

Qi:2024:MNB

[Qi24] Xiaofei Qi. Mapping the national blue economic system space: a case study of China. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2300492X.

Quevedo:2024:SRC

[QK24] Jay Mar D. Quevedo and Ryo Kohsaka. A systematic review of cultural ecosystem services of blue carbon ecosystems: Trends, gaps, and challenges in Asia and beyond. Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004311.

Qin:2024:IPS

[QL24] Man Qin and Weixiang Lu. The influence of public sentiment on the price of aquatic products under the influence of nuclear wastewater discharge in Japan. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003828.

Quevedo:2023:ADF

[QLU^{+}23] Jay Mar D. Quevedo, Kevin Muhamad Lukman, Yaya Ihya Ulumuddin, Yuta Uchiyama, and Ryo Kohsaka. Applying the DPSIR framework to qualitatively assess the globally important mangrove ecosystems of Indonesia: a review towards evidence-based policymaking approaches. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004018.

Qi:2024:IGG

[QLZ^{+}24] Xinli Qi, Zhenfu Li, Changping Zhao, Jiaqi Li, and Qiqi Zhang. International governance of Arctic shipping black carbon emissions: Current situation and impediments to progress. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000411.

Quinio:2023:EEB

[QRK^{+}23] Louise Quinio, Malena Ripken, Thomas Klenke, Brice Trouillet, Henning Sten Hansen, and Lise Schrøder. Exploring ecosystem-based approaches in MSP through actordriven perceptual mapping. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001318.

Qin:2021:EMR

[QS21] Man Qin and Mingxue Sun. Effects of marine ranching policies on the ecological efficiency of marine ranching — Based on 25 marine ranching in Shandong Province. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003997.

Qian:2024:IIC

[QSC24] Jiangtao Qian, Kangjie Sun, and Yen-Chiang Chang. The impact of the ITLOS climate change advisory opinion on the development of international law. Marine Policy, 170(??): ??, December 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24004044.

Qu:2023:ACS

[QTBL23] Zoe Qu, Simon Thrush, Caitlin Blain, and Nicolas Lewis. Assessing the carbon storage value of kelp forest restoration in the Hauraki Gulf Marine Park, New Zealand: Lessons from no-take Marine Protected Areas. Marine Policy, 154(??):??, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002099.

Qin:2021:PIP

[QTD21] Man Qin, Qiaoan Tao, and Yuanwei Du. Policy implementation and project performance: a qualitative comparative analysis based on 29 National Marine ranchings in China. Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100138X.

Qu:2021:EVS

[QTPL21] Zoe Qu, Simon Thrush, Darren Parsons, and Nicolas Lewis. Economic valuation of the snapper recruitment effect from a well-established temperate no-take marine reserve on adjacent fisheries. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004036.

Quevedo:2021:BCE

[QUK21] Jay Mar D. Quevedo, Yuta Uchiyama, and Ryo Kohsaka. A blue carbon ecosystems qualitative assessment applying the DPSIR framework: Local perspective of global benefits and contributions. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000737.

Qiu:2021:VRF

[QWBY21] Yingying Qiu, Erda Wang, Yan Bu, and Yang Yu. Valuing recreational fishery attributes, opportunities and associated activities in China from the tourists’ satisfaction perspectives. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100227X.

Qin:2020:ECM

[QYD20] Man Qin, Caixuan Yue, and Yuanwei Du. Evolution of China’s marine ranching policy based on the perspective of policy tools. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308048.

Quan:2021:YRD

[QYX21] Yongbo Quan, Xiao Yu, and Weiwei Xu. The Yangtze River Delta integration and regional development of marine economy: Conference report. Marine Policy, 127(??): ??, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000300.

Qi:2020:WDP

[QZZ20] Xiaofei Qi, Bingxin Zhao, and Jiang-Hua Zhang. What does the product space mean for firms in the development of marine-related products? Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19306864.

Qi:2020:DNB

[QZZX20] Xiaofei Qi, Bingxin Zhao, Jianghua Zhang, and Wenwen Xiao. The drawing of a national blue product space and its evolution. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304142.

Rasheed:2020:ESP

[RA20] A. Rifaee Rasheed and Ameer Abdulla. Evaluating stakeholder participatory processes in policy development for Marine Protected Areas. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19301496.

Roll:2022:EIF

[RAB22] Kristin Helen Roll, Frank Asche, and Trond Bjørndal. The effect of introducing fuel tax to the Norwegian fishery industry. Marine Policy, 135(??):??, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004401.

Rey-Ares:2021:IWC

[RAFLRP21] Lucía Rey-Ares, Sara Fernández-López, and David Rodeiro-Pazos. Impact of working capital management on profitability for Spanish fish canning companies. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001949.

Radomska:2024:GSF

[RAMEG24] Aleksander Kozinski Radomska, Maria Elena Aramendia-Muneta, and Amaya Erro-Garcés. Spanish fishing industry within the common fishery policy. Marine Policy, 163(??): ??, May 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24001283.

Reis:2020:CCL

[RB20] Jeanette Reis and Rhoda C. Ballinger. Creating a climate for learning-experiences of educating existing and future decision-makers about climate change. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18305189.

Rozynski:2022:SQP

[RB22] Grzegorz Różyński and Małgorzata Bielecka. Sediment quality in the Polish part of the trans-boundary Vistula Lagoon: implications for deposition management. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003359.

R:2023:SLS

[RBA23] Sheeja S. R., Rahi. T. B., and Athira Ajay. State led social security and inclusion of marine fisherfolk: Analyzing the case of Kerala, India. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004390.

Ramieri:2024:DIM

[RBB^{+}24a] Emiliano Ramieri, Martina Bocci, Daniele Brigolin, Pierpaolo Campostrini, Fabio Carella, Amedeo Fadini, Giulio Farella, Elena Gissi, Fabrizio Madeddu, Stefano Menegon, Micol Roversi Monaco, Francesco Musco, Folco Soffietti, Laura Barberi, and Andrea Barbanti. Designing and implementing a multi-scalar approach to Maritime Spatial Planning: the case study of Italy. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2300444X.

Riechers:2024:CSD

[RBB^{+}24b] Maraja Riechers, Lilly Baumann, Marjan Braun, Annie Carew, Matthew Chinappa, Jasha Dehm, Alexandre Ganachaud, Elisabeth Holland, Heath Kelsey, Shilpa Lal, Mathilde Landemard, Nicolas Rocle, and Brian L. Stockwell. Contributing to sustainable development pathways in the South Pacific through transdisciplinary research: Conference report. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001222.

Razak:2022:CRR

[RBEA^{+}22] Tries B. Razak, Lisa Boström-Einarsson, Cut Aja Gita Alisa, Rindah Talitha Vida, and Timothy A. C. Lamont. Coral reef restoration in Indonesia: a review of policies and projects. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005510.

Rickard:2020:FMF

[RBNE20] Laura N. Rickard, Kofi Britwum, Caroline L. Noblet, and Keith S. Evans. Factory-made or farm fresh? Measuring U.S. support for aquaculture as a food technology. Marine Policy, 115(??):??, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302209.

Rahayu:2024:MTG

[RBTW24] Lisna Rahayu, Tim Busscher, Taede Tillema, and Johan Woltjer. Maritime transport governance challenges in the Global South. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001453.

Richmond:2022:RSC

[RC22] Laurie Richmond and Laura Casali. The role of social capital in fishing community sustainability: Spiraling down and up in a rural California port. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005455.

Roessger:2022:TTP

[RCH22] Julia Roessger, Joachim Claudet, and Barbara Horta e Costa. Turning the tide on protection illusions: the underprotected MPAs of the ‘OSPAR Regional Sea Convention’. Marine Policy, 142(??):??, August 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001567.

RodriguesdeFreitas:2022:CFE

[RCM^{+}22] Rodrigo Rodrigues de Freitas, Ronaldo Cataldo Costa, Carolina Mattosinho De Carvalho Alvite, Deisi Cristiane Balensiefer, Cecil Roberto de Maya Brotherhood de Barros, and Jonatas Henrique Prado. Challenges for fishers’ engagement in Marine Protected Areas: Lessons from Right Whale Environmental Protection Area, Southern Brazil. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001828.

Rijkure:2024:ALO

[RCM24] Astrida Rijkure, Kristine Cerbule, and Janis Megnis. Abandoned, lost, or otherwise discarded fishing gear in coastal fisheries: a case study in the Baltic Sea coastal waters of Latvia. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001854.

Rahbek-Clemmensen:2020:HAC

[RCT20] Jon Rahbek-Clemmensen and Gry Thomasen. How has Arctic coastal state cooperation affected the Arctic Council? Marine Policy, 122(??):??, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030885X.

Ryan:2021:III

[RDD^{+}21] Karina L. Ryan, Cameron J. Desfosses, Ainslie M. Denham, Stephen M. Taylor, and Gary Jackson. Initial insights on the impact of COVID-19 on boat-based recreational fishing in Western Australia. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002578.

Ruiz-Diaz:2020:SEV

[RDLA^{+}20] Raquel Ruiz-Díaz, Xiaozi Liu, Alba Aguión, Gonzalo Macho, Maite deCastro, Moncho Gómez-Gesteira, and Elena Ojea. Social-ecological vulnerability to climate change in small-scale fisheries managed under spatial property rights systems. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20303766.

Rocha:2020:MTS

[RDM^{+}20] Diana Rocha, Benjamin Drakeford, Sarah A. Marley, Jonathan Potts, Michelle Hale, and Angie Gullan. Moving towards a sustainable cetacean-based tourism industry — a case study from Mozambique. Marine Policy, 120(??): ??, October 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19305767.

Reimer:2023:CRM

[RDT^{+}23] Julie M. Reimer, Rodolphe Devillers, Brice Trouillet, Natalie C. Ban, Tundi Agardy, and Joachim Claudet. Conservation ready marine spatial planning. Marine Policy, 153 (??):??, July 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001823.

Rajapakse:2021:UFW

[RE21] Asanka Rajapakse and Gholam Reza Emad. Underlying factors which cause task deviation leading to dangerous situations at sea. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001597.

Rajapakse:2023:FUP

[RE23] Asanka Rajapakse and Gholam Reza Emad. Fatigue, an unsolved puzzle that continues contributing to accidents at sea. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002786.

Reedy:2020:NAS

[Ree20] Katherine Reedy. Neoliberal Aleutians: Seeing Like a fishing company, Seeing Like a coastal community. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20301020.

Ren:2021:HRS

[Ren21a] Wenhan Ren. How to realize the sustainable development of mariculture industry — re-examine from the perspective of biased technological progress in China. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004024.

Ren:2021:RDC

[Ren21b] Wenhan Ren. Research on dynamic comprehensive evaluation of allocation efficiency of green science and technology resources in China’s marine industry. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002487.

Rector:2025:DCD

[RF25] Megan E. Rector and Ramón Filgueira. Different coasts for different folks: Place-based community values and experience mediate social acceptability of low-trophic aquaculture. Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004457.

Rector:2024:RSA

[RFG24] M. E. Rector, R. Filgueira, and J. Grant. The role of salmon aquaculture eco-certification in corporate social responsibility and the delivery of ecosystem services and disservices. Marine Policy, 160(??):??, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004815.

Ribeiro:2020:STL

[RFPS20] Marta Chantal Ribeiro, Rui Ferreira, Eliana Pereira, and Joana Soares. Scientific, technical and legal challenges of deep sea mining. A vision for Portugal — Conference report. Marine Policy, 114:??, April 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18307346.

Ruslan:2022:MES

[RGH^{+}22] Nur Fatin Nabilah Ruslan, Hong Ching Goh, Caroline Hattam, Andrew Edwards-Jones, and Heng Hing Moh. Mangrove ecosystem services: Contribution to the well-being of the coastal communities in Klang Islands. Marine Policy, 144(??):??, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200269X.

Robertson:2020:LMM

[RGK^{+}20] Thomas Robertson, Suzie Greenhalgh, Isoa Korovulavula, Tomasi Tikoibua, Pio Radikedike, and Philip Stahlmann-Brown. Locally managed marine areas: Implications for socio-economic impacts in Kadavu, Fiji. Marine Policy, 117 (??):??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19309601.

Roberts:2020:EPM

[RHC20] Kelsey E. Roberts, Olivia Hill, and Carly N. Cook. Evaluating perceptions of marine protection in Australia: Does policy match public expectation? Marine Policy, 112(??): ??, February 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19300375.

Rangel:2025:ELC

[RHG^{+}25] Mafalda Rangel, Barbara Horta e Costa, Ma. Helena Guimarães, Adriana Ressurreição, Pedro Monteiro, Frederico Oliveira, Luís Bentes, Nuno Sales Henriques, Inês Sousa, Sofia Alexandre, João Pontes, Carlos ML Afonso, Adela Belackova, Ana Marçalo, Mariana Cardoso-Andrade, António Cortês, António José Correia, Vanda Lobo, Emanuel J. Gonçalves, Tiago Pitta e Cunha, and Jorge MS Gonçalves. Engaging and legitimizing communities: codesigning a community-based Marine Protected Area. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001101.

Rutherford:2020:CAA

[RHL20] V. E. Rutherford, J. M. Hills, and M. D. A. Le Tissier. Comparative analysis of adaptation strategies for coastal climate change in North West Europe. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X16304262.

Rubio:2024:APC

[RHO24] Iratxe Rubio, Alistair J. Hobday, and Elena Ojea. Adaption planning to climate change in industrial fisheries: Progress in the Basque tropical tuna fishery. Marine Policy, 161(??): ??, March 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23005341.

Rybicki:2021:STI

[RHST21] Sandra Rybicki, Katell G. Hamon, Sarah Simons, and Axel Temming. Spatio-temporal interactions between Northeast Atlantic Mackerel and its fishery — Simulating different futures. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003511.

Ren:2021:HDE

[RJ21] Wenhan Ren and Jianyue Ji. How do environmental regulation and technological innovation affect the sustainable development of marine economy: New evidence from China’s coastal provinces and cities. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21000798.

Rodriguez:2023:APO

[RJAT23] Richard Castillo Rodríguez, Jorge Jiménez, Katherine Arroyo Arce, and Katie R. Thompson. An analysis of policy options available to the International Maritime Organization to protect the Costa Rica Thermal Dome: Building the case for a particularly sensitive sea area. Marine Policy, 148(??):??, February 2023. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004225.

Raabe:2024:WAY

[RKC^{+}24] Nyree L. Raabe, Tanya J. King, Adam P. A. Cardilini, Kelly K. Miller, and William T. Borrie. Who asked you? A systematic review exploring the origin of primary research into attitudes to marine protected areas in Australia. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002136.

Ripken:2024:CMS

[RKKM24] Malena Ripken, Xander Keijser, Thomas Klenke, and Igor Mayer. Conceptualizing maritime spatial planning — a pilot study on viewpoints across Northern European Seas. Marine Policy, 166(??):??, August 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001799.

Rahman:2020:ITM

[RKR20] Md Sadique Rahman, Mohammad Mizanul Haque Kazal, and Shah Johir Rayhan. Impacts of the training of mud crab farmers: an adaptation strategy to cope with salinity intrusion in Bangladesh. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20305856.

Ruskule:2023:TCG

[RKS^{+}23] Anda Ruskule, Jonne Kotta, Champa Rani Saha, Philipp Arndt, Didzis Ustups, Solvita Strāķe, and Lena Bergström. Testing the concept of green infrastructure at the Baltic Sea scale to support an ecosystem-based approach to management of marine areas. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004213.

Roozee:2024:MIO

[RKS^{+}24] Evelyn Roozee, Dongkyu Kim, Antonia Sohns, Jasper R. de Vries, Owen F. Temby, and Gordon M. Hickey. Managing inter-organizational trust and risk perceptions in transboundary fisheries governance networks. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004608.

Ruff:2024:LST

[RL24] Elizabeth O. Ruff and Sarah E. Lester. Leaving seafood on the table: Pilot pathways a missed opportunity for U.S. mariculture. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400280X.

Rojon:2021:ICP

[RLRS21] Isabelle Rojon, Nicholas-Joseph Lazarou, Nishatabbas Rehmatulla, and Tristan Smith. The impacts of carbon pricing on maritime transport costs and their implications for developing economies. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002645.

Reed:2020:DSM

[RLS20] Jodie R. Reed, Amanda T. Lombard, and Kerry J. Sink. A diversity of spatial management instruments can support integration of fisheries management and marine spatial planning. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308292.

Reid:2021:SNZ

[RM21a] Oscar Schouten Reid and Fabien Medvecky. Sharks and New Zealand news media. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003626.

Romero:2021:CTU

[RM21b] Pedro Romero and Oscar Melo. Can a Territorial Use Right for Fisheries management make a difference for fishing communities? Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310101.

Rosas-Munoz:2022:WCW

[RMBR22] Juan Rosas-Muñoz and Marjorie Baquedano-Rodríguez. What can we learn about the natural disasters and multidimensional poverty link for artisanal fishery in Chile? Marine Policy, 144(??):??, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200272X.

Rechberger:2025:PPC

[RMBS25] Kristin D. Rechberger, Juan Mayorga, Mara Booth, and Enric Sala. A pathway to protect 30% of coastal waters by 2030. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001885.

Racine:2021:CSA

[RMF^{+}21] Phoebe Racine, AnnaClaire Marley, Halley E. Froehlich, Steven D. Gaines, Ian Ladner, Ilan MacAdam-Somer, and Darcy Bradley. A case for seaweed aquaculture inclusion in U.S. nutrient pollution management. Marine Policy, 129 (??):??, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001172.

Roberts:2022:PCR

[RMU^{+}22] Nicky Roberts, Buchari Mengge, Muh. Rifadly Utina, Farhan Muhatar, Anugerah, Arham Icwardanhi, R. Muhammad Zulkifli, and Austin Humphries. Patron-client relationships shape value chains in an Indonesian island-based fisheries system. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001890.

Rojas-Nazar:2022:VMR

[RNGRC22] Ursula A. Rojas-Nazar, Jonathan P. A. Gardner, Agnes Rouchon, and Ross Cullen. Valuing marine reserves: a case study from two locations in central New Zealand. Marine Policy, 140(??):??, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000987.

Rogiers:2024:NCW

[Rog24] Frederik Rogiers. A new cold war at sea: Fishing rights vs. the proliferation of military activities in the EEZ. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000757.

Rosello:2022:RFM

[Ros22] Mercedes Rosello. Regional fishery management organisation measures and the imposition of criminal and administrative sanctions in respect of high seas fishing. Marine Policy, 144(??):??, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002603.

Rossiter:2025:CRR

[Ros25] Ash Rossiter. Cable risk and resilience in the age of uncrewed undersea vehicles (UUVs). Marine Policy, 171(??): ??, January 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24004342.

Relano:2023:CLP

[RP23a] Veronica Relano and Daniel Pauly. Corrigendum to “The ‘Paper Park Index’: Evaluating marine protected area effectiveness through a global study of stakeholder perceptions” [Mar. Policy 151 (2023) 105571]. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2300235X. See [RP23b].

Relano:2023:LPP

[RP23b] Veronica Relano and Daniel Pauly. The ‘Paper Park Index’: Evaluating Marine Protected Area effectiveness through a global study of stakeholder perceptions. Marine Policy, 151 (??):??, May 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000982. See cor-

rigendum [RP23a].

Raut:2023:AAF

[RPK23] Shrutika Raut, S. Prakash, and Amit Kumar. Assessing artisanal fishers’ attitude and perception towards electric rays conservation along Indian coast. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003597.

Rossi:2024:POI

[RPY^{+}24] Valentina Rossi, Carlo Pipitone, Katherine L. Yates, Fabio Badalamenti, Giovanni D’Anna, Cristina Pita, Fátima L. Alves, Jesús E. Argente-García, Jelena Basta, Joachim Claudet, Karsten Dahl, Simonetta Fraschetti, Ioannis Giovos, Peter Mackelworth, Mairi Maniopoulou, Vasiliki Markantonatou, Márcia Marques, Pedro Noguera-Méndez, Joanna Piwowarczyk, Violin Raykov, Gil Rilov, Bob Rumes, Alicia Said, María Semitiel-García, Yael Teff-Seker, Tomás Vega Fernández, and David Goldsborough. Poor online information on European marine protected areas impairs public participation under the Aarhus Convention. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000101.

Roman:2020:BRD

[RQB^{+}20] Catalina Román, Dante Queirolo, Claudio Bernal, Marcelo A. San Martín, and Victoria Escobar. Bycatch reduction device in the Chilean hake (Merluccius gayi) bottom trawl fishery, using a rigid separation grid. Marine Policy, 116 (??):??, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X1830873X.

Reid:2020:IEB

[RR20] John Reid and Matthew Rout. The implementation of ecosystem-based management in New Zealand — A Māori perspective. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305688.

Rout:2024:IBE

[RRM^{+}24] Matthew Rout, John Reid, Jason Paul Mika, Jay Whitehead, Annemarie Gillies, Fiona Wiremu, Georgia McLellan, and Corey Ruha. Indigenising the blue economy in Aotearoa New Zealand. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005201.

Robertson:2023:PSN

[RROP23] Kerrie Robertson, Georgina Rood, Metua Okotai, and Mariata Pittman. Purse seiner non grata: Privy Council wades into Pacific fisheries. Marine Policy, 156(??):??, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003342.

Rakotomahazo:2023:EPI

[RRR^{+}23] Cicelin Rakotomahazo, Nirinarisoa Lantoasinoro Ranivoarivelo, Jacqueline Razanoelisoa, Gildas Georges Boleslas Todinanahary, Eulalie Ranaivoson, Mara Edouard Remanevy, Lalao Aigrette Ravaoarinorotsihoarana, and Thierry Lavitra. Exploring the policy and institutional context of a Payment for Ecosystem Services (PES) scheme for mangroves in southwestern Madagascar. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004973.

Reimer:2025:MAC

[RRS25] Matthew N. Reimer, Anthony Rogers, and James N. Sanchirico. Managing for adaptive capacity in climateready fisheries. Marine Policy, 174(??):??, April 2025.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000168.

RoseteVerges:2024:CLE

[RS24] Fernando A. Rosete Verges and F. Javier Sanz Larruga. Coastal laws evolution in Spain: an analysis from the public policy cycle. Marine Policy, 162(??):??, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000502.

Ramadhan:2022:HDD

[RSAP22] Andrian Ramadhan, Wilmar A. Salim, Teti Armiati Argo, and Puji Prihatiningsih. The human dimension dilemma in marine spatial planning. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200344X.

Riisager-Simonsen:2022:MNB

[RSFvH^{+}22] Christian Riisager-Simonsen, Gianna Fabi, Luc van Hoof, Noél Holmgren, Giovanna Marino, and Dennis Lisbjerg. Marine nature-based solutions: Where societal challenges and ecosystem requirements meet the potential of our oceans. Marine Policy, 144(??):??, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002457.

Ryan:2025:RFA

[RSH^{+}25] Karina L. Ryan, Clinton Syers, Kane Holtom, Timothy Green, Jeremy M. Lyle, Kate E. Stark, and Sean R. Tracey. Recreational fishers’ attitudes to fisheries management and compliance. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004834.

R:2025:ICP

[RSK^{+}25] Anuja A R, Shyam S. Salim, Akhila Kumaran, Narayanakumar R., Jayasankar J., and Aswathy N. The impact of the COVID-19 pandemic on the marine fisheries sector in India. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001915.

Rincon-Sandoval:2024:PIR

[RSLR24] Luis A. Rincón-Sandoval and Jorge A. López-Rocha. Performance indicators of the red grouper (Epinephelus morio) fishery along the Yucatan coast, Southeast Mexico. Marine Policy, 169(??):??, November 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003312.

Rehmatulla:2025:RRP

[RSN25] Nishatabbbas Rehmatulla, Kelly Schwarz, and Fatemeh Habibi Nameghi. Revisiting the role of private voluntary standards and initiatives in decarbonising shipping. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001228.

Rees:2020:ETS

[RSS^{+}20] Siân E. Rees, Emma V. Sheehan, Bryce D. Stewart, Robert Clark, Thomas Appleby, Martin J. Attrill, Peter J. S. Jones, David Johnson, Natasha Bradshaw, Simon Pittman, Jenny Oates, and Jean-Luc Solandt. Emerging themes to support ambitious UK marine biodiversity conservation. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930199X.

Romero-Tenorio:2022:IBS

[RTMCVMDE22] Arturo Romero-Tenorio, Manuel Mendoza-Carranza, Javier Francisco Valle-Mora, and Alberto Delgado-Estrella. Interactions between small-scale fisheries and marine mammals from the perspective of fishers in the Mexican tropical Pacific Coast. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000306.

Raubenheimer:2020:RGG

[RU20] Karen Raubenheimer and Niko Urho. Rethinking global governance of plastics — The role of industry. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309698.

Ruban:2024:BRA

[Rub24] Dmitry A. Ruban. Book review: Amanda Reichelt-Brushett (Ed.). Marine Pollution — Monitoring, Management and Mitigation, Springer, Cham (2023), xv + 369 pp., ISBN 978-3-031-10126-7, Hardback, EUR 49.99. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23005237.

Runolfsson:2024:MQP

[Run24] Birgir Th. Runolfsson. Measuring quality of property rights: Development of user rights quality in the Icelandic fisheries. Marine Policy, 160(??):??, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004955.

Russi:2020:GSS

[Rus20] Daniela Russi. Governance strategies for a successful marine protected area — The case of Torre Guaceto. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19301770.

Richardson:2021:CMA

[RWVH21] Kelsey Richardson, Chris Wilcox, Joanna Vince, and Britta Denise Hardesty. Challenges and misperceptions around global fishing gear loss estimates. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001330.

Ren:2023:DHT

[RXN23] Wenhan Ren, Yuhan Xu, and Jing Ni. Development, hotspots and trend directions of marine ecological compensation in China: a bibliometric-based assessment. Marine Policy, 156(??):??, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003482.

Ren:2024:RIM

[RXX24] Wenhan Ren, Yuhan Xu, and Han Xiao. Research on the impact of marine ecological civilization demonstration zone policies on the green development level of China’s marine economy: a quasi natural experiment based on coastal cities. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000460.

Ren:2025:MHQ

[RYL25] Wei Ren, Bei Yuan, and Qixu Liang. Mariculture highquality development process and spatial differences: a case study of China. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004263.

Retnoningtyas:2021:SPM

[RYS^{+}21] Heidi Retnoningtyas, Irfan Yulianto, Arisetiarso Soemodinoto, Yudi Herdiana, Tasrif Kartawijaya, Mohamad Natsir, and Joko Tri Haryanto. Stakeholder participation in management planning for grouper and snapper fisheries in West Nusa Tenggara Province, Indonesia. Marine Policy, 128(??):??, June 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000634.

Rajagopal:2021:HWU

[RZ21a] Sriram Rajagopal and Pengfei Zhang. How widespread is the usage of the Northern Sea Route as a commercially viable shipping route? A statistical analysis of ship transits from 2011 to 2018 based on empirical data. Marine Policy, 125(??):??, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309453.

Ren:2021:GTP

[RZ21b] Wenhan Ren and Qi Zeng. Is the green technological progress bias of mariculture suitable for its factor endowment? — — Empirical results from 10 coastal provinces and cities in China. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2030988X.

Scherer:2021:MEC

[SA21] Marinez E. G. Scherer and Milton L. Asmus. Modeling to evaluate coastal governance in Brazil. Marine Policy, 129 (??):??, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001123.

Standal:2025:TCF

[SA25] Dag Standal and Ina Helene Ahlquist. Transforming coastal fisheries in Norway: Institutional implications of green technology implementation. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24004718.

Santos:2021:MCD

[SAA^{+}21] Catarina Frazão Santos, Tundi Agardy, Francisco Andrade, Larry B. Crowder, Charles N. Ehler, and Michael K. Orbach. Major challenges in developing marine spatial planning. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306213.

Stokes:2020:BUP

[SAB^{+}20] Debra Stokes, Kirin Apps, Paul A. Butcher, Betty Weiler, Hanabeth Luke, and Andrew P. Colefax. Beach-user perceptions and attitudes towards drone surveillance as a sharkbite mitigation tool. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2030347X.

Suradja:2024:ALC

[SAB^{+}24] Ivana Suradja, Theresia Susanti Asih, Basir, Dysi Polite Dyspriani, Hilda Lionata, Aby Marsiyono, Bambang Adhitya Nugraha, Tiara Intan Palupi, Indra Adi Putra Salam, Liliana Sierra Castillo, Asri Toldo, Andi Trisnawati, and Erendira Aceves-Bueno. Alternating livelihoods and coping with shocks: an examination of coastal tourism in Indonesia amidst COVID-19. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24003774.

Sole:2025:NES

[SAC25] Liliana Solé, Eduard Ariza, and Antònia Casellas. Navigating ecosystem services and access inequalities. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25002143.

Sun:2025:CAS

[SADP25] Ruikai Sun, Wessam Abouarghoub, Emrah Demir, and Andrew Potter. A comprehensive analysis of strategies for reducing GHG emissions in maritime ports. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2400455X.

Saha:2022:EEA

[Sah22] Choyon Kumar Saha. Emergence and evolution of aquaculture sustainability certification schemes. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002433.

Sakita:2021:CUD

[Sak21] Seishiro Sakita. Centralization under decentralization: the development of fishery clubs in Lesvos under the administrative reforms of Greece. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002669.

Sampson:2022:LBS

[Sam22] Helen Sampson. ‘Beyond the state’: the limits of international regulation and the example of abandoned seafarers. Marine Policy, 140(??):??, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000938.

Sanlier:2020:APS

[Şan20]  Şengül Şanlıer. Analysis of port state control inspection

data: the Black Sea Region. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19305469.

Santos:2025:IMS

[San25] Thauan Santos. Integrating marine spatial planning into blue economy policies in Latin America. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25002726.

Sagoe:2021:CPA

[SAO^{+}21] Alberta Ama Sagoe, Denis Worlanyo Aheto, Isaac Okyere, Richard Adade, and Justice Odoi. Community participation in assessment of fisheries related ecosystem services towards the establishment of marine protected area in the Greater Cape Three Points area in Ghana. Marine Policy, 124(??): ??, February 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309854.

Straume:2024:SGN

[SAO^{+}24] Hans-Martin Straume, Frank Asche, Atle Oglend, Ivar Gaasland, Ingrid K. Pettersen, and Geir Sogn-Grundvåg. The structure of Norwegian seafood trade. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X23004542.

Sar:2023:CAR

[Sar23] Aysegul Bugra Sar. Considerations on assistance and rescue at sea in the light of the increasing autonomy in shipping. Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001665.

Simantiris:2024:YMI

[SAS24] Nikolaos Simantiris, Konstantina Andreanidou, and George Sampson. Over 30 years of monitoring and implementing the Bern Convention’s recommendations for the protection of Mediterranean sea turtles. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24003178.

Sasaki:2025:SCJ

[Sas25a] Takafumi Sasaki. The significance and challenges of the Japan–Taiwan fisheries agreement in avoiding conflicts and promoting sustainable fisheries in the East China sea: Focusing on the background, negotiation process, and operational realities of the agreement. Marine Policy, 180(??): ??, October 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25001599.

Sistiaga:2025:EGO

[SAS^{+}25b] Manu Sistiaga, Neil Anders, Jostein Saltskår, Hermann Pettersen, Michael Breen, Odd-Børre Humborstad, Hector Peña, Pau Muñoz-Benavent, Joaquín Martínez-Peiró, Gabriela Andreu-García, Victor Espinosa, Stein Harris Olsen, Tonje Kristin Jensen, Keno Ferter, Leif Nøttestad, and Svein Løkkeborg. Exploiting the “gold of the ocean”: Can live storage solve the paradox of the purse seine fishery for Atlantic bluefin tuna in Norway? Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24005062.

Sylvester:2020:PAT

[SB20] Zephyr T. Sylvester and Cassandra M. Brooks. Protecting Antarctica through co-production of actionable science: Lessons from the CCAMLR marine protected area process. Marine Policy, 111(??):??, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305354.

Sam:2025:CFS

[SB25] Theodora Sam and Melania Borit. Conceptualizing fishery systems: an analysis of definitions. Marine Policy, 181(??): ??, November 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25002362.

Schwoerer:2023:FFF

[SBD^{+}23] Tobias Schwoerer, Kevin Berry, Darcy G. Dugan, David C. Finnoff, Molly Mayo, Jan Ohlberger, and Eric J. Ward. Fish or not fish — fisheries participation and harvest diversification under economic and ecological change. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003664.

Siddiqi:2024:UPC

[SBDC24] Aliya C. Siddiqi, Brock J. Bergseth, Amy Diedrich, and Andrew Chin. Understanding the perceived conservation benefits of shark-marine tourism in the Global South. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005286.

Seminara:2023:AFK

[SBFHD^{+}23] Cecilia Inés Seminara, Márcio Luiz Vargas Barbosa-Filho, Rachel Ann Hauser-Davis, Gabriel Barros Gonçalves Souza, Marcela Alvarenga, Christiana Cabicieri Profice, and Salvatore Siciliano. Artisanal fisher knowledge on the impacts of destructive and illegal practices on the Southern Coast of Bahia, Brazil. Marine Policy, 158(??):??, December 2023.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003950.

Seguigne:2024:SPD

[SBG^{+}24] Clémentine Séguigne, Mark E. Bond, Adeline Goyaud, Michael R. Heithaus, Gilles Siu, Frédéric Torrente, and Eric Clua. Stakeholder perception of the danger posed by sharks in the world’s largest shark sanctuary. Marine Policy, 164 (??):??, June 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24001258.

Seguigne:2025:KPO

[SBG^{+}25] Clémentine JM Séguigne, Mark E. Bond, Adeline Goyaud, Michael R. Heithaus, Gilles Siu, Zenya Rowe, Frédéric Torrente, and Eric Clua. Knowledge, perception and ownership by local stakeholders among the world’s largest shark sanctuary a decade after its inception. Marine Policy, 173(??): ??, March 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2400527X.

Schulz:2024:NAC

[SBGR24] Veronica Schulz, Michael G. H. Bell, D. Glenn Geers, and John Matthew Rose. A new age for coal ports: an historical perspective on diversification and a circular economy future. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004007.

Shumway:2021:PSF

[SBJF^{+}21] Nicole Shumway, Justine Bell-James, James A. Fitzsimons, Rose Foster, Chris Gillies, and Catherine E. Lovelock. Policy solutions to facilitate restoration in coastal marine environments. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004000.

Schoderer:2025:OSO

[SBM^{+}25a] Mirja Schoderer, Henry Bittig, Karel Castro Morales, Leticia Cotrim da Cunha, Claas Faber, Ramona Hägele, Birgit Klein, Arne Körtzinger, Carlos Musetti, Raquel Oliveira, Tobias Steinhoff, Cathy Wimart-Rousseau, Matthias Wunsch, and Anna-Katharina Hornidge. Open science for the ocean. Recommendations from the perspective of marine carbon observations in Germany, Brazil, and beyond. Marine Policy, 178(??):??, August 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001514.

Seijo:2025:FSM

[SBM^{+}25b] Javier Seijo, Carlota Barañano, Yolanda Molares, Celia Olabarría, María Elsa Vázquez Otero, and Sebastián Villasante. Fishing with science: Measures to mitigate the impacts of climate change on small-scale fisheries in Galicia (Spain). Marine Policy, 175(??):??, May 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004585.

Schwartz-Belkin:2024:EBI

[SBP24] Inbar Schwartz-Belkin and Michelle E. Portman. Exploring barriers to the implementation of geospatial technologies in marine spatial planning: Reports from practitioners. Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004244.

SalejHiggins:2025:SCW

[SBP^{+}25] Silvio Salej Higgins, Jorge Alexandre Barbosa Neves, Fernanda Pacheco Huguenin, Jucimar Eugenio Da Costa, and Geraldo Timóteo. Social capital and women’s earnings in traditional fisheries: Evidence from the Campos Basin (State of Rio de Janeiro, Brazil). Marine Policy, 172(??): ??, February 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24005049.

Sinan:2021:DPI

[SBS21] Hussain Sinan, Megan Bailey, and Wilf Swartz. Disentangling politics in the Indian Ocean Tuna Commission. Marine Policy, 133(??):??, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003924.

Smythe:2021:ORI

[SBT21] Tiffany Smythe, David Bidwell, and Grant Tyler. Optimistic with reservations: the impacts of the United States’ first offshore wind farm on the recreational fishing experience. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000506.

Soares:2023:DVN

[SBT^{+}23] Marcelo O. Soares, Daysiane Barbosa Brandão, Carlos Eduardo Peres Teixeira, Rivelino Martins Cavalcante, André Henrique Barbosa de Oliveira, Luis Ernesto Arruda Bezerra, and Eduardo Lacerda Barros. Déjà vu: New oil spill poses cumulative risks to protected coastal environments in the South Atlantic. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2300297X.

Solis:2021:MEA

[SBW^{+}21] Daniel Solís, Kathryn Bisack, John Walden, Paul Richards, and Juan Agar. Measuring the economic abatement cost of sea turtle bycatch in the Northwest Atlantic commercial pelagic longline fishery. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003456.

Schaar:2021:FHM

[SC21] Siena I. Schaar and Linda J. Cox. The future for Hawai’i’s marine aquarium fishery: a cost benefit analysis compared to an environmental impact assessment. Marine Policy, 127 (??):??, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000397.

Shen:2022:IFS

[SC22] Chuyi Shen and Tinggui Chen. Impact of fuel subsidies on bottom trawl fishery operation in China. Marine Policy, 138 (??):??, April 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000240.

Song:2024:SVL

[SC24] Annie Young Song and Hubert Cheung. Seizing a venue linking opportunity: China’s strategy to advance its sea cucumber interests in global environmental governance. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0308597X24003439. See corrigendum [SC25].

Song:2025:CSV

[SC25] Annie Young Song and Hubert Cheung. Corrigendum to “Seizing a venue linking opportunity: China’s strategy to advance its sea cucumber interests in global environmental governance” [Mar. Policy 169 (2024) 106345]. Marine Policy, 177(??):??, July 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001137. See [SC24].

Siano:2020:CPM

[SCA^{+}20] R. Siano, A. Chapelle, V. Antoine, E. Michel-Guillou, F. Rigaut-Jalabert, L. Guillou, H. Hégaret, A. Leynaert, and A. Curd. Citizen participation in monitoring phytoplankton seawater discolorations. Marine Policy, 117(??): ??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X16306169.

Stratoudakis:2022:OSD

[SCA^{+}22] Yorgos Stratoudakis, Miguel Caetano, Cláudia Afonso, Cátia Bartilloti, Antonina dos Santos, Pedro Terrinha, Ivone Figueiredo, Miguel Miranda, José Lino Costa, Isabel Caçador, Vanessa Fonseca, Paula Chainho, Ricardo Melo, Bernardo Quintella, Maria José Costa, André B. Fortunato, Paula Freire, Marta Rodrigues, Luís Portela, Miguel Bessa Pacheco, Joaquim Pombo, Carla Palma, Anabela Oliveira, António Gonçalves Tavares, Filipe Porteiro, Inês Metelo, José Sá Fernandes, Maria da Conceição Freitas, Henrique Cabral, Marina Cunha, Christopher Pham, and Inês Carvalho. One step from the deep ocean: Linking research and governance for the Tagus river estuary and the Lisbon canyon system — a conference report. Marine Policy, 144(??):??, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002718.

Smythe:2025:CCS

[SCB25] Tiffany Smythe, Letty Cass, and David Bidwell. Conflict and coexistence: Small operators’ multi-use experiences in Rhode Island’s blue economy. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24005451.

Seidu:2022:EFS

[SCBB^{+}22] Issah Seidu, Francoise Cabada-Blanco, Lawrence K. Brobbey, Berchie Asiedu, Paul Barnes, Moro Seidu, and Nicholas K. Dulvy. “Every fish in the sea is meat and so are guitarfishes”: Socio-economic drivers of a guitarfish fishery in Ghana. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002068.

Stratoudakis:2020:IPM

[SCBdA20] Yorgos Stratoudakis, C. Correia, A. F. Belo, and P. R. de Almeida. Improving participated management under poor fishers’ organization: Anadromous fishing in the estuary of Mondego River, Portugal. Marine Policy, 119(??): ??, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20300427.

Solandt:2025:DEE

[SCC25] Jean-Luc Solandt, Robert Clark, and Sarah Coulthard. Delivering effective and equitably governed marine protected area networks in the UK: the role of Inshore Fisheries and Conservation Authorities (IFCAs). Marine Policy, 171(??): ??, January 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24004494.

Sjostrom:2021:GLE

[SCCW21] Anja J. C. Sjostrom, Lorenzo Ciannelli, Flaxen Conway, and W. Waldo Wakefield. Gathering local ecological knowledge to augment scientific and management understanding of a living coastal resource: the case of Oregon’s nearshore groundfish trawl fishery. Marine Policy, 131(??): ??, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002281.

Sbragaglia:2021:RUS

[SCD21] Valerio Sbragaglia, Ricardo A. Correia, and Enrico Di Minin. Responsible use of social media data is needed: a reply to Maya-Jariego et al. “Plenty of black money: Netnography of illegal recreational underwater fishing in southern Spain”. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003912.  See

[MJMAA21, MJMAA22].

Snook:2022:JBY

[SCF^{+}22] Jamie Snook, Ashlee Cunsolo, James Ford, Chris Furgal, Andria Jones-Bitton, and Sherilee Harper. “Just because you have a land claim, that doesn’t mean everything’s going to fall in place”: an Inuit social struggle for fishery access and well-being. Marine Policy, 140(??):??, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200118X.

Shih:2020:COF

[SCGC20] Yi-Che Shih, Yen-Chiang Chang, Warwick Gullett, and Wen-Yan Chiau. Challenges and opportunities for fishery rights negotiations in disputed waters — a Taiwanese perspective regarding a fishing boat case incident. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301289.

Schatz:2024:PNC

[Sch24a] Valentin Schatz. Provisions for nullification of conservation and management measures in RFMO objection procedures. Marine Policy, 166(??):??, August 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002288.

Smith:2024:DPC

[SCH^{+}24b] M. Luke Smith, Guangqing Chi, Hannah Hennighausen, Davin Holen, and E. Lance Howe. Differences in perceptions of COVID-19 risks in a fishing community in Alaska, 2020–2021. Marine Policy, 161(??):??, March 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000435.

Schreiber:2022:BJC

[SCJ22] Milena Arias Schreiber, Ratana Chuenpagdee, and Svein Jentoft. Blue Justice and the co-production of hermeneutical resources for small-scale fisheries. Marine Policy, 137 (??):??, March 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000069.

Soto:2021:CSF

[SCLM^{+}21] Doris Soto, Carlos Chávez, Jorge León-Muñoz, Carol Luengo, and Yuri Soria-Galvarro. Chilean salmon farming vulnerability to external stressors: the COVID 19 as a case to test and build resilience. Marine Policy, 128(??): ??, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100097X.

Sherley:2025:RCA

[SCM^{+}25] Richard B. Sherley, Robert J. M. Crawford, Alistair M. McInnes, Lynne J. Shannon, William J. Sydeman, and Lauren J. Waller. Rejoinder to: Cochrane et al., Errors and bias in marine conservation and fisheries literature: Their impact on policies and perceptions [Mar. Policy 168 (2024) 106329]. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005165. See [CBH^{+}24a] and re-

sponse [CBH^{+}25].

Sierra-Correa:2024:ICZ

[SCRCI24] Paula Cristina Sierra-Correa, Ángela Cecilia López Rodríguez, David Alejandro Alonso Carvajal, and Francisco Armando Arias Isaza. Integrated coastal zone management in Colombia: 25 years of implementing an environmental planning policy. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001209.

Scott:2022:WCE

[SCSR^{+}22] Molly Scott, Edward Cardona, Kaylee Scidmore-Rossing, Mark Royer, Jennifer Stahl, and Melanie Hutchinson. What’s the catch? Examining optimal longline fishing gear configurations to minimize negative impacts on nontarget species. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002330.

Scott:2023:CWC

[SCSR^{+}23] Molly Scott, Edward Cardona, Kaylee Scidmore-Rossing, Mark Royer, Jennifer Stahl, and Melanie Hutchinson. Corrigendum to “What’s the catch? Examining optimal longline fishing gear configurations to minimize negative impacts on non-target species” [Mar. Policy 143 (2022) 105186]. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22005164.

Sun:2024:GCC

[SCY24] Ruihong Sun, Yue Chu, and Xinliang Ye. China’s cruise tourism policy evolution and effectiveness analysis. Marine Policy, 159(??):??, January 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004268.

Sun:2024:CPD

[SCZ^{+}24] Peng Sun, Cailing Cai, Yifan Zhang, Daniel WanKit Yip, and Aysan Esmradi. A confidentiality preserved data sharing framework for decision support in the maritime law enforcement. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002422.

Santos:2024:GRJ

[SdACdSLdAS24] Thauan Santos, Joilson de Assis Cabral, Paulo Vitor dos Santos Lima, and Matheus de Andrade Santos. Rio de Janeiro’s ocean economy as a key vector for sustainable development in Brazil. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004098.

Soares:2022:ICV

[SdCMN^{+}22] Júlia Benevenuti Soares, Marcus Rodrigues da Costa, Cassiano Monteiro-Neto, Luciana Loto, Maurício Düppré de Abreu, and Rafael de Almeida Tubino. Impacts of COVID-19 on the value chain of a small-scale fishery system in a tropical metropolitan city. Marine Policy, 140 (??):??, June 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001154.

Sletten:2021:BBH

[SDG^{+}21] Jennifer Sletten, Mimi D’Iorio, Mary G. Gleason, Alex Driedger, Timothé Vincent, Claire Colegrove, Dawn Wright, and Virgil Zetterlind. Beyond the boundaries: How regulation-centered marine protected area information improves ocean protection assessments. Marine Policy, 124(??):??, February 2021. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309908.

Sletten:2022:LCB

[SDG^{+}22] Jennifer Sletten, Mimi D’Iorio, Mary G. Gleason, Alex Driedger, Timothé Vincent, Claire Colegrove, Dawn Wright, and Virgil Zetterlind. Corrigendum to “Beyond the boundaries: How regulation-centered marine protected area information improves ocean protection assessments” [124, 2021, 104340]. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002500.

Sala:2023:TIR

[SDG^{+}23] Antonello Sala, Jochen Depestele, Aysun Gümüş, Pascal Laffargue, J. Rasmus Nielsen, Hans Polet, Chris J. Smith, Mustafa Zengin, Francois Bastardie, Ole R. Eigaard, Katell G. Hamon, Frank Jensen, Alessandro Lucchetti, Sonia Méhault, Emilio Notti, Nadia Papadopoulou, Andrea Petetta, Mattias Sköld, Benoit Vincent, and Adriaan D. Rijnsdorp. Technological innovations to reduce the impact of bottom gears on the seabed. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003949.

Saviolidis:2020:RBG

[SDI^{+}20] Nína M. Saviolidis, Brynhildur Davídsdóttir, Leena Ilmola, Anastasia Stepanova, Matilda Valman, and Elena Rovenskaya. Realising blue growth in the fishing industry in Iceland and Norway: Industry perceptions on drivers and barriers to blue growth investments and policy implications. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308760.

Sgardeli:2022:ASD

[SDL^{+}22] Vasiliki Sgardeli, Dimitrios Damalas, Angelos Liontakis, Christos D. Maravelias, Danai Mantopoulou-Palouka, and George Tserpes. The Aegean Sea demersal fishery under four climatic and socio-political futures. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200241X.

Setyawan:2022:HAM

[SEG^{+}22] Edy Setyawan, Mark Erdmann, Nikka Gunadharma, Tiene Gunawan, Abdi Hasan, Muhamad Izuan, Meidiarti Kasmidi, Yusdi Lamatenggo, Sarah Lewis, Nugraha Maulana, Ronald Mambrasar, Meity Mongdong, Alberth Nebore, Mochamad Iqbal Herwata Putra, Abraham Sianipar, Kristian Thebu, Syafri Tuharea, and Rochelle Constantine. A holistic approach to manta ray conservation in the Papuan Bird’s Head Seascape: Resounding success, ongoing challenges. Marine Policy, 137(??):??, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005649.

Sellheim:2020:CAI

[Sel20] Nikolas Sellheim. The CITES appendix II-Listing of mako sharks — Revisiting counter arguments. Marine Policy, 115 (??):??, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19307742.

Senghor:2025:PRS

[Sen25] Khadidiatou Senghor. Power relations in Sub-Saharan African small scale fisheries and conservation: a multidimensional review. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25001472.

Sepulveda:2025:ISA

[Sep25] Paulo César Sepúlveda Sepúlveda. The importance of sustainability approach for marine resources policy-making in Chile. Marine Policy, 180(??):??, October 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002234.

Seta:2023:CIC

[Set23] Makoto Seta. Compulsory insurance for cruise vessels as a preparation for the next pandemic: Law of the sea perspective. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001136.

Seung:2022:DGV

[Seu22] Chang K. Seung. Decomposing global value chain (GVC) income for world fisheries. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005613.

Song:2022:CCM

[SF22] Annie Young Song and Michael Fabinyi. China’s 21st century maritime silk road: Challenges and opportunities to coastal livelihoods in ASEAN countries. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21005340.

Song:2024:GCS

[SFB24] Annie Young Song, Michael Fabinyi, and Kate Barclay. China and seafood trade after coronavirus: the role of state actors. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005250.

Sanz-Fernandez:2022:PGC

[SFGEPC22] Víctor Sanz-Fernández, Juan Carlos Gutiérrez-Estrada, and Inmaculada Pulido-Calvo. Pagellus genus catches time series in the FAO Major Fishing Areas 27 and 34: Analysis of fishery behaviour. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005236.

Stein:2021:CEP

[SFN^{+}21] Florian Martin Stein, Jens Frankowski, Vincent Nijman, Christine Absil, Irene Kranendonk, and Willem Dekker. Chinese eel products in EU markets imply the effectiveness of trade regulations but expose fraudulent labelling. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002621.

Standal:2020:INB

[SG20a] Dag Standal and Eduardo Grimaldo. Institutional nuts and bolts for a mesopelagic fishery in Norway. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309649.

Stojanovic:2020:GFT

[SG20b] Tim Stojanovic and Kira Gee. Governance as a framework to theorise and evaluate marine planning. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19307560.

Suuronen:2020:MMF

[SG20c] Petri Suuronen and Eric Gilman. Monitoring and managing fisheries discards: New technologies and approaches. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18308716.

Standal:2021:LTP

[SG21] Dag Standal and Eduardo Grimaldo. Lost in translation? Practical- and scientific input to the mesopelagic fisheries discourse. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003961.

Shilland:2021:QSA

[SGA^{+}21] Robyn Shilland, Gabriel Grimsditch, Mohamed Ahmed, Salomão Bandeira, Hilary Kennedy, Maria Potouroglou, and Mark Huxham. A question of standards: Adapting carbon and other PES markets to work for community seagrass conservation. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001858.

Sakib:2021:APC

[SGB21] Nazmus Sakib, Elena Gissi, and Hermanni Backer. Addressing the pollution control potential of marine spatial planning for shipping activity. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002591.

Solo-Gabriele:2021:TIM

[SGFM^{+}21] Helena M. Solo-Gabriele, Tom Fiddaman, Cecilie Mauritzen, Cameron Ainsworth, David M. Abramson, Igal Berenshtein, Eric P. Chassignet, Shuyi S. Chen, Robyn N. Conmy, Christa D. Court, William K. Dewar, John W. Farrington, Michael G. Feldman, Alesia C. Ferguson, Elizabeth Fetherston-Resch, Deborah French-McCay, Christine Hale, Ruoying He, Vassiliki H. Kourafalou, Kenneth Lee, Yonggang Liu, Michelle Masi, Emily S. Maung-Douglass, Steven L. Morey, Steven A. Murawski, Claire B. Paris, Natalie Perlin, Erin L. Pulster, Antonietta Quigg, Denise J. Reed, James J. Ruzicka, Paul A. Sandifer, John G. Shepherd, Burton H. Singer, Michael R. Stukel, Tracey T. Sutton, Robert H. Weisberg, Denis Wiesenburg, Charles A. Wilson, Monica Wilson, Kateryna M. Wowk, Callan Yanoff, and David Yoskowitz. Towards integrated modeling of the long-term impacts of oil spills. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001652.

Sogn-Grundvaag:2022:QEF

[SGH22] Geir Sogn-Grundvåg and Øystein Hermansen. Qualityenhancing fishing in the coastal fishery for Atlantic cod in Norway. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200238X.

Shao:2021:ERG

[SGK21] Qinglong Shao, Junjie Guo, and Peng Kang. Environmental response to growth in the marine economy and urbanization: a heterogeneity analysis of 11 Chinese coastal regions using a panel vector autoregressive model. Marine Policy, 124(??):??, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310009.

Standal:2020:GII

[SGL20] Dag Standal, Eduardo Grimaldo, and Roger B. Larsen. Governance implications for the implementation of biodegradable gillnets in Norway. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20308848.

Stacey:2021:DSS

[SGL^{+}21] Natasha Stacey, Emily Gibson, Neil R. Loneragan, Carol Warren, Budy Wiryawan, Dedi S. Adhuri, Dirk J. Steenbergen, and Ria Fitriana. Developing sustainable smallscale fisheries livelihoods in Indonesia: Trends, enabling and constraining factors, and future opportunities. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21002657.

Sguotti:2023:TAC

[SGL^{+}23] Camilla Sguotti, Sanmitra Gokhale, Tin-Yu Lai, Esther Schuch, Christian Möllmann, and Andries Richter. Telecouplings in Atlantic cod — the role of global trade and climate change. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003512.

Serra-Goncalves:2023:SMD

[SGLF^{+}23] Catarina Serra-Gonçalves, Jennifer L. Lavers, Andrew M. Fischer, Heidi L. Tait, and Alexander L. Bond. Stopping marine debris at the source: Effectiveness of source reduction plans in Australia. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003093.

Shadrack:2021:SAS

[SGPF21] Ronick Spenly Shadrack, Sompert Gereva, Tim Pickering, and Marta Ferreira. Seasonality, abundance and spawning season of milkfish Chanos chanos (Forsskål, 1775) at Teouma Bay, Vanuatu. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001986.

Schuch:2021:IIE

[SGR21] Esther Schuch, Silke Gabbert, and Andries P. Richter. Institutional inertia in European fisheries — Insights from the Atlantic horse mackerel case. Marine Policy, 128(??): ??, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000750.

Strating:2022:CCM

[SGRW22] Rebecca Strating, Douglas Guilfoyle, Steven Ratuva, and Joanne Wallis. Commentary: China in the Maritime Pacific. Marine Policy, 141(??):??, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001397.

Song:2021:EDO

[SGWL21] Weiling Song, Jing Guo, Yue Wang, and Linlin Li. An evaluation of the development of the Ocean Economy in China using an Ocean Economic Development Index. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100302X.

Sogn-Grundvaag:2021:FQM

[SGZH^{+}21] Geir Sogn-Grundvåg, Dengjun Zhang, Edgar Henriksen, Sjurdur Joensen, Bjørn-Inge Bendiksen, and Øystein Hermansen. Fish quality and market performance: the case of the coastal fishery for Atlantic cod in Norway. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000592.

Standal:2022:CED

[SH22] Dag Standal and Bjørn Hersoug. Caught in the environmental discourse — coastal fisheries under pressure. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002305.

Standal:2023:IFC

[SH23] Dag Standal and Bjørn Hersoug. Illegal fishing: a challenge to fisheries management in Norway. Marine Policy, 155(??): ??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2300283X.

Standal:2024:WSM

[SH24a] Dag Standal and Bjørn Hersoug. Walking sideways? Management of the Norwegian snow crab fishery. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001787.

Stephenson:2024:BBE

[SH24b] Robert L. Stephenson and Alistair J. Hobday. Blueprint for Blue Economy implementation. Marine Policy, 163(??): ??, May 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24001271.

Sun:2025:CME

[SH25] Ruiheng Sun and Desheng Hu. China’s marine ecoenvironment law enforcement from the perspective of the responsibility system: Dilemmas and solutions. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004378.

Shao:2020:NEM

[Sha20] Qinglong Shao. Nonlinear effects of marine economic growth and technological innovation on marine pollution: Panel threshold analysis for China’s 11 coastal regions. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308371.

Shan:2022:EFO

[Sha22] Desai Shan. Enforcement of fishing Occupational Health and Safety (OHS) standards: Challenges in Atlantic Canada. Marine Policy, 145(??):??, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003293.

Sprogis:2023:EWW

[SHAC23] Kate R. Sprogis, Dirk Holman, Patricia Arranz, and Fredrik Christiansen. Effects of whale-watching activities on southern right whales in Encounter Bay, South Australia. Marine Policy, 150(??):??, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000520.

Sumby:2021:HFR

[SHFP21] Jonathan Sumby, Marcus Haward, Elizabeth A. Fulton, and Gretta T. Pecl. Hot fish: the response to climate change by regional fisheries bodies. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309301.

Shi:2020:SDB

[Shi20a] Yubing Shi. Settlement of disputes in a BBNJ agreement: Options and analysis. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20302621.

Shih:2020:TPT

[Shi20b] Yi-Che Shih. Taiwan’s progress towards becoming an ocean country. Marine Policy, 111(??):??, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302180.

Shields:2022:CWW

[Shi22] Monika W. Shields. Commercial whale-watching reduces vessel incidents in the vicinity of killer whales in Washington State. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003372.

Shi:2023:RBG

[Shi23a] Xinxiang Shi. The relationship between general principles of international law and Article 38(1) of the ICJ statute: a law of the sea perspective. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004742.

Shi:2023:IAC

[Shi23b] Xuefei Shi. International advocacy and China’s distant water fisheries policies. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001628.

Shi:2024:DBL

[Shi24a] Xinxiang Shi. Distinction between law enforcement activity and military activity in article 298(1)(b) of UNC-LOS: Commentary on recent international decisions. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002859.

Shih:2024:OGP

[Shi24b] Yi-Che Shih. Ocean governance in practice: a study of the application of marine science and technology research techniques to maritime law enforcement in Taiwan. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000794.

Smith:2021:EPI

[SHP^{+}21] David C. Smith, Malcolm Haddon, André E. Punt, Caleb Gardner, L. Richard Little, Stephen Mayfield, Michael F. O’Neill, Thor Saunders, John Stewart, Brent Wise, Elizabeth A. Fulton, and Simon Conron. Evaluating the potential for an increased and sustainable commercial fisheries production across multiple jurisdictions and diverse fisheries. Marine Policy, 124(??):??, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310034.

Schutter:2021:BEB

[SHPW21] Marleen S. Schutter, Christina C. Hicks, Jacob Phelps, and Claire Waterton. The blue economy as a boundary object for hegemony across scales. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002840.

Suherman:2025:TIC

[SHS^{+}25] Agus Suherman, Yayan Hernuryadin, Putuh Suadela, Ukon Ahmad Furkon, and Tono Amboro. Transformation of Indonesian capture fisheries governance: Review and prospects. Marine Policy, 174(??):??, April 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2500034X.

Song:2021:UAO

[SHZ^{+}21] Lan Song, Zhiqiu Huang, Huahan Zhang, Kaiwei Tian, Ning Yin, Yanmin Xu, Langxiong Gan, Jiangchao Wang, Jihong Chen, Yaqing Shu, and Chunmiao Zheng. The urgency to address the occupational health of Chinese seafarers for sustainable development. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001299.

Sun:2021:CGR

[SHZC21] Yue Sun, Yang Han, Liangfu Zhang, and Yen-Chiang Chang. The construction and governance of regional maritime security order in the South China Sea: Conference report. Marine Policy, 126(??):??, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000488.

Sakai:2022:SCE

[SII22] Yutaro Sakai, Hiroe Ishihara, and Mitsuhiro Ishino. Sharing in the commons: Evaluating the pooling system in a Danish seine fishery in Japan. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000641.

Singh:2021:TYD

[Sin21] Pradeep A. Singh. The two-year deadline to complete the International Seabed Authority’s Mining Code: Key outstanding matters that still need to be resolved. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21004152.

Sari:2021:MSS

[SIW^{+}21] Irna Sari, Muhammad Ichsan, Alan White, Syahril Abdul Raup, and Sugeng Hari Wisudo. Monitoring small-scale fisheries catches in Indonesia through a fishing logbook system: Challenges and strategies. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100381X.

Singer:2021:LBM

[SJ21] Rebecca Singer and Peter J. S. Jones. Lyme Bay marine protected area: a governance analysis. Marine Policy, 127 (??):??, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X18304317.

Samy-Kamal:2020:EFM

[SK20a] Mohamed Samy-Kamal. Estonian fisheries management system under the lens of the Marine Stewardship Council (MSC) standard. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19308826.

Samy-Kamal:2020:OFP

[SK20b] Mohamed Samy-Kamal. Outlook on the fisheries policy reform in Egypt and the draft of the new fisheries law. Marine Policy, 120(??):??, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302700.

Shiraishi:2024:EWU

[SK24] Hiromi Shiraishi and Kenzo Kaifu. Early warning of an upsurge in international trade in the American eel. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004712.

Szymanowska:2023:SIT

[SKDK23] Beata Barbara Szymanowska, Arkadiusz Kozłowski, Janusz Dabrowski, and Hanna Klimek. Seaport innovation trends: Global insights. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001124.

Shimose:2022:EOR

[SKE22] Tamaki Shimose, Minoru Kanaiwa, and Akihiko Ebisawa. Estimation of the optimal regulation size considering bid price increase with growth in size selectable coral reef fisheries. Marine Policy, 137(??):??, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000082.

Skerritt:2024:SCT

[Ske24] Daniel J. Skerritt. Seeking clarity on transparency in fisheries governance and management. Marine Policy, 165(??): ??, July 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24002197.

Simon:2020:UAV

[SKL^{+}20] Julien Simon, Dorothée Kopp, Pascal Larnaud, Jean-Philippe Vacherot, Fabien Morandeau, Gaël Lavialle, and Marie Morfin. Using automated video analysis to study fish escapement through escape panels in active fishing gears: Application to the effect of net colour. Marine Policy, 116 (??):??, June 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X18308078.

Svorken:2023:UOS

[SKL23] Marianne Svorken, Ingrid Kvalvik, and Nicholas Lord. Understanding the organisational structure of fisheries crime in well-regulated fisheries. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003937.

Smith:2021:DAS

[SKS21] William E. Smith, Gerard T. Kyle, and Steven Sutton. Displacement and associated substitution behavior among Texas inshore fishing guides due to perceived spotted seatrout declines. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002359.

Sorensen:2022:FGA

[SKW^{+}22] JA Sorensen, L. Kincl, R. Weil, J. Dzugan, and D. Christel. Fisheries governance and associated health implications: Current perspectives from US commercial fishermen. Marine Policy, 141(??):??, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200166X.

Shan:2020:VME

[SL20] Jingzhu Shan and Jingmei Li. Valuing marine ecosystem service damage caused by land reclamation: Insights from a deliberative choice experiment in Jiaozhou Bay. Marine Policy, 122(??):??, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308952.

Santos:2023:CNA

[SL23] Mateus Rennó Santos and Michael J. Lynch. The crossnational association between institutional imbalance, national culture, and illegal fishing. Marine Policy, 151(??): ??, May 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001288.

Schluter:2021:IBT

[SLB21] Maja Schlüter, Emilie Lindkvist, and Xavier Basurto. The interplay between top-down interventions and bottom-up self-organization shapes opportunities for transforming selfgovernance in small-scale fisheries. Marine Policy, 128(??): ??, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000968.

Solarin:2024:PFP

[SLGAG24] Sakiru Adebola Solarin, Carmen Lafuente, Luis A. Gil-Alana, and María Goenechea. Persistence of fisheries production: a disaggregated analysis in 31 OECD countries. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001647.

Strachan:2022:RIF

[SLH22] Laura L. Strachan, Richard J. Lilley, and Sebastian J. Hennige. A regional and international framework for evaluating seagrass management and conservation. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003530.

Suuronen:2023:RMC

[SLK^{+}23] Petri Suuronen, Sven-Gunnar Lunneryd, Sara Königson, Nelson F. Coelho, Åsa Waldo, Viktor Eriksson, Kristina Svels, Esa Lehtonen, Iwona Psuty, and Markus Vetemaa. Reassessing the management criteria of growing seal populations: the case of Baltic grey seal and coastal fishery. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002117.

Scherrer:2024:SRI

[SLL^{+}24] Kim J. N. Scherrer, Tom J. Langbehn, Gabriella Ljungström, Katja Enberg, Sara Hornborg, Gjert Dingsør, and Christian Jørgensen. Spatial restrictions inadvertently doubled the carbon footprint of Norway’s mackerel fishing fleet. Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000125.

Soto:2021:SWC

[SLMG^{+}21] Doris Soto, Jorge León-Muñoz, René Garreaud, Renato A. Quiñones, and Francisco Morey. Scientific warnings could help to reduce farmed salmon mortality due to harmful algal blooms. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100316X.

Srebaliene:2024:EWS

[SLO^{+}24] Greta Srėbalienė, Maiju Lehtiniemi, Henn Ojaveer, Stephan Gollasch, Okko Outinen, Manuel Sala Perez, Aleksas Narščius, and Sergej Olenin. Early warning system on harmful aquatic organisms at a regional sea scale: Components and mechanisms. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24003324.

Strongin:2022:PFT

[SLP^{+}22] Kyle Strongin, Alana Malinde S. N. Lancaster, Beth Polidoro, Alfonso Aguilar-Perera, Leah Gerber, Patricia González Díaz, Juliett González-Méndez, Larry McKinney, Héctor Espinosa Pérez, Daniel Pech, Dorka Cobián Rojas, Steven Saul, and Susana Perera Valderrama. A proposal framework for a Tri-National Agreement on Biological Conservation in the Gulf of Mexico Large Marine Ecosystem. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000884.

Suyo:2020:SNA

[SLS^{+}20] Jee Grace B. Suyo, Virginie Le Masson, Louise Shaxson, Maria Rovilla J. Luhan, and Anicia Q. Hurtado. A social network analysis of the Philippine seaweed farming industry: Unravelling the web. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19309571.

Suyo:2021:NRU

[SLS^{+}21] Jee Grace B. Suyo, Virginie Le Masson, Louise Shaxson, Maria Rovilla J. Luhan, and Anicia Q. Hurtado. Navigating risks and uncertainties: Risk perceptions and risk management strategies in the Philippine seaweed industry. Marine Policy, 126(??):??, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100018X.

Shrestha:2025:OWF

[SLS25] Kripa Shrestha, Pankaj Lal, and Meghann Smith. Offshore wind farms in New Jersey, USA: Stakeholder concerns and anticipated impacts on recreational fishing. Marine Policy, 181(??):??, November 2025. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25002702.

Sonvisen:2025:DSF

[SLSV25] Signe A. Sønvisen, Grethe Lilleng, Tore Syversen, and Dorthea Mathilde Kristin Vatn. Data sharing in the fisheries: Exploring the willingness to share data in the Norwegian fishing fleet. Marine Policy, 175(??):??, May 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000351.

Sidik:2023:BCN

[SLW^{+}23] Frida Sidik, Anissa Lawrence, Tonny Wagey, Franky Zamzani, and Catherine E. Lovelock. Blue carbon: a new paradigm of mangrove conservation and management in Indonesia. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004353.

Sheng:2021:IST

[SLY21] X. U. Sheng, Binbin Lu, and Qingde Yue. Impact of scitech finance on the innovation efficiency of China’s marine industry. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003195.

Sun:2025:DGM

[SLY25] Yu Sun, Feng Lian, and Zhongzhen Yang. Deciphering global marine product trade dynamics: Patterns, drivers, and policy insights. Marine Policy, 177(??):??, July 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2500096X.

Schinas:2023:CSC

[SM23] Orestis Schinas and Daniel Metzger. Cyber-seaworthiness: a critical review of the literature. Marine Policy, 151(??): ??, May 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001197.

Smith:2025:AMU

[SMB^{+}25] Sarah Lindley Smith, Jennifer McCann, Julia A. Bingham, Sereno Diederichsen, Fredrik Gröndahl, Josselin Guyot, Céline Rebours, Jean-Baptiste E. Thomas, Carlos V. C. Weiss, and John P. Walsh. Advancing multi-use in offshore wind energy planning: Perceived opportunities and barriers in southern New England, U.S. Marine Policy, 181(??): ??, November 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25002672.

Simmons:2021:SSA

[SMC^{+}21] Peter Simmons, Michael Mehmet, Belinda Curley, Nicola Ivory, Kane Callaghan, Kim Wolfenden, and Gang Xie. A scenario study of the acceptability to ocean users of more and less invasive management after shark-human interactions. Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100169X.

Stamatiadou:2025:BAA

[SMC^{+}25] Valentini Stamatiadou, Antonios D. Mazaris, Theodoros Chalazas, Adonis F. Velegrakis, and Stelios Katsanevakis. Are beaches accessible to all? Economic valuation and mapping of recreational ecosystem services for wheelchair users in Greece. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002209.

Sparks:2022:WLS

[SMCW22] Jessica L. Decker Sparks, Lee Matthews, Daniel Cárdenas, and Chris Williams. Worker-less social responsibility: How the proliferation of voluntary labour governance tools in seafood marginalise the workers they claim to protect. Marine Policy, 139(??):??, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000914.

Sahri:2020:CRM

[SMDM20] Achmad Sahri, Putu Liza Kusuma Mustika, Hadi Yoga Dewanto, and Albertinka J. Murk. A critical review of marine mammal governance and protection in Indonesia. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304397.

Schwoerer:2023:BHO

[SMF^{+}23] Tobias Schwoerer, Aaron Martin, Ginny Fay, Erik R. Schoen, and Michael Buntjer. Batten down the hatches: Opportunities to protect Alaska from biological invasions through watercraft trade and traffic. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200495X.

Santana-Morales:2021:HMT

[SMHPMO^{+}21] Omar Santana-Morales, E. Mauricio Hoyos-Padilla, Alfonso Medellín-Ortíz, Chugey Sepulveda, Rodrigo Beas-Luna, Marc Aquino-Baleytó, Edgar E. Becerril-García, Daniel Arellano-Millán, Luis Malpica-Cruz, Julio Lorda, and José Leonardo Castillo-Géniz. How much is too much? A carrying capacity study of white shark cage diving in Guadalupe Island, Mexico. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001998.

Smith:2020:BRC

[Smi20] Hance Smith. Book review: Climate Change and Ocean Governance: politics and policy for threatened seas, Paul G. Harris (Ed.). Cambridge University Press, Cambridge (2019), xiii + 450. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19302441.

Smith:2022:BRP

[Smi22] Hance Smith. Book review: The Poseidon Project: the struggle to govern the world’s oceans, David Bosco. Oxford University Press (2022), Notes, Selected Bibliography, Index, pp. x + 305. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002044.

Smith:2023:BRF

[Smi23a] Hance Smith. Book review: Fishing Europe’s Troubled Waters: fifty years of fisheries policy, David Symes. Earthscan from Routledge, New York (2023). Marine Policy, 155(??): ??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002774.

Smith:2023:BRO

[Smi23b] Hance Smith. Book review: Ocean Geopolitics: Marine Resources, Maritime Boundary Disputes and the Law of the Sea, Andreas Osthagen. Edward Elgar, Cheltenham (2022), Bibliography,Index. Pp. ix + 160. Marine Policy, 148(??): ??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2200522X.

Simard:2022:SWP

[SMKS22] Nittya S. M. Simard, Thane A. Militz, Jeff Kinch, and Paul C. Southgate. Shocks within a Pacific island fishery: a historic study of events impacting the mother-of-pearl fishery in Papua New Guinea. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002202.

Shalli:2025:CCP

[SML^{+}25] Mwanahija S. Shalli, Marcelina Mushi, Blandina R. Lugendo, Yohana W. Shaghude, January Wegoro, and Johan Hollander. Coastal communities’ perceptions on coastal erosion and the protective role of seagrass meadows in Dar es Salaam, Tanzania. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24004287.

Salvadeo:2021:FPI

[SMLRB^{+}21] Christian Salvadeo, Hem Nalini Morzaria-Luna, Héctor Reyes-Bonilla, Antonina Ivanova-Bonchera, David Petatán Ramírez, and Eduardo Juárez-León. Fisher’s perceptions inform adaptation measures to reduce vulnerability to climate change in a Mexican natural protected area. Marine Policy, 134(??):??, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004048.

Shiffman:2021:MSC

[SMW^{+}21] D. S. Shiffman, C. C. Macdonald, J. N. Wester, M. B. Walsh, A. Chevalier, D. Kachelriess, and K. J. Friedman. Marine species conservation at CITES: How does media coverage inform or misinform? Marine Policy, 134(??): ??, December 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004243.

Smythe:2024:CCA

[Smy24] Tiffany C. Smythe. Conflict or coordination? An analysis of the Southern New England offshore wind — fisheries policy network. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001040.

Seto:2022:TTG

[SMYH22] Katherine Seto, Nathan Miller, Mark Young, and Quentin Hanich. Toward transparent governance of transboundary fisheries: the case of Pacific tuna transshipment. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309133.

Suh:2023:CCD

[SN23] Neville N. Suh and Richard A. Nyiawung. Climate change dynamics and youth participation decisions in aquatic food systems: Case of the oyster sector in The Gambia, West Africa. Marine Policy, 156(??):??, October 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003378.

Steenbergen:2020:CRA

[SNK^{+}20] D. J. Steenbergen, P. T. Neihapi, D. Koran, A. Sami, V. Malverus, R. Ephraim, and N. Andrew. COVID-19 restrictions amidst cyclones and volcanoes: a rapid assessment of early impacts on livelihoods and food security in coastal communities in Vanuatu. Marine Policy, 121(??): ??, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20306898.

Sebe:2021:UCE

[SNP21] Maxime Sèbe, Abdelhak Nassiri, and Linwood Pendleton. Using choice experiment designs to evaluate mitigation solutions to reduce whale-ship collisions. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20310198.

Shan:2023:ISI

[SNS23] Desai Shan, Barb Neis, and Julie Sorensen. Introduction to special issue on governance and fishing health and safety in the North Atlantic. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002622.

Stoner-Osborne:2020:EMP

[SO20] Blake Stoner-Osborne. The effects of marine protected areas on populations of commercial reef fishes in Moorea, French Polynesia. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309625.

Sugimura:2022:NPC

[SOK^{+}22] Yoshihisa Sugimura, Tomonari Okada, Tomohiro Kuwae, Yugo Mito, Ryoji Naito, and Yasuyuki Nakagawa. New possibilities for climate change countermeasures in ports:

Organic carbon containment and creation of blue carbon ecosystems through beneficial utilization of dredged soil. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001191.

Solino:2023:SPL

[Sol23a] Mario Soliño. Social preferences for large marine protected areas in NW Spain. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003147.

Solomon:2023:CCT

[Sol23b] Preethi Solomon. Can cooperatives be a tool for poverty reduction? Social capital perspectives of fisher’s cooperatives in India. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004201.

Solarin:2024:CAP

[Sol24] Sakiru Adebola Solarin. Convergence of aquatic products consumption: a disaggregated analysis in European countries. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001908.

Soto-Onate:2020:ULA

[SOLN20] David Soto-Oñate and Ana C. Lemos-Nobre. Unreported landings across the EU Exclusive Economic Zones: the role of regional governance and social capital. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19303434.

Soto-Onate:2021:EUL

[SOLN21] David Soto-Oñate and Ana C. Lemos-Nobre. The European Union landing obligation: the compliance problems derived from its multilevel approach. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002773.

Sheridan:2021:RPL

[SON21] Kate Sheridan, M. Justin O’Riain, and Mark D. Needham. Recreationist perceptions of lethal and non-lethal management of sharks in two of South Africa’s marine areas. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100244X.

Brien:2022:MPD

[SONB^{+}22] Aleke Stöfen-O’Brien, Abolfazl Naji, Amy L. Brooks, Jenna R. Jambeck, and Farhan R. Khan. Marine plastic debris in the Arabian/Persian Gulf: Challenges, opportunities and recommendations from a transdisciplinary perspective. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005200.

Salgueiro-Otero:2020:BUS

[SOO20] Diego Salgueiro-Otero and Elena Ojea. A better understanding of social-ecological systems is needed for adapting fisheries to climate change. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20300440.

Seara:2022:LLN

[SOP^{+}22] Tarsila Seara, Adrien Owens, Richard Pollnac, Robert Pomeroy, and Christopher Dyer. Lessons learned from a natural resource disaster: the long-term impacts of the Long Island Sound lobster die-off on individuals and communities. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005546.

Snyder:2020:HCE

[SOS20] Hunter T. Snyder, Natuk Lund Olsen, and Andrew M. Song. How cybernetics explains institutional failure: a case of Greenland’s open-air fish markets. Marine Policy, 117 (??):??, July 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19301058.

Sosin:2021:BRG

[Sos21a] Claudia Sosin. Book review: Governing Marine Living Resources in the Polar Regions, Nengye Liu, Cassandra M. Brooks, Tianbao Qin (Eds.). Edward Elgar Publishing (2019). Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001639.

Storkersen:2021:FPD

[SOS^{+}21b] Kristine Vedal Størkersen, Tonje C. Osmundsen, Lars Helge Stien, Christian Medaas, Marianne Elisabeth Lien, Brit Tørud, Tore S. Kristiansen, and Kristine Gismervik. Fish protection during fish production. Organizational conditions for fish welfare. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100141X.

Sanada:2025:JWI

[SOS25] Yasuhiro Sanada, Ayako Okubo, and Isao Sakaguchi. Japan’s withdrawal from the IWC: an explanation. Marine Policy, 175(??):??, May 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25000107.

Saito:2022:TAN

[SOT^{+}22] Rui Saito, Yoshioki Oozeki, Kenji Takasaki, Tsutomu Saito, Shinji Uehara, and Masanori Miyahara. Transshipment activities in the North Pacific Fisheries Commission Convention Area. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003463.

Sreya:2021:EVS

[SPA^{+}21] P. S. Sreya, Chitra Parayil, N. Aswathy, Binoo P. Bonny, T. P. Aiswarya, and P. O. Nameer. Economic vulnerability of small-scale coastal households to extreme weather events in Southern India. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002190.

Smith:2021:MRM

[SPCCvT21] Christopher J. Smith, K. Nadia Papadopoulou, Eira Carballo-Cárdenas, and Jan P. M. van Tatenhove. Marine restoration in the Mediterranean: red coral and fan mussel discourses, uncertainty and reaching restoration targets. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000993.

Said:2020:SSF

[SPFA^{+}20] Alicia Said, Jose Pascual-Fernández, Vanessa Iglésias Amorim, Mathilde Højrup Autzen, Troels Jacob Hegland, Cristina Pita, Johanna Ferretti, and Jerneja Penca. Smallscale fisheries access to fishing opportunities in the European Union: Is the Common Fisheries Policy the right step to SDG14b? Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20302475.

Sanchez:2020:EPP

[SPG20] José L. Fernández Sánchez, José M. Fernández Polanco, and Ignacio Llorente García. Evidence of price premium for MSC-certified products at fishers’ level: the case of the artisanal fleet of common octopus from Asturias (Spain). Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304592.

Stebbings:2020:MEU

[SPH^{+}20] Emily Stebbings, Eleni Papathanasopoulou, Tara Hooper, Melanie C. Austen, and Xiaoyu Yan. The marine economy of the United Kingdom. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19307390.

Senghor:2023:CGR

[SPHO23] Khadidiatou Senghor, Stefan Partelow, Cesar Giraldo Herrera, and Isimemen Osemwegie. Conflicting governance realities: Aligning historical and cultural practices with formal marine protected area co-management in Senegal. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002397.

Stockin:2022:ERH

[SPJB22] Karen A. Stockin, Matthew D. M. Pawley, Rebecca M. Jarvis, and Rebecca M. Boys. Examining the role of human perceptions during cetacean stranding response in New Zealand. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200330X.

Speir:2023:MPL

[SPMN23] Cameron Speir, Amanda Phillips, Aaron Mamula, and Karma Norman. A measure of port-level resilience to shocks in commercial fisheries. Marine Policy, 151(??):??, May 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001021.

Sliskovic:2021:NIS

[SPN^{+}21] Merica Slišković, Marina Piria, Vedrana Nerlović, Katarina Pavičić Ivelja, Ana Gavrilović, and Gorana Jelić Mrčelić. Non-indigenous species likely introduced by shipping into the Adriatic Sea. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001275.

Sosnowski:2024:CSE

[SPNP24] Monique Sosnowski, Gohar Petrossian, Thanaphon Nunphong, and Eric Piza. Crimes at sea: Exploring the nexus of maritime crimes across global EEZs. Marine Policy, 166(??):??, August 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001593.

Stavroulakis:2020:ESA

[SPT^{+}20] Peter J. Stavroulakis, Stratos Papadimitriou, Vangelis Tsioumas, Ioannis G. Koliousis, Elena Riza, and Freideriki Tsirikou. Exploratory spatial analysis of maritime clusters. Marine Policy, 120(??):??, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309042.

Squalli:2020:EPE

[Squ20] Jay Squalli. Evaluating the potential economic, environmental, and social benefits of orange-spotted grouper aquaculture in the United Arab Emirates. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X20301809.

Squillace:2021:BRP

[Squ21] Mark Squillace. Best regulatory practices for deep seabed mining: Lessons learned from the U.S. Surface Mining Control and Reclamation Act. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2030974X.

Spaniol:2022:AIB

[SR22] Matthew J. Spaniol and Nicholas J. Rowland. Anticipated innovations for the blue economy: Crowdsourced predictions for the North Sea Region. Marine Policy, 137(??): ??, March 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004851.

Sugimoto:2022:HCN

[SRH^{+}22] Aoi Sugimoto, Raphael Roman, Juri Hori, Norie Tamura, Shingo Watari, and Mitsutaku Makino. How has the ‘customary nature’ of Japanese fisheries reacted to Covid-19? An interdisciplinary study examining the impacts of the pandemic in 2020. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000525.

Stoll:2023:RSS

[SRH23] Joshua S. Stoll, Sarah C. Risley, and Paloma R. Henriques. A review of small-scale marine fisheries in the United States: Definitions, scale, drivers of change, and policy gaps. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004560.

Suarez-Rojas:2025:SRF

[SRL25] Chaitanya Suárez-Rojas and Carmelo J. León. A scientometrics review of fifty years of whale-watching tourism research. Marine Policy, 176(??):??, June 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000739.

Suarez-Rojas:2023:WDY

[SRLLG23] Chaitanya Suárez-Rojas, Carmelo J. León, and Yen E. Lam-González. What drives you to the sea? Animal rights, environmental protection and sensation seeking. Marine Policy, 147(??):??, January 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003955.

Sarker:2021:MZI

[SRR^{+}21] Subrata Sarker, Md Jalilur Rahman, Muhammad Mizanur Rahman, Morgina Akter, Md Shajjadur Rahman, and Md Abdul Wahab. MPA zoning integrating socioecological data in the South East coast of Bangladesh. Marine Policy, 133(??):??, November 2021. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100347X.

Strand:2023:CEC

[SRS23] M. Strand, N. Rivers, and B. Snow. The complexity of evaluating, categorising and quantifying marine cultural heritage. Marine Policy, 148(??):??, February 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004961.

Stancheva:2025:EIL

[SRS^{+}25] Margarita Stancheva, Emiliano Ramieri, Hristo Stanchev, Martina Bocci, Marina Markovic, and Timothy Roberts. Exploring integration of land-sea interactions in maritime spatial planning in the Bulgarian context, Black Sea. Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004160.

Suris-Regueiro:2021:AFE

[SRSGMGG21a] Juan C. Surís-Regueiro, José L. Santiago, Xosé M. González-Martínez, and M. Dolores Garza-Gil. An applied framework to estimate the direct economic impact of Marine Spatial Planning. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000531.

Suris-Regueiro:2021:EEI

[SRSGMGG21b] Juan C. Surís-Regueiro, José L. Santiago, Xosé M. González-Martínez, and M. Dolores Garza-Gil. Estimating economic impacts linked to Marine Spatial Planning with input-output techniques. Application to three case studies. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001524.

Sabela-Rikhotso:2021:CAL

[SRvNN21] Phindile Tiyiselani Zanele Sabela-Rikhotso, Dewald van Niekerk, and Livhuwani David Nemakonde. A critical analysis of the legal frameworks governing oil spill management in South Africa. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000439.

Strating:2024:HRS

[SRY24] Rebecca Strating, Sunil Rao, and Sallie Yea. Human rights at sea: the limits of inter-state cooperation in addressing forced labour on fishing vessels. Marine Policy, 159(??): ??, January 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004670.

Skerritt:2021:BGD

[SS21a] Daniel J. Skerritt and U. Rashid Sumaila. Broadening the global debate on harmful fisheries subsidies through the use of subsidy intensity metrics. Marine Policy, 128(??): ??, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001184.

Sowman:2021:JTN

[SS21b] M. Sowman and Jackie Sunde. A just transition? Navigating the process of policy implementation in small-scale fisheries in South Africa. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002943.

Silver:2022:FIC

[SS22] Jennifer J. Silver and Joshua S. Stoll. A framework for investigating commercial license and quota holdings in an era of fisheries consolidation, concentration and financialization. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002263.

Silva:2023:EBL

[SS23] Luana C. C. Silva and Bráulio A. Santos. Evolution of Brazilian legislation regarding the marine fauna: Advances and shortcomings between 1960 and 2020.

Marine Policy, 153(??):??, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23001653.

Scotts:2023:SCR

[SSA^{+}23] Gemma L. Scotts, Michael J. Scales, Gonzalo Araujo, Hollie Booth, and Sarah A. Marley. Socio-cultural relationship between recreational sea anglers and blue sharks (Prionace glauca) in the United Kingdom. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003640.

Sullivan-Stack:2024:AEM

[SSAAB^{+}24] Jenna Sullivan-Stack, Gabby N. Ahmadia, Dominic A. Andradi-Brown, Alexandra Barron, Cassandra M. Brooks, Joachim Claudet, Barbara Horta e Costa, Estradivari, Laurel C. Field, Sylvaine Giakoumi, Emanuel Gonçalves, Natalie Groulx, Jean Harris, Sabine Jessen, Steven Mana’oakamai Johnson, Jessica M. C. MacCarthy, Guilherme Maricato, Lance Morgan, Katharine Bear Nalven, Emily S. Nocito, Elizabeth P. Pike, Enric Sala, Rodrigo Tardin, Angelo Villagomez, Kendyl Wright, and Kirsten Grorud-Colvert. Assessments of expected MPA outcomes can inform and improve biodiversity conservation: Case studies using The MPA Guide. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003622.

Snyder:2024:GGI

[SSAP24] Hunter T. Snyder, Simon Stone, Mary Albert, and Chris Polashenski. Are Greenland’s inshore halibut fishers ready for individual transferable quotas? Marine Policy, 159(??): ??, January 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004657.

Silva:2022:EFL

[SSB^{+}22] Monalisa R. O. Silva, André B. Silva, Jaciana C. Barbosa, Cássia Amaral, and Priscila F. M. Lopes. Empowering fisherwomen leaders helped reduce the effects of the COVID-19 pandemic on fishing communities: Insights from Brazil. Marine Policy, 135(??):??, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100453X.

Short:2023:MZM

[SSB^{+}23] Charlie Short, Joanna L. Smith, John Bones, Steve Diggon, Aaron Heidt, Chris Mcdougall, and Kylee A. Pawluk. Marine zoning for the Marine Plan Partnership (MaPP) in British Columbia, Canada. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000519.

Shannon:2020:GBB

[SSDQ20] Caitriona Shannon, Paul D. Stebbing, Alison M. Dunn, and Claire H. Quinn. Getting on board with biosecurity: Evaluating the effectiveness of marine invasive alien species biosecurity policy for England and Wales. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309210.

Szymkowiak:2023:ICA

[SSF23] Marysia Szymkowiak, Andrew Steinkruger, and Kelsi Furman. Integrating climate adaptation into Gulf of Alaska fishing community planning. Marine Policy, 156(??):??, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003354.

Sherman:2023:GPU

[SSH^{+}23] C. Samantha Sherman, Colin A. Simpfendorfer, Alifa B. Haque, Eric D. Digel, Patrick Zubick, Jonathan Eged, Jay H. Matsushiba, Glenn Sant, and Nicholas K. Dulvy. Guitarfishes are plucked: Undermanaged in global fisheries despite declining populations and high volume of unreported international trade. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002865.

Shimotsuura:2023:FHS

[SSK23] Taiga Shimotsuura, Tomomi Shoda, and Shigemi Kagawa. Firm heterogeneity in sources of changes in CO_{2} emissions from international container shipping. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003925.

Suryawan:2025:CIP

[SSL25] I. Wayan Koko Suryawan, Imelda Masni Juniaty Sianipar, and Chun-Hung Lee. Community importance-performance preferences and policy adaptiveness in marine debris management: a case study from the Komodo Subdistrict, Indonesia. Marine Policy, 174(??):??, April 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000077.

Solaun:2021:SOB

[SSMG21] Oihana Solaun, Yolanda Sagarminaga, Iratxe Menchaca, and Ibon Galparsoro. Spatial overlap between human activities and seabed habitats in European Seas: Insights into EMODnet’s data for management purposes. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21001627.

Serra-Sogas:2021:UAS

[SSOP^{+}21] Norma Serra-Sogas, Patrick D. O’Hara, Kim Pearce, Leh Smallshaw, and Rosaline Canessa. Using aerial surveys to fill gaps in AIS vessel traffic data to inform threat assessments, vessel management and planning. Marine Policy, 133(??):??, November 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003766.

Sowman:2021:UGF

[SSP^{+}21a] M. Sowman, J. Sunde, T. Pereira, B. Snow, P. Mbatha, and A. James. Unmasking governance failures: the impact of COVID-19 on small-scale fishing communities in South Africa. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003249.

Sunny:2021:AIC

[SSP^{+}21b] Atiqur Rahman Sunny, Sharif Ahmed Sazzad, Shamsul Haque Prodhan, Md. Ashrafuzzaman, Gopal Chandra Datta, Ashoke Kumar Sarker, Mizanur Rahman, and Mahmudul Hasan Mithun. Assessing impacts of COVID-19 on aquatic food system and small-scale fisheries in Bangladesh. Marine Policy, 126(??):??, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000324.

Sornia:2025:BRR

[SSSL25] Muti’ah Dewi Sornia, Sebastianus Kia Suban, Amelia Fransiska Siregar, and Meitin Lolon. Book review: Research Handbook on International Marine Environmental Law (Second Edition), edited by Rosemary Rayfuse, Aline Jaeckel, and Natalie Klein, Part of the book series: Research Handbooks in Environmental Law, Edward Elgar (2023), 492pp. ISBN: 978-1-78990-907-4 (hardback). Marine Policy, 180(??):??, October 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001861.

Skerritt:2023:MUG

[SSV^{+}23] Daniel J. Skerritt, Anna Schuhbauer, Sebastian Villasante, Andrés M. Cisneros-Montemayor, Nathan J. Bennett, Tabitha G. Mallory, Vicky W. L. Lam, Robert I. Arthur, William W. L. Cheung, Louise S. L. Teh, Katina Roumbedakis, Maria L. D. Palomares, and U. Rashid Sumaila. Mapping the unjust global distribution of harmful fisheries subsidies. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001380.

Sparks:2025:ISI

[SSW^{+}25] Jessica L. Decker Sparks, Natalie Somers, Chris Williams, Michael O’Brien, and Bethany Jackson. The illusion of social improvements: a case study of the role of Fishery Improvement Projects (FIPs) in fairwashing the UK Nephrops fishery. Marine Policy, 175(??):??, May 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000442.

Stancheva:2022:SMU

[SSZ^{+}22] Margarita Stancheva, Hristo Stanchev, Jacek Zaucha, Emiliano Ramieri, and Timothy Roberts. Supporting multiuse of the sea with maritime spatial planning. The case of a multi-use opportunity development — Bulgaria, Black Sea. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005388.

Seminara:2024:OCS

[SSZ^{+}24] Cecilia Inés Seminara, Gabriel Barros Gonçalves Souza, Cleverson Zapelini, Márcio Luiz Vargas Barbosa-Filho, and Christiana Cabicieri Profice. Opportunities and challenges for small-scale fisheries co-management: a case study of the unofficial landing records system in a Brazilian fishing community. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005249.

Soeparna:2024:TPB

[ST24] Intan Innayatun Soeparna and Ankiq Taofiqurohman. Transversal policy between the protection of marine fishery resources and fisheries subsidies to address overfishing in Indonesia. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001106.

Stadtlander:2023:BRP

[Sta23a] Christian T. K.-H. Stadtländer. Book review: S. Partelow, M. Hadjimichael, A.-K. Hornidge, Ocean Governance: knowledge systems, policy foundations and thematic, 2023 Springer Nature Publishing Cham, Switzerland xx + 436 pp., ISBN: 978-3-031-20739-6 (Hardcover), ISBN: 978-3-031-20742-6 (Softcover), ISBN: 978-3-031-20740-2 (eBook). Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001720.

Starr:2023:TAC

[Sta23b] Antony Starr. Total allowable catches and fishing opportunity decisions of the Council of the EU — a voting power index approach. Marine Policy, 156(??):??, October 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003317.

Standen:2025:DMD

[Sta25] Sophie Standen. Developing the market, developing the fishery? Post-harvest associations in the making of the fish market in Ghana. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24005360.

Stepien:2023:CSO

[Ste23] Barbara Stepień. Can a ship be its own captain? Safe manning of autonomous and uncrewed vessels. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004985.

Stenek:2025:BRU

[Ste25] Ariel Stenek. Book review: Understanding Maritime Security, Christian Bueger, Timothy Edmunds. Oxford University Press (2024), 240, ISBN: 978-0-19-776715-3 (paperback). Marine Policy, 175(??):??, May 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000405.

Syddall:2021:TAP

[STF21] Victoria Syddall, Simon Thrush, and Karen Fisher. Transdisciplinary analysis of Pacific tuna fisheries: a research framework for understanding and governing oceans as social–ecological systems. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003948.

Stien:2020:GWN

[STG^{+}20] Lars Helge Stien, Brit Tørud, Kristine Gismervik, Marianne Elisabeth Lien, Christian Medaas, Tonje Osmundsen, Tore S. Kristiansen, and Kristine Vedal Størkersen. Governing the welfare of Norwegian farmed salmon: Three conflict cases. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309108.

Schug:2020:UOD

[STIS20] Donald M. Schug, Peter H. Taylor, Suzanne Iudicello, and Jill H. Swasey. Using online data visualization and analysis to facilitate public involvement in management of catch share programs. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309180.

Su:2022:UST

[STKC22] Shu Su, Yi Tang, Jacob P. Kritzer, and Yong Chen. Using systems thinking to diagnose science-based fisheries management in China. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000215.

Storkersen:2021:SMR

[Stø21] Kristine Vedal Størkersen. Safety management in remotely controlled vessel operations. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309994.

Scuderi:2023:SCG

[STS^{+}23] Alessia Scuderi, Rocco Tiberti, Javier García Sanabria, Lucia Merino, Cristina Otero-Sabio, André Pedrosa, and Filomena Cardoso Martins. Sustainability as a common goal: Regulatory compliance, stakeholder perspectives, and management implications of whale and dolphin watching in the Strait of Gibraltar. Marine Policy, 154(??):??, August 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001902.

Socrate:2022:AUA

[SV22] Juliana Socrate and Eleonora Verón. Analysis of uses and activities in the Argentine Sea. Bases for a marine spatial planning in the North Argentina Basin. Marine Policy, 139 (??):??, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000616.

Schluter:2020:CRS

[SVF20] Achim Schlüter, Colin Vance, and Sebastian Ferse. Coral reefs and the slow emergence of institutional structures for a glocal land- and sea-based collective dilemma. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18303579.

Socrate:2024:CPA

[SVG24] Juliana Socrate, Eleonora Verón, and Germán García. Contributions to the planning of Argentine maritime spaces: the Northern Patagonian socioecological system as a case study. Marine Policy, 168(??):??, October 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003208.

Schluter:2020:LSI

[SVHV20] Achim Schlüter, Kristof Van Assche, Anna-Katharina Hornidge, and Nataşa Văidianu. Land-sea interactions and coastal development: an evolutionary governance perspective. Marine Policy, 112(??):??, February 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309273.

Steins:2021:COW

[SVKP21] Nathalie A. Steins, Jeroen A. Veraart, Judith E. M. Klostermann, and Marnix Poelman. Combining offshore wind farms, nature conservation and seafood: Lessons from a Dutch community of practice. Marine Policy, 126(??):??, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310228.

Schadeberg:2024:SER

[SvLGK24] Amanda Schadeberg, Judith van Leeuwen, Rolf A. Groeneveld, and Marloes Kraan. Science is not enough: the role of legitimacy in the governance of marine activities. Marine Policy, 169(??):??, November 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400335X.

Sordo:2023:RHW

[SVP^{+}23] Laura Sordo, Paulo Vasconcelos, David Piló, André N. Carvalho, Fábio Pereira, and Miguel B. Gaspar. Recreational harvesting of the wedge clam (Donax trunculus) in southern Portugal: Characterization of the activity based on harvesters’ perception and local ecological knowledge. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300221X.

Seixas:2024:HCP

[SVV24] Sónia Seixas, Tiago Verdelhos, and Helena Veríssimo. How COVID-19 pandemic affected fisheries (catch volume and price): a case study in Europe. Marine Policy, 159(??): ??, January 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004293.

Strating:2022:MST

[SW22] Rebecca Strating and Joanne Wallis. Maritime sovereignty and territorialisation: Comparing the Pacific Islands and South China Sea. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001579.

Shelton:2024:DRG

[SWF^{+}24] Clare Shelton, Carole S. White, Johanna Forster, Susan Conlon, Georg H. Engelhard, and John K. Pinnegar. Disaster risk in Caribbean fisheries: How vulnerability is shaped and how it can be reduced in Dominica and Antigua and Barbuda. Marine Policy, 160(??):??, February 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004840.

Song:2022:MCO

[SWGZ22] Runxi Song, Lei Wu, Matt Geraci, and Hui Zhong. Maritime Cooperation and Ocean Governance 2021: Symposium report. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003499.

Sjafrie:2023:NAR

[SWH^{+}23] Nurul Dhewani Mirah Sjafrie, Pramaditya Wicaksono, Udhi E. Hernawan, Triyono, Doddy M. Yuwono, Muhammad Hafizt, Novi Susetyo Adi, Rohani Ambo-Rappe, Bayu Prayudha, Muhammad Banda Selamat, Sofia Yunia Sani, Setiawan Djody Harahap, Huwaida Nur Salsabila, Jennifer Wijaya, and Chris Roelfsema. Network analysis reveals overlapping roles of stakeholders related to seagrassdata provisioning in Indonesia. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003706.

Shiiba:2022:HBF

[SWHT22] Nagisa Shiiba, Hsing Hao Wu, Michael C. Huang, and Hajime Tanaka. How blue financing can sustain ocean conservation and development: a proposed conceptual framework for blue financing mechanism. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2100186X.

Sari:2022:TEA

[SWI^{+}22] Irna Sari, Alan White, Muhammad Ichsan, Jason Cope, Josh Nowlis, Christovel Rotinsulu, Stephanus Mandagi, Enos Menai, Zulkifli Henan, Rishi Sharma, Syafri Tuharea, Roger Tabalessy, and Melisa Masengi. Translating the ecosystem approach to fisheries management into practice: Case of anchovy management, Raja Ampat, West Papua, Indonesia. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002093.

Su:2021:TPM

[SWW^{+}21] Ming Ming Su, Geoffrey Wall, Bangle Wu, Huangxuan Xu, Xiaohan Fu, and Yuncheng Deng. Tourism place making through the bioluminescent “Blue Tears” of Pingtan Islands, China. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003559.

Su:2021:ATC

[SWXM21] Meng Su, Lingling Wang, Jiahao Xiang, and Yuxia Ma. Adjustment trend of China’s marine fishery policy since 2011. Marine Policy, 124(??):??, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309696.

Shi:2025:MIO

[SX25] Zhaoyuan Shi and Desheng Xue. Measuring the international ocean economy trade: Method and application. Marine Policy, 175(??):??, May 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2500051X.

Sagheer:2025:DIG

[SYA25] Muneeb Sagheer, Zhengyong Yang, and Mohd Alsaleh. Determinants influencing green cost efficiency in the aquaculture sector: New insights from Asian countries. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001988.

Sun:2024:NTC

[SYY24] Caizhi Sun, Zhennan Yang, and Yudi Yang. Navigating the tides of change: the local dynamics of China’s marine economic policy. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002057.

Sikveland:2020:DCS

[SZ20] Marius Sikveland and Dengjun Zhang. Determinants of capital structure in the Norwegian salmon aquaculture industry. Marine Policy, 119(??):??, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X1930764X.

Sun:2024:WGO

[SZ24] Dong Sun and Linjie Zhang. Where are the global ocean priority conservation areas under the most severe threat from microplastics? Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003245.

Spijkerboer:2020:PMS

[SZBA20] R. C. Spijkerboer, C. Zuidema, T. Busscher, and J. Arts. The performance of marine spatial planning in coordinating offshore wind energy with other sea-uses: the case of the Dutch North Sea. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19300703.

Szymkowiak:2020:GFA

[Szy20] Marysia Szymkowiak. Genderizing fisheries: Assessing over thirty years of women’s participation in Alaska fisheries. Marine Policy, 115(??):??, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307936.

Tang:2022:SMH

[TAK^{+}22] Lijun Tang, Sanley Abila, Momoko Kitada, Serafin Malecosio, and Karima Krista Montes. Seafarers’ mental health during the COVID-19 pandemic: an examination of current supportive measures and their perceived effectiveness. Marine Policy, 145(??):??, November 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003232.

Tang:2022:IFC

[Tan22] Lijun Tang. The impact of 2008 financial crisis and COVID-19 pandemic on the demand and supply of seafarer officers: Evidence from China. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22003104.

Tan:2023:BRE

[Tan23a] Ying Tan. Book review: Emerging Technology and the Law of the Sea, James Kraska, Young Kil Park (Eds.). Cambridge University Press (2022), 369, ISBN 978-1-316-51742-0. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001495.

Tan:2023:RHI

[Tan23b] Ying Tan. Book review: Research Handbook on International Marine Environmental Law, 2nd Edition, Rosemary Rayfuse, Aline Jaeckel, Natalie Klein (Eds.). Edward Elgar Publishing, Inc. (2023), 492, ISBN 978-1-78990-907-4. Marine Policy, 157(??):??, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300386X.

Tang:2024:FFF

[Tan24] Jianye Tang. Form follows function: an initial evaluation of the BBNJ Agreement’s achievements regarding the “not undermining” proviso. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004852.

Tanaka:2025:JLE

[Tan25] Yoshihiro Tanaka. Japan’s law enforcement issues to combat illegal fisheries. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2500123X.

Turrell:2023:CRI

[TAP^{+}23] W. R. Turrell, W. E. N. Austin, S. P. Philbrick, C. Tilbrook, and H. Kennedy. Clarifying the role of inorganic carbon in blue carbon policy and practice. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23004062.

Tarkun:2025:DRB

[Tar25] Savaş Tarkun. Dynamic relationships between inflation, industrial production, and freight rates in the EU: a wavelet coherence and causality analysis. Marine Policy, 172(??): ??, February 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24005347.

Toro:2025:HRS

[TBAR^{+}25] Frederick Toro, Susannah J. Buchan, Mario Alvarado-Rybak, Luis Bedriñana-Romano, Ana M. Garcia-Cegarra, Rodrigo Hucke-Gaete, Carlos Olavarria, Maritza Sepúlveda, Barbara Galletti-Vernazzani, Walter Sielfeld, Anelio Aguayo, Juan Capella, Maria José Perez-Alvarez, Francisco Viddi, Benjamin Caceres, Jorge Gibbons, Jorge Acevedo, Hector Guzman, Jorge Guerra, Ricardo Saez, Mauricio Seguel, and M. Michael Fishbach. High rate of ship strike to large whales off Chile: Historical data and proposed actions to reduce risk. Marine Policy, 174(??):??, April 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24005773.

Thompson:2025:MPA

[TBCB25] Katie R. Thompson, Fernando Bretos, Puri Canals, and Charles Besancon. Marine protected area networks as tools for Ocean Science Diplomacy: Global lessons from the Gulf of Mexico. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001241.

Teschke:2021:PMP

[TBH^{+}21] Katharina Teschke, Patricia Brtnik, Stefan Hain, Heike Herata, Alexander Liebschner, Hendrik Pehlke, and Thomas Brey. Planning marine protected areas under the CCAMLR regime — The case of the Weddell Sea (Antarctica). Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20310216.

Turschwell:2020:CBR

[TBPC20] Mischa P. Turschwell, Christopher J. Brown, Ryan M. Pearson, and Rod M. Connolly. China’s Belt and Road Initiative: Conservation opportunities for threatened marine species and habitats. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19307675.

Tran:2023:ROV

[TBV^{+}23] Trung Tien Tran, Thomas Browne, Brian Veitch, Mashrura Musharraf, and Dennis Peters. Route optimization for vessels in ice: Investigating operational implications of the carbon intensity indicator regulation. Marine Policy, 158(??): ??, December 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003913.

Tong:2023:NPG

[TC23] Helong Tong and Shishi Chen. A new pattern of green, intelligent and resilient international shipping: Conference report. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001148.

Tran:2022:PFS

[TCC^{+}22] Nhuong Tran, Long Chu, Chin Yee Chan, Jeffrey Peart, Ahmed M. Nasr-Allah, and Harrison Charo-Karisa. Prospects of fish supply–demand and its implications for food and nutrition security in Egypt. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003803.

Turner:2020:MMP

[TCD^{+}20] Phillip J. Turner, Sophie Cannon, Sarah DeLand, James P. Delgado, David Eltis, Patrick N. Halpin, Michael I. Kanu, Charlotte S. Sussman, Ole Varmer, and Cindy L. Van Dover. Memorializing the Middle Passage on the Atlantic seabed in Areas Beyond National Jurisdiction. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309003.

Tett:2025:LNR

[TCF^{+}25] Paul Tett, George Charalambides, Sofia C. Franco, Adam D. Hughes, Eirik Mikkelsen, Kåre Nolde Nielsen, Eric Arthur Bastos Routledge, Pernille Nielsen, and Philip James. Leaving the niche: Recommendations for mainstreaming Low Trophic Aquaculture in countries around the Atlantic basin. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004755.

Tran:2020:PPR

[TCS^{+}20] Nhuong Tran, Quyen Le Cao, Kelvin Mashisia Shikuku, Thanh Phuong Phan, and Lauren K. Banks. Profitability and perceived resilience benefits of integrated shrimptilapia-seaweed aquaculture in North Central Coast, Vietnam. Marine Policy, 120(??):??, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300725.

Ting:2021:IPP

[TCT21] Kuo-Huan Ting, Chung-Ting Cheng, and Hou-Yi Ting. Introducing the problem/project based learning as a learning strategy in University Social Responsibility Program — a study of local revitalization of Coastal Area, Yong-An District of Kaohsiung City. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001573.

Tsai:2024:PPM

[TCT24] Chi-Heng Tsai, Huey-Shian Elly Chung, and Huan-Sheng Tseng. Preventing poaching in marine protected areas: a crime script analysis to inform interventions. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24001246.

Tang:2021:ICC

[TCZ21] Yi Tang, Wenjin Chen, and Yanxuedan Zhang. International cooperation and coordination in the Global Legislation of High Seas ABMTs including MPAs: Taking OSPAR practice as reference. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2100378X.

Thiebot:2021:PMB

[TD21] Jean-Baptiste Thiebot and Magali Dreyfus. Protecting marine biodiversity beyond national jurisdiction: a penguins’ perspective. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002517.

Tortolini:2021:FCS

[TDC21] Virginia Mailen Tortolini, Mariana Degrati, and Mariano Alberto Coscarella. Framing and communicating southern right whale — kelp gull biological interaction in Peninsula Valdés, Argentina: the effects of attribute frames on human’s perceptions and decision-making policies. Marine Policy, 124(??):??, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309611.

Tsikliras:2020:AUT

[TDP20] Athanassios C. Tsikliras, Donna Dimarchopoulou, and Androniki Pardalou. Artificial upward trends in Greek marine landings: a case of presentist bias in European fisheries. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19307377.

Teampau:2020:TPC

[Tea20] Petruţa Teampău. Trouble in paradise: Competing discourses and complex governance in the Romanian Danube delta. Marine Policy, 112(??):??, February 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303737.

Tranter:2022:IFT

[TEA^{+}22] Sylvie N. Tranter, Estradivari, Gabby N. Ahmadia, Dominic A. Andradi-Brown, Dominic Muenzel, Firdaus Agung, Amkieltiela, Amanda K. Ford, Abdullah Habibi, Christian N. Handayani, Mohamad Iqbal, Nils C. Krueck, Muhammad E. Lazuardi, Umi Muawanah, Renoldy L. Papilaya, Tries B. Razak, Agus Sapari, Fikri F. Sjahruddin, Laura Veverka, Safran Yusri, and Maria Beger. The inclusion of fisheries and tourism in marine protected areas to support conservation in Indonesia. Marine Policy, 146(??): ??, December 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003487.

Teijo:2023:CSI

[Tei23] Carlos Teijo. Criminal sanctions and IUU fishing: the case of Spain. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003888.

Telesetsky:2023:CMC

[Tel23a] Anastasia Telesetsky. Cellular mariculture: Challenges of delivering sustainable protein security. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X2200447X.

Telesetsky:2023:CRU

[Tel23b] Anastasia Telesetsky. Cross regime UNCLOS and UNFCCC cooperation to address loss and damage from climateshifted transboundary fisheries. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004730.

Teo:2023:OMO

[Teo23] Jotham Teo. Ocean materialities and the ontologies of ocean conservation in Areas Beyond National Jurisdiction: the Costa Rica Thermal Dome. Marine Policy, 153(??):??, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001732.

Testa:2020:NPD

[Tes20] David Testa. A note on the potential designation of the Mediterranean Sea as a sulphur emission control area. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20305765.

Trueman:2022:TCN

[TFF22] Justin D. Trueman, Ramón Filgueira, and Lucia Fanning. Transparency and communication in Norwegian and Nova Scotian Atlantic salmon aquaculture industries. Marine Policy, 138(??):??, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000057.

Teneva:2023:SIN

[TFH^{+}23] Lida Teneva, Christopher M. Free, Andrew Hume, Vera N. Agostini, Carissa J. Klein, Reg A. Watson, and Steven D. Gaines. Small island nations can achieve food security benefits through climate-adaptive blue food governance by 2050. Marine Policy, 151(??):??, May 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001045.

Trathan:2025:FAK

[TFJR25] Philip Trathan, Ari Friedlaender, Chris Johnson, and Ryan Reisinger. The fishery for Antarctic krill — conflicts between industrial production, protection of biodiversity, and legal governance. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002027.

Turschwell:2023:SAS

[TFMT^{+}23] M. P. Turschwell, E. A. Fulton, J. Melbourne-Thomas, M. Lacharité, K. R. Hayes, G. Wood, K. Evans, D. Hatton MacDonald, J. Dambacher, R. H. Bustamante, R. Abbassi, P. Fidelman, and C. J. Brown. A synthesis of approaches to support integrated assessments of hazards for the emerging blue economy. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002233.

Tsai:2025:NGW

[TFP25] Chi-Heng Tsai, MoSiao Fang, and Gohar Petrossian. Netting the global workforce: Unraveling the networks of laborers on distant water fishing fleets. Marine Policy, 181(??): ??, November 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25002386.

Tian:2025:PIB

[TG25a] Haoyu Tian and Jianping Guo. The potential interactions between the BBNJ Agreement and RFMOs in the establishment of ABMTs: Implications for RFMOs. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004779.

Tsuchiya:2025:UCE

[TG25b] Motohiro Tsuchiya and Kristi Govella. Undersea cables and the extension of empire: the rise of Britain, Japan, and the United States and the competition to connect Hawai’i. Marine Policy, 181(??):??, November 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2500260X.

Tidd:2025:EPS

[TGF^{+}25] Alex Tidd, Patrice Guillotreau, Dan Fu, Iago Mosqueira, Laurent Dagorn, and Manuela Capello. Equitable pathways for sustainable tuna fisheries management in the Indian Ocean. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005633.

Touron-Gardic:2022:SPV

[TGHF^{+}22] Gregoire Touron-Gardic, Øystein Hermansen, Pierre Failler, Abdou Daïm Dia, Mohamed Ould Lemine Tarbia, Khallahi Brahim, Andy Thorpe, El Hadj Bara Dème, Eli Beibou, Elimane Abou Kane, Moustapha Bouzouma, and Juliana Arias-Hansen. The small pelagics value chain in Mauritania — recent changes and food security impacts. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002378.

Tudela:2022:RPB

[TGM22] Alberto Tudela, Vincent Giroud, and Llibori Martinez. Rebuttal to “Plenty of black money: Netnography of illegal recreational underwater fishing in southern Spain”

Maya-Jariego et al., Marine Policy 126 (2021). Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0308597X22002822. See response [MJ23].

Trueb:2024:IIM

[TGP^{+}24] Laura Trueb, Jordan Gacutan, Teerapong Praphotjanaporn, Maria Alarcon Blazquez, and Ben Milligan. Identifying the institutional modalities for commencing Environmental–Economic Accounting for the ocean. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004081.

Titterton:2022:KPI

[TH22] Haydn Titterton and Marcus Haward. The Kerguelen Plateau: Interactions between the Law of the Sea and the Antarctic Treaty. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000409.

Trebilco:2025:LSE

[THF^{+}25] Rowan Trebilco, Cass Hunter, Mibu Fischer, Karen Evans, Alistair J. Hobday, and Linda Thomas. ‘State of the Environment’ reports require integrated and holistic approaches — lessons from Australia’s marine experience. Marine Policy, 181(??):??, November 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25002696.

Thayer:2020:IEC

[THGR^{+}20] J. A. Thayer, E. L. Hazen, M. García-Reyes, A. Szoboszlai, and W. J. Sydeman. Implementing ecosystem considerations in forage fisheries: San Francisco Bay herring case study. Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304051.

Thompson:2022:BBM

[Tho22] Benjamin S. Thompson. Blue bonds for marine conservation and a sustainable ocean economy: Status, trends, and insights from green bonds. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22002664.

Thompson:2025:ISO

[Tho25] Ruqayyah Thompson. Implementing SDG 14: Opportunities and challenges in a Caribbean Small Island Developing State. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001289.

Teniwut:2022:TSG

[THP22a] Wellem Anselmus Teniwut, Cawalinya Livsanthi Hasyim, and Frischilla Pentury. Towards smart government for sustainable fisheries and marine development: an intelligent web-based support system approach in small islands. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002056.

Thorpe:2022:UTT

[THP^{+}22b] Andy Thorpe, Oystein Hermansen, Iain Pollard, John Isaksen, Pierre Failler, and Grégoire Touron-Gardic. Unpacking the tuna traceability mosaic — EU SFPAs and the tuna value chain. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000847.

Tsvetkova:2024:TTC

[THS^{+}24] Anastasia Tsvetkova, Magnus Hellström, Henry Schwartz, Rodrigo Rabetino, and Hammad Syed. A transition towards clean propulsion in shipping: the role of PESTLE drivers and implications for policy. Marine Policy, 161(??): ??, March 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23005353.

Teniwut:2023:DCC

[THT^{+}23] Wellem Anselmus Teniwut, Syahibul Kahfi Hamid, Roberto Mario Kabi Teniwut, Maimuna Renhoran, and Christy Desta Pratama. Do coastal communities in small islands value marine resources through marine protected areas?: Evidence from Kei Islands Indonesia with choice modelling. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003718.

Thurlow:2021:GAS

[TJ21a] Grace Thurlow and Peter J. S. Jones. A governance analysis of Soufriere Marine Management Area, Saint Lucia: Previously effective but increasingly challenged by driving forces. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308666.

Trouillet:2021:CRB

[TJ21b] Brice Trouillet and Stephen Jay. The complex relationships between marine protected areas and marine spatial planning: Towards an analytical framework. Marine Policy, 127 (??):??, May 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000518.

Tyabji:2022:UTS

[TJS22] Zoya Tyabji, Rima W. Jabado, and Dipani Sutaria. Utilization and trade of sharks and rays in the Andaman Islands, India. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003426.

Tigchelaar:2025:DWF

[TJS^{+}25] Michelle Tigchelaar, Bethany Jackson, Elizabeth R. Selig, Adrianna Davis, Emily O’Regan, Trond Kristiansen, Shinnosuke Nakayama, Doreen Boyd, William W. L. Cheung, Edgar Rodriguez-Huerta, Chris Williams, and Jessica L. Decker Sparks. Decent work in fishing in a changing climate. Marine Policy, 181(??):??, November 2025.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002623.

Twumasi:2021:IGF

[TJZ^{+}21] Martinson Ankrah Twumasi, Yuansheng Jiang, Xiaoshi Zhou, Bismark Addai, Kwabena Nkansah Darfor, Selorm Akaba, and Prince Fosu. Increasing Ghanaian fish farms’ productivity: Does the use of the Internet matter? Marine Policy, 125(??):??, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310368.

Tseng:2022:TRF

[TK22] Huan-Sheng Tseng and Shih-Ming Kao. Towards responsible fisheries: Management of gillnet fishing in Taiwan. Marine Policy, 141(??):??, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001452.

Tseng:2023:ODM

[TK23] Hui-Yi Katherine Tseng and Shih-Ming Kao. Oceanographic data management in Taiwan: an appraisal under the transnational legal ordering perspective. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003159.

Tian:2023:MTL

[TKC23] Huimin Tian, Mehran Idris Khan, and Yen-Chiang Chang. Maritime traffic law of the People’s Republic of China and its implications in international law. Marine Policy, 151 (??):??, May 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000970.

Tokunaga:2023:IFA

[TKP23] Kanae Tokunaga, Lisa A. Kerr, and Andrew J. Pershing. Implications of fisheries allocation policy on anticipated climate change impacts. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004493.

Tian:2022:AUN

[TLT^{+}22] Hao Tian, Yang Liu, Yongjun Tian, Youfu Jing, Shuhao Liu, Xudong Liu, and Yong Zhang. Advances in the use of nighttime light data to monitor and assess coastal fisheries under the impacts of human activities and climate and environmental changes: a case study in the Beibu Gulf. Marine Policy, 144(??):??, October 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002743.

Tang:2020:CPD

[TLYC20] Jianye Tang, Yunzhou Li, Xiaoming Yang, and Yong Chen. Can the presidential diplomacy of China and France promote the East Antarctic marine protected area? Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301299.

Tanaka:2023:EEM

[TM23a] Hajime Tanaka and Mitsutaku Makino. Economic evaluation of MSY-based fishery policy using input–output table: a case study of squid-related industries in Hakodate City, Japan. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003767.

Tiller:2023:IBA

[TM23b] Rachel Tiller and Elizabeth Mendenhall. And so it begins — the adoption of the ‘Biodiversity Beyond National Jurisdiction’ treaty. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300369X.

Teng:2024:BSL

[TMAS24] Fangfei Teng, Usman Mehmood, Hind Alofaysan, and Yunpeng Sun. Bridging shores: Leveraging green finance, financial globalization, and human capital for a cleaner environment in G-20 marine ecosystems using the fishing ground capability curve theory. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2400352X.

Turgeon:2025:AIS

[TMBM^{+}25] Samuel Turgeon, Cristiane C. A. Martins, Camille Bégin-Marchand, Chloé Chartrand, and Nadia Ménard. Automatic Identification System vessel tracking data reports to promote conservation and restoration of the St. Lawrence Estuary beluga habitat. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24004949.

Tiller:2023:SIN

[TMDN23] Rachel Tiller, Elizabeth Mendenhall, Elizabeth De Santo, and Elizabeth Nyman. Shake it off: Negotiations suspended, but hope simmering, after a lack of consensus at the fifth intergovernmental conference on biodiversity beyond national jurisdiction. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22005048.

Tola:2024:DES

[TMG24] Francesco Tola, Enrico Maria Mosconi, and Mattia Gianvincenzi. Demolition of the European ships fleet: a scenario analysis. Marine Policy, 166(??):??, August 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002203.

Tunnicliffe:2020:SEG

[TML^{+}20] Verena Tunnicliffe, Anna Metaxas, Jennifer Le, Eva Ramirez-Llodra, and Lisa A. Levin. Strategic Environmental Goals and Objectives: Setting the basis for environmental regulation of deep seabed mining. Marine Policy, 114:

??, April 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X1830321X.

Toupin:2023:MES

[TMM^{+}23] Rémi Toupin, Kory Melnick, Tamanna Moharana, Blake Curry, Catherine Gracey, Naomi Richards, Bertrum H. MacDonald, and Philippe Mongeon. Measuring engagement with shark research in science, news and social media, and policy: a bibliometric analysis. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003627.

Trialfhianty:2022:SMC

[TMS^{+}22] Tyas Ismi Trialfhianty, Fajrun Wahidil Muharram, Suadi, Claire Helen Quinn, and Maria Beger. Spatial multicriteria analysis to capture socio-economic factors in mangrove conservation. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001415.

Tissiere:2020:FES

[TMT20] Laurie Tissière, Stéphanie Mahévas, and Brice Trouillet. Findings from an exploratory study on the governance of a French fishery. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X16307072.

Trebilco:2020:PRS

[TMTC20] Rowan Trebilco, Jess Melbourne-Thomas, and Andrew John Constable. The policy relevance of Southern Ocean food web structure: Implications of food web change for fisheries, conservation and carbon sequestration. Marine Policy, 115 (??):??, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X1930212X.

Torabi:2021:LRR

[TN21] Foroogh Torabi and Seyed Mohammad Tababaye Nejad. Legal regime of residual liability in decommissioning: the importance of role of states. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003389.

Takyi:2024:CSW

[TNE^{+}24] Richard Takyi, Francis Kofi Ewusie Nunoo, Badr El Mahrad, Cynthia Addo, and John Essandoh. Charting sustainable waters: Governance challenges and opportunities for fisheries and coastal beach resources in a West African country. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003968.

Toygar:2023:WFC

[TNY23] Arda Toygar, Senem Nart, and Umut Yıldırım. Workfamily conflict and work alienation among seafarers: the mediating role of emotional exhaustion. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002920.

Tanner:2021:SDS

[TOAPJ21] Michael K. Tanner, Marcelo Olivares-Arenas, Luciana Puebla, and Jose R. Marin Jarrin. Shifting demand to sustainable fishing practices in Darwin’s Archipelago: a discrete choice experiment application for Galapagos’ certified yellow-fin tuna. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002761.

Todorov:2020:RMA

[Tod20] Andrey Todorov. Russia in maritime areas off Spitsbergen (Svalbard): Is it worth opening the Pandora’s Box? Marine Policy, 122(??):??, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309106.

Todorov:2022:DSR

[Tod22a] Andrey Todorov. Dire straits of the Russian Arctic: Options and challenges for a potential US FONOP in the Northern Sea Route. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000677.

Todorov:2022:SGB

[Tod22b] Andrey Todorov. Shipping governance in the Bering Strait Region: Protecting the Diomede Islands and adjacent waters. Marine Policy, 146(??):??, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003360.

Todorov:2025:GUN

[Tod25] Andrey Todorov. Governing underwater noise from shipping in Antarctica: Current state of affairs and paths forward. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000016.

Tsakiridis:2020:CEE

[TOHK20] Andreas Tsakiridis, Cathal O’Donoghue, Stephen Hynes, and Kevin Kilcline. A comparison of environmental and economic sustainability across seafood and livestock product value chains. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305561.

Tindall:2022:IMM

[TOL^{+}22] Charlotte Tindall, Oluyemisi Oloruntuyi, Samantha Lees, Catherine S. Longo, David Schley, and Rohan J. C. Currey. Illuminating the mechanisms to mitigate forced and child labour risks within Marine Stewardship Council certified fisheries. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001877.

Tong:2022:FDR

[Ton22a] Helong Tong. The future development and restructuring of the international shipping industry: Conference report. Marine Policy, 137(??):??, March 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000033.

Tong:2022:SDW

[Ton22b] Helong Tong. Sustainable development and win–win cooperation in Asian marine tourism: Conference report. Marine Policy, 146(??):??, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003517.

Tong:2024:CBN

[Ton24] Helong Tong. Co-building a new landscape of global shipping based on openness, cooperation, and innovation: Conference report. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24000241.

Toledo:2024:DNN

[TPAC24] Ignacio Toledo, José Ignacio Pagán, Luis Aragonés, and Manuel Benito Crespo. Doing nothing is no solution: Coastal erosion management in Guardamar del Segura (Spain). Marine Policy, 169(??):??, November 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003385.

Tukana:2023:BSF

[TPG^{+}23] Max Tukana, Jeremy Prince, Kerstin B. J. Glaus, Kalisiana Marama, and Cherie Whippy-Morris. A baseline study of Fiji’s small-scale lobster fishery using value chain analysis and size at maturity thresholds. Marine Policy, 149(??): ??, March 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000404.

Trathan:2023:FSG

[Tra23a] Philip N. Trathan. The future of the South Georgia and South Sandwich Islands marine protected area in a changing environment: the choice between industrial fisheries, or ecosystem protection. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23003068.

Trathan:2023:WNI

[Tra23b] Philip N. Trathan. What is needed to implement a sustainable expansion of the Antarctic krill fishery in the Southern Ocean? Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003032.

Thompson:2025:OWP

[TRL^{+}25] Katie R. Thompson, Steve Rocliffe, Alejandro Castillo López, Tinah Martin, S. Hoyt Peckham, Christian Portillo, Xavier Vincke, and Lekelia D. Jenkins. One well-planned and executed exchange is better than two: Lessons learned from a reciprocal fisheries learning exchange between Madagascar and Mexico. Marine Policy, 174(??):??, April 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2500017X.

Toiba:2024:ULB

[TRN^{+}24] Hery Toiba, Moh Shadiqur Rahman, Tri Wahyu Nugroho, Moh Wahyudi Priyanto, Arif Yustian Maulana Noor, and Mohammad Ilyas Shaleh. Understanding the link between climate change adaptation and household food security among shrimp farmers in Indonesia. Marine Policy, 165 (??):??, July 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24002045.

Trudo:2023:PMP

[Tru23] Karin Otsuka Trudo. Perceptions of marine pollution when intertwined with displacement: the case of Miyakojima, Okinawa Prefecture, Japan. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002828.

Teneva:2022:EPO

[TSA^{+}22] Lida Teneva, Aaron L. Strong, Vera Agostini, Kenneth J. Bagstad, Evangelia G. Drakou, Zachary Ancona, Kristina Gjerde, Andrew C. Hume, and Nicholas Jickling. Estimating the pelagic ocean’s benefits to humanity can enhance ocean governance. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005170.

Tseng:2021:MLE

[Tse21] Huan-Sheng Tseng. Maritime law enforcement on bordercrossing fishing vessels: Experience of Taiwan’s practices under its unique political relationship. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21003602.

Temple:2022:IUU

[TSH^{+}22a] Andrew J. Temple, Daniel J. Skerritt, Philippa E. C. Howarth, John Pearce, and Stephen C. Mangi. Illegal, unregulated and unreported fishing impacts: a systematic review of evidence and proposed future agenda. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200080X.

Thorvaldsen:2022:IFM

[TSH22b] Trine Thorvaldsen, Signe A. Sønvisen, and Ingunn M. Holmen. The impact of fisheries management on fishers’ health and safety: a case study from Norway. Marine Policy, 140 (??):??, June 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001130.

Thorbjornsen:2023:HCC

[TSLK23] Susanna Huneide Thorbjørnsen, Ann-Elin Wårøy Synnes, Ingrid Disch Løset, and Alf Ring Kleiven. Hazard and catch composition of ghost fishing gear revealed by a citizen science clean-up initiative. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200478X.

Taniu:2024:IEC

[TSS^{+}24] Silviana Taniu, Dyah Wulan Sari, Dias Satria, Tri Haryanto, and Wahyu Wisnu Wardana. Impact evaluation of cooperative membership on welfare: Evidence from captured fishery households in Indonesia. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004566.

Tsai:2023:OLA

[TSWC23] Liang-Ting Tsai, Tsuyoshi Sasaki, Chin-Kuo Wu, and Cheng-Chieh Chang. Ocean literacy among Taiwanese and Japanese high school students. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000829.

Tirumala:2022:IFM

[TT22] Raghu Dharmapuri Tirumala and Piyush Tiwari. Innovative financing mechanism for blue economy projects. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20303262.

Teh:2020:SSF

[TTA^{+}20] Lydia C. L. Teh, Louise S. L. Teh, Keita Abe, Gakushi Ishimura, and Raphael Roman. Small-scale fisheries in developed countries: Looking beyond developing country narratives through Japan’s perspective. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309209.

Tsirimokos:2025:IID

[TTL^{+}25] Christos Tsirimokos, Irene Tzouramani, Angelos Liontakis, Stamatis Mantziaris, and Alexandra Sintori. Investigating the interindustry dynamics of Greece’s fisheries and aquaculture sector: Effects on production and employment. Marine Policy, 175(??):??, May 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000363.

Talib:2022:TCM

[TUBA22] Naimah Lutfi Talib, Ariane Utomo, Jon Barnett, and Dedi Supriadi Adhuri. Three centuries of marine governance in Indonesia: Path dependence impedes sustainability. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002184.

Tolochko:2021:USD

[TV21] Petro Tolochko and Alice B. M. Vadrot. The usual suspects? Distribution of collaboration capital in marine biodiversity research. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309659.

Tran:2023:AHI

[TVN23] Tuyen Quang Tran, Huong Van Vu, and Thanh Viet Nguyen. Aquaculture, household income and inequality in Vietnam’s coastal region. Marine Policy, 153(??):??, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001616.

Townsend:2022:EDV

[TW22] Ralph E. Townsend and Scott Walker. The economics of deemed values. Marine Policy, 142(??):??, August 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200152X.

Tseng:2025:IMI

[TWK25] Hui-Yi Katherine Tseng, Shu-Chane Wu, and Shih-Ming Kao. ICCAT modernization and its implications for Taiwan’s future participation. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2400438X.

Tsujita:2022:IIC

[TYS22] Takahiro Tsujita, Nobuyuki Yagi, and Yutaro Sakai. Industry-initiated catch limit management: the case of purse seine fishery in Japan’s EEZ waters of the North Pacific Ocean. Marine Policy, 140(??):??, June 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001002.

Tajima:2024:CBF

[TYS24] Satoshi Tajima, Kensuke Yamaguchi, and Hideaki Shiroyama. Consensus building with fishermen on offshore wind farms in Japan: Current status and policy recommendations. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23005080.

Tang:2023:ECB

[TZC23] Hao Tang, Khadija Zulfiqar, and Yen-Chiang Chang. Ecological civilization building and the law of the sea: Conference report. Marine Policy, 157(??):??, November 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004049.

Tanaka:2022:SDM

[TZM^{+}22] Kotaro Tanaka, Mengyao Zhu, Kohei Miyaji, Tadayuki Kurokawa, and Tomonari Akamatsu. Spatial distribution maps of real-time ocean observation platforms and sensors in Japanese waters. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200149X.

Teng:2021:IMF

[TZZ^{+}21] Xin Teng, Qiwei Zhao, Panpan Zhang, Liang Liu, Yue’e Dong, Heng Hu, Qi Yue, Ling Ou, and Wei Xu. Implementing marine functional zoning in China. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18304093.

Uzar:2025:RII

[UE25] Umut Uzar and Kemal Eyuboglu. The role of income inequality in shaping fishing ground footprint in Indonesia: Insights from the Fourier augmented ARDL approach. Marine Policy, 176(??):??, June 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000508.

Uriarte:2023:LLM

[UISM^{+}23] Andrés Uriarte, Leire Ibaibarriaga, Sonia Sánchez-Maroño, Pablo Abaunza, Marga Andrés, Erwan Duhamel, Ernesto Jardim, Lionel Pawlowski, Raúl Prellezo, and Beatriz A. Roel. Lessons learnt on the management of short-lived fish from the Bay of Biscay anchovy case study: Satisfying fishery needs and sustainability under recruitment uncertainty. Marine Policy, 150(??):??, April 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000398.

Uduji:2020:DCS

[UOO20] Joseph Ikechukwu Uduji and Elda Nduka Okolo-Obasi. Does corporate social responsibility (CSR) impact on development of women in small-scale fisheries of sub-Saharan Africa? Evidence from coastal communities of Niger Delta in Nigeria. Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18306195.

Umana:2025:MPC

[UPBB25] Juan M. Zúñiga Umaña, Sergio A. Paniagua, Rebeca Corrales Brenes, and José R. Vega Baudrit. Microplastic pollution in Costa Rican marine ecosystems: Origins, ecotoxicological impacts, and mitigation strategies. Marine Policy, 179(??):??, September 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001873.

Urbina:2022:TIL

[Urb22] Julio Jorge Urbina. Towards an international legal definition of the notion of fisheries crime. Marine Policy, 144(??): ??, October 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22002615.

Vis:2024:BCM

[VAG^{+}24] Chantal Vis, Karla Abbott, Niisii Guujaaw, Cavan Harpur, Lynn Lee, Pippa Shepherd, Douglas P. Tate, Samuel Turgeon, and Blythe Beynon. Building collaborative monitoring programs to manage for ecological sustainability in multi-use marine protected areas: Parks Canada’s national marine conservation areas. Marine Policy, 167(??): ??, September 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24002471.

Vilas:2020:UCV

[VAMR^{+}20] C. Vilas, L. T. Antelo, F. Martin-Rodriguez, X. Morales, R. I. Perez-Martin, A. A. Alonso, J. Valeiras, E. Abad, M. Quinzan, and M. Barral-Martinez. Use of computer vision onboard fishing vessels to quantify catches: the iObserver. Marine Policy, 116(??):??, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308340.

Vieira:2025:PII

[VAS^{+}25] Helena Vieira, Mariana Almeida, Muhammad Noman Shafique, Miguel C. Leal, and Ana I. Lillebø. Policy impacts of intellectual property and academic research trends in marine-derived collagen and chitin/chitosan value chains. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400575X.

Visalli:2020:DDA

[VBC^{+}20] Morgan E. Visalli, Benjamin D. Best, Reniel B. Cabral, William W. L. Cheung, Nichola A. Clark, Cristina Garilao, Kristin Kaschner, Kathleen Kesner-Reyes, Vicky W. Y. Lam, Sara M. Maxwell, Juan Mayorga, Holly V. Moeller, Lance Morgan, Guillermo Ortuño Crespo, Malin L. Pinsky, Timothy D. White, and Douglas J. Mc-Cauley. Data-driven approach for highlighting priority areas for protection in marine areas beyond national jurisdiction. Marine Policy, 122(??):??, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309194.

VargasBarbosaFilho:2021:RFP

[VBd^{+}21] Márcio Luiz Vargas Barbosa Filho, Gabriel Barros Gonçalves de Souza, Sérgio de Faria Lopes, Rachel Ann Hauser-Davis, Salvatore Siciliano, and José da Silva Mourão. Reef fisher perceptions acknowledge the socioenvironmental effectiveness of a 20-year old Brazilian Marine Protected Area. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004085.

Verling:2023:ARB

[VBH^{+}23] Emma Verling, Cátia Bartilotti, Claudia Hollatz, Miriam Tuaty-Guerra, Jorge Lobo-Arteaga, and Tim O’Higgins. Applying risk-based approaches to implementation of the Marine Strategy Framework Directive in the North-East Atlantic: Learning lessons and moving forward. Marine Policy, 153(??):??, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2300194X.

VanDover:2020:RNI

[VCC^{+}20] C. L. Van Dover, A. Colaço, P. C. Collins, P. Croot, A. Metaxas, B. J. Murton, A. Swaddling, R. E. Boschen-Rose, J. Carlsson, L. Cuyvers, T. Fukushima, A. Gartman, R. Kennedy, C. Kriete, N. C. Mestre, T. Molodtsova, A. Myhrvold, E. Pelleter, S. O. Popoola, P.-Y. Qian, J. Sarrazin, R. Sharma, Y. J. Suh, J. B. Sylvan, C. Tao, M. Tomczak, and J. Vermilye. Research is needed to inform environmental management of hydrothermally inactive and extinct polymetallic sulfide (PMS) deposits. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20302943.

Vilas:2021:CPC

[VCC^{+}21] Daniel Vilas, Marta Coll, Xavier Corrales, Jeroen Steenbeek, Chiara Piroddi, Diego Macias, Alessandro Ligas, Paolo Sartor, and Joachim Claudet. Current and potential contributions of the Gulf of Lion Fisheries Restricted Area to fisheries sustainability in the NW Mediterranean Sea. Marine Policy, 123(??):??, January 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309416.

vandenBurg:2023:MEM

[vdBSvdV^{+}23] S. W. K. van den Burg, M. Skirtun, O. van der Valk, W. Rossi Cervi, T. Selnes, T. Neumann, J. Steinmann, G. Arora, and P. Roebeling. Monitoring and evaluation of maritime spatial planning — a review of accumulated practices and guidance for future action. Marine Policy, 150 (??):??, April 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000568.

vanderGrient:2021:PSI

[vdGD21] J. M. A. van der Grient and J. C. Drazen. Potential spatial intersection between high-seas fisheries and deep-sea mining in international waters. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001755.

Vuong:2024:OEC

[VDN^{+}24] Quan-Hoang Vuong, Minh-Phuong Thi Duong, Quynh-Yen Thi Nguyen, Viet-Phuong La, Phuong-Tri Nguyen, and Minh-Hoang Nguyen. Ocean economic and cultural benefit perceptions as stakeholders’ constraints for supporting conservation policies: a multi-national investigation. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001325.

Verfuss:2024:EMA

[VDOE^{+}24] U. K. Verfuss, A. K. Darias-O’Hara, C. Erbe, D. Houser, V. M. Janik, D. Ketten, K. Lucke, M. Morell, A. Pacini, C. Reichmuth, and C. G. Booth. Eliciting the magnitude of auditory threshold shift considered injury in Antarctic marine mammals. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004529.

Vecchio:2022:PHR

[VDP^{+}22] Y. Vecchio, J. Di Pasquale, G. Pauselli, M. Masi, and F. Adinolfi. Public health risk management during the Covid-19 pandemic, new amendments in the European Maritime and Fisheries Fund to meet fishers’ needs. Marine Policy, 135(??):??, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100484X.

Vandenberg:2025:ISI

[VFOA25] Jessica M. Vandenberg, Trisia Farrelly, Yoshitaka Ota, and Holly Amos. Introduction to the special issue: Marine plastic pollution is an equity issue. Marine Policy, 171(??): ??, January 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24005013.

Villamizar:2025:AEC

[VGA25] Javier Seijo Villamizar and Antonio García-Allut. Analysing the epistemic consensus of fisheries co-management systems as a formula for democratic innovation and open government. Marine Policy, 176(??):??, June 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25000661.

Villegas:2021:CCS

[VGAH^{+}21] Ciera Villegas, Nicolás X. Gómez-Andújar, Michael Harte, Sarah M. Glaser, and James R. Watson. Cooperation and conflict in the small-scale fisheries of Puerto Rico. Marine Policy, 134(??):??, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004206.

Villasante:2022:SES

[VGC^{+}22] Sebastian Villasante, Ignacio Gianelli, Mauricio Castrejón, Laura Nahuelhual, Leonardo Ortega, U. Rashid Sumaila, and Omar Defeo. Social-ecological shifts, traps and collapses in small-scale fisheries: Envisioning a way forward to transformative changes. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005443.

Vukic:2023:APE

[VGK23] Luka Vukić, Giambattista Guidi, and Iven Krämer. Air pollutant emission calculation and shipping costs: the case of yacht transport between the North Sea and Mediterranean ports. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004948.

Valdor:2020:GAM

[VGS^{+}20] Paloma F. Valdor, Aina G. Gómez, Peter Steinberg, Edwina Tanner, Antony M. Knights, Rochelle D. Seitz, Laura Airoldi, Louise B. Firth, Christos Arvanitidis, Massimo Ponti, Eva Chatzinikolaou, Paul R. Brooks, Tasman P. Crowe, Alison Smith, Gonzalo Méndez, Aida Ovejero, Abilio Soares-Gomes, John A. Burt, Catriona MacLeod, and José A. Juanes. A global approach to mapping the environmental risk of harbours on aquatic systems. Marine Policy, 119(??):??, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302684.

Vidal-Hernandez:2024:PPI

[VHCJdYC^{+}24] L. Vidal-Hernández, A. Cuevas-Jiménez, D. de Yta-Castillo, V. S. Ávila-Foucat, and Ileana Espejel. Public policy instruments coherence analysis to address coastal risk in Yucatan, Mexico. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24002781.

Vierros:2020:CBT

[VHD20] Marjo K. Vierros and Harriet Harden-Davies. Capacity building and technology transfer for improving governance of marine areas both beyond and within national jurisdiction. Marine Policy, 122(??):??, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301469.

Vierros:2020:CIP

[VHS^{+}20] Marjo K. Vierros, Autumn-Lynn Harrison, Matthew R. Sloat, Guillermo Ortuño Crespo, Jonathan W. Moore, Daniel C. Dunn, Yoshitaka Ota, Andrés M. Cisneros-Montemayor, George L. Shillinger, Trisha Kehaulani Watson, and Hugh Govan. Considering Indigenous Peoples and local communities in governance of the global ocean commons. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309212.

vanHoof:2020:CPC

[vHSSK20] L. van Hoof, N. A. Steins, S. Smith, and M. Kraan. Change as a permanent condition: a history of transition processes in Dutch North Sea fisheries. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20308915.

VanAssche:2020:GCC

[VHSV20] Kristof Van Assche, Anna-Katharina Hornidge, Achim Schlüter, and Nataşa Vaidianu. Governance and the coastal condition: Towards new modes of observation, adaptation and integration. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18303865.

Veneroni:2024:PRD

[VJ24] Benedetta Veneroni and Rikke Becker Jacobsen. The price of regionalisation: Discursive dominance and stakeholders coalitions in the Northern Adriatic Sea fishery governance arrangement. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001118.

Vescovo:2021:SCD

[VJL^{+}21] Victor L. Vescovo, Alan J. Jamieson, Patrick Lahey, Rob McCallum, Heather A. Stewart, and Casey Machado. Safety and conservation at the deepest place on Earth: a call for prohibiting the deliberate discarding of nondegradable umbilicals from deep-sea exploration vehicles. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000749.

Vigfusson:2023:SIO

[VJO23] Kristjan Reykjalin Vigfusson, Lara Johannsdottir, and Snjolfur Olafsson. Strategy implementation obstacles: Iceland fishery CEO perspectives. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22004006.

Vigfusson:2025:IRB

[VJÓK25] Kristján Vigfússon, Lára Jóhannsdóttir, Snjólfur Ólafsson, and Mehmet Ali Köseoglu. Integrating resource-based and market-based views in the fisheries industry – CEO perspectives. Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004548.

Gerven:2025:FIA

[VKL^{+}25] A. Van Gerven, T. R. H. Kerkhove, M. Lago, A. Araujo, M. Berge, O. Mashkina, A. Ziemba, G. El Serafy, J. Tamis, R. Jongbloed, G. Piet, M. J. C. Rozemeijer, and S. Degraer. Facilitating an integrated assessment of impacts in marine multi-use: the Ocean Multi-use Assessment Framework (OMAF). Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005700.

Voldnes:2020:TCH

[VKN20] Gøril Voldnes, Ingrid Kvalvik, and Bjørg Nøstvold. Taking care of a highly valuable resource throughout the value chain — Lack of market orientation in red king crab export? Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303914.

Vinata:2023:CCR

[VKS23] Ria Tri Vinata, Masitha Tismananda Kumala, and Cita Yustisia Serfiyani. Climate change and reconstruction of Indonesia’s geographic basepoints: Reconfiguration of baselines and Indonesian Archipelagic Sea lanes. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004900.

Vye:2024:AFC

[VLG^{+}24] Siobhan R. Vye, Janine Lavarello, Trevor Glass, James Glass, Jéssica Escobar-Porras, and Andy Schofield. Adaptive, flexible and community-led: Management support for the largest fully protected marine area in the Atlantic. Marine Policy, 167(??):??, September 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002823.

Vogel:2024:FFB

[VLL^{+}24] Jacqueline M. Vogel, Arielle Levine, Catherine Longo, Rod Fujita, Catherine L. Alves, Gemma Carroll, J. Kevin Craig, Kiley Dancy, Melissa Errend, Timothy E. Essington, Nima Farchadi, Sarah Glaser, Abigail S. Golden, Olaf P. Jensen, Monica LeFlore, Julia G. Mason, Katherine E. Mills, Juliano Palacios-Abrantes, Anthony Rogers, Jameal F. Samhouri, Matthew Seeley, Elizabeth R. Selig, Ashley Trudeau, and Colette C. C. Wabnitz. Fisheries in flux: Bridging science and policy for climateresilient management of US fisheries under distributional change. Marine Policy, 170(??):??, December 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400383X.

vanLeeuwen:2022:DSS

[vLM22] Judith van Leeuwen and Jason Monios. Decarbonisation of the shipping sector — time to ban fossil fuels? Marine Policy, 146(??):??, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003578.

vanLeeuwen:2021:MMS

[vLPBK21] Wessel M. A. van Leeuwen, Claire Pekcan, Mike Barnett, and Göran Kecklund. Mathematical modelling of sleep and sleepiness under various watch keeping schedules in the maritime industry. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309234.

Vogel:2023:DCR

[VLS^{+}23] Jacqueline M. Vogel, Catherine Longo, Jessica Spijkers, Juliano Palacios-Abrantes, Julia Mason, Colette C. C. Wabnitz, William Cheung, U. Rashid Sumaila, Gordon Munro, Sarah Glaser, Johann Bell, Yongjun Tian, Nancy L. Shackell, Elizabeth R. Selig, Philippe Le Billon, James R. Watson, Cullen Hendrix, Malin L. Pinsky, Ingrid van Putten, Kendra Karr, Eva A. Papaioannou, and Rod Fujita. Drivers of conflict and resilience in shifting transboundary fisheries. Marine Policy, 155(??):??, September 2023.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002737.

vanLeeuwen:2024:NPE

[vLS24] Judith van Leeuwen and Ilham R. F. Surya. Network power and exclusion of informal waste pickers when plastic flows change: a case study of community waste banks in Klaten Municipality in Indonesia. Marine Policy, 167(??): ??, September 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24002835.

Vandergeest:2021:BSS

[VM21] Peter Vandergeest and Melissa Marschke. Beyond slavery scandals: Explaining working conditions among fish workers in Taiwan and Thailand. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21002967.

Venturini:2021:RBP

[VMC^{+}21] S. Venturini, F. Massa, M. Castellano, G. Fanciulli, and P. Povero. Recreational boating in the Portofino Marine Protected Area (MPA), Italy: Characterization and analysis in the last decade (2006–2016) and some considerations on management. Marine Policy, 127(??):??, May 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X17306942.

Vecchio:2024:TIF

[VMD^{+}24] Yari Vecchio, Margherita Masi, Teresa Del Giudice, Marcello De Rosa, and Felice Adinolfi. Technological innovation in fisheries and aquaculture: What are the “discourses” of the Italian policy network? Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004803.

Veylit:2025:SIC

[VMJ25] Lara Veylit, Magnus Stoud Myhre, and Sepideh Jafarzadeh. Sustainability implications of COVID-19 related seafood supply chain disruptions in Norway. Marine Policy, 172(??): ??, February 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24004883.

Vandergeest:2021:SFY

[VMM21] Peter Vandergeest, Melissa Marschke, and Mallory Mac-Donnell. Seafarers in fishing: a year into the COVID-19 pandemic. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004073.

Villasante:2024:ICP

[VMP^{+}24] Sebastian Villasante, Arantza Murillas, Pablo Pita, Ana Tubío, Jose J. Pascual-Fernández, Guillherme Arangao, Joan Moranta, Marta Coll, Andrés Ospina-Alvarez, Silvia de Juan, Ana Ruiz-Frau, Raul Prellezo, José María Da Rocha, Jeroen Steenbeek, Javier Seijo, Maria Grazia Pennino, and Gillian B. Ainsworth. Impacts of the COVID-19 pandemic on the Spanish seafood sector. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002914.

VanRoy:2024:PRI

[VMV^{+}24a] Ward Van Roy, Jean-Baptiste Merveille, Annelore Van Nieuwenhove, Kobe Scheldeman, and Frank Maes. Policy recommendations for international regulations addressing air pollution from ships. Marine Policy, 159(??):??, January 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004463.

Vasquez:2024:BSF

[VMV24b] Julio A. Vásquez, Cristian Morales, and Andrés Vallone. Brown seaweeds fishery and copper mining production: Two distant economic industries connected by socioecological impacts in Northern Chile. Marine Policy, 165(??):??, July 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24001891.

vanNorren:2025:OMR

[vNdB25] Dorine Eva van Norren and Chris de Blok. Ocean management and rights of nature: the case of the Galapagos in Ecuador and beyond. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24005645.

Vakili:2020:DPFa

[VÖB20a] S. V. Vakili, A. I. Ölcer, and F. Ballini. The development of a policy framework to mitigate underwater noise pollution from commercial vessels. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19308942.

Vakili:2020:DPFb

[VÖB20b] S. V. Vakili, A. I. Ölçer, and F. Ballini. The development of a policy framework to mitigate underwater noise pollution from commercial vessels: the role of ports. Marine Policy, 120(??):??, October 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309480.

Virgili:2024:EFS

[VPB^{+}24] M. Virgili, A. Petetta, G. Barone, D. Li Veli, G. Bargione, and A. Lucchetti. Engaging fishers in sea turtle conservation in the Mediterranean Sea. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23005146.

Vilela:2021:USM

[VPRR^{+}21] Raul Vilela, Maria Grazia Pennino, Gonzalo Rodriguez-Rodriguez, Hugo M. Ballesteros, and Jose Maria Bellido. The use of a spatial model of economic efficiency to predict the most likely outcomes under different fishing strategy scenarios. Marine Policy, 129(??):??, July 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100110X.

vanRensburg:2024:UPP

[vRB24] Thomas M. van Rensburg and Noreen Brennan. Understanding public preferences towards Ireland’s offshore wind sector: a study on renewable energy trade, public involvement, and setback distance. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23005213.

Vazquez-Rowe:2024:ABB

[VRBBM^{+}24] Ian Vázquez-Rowe, Karen Biberos-Bendezú, Luis Moya, Alejandro Deville, Diana Ita-Nagy, and Ramzy Kahhat. Analyzing the behavior of beachgoers in the city of Lima and their relationship with potential plastic emissions. Marine Policy, 170(??):??, December 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004019.

Verling:2021:ARB

[VRBC^{+}21] Emma Verling, Ramon Miralles Ricós, Manuel Bou-Cabo, Guillermo Lara, Maria Garagouni, Jean-Marc Brignon, and Tim O’Higgins. Application of a risk-based approach to continuous underwater noise at local and subregional scales for the Marine Strategy Framework Directive. Marine Policy, 134(??):??, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003973.

Vaughan:2023:RFT

[VSD^{+}23] Duncan Vaughan, Daniel J. Skerritt, James Duckworth, U. Rashid Sumaila, and Mark Duffy. Revisiting fuel tax concessions (FTCs): the economic implications of fuel subsidies for the commercial fishing fleet of the United Kingdom. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002968.

Villasante:2022:SEU

[VSDR^{+}22] Sebastián Villasante, U. Rashid Sumaila, Jose María Da-Rocha, Natacha Carvalho, Daniel J. Skerritt, Anna Schuhbauer, Andrés M. Cisneros-Montemayor, Nathan J. Bennett, Quentin Hanich, and Raúl Prellezo. Strengthening European Union fisheries by removing harmful subsidies. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004954.

vonSchuckmann:2020:OSD

[vSHHT20] Karina von Schuckmann, Elisabeth Holland, Peter Haugan, and Peter Thomson. Ocean science, data, and services for the UN 2030 Sustainable Development Goals. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20302050.

Vergara-Solana:2022:VVM

[VSPMS^{+}22] Francisco Vergara-Solana, Daniel Peñalosa-Martinell, Daniel Skerritt, Annie Mejaes, German Ponce-Diaz, Fernando Aranceta-Garza, Fernando González-Laxe, Juan Carlos Seijo, and U. Rashid Sumaila. Volatility and vulnerability in Mexican fisheries and aquaculture: Enhancing resilience via public policy. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004991.

VanRoy:2023:CPD

[VSV^{+}23] Ward Van Roy, Kobe Scheldeman, Annelore Van Nieuwenhove, Jean-Baptiste Merveille, Ronny Schallier, and Frank Maes. Current progress in developing a MARPOL Annex VI enforcement strategy in the Bonn Agreement through remote measurements. Marine Policy, 158(??):??, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23004153.

vonThenen:2021:GMP

[vTHS21] Miriam von Thenen, Henning Sten Hansen, and Kerstin S. Schiele. A generalised marine planning framework for site selection based on ecosystem services. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309738.

Vaidianu:2020:MCP

[VTRS20] Nataşa Văidianu, Florin Tătui, Mădălina Ristea, and Adrian Stănică. Managing coastal protection through multi-scale governance structures in Romania. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X18304123.

Virdin:2022:CIF

[VVR^{+}22] John Virdin, Tibor Vegh, Blake Ratcliff, Elizabeth Havice, Jack Daly, and Jack Stuart. Combatting illegal fishing through transparency initiatives: Lessons learned from comparative analysis of transparency initiatives in seafood, apparel, extractive, and timber supply chains. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22000318.

Vrooman:2022:UDR

[VvSvH^{+}22] Jip Vrooman, Christiaan van Sluis, Floris van Hest, Han Lindeboom, and Albertinka Murk. Unambiguously defined and recognized seabed protection targets are necessary for successful implementation of MPAs. Marine Policy, 140 (??):??, June 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001038.

VanDiggelen:2022:CLL

[VWFW22] Amanda D. Van Diggelen, Sara E. Worden, Adam J. Frimodig, and Stephen P. Wertz. California’s lessons learned and recommendations for effective marine protected area network management. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2100539X.

Vye:2020:EPM

[VWJMAJ20] Siobhan R. Vye, Sophie Wynne-Jones, Patricia Masterson-Algar, and Stuart R. Jenkins. Exploring perceptions of marine biosecurity interventions: insights from the commercial marina sector. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302222.

Vylegzhanin:2020:CAO

[VYB20] Alexander N. Vylegzhanin, Oran R. Young, and Paul Arthur Berkman. The Central Arctic Ocean Fisheries Agreement as an element in the evolving Arctic Ocean governance complex. Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20301780.

Wang:2024:GAM

[WA24] Weishan Wang and Claudio Aporta. Arctic marine shipping development and governance in Canada: a historical overview. Marine Policy, 160(??):??, February 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004918.

Weaver:2022:API

[WABR^{+}22] P. P. E. Weaver, J. Aguzzi, R. E. Boschen-Rose, A. Colaço, H. de Stigter, S. Gollner, M. Haeckel, C. Hauton, R. Helmons, D. O. B. Jones, H. Lily, N. C. Mestre, C. Mohn, and L. Thomsen. Assessing plume impacts caused by polymetallic nodule mining vehicles. Marine Policy, 139(??): ??, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000586.

Wang:2023:SCM

[Wan23a] Chuanliang Wang. The significance of the concept of a maritime community with a shared future in the international law-making process: Taking the BBNJ Agreement as an example. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000362.

Wang:2023:ILM

[Wan23b] Sen Wang. International law-making process of combating plastic pollution: Status Quo, debates and prospects. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22004237.

Wang:2023:URG

[Wan23c] Weishan Wang. UNDRIP rights to guide the governance of the Northern Low-impact Shipping Corridors initiative. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002701.

Wang:2024:CGC

[Wan24a] Qiuwen Wang. Charting a greener course: China’s recent legal strategies in shore power promotion for achieving maritime emission targets. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24003841.

Wang:2024:MLE

[Wan24b] Qiuwen Wang. Maritime law enforcement concerning offshore energy platforms: Navigating international law constraints and challenges. Marine Policy, 170(??):??, December 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24003683.

Wang:2024:ISF

[Wan24c] Wei Wang. Innovative strategies and forward thinking on China’s digital maritime law enforcement. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24003671.

Warner:2020:IEL

[War20] Robin Warner. International environmental law principles relevant to exploitation activity in the Area. Marine Policy, 114:??, April 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X18303609.

Ward:2023:FAP

[War23] Sarah Ward. The Fourth Asia Pacific Regional Conference on Underwater Cultural Heritage: a marine policy perspective. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004055.

Waters:2022:EBD

[Wat22] Yolanda L. Waters. Establishing a baseline decision context for restoring the Great Barrier Reef — implications for longterm management. Marine Policy, 141(??):??, July 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001531.

Watson:2025:UCO

[Wat25] Amanda H. A. Watson. Undersea cables: the official perspectives expressed in the Pacific region. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001502.

Wei:2024:SRB

[WB24] Jifei Wei and Junaid Bilal. Study on the relationship between economic input growth in Marine Protected Areas and environmental governance in the North Pacific Coast of Costa Rica: Based on reef fish assemblages. Marine Policy, 169(??):??, November 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003117.

Webb:2025:RPI

[WBB^{+}25] Thomas J. Webb, Joana Beja, Salvador Jesús Fernández Bejarano, Elvira Ramos, Samuel Sainz-Villegas, Karline Soetaert, Willem Stolte, Charles Troupin, and Benjamin Weigel. Realising the potential of interoperable data products to improve the outlook for marine biodiversity: Lessons from the European marine observation and data network. Marine Policy, 173(??):??, March 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005785.

Willis:2023:HCG

[WBH^{+}23] Sam Willis, Despoina Andrioti Bygvraa, Md. Sazedul Hoque, Emily S. Klein, Cagri Kucukyildiz, Joseph Westwood-Booth, and Eric Holliday. The human cost of global fishing. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004870.

Wilson:2020:BPF

[WBPG20] Jono R. Wilson, Darcy Bradley, Kristina Phipps, and Mary G. Gleason. Beyond protection: Fisheries co-benefits of no-take marine reserves. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20308708.

Wang:2020:IEI

[WCC^{+}20] Grace Wang, Wen-Huei Chang, Yue Cui, Guanqiu Qi, and Kevin X. Li. Introducing an economic impact platform to navigate cruise value-added chain with environmental considerations. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19303331.

Weiss:2021:MAE

[WCF^{+}21] Sarah G. Weiss, Danielle Cholewiak, Kaitlin E. Frasier, Jennifer S. Trickey, Simone Baumann-Pickering, John A. Hildebrand, and Sofie M. Van Parijs. Monitoring the acoustic ecology of the shelf break of Georges Bank, Northwestern Atlantic Ocean: New approaches to visualizing complex acoustic data. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001810.

Wei:2025:AME

[WCL25] Youzhou Wei, Qing Cui, and Yihan Liu. An assessment for marine economic resilience and exploration of enhancement strategies: a case study of China. Marine Policy, 174(??): ??, April 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25000120.

Wood:2024:MNI

[WCM^{+}24] Louisa E. Wood, Stacey A. Clarke, Debbie Murphy, Phil I. Davison, Mark A. Thrush, David Bass, Silvana N. R. Birchenough, Edmund J. Peeler, and Hannah J. Tidbury. Monitoring of non-indigenous marine species for legislative and policy goals in the UK. Marine Policy, 162(??):??, April 2024. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24000253.

Wilson:2025:FPI

[WCN^{+}25] David Wilson, Elias Chirwa, Bryson Nkhoma, Milo Gough, Charles W. Knapp, Tracy Morse, and Wapulumuka Mulwafu. Fishing (in) the past to inform the future: Lessons from the histories of fisheries management in Lake Malawi and Mbenji Island. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000041.

Wilson:2020:ERI

[WCT^{+}20] Susan C. Wilson, Imogen Crawford, Irina Trukhanova, Lilia Dmitrieva, and Simon J. Goodman. Estimating risk to ice-breeding pinnipeds from shipping in Arctic and sub-Arctic seas. Marine Policy, 111(??):??, January 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302039.

Wang:2021:MEW

[WCZS21] Shuhong Wang, Suisui Chen, Hongyan Zhang, and Malin Song. The model of early warning for China’s marine ecology-economy symbiosis security. Marine Policy, 128 (??):??, June 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21000877.

Ware:2020:CHM

[WD20] Suzanne Ware and Anna-Leena Downie. Challenges of habitat mapping to inform marine protected area (MPA) designation and monitoring: an operational perspective. Marine Policy, 111(??):??, January 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19301022.

Wood-Donnelly:2022:IS

[WD22] Corine Wood-Donnelly. Iceberg sovereignty. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001865.

Walden:2023:BEP

[WD23] John B. Walden and Geret DePiper. Benchmarking ecosystem performance using index numbers: a case study of the northeast large marine ecosystem. Marine Policy, 155(??): ??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002415.

Williams:2020:ECS

[WDC^{+}20] Chris Williams, William Davies, Robert E. Clark, Angela Muench, and Kieran Hyder. The economic contribution of sea angling from charter boats: a case study from the south coast of England. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X1930795X.

Wang:2020:FNS

[WDG^{+}20] Dan Wang, Renke Ding, Yu Gong, Rui Wang, Jie Wang, and Xiaoling Huang. Feasibility of the Northern Sea Route for oil shipping from the economic and environmental perspective and its influence on China’s oil imports. Marine Policy, 118(??):??, August 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301792.

Watterson:2023:CSL

[WdSRY23] Christopher J. Watterson, Jean-Annet de Saint Rapt, and Eunha Yun. China’s state-led international port development: Challenges for port states. Marine Policy, 155(??): ??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002245.

Willette:2021:LMC

[WEF^{+}21] Demian A. Willette, Sofia C. Esteves, Ben Fitzpatrick, Marie L. Smith, Kesterlyn Wilson, and Xiaoya Yuan. The last mile challenge: Certified seafood and federal labeling laws out of sync at the end of the supply chain in Los Angeles, California. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310319.

Wendebourg:2020:SOF

[Wen20] Mara R. Wendebourg. Southern Ocean fishery management — is CCAMLR addressing the challenges posed by a changing climate? Marine Policy, 118(??):??, August 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18308121.

Wang:2025:TKA

[WF25] Jinpeng Wang and Xiaohan Fan. Traditional knowledge of Arctic Indigenous Peoples and the establishment of area-based management tools beyond national jurisdiction. Marine Policy, 174(??):??, April 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000193.

Weitzman:2022:IKF

[WFG22a] Jenny Weitzman, Ramón Filgueira, and Jon Grant. Identifying key factors driving public opinion of salmon aquaculture. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002226.

Wichmann:2022:PPE

[WFG^{+}22b] Carla-Sophie Wichmann, Daniel Fischer, Sonja Maria Geiger, Daniela Honorato-Zimmer, Katrin Knickmeier, Katrin Kruse, Anna Sundermann, and Martin Thiel. Promoting pro-environmental behavior through citizen science? A case study with Chilean schoolchildren on marine plastic pollution. Marine Policy, 141(??):??, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000823.

Wrobel:2022:TME

[WFG22c] Krzysztof Wróbel, Kamil Formela, and Mateusz Gil. Towards maritime employment and retention policy — a longitudinal study on entrants’ perspectives. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002792.

Wahju:2024:PSA

[WFK^{+}24] Ronny Irawan Wahju, Faridz Rizal Fachri, Mohammad Mukhlis Kamal, Yu-Jia Lin, Achmad Mustofa, Teo Andri Saputra, Endan Sutendi, Makoto Yoshida, Shuhei Uematsu, and Kenzo Kaifu. Participatory stock assessment in West Java contributes to the management of glass eel fisheries in Indonesia. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001015.

Wenjun:2025:CRC

[WFP25] Li Wenjun, Gao Fei, and Guo Peiqing. 4China–Russia cooperation in Arctic governance and sustainable development: a conference report from the 13th China–Russia Arctic Workshop. Marine Policy, 175(??):??, May 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000648.

Wahler:2023:VBB

[WFR23] Hannah Wahler, Lucia Fanini, and Maraja Riechers. Valuing beaches for beauty and recreation only? Uncovering perception bias through a hashtag analysis. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23002403.

White:2025:WTL

[WGBS25] Shelby B. White, Barbara Garrity-Blake, and Andrew M. Scheld. What they live for: Adaptive strategies of Virginia’s small-scale commercial fishermen in a changing industry. Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004512.

Wright:2021:MSP

[WGJ^{+}21] Glen Wright, Kristina M. Gjerde, David E. Johnson, Aria Finkelstein, Maria Adelaide Ferreira, Daniel C. Dunn, Mariamalia Rodriguez Chaves, and Anthony Grehan. Marine spatial planning in areas beyond national jurisdiction. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304408.

Wongbusarakum:2021:MCA

[WGP^{+}21] Supin Wongbusarakum, Matt Gorstein, Robert Pomeroy, Cheryl L. Anderson, and Alexander Mawyer. Mobilizing for change: Assessing Social adaptive capacity in Micronesian fishing communities. Marine Policy, 129(??):??, July 2021.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001196.

Wedding:2022:LMS

[WGR^{+}22] L. M. Wedding, S. J. Green, S. Reiter, K. R. Arrigo, L. Hazen, M. Ruckelshaus, J. M. A. van der Grient, R. M. Bailey, M. A. Cameron, J. Leape, M. Levi, A. Merkl, M. M. Mills, S. Monismith, N. T. Ouellette, G. van Dijken, and F. Micheli. Linking multiple stressor science to policy opportunities through network modeling. Marine Policy, 146(??):??, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003542.

Winter:2020:IFR

[WH20] Anna-Marie Winter and Jeffrey A. Hutchings. Impediments to fisheries recovery in Canada: Policy and institutional constraints on developing management practices compliant with the precautionary approach. Marine Policy, 121(??): ??, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19308358.

White:2025:MMM

[WHAD^{+}25] J. Wilson White, Jess K. Hopf, Nur Arafeh-Dalmau, Natalie C. Ban, Amanda E. Bates, Joachim Claudet, Cori Lopazanski, Jennifer M. Sunday, and Jennifer E. Caselle. Measurements, mechanisms, and management recommendations for how marine protected areas can provide climate resilience. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004196.

Wang:2022:ICP

[WHL22] Qiang Wang, Rui Huang, and Rongrong Li. Impact of the COVID-19 pandemic on research on marine plastic pollution — a bibliometric-based assessment. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22003323.

Whomersley:2023:UWA

[Who23] Chris Whomersley. UNCLOS at 40: What about human rights? Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22004717.

Wang:2020:EBF

[WHPF20] Ying Wang, Jianfeng Hu, Haoran Pan, and Pierre Failler. Ecosystem-based fisheries management in the Pearl River Delta: Applying a computable general equilibrium model. Marine Policy, 112(??):??, February 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19306001.

Willaert:2020:CPD

[Wil20] Klaas Willaert. Crafting the perfect deep sea mining legislation: a patchwork of national laws. Marine Policy, 119(??): ??, September 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19308954.

Willaert:2021:DSM

[Wil21a] Klaas Willaert. Deep sea mining and the United States: Unbound powerhouse or odd man out? Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20309891.

Willaert:2021:ESA

[Wil21b] Klaas Willaert. The Enterprise: State of affairs, challenges and way forward. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21002013.

Willaert:2022:TFD

[Wil22] Klaas Willaert. Transparency in the field of deep sea mining: Filtering the murky waters. Marine Policy, 135(??):

??, January 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21004516.

Waldo:2020:LAT

[WJB^{+}20] Åsa Waldo, Maria Johansson, Johan Blomquist, Torbjörn Jansson, Sara Königson, Sven-Gunnar Lunneryd, Anders Persson, and Staffan Waldo. Local attitudes towards management measures for the co-existence of seals and coastal fishery — a Swedish case study. Marine Policy, 118(??): ??, August 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19306098.

Watson:2021:CMP

[WJLSH21] Maryann S. Watson, Anne-Marie Jackson, Georgia Lloyd-Smith, and Christopher D. Hepburn. Comparing the Marine Protected Area Network Planning Process in British Columbia, Canada and New Zealand — Planning for cooperative partnerships with Indigenous communities. Marine Policy, 125(??):??, March 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2031037X.

Wilhelm:2020:PGH

[WKBS20] Miriam Wilhelm, Alin Kadfak, Vikram Bhakoo, and Kate Skattang. Private governance of human and labor rights in seafood supply chains — The case of the modern slavery crisis in Thailand. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304403.

Walton:2022:CGT

[WKH22] Grant W. Walton, Meg Keen, and Quentin Hanich. Can greater transparency improve the sustainability of Pacific fisheries? Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308976.

Wang:2025:NDC

[WKK^{+}25] Yixiao Wang, Sufyan Ullah Khan, Dequan Kong, Ruijing Xu, and Yao Zhang. Navigating the dual currents: Advanced insights into carbon emissions and economic growth in China’s coastal marine fisheries. Marine Policy, 177(??): ??, July 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25000867.

Warner:2025:SRM

[WKN^{+}25] Pedro Warner, Chris Kerry, Ana Nuno, Nathan A. Miller, Mollie Rickwood, and Kristian Metcalfe. A synthesis of research into marine small-scale fishers’ operational behaviour. Marine Policy, 179(??):??, September 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001551.

Wakita:2022:PSP

[WKO^{+}22] Kazumi Wakita, Hisashi Kurokura, Zaida A. Ochavo, Reyda I. Inolino, Hiroshi Fushimi, and Satoshi Ishikawa. Potential signals promoting behavior for coastal conservation: Conformity in small-scale fishing communities in the Philippines. Marine Policy, 146(??):??, December 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003396.

Wang:2020:AIA

[WL20] Leifan Wang and Fenghua Li. Alternative interpretation to Article VIII of the International Convention for the Regulation of Whaling: Implications from Article XX of the GATT. Marine Policy, 117(??):??, July 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19306463.

Wang:2022:SDC

[WL22a] Leifan Wang and Fenghua Li. State duty to cooperate on the Fukushima nuclear contaminated water release. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004899.

Wu:2022:STE

[WL22b] Jinghui Wu and Bo Li. Spatio-temporal evolutionary characteristics of carbon emissions and carbon sinks of marine industry in China and their time-dependent models. Marine Policy, 135(??):??, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004905.

Wu:2024:CES

[WL24a] Xu Wu and Xiaowei Lu. Can employment salvage payments be entitled to a maritime lien and allowed in the general average? A comparative analysis of the UK, US, and China. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002380.

Wu:2024:IIC

[WL24b] Xu Wu and Xiaowei Lu. Implementation of the international convention on the removal of wrecks: Legal obligations, model selections, and perspectives from China. Marine Policy, 170(??):??, December 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400366X.

Wilde:2023:ESD

[WLCC23] Daniel Wilde, Hannah Lily, Neil Craik, and Anindita Chakraborty. Equitable sharing of deep-sea mining benefits: More questions than answers. Marine Policy, 151 (??):??, May 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000994.

Williams:2023:SPC

[WLCG23] Lacey Williams, Serena Lucrezi, Paul Cowley, and Enrico Gennari. Stakeholders’ perceptions of the conservation and management of elasmobranchs in South Africa. Marine Policy, 157(??):??, November 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23003809.

Wan:2021:HDC

[WLQD21] Xiaole Wan, Qianqian Li, Lulian Qiu, and Yuanwei Du. How do carbon trading platform participation and government subsidy motivate blue carbon trading of marine ranching? A study based on evolutionary equilibrium strategy method. Marine Policy, 130(??):??, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001780.

Wilson:2020:STM

[WLRH^{+}20] Margaret W. Wilson, Julia M. Lawson, Maria Ignacia Rivera-Hechem, Juan Carlos Villaseñor-Derbez, and Steven D. Gaines. Status and trends of moored fish aggregating device (MFAD) fisheries in the Caribbean and Bermuda. Marine Policy, 121(??):??, November 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20304796.

Wang:2025:SPD

[WLS^{+}25] Bei Wang, Kaiyu Liu, Watchara Sakornwimon, Wenhao Huang, Tingting Li, Xianshui Lai, Chengcheng Li, Linlin Zhao, Bailin Cong, and Shenghao Liu. Spatial planning for dugong conservation: Assessing habitat suitability and conservation gaps in Indo–Pacific Convergence Zone. Marine Policy, 180(??):??, October 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001927.

Wu:2022:BMR

[WLW^{+}22] Haiping Wu, Qiuhui Li, Chunyan Wang, Qianwen Wu, Chongwei Peng, Thomas A. Jefferson, Zhizhen Long, Fangqiang Luo, Youhou Xu, and Shiang-Lin Huang. Bycatch mitigation requires livelihood solutions, not just fishing bans: a case study of the trammel-net fishery in the northern Beibu Gulf, China. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000653.

Wang:2021:PPE

[WLX21] Na Wang, Jing-Mei Li, and Zhi-Hua Xu. Public preference for the ecological restoration of coastal wetlands in Jiaozhou Bay in China based on a choice experiment. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000981.

Wu:2022:CET

[WLXY22] Min Wu, Kevin X. Li, Yi Xiao, and Kum Fai Yuen. Carbon Emission Trading Scheme in the shipping sector: Drivers, challenges, and impacts. Marine Policy, 138(??):??, April 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000367.

Wang:2021:FDP

[WLY21] Shuhong Wang, Binbin Lu, and Kedong Yin. Financial development, productivity, and high-quality development of the marine economy. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001640.

Wakamatsu:2021:ERS

[WM21] Hiroki Wakamatsu and Tsutom Miyata. Effects of radioactive safety information on consumer fears of radioactive contamination from oyster products in Japan. Marine Policy, 126(??):??, April 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X21000117.

Ward:2022:PUC

[WM22a] Sarah Ward and Mingfei Ma. Protection of the Underwater Cultural Heritage in Asia: Policy, law and practice. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002408.

Wuor:2022:DLF

[WM22b] Mehnwon Wuor and Leslie Mabon. Development of Liberia’s fisheries sectors: Current status and future needs. Marine Policy, 146(??):??, December 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003724.

Wiber:2021:RBE

[WMKG21] Melanie G. Wiber, Charles Mather, Christine Knott, and María Andrée López Gómez. Regulating the Blue Economy? Challenges to an effective Canadian aquaculture act. Marine Policy, 131(??):??, September 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003110.

Wawrzynkowski:2025:API

[WML25] Paul Wawrzynkowski, Climent Molins, and Josep Lloret. Assessing the potential impacts of floating Offshore Wind Farms on policy-relevant species: a case study in the Gulf of Roses, NW Mediterranean. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24005189.

Wang:2024:SUR

[WMM24] Nannan Wang, Wenjian Mu, and Rutai Ma. A systematic understanding of the risk development process for port authority. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002410.

White:2022:BRN

[WMRB22] Casey M. White, Sangeeta Mangubhai, Lukas Rumetna, and Cassandra M. Brooks. The bridging role of nongovernmental organizations in the planning, adoption, and management of the marine protected area network in Raja Ampat, Indonesia. Marine Policy, 141(??):??, July 2022.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001427.

Witter:2021:CSP

[WMS21] Allison Witter, Grant Murray, and U. Rashid Sumaila. Consumer seafood preferences related to alternative food networks and their value chains. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003055.

Widiastutie:2025:RAM

[WMS^{+}25] Sophiana Widiastutie, Dairatul Maarif, Dini Putri Saraswati, Imelda Masni Juniaty Sianipar, Thi Thanh Thuy Phan, and Van Viet Nguyen. Residents’ adaptive marine plastic litter management in Galang Island, Indonesia: an importanceperformance analysis. Marine Policy, 172(??):??, February 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24005086.

Wirajing:2025:ESF

[WN25] Muhamadu Awal Kindzeka Wirajing and Roger Tsafack Nanfosso. Establishing a sustainable fishing industry: the role of environmental stringency, fishery training and education infrastructure in the Maritime Cameroon. Marine Policy, 178(??):??, August 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X25001459.

Wootton:2024:BUO

[WNBHG24] Nina Wootton, Melissa Nursey-Bray, Shane Holland, and Bronwyn M. Gillanders. Better understanding ocean awareness: Insights from young people. Marine Policy, 164(??): ??, June 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2400157X.

Wootton:2022:PPP

[WNBRSG22] Nina Wootton, Melissa Nursey-Bray, Patrick Reis-Santos, and Bronwyn M. Gillanders. Perceptions of plastic pollution in a prominent fishery: Building strategies to inform management. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004577.

Wang:2023:BRM

[WNDZ23] Lei Wang, Wenzheng Niu, Lili Ding, and Zhongchao Zhao. A bibliometric review on marine economy: Current status, development and future directions. Marine Policy, 155(??): ??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002385.

Waldo:2025:HCW

[WNJ25] Jennifer L. Waldo, Mark D. Needham, and Megan S. Jones. How can we sea change? Audience subgroups and psychological cognitions to target in action-oriented ocean change communication. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005852.

Wakamatsu:2022:VWS

[WNM22] Mihoko Wakamatsu, Hiroki Nakamura, and Shunsuke Managi. The value of whaling and its spatial heterogeneity in Japan. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004632.

Willis-Norton:2024:SSS

[WNMS^{+}24] Ellen Willis-Norton, Tracey Mangin, Donna M. Schroeder, Reniel B. Cabral, and Steven D. Gaines. A synthesis of socioeconomic and sociocultural indicators for assessing the impacts of offshore renewable energy on fishery participants and fishing communities. Marine Policy, 161(??): ??, March 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000113.

Watterson:2020:ORE

[WOG20] Christopher J. Watterson, Stephen Osborne, and Samuel Grant. Open registries as an enabler of maritime sanctions evasion. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300385.

Weiss:2025:UMU

[WOR^{+}25] Carlos V. C. Weiss, Bárbara Ondiviela, Elvira Ramos, Lucía Meneses, Raúl Guanche, and José Juanes. Unlocking multiuse synergies: Spanish industry perspectives on offshore wind and aquaculture integration. Marine Policy, 180(??): ??, October 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X25002015.

Wang:2023:CCS

[WP23] Liang Wang and Haoxin Peng. Collaboration or competition? A science map for promoting the sustainable transition of port relationships. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23003287.

Whittaker:2024:OSY

[WPvO24] Geraint Rhys Whittaker, Kimberley Peters, and Ilse van Opzeeland. Oceans sing, are you listening? Sounding out potentials for artistic audio engagements with science through the Polar Sounds project. Marine Policy, 169(??): ??, November 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003452.

Ward-Paige:2022:FMM

[WPWM^{+}22] C. A. Ward-Paige, E. R. White, E. M. P. Madin, G. J. Osgood, L. K. Bailes, R. L. Bateman, E. Belonje, K. V. Burns, N. Cullain, P. Darbyshire-Jenkins, R. S. de Waegh, A. M. Eger, L. Fola-Matthews, B. M. Ford, C. Gonson, C. J. Honeyman, J. E. House, E. Jacobs, L. K. Jordan, J. J. Levenson, K. Lucchini, Marti-PuigP Martí-Puig, L. A. H. McGuire, C. Meneses, P. H. Montoya-Maya, R. A. Noonan, P. A. Ruiz-Ruiz, P. E. Ruy, R. A. Saputra, G. Shedrawi, B. Sing, M. D. Tietbohl, A. Twomey, D. C. Vergara Florez, and L. Yamb. A framework for mapping and monitoring human-ocean interactions in near real-time during COVID-19 and beyond. Marine Policy, 140(??):??, June 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001014.

Wang:2025:SIS

[WQ25] Jiajia Wang and Yong Qian. Book review: Small Island States and International Law: the Challenge of Rising Seas, Carolin König. Routledge Research in International Law. Routledge, Taylor & Francis, Francis, Abingdon, Oxon, and New York (2023), ISBN 978-1-032-20442-0. Marine Policy, 177(??):??, July 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25001083.

Wang:2023:AIG

[WRL23] Qiang Wang, Feng Ren, and Rongrong Li. Assessing the impact of geopolitics on international scientific cooperation — the case of us–china marine pollution research. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002567.

Watzel:2020:MMR

[WRV20] Ralph Watzel, Carsten Rühlemann, and Annemiek Vink. Mining mineral resources from the seabed: Opportunities and challenges. Marine Policy, 114:??, April 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300622.

Wakamatsu:2021:CJF

[WS21] Hiroki Wakamatsu and Yutaro Sakai. Can the Japanese fisheries qualify for MSC certification? Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0308597X21003614.

Williams:2023:FIP

[WS23] Chris Williams and Jessica L. Decker Sparks. Fishery improvement projects: a voluntary, corporate “tool” not fit for the purpose of mitigating labour abuses and guaranteeing labour rights for workers. Marine Policy, 147(??): ??, January 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003876.

Wang:2024:IBB

[WS24] Jinpeng Wang and Xiaohan Sun. Interplays between the BBNJ agreement and the Antarctic Treaty System on the Southern Ocean bioprospecting. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003361.

Wang:2021:ERA

[WSD21] Juying Wang, Xiao Shi, and Yuanwei Du. Exploring the relationship among marine science and technology innovation, marine finance, and marine higher education using a coupling analysis: a case study of China’s coastal areas. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003171.

Woker:2020:LSC

[WSDB20] Hilde Woker, Bernhard Schartmüller, Knut Ola Dølven, and Katalin Blix. The law of the sea and current practices of marine scientific research in the Arctic. Marine Policy, 115 (??):??, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19304130.

Wang:2025:ENE

[WSZ^{+}25] Chang Wang, Xiaojie Shu, Siyuan Zhou, Huiling Song, and Qi He. Embracing a new era of deep-sea mining: Research progress and prospects. Marine Policy, 180(??):??, October 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25001939.

Williamson:2021:CSI

[WtDD^{+}21] Michael J. Williamson, Mariel T. I. ten Doeschate, Rob Deaville, Andrew C. Brownlow, and Nicola L. Taylor. Cetaceans as sentinels for informing climate change policy in UK waters. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002451.

Wiranthi:2024:FTB

[WTO24] Puspi Eko Wiranthi, Hilde M. Toonen, and Peter Oosterveer. Flexibility through bundles of capital: the capabilities of Indonesian small-scale handline tuna fishers under voluntary sustainability certification. Marine Policy, 168(??): ??, October 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2400294X.

Wu:2022:STR

[Wu22] Hsing-Hao Wu. A study on transnational regulatory governance for marine plastic debris: Trends, challenges, and prospect. Marine Policy, 136(??):??, February 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301822.

Wagner:2021:SGN

[WvdMG^{+}21] Daniel Wagner, Liesbeth van der Meer, Matthias Gorny, Javier Sellanes, Carlos F. Gaymer, Eulogio H. Soto, Erin E. Easton, Alan M. Friedlander, Dhugal J. Lindsay, Tina N. Molodtsova, Ben Boteler, Carole Durussel, Kristina M. Gjerde, Duncan Currie, Matthew Gianni, Cassandra M. Brooks, Marianne J. Shiple, T. ‘Aulani Wilhelm, Marco Quesada, Tamara Thomas, Piers K. Dunstan, Nichola A. Clark, Luis A. Villanueva, Richard L. Pyle, Malcolm R. Clark, Samuel E. Georgian, and Lance E. Morgan. The Salas y Gómez and Nazca ridges: a review of the importance, opportunities and challenges for protecting a global diversity hotspot on the high seas. Marine Policy, 126(??): ??, April 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20310289.

Walters:2021:DEA

[WW21a] David Walters and Emma Wadsworth. Determinants of effective action on workplace safety and health in global companies — The case of global network container terminal operators. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310253.

Wang:2021:MFP

[WW21b] Sun Ling Wang and John B. Walden. Measuring fishery productivity growth in the Northeastern United States 2007–2018. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000786.

Wang:2022:CEE

[WW22] Qiang Wang and Shasha Wang. Carbon emission and economic output of China’s marine fishery — a decoupling efforts analysis. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004425.

Wang:2025:PBA

[WW25] Chuanliang Wang and Min Wu. On the provisions of the BBNJ Agreement addressing the impacts of climate change. Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004299.

Wu:2023:ALH

[WWC23] Ling-I Wu, Chia-Hui Wang, and Chih-Shin Chen. Applying life history traits to eco-labelling schemes for fisheries management: a case study in Taiwan. Marine Policy, 148(??): ??, February 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22004596.

Wan:2021:SSR

[WWCS21] Zheng Wan, Likun Wang, Jihong Chen, and Daniel Sperling. Ship scrappage records reveal disturbing environmental injustice. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001536.

Waters:2023:PAD

[WWD23] Yolanda L. Waters, Kerrie A. Wilson, and Angela J. Dean. Plastic action or distraction? Marine plastic campaigns influence public engagement with climate change in both general and engaged audiences. Marine Policy, 152(??):??, June 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23001070.

Wally:2024:CRC

[WWD24] Kasen Wally, Ryan Williamson, and Kelly Dunning. Coral reefs and climate change: Examining two institutional approaches to managing a novel marine ecosystem. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001052.

Ward:2023:LSR

[WWG^{+}23] Tim M. Ward, Barrett W. Wolfe, Gretchen L. Grammer, Alex R. Ivey, Edward King, Andreas Schiller, Karlie S. McDonald, and Jeffrey M. Dambacher. Large sardine resource discovered off south-eastern Australia: Potential risks, challenges and benefits of establishing a new fishery. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002725. See corrigen-

dum [WWG^{+}25].

Ward:2025:CLS

[WWG^{+}25] Tim M. Ward, Barrett W. Wolfe, Gretchen L. Grammer, Alex R. Ivey, Edward King, Andreas Schiller, Karlie S. McDonald, and Jeffrey M. Dambacher. Corrigendum to Large sardine resource discovered off southeastern Australia: Potential risks, challenges and benefits of establishing a new fishery [Mar. Policy 155 (2023) 105739]. Marine Policy, 175(??):??, May 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000697. See [WWG^{+}23].

Wolters:2025:SAO

[WWS^{+}25] Erika Allen Wolters, Melissa Ward, Ana K. Spalding, Arielle Levine, Kristen Marie Green, Ryan Hasert, Charlotte M. Regula-Whitefield, and Kevin M. Johnson. Shellfisheries’ adaptation to ocean acidification in OR and CA: Linking strategies to strategic policy action. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24004846.

Wan:2021:EPT

[WXLD21] Xiaole Wan, Shuwen Xiao, Qianqian Li, and Yuanwei Du. Evolutionary policy of trading of blue carbon produced by marine ranching with media participation and government supervision. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309477.

Wang:2022:HDE

[WYD^{+}22] Han Wang, YanChun Yao, XiaoLing Dai, ZhouPeng Chen, Jianxian Wu, GuangLong Qiu, and Ting Feng. How do ecological protection policies affect the restriction of coastal development rights? Analysis of choice preference based on choice experiment. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21005169.

Walther:2021:SAP

[WYH21] Bruno Andreas Walther, Ning Yen, and Chieh-Shen Hu. Strategies, actions, and policies by Taiwan’s ENGOs, media, and government to reduce plastic use and marine plastic pollution. Marine Policy, 126(??):??, April 2021.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000014.

Wang:2025:LIS

[WZ25] Chuanliang Wang and Qian Zhao. On the legal issues of the Shiprider Program. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24004366.

Wang:2025:FNW

[WZL25] Chao Wang, Yanting Zhao, and Nicholas Lassi. Fukushima nuclear wastewater discharge: an analysis of the international regulatory framework. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2400441X.

Wang:2024:CLP

[WZX24] Qiuwen Wang, Hu Zhang, and Shutong Xi. China’s law and policy framework for maritime safety regulation of alternative fuel ships in the decarbonization transition. Marine Policy, 163(??):??, May 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001404.

Wan:2021:ISE

[WZZC21] Zheng Wan, Xujun Zhou, Qiang Zhang, and Jihong Chen. Influence of sulfur emission control areas on particulate matter emission: a difference-in-differences analysis. Marine Policy, 130(??):??, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001950.

Xu:2022:REE

[XHY22] Jie Xu, Limin Han, and Wei Yin. Research on the ecologicalization efficiency of mariculture industry in China and its influencing factors. Marine Policy, 137(??):??, March 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21005467.

Xu:2021:RCM

[XHZ^{+}21] Xin Xu, Yingying Hou, Changping Zhao, Lei Shi, and Yu Gong. Research on cooperation mechanism of marine plastic waste management based on complex network evolutionary game. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003857.

Xiang:2025:ICM

[XL25] Linhan Xiang and Tiansheng Li. Institutional consequences and mechanism adjustments to China’s international ship registration policy — taking the Red Sea crisis as an example. Marine Policy, 171(??):??, January 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004573.

Xu:2021:ECR

[XLLS21] Zhihua Xu, Shuqin Li, Jingmei Li, and Jingzhu Shan. The effect of consequentiality on the residents’ willingness to pay for the governance of Ulva prolifera bloom in Qingdao, China. Marine Policy, 126(??):??, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000142.

Xuan:2024:IIP

[XLW24] Huihui Xuan, Qing Liu, and Lei Wang. The impact of incentive policies on shipowners’ adoption behavior of clean energy technologies: Evidence from China. Marine Policy, 167(??):??, September 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24002756.

Xu:2021:EEC

[XSCL21] Lang Xu, Jia Shi, Jihong Chen, and Li Li. Estimating the effect of COVID-19 epidemic on shipping trade: an empirical analysis using panel data. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003791.

Xu:2023:BMC

[XT23] Qi Xu and Ziyue Tan. Building a maritime community with a shared future: Scholarly reflections on China’s new ocean vision. Marine Policy, 149(??):??, March 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000350.

Xixi:2022:OSS

[XTY22] Zhang Xixi, Qu Tongkun, and Wang Yecheng. Optimal strategies for stakeholders of Fukushima nuclear waste water discharge in Japan. Marine Policy, 135(??):??, January 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21004929.

Xu:2022:MCR

[XWH^{+}22] Nan Xu, Yuqing Wang, Conghong Huang, Shuai Jiang, Mingming Jia, and Yue Ma. Monitoring coastal reclamation changes across Jiangsu Province during 1984–2019 using Landsat data. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100498X.

Xiao:2020:ENI

[XWL^{+}20] Yi Xiao, Grace Wang, Kun-Chin Lin, Guanqiu Qi, and Kevin X. Li. The effectiveness of the New Inspection Regime for Port State Control: Application of the Tokyo MoU. Marine Policy, 115(??):??, May 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19302659.

Xu:2024:IFR

[XXLW24] Zhihua Xu, Juan Xu, Shuqin Li, and Chuwei Wang. The influencing factors of residents’ willingness to pay in marine ecological restoration: the integration of the theory of planned behavior and social capital theory.

Marine Policy, 161(??):??, March 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X24000290.

Xu:2023:URP

[XXMZ23] Zhihua Xu, Juan Xu, Wenchao Meng, and Gang Zhou. Understanding residents’ policy preferences to construct effective marine debris governance policies. Marine Policy, 158(??):??, December 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23004050.

Xing:2023:ELG

[XZ23] Wangwang Xing and Ling Zhu. Exploring legal gaps and barriers to the use of unmanned merchant ships in China. Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001896.

Yadav:2024:DSI

[Yad24] Dhanwati Yadav. Deciphering Sino–Indian blue mining competencies in the Indo–Pacific. Marine Policy, 170(??): ??, December 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003762.

Yuen:2020:SBS

[YBW20] Kum Fai Yuen, Xiwen Bai, and Xueqin Wang. Safety behaviour at sea: Policy implications for managing seafarers through positive psychology. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20306953.

Yuan:2021:MCO

[YC21a] Wei Yuan and Yen-Chiang Chang. Maritime cooperation and ocean governance: Conference report. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20310393.

Yunitawati:2021:GES

[YC21b] Dinah Yunitawati and Julian Clifton. Governance in the early stages of marine protected area development: a case study of Nusa Penida District Marine Conservation Area, Indonesia. Marine Policy, 127(??):??, May 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19305172.

Yuen:2021:PCS

[YCBW21] Kum Fai Yuen, Yangyi Cao, Xiwen Bai, and Xueqin Wang. The psychology of cruise service usage post COVID-19: Health management and policy implications. Marine Policy, 130(??):??, August 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21001974.

Yuan:2024:CJL

[YCF24] He Yuan, Leïla Choukroune, and Pierre Failler. Centring justice for labour in the new blue economy: Principles for applying emerging evidence and theoretical critiques to policy and practice. Marine Policy, 168(??):??, October 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003257.

Yang:2022:SSF

[YCL22a] Tsung-Yu Yang, Tsun-Feng Chiang, and Wen-Hong Liu. Small-scale fishers’ catch production in Taiwanese coastal areas. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002299.

Yu:2022:EGR

[YCL22b] Derek Chin-Yun Yu, Wen-Yan Chiau, and Hsueh-Jung Lu. Effective governance of a remote marine protected area: the case of Dongsha Atoll National Park. Marine Policy, 139 (??):??, May 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22000604.

Yang:2025:RDB

[YCL25] Weiming Yang, Xidi Chen, and Yu Liu. Recent developments in building sustainable marine fisheries in China: Reflections on the 2023 Revision of the Marine Environmental Protection Law. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004391.

Yu:2022:LPM

[YD22] Xiao Yu and Yichong Dong. Local practice of marine protected areas legislation in China: the case of Zhoushan. Marine Policy, 141(??):??, July 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001312.

Yadav:2021:SFC

[YFD^{+}21] Shreya Yadav, Abdulla Fisam, Rachel Dacks, Joshua S. Madin, and Alexander Mawyer. Shifting fish consumption preferences can impact coral reef resilience in the Maldives: a case study. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21003845.

Yadav:2020:OCC

[YG20] Siddharth Shekhar Yadav and Kristina Maria Gjerde. The ocean, climate change and resilience: Making ocean areas beyond national jurisdiction more resilient to climate change and other anthropogenic activities. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X20302074.

Yu:2020:FSM

[YH20] Jinkai Yu and Qingchao Han. Food security of mariculture in China: Evolution, future potential and policy. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X1930675X.

Yang:2022:SAF

[YHU22] Xue Yang, Ingunn Marie Holmen, and Ingrid Bouwer Utne. Scenario analysis of fish escapes in Norwegian aquaculture for implementation of barrier management and improved learning from accidents. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200255X.

Ye:2024:STG

[YHX24] Zhou Ye, Xisheng Huang, and Zhuojun Xie. Sustainable transformation of global ocean governance: Can a global public goods approach lead the way? Marine Policy, 165 (??):??, July 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24002112.

Yang:2024:AED

[YJG24] Nina Yang, Di Jin, and Annette F. Govindarajan. Applying environmental DNA approaches to inform marine biodiversity conservation: the case of the Ocean Twilight Zone. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001490.

Yi:2020:POA

[YK20] Sangchoul Yi and Wonkook Kim. Potential for offshore aquaculture development in North Korea: Focusing on Atlantic salmon farming. Marine Policy, 119(??):??, September 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20302323.

Young:2021:NSA

[YK21] Oran R. Young and Jong-Deog Kim. Next steps in Arctic Ocean Governance Meeting the challenge of coordinating a dynamic regime complex. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003377.

Yen:2021:RML

[YL21] Kuo-Wei Yen and Li-Chuan Liuhuang. A review of migrant labour rights protection in distant water fishing in Taiwan: From laissez-faire to regulation and challenges behind. Marine Policy, 134(??):??, December 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004164.

Yu:2023:TTS

[YLSC23] Xuan Yu, Sen Lin, Weiteng Shen, and Qi Chen. The tiered transaction system for sea use rights in China: Theoretical basis and market construction. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X2200450X.

Yanda:2023:CCL

[YMG23] Pius Z. Yanda, Anselm R. Mwajombe, and Brown Gwambene. Coastal communities’ livelihoods conditioned by resource use dynamics and changing environments in Lindi Region, South-Eastern Tanzania. Marine Policy, 150 (??):??, April 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000593.

Yu:2021:LMP

[YML^{+}21] Lance Yu, Linna Ma, Vicky Lam, Xiaomu Guan, Yujie Zhao, Sheng Wang, Yongtong Mu, and Rashid Sumaila. Local marine policy whacking the national Zhikong scallop fishery. Marine Policy, 124(??):??, February 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20310022.

Yosef:2022:API

[YNZ22] Reuven Yosef, Tal Nachshonov, and Piotr Zduniak. Anthropopause positively influenced Red Sea Clownfish (Amphiprion bicinctus) populations but not the host sea anemone (Actiniaria spp.) in Eilat, Israel. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2200327X.

Yea:2021:CVC

[YS21] Sallie Yea and Christina Stringer. Caught in a vicious cycle: Connecting forced labour and environmental exploitation through a case study of Asia–Pacific. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X2100436X.

Yea:2023:IPW

[YS23] Sallie Yea and Christina Stringer. The informalisation of precarious work in fishing crew: Experiences of Fijian fishers on distant water vessels. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002427.

Yosef:2022:ABA

[YSEK22] Reuven Yosef, Veronica Spivak, Shalev Edelman, and Jakub Z. Kosicki. Anthropogenic beach-activities influences the endemic Red Sea Ghost Crab (Ocypode saratan). Marine Policy, 138(??):??, April 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000331.

Yuan:2025:IOM

[YSL^{+}25] Bin Yuan, Wen Sun, Wei Liu, Jingyi Yang, and Jiawei Zhang. Impact and optimization of marine ecological protection compensation mechanisms on fishermen’s income. Marine Policy, 181(??):??, November 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25002374.

Yi:2021:CED

[YSWW21] Xiaojing Yi, Kun Sheng, Yuanyue Wang, and Shuhong Wang. Can economic development alleviate storm surge disaster losses in coastal areas of China? Marine Policy, 129 (??):??, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001421.

Yin:2020:CAM

[YT20] Mengmeng Yin and Erika J. Techera. A critical analysis of marine protected area legislation across state and territory jurisdictions in Australia. Marine Policy, 118(??): ??, August 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19302581.

Yandle:2022:AVC

[YTBM^{+}22] Tracy Yandle, Jennifer Sweeney Tookes, Paulita Bennett-Martin, Sherry Larkin, and Michael Page. An assessment of the viability of a commercial lionfish fishery in the US Virgin Islands. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22001506.

Yildirim:2022:CEW

[YTÇ22] Umut Yildirim, Arda Toygar, and Cavit Çolakoğlu. Compensation effect of wages on decent work: a study on seafarers attitudes. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22002020.

Yu:2023:VCD

[YTG^{+}23] Lance Yu, Xiaolan Tan, Xiaomu Guan, Yongtong Mu, Vicky Lam, and Rashid Sumaila. Value chain of the datapoor Chinese bay scallop aquaculture. Marine Policy, 150 (??):??, April 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23000830.

Yildiz:2020:AGP

[YU20] Taner Yildiz and Aylin Ulman. Analyzing gaps in policy: Evaluation of the effectiveness of minimum landing size (MLS) regulations in Turkey. Marine Policy, 115(??): ??, May 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19301587.

Yu:2022:OOD

[Yu22] Chuxiao Yu. Operational oceanography as a distinct activity from marine scientific research under UNCLOS?

— An analysis of WMO Resolution 45 (Cg-18). Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001786.

Yu:2020:EGO

[YW20] Jin-Kai Yu and Yan-Ling Wang. Exploring the goals and objectives of policies for marine ranching management: Performance and prospects for China. Marine Policy, 122(??): ??, December 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X20309015.

Yin:2024:AJL

[YW24] Miaomiao Yin and Sen Wang. Australia’s jurisdiction and law enforcement in combating IUU fishing in the Southern Ocean. Marine Policy, 169(??):??, November 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003658.

Yang:2020:MBS

[YWY20] Lijing Yang, LingLing Wu, and Shuqing Yan. Marketbased sea use management in China: Features and countermeasures. Marine Policy, 120(??):??, October 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309182.

Yu:2022:PNA

[YWYL22] Huijuan Yu, Yezhi Wang, Wei Yin, and Dahai Li. Policy network analysis of China’s ocean ranching policy: Network structure, actors and interaction. Marine Policy, 140(??): ??, June 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001178.

Yuan:2024:AIG

[YXDL24] Yingyi Yuan, Tianyuan Xie, Xuxuan Dong, and Jiaguo Liu. Analysis of the influence of Japan’s nuclear waste water release on South Korean fishery people from the perspective of news reports. Marine Policy, 163(??):??, May 2024. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001234.

Yuan:2024:APD

[YXX^{+}24] Yuan Yuan, Gangchun Xu, Pao Xu, Hongxia Li, Lin Zhang, Zhijuan Nie, Pierre Murekezi, Xiao Liang, Zhengyong Yang, and Xinhua Yuan. Aquaculture PPP development in China — case study from Hani Terrace. Marine Policy, 163(??): ??, May 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24000733.

Yu:2023:CEH

[YYD23] Wencheng Yu, Miao Yu, and Lili Ding. Characteristics and evolution of hierarchical fishery policies in China — a textual analysis based on 5311 policies from 2003 to 2022. Marine Policy, 155(??):??, September 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002269.

Yu:2020:EMP

[YYL20] Jinkai Yu, Wang Yin, and Dahai Liu. Evolution of mariculture policies in China: Experience and challenge. Marine Policy, 119(??):??, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19304798.

Yu:2025:ESP

[YYM25] Jinkai Yu, Tingjun Yan, and Xingyun Ma. Evolution of support policies for the development of distant-water fishing bases in China: Review, performance, challenges and prospects. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004895.

Yu:2023:GER

[YZG23] Lianghong Yu, Shan Zheng, and Qiang Gao. Government environmental regulation strategy for new pollutants control in mariculture. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23000726.

Yin:2022:ISM

[YZH22] Kedong Yin, Kai Zhang, and Chong Huang. Institutional supply, market cultivation, and the development of marine strategic emerging industries. Marine Policy, 139(??):??, May 2022. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X22000926.

Yu:2024:TDA

[YZL24] Wencheng Yu, Benshuo Zhu, and Kangkang Lin. Threedimensional analysis of China’s fisheries insurance policy: a tool-direction-subject perspective. Marine Policy, 168(??): ??, October 2024. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24003282.

Zurel:2025:AEY

[ZABI25] Dror Zurel, Fred Arzoine, and Zohar Barnett-Itzhaki. Assessing the effectiveness of 18 years of Israel’s “Clean Coast” program. Marine Policy, 176(??):??, June 2025. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X25000703.

Zuercher:2022:ECE

[ZBF^{+}22] Rachel Zuercher, Natalie C. Ban, Wesley Flannery, Anne D. Guerry, Benjamin S. Halpern, Rafael Almeida Magris, Shauna L. Mahajan, Nicole Motzer, Ana K. Spalding, Vanessa Stelzenmüller, and Jonathan G. Kramer. Enabling conditions for effective marine spatial planning. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22001889.

Zimmerhackel:2023:FIA

[ZCA^{+}23] J. S. Zimmerhackel, J. Clifton, F. Ackermann, M. P. Burton, C. E. Elrick-Barr, G. Hill, and E. S. Harvey. A framework for the integrated assessment of social and economic values associated with man-made marine structures. Marine Policy, 152(??):??, June 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001392.

Zumpano:2024:IBC

[ZCC^{+}24] Francisco Zumpano, Sofía Copello, Victoria Cabral, Marco Favero, and Germán García. Interaction between coastal birds and recreational fisheries in Argentine: Incorporating the social dimension to improve effectiveness of conservation and management measures in a key estuarine reserve. Marine Policy, 167(??):??, September 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24002768.

Zhao:2022:SCR

[ZCCQ22a] Qilei Zhao, Xinjun Chen, Guiying Chen, and Siqiong Qi. The socioeconomic contribution of the recreational fishery based on input–output analysis: the case of China. Marine Policy, 143(??):??, September 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2200224X.

Zhao:2022:UDR

[ZCCQ22b] Qilei Zhao, Xinjun Chen, Guiying Chen, and Siqiong Qi. Understanding the development of recreational fisheries with an industry competitiveness framework: Evaluation of a case study from China. Marine Policy, 143(??):??, September 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22001683.

Zhao:2023:MCE

[ZCG^{+}23] Qiqi Zhao, Yan Chen, Keshava Pallavi Gone, Emily Wells, Keahna Margeson, and Kate Sherren. Modelling cultural ecosystem services in agricultural dykelands and tidal wetlands to inform coastal infrastructure decisions: a social media data approach. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300060X.

Zhou:2023:ICC

[ZCS^{+}23] Jingen Zhou, Shu-Ling Peggy Chen, Wenming Wendy Shi, Maneerat Kanrak, and Jiawei Ge. The impacts of COVID-19 on the cruise industry based on an empirical study in China. Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001586.

Zhou:2023:CSC

[ZCSK23] Jingen Zhou, Shu-Ling (Peggy) Chen, Wenming (Wendy) Shi, and Maneerat Kanrak. Cruise supply chain risk mitigation strategies: an empirical study in Shanghai, China. Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001276.

Zhang:2020:OCS

[ZCZ20] Weibin Zhang, Yen-Chiang Chang, and Liangfu Zhang. An ocean community with a shared future: Conference report. Marine Policy, 116(??):??, June 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20301081.

Zhang:2020:FSE

[ZD20a] Pengfei Zhang and Lukas Drumm. The flagging-out strategy: an examination of the impacts on the decreasing German national fleet. Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X19309832.

Zhang:2020:GSF

[ZD20b] Pengfei Zhang and Lukas Drumm. The German Shipping Foundation: Has it been effective in maintaining maritime expertise in Germany? Marine Policy, 115(??):??, May 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20300117.

Zhao:2020:EFB

[ZFGL20] Changping Zhao, Chao Fang, Yu Gong, and Zudi Lu. The economic feasibility of Blue Carbon cooperation in the South China Sea region. Marine Policy, 113(??):??, March 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19304725.

Zaucha:2025:CIE

[ZGR^{+}25a] Jacek Zaucha, Kira Gee, Emiliano Ramieri, Leila Neimane, Neil Alloncle, Nerijus Blažauskas, Helena Calado, Cristina Cervera-Núñez, Vesna Marohnić Kuzmanović, Margarita Stancheva, Joanna Witkowska, Sigrid Eskeland Schütz, Juan Ronco Zapatero, and Charles N. Ehler. Corrigendum to “Implementing the EU MSP Directive: Current status and lessons learned in 22 EU Member States” [Mar. Policy 171 (2025) 106425]. Marine Policy, 173(??):??, March 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24005359. See [ZGR^{+}25b].

Zaucha:2025:IEM

[ZGR^{+}25b] Jacek Zaucha, Kira Gee, Emiliano Ramieri, Leila Neimane, Neil Alloncle, Nerijus Blažauskas, Helena Calado, Cristina Cervera-Núñez, Vesna Marohnić Kuzmanović, Margarita Stancheva, Joanna Witkowska, Sigrid Eskeland Schütz, Juan Ronco Zapatero, and Charles N. Ehler. Implementing the EU MSP Directive: Current status and lessons learned in 22 EU Member States. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X24004251. See corrigen-

dum [ZGR^{+}25a].

Zhou:2021:MFR

[ZGY21] Qin Zhou, Ruobin Gao, and Kum Fai Yuen. Marine fuel refining technology improvement trade-offs: a game theoretic approach. Marine Policy, 132(??):??, October 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21002888.

Zhang:2021:UTC

[Zha21] Hongzhou Zhang. Understanding the tragedy of the commons in the South China Sea fisheries: the prisoner’s dilemma model revisited. Marine Policy, 125(??):??, March 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20310277.

Zhao:2022:ICN

[Zha22] Na Zhao. The impact of China’s new-type urbanization on individual sustainable development: a comparison between coastal and inland areas. Marine Policy, 136(??):??, February 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21005492.

Zhang:2025:LEU

[Zha25] Shi’ao Zhang. Law enforcement or use of force: the legal nature of activities conducted by CCG and JCG in the disputed waters. Marine Policy, 171(??):??, January 2025. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24004408.

Zhao:2023:RCN

[ZHL^{+}23] Changping Zhao, Yingying Hou, Mengru Liu, Yu Gong, and Junbin Wang. Research on the cooperative network game model of marine plastic waste management. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23000313.

Zhou:2023:PVE

[Zho23] Hang Zhou. Provincial variations and entrepreneurialism in the development of China’s Distant Water Fisheries (2011–2020). Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003918.

Zhu:2022:QBS

[Zhu22] Ling Zhu. A questionnaire based study on the status quo and future action over the regulation of marine greenhouse gas emissions in Hong Kong. Marine Policy, 145(??):??, November 2022. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X22003256.

Zhu:2023:STA

[Zhu23] Ling Zhu. Some thoughts on application of the polluter pays principle for controlling marine greenhouse gas emissions. Marine Policy, 158(??):??, December 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23004104.

Zhao:2020:TSS

[ZJ20] Xin Zhao and Peihong Jia. Towards sustainable smallscale fisheries in China: a case study of Hainan. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19305998.

Zaucha:2022:EMS

[ZJ22] Jacek Zaucha and Stephen Jay. The extension of marine spatial planning to the management of the world ocean, especially areas beyond national jurisdiction. Marine Policy, 144(??):??, October 2022. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X22002652.

Zaucha:2021:ESM

[ZK21] Jacek Zaucha and Anja Kreiner. Engagement of stakeholders in the marine/maritime spatial planning process. Marine Policy, 132(??):??, October 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X18304081.

Zhang:2020:FGS

[ZL20] Hongzhou Zhang and Mingjiang Li. Fisheries governance in the south China sea: Problems, progress, and prospects. Marine Policy, 121(??):??, November 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20308836.

Zhang:2024:FBA

[ZL24] Jiefang Zhang and Huirong Liu. Feasibility of the BBNJ Agreement to regulate bioprospecting in the Southern Ocean. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2400201X.

Zhu:2021:TTA

[ZLD^{+}21] Wenbin Zhu, Zhanhui Lu, Qian Dai, Kaner Lu, Zhenhua Li, Yongdong Zhou, Yazhou Zhang, Ming Sun, Yunzhou Li, and Wei Li. Transition to timely and accurate reporting: an evaluation of monitoring programs for China’s first Total Allowable Catch (TAC) pilot fishery. Marine Policy, 129 (??):??, July 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21001147.

Zejnilovic:2024:ECR

[ZLG^{+}24] Leid Zejnilović, Susana Lavado, Catarina Grilo, Rita Sá, and Sérgio Faias. The effects of COVID-19 restrictions on Portuguese marine small-scale fisheries: a regional analysis. Marine Policy, 165(??):??, July 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24001544.

Zhao:2020:SAQ

[ZLGX20] Changping Zhao, Rui Li, Yu Gong, and Xin Xu. Study on the asymmetric QRE network game simulation of the South China Sea route trade cooperation. Marine Policy, 111(??): ??, January 2020. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X19300302.

Zhu:2021:SAR

[ZLH21] Wenchao Zhu, Bo Li, and Zenglin Han. Synergistic analysis of the resilience and efficiency of China’s marine economy and the role of resilience policy. Marine Policy, 132(??): ??, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X21003146.

Zhang:2023:MSP

[ZLPT23] Yiqi Zhang, Christine Loh, Gerald R. Patchell, and Kellee S. Tsai. Multi-scale policy diffusion of marine emissions governance. Marine Policy, 153(??):??, July 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001641.

Zhou:2022:ICS

[ZLWY22] Yusheng Zhou, Xue Li, Xueqin Wang, and Kum Fai Yuen. Intelligent container shipping sustainability disclosure via stakeholder sentiment views on social media. Marine Policy, 135(??):??, January 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004644.

Zhou:2023:SSC

[ZLY23] Yusheng Zhou, Xue Li, and Kum Fai Yuen. Sustainable shipping: a critical review for a unified framework and future research agenda. Marine Policy, 148(??):??, February 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000052.

Zheng:2021:HDF

[ZLZ21] Hui Zheng, Jie Li, and Xin Zhao. How does financial policy support the development of China’s fishery? Characteristics, experience and prospects. Marine Policy, 132(??): ??, October 2021. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X2100289X.

Zudaire:2023:BDF

[ZMM^{+}23] Iker Zudaire, Gala Moreno, Jefferson Murua, Paul Hamer, Hilario Murua, Mariana T. Tolotti, Marlon Roman, Martin Hall, Jon Lopez, Maitane Grande, Gorka Merino, Lauriane Escalle, Oihane C. Basurko, Manuela Capello, Laurent Dagorn, Maria Lourdes Ramos, Francisco J. Abascal, José Carlos Báez, Pedro J. Pascual-Alayón, Santiago Déniz, and Josu Santiago. Biodegradable drifting fish aggregating devices: Current status and future prospects. Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001860.

Zambrano-Monserrate:2020:EDC

[ZMR20] Manuel A. Zambrano-Monserrate and Maria Alejandra Ruano. Estimating the damage cost of plastic waste in Galapagos Islands: a contingent valuation approach. Marine Policy, 117(??):??, July 2020. CODEN ???? ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X19309820.

Zhang:2025:ASF

[ZMT^{+}25] Lei Zhang, Shengwei Ma, Yi Tang, Yanbo Zhou, and Qiaer Wu. Acceptance of the summer fishing moratorium among small-scale fishers in China. Marine Policy, 173(??): ??, March 2025. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X24005566.

Zhang:2023:CMS

[ZMZ23] Shiyang Zhang, Zhengkai Mao, and Zhijun Zhang. China’s maritime strategy in the new vision. Marine Policy, 155(??):??, September 2023. CODEN ????

ISSN 0308-597X (print), 1872-9460 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0308597X23003020.

Zhu:2023:SEI

[ZNA^{+}23] Shengda Zhu, Adolf K. Y. Ng, Mawuli Afenyo, Roozbeh Panahi, and Michael G. H. Bell. Socio-economic impacts of shipping along the Northwest Passage: the cost to locals. Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001744.

Zou:2023:CIR

[Zou23] Yizheng Zou. China and Indonesia’s responses to maritime disputes in the South China Sea: forming a tacit understanding on security. Marine Policy, 149(??):??, March 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000295.

Zhao:2022:EME

[ZPX^{+}22] Yiyi Zhao, Ellen K. Pikitch, Xiangxin Xu, Timothy Frankstone, John Bohorquez, Xu Fang, Jie Zheng, Yunyi Li, Zhu Chen, Wuying Lin, Wenjia Hu, and Guifang (Julia) Xue. An evaluation of management effectiveness of China’s marine protected areas and implications of the 2018 Reform. Marine Policy, 139(??):??, May 2022. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22000872.

Zafimahatradraibe:2025:ACP

[ZRR^{+}25] Jean Aimé Zafimahatradraibe, Lala N. J. Ranaivomanana, Cicelin Rakotomahazo, Bemahafaly Randriamanantsoa, Aaron C. Hartmann, and Gildas G. B. Todinanahary. Authority, capacity, and power to govern: Three marine protected areas co-managed by resource users and nongovernmental organizations. Marine Policy, 177(??):??, July 2025. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X25000624.

Zhou:2020:KCC

[ZSLY20] Yusheng Zhou, Ying Shan Soh, Hui Shan Loh, and Kum Fai Yuen. The key challenges and critical success factors of blockchain implementation: Policy implications for Singapore’s maritime industry. Marine Policy, 122(??):??, December 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309118.

Zuidema:2023:PBS

[ZSWJ23] Jason Zuidema, Desai Shan, Kevin Walker, and Alireza Jahanbakhsh. Port-based seafarers’ welfare in Canada: resources and gaps. Marine Policy, 147(??):??, January 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X22003839.

Zheng:2020:CPF

[ZSZ20] Hui Zheng, Min Shang, and Xin Zhao. Chinese policy on fishery insurance: Evolution, characteristics and challenges. Marine Policy, 119(??):??, September 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20300026.

Zheng:2021:COT

[ZT21] Li Zheng and Kailan Tian. The contribution of ocean trade to national economic growth: a non-competitive inputoutput analysis in China. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001706.

Zhao:2023:ICM

[ZT23a] Zhiwei Zhao and Lijun Tang. The impact of COVID-19 on maritime pilots: Evidence and lessons. Marine Policy, 153 (??):??, July 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23001914.

Zou:2023:UCD

[ZT23b] Keyuan Zou and Anastasia Telesetsky. UNCLOS and its contributions to the development of international law: an editorial note. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002361.

Zhu:2023:VAT

[ZTA23] Mengyao Zhu, Kotaro Tanaka, and Tomonari Akamatsu. Visualizing the annual transition of ocean policy in Japan using text mining. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23002877.

Zhao:2023:FSD

[ZTM^{+}23] Zhiwei Zhao, Lijun Tang, Yunlei Ma, Yueyan Wu, Shiqi Lin, Zefan Wu, Zheng Zhou, Xinyi Wang, and Xinjian Wang. Fatigue at sea during and after the COVID-19 pandemic: a comparative study of two matched samples of seafarers. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23002634.

Zhao:2023:FDC

[ZTW23] Zhiwei Zhao, Lijun Tang, and Yueyan Wu. Fatigue during the COVID-19 pandemic: the experiences of Chinese seafarers. Marine Policy, 153(??):??, July 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X23001707.

Zhang:2021:ICP

[ZTZ^{+}21] Yanxuedan Zhang, Yi Tang, Yu Zhang, Yuqiong Sun, and Haoran Yang. Impacts of the COVID-19 pandemic on fish trade and the coping strategies: an initial assessment from China’s perspective. Marine Policy, 133(??):??, November 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21003596.

Zhang:2020:CPB

[ZV20] Xiong Zhang and Amanda C. J. Vincent. China’s policies on bottom trawl fisheries over seven decades (1949– 2018). Marine Policy, 122(??):??, December 2020. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309027.

Zhang:2021:IMI

[ZW21] Ying Zhang and Shuhong Wang. Influence of marine industrial agglomeration and environmental regulation on marine innovation efficiency — From an innovation value chain perspective. Marine Policy, 134(??):??, December 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21004188.

Zheng:2023:PEA

[ZWC^{+}23] Hui Zheng, Chong Wang, Jingya Chen, Xuan Liu, and Xin Zhao. Policy evolution and analysis of characteristics and values of China’s marine seed industry development under the food security strategy. Marine Policy, 155(??):??, September 2023. CODEN ???? ISSN 0308-597X (print),

1872-9460 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0308597X23002087.

Zhao:2021:EPN

[ZWG^{+}21] Changping Zhao, Yecheng Wang, Yu Gong, Steve Brown, and Rui Li. The evolution of the port network along the Maritime Silk Road: From a sustainable development perspective. Marine Policy, 126(??):??, April 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X21000361.

Zhao:2020:CPF

[ZWJvL20] Zhiwei Zhao, Emma Wadsworth, Jørgen Riis Jepsen, and Wessel M. A. van Leeuwen. Comparison of perceived fatigue levels of seafarers and management approaches in fatigue mitigation: Case studies from two Chinese and two European shipping companies. Marine Policy, 116(??):??, June 2020. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X19307638.

Zheng:2021:RJT

[ZWXL21] Sining Zheng, Shuqin Wang, Wei Xu, and Qiang Liu. Research on the job transfer pathway of Chinese marine fishermen and its driving factors. Marine Policy, 129(??):??, July 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X21001834.

Zhu:2024:RSF

[ZX24] Jiaxin Zhu and Qi Xu. Reflections on the settlement of fisheries disputes between the EU–UK in the post-Brexit era: Lessons for China’s fishery enforcement disputes settlement. Marine Policy, 169(??):??, November 2024. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003634.

Zhao:2021:PSF

[ZXGL21] Changping Zhao, Xiya Xie, Yu Gong, and Beili Liu. The propagation of sustainable fishery by Arctic shipping route stakeholders. Marine Policy, 131(??):??, September 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100230X.

Zhou:2021:LSF

[ZY21] Qin Zhou and Kum Fai Yuen. Low-sulfur fuel consumption: Marine policy implications based on game theory. Marine Policy, 124(??):??, February 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309490.

Zou:2023:RBU

[ZY23] Keyuan Zou and Qiang Ye. The relationship between UNCLOS and Customary International Law: Some reflections. Marine Policy, 154(??):??, August 2023. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2300218X.

Zhou:2021:EMK

[ZYTT21a] Yusheng Zhou, Kum Fai Yuen, Beverly Tan, and Vinh V. Thai. The effect of maritime knowledge clusters on maritime firms’ performance: an organizational learning perspective. Marine Policy, 128(??):??, June 2021. CO-DEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X2100083X.

Zhou:2021:MKC

[ZYTT21b] Yusheng Zhou, Kum Fai Yuen, Beverly Tan, and Vinh V. Thai. Maritime knowledge clusters: a conceptual model and empirical evidence. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20309441.

Zheng:2023:PDA

[ZZD^{+}23] Li Zheng, Hongyang Zou, Xiaofeng Duan, Zhongguo Lin, and Huibin Du. Potential determinants affecting the growth of China’s ocean economy: an input-output structural decomposition analysis. Marine Policy, 150(??):??, April 2023. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X23000477.

Zhao:2021:PBW

[ZZH21] Minghua Zhao, Pengfei Zhang, and Gaochao He. Portbased welfare services for seafarers in Chinese ports: Their roles, changes and challenges. Marine Policy, 130(??):??, August 2021. CODEN ???? ISSN 0308-597X (print), 1872-

9460 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0308597X20300701.

Zaldua:2024:CFN

[ZZMG^{+}24] N. Zaldúa, Z. Zajková, A. L. Machado-Gaye, V. Franco-Trecu, M. Cosse, Y. Ropert-Coudert, A. Kato, and A. Soutullo. Changing the focus: the need for cross-scale dynamic management in the Southern Ocean and implications for holistic conservation of Antarctic marine living resources. Marine Policy, 170(??):??, December 2024.

CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X24003592.

Zheng:2021:ACP

[ZZW^{+}21] Hui Zheng, Li Zhang, Shuyi Wang, Juan Xu, and Xin Zhao. The affecting channels and performances of financial development and poverty reduction: New evidence from China’s fishery industry. Marine Policy, 123(??):??, January 2021. CODEN ???? ISSN 0308-597X (print), 1872-9460 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0308597X20309714.